A Complete Bibliography of Publications in Studies in History and Philosophy of Science

Part A

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

17 April 2025 Version 1.71

Title word cross-reference

$10.00 [Eas72, Hei75, Osl72]. $11.00 [Ano73a]. $12.50 [Ege70a]. $14.25 [Vea72]. $14.50 [Bro71]. $15.00 [Elk74, Jan73, McR72]. $16 [Car73]. $16.00 [Gab73]. $16.95 [Rhe94b]. $169.00 [Sno04a]. $17.50 [Gro90]. $19.50 [Ait70]. $22.50 [Bus71]. $224 [Hac89]. $24.50 [Ano91b]. $24.95 [Ere91]. $25.50 [McE76]. $27.50/£ [Hop94]. 3 [Grü78]. $3.95 [Gal71]. $35.00 [Ful93]. $45.00 [ML18]. $49.50 [PR90]. $5.95 [Luc71]. $50.00 [Gui04a]. $6.00 [Osl71]. $7.50 [Ege71]. $73.50 [Ere91]. $74.75/£ [Hop94]. $8.50 [Kni71]. $8.95 [Bau71]. < [Sto12]. > [Sto12]. ^{th} [FSJ08]. _{2} [Mau13]. E [Mül12]. f(R) [Due21]. H [Kas14]. Λ [Jac21]. S [vL24]. X [Hen16].

-c [Ryc92]. -matrix [vL24]. -space [Grü78]. -theorem [Kas14].

/£ [Rhe94b].

0 [Cul04, Ful94, Jas04, Sie04, Sno04a, Ste04]. 0-19-504628-5 [Ful94]. 0-19-506447-X [Kam94]. 0-19-514583-6 [Jas04]. 0-19-875138-9 [Mul94]. 0-226-31881-8 [Hop94]. 0-226-31882-6 [Hop94]. 0-226-53228-3 [Rhe94a]. 0-226-53229-1 [Rhe94a]. 0-262-02477-2 [Gui04a]. 0-262-08086-9 [Ful93]. 0-262-62076-6 [Rhe94b]. 0-300-09297-0 [Cul04]. 0-300-10160-0 [Cul04]. 0-415-04056-6 [Mis94]. 0-415-04057-4 [Mis94]. 0-415-05886-4 [Ach92]. 0-520-08816-6 [Ste04]. 0-520-08817-4 [Ste04]. 0-674-60483-0 [Liv93]. 0-674-65760-8 [Sch91]. 0-7914-0211-8 [Ere91]. 0-7923-1276-7 [Zol93]. 0-7923-4924-5 [Sie04]. 0-7923-5033-2 [Sie04]. 0-7923-5744-2 [Sno04a]. 0-8133-0611-6 [Gie91b]. 0-86078-297-2 [Bia94]. 0-942299-58-2 [Ano91b]. 0252 [Spa91]. 0252-7960 [Spa91]. 0495895 [Cul06].

1 [Bru99a, CL05b, Cun00, ML18, O’F10a, Pes13, Rap11b, Ric09a]. 1.87 [Nic72]. 11 [vdLBV22]. 11.95 [Mul94]. 1214/1220 [KH17]. 1391 [CS78]. 14.95 [Rhe94b]. 1592 [Par12]. 160 [Whi74]. 1632 [Gia94]. 16th [Dav20]. 17.95 [Hop94, Sha90a]. 175.00/$99.00/£ [Zol93]. 1750s [Sto25]. 1770s [Sto25]. 17th [Dav20]. 180.00/$99.00/£ [Ere91]. 1800 [BCC87]. 1835-1900 [Ste22c]. 1850s [CB92]. 1860 [Gab73]. 1870s [Cam18b]. 1875 [Hol21a]. 1878 [Car73]. 1897 [Ano73a]. 18th [Fri24a, Rei18, Ano89a]. 1914 [Cra93]. 1920s [Mil19, Ram92, Ric03]. 1930s [DB96]. 1933 [Hop94]. 1956/57 [Pie97]. 1966 [Pal71, Elk74]. 1967 [Bus71]. 1968 [Bri73]. 1970 [Kir73]. 1970s [Aro18]. 1972 [Ano72d]. 1978 [Ano78b]. 1980s [Cam18b]. 1984 [Ano84g]. 1996 [Ano95c, Ano96a]. 19th [Ano90f, Fay24b, Gay76, McO96, Rup76, Wra19].

2 [Bru99b, CL05a, Cul06, O’F10b, Ric09b, Sno04a]. 2-10-0495895 [Cul06]. 2.75 [Rog72]. 2004 [Ano04b]. 2006 [Ano07a]. 2009 [Ano09a]. 201 [Aba08b]. 2011 [Ano11a, May14]. 2012 [Ano12a]. 2014 [Ano14a]. 2022 [Ano22k, Ano22h, Ano22g, Ano22i, Ano22j]. 2023 [Ano23l, Ano23g, Ano23j, Ano23h, Ano23k, Ano23i]. 2024 [Ano24i, Ano24l, Ano24g, Ano24k, Ano24j, Ano24h]. 2025 [Ano25c, Ano25d]. 20th [Ano90f, JW22, Mir04]. 20th-century [Mir04]. 21st [Dur13]. 2211 [Zam18]. 2211-4548 [Zam18]. 23.95 [Sch91]. 27.50 [Mye89].

3.00 [Mun74]. ’30s [Ram92]. 31.95 [Liv93]. 32 [Ano01b]. 35.00 [Ach92, Kam94, Mor90b]. 39 [Aba08b]. 3rd [Ano95c].

4 [Ste04]. 40.00 [Mis94]. 419pp [Heu14]. 45.00 [Bia94, Hac89]. 4548 [Zam18]. 48 [Gam21b].

5 [Ful94, Sie04]. 5.95 [Blo75]. 50 [May71]. 50s [Ege70b]. 581 [Ano01b]. 59.00 [Ere91, Zol93].

6 [Jas04]. 60 [Blu21a]. 60.00 [Cra93, Mor90a]. 64 [Gam21a].

7 [Heu14]. 7.75 [Gra72]. 7960 [Spa91].

8 [Gur19]. 8.95 [Rhe94a, Sha90b].

9 [Mel04]. 90-247-3778-8 [Ere91]. 96b15 [Gol04]. 96b15-25 [Gol04]. 978 [Gur19, Heu14, ML18, Pes13]. 978-0-226-40188-1 [ML18]. 978-0-7391-8032-7 [Heu14]. 978-0-8229-4450-8 [Gur19]. 978-1-118-62077-9 [Van18]. 978-3-0340-1006-1 [Pes13]. 981 [Mel04]. 981-02-3920-3 [Mel04]. 981-02-4044-9 [Mel04].

= [Red87].

A. [O’G78]. A.D. [CS78]. A1 [Ano23l, Ano23g]. A14 [Ano23g]. A2 [Ano23l]. Abduction [Nii11, BP20, Bir10]. abductive [Car05, Kle83, Tur04, Woo11]. Abi [Edi22, AR22, Rob22]. Abi-Rached [Edi22, AR22, Rob22]. Abiding [Mor01]. abilities [Bro11b]. ability [Sta18]. abolished [Blu21a]. Abraham [Hei24]. absence [Ric92, vD21]. absent [Gin09]. absolute [FK90, HK21, Mit93, Tho15]. Absorbing [HS23c]. Abstract [Ano98a, Ano98b, Ili04, PvF14]. abstraction [Kid11a, Ord16]. abundant [Bro16c]. academia [Leu15]. Academic [Ere91, Sno04a, EHPL22, Sto25]. academics [PS24]. academies [Har96]. Academy [Cra93, Cro92, Cro11]. Accelerating [WL22]. Accentuate [Len21]. accept [ZZW20]. Acceptance [Ste15, Bro15, Daw21]. access [LW13a]. accessible [Mar22]. accommodate [Fei14]. accommodation [Wor14]. according [OS14]. Account [Mil14, Pin22, Bia20, Bou21, De 11, Eng18, Eve18, Kri18, Met91, NI03, Sch09b, Smi13b]. Accountability [WHK14]. accounts [Dar21b, San16b, ST20, Vos07]. accumulation [Bir08]. accuracy [Cam21, Gré19, Mat22b, Sib95]. Achieving [Eva10]. Achinstein [Che90]. Acidity [RC17]. acquiring [Pla20]. across [Hum19]. act [Le 58, PMG21]. Actants [Kho03]. action [Bro16b, Cat12, Duc11, Goo78, Hen11, Hen14a, Mer92, Sch11]. active [Mat80, Par20a, Wal16]. activities [Bog08, Lac15]. activity [Baz88, Gro90]. acts [Bel20, Gil83]. actual [Fal17, Hat88, Wan22, dC10]. Ad [Kep68, KR08, Car15, Lep75, Ray92, Sch18a, Sch24, Fie86]. Adams [SOC19]. Adaptation [Bra78, Cam83, Gre21, Shr07]. Adaptationist [Res89]. adapted [Cur19b]. addressed [CS78]. Addressing [Daw22]. adequacy [Bue97]. adequate [Cob11]. adequately [Bou21]. ADHD [Koi21]. Adjusting [Mil14]. adjustment [Gré19]. Adolphe [CM00]. adsorption [Liv09]. advancement [Ful94, Kit93, Sol95]. adverse [O’G78]. advice [Qui72]. advise [Fra93]. Aeronautical [HLM13]. Aesthetic [MS24, Bra08, IRDV24, Mar09]. aestheticism [Mor93]. aesthetics [Guy08]. Aether [Ner84]. Aether/or [Ner84]. Aetna [Hin12]. affair [Fin89, Sha90b]. Affect [Nik19]. affinities [Whi90]. afraid [Dou21]. Africa [Coe22b, Van95]. African [Coe22b]. after [Bai87, Bal92, Car13, Dre20, Fra84, Run94, SG13, Wra05]. afterimage [Web22]. afterlives [Rad13]. aftermath [Bla24]. Afterward [Wis18]. again [Col18]. Against [Ayl19, BRdS96, GK88, Non94, Pic84, Fra93, Jar79, Sha20, dB09, CF21b]. Agassi [Aga71, Bec71]. Age [Mia01, Gio24a, Hac91, Mei21, Van04a, Wra06, dC10]. agencies [Hau13]. agency [Cat12]. agenda [Fri98, Rap11b]. Agendas [May00]. agens [Sie07]. agent [Anz19, Bha17]. agent-based [Anz19]. agentive [CP18]. agents [Coo05]. Ages [PR90, Fun86]. aggregation [MN20]. aggregations [Ord16]. Aggression [Lon01]. aging [Kuu17]. agissent [Le 58]. agnosticism [CCB24]. agree [Fau08]. agreement [Cle23, Goo86, JB16]. agricultural [BK21, Cha13, Fri24a, WL22]. agriculture [WL22]. agroecological [Lac21]. Ahlborn [Hin15]. AI [Bec23, LF25, Tam23]. Aim [Lau90, Tam15]. Aim-less [Lau90]. aims [Buc88, Kid11b, Pot15, Ser23]. ain’t [Dou21]. airplane [Mac13]. airs [Boa07]. Akaike [Lee13]. Al [Gal71, GS22]. Al-Azm [Gal71]. al-Din [GS22]. Alan [Sch91]. Albany [Ere91]. Albert [Dra75]. Alchemical [Cle97, Ral06]. alchemist [Kas04]. alchemists [Ram12a]. Alchemy [McC06, Mor00, Coo96, Kir09, Ral06, Ram12a, Why21]. alcoholism [PM23]. Aleksandr [Gar00]. Alex [McE76]. Alexander [Gol75]. Alexandria [Ose22, Tyb03]. Alfonso [Fra93]. Alfred [Kam94]. algebra [CW68, Hen84, Neu71, Eas72]. algebraic [Cal24, CF21a, Kit75, Men24]. algebras [Man22]. Algorithmic [McA03, McA05, JW21]. algorithmically [TGD05]. Alhazen [Cro90]. alien [Cha19b]. all-pervasive [De 16]. alleged [Aga71]. Allen [Sha90a]. almost [DBM22]. Alois [Kwa12]. alterations [Gee14]. alternative [Bro15]. Alternatives [Men19]. Alvin [Kus01]. Always [Mia01]. am [Kuu11]. ambiguities [Yea82]. ambiguous [Loi22]. ambivalent [BS13, Whi90]. America [Dai19, Ras96]. American [Car73, Tur22b, GR23, Har19a, Mir04, Rei96]. amis [Des49]. amodern [Lat90]. Amos [PR90]. Ampére [Gur19]. ample [Suá12a]. Anachronism [Tos03, Hal05]. Analog [Pap20, Sha10]. analog-model [Sha10]. Analogical [KL14, Nas16, Gen20]. Analogies [Car16, Bec23, FG22, Nap21, Rod21]. Analogue [Pap20, SHB94, Wat23]. analogues [Gre21, Tur05a]. Analogy [Fis18, Man12b, Bar20b, Bok15, Cam18a, Cur16, McG70, vdB18, vdB24]. analyses [Nor22, Zyn96]. Analysis [Cha20a, Cha20b, Aro71, BH19, CP10, ES06, EG80, Fal24, Fra79a, Gau76, Hin15, IRDV24, Kim00, Kit75, Lau23, Liv09, Maz22, Pan06, Raf03, Ric21, Sch85, Vez16, VS12, vdLBV22, Han21]. analyst [She87, Mor22]. analytic [Fal18, dJ97]. analytic-synthetic [dJ97]. Analytical [Fer01, Fra90, Kra81, Sol08b]. Analytics [Pat69, Gol04, Vea72]. Analyzing [Lus21]. anarchism [vS20]. anarchist [Sha17]. Anatomical [Kut22, Ose22]. anatomists [Tur22a]. Anatomy [Wal75, Ols16, vdB24]. Anaxagoras [Alt12]. ancestral [Tur22a]. ancienne [CL05c, Cul06]. Ancient [Fek24, Rob03, Bak16, Con00, Cul06, DFT12, Hay04, Hei16, Leh04a, Nor22, Smi15, Asp13]. Andrew [Mye89, Bai92, Dea01b, Hen14a]. android [Vos07]. anecdotal [Buc89]. Angeles [Ege71, Ste04]. angle [Ekb24]. Anglo [FM04]. Angular [GG21]. Animal [Col21, FL21, Vic21, vdB22, Bae24, Dav21b, Hol21a, KM22, Loh21, MHW22, Nob23, PMG21, Pol21, Pri22, SL21]. animal-based [Loh21]. animals [Aro18, Atr85, Coe22b, tF18, PMG21]. Animism [Koc24]. Anna [Eig19]. Anne [Ste04]. Anniversary [Ano89b]. annotated [Ano03f]. Announcement [Ano70a, Ano71a, Ano71b, Ano77a, Ano79a, Ano80a, Ano82a, Ano89c, Ano92a, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b]. Announcements [Ano74a, Ano85a, Ano85b, Ano86a, Ano90a, Ano91a]. Annus [Elk74, Pal71]. anomalies [Bîg13, Kem07, Van79]. anomalous [Fis06]. anomaly [Kal75, Sch87]. Anschauung [De 16, LJ11, Kra72]. Anschluss [Pes13, Sie10]. Anstey [Cha02]. answer [Coh08b, Zam08]. antecedents [Mun16]. anthropic [McM93]. anthropogenic [Vez16]. anthropological [Bec23]. Anthropology [Ham23, Lou08, Bia90, Bra98, Coh08b, Lat90, Stu08]. Anti [Blo99, Row23, TA23, Ara23, Bra98, Cok15, Díe11, Edg08, Hen23, Mor22, Pap89, Pre08b, Ueb25, Waj21, Wat13]. anti- [Wat13]. anti-anthropology [Bra98]. anti-atomism [Cok15]. anti-exceptionalism [Ueb25]. anti-inductivism [Díe11]. Anti-Latour [Blo99]. anti-mathematical [Mor22]. anti-psychologism [Edg08, Pre08b]. Anti-Realism [Row23, TA23, Ara23, Hen23, Pap89]. anti-reductionist [Waj21]. anticipates [Obe16]. antiken [Neu71]. antiquity [Dea03, Dun12, Mou20, Way21, Olm12]. Antirealism [TA23]. antirealisms [Kuk95]. antithetical [Fis85]. Antoine [Bas92]. Antoni [Dav20]. Antonio [Cam12]. any [Grü78, Joh22, Le 58]. anything [Sha17]. apart [Riv21, dR13]. apocalyptic [Clu12a]. Apology [Ano97f]. apparently [O’G78]. appeal [Bas17]. appearances [PEK06]. application [And94, Bor20, Cha73, Fra79a, Gre22, Mar13, Tei22]. applications [Jhu22]. applied [Gro87, Mer82, O’G77, Ver07a]. apply [Dic08]. Applying [MB22]. appraisal [CB92, Jon78, Mic71a, NRB04, Qui72, Whi88]. Approach [Bat23, Dar23, FR23, Heu14, Mil23, Pal23, PK14, AS16, Bau98, BZ81, BZ86, Bîg13, Bog21, Bue97, CF21a, Con00, Dar07, DZ14, EZ13, FH85, FH88, GR23, Har19b, Joh21a, Knu11, Mat11, Pal70, Ree12, Res00, Spa23, Tho83, Wag24, dC18]. approaches [AB23, Car03b, LW13a, Tam20, Wat21]. Approaching [SFT23]. approximation [Kui14]. approximations [dOFGB22]. April [Ano22k, Ano23j, Ano24i, Ano25c]. Arago [Tre07]. Aratus [Gee14]. arbitrariness [Aga71, vL24]. archaeological [Cur16]. archaeology [Kos93]. archetype [PB96]. Archimedes [Cam18a, Goe72]. architecture [For89]. Archive [Mot02]. Archives [Sno04a, dC13]. argument [Bar21b, Car91, Car93, Dar02, GG94, Goo80, How15, Kle83, LM12, Len21, Mee92, Men14, Mit17, Miz12, O’N98, San12, Sch02a, Smi13b, Wis84, How15, Men19, Wra18]. argumentation [Fin11, GG02]. arguments [Cal10, Car05, Har10, Kos99, Tur04, Wei10a]. Ariew [Ano01b]. Aristarchos [Wal75]. Aristotelian [Atr85, Cox12, De 00, Ede09, HH00, Man92, Mar04, Nee96, Nee02, Rap11b].

Aristotelian-inspired [Rap11b]. Aristotle [Con00, Föl12, Gau81, Gla20, Gra04, Kri80, Mar97, Mau90, Mia01, Pat69, SG13, Ver82, Wal91, WW04, Bia94, Vea72]. arithmetic [McC06, Kno76]. arithmetical [Qui18]. Armstrong [Jac95]. Arnold [Set08]. arrive [Dor70]. arrow [HS23c, Ama05]. Ars [Cam18a]. art [CW68, Coc14, Eas72, EC25, Kes07, Kwa12, Low93]. Arte [Dun12]. artefact [Hou06, McL06a, Pre06, Sch09b, dR06]. artefacts [Bak06, Fra06, HM06, HKMV11, Kro06, KM06, Qui17, Sch06, Vae11, Ver06, VH06]. artefactual [Knu11]. Arthur [Cha20a, Cha20b]. Article [Tal25, Baz88, Gro90]. Articulations [lC00, Pen22c, SFF23]. artifact [De 11]. artifacts [PE24, Ras93]. artifactual [CK22]. Artificial [Bec23, Ric97, San16b]. Artisanal [Jac94, Ral06]. artisans [Mor05, Ral06]. arts [Dea03]. Arturo [Luc71]. Ascent [Nik19]. ascription [Sch06]. ascriptions [Ver06]. ’Asfūriyyeh [Edi22, AR22, Rob22]. Ashmole [Feo12]. Asian [Low93]. Aspects [De 00, Pan99, Atr85, BO24, Buc72, Kai92, Red80b, Rog72, Sim14, Tie88, Woo71, Woo94]. Aspirations [Kus01]. assault [Obe05]. Assaying [Cor97]. assent [Cro90, Cro91]. assess [Col21]. Assessing [Gré19, Axt93, Maz22]. assessment [Bue15, Cas24, Daw21, LT20, Wol24, vD21]. assessments [Rit21]. ASSET [Ano84a, VFHS23]. associated [AD12]. Association [Ano95c, Hee21, Ano84a]. Associationism [Dac15]. associationist [Dac15]. associative [Dac15]. assumptions [Den00, Pic22, Sch20b, dC10]. astrolabe [CS78, Eag07, ES06]. Astrological [Van04b]. astrology [Leh04a]. Astronomer [Mos01, Rei08, Hun03]. Astronomia [Miy15]. astronomiae [KR08]. astronomical [Grü78, Pan96, Rot11]. Astronomy [Gre96, Mou81, Abe22, Ben87, CD15, GT05, Hac89, Mit72b, Pal70, Rem07, Roc02, Spe17, Tay09, Van04b, Wri73]. astrophysics [Eld24]. Astrostereon [Bie20]. asymmetric [Gay78]. asymmetry [Fee06]. Athanasius [Ver07a]. atheism [O’F10b]. Athens [Sto12]. Atlas [Rem07]. atmosphere [Mar15a]. atom [Gay78, Sut71]. atomic [Nye76]. atomisation [Mol82]. Atomism [Kri86, Bog75, Cok15, Daw21, Gab73, Sco70, Wil82]. atomist [Ros19]. Atoms [BR23, Cle23, Eat23, Gol23, Gol04, McG70, Whi90, Big08, Jan00, KC94, Qui17]. attachment [Ryc92]. attack [Cat95]. attempt [Zam08]. attempts [Dor70]. attitude [Hen92]. attitudes [ME15, SGO17]. attributes [Pal02, PEK06]. attribution [GG94, WOL20a]. auctions [Gua01]. audiences [How04]. Aufbau [PY06]. Aufbau/Bauhaus [PY06]. August [Ano22h, Ano23h, Ano24l]. Augustan [FM04]. Auguste [McC01]. Augustus [Phi05]. Australasian [HBR24]. Australia [Max22]. Australian [Tur22a, Way21]. Austrian [Ueb91b, Ueb94]. Austro [Ueb25]. authentication [dC02]. Author [Ano92b, Ano02a, Ano05a, Ano06a, Bro21, Gol22, McH21, Miz21, Rhe25, She22, Val22, AR22, Ano80g, Ano81d, Ano82d, Ano83d, Ano84e, Ano85g, Ano87e, Ano88d, Ano89h, Ano90g, Ano91g, Ano95b, Ano01c, Ano04a, CM00, BR23, Bat23, Nor21a, Ric22, Row23, Sol22]. authorial [Tot12]. Authority [Bla96, Net93, Sor16, dC12]. Authors [MM24a]. authorship [Dea03, Wra06]. autism [Koi21]. autobiographical [DD69a, DD72]. autobiography [tH02]. autoinhibition [BB22b]. Automata [Nob23, Gei23, Vos07]. automated [Ull24]. Autonomy [FR23, Pal23, BR19, BC92, Coa02, Düp15, Fis06, Hoc16, LM12]. autres [Le 58]. avant [Bra08]. avant-gardes [Bra08]. Average [Leh18]. Averroës [EB11]. aviation [HL13]. avoid [Fra94a]. avoidance [Sch24]. away [Blu21a, Chi16, WOL20a]. awry [Gol75]. axiomatic [Mit22, Stö15]. Axiomatics [Díe97a, Stö15]. axioms [Cal24]. Azm [Gal71]. AZT [Wei07].

B [ALR21, Bid11, Sla02, Whi74, Gay78]. B. [Car25]. B7 [Ano79d]. Babbage [Coo05, DS21]. Babel [Kap16]. Babylonian [Roc02, Neu71]. Bachelard [BV05, CL04, Chi08, Fab19, Gau76]. back [Dra87, LC24b, Mat22a, McK16, RGD15]. Background [Hud94, HG05]. backward [Tro94]. Bacon [Osl71, Har74, Hor73, KH17, Lux67, Sar12]. Baconian [Sar86, Woo75]. bad [Fin02, Del17]. Baigrie [Lug92]. balance [HG05, Le 58, Ros17, Row13, Smi13b]. Balances [Dör94]. Baldini [Gia94]. ball [Sla17]. Baltimore [Ano73a]. ban [Her15]. Banchetti [BR23, Cle23, Eat23, Gol23]. Banchetti-Robino [Cle23, Gol23]. Banchetti-Robino-Robino [Eat23]. Band [Neu71]. Banks [Sar12]. barnacle [Jar09]. Barnes [Sce05, Wor05]. Barrovian [Len07]. Bas [Kuk95]. base [Dic13, Men14, Pie14, Sch17]. base-rate [Dic13]. based [Anz19, Bîg13, Bou16, Gel11b, Gel11a, HS23b, Hin12, Joh21a, KH25, Knu11, KD19, Loh21, MR19, PG24, Pot12, Qui17, Yli19]. bases [Goo78]. basic [Cov97]. Basil [Rog72]. Basin [Boe22]. basis [Dup94, Pal08a]. Bateson [Rad13]. Bathybius [Rup76]. Batterman [Dar23, FR23, Mil23, Pal23, Woo24]. Battles [Hau22, Lin22a, Rot22, Sol22, Tur22b, Bor92, De 96a]. baubles [Pet15]. Bauhaus [PY06]. Bavarian [Jac94]. Bay [CSD21]. Baye [Ear90]. Bayesian [Dor79, FH85, FH88, Len22, Nun84, Ped20, Red80a, Sch14b]. Bayesianism [Ear90, Hes75]. Bazerman [Gro90]. Be [Nik19, Ahn20, And13, Bar18, BR19, Bro90a, Cor21, Dic06, Dri24, Gro20, Hee19, Hol22, Lan21, MJ23, Mee92, Nel94, Qui17, Ram92, SLH23, vdLBV22]. Beagle [Jar09]. bearing [Sie04]. beat [San22]. beautiful [Cuo06, Kon24a, Pal05]. Beauty [IRDV24, Mou20, LJ11]. became [Blu21a]. Becher [Coo96]. Becker [Ric03]. become [Lau23]. Beef [Rot22]. Been [Mia01, Bon94, Coh97, Lat93]. Beer [Ege70b]. before [Ari87, Bai87, Fra84, Hen83, KH20, Mel21]. begets [How15]. beginning [Fri24a]. beginnings [Wan12, Zin11]. Begriff [Fey16b]. Behavior [Sha94, KT21]. Behavioral [Ano80c, Fed23, Lis18, MC25, Meu19, Rom16]. behavioural [Tra22]. behaviourism [Tam05]. behaviours [Waj21]. behind [Bha15, Bol15, Gie09]. being [Mar24, Ueb94, Wal16, Yee24]. beings [Pal09a]. Beitrag [Car21]. belief [Bro15, Van91, Wal06, dL09]. beliefs [Bas92]. Believing [Gon25, Cok15]. Bell [Bra24a]. Bénard [Aub08]. Benedetti [DD69b, Dra76]. beneficial [Ahn20]. benefit [LT20, Wys23]. benefits [Pla20, dG20]. Benjamin [Mye89]. Bennett [Hac89]. Bentley [Bro16b, Cal10]. benzene [Bru99a, Bru99b]. Benzer [Meu19]. Bergmann [Pit22]. Berkeley [Ege71, Sha90b, Ste04, Swo74, Bau72, Har24, Mok71, Mor22, She87, She93, Win86]. Berlin [Neu71, Düp15]. Bernard [Gui04a, Ste04, Van18, Sch20b]. Bernoulli [Cam18a, Gow87]. Berry [Lyr14]. Berzelian [Gro05]. Berzelius [Bro73]. best [Ach92, CDC24, Dav21a, Gel10, Hal05, Hen22, Kle79, Lip91, Oka00, Pes10, Wal12, Wat12]. better [Cra22, LGC21]. Between [Fer12, Hel16, Jar09, McK95, Rei96, SBM09, tH02, Acu25, Ara92, Asp13, Bas92, Bor15a, Che03, Chi01, Cox12, DK18, Dio13, Düp15, EHPL22, Fur22, Gab73, GL24, Gla83, Gla89b, Hen14a, JM11, KR22, Kva14, Los83, Mar97, Mit17, Nee08, Nel94, Pan06, Pri22, RK06, San91, Sch16, Sco70, Sha97, VH06, Yea82]. Beyond [Aba08a, Aba08b, Bac20, Cam21, EHPL22, Man12a, Rad97, Sha21, BDR13, Ede09, tF18, Gre24b, Hol21a, Lan20, She06, Tul16, WH14, Wit13, dRM23, Bel15]. Bias [Bue22, Wil09a, BH19, Vae11, Ven22]. biases [Leu15]. biblical [How98]. bibliographical [Ege70b]. Bicentenary [Spa91, Ano84d]. Big [Gre24a, Nor20a, PP23]. Bigger [Gre24a]. Bio [tF18, Gla83, Tam23]. bio-robotics [Tam23]. Bio-Science [Gla83]. Bio-sentinels [tF18]. BioBike [Shr07]. biocomputing [Shr07]. biocultural [GR23]. Biogeography [Fra84]. biographical [Mor90a, SW89]. Biographies [Ara03]. Biography [Nor00, Car89, Mor90b]. Biological [Ano97g, Bra81, Gam21b, GVD24, Gre21, Vet18, Aro18, BB22b, Dup94, Fag15, Föl12, Gam21a, GKS24, Lin98, Nas16, Pie97, Ras96, Yos23]. Biology [Dar96, Ere91, Kam94, McO01, Ano90e, BB23, BP25, Cam21, CP18, Cra22, Fis23, Gam21b, GN22, JW21, KL14, Loi22, Man80b, PE24, Rus73, Rus89, Sim21, Sto09a, Mun74]. Biomedical [Ano97g, Gam21b, Cho22, Gam21a, SBM09]. biomedicine [FBW22, Loh21]. Biometry [Sha21]. biomimetic [Tam23]. biomolecular [Bol23]. biomonitoring [Cre18]. Biopolitics [Mel21]. biopsychosocial [Kon24b]. biosemantics [Gre21]. biosocial [Pen22c]. biotech [Pet15]. biotechnology [Kir09]. Bird [Pre04]. birth [Bel92, BJ22, Bra24a, Bre04b, CM00, Kob24, Luc20, ND22, Rog95, Vor14]. birthday [Buc88]. bit [McL05, McL06b]. Bjorken [KG13]. Black [Mer04, Fur22, Maa13, Ram92, Sch02b, Wei21a, Wei21b, Mah99]. Blackmore [Swo74]. Blackwell [McR72, Rog72, Van18]. Blaise [Jon01]. blame [CP07]. Bland [San14]. blood [Pri22]. Bloomington [Hei75]. Bloor [Buc82, Fre79, Hes82, Kem05, Kem07, Luk82, Seg00, Smi84]. Blumenbach [HN16]. Board [Ano18a, Ano18b, Ano18c, Ano17e, Ano17b, Ano17c, Ano17d, Ano70e, Ano71g, Ano72e, Ano73d, Ano74e, Ano75c, Ano76a, Ano77d, Ano78c, Ano79e, Ano80e, Ano81b, Ano82c, Ano83b, Ano84f, Ano85e, Ano86e, Ano87c, Ano88b, Ano89f, Ano90c, Ano91e, Ano92f, Ano93f, Ano94d, Ano95a, Ano96c, Ano97e, Ano98p, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano03e, Ano04c, Ano04d, Ano04e, Ano04f, Ano05f, Ano05c, Ano05d, Ano05e, Ano06b, Ano06c, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13a, Ano13b, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c]. Board [Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano18f, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano25a, Ano25b]. Bodies [FJ24, Bar21c, Bog75, Cre18, Gue17, Hen14a, Jac20, KC12, Le 58, Mat80, Mor05, Wil77]. bodily [Tur22a]. Body [Bat23, Dar23, FR23, Mil23, Pal23, De 02, Dem06, Dou95, FBW22, Le 58, Sch03a]. body-body [Sch03a]. Bogdanov [Gar00]. Bohm [Tal25]. Bohmian [DE21, Bar21a, Fei14]. Bohr [Bel92, Fav91, Fay88, FJ21, Hon82, Toa24]. Bolstering [KH25]. bolt [Tos07]. bolt-holes [Tos07]. Boltzmann [Fla83, Wil89, dC10]. Bolzano [Kit75]. Bonaventure [Mou20]. bond [Har08]. bonding [Nee13, Nee14]. BonJour [Vog10]. Bonnevie [Tei22]. Book [Ach92, Ait70, Ano73a, Ano91b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano00a, Ara23, Bak16, BR23, Bat23, Bau71, Bia94, Blo75, Boo23, Bri73, Bro71, Bro21, Bus71, Cam22, Car73, Chu22, Cle23, Cra93, Cul04, Dar23, Die24, Eas72, Eat23, Ege70a, Ege70b, Ege71, Elg22, Elk74, Ere91, Fin73, FR23, Ful94, Ful93, Gab73, Gal71, Gia94, Gie91b, Gol23, Gol22, Gra72, Gro90, Gui04a, Hac89, Hau22, Hav21, Hei75, Hen23, Heu14, Hop94, Hus24, Jan73, Jas04, Jhu22, Joh22, Kam94, Ken23, Kir73, Kni71, Kra22a, Kuk22, Lea22, Lin22a, Liv93, Luc71, MM24a, ML18, May71]. Book [McE76, McH21, McR72, Mel04, Mil23, Mis94, Miy23, Mor90a, Mor90b, Mul94, Mun74, Mye89, Myk22, Neu71, Nic72, Ode22, Osl71, Osl72, Pal23, PR90, Pin22, Rhe94a, Rhe94b, Ric22, Rob22, Rog72, Rot22, Row23, Rup22, Sch91, Sha90a, Sha90b, Sie04, Slo24, Sno04a, Sol22, Spa91, Ste04, Swo74, Tur22b, TA23, Ups22, Val22, Van18, Vea72, Wes23, Whi74, Wie21, WV22, Zam18, Zol93, Fin90, Fra05, Gin09, Gio16, Gro87, Joh98, Pat69]. bookmaking [Nor22]. Books [Ano70b, Ano70c, Ano70d, Ano71c, Ano71d, Ano71e, Ano72d, Ano72a, Ano72b, Ano72c, Ano73b, Ano73c, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano77b, Ano78a, Ano79d, Ano79b, Ano79c, Ano80b, Ano81a, Ano82b, Ano83a, Ano84b, Ano85c, Ano85d, Ano86b, Ano86c, Ano87a, Ano87b, Ano88a, Ano89d, Ano89e, Ano90b, Ano91c, Ano91d, Ano91b, Ano92c, Ano92d, Ano93d, Ano93e, Ano94c, Ano96a, Ano97a, Ano97b, Ano01a, Ano02b, Ano03a, Ano04b, Ano07a, Ano09a, Ano11a, Ano12a, Ano14a, Heu14, Ano03f, Ede09, Rob70, Tot12]. Boole [Coo05]. Boon [Mas24]. bootstrap [Cus85, Fre80b]. bootstrapped [Sch09b]. Border [Düp15]. borderlands [PMG21]. borderline [Bor15a].

Born [Bel90]. Boscovich [Koz15, Ros22]. Boscovichean [Sut71]. Bose [Tom22]. Boston [Ano78b, Zol93, Ano78b]. Boswell [HL13]. both [Cha16, HK21, Let21]. bottom [Row11b, Sch07b]. bottom-up [Sch07b]. Bottoms [BCK^{+}22]. Boulder [Gie91b]. boundary [Hol21a]. bounded [DH22, Sim01]. box [Maa13, Ram92]. Boxes [tH00, Mer04]. Boyd [Mei21]. Boyle [BR23, Cle23, Eat23, Gol23, Aga77, Ans02, Car12, Cha93, Cha02, Cha10c, Cha16, Eat23, Gol97, Mat19, Pyl02, Sar86, Why21]. Brahe [How98, Mos01]. brain [Eli10, HS23a, Luc71, Ros70, Cah97]. Brains [Sha10]. Brandeis [vE14]. Branestawm [Out96]. Brazil [Gra13]. Brazilian [EHPL22]. breaking [Leu12]. breeding [Bru20, CR13, WL22]. Breuer [Nir23]. Brewer [HL13]. Brewster [CB92]. Brian [Lug92]. Bridging [DK18, Nee08, Koc06]. Bridgman [Moy91a, Moy91b]. brief [Mal01b]. brightness [Car15, McG21]. Bringing [LC24b, RGD15]. Britain [Che95, HLM13, How04, KM22, Max22, Ric86, Top00]. British [Car73, AD12, Ano84c, Her69a, Her69b, HL13, Jac94, May04]. broad [Bol14]. Broglie [Bon94]. broom [Mar15b]. brother [HL13]. Brougham [Can71, CB92]. Brouwer [BO24]. Brown [McH21, Bro21, Dac15, Hav21, Pao11, Wie21]. Brownian [Big08]. Bruno [Sch91, Coh97, Gow97, Gra04, Ros19]. Brunonian [PM23]. Buchdahl [McR72, Ano01d, Ano03f, Blo82b, Jar03b]. Buchwald [Gui04a]. Bueno [Bol14]. build [Nor20b]. Building [Gua01, Kai19, Lyn95, Cha13, Hin15, Mat11, Ram08]. buildings [Sum13]. bull [Bru20]. bullshit [Mic24]. Bulzoni [Gia94]. bumpy [Kra24]. bundles [Jac23]. Bureau [Vet18]. Bürgerliche [Cah97, Bra95]. Buridan [Dra76, Kin87]. Burtt [Cha20a, Cha20b]. bus [BM14]. Busch [Saa08]. business [Mac13]. Butterfield [Cab96]. By-whom [Del23]. Bycatch [Jen07].

C [Ait70, Ano97g, Gam21a, Gam21b, McE76, Mor90a, Ahn04, Ast04, Cam12, KH17, ND22, Ryc92, Tur22a, ALR21]. C. [Fra20, Kni71]. ca [HM24, KC94, Rei08, Sch09a]. cabala [Man12a]. cake [Bus08, Saa05]. calculus [Fer01, MB22]. Calibrating [Det22]. Calibration [Tal17]. California [Ege71, Sha90b, Ste04, Swo74, Aro18]. caloric [Psi94]. Calvinism [Van04b]. Cambridge [Bau71, Bus71, Cra93, Eas72, Elk74, Fin73, Ful93, Goo04, Gui04a, Jan73, Kir73, Liv93, Luc71, McR72, Mor90a, Osl72, Rhe94b, Sch91, Coo05, May00, War92, War93]. camera [Sch09a]. campaign [Rad13]. Campbell [War92, War93]. Can [Bid14, Che22, Cul06, Dar21b, Dic06, Dri24, tF18, Fau08, Fei14, Rom16, Ahn20, Has97, Hol22, Kin15a, Kle79, Mac10, MJ23, Rat20, RS21a, Saa05, Szn16]. cancer [PP23]. Canguilhem [Ano91b, Chi15, Mét13]. cannot [Sei13a, Sei13b]. canon [BM21, Fyf02]. canonicity [BM20]. Can’t [Nik19, Cur18, Fre05, Han05, vdLBV22]. Cantor [Mye89, BT09]. capacity [Mum06, POW16]. capital [VFHS23]. CAPM [VFHS23]. carbon [Gay78]. Cardano [Eas72, Wil05]. Cardwell [Mor90b]. care [KM22]. career [Fra79a, Ham00]. Careful [CF21b]. Caring [Pen22c]. Carl [Mäk97, Web07]. Carnap [McE76, Axt93, Cir93, FJ21, Fre15, Kle16, Mic71b, Psi00, Ric03, Ric09a, Stö15, Sto10, Szn16, Tam05, Ueb10]. carnival [Spr05]. Carolyn [Ano91b]. Carrier [Mee92]. Cartesian [Raf03, Sch03a, Cur18, Fra90, GS02, MD13, Sar86]. Cartesianism [Van04b]. Cartwrightian [dC18]. Case [SGO17, YZ19, Alt12, Ara92, Atr85, Bar18, Bas03, Bel90, BM89, BM20, Bra15, Cam12, CP18, Chi11, Coh85, Cul06, DD13, Ekb24, Fal24, Fri24b, Gay77, Gay78, Gei23, Gla86, Gla02, Gök90, Gro87, Hal12, Han21, HK22, Hei80, Hon87a, Hur17, Jen10, Kai19, KG13, Kin15a, Koe95, Kra81, Lis18, Llo15, LA22, LW13b, MM09, MN20, Mei21, MBG21, Mor19, Mor92, Mow85, Nor20a, Ord16, Rat20, Ree86, Rib07, Sch13c, Sie04, Sim21, Van16, VFHS23, War22, Wel05, Wor82, Yli19, Zah19, vE19, vH91, Len07]. case-study [Hon87a, Kra81]. cases [Blo08, CM00, De 02, Düp15, Kon24b, Mor19, PMG21, dOFGB22]. Casimir [Sim14]. Cassirer [Bia20, Cah86, Fer12, Gio22, Gio24a, Hei14]. Castile [Fra93]. Casting [Rap11a]. Castle [Ano91k]. catalogue [Kut22]. Catalyzing [POW16]. Catching [Mil14]. categories [Gre21, Win12]. Catherine [Mas24]. cathode [Hon87a]. Catholic [Rap11a]. cattle [Bru20]. causa [Sch20b]. Causal [EHPL22, Let22, Sch13c, Cal24, Car03b, Eag04, GFE23, Hea94, Hen14b, Kin21, KVJ22, Lem21, LL19, LB22, Mer80, Mil19, Sta18]. causal-deflationary [LL19]. causality [BM89, Goo87, LT20, Pad15, Pao11]. Causally [Bog08]. Causation [Con21, Aro71, Aus78, Hen14a, Hud94, Hut21, Jan13a, KT21, Kno09, Oft22, San16b, Sch03a, Whi71, Zyn96]. cause [Duc11, FJ24, Koc11b, Lar03, LB94, Sch15]. cause-of-death [FJ24]. Causes [Pao11, Car03b, Cha12, For88, GVD24, Pol15, Ryn95a, Ryn95b, San16b, Tuc20, Wan22, Wat03]. Cavendish [Bro07, Bro11a, Dor74, Fal17, War92, War93]. CDM [Jac21]. Cecil [Bus71]. celestial [Mar99, Sha74, Tay09, dS09]. cell [Bec84, Fis23]. cells [All96a, Aub08, Fag16]. Cellular [All96a]. censoring [May11]. censorship [Rap11a]. center [Car05]. central [Cam21]. centrality [Gra04]. centred [Tos03]. centric [Cor21]. centripetal [Pou07]. centuries [Ano90f, Rem07]. century [Bai04, Ben74, Bru20, Buc09, Bur02, Car03a, Cha15a, Che95, Chi08, Cun00, Das86, Dav20, De 02, Dea87, DD69b, Dur13, Edg08, Ege70a, FK90, Fay24b, Fer01, Fis06, Fri24a, Fyf02, Gab71, Gay76, Gre92a, Har96, JW22, Jes89, Lor20, Mal91, Man92, Mat22a, McG69, McO96, Mir04, Mor07, Nye76, Péc11, Pre08b, Pri22, Rei18, Ric92, Rup76, Sar89, Sha02, Sla02, Slo77, Spe17, Ter17, Top00, Van16, Vos07, Wil09b, Wra19, dC02, dC93, vdB22, Cun01b, Fun86, GW73, PR90, Hei75, Liv93, Ano73a]. certainty [AP06, Zel24]. certitudine [Man92]. cervimetry [JW22]. ceteris [Wri17]. Châtelet [Hut04, Ilt77, Jac20, Rei18]. chains [Miy23]. Challenge [Cal21, Nee02, O’F10b, SBM09]. Challenger [CP07]. challenges [Ueb15]. Challenging [Sor16]. Chalmers [New10, Pyl02]. Chambers [Ege70b].

Chance [Vas18, Eag04, Kil99, Nor22, Pen15, Van91, Zyn96, ML18]. chances [Swi86]. change [AL16, Bru99a, Bru99b, Bue00, Che10, Dar76, Eve18, Fee06, Gla86, Gla89a, Gla02, Goo21, Gut73, Hei80, Kit87, LL16, Mac75, Mow85, RT95, Sta08, Tam05, Vez16, Vor14, Wra07, Zan88]. changeability [LAL21]. changes [Che95, Dic25, Lin22b, Sch13b]. Changing [SFF23, Gre22, KM22, Kle05, MM06, Sch70, Sch88d, Sib08]. chaos [Cat12]. chapitres [CL05c, Cul06]. chapter [DD69a, DD72]. character [Woo75]. characterization [Col18, RC17]. Characterizing [JW21]. charge [FR24]. Charles [Gro90, ML18, May71, Zam18, Ari22, Bra24a, Dar76, DD69a, DD72, Hes21, Rus75]. Charybdis [Sei13a]. Chaucer [ES06, Eag07]. Checking [Har02]. Cheiron [Ano80c]. Chemical [BR23, Cle23, Eat23, Gol23, Har08, Kim00, Bas92, Col21, Cre18, McN15, Nee08, Nee13, Nee14, Ram92, Wol20b, Cha15b, Kus15]. chemist [Rob95]. chemistries [Kra22b]. Chemistry [And13, BV05, For75, AG09, Bas03, Bor20, Bro73, Bru99a, Bru99b, Cha10c, Dai19, Fis06, Gau16, Gay76, Gla78, Kim95, Kni71, Kra24, Len14, Mor00, Nee96, Pri19, Rob95, Sch03b, Sla02, Zan88, Zwi11]. chemists [Chr94]. Chemla [Cul06]. Chen [Ach90]. Chernyak [Kam94]. chess [WH14]. Chew [dOFGB22]. Cheyne [Con21]. Chicago [Bak16, Hop94, ML18, Rhe94a, Alm21, Sch12]. Chichester [Van18]. child [BD02, LSH24]. Childhood [MC25]. Chilean [WHMO22]. chimeras [All96a]. Chimie [Hom02, Kim00]. China [Cul04, Cul06, tF18, Kim12, LS02, Wu14]. Chine [CL05c, Cul06]. Chinese [Spr05, ZZW20]. chloride [Cor21]. Chlorine [Bro73]. choice [AS16, CL05a, Dav13, Fum20, Isa14, Iva10, Lic21, Mje24, Sch17, Wel13]. choice-semantical [AS16]. choices [Fum20]. Chomsky [Ter21]. Choosing [Bau98]. Christian [Cab96, Dun13]. Christie [Mye89]. Christie’s [Hac89]. chromatographic [Liv09]. chromosomes [Pen22a]. chronic [vdLBV22]. chronogeometry [Acu25]. Chronos [Pes13]. Churchill [Ege70b]. Chymical [BR23, Cle23, Eat23, Gol23]. chymists [Cle23]. Cicero [Cab20, FB20, Leh20]. Cigoli [Fin11]. cinema [Alg08]. cinematography [Aub08]. circa [Bra08, Noa08, Sib08]. circadian [BA10]. circle [Ben87, Gin88, Hac89, Leu12, MH84, PM23, Ben24, Ric09a, Ric09b, Ueb91b, Ueb05a, Ueb10, Zol93]. circles [Wil24a]. circular [Kra22b]. Circularity [Bak21, Eig19, PE24, RS21a]. circulating [Koc15b]. circulation [DK09]. Circuli [Cam18a]. citizen [Qui21, San22]. civilization [Alg08, Bau71, Nas68, Nas83]. Clagett [Gra72]. Claiming [MR13]. claims [Kin15a, Lac15, Mel91]. Clarifying [Bak23, God14]. Clarke [Bro16b]. Classical [Mul24, Sam10, Buc69a, Buc69b, Cal10, Fey70, McR72, RT95, Sta08, SF21, Vas18, Vor14, Wat04]. classics [Fyf02]. Classification [Blo82a, Ahn20, Atr85, Bha15, Con20, Dio13, Dio18, Dup94, Gua13, Pap24, Pra77, San16a, Smi84, vdLBV22]. classifications [Kle05]. classificatory [Sla17]. classifies [Pol15]. classique [CL05c, Cul06]. Claude [Sch20b]. Clavius [Jar79, Pan96]. claw [Pie14]. climate [Det22, Don21, Gal21, Gra11, JC20, JC21, KB20, Leh18, Llo12, Llo15, MJ23, O’L20, O’L21, Par14, Vez16, Wat23]. clinic [SFT23]. Clinical [MBC17, PM22, Fed23, Hur17, MBG21]. clocks [Aro18, Gio14]. close [Bam96, Cox12]. closed [Böh80, Zan82]. Closing [Kuu09]. closure [Sha97]. Cloth [Ach92, Ano91b, Cra93, Ful93, Hac89, Hop94, Kam94, Mye89, PR90, Sch91, Zol93, Liv93, ML18, Gie91b]. Clothing [Ser23]. cloud [Kuu09]. cloud-cuckoo [Kuu09]. cluster [Lip20]. Coalgebra [Cor11]. coastal [Ale98]. coat [Gol97]. coded [vdLBV22]. codependent [Hoc16]. Cognition [Ada18, Mak03, Pri18, Sch08c, Spr10c, Wei10b, Wra07, vEdJ16]. cognition-knowledge [Mak03]. Cognitive [Ara23, Cam22, CB92, Kra22a, ME15, Rup22, She22, Aiz10, All96b, BA10, Bro11b, Dhe23, PS10, Sam10, SOC19, Spr10a, WHC23, Yea82, vEdJ16, vE19]. Cohen [Ano91b, Gui04a, Ste04, Gio16]. Coherence [Dip06, Kro06, Sch14a]. coherentism [Pap24]. coherentist [Kuu07, Sch18a, SS09]. Coincidence [Wei21a]. Cold [Lin22a, Aro18, Bid11, Pet16]. Collaboration [And16, Ano91b, DB96, FA13]. collaborations [MN20]. Collaborative [Fag15, Lov15, AL16, LGC21, WHK14, Wra06]. collapse [Gao21]. collected [PEK06]. Collecting [Boa07, Win99]. collections [Feo12, Jar13, LW13b]. collectivist [HKMV11, Sch09b]. Colligation [CSD21]. Collin [Elg22, Jhu22, Ode22, Pin22]. Collins [Sor16, Gin09, Weh19b]. Colloquies [dC12]. Colloquium [Ano78b]. Colonial [Leh18]. colonies [Way21]. Colonising [Spa95]. Color [Coc14, Jac95, Cat14]. colors [Leh18, Tur16]. colour [Lau87]. colours [Gol97, Top90]. column [Hau21]. Combative [Goo13]. combinatorial [Kim24]. combustion [Gök90]. comes [Hac91]. comets [BG88]. coming [Gio24a, dC10]. Commandino [Mel89]. Commemoration [Ano84d, RZ21]. commensuration [D’A00]. Comment [Coe22a, Bar05, Sei13b]. commentaires [CL05c, Cul06]. commentary [Gra72, Osl72, Pal09b, Yol70]. commentators [Asp13, Bam96]. Comments [Hes82, Luk82, Car05]. Commission [Hol21a]. commitment [Con06a, Dic06, vL24]. commitments [SFF23]. Committee [May00]. common [Atr85, BC00, Blo08, Fri08, Red02, Sar12, Tuc20]. communicating [Har96]. communication [LL19, PS24]. communicators [PS24]. communist [tF18]. communitarian [Kus02b]. communities [Har19b, Pla20]. Community [Shr07, Aya25, EHPL22, Har93, Hop94, KH25, Kov21, Rot11, Sto25]. community-based [KH25]. Companion [Van18]. Company [Car73]. Comparative [Par20b, Car14, Cur16, Llo97, Lyo16, Möß11a, Möß11b, vdB24]. Comparing [Nor86, War93, dCR24]. comparison [De 14, Gay76, HKMV11]. comparisons [O’L21]. compass [Osl72, Ros19, Yol70]. compatibility [GL24]. compatible [Grü78]. Compendious [Par12]. competition [Dav21b, Tre87]. complementarity [Bel92, Kai92, Mor11, Ueb15]. complementary [ST20]. complete [Dar02]. completion [Mit22]. complex [Bog75, BDB21, Llo12, O’N86, Rie14]. complexes [Gol04, MC25]. complexity [Pen22c, Tab14, Zuc19]. complicating [Hol21a]. complications [DHL21]. components [Ver06]. Composite [Mia01, Gla97]. composition [For88, WW04]. compositional [KD19]. compossibility [Kim24]. compounds [Gay77]. comprehension [Wei10b]. compressible [TGD05]. compression [McA05]. compromise [Coc14, Sha97]. computability [CP10, Res12]. Computation [Aiz10, Len22, LSH24, PS10, Spr10a, Spr10b, Eli10]. Computational [Ega10, Anz19, BA10, Bec16, HS23a, Len14]. computationalism [Pic04, Sam10]. Computer [Fal24, Lus16, Vez16, Bog20, KSZ24]. computers [Len14, Sha10]. Computing [Pap20, Tha93]. Comte [Fra90, McC01, Sch20a]. Comtean [Sch85]. Conant [Bid11]. Conatus [Ber80]. concept [BT09, Chi15, Fey16a, HG21, Hüb80, Jac02, KH17, Kra81, Kub16, Kwa12, Lep75, Lev11, Loi22, Mer82, Nee13, PB96, Sch93, Sha03, Tab14, VABF24, Wag24, Whi71]. conception [Coo20, Gie92, Gio22, HH90, Kra72, Lor13, Lou08, Mäk97, Nee02, RK06, Riv21, Sch88b, Sch18a, Sch88c, Spa90, Wri15]. conceptions [VS12]. Concepts [Kem07, Psi00, And94, Bel90, Che03, Che10, Cor21, De 24, Fra21c, LAL21, Goo21, Hen14a, HS22, Jan13a, Mag96, Mal01b, Mét13, Ner84, PG20, Pip79, Sch70, Sch88d, Sha98, Smi20]. concepts-centric [Cor21]. Conceptual [Goo78, Gut73, Lau88, Vor14, Whi88, Dic25, Gla89a, Goo21, Gra71, Hau21, Jan73, Mer92, Obe05, RT95, Szn16, Van86]. conceptualisation [Cor01]. conceptualization [Cun21]. concerning [Chr94, Rog72, Tho09, Woo71, Woo94]. concerns [VK19]. concrete [Boe22, PvF14]. condensation [Tom22]. conditioned [Gla02]. Conditions [Dic08, Elw12, Sch06]. Conference [Ano86d, Ano92e, Ano95c, Bus71, May14]. confidence [DH22]. configurationibus [Cla68, Gra72]. confinement [Bla22]. Confirmation [Mah99, Rui23, Tre87, All21, Col94, CT14, Daw22, Oni17, Pad15, Pru89, Sch14b, Vot14]. Confirmational [Per08]. confirmatory [Due22, Llo15]. confirming [Pin09]. conflation [ES06]. conflict [Gab73, Sco70]. Conflicting [May00]. conformity [Ste21]. confusion [Hen06]. Congress [Ano89a, Ano91m]. congruences [Ver82]. conic [Pou07]. Conjectandi [Cam18a]. conjuncts [KS19]. conjuncture [Cuo95]. Connecting [Aba10, SFT23, Noa08]. connection [Acu25, Alm12, Axt93, Ver06]. connects [PE24]. conquest [Bal92]. consciousness [Bae24, Man02]. consequences [HM88, Ryc92]. conservation [For75, Gab73, JW21, Nir23, Sco70]. conservatism [Blu21b, Del19, Fur22, Gay77]. Conservative [Cra84, O’C19, Pos71]. conservativeness [Cur19a]. conservativism [Obe05]. consideration [Roc02, Rog72, Woo71, Woo94]. Considerations [MS24, Ili04]. consistency [Dic06, Lau82a]. Consistent [Ruy23, OS14]. constant [Cur92, MM24b, Ray90, Sch23]. constants [Gré19, Qui17]. Constituting [Hol21a, PMG21, Sak22]. constitutions [Dav21b, KM22]. Constitutive [vE19, Kri18]. constraint [Van86]. constraint-structure [Van86]. Constraints [Rit21, Wei99]. constructed [Car15, Nel94]. Constructing [Bow93, dC18]. Construction [Kus02a, Win81, Bie20, Blo08, Cob11, Dom09b, Jac03, KM14, Mor91, Sch07b, Sha98, SL21, Tay09]. Constructive [Con06a, Dic06, HKMV11, Pru89]. Constructivism [Mat98, Mas15, Sie04, Sie10, Sie04]. constructivist [Ban13]. constructivists [Nel94]. constructs [Hur17]. Contemporary [Fre21, And16, GN22, GMC22, Gra71, HS23a, Jan73, Obe05, Pre08b, SBM09, Wit13]. Content [Rup22, Ega10, FL21, Hat88, Neu71, Pre08b]. Contents [Ano84e, Ano96b, Ano01c, Ano02a, Ano03g, Ano04a, Ano05b, Ano05a, Ano06a, Ano80g, Ano81d, Ano82d, Ano83d, Ano85g, Ano87e, Ano88d, Ano89h, Ano90g, Ano91g, Ano95b]. Context [HH87, Ari84, Bel92, Car14, Dav20, DvdB22, Fis18, FP99, Gre24b, Hen22, Iva11, Lee19, O’F10a, Pre06, Ryn95b, Sno04a, SR05, Ueb25]. contexts [HL19]. contextual [Lar24, Pin14, Pol21, SR05, dRM23]. contextualizing [Koc11b]. continental [Fra79a, Nun84, SW12]. Contingency [BM20, Fra08, Tri08, Ayl19, Joy97, Kid16b, Swa24, dCR24, Cus85]. contingency/inevitability [Ayl19]. contingent [Fis85, Kin15b, Sol08a]. Contingentism [Bou22b]. contingentist [Sol08b]. contingentist/inevitabilist [Sol08b]. continua [Koi21]. continually [Zam08]. continuity [Eva10, Sce06]. Continuum [Moy78]. contrasted [Cah86]. contrastive [Cha10b]. contribution [Bai04, Bra15, Car21, FK90, Fay24b, NRB04]. contributions [Aba11]. contributors [Ano71i, Ano72h, Ano73f, Ano73g, Ano74g, Ano74h, Ano74i, Ano74j, Ano75f, Ano75g, Ano75h, Ano76b, Ano76c]. Contributory [CEW16, God14, RGD15]. control [BB22a, BB22b, Cra93, Cro92, Fri24b, Mor91]. controlled [Rup76]. controversial [GN22]. Controversy [Mor22, Bha24, CB92, Díe07, Fre02, Ilt70, MM24b, Mic71b, Miy17, Pap77, Rit21, McE76]. Convention [Ker12]. Conventionalism [Gie92, Iva15, Slo03, Due21, DBM22, Fab19, Gie77, Gie91a, O’G77, Sch20a]. Conventions [Gow97]. Convergence [Cas24, Hoc16, Sch17]. conversation [Ayl19]. conversion [Gor91]. conversions [De 96a]. convex [Acu25, KC12]. Cool [Gre24a, Wes23, Slo24, Wes23]. Cooper [WW04]. Coordination [Ohn21, Luc20, Pad15, Pap24]. Copeland [Res12]. Copenhagen [Osl71]. Copenhagenism [Fre78]. Copernican [Ant03, BG88, Bie20, Van04b]. Copernicanism [Dra87, Pan99, Sch15]. Copernicus [Gra04, GT05, LoL15]. Copying [ES06]. corkscrew [Mor06]. corporate [Boa11]. Corporation [Bri73, May71]. corps [Le 58]. corpuscular [Hal05]. corpuscularian [Mok71]. Corral [Mir05]. Correct [Ups22, Rom16]. correction [BH19]. Correlation [Mor91]. Correspondence [Pos71, Bro73, Bro16b, Clu12a, Hun03, Kuu07, Toa24]. Corrigenda [Ano77c]. Corrigendum [Ano71f, Ano97c, Ano97d, Ano00b, Gam21a, Gam21b]. corroboration [Row08b]. cortical [Hau21]. cosmic [LC21]. cosmogony [Pal87a]. cosmography [Pet11]. Cosmological [Mer21, Wol24, Cur92, Ray90, Sch23]. cosmology [How98, LC24a, McM93, Sch21a, Wen23]. cosmopolitan [Lou08, Vet18]. cosmos [Web07]. could [Lan21, Nel94, VJ19]. counseling [Sti13]. counter [De 96a, Koc06, Mee92]. counter-examples [Mee92]. counter-reflections [Koc06]. counter-reformation [De 96a]. Counterfactual [Pin22]. Countering [UG20]. Counting [Koc09, Sch11, Koc11b]. Countries [Van04b]. counts [BB23, Lar24]. course [Lac15]. Court [DK09]. Courts [De 96a]. covariance [EG78]. cowpath [HS23b]. Cowperthwaite [Jan73]. Crafting [Set08]. Craver [Lan18]. create [Kal96]. creation [Ast04, Cha44, Cha69, Ner84, Wra19, Ege70b]. Creativity [Cur19a, SD23, Cur19b]. Creature [Cun01b, Cun00]. credit [Hee19]. Criminal [Hag01]. crisis [Sno04b]. criteria [Fag24, FB20]. Criterion [Mul24, San11, San12]. Critical [Wat03, AB23, Fab19, Gay76, Kra22b, Lar24, Lau23, Pin14, Pit22, Rab75, Rod21, Sha97, Sha20, SR05]. criticality [Fri03b]. criticality-what [Fri03b]. Criticism [LM70, Daw22, Hor73, Lar24, Pri18, Row11a, Row13, Fin73]. criticisms [Blo07, Lor13]. criticize [DK22]. critics [Bar18, Hor73, McH21]. Critique [Myk22, Ahn04, Bre04a, Coh97, Feh07, GVD24, Goe72, Kem05, Kus10, Lor13, Nob23, NI03, Oka00, Pic22, Rib85, Sce05, Waj21, Nik19]. Critiquing [MHW22]. critters [Cur18]. Crosland [Cra93]. crossroads [Dar03]. crowd [GID21]. crucial [Lak74, Qui74, Wor82]. crucis [Lay78]. Cry [Goo04]. cryptanalysis [Pan06]. cryptographic [Mat19]. cuckoo [Kuu09]. Cultural [Gee14, BS13, Joh98, Mak08, bAM11]. culture [Bre04b, Dav21b, Dea03, Kid16a, KM22, Ort08, SBM09, Vic21, WG13]. cultures [BK17, Kle01, O’C05, Spa95]. cum [KR08]. Cunningham [Dea01b, Dea01a, War92, War93]. Curie [Kin21]. curious [Ekb24, Mei21, San09a, Wil09c]. current [Hon87b, Mir05, Pic84]. currents [Cam18b]. Curry [Ray92]. curvature [Bar21b]. curve [Pic22]. Cusa [Mou20]. Cutting [McO01]. Cuvier [McC01, vdB24].

D [Bro71, Car05, Hin12, Mor90b, Des49]. D. [Car05]. DADTA [Ull24]. daguerreotype [Tre07]. dairy [Bru20]. d’Alembert [Bri73, Ilt70, Swi86]. Dalitz [Mel04]. Dalton [Zwi11]. Daltonian [Whi90]. Daniel [Gow87]. Daniell [Sut71]. dans [Le 58]. daring [Fur22]. dark [DW23, MSS^{+}22, McG21]. Darrell [Ara23, Hen23, Miy23, TA23]. Darstellung [Sch18b]. Darwin [Hes21, May71, Ari22, Buc21, Dar76, DD69a, DD72, Dou22, HBR24, Jar09, Kle79, Len21, Liu22, Man80a, Mei21, Ric97, Roc88, Rud74, Rus75, Rus78, Ste22c, Ste24b, Tha77, Yea85, May71]. Darwinian [Ege71, Bar22c, Ghi69a, Ghi69b, Ghi72, Ghi84, Ghi03, Kan96, Tur22a]. Darwinism [Cap79, Fre21, Mor93]. Darwinnovation [Lew02]. Data [Gre24a, WL22, Zah19, CL22, Fal24, KH25, KB20, Leh18, Llo12, Lus16, McA03, McA05, Nor20a, Rap80, RM16, Rup76, TGD05, vL24]. data-driven [KB20]. Data-intensive [WL22, CL22]. David [Ere91, Buc82, CB92, Hes82, How15, Kem05, Luk82, Rus89, Tal25].

Davidson [Bro71]. Davies [Car73]. Dawkins [HM88, Mei21]. dawned [Cra22]. day [Cra22, Rap11b]. De-synthesizing [Ueb12]. Dead [Rot96]. Dear [Cun01a]. death [Ano84d, CM00, FJ24, Hau21, Kon24a, Rob95]. debate [Ach92, Ano91k, AP06, Bor15a, CD15, Cat95, Che95, Cob11, Cra84, Dhe23, Dil90, DW23, Fra81, Gat16, Jac21, Men18, Par95, Rui23, Sny07, Tur21, Ver07b, WHC23, Wei07, Hen23]. debates [Lip20, Llo12, Löw08, Nye76, SL21]. debt [Mas11, Mil08, Rus75]. Debunking [LK20]. decay [Car73, Her69a, Her69b]. December [Ano84g, Ano22g, Ano23k, Ano24g]. deception [FL21, Her15]. Decision [Rus23, Wei07, WS21]. decision-making [Wei07]. decisions [Maz22, Ray92]. decisive [Lau87]. decline [Jes99]. decolonial [Har19a]. decomposition [Bor25, Dou95, Miy15]. decrees [Pal09a]. Dedicated [Ere91, Cha10a, Rus89]. dedicatory [GT05]. deduction [Dor74, She06, Zel24]. deductive [Wat12]. deductivist [Cap79]. Dee [Alm12, AD12, Alm12, Cam12, Clu12a, Clu12b, Feo12, Joh12b, Man12a, Par12, Pum12, Ram12a, Ram12b]. Dee-Pedro [Alm12]. deep [Cam18b, CT16, Tam20, Wit24]. defaultism [Ayl19]. defects [ND22]. defence [Cox75, CS17, Fey70, Fre78, Har74, Hor73, Run94, Sar83, Tos03, Wal12]. Defending [Fre19, Oni17]. Defense [Ode22, Fri14, Mar13, PSS18, Ris90, Wra15]. Defining [Han06, Mal01b, Rot11]. definition [Atr85, Bla24, LAL21]. definitions [McO96]. Deflating [Pet16]. Deflationary [Suá15, LL19]. degenerate [Ant25]. degeneration [Mor93]. degree [Wal06]. deity [PEK06, Pal02]. Delayed [Mje24]. Delayed-choice [Mje24]. Delboeuf [Fay24b]. Delegation [Boh00]. delicate [Mac13]. Delirium [FJ24]. della [Man12a]. Delphi’ [Mül12]. delusive [Wys23]. Demarcating [Ful24, Cho22]. Demarcation [Yea85, FB20, HW22, Joh21a, Res00, Reu22, Wag22, vEdJ16]. democracy [Kit01, Koc04, Lus21, Sor16, Jas04]. Democratic [Pin16, Por01, Chu22, Koc04]. democratization [LF25]. demon [Koz15, Col90, Dau70, Hei70]. demonstrated [Le 58]. demonstration [Gol04]. demonstrative [Feh82, Jar76]. démontrées [Le 58]. denial [Goo78, Han17]. Dénis [Osl71, Ano84d, Lux67]. Dennett [Mul24]. denotation [Gel11a]. dependence [BM21, Dar23, Mas15, PM23, Wan22, Wer80]. depiction [Ste05]. deployment [Lyo16]. Depression [Kon24b]. depths [Far07]. derangement [Rou05]. derivations [Ach86]. derivatives [Fri21]. derived [De 24]. Descartes [McR72, Abr11, Bab25, Buc69a, Buc69b, Cla79, Cro91, Des49, Dom09b, Fun75, Hat79, Hat07, Jan15, MM06, Man12b, Nob23, Non94, Osl92, Sch15, Sha74, Slo77, Tie88, Zel24]. describe [Eli10]. Describing [VFHS23]. descriptions [Kro06, Kut22]. descriptive [Cam21]. Desert [Blu21b]. Design [Fan24, Wis13, BB22b, Buo24, Dip06, DS21, Meu19, Nii14b]. designs [Mac13]. Detail [Mat20, KM14, Pen22a]. Detecting [Kos06]. detection [Eld25]. determinacy [Cat12]. determination [Bol23, Lud16]. determinism [Bor15a, Koz15, Sar15b, Vas18, vS14]. Development [Ano73a, MS24, Ull24, Ano90f, Eve18, Fek12, Gau98, Gla93, Goo87, Hen11, Krü80, Mag96, Mak08, McG69, Moy91a, NRB04, Rus75, Sha02, Tho15, Van95, WL22, Zap24, Zem06]. Developmental [Pen22c, KT21, Met91]. developments [Man92, Woo11]. Deviant [CP10]. device [Jen10]. devices [Ano91k]. Dewey [Bro15, DB96]. Dfl [Ere91, Zol93]. d’Histoire [Sig90, Spa91]. diabetes [Sch93]. Diagnosing [Buc21]. diagnosis [Wel13]. diagnostic [Cal24, SFT23]. diagrammatic [Geo14]. Diagrams [Cha19b]. dialectic [Ede09]. Dialogic [Mül12]. dialogue [Ano91l, LGC21, MHW22, Mar13]. dialogues [Bel92]. Dice [vE14]. dichotomy [Bro15]. dictatorship [Joh19]. Did [Bro16b, CD15, Fra93, Tom22, Che03, Chi16, Gin07, Lar03, Lau23, Pal05, Sch02b, Stu08]. Diderot [Osl71, Ano84d, Lux67]. difference [Con00, PS10, Rhe92a, Rhe92b]. differences [Cal10, Che03, Tra22, War23]. different [Che10]. differentiation [SD25]. difficulties [Sol08b]. difficulty [Kla00]. difformity [Cla68, Gra72]. diffraction [Sch09a]. diffusion [HD19]. dilemma [Gré19, Nye76, Wes72]. dilemmas [VABF24]. Dilthey [Ham23, Mak03]. dimension [Geo14, SG85, Tei06, Ueb05b, Wra07]. Dimensione [Cam18a]. dimensions [Mir04, Whi88]. Din [GS22]. dinosaurs [Car05, Tur16]. Dioptrice [KR08]. dip [Jar89]. Dirac [FG22, Hen83]. direct [BZ86, Eld25]. directionality [CP17, Hol21b, Lan18]. directions [SLH23]. Disadvantage [Por01]. disagree [Fau08]. disagreement [Cle23, DD13, Lic21, San09a, WB20]. disagreements [Dou10]. disasters [RZ21]. disavow [Bro16b]. discernibility [Brä24b]. Disciplinary [Bus71, Ben18, Ric09b, Sta08]. Discipline [May00, BK17, Kal96, May04, SFF23, War89, dC93]. Disciplines [Len14, Hum19, Ili04, Vor14]. Disciplining [Bru20]. discontents [Ful89, Gie91b]. discorsi [Pru89]. Discourse [Bie20, PG20, Alt12, Kau08, Yeo89]. Discourses [Coo95]. discover [Bec16, Fra94b]. discovered [Aga77]. discoveries [Kal75, TC21, Van86]. Discovering [BB22b, GW15]. discovery [Ara92, Bam96, Bra24a, Dar76, Fis18, Fra79b, FWW09, Gay78, Gla01a, HH87, Pic84, RS21b, Rup76, Sch74, Sha89]. discrete [KC94]. Discussion [Aga71, Ano98n, Ano98o, Ano99l, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano00c, Ano00d, Ano00e, Mic71a, Sch11, WH14]. discussions [Woo15]. Disease [Pen22c, Bla22, CCB24, Ell21, GMC22, PM23, Ric05, vH91]. Disentangling [Luc20]. Disjunctive [McL06b]. disjunctively [Cra84, Fra81]. diskussionen [KMS73, Whi74]. disorders [Kon24b]. Disparities [Gre24a]. dispersed [Jar13]. Displayers [Mer04]. disposition [Mal01b]. dispositions [Cha19a]. dispute [Nel94]. disputes [Alm21]. Disregarding [Ach23]. Dissertatio [KR08]. dissimilarities [Bas92]. dissipative [CN83]. distance [Bro16b, Duc11, Goo78, Hen11, Hen14a]. distant [Mir04, Sch11]. distinct [Gun18, Ueb05b]. Distinction [Win99, Mak03, Nee08, RK06, dJ97].

Distinctively [Hol21b, CP17, Lan18, Lan21]. distinguishability [Fri14]. Distinguishing [KR22]. Distortions [Elg22, Ode22, Pin22, Ric22, Jhu22]. distributed [Cra84, Fra81, Tha93]. Divergence [DCJ^{+}21, Kle77, Pic03]. divergent [Rit21]. Diverse [Tro95, Pot15]. Diversity [SFF23, Lip20]. Divide [Cuo00, EHPL22]. divided [Ben87, Hac89, Pan99]. Divination [Cab20]. Divine [Kim24, BMBM20, Fun75, Pal09a]. divinity [Sch88a]. divisibility [Jan00]. DM [Whi74]. Do [BM23, Lon01, Pap77, Spe04, Col84, FH84, Goo86, Gra11, Hau13, MR19, Mir89, Ose22, Reu22, RM16, Szn16, Wil01]. Dobzhansky [Kon24a]. Doctrine [Isa19, Pri18]. document [Hut04]. documentary [Fin89, Sha90b]. Documenting [Dur13]. Documents [Pal87b]. Does [GID21, Lus21, Men14, Smi13b, CT13, CS21, MN18, Man72, ZZW20]. doesn’t [vdLBV22]. dogma [Cam21]. dogmatic [Ahn04]. dogmatism [Row11a, Row13, Zan88]. doing [Gal21]. dolphins [Jen07]. domains [Kim00, PE24, Pol15]. Domenico [Car03a]. Donkey [Bau15]. donne [Le 58, Mos01]. Don’t [She93]. door [Gie09, Kuu09]. Dordrecht [Ere91, Sie04, Sno04a, Zol93]. Doris [Mis94]. double [Dai19, Fal98, Nap21, Web22, Cho22]. double-vision [Web22]. doubt [Wys23]. Douglas [Ell13]. Dowe [McA05]. Down [Hes21, HBR24, Tur05b]. Dr. [Fre79]. Drabkin [Ege70a]. draft [HH95, Mor22]. Drake [Ege70a]. dramatic [Buc09]. drawbridge [VH06]. Drawing [Lor20, Riv21, Nas11]. dreamed [Sch02b]. dreams [Web07]. Dresden [DK09]. drift [Fra79a, Nun84, SW12]. drive [CS07]. driven [Don21, HS23b, JW21, KB20, WHMO22]. drivers [KM14, Rei20]. Drugs [dC12]. Dual [HKMV11, TC25, KM06, Vae11]. Dual-Nature [HKMV11]. duck [Sla17]. due [Joh21b]. Duhem [BV05, Bha15, Cok15, Dio13, Dio18, Dor79, Iva10, Iva15, KS19, Kid11b, Kri82, Nee98, Nee02, Stu07]. Duhemian [Bha17, Dar02, Nee96]. dummy [Tre87]. d’une [Le 58]. Dunod [Cul06]. durée [Edi22, Mel21]. during [Car15, CB92, Dil90, Ram92, Ste22c]. Durkheim [Blo82a, Sch85, Tur11, Ver82]. duro [Pal05]. Dusseldorf [May14]. Dyck [Pal09b]. Dynamic [BA10, Nea24]. Dynamical [TC25, Gio14, Kav24, Mas11, Sim70]. Dynamics [Bru99a, Bru99b, Acu25, Bec84, De 00, Fee16, Fri12, Gab71, GS02, Gla86, Le 58, Mor12, Moy77, Ste03]. dynamique [Le 58, Le 68, Bri73]. dystopian [dC18].

E. [Ber85, Dou95, Par95]. E.-J. [Dou95]. each [Le 58]. eagle [LM12]. EAHP [Ano95c]. Early [Abl15, Bai04, Bel20, Gol23, OHH00, Oos12, Ahn04, Bie20, Boa11, Bor15b, Bur02, Cam97, Cam12, CB92, Che95, Chi08, Cul04, Das86, De 96b, Dea87, Dom09b, Fis06, Gal21, Gio24a, Gla20, Gue17, Ham23, Har96, Jac97, Ker87, Kob24, LS02, Pal03, Pal70, Pen15, Pou07, Pre08b, Ram12b, Rap11a, Rot11, Ser99, Sha02, Sib95, Tur22a, Wat03, Wat13, dL09, Win99]. early-modern [Ser99]. early-stage [Gal21]. Earth [Car73, CN83, FG22, O’C05, Ros18, Car05, Her69a, Her69b, Ohn21].

Earthquake [Aro18]. earthquakes [tF18]. East [Ano91k, Low93, AR22, Edi22, Rob22]. easy [Räz24]. EAT [Saa05]. echo [Ueb05b]. echoes [MD13, Mir04]. Eckart [Hun21]. Eclectic [Bus08, Saa08]. eclipse [Mor93]. ecofeminist [RZ21]. ecological [Alm21, BZ81, Dou20, Kwa12]. Ecologist [Cun01b, Cun00]. ecology [DS19, MC25, Tra22]. Economic [Düp15, Han21, Maa99, Coo05, Eva07, Gua01, Her15]. economics [BK17, Han21, KM14, Lar24, Lis18, Mäk97, MM09, Mir89, VFHS23, Jhu22]. economy [Bak22, Car03a, Den00, Hol83, Tre07, Wil77]. Ed [Sie04, Van18, Bus71, Elk74, Ere91, Gra72, May71, Mis94, Whi74, Zol93]. Edge [PMG21]. edged [Fal98]. Edgeworth [Yee24]. Edison [Wil09c]. edited [Eas72, Gra72, Sha90b]. edition [Bri73]. Editor [MTC17, Ano75d, Ano01d]. Editorial [Ank23, Ano80d, Ano17e, Ano17b, Ano17c, Ano17d, Ano18a, Ano18b, Ano18c, Buc82, Jar03a, Ano70e, Ano71g, Ano72e, Ano73d, Ano74e, Ano75c, Ano76a, Ano77d, Ano78c, Ano79e, Ano80e, Ano81b, Ano82c, Ano83b, Ano84f, Ano85e, Ano86e, Ano87c, Ano88b, Ano89f, Ano90c, Ano91e, Ano92f, Ano93f, Ano94d, Ano95a, Ano96c, Ano97e, Ano98p, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano03e, Ano04c, Ano04d, Ano04e, Ano04f, Ano05f, Ano05c, Ano05d, Ano05e, Ano06b, Ano06c, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13a, Ano13b, Ano14b, Ano14c, Ano14d, Ano14e]. Editorial [Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano18f, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano25a, Ano25b]. editors [Ano89g, Ano09f, Mye89, Ano70f, Ano70g, Ano72f]. Eds [Gui04a, ML18, Mel04, Sno04a, Bro71, Hei75, Rhe94b]. Eduard [Sch93]. education [Gla93, Hei24, Kid16a, Mat98, Sie04]. Edward [Bie20]. Edwardes [Clu12a]. Edwin [Cha20a, Cha20b]. Effect [Sim14, Ara92, Bou21, CN83, DJ14, Gra24, SHB94, Str06]. Effective [BCK^{+}22, KF23, Riv21]. effectiveness [Bok15, HK21, Ste22a]. effectivize [Che22]. effects [Car03b, DH22, GVD24, Ryn95a, Ryn95b, SBM09]. efficacy [Maz22]. egg [Wit13]. Ehrlich [vdBG90]. EHT [Wei21a, Wei21b]. eighteenth [Bai04, Car03a, Fer01, Ric92, Spe17, Ter17, Vos07, dC02, vdB22]. eighteenth-century [Bai04, Car03a, Fer01, Vos07, vdB22]. Einselection [Kas14]. Einstein’ [Hen85, Dar21b, EG78, Fey84, Gio14, Hen92, Hoe94, How85, Kat92, Mau90, Nor85, Obe16, Ray90, She88, Sla22, Tom22, War92, War93, Wil77, vD21]. Eisenthal [Coe22a]. either [Hon87a]. Electrical [AG13, Ker07, Mit17]. electricity [Ara13, Sim70]. electrodynamics [For22, Wil77, Gur19]. electromagnetic [And94, Cha86, Hon87a, HS23c, Joh21b, Mit17, Mor92, Sim70]. electromagnetism [Cha73, Sha89]. electron [Ras93, Ras96]. electrons [Car05]. electrostatic [Dor74, Goo78, Hon87a]. electrostatics [Fal17]. elemental [WW04]. elementary [Mit17]. elements [Bau72, Bro76, MS72, Mit72b, Sce06]. elephant [Tam23]. Elias [Heu14, Alg08, Feo12]. eliminating [Nis93]. Eliminative [Bir10, Fre19, McC21]. eliminativism [Wol19]. Elizabethan [Ale95]. ellipse [Bai90]. Elsevier [Car73, Gab73]. elusive [CL05b, Ste03]. embarrassment [Ter05]. Embodied [Bud13, Wei10b, Tam23]. embodiment [SDC07]. embody [Men14]. embrace [Del19, Pee17]. embryonic [Fag16]. Emergence [Col21, Eat23, Plu78, Cur89, Edg08, Gla78, Hen21, LL16, Nor21b, Sci21, dCR24, BR23, Cle23, Eat23, Gol23]. Emergent [Sar15b]. Émilie [Hut04, Rei18]. emission [CB92]. Emotion [Cam97, Koc13]. emotions [Bra24a]. Empedocles [Alt12]. Empire [Mor90a, Cha13, Jac94, SW89, Ram12a]. empire-building [Cha13]. empiric [BC92]. Empirical [Bue97, Cha04, Fri08, JB16, Kui14, Mat22b, TGD05, All21, Bar18, Blu21b, Daw21, Daw22, Eng18, Fri02, GN16, HC23, Lus16, McA03, McA05, McC21, McN15, Nic72, Pal08a, PS24, Pip79, Prz69, Rit21, Rui23, RDM25, Sch24, Sch08c, Wol24, vD21]. Empiricism [Bue00, Dil90, Con06a, Gla01b, Lar24, Llo12, Pin14, Ryc92, SU15, Sch11, SR05, Ueb05a]. empiricist [Di 11, Dic06, Ueb25]. emptiness [Non94]. Empty [Fin90, Sch23]. enabling [Ano91i]. enclose [Has97]. encodings [CP10]. encounter [MBG21]. encounters [Pri22, Sim17]. encyclopaedias [Olm12]. End [Win99, FK90, May04, Tot12, Yea05]. ending [Tur11]. endophenotypic [Koi21]. endorsing [Lac15]. ends [McO96, Rot96, Tur13]. Energies [Isa19, Wal22c]. Energy [Moy77, SW89, All96a, For75, Joh21b, KSZ24, Nir23, Mor90a]. enframing [Kho03]. engage [Pla20]. engagement [LF25]. engineer [Mor21]. engineering [Dip06, Gal21, Ver06, Wan22]. England [Bur02, Pum12, Sar89, Ser99, Sib95]. English [Gra72, Mor21, Ram12a, Sha02]. enigmas [Wis13]. enlightened [Car03a, Lau87]. Enlightenment [Van04a, Dea12, Stu91b, Ahn04, You04, Zin11]. ensembles [JC20]. Entanglement [FL22, Brä24b, Con20, Fri14, Mje24, RE19, dRM23]. enterprise [Cha13]. enthusiasm [You04]. entirely [DBM22]. entities [Mac72, Mar04, Rhe08]. entrepreneurship [Bak22, Goo13]. environment [BB23, CL18, CL22, Fra21b, Loi22, Mel21, Pen22c]. environmental [Cre18, Fra21b, GID21, GMC22, LGC21, MN18, PP21]. Environmentality [FBW22]. environments [Mel21, ND22, Sie10]. Enzyme [Con20]. Epicurean [Bog75, Sch13b]. Epicurus [LoL15]. Epigenesis [Zam16, HN16, Zam03]. epigenetic [LAL21, Smi83]. Epimutation [LAL21]. Epinomis [Leh20]. Epistemic [Bol23, BK21, Ell12, Man02, Pic03, San11, Wil22a, dG20, Ahn20, Bal20, Bar22a, Bla13, Bol16, Bue22, Cam15, Car14, Con20, DD13, Fai12, GID21, Käs18, Kid11b, Kid16b, Nor98, PE24, Qui74, Reu22, Rom16, SD23, San13, Sch05, SFT23, Sei13a, Sei13b, SFF23, Zah19, dB09]. epistemically [Ahn20, Reu22]. Epistemological [Kri82, Rou05, Gau76, Gie91a, GK88, Hon89, MCGP17, Mot02, Van09, Yea82, dC10, dCG17]. epistemologically [Let21, Ram92]. epistemologists [Vas00]. Epistemology [Cok15, Den00, And16, Ano91h, Bid14, Byc12, Cer19, Chi15, Col02a, Cur19a, De 18, Esc08, FJ24, FH88, Fre05, Fre15, Gel10, Haz23, Hoo13, Kit87, KF23, Kow11, Kus01, Kus02b, Kus10, Kuu07, Kva14, Lau90, Lau87, Lip98, Loi16, LJ11, Mag96, Mét13, PG20, Ras93, Res96, SU15, SS09, She88, Sim17, Smi83, Stu20, Stu07, Tuc12, Ueb91a, Vez16]. Epperson [Heu14]. EPR [GG94]. equably [Art95]. equalities [Cat22]. equilibration [Kit87]. équilibre [Le 58]. Equilibrium [Val21, dB09, Bol15, Hep10, Pal08a]. equiprobable [KS19]. Equivalence [Due21, Boy14, Eve18, Nor85]. equivalent [Aiz10]. era [vdBG90, Ped20, Myk22, Ups22, Val22, Kuk22]. eretico [Red87, Sha90a]. Erhaltung [Hei74]. Ernst [Bak23, Bia20, Bla72, Hei14, Swo74]. erotetic [Hol21b]. Errata [Ano88c, Ano04g]. Erratum [Ano78d, Ano83c, Ano85f, Ano86f, Ano91f, Blu21a]. error [Chi16, Hon87a, Sch05, She97]. errors [Cat14, Hon89, Sch05]. escrits [Des49]. especially [Mah77]. Essais [Hom02, Kim00]. Essay [Bar20a, Cer19, Cul06, Gur19, Nik19, Pes13, Sha76, Wra19, Yol70, Bai87, Rog72, Sol95, Ver07a, Woo75, Woo71, Woo94, Osl72]. Essays [Bro71, DH69, Ere91, Tru68, Cha10a, Rus89, Wig67, Wig70, Wig78, Kir73, Ait70]. essence [McK95]. essential [HHL11, Sim21]. essentialist [Lip20]. Essentially [Rot17]. est [Jan13b]. Establishing [Car93]. establishment [War89]. estimation [EC25]. estimative [Qua75]. eternal [Fun75, Sch03a]. ethical [Coe22b, Don21, Gal21]. ethically [JW21]. ethically-driven [JW21]. ethico [D’A00]. ethico-political [D’A00]. ethics [Fra21b, Res96]. ethnobiology [GR23]. Ethnographic [Cur16]. ethnography [Wil22b]. ethological [Kal75]. ethos [Chu22]. Étienne [Hof13]. Étienne-Jules [Hof13]. ETTLG [Lee19]. Euclid [Kno76, O’G78]. Euclidean [Ekb24, Gri91, Joh12b, Man92]. Euclideanism [Grü78]. Eudoxus [Het96, Wri73]. Eugene [CF21b, Wig67, Wig70, Wig78]. Eugenics [Max22]. Euler [Gio24a, Hep10]. EUR [Pes13]. Europe [Dai19, Gue17, Pan99, Rot11, Vos07, Win99]. European [Ano95c, Koc24]. evaluated [Pal87a]. Evaluating [Kov21]. evaluation [Fre80b, Leu15, Lyo16, Mac13, dCG17]. Evans [Sor16, Weh19b]. eve [Pan96]. Even [TA23]. event [Che03, WOL20a]. events [Pao11]. ever [Pal05]. Everything [Gin07]. Evidence [Cab20, GY11, PG24, Ste22a, Abe22, Ach23, Bas03, Ben22, Buo24, Bur02, Eld24, FJ24, Kin15a, LT20, MR19, Maz22, Mje24, PEK06, Par95, Pas11, Swa24, Tho09, Tor24, Vez16, Mil14]. evidence-based [MR19, PG24]. Evidences [Pal02]. Evident [Big08]. evidential [McA96]. evolution [Abe24, Ano92e, Bec84, Ghi94, Gla20, Hen84, Hul89, Jen10, Kal75, Obe15, Pen15, Shr07, Zap24, ZZW20, Rus75, ML18, Ere91]. Evolutionary [Lew02, Ano90e, Bou22a, Bra78, Byc12, Cam97, Col02a, GY11, KN89, Kil99, Tho83, Van86]. evolutionism [Sar15b]. evolutionist [Los77]. Evolving [Ghi94]. ex [Mer82, Wis84]. exact [Mäk97, Qui18, Ris93]. examination [Cox12, Mäk97, Mat98, Mit17, Rus75, Sha97, Sie04]. examine [RZ21]. Examining [Mee21, Smi20]. example [Boe22, Cor21, Hei80, Van86]. examples [Bec71, Bir10, Mee92, Wal22a, Wil24b]. excellency [Cha93]. exceptionalism [Ueb25]. exchange [And94, GY11]. exchanges [GYM14]. excitability [PM23]. Excited [FJ24]. excluder [Jen10]. exclusion [Leu15, PSS18]. exegetical [Bie16]. exemplar [BA10]. exemplars [Goo21]. Exemplification [Mor19]. exist [vdLBV22]. Existence [Gao21, Ell76, LW94, Pal02, PEK06]. Existential [Cur19b]. Exner [Nir23]. Expanding [Kon24b]. expansion [Pet11, dC12]. expect [BR19]. Expectation [Cro90, Cro91]. experience [BC92, CL18, Dea87, Kra72, Kra73, Leh18, Mac73, Red02, Sar89, Sta08, dC18]. experiences [IRDV24]. Experiencing [Cam18b, CL18]. Experiential [Vet18, GP10]. Experiment [Cha02, Rhe92a, Rhe92b, Van16, Ara92, Cat12, Cha16, Dor74, Dör94, Düp15, FH88, Gla83, Har96, IRDV24, McA96, Nea24, Pal05, Pru89, Qui74, Sar89, Sch18b, Wei21a, Wei21b, Wor82]. Experimental [Her15, KSZ24, Nor00, Sto09a, Baz88, Boa07, Boa11, Bra08, Cos98, Fis18, FH88, Goo78, Gro90, Hon87a, Hon89, Jac94, Kle01, Mét13, Mor21, PvF14, RH95, Sch88a, Sto09b, WHMO22]. experimentalism [Mas11]. Experimentation [LC21, LC24a, MW25, Weh19a, Gla99, Gur19, Mor21, Pfe12, Sch16, Sch13a, Rhe25]. experimenter [Col94, Col02b, Fee16, Fra94a, GG02]. Experiments [Cha13, CERH06, Bak23, Cam15, Col84, Due22, FH84, GW15, Gur19, Hac91, Hen92, Hon87b, Lak74, LC24b, Mil19, Mje24, Sch08a, Stu16, Stu20]. experimentum [Lay78]. Expert [MJ23, JC20]. Expertise [Chu22, CE15, CEW16, Gol22, Joh22, Lea22, PK14, Sor16, WV22, Ara13, CERH06, CS07, CEG07, God14, Jen07, KM22, MN20, Pla20, RGD15, Rib07, Sar89, Sch07a, SDC07]. expertises [CS07]. experts [Gun18]. explain [CT13, Fre05]. Explaining [Mar88, Pra74, Rod21, War23, dR98, BG23, Lov15]. explains [Fum20]. explanandum [Car15]. explanantem [Car15]. Explanation [Asp13, Cha10a, Elg22, Hin13, Koe95, Ode22, Pin22, Ric22, vdB24, Ach92, Bak21, Bea17, BA10, Cab20, CM22, Cel08, Cha10b, CDC24, Dav21a, De 11, Dea12, ET20, Fag15, Fis06, For88, Fra21a, GFE23, Gel10, Gij13, Gri10, Hal05, Hea94, Hen22, Hol21b, Hun18, JSZ23, Jhu22, Ken12, KVJ22, Lip91, Lov15, Mat20, McL06a, Mor17, Mou81, Oka00, Pes10, Ris98, Rot96, Str13, Tos03, Tur13, Vog10, Wal12, Woo15, Wri15, dR13, dR06, vEdJ16, Pin22]. explanations [Bog20, Bra81, Cha12, CP17, EHPL22, Ger14, HH00, Käs18, Lan18, Lan21, PP23, Res89, Rot17, Stu08, VJ19, Waj21]. Explanatory [Wil24a, Cam97, Dou20, Ger14, Hol22, Jan13b, KG13, Lov15, MM09, Ord16]. Explicating [Fre15]. Explicit [Gua13]. explorations [Kwa12].

Exploratory [Gur19, Sch16, Due22, Mor21, Sal17, Gur19]. explorer [Web07]. explosion [CP07]. Exposing [Leu12]. expressibility [Per91]. expression [Bol23]. extended [Spr10c, vEdJ16]. Extension [Ban13, BK21, Gra04, Jac20]. Extensional [Par16]. extracts [Pri22]. extraordinary [dC02]. Extrapolating [Bae24]. extrasolar [Kos06]. extraterrestrial [Kuk01]. eye [vE19].

F [Rus75, Sut71, Whi74]. F. [Hin12]. Fabri [Dra75]. Fabrizio [Ros19]. Face [Spa91, Sig90]. faces [Col90, Han21, Sak22]. facet [Mol82]. Facing [Sch17]. fact [Den00]. factor [Lin98, Pie97]. Facts [Ras93, May14, Mur89, Nel94, Nun84, UG20, dC02]. factual [WB20]. fail [Che03]. failed [Pou00]. fails [Miz12]. Failure [WV22, Fra08, Ker07, Kuu16, Rud74, San91, Wan22]. fairness [DZ14]. fall [Dra75, Jes99]. fallacy [Dic13, Men14, Sch17]. Fallibilism [Stu91a]. fallibilist [Gla01a]. Falling [Bie20]. falsa [Wis84]. falsification [Man12b]. Falsificationism [Díe07]. family [Man80b, VFHS23]. Faraday [Gur19, And94, Goo78]. Farewell [Lan06]. Fascimile [Ege70b]. fashion [How04]. fate [Mit72a]. favor [Hau13]. favour [dC10]. favourable [Cuo95]. Fawcett [Ano91b]. FCC [Gua01]. Feature [Wan22]. features [LA22, Lus16, Spa90]. February [May14, Ano23l, Ano24k, Ano25d]. Federico [Mel89]. feeble.’ [Hon87a]. Feenberg [Koc06]. Fehér [Kem07]. Felix [Gla93]. Feminist [Pot12, Lar24, Ric09b]. fermentation [Tem86]. Ferrari [Fri12]. fertility [Rit21]. Fetal [SL21]. fever [Bla22]. Feyerabend [Bro16c, BK16, Col16, Gat16, Gie16, Hei16, Hel16, Hen85, HH95, HH06, Kid11a, Kid16a, Kle79, Kub16, Kus16, Mac73, Mar16, Mun16, Net91, Net93, OHH00, Obe05, Obe16, Pre98, Pre16, Sha17, Sha20, Sor16, Tam15, vS20]. Feynman [For22]. fibre [Jac23]. Fichte [Sie07, Ste03]. fictionalism [GL24]. Field [BCK^{+}22, DM13, Bid13, Blu21a, Cal24, CT14, Gli16, Har19a, Jar17, Joh21b, Kin15b, KF23, Koc15a, Mit17, Moy78, Riv21, Seg00, Sim70, Sto09a, Vet18, Wal22b]. fields [CF21a, Gök90]. fieldwork [Ben18]. fiend [Sny07]. figs [Ait70, Gab73]. Figural [BCC87, Mye89]. figures [Fer01, Lor20]. filling [Bus08]. filosofia [Bal92, Gia94]. filter [Pas05]. financial [VFHS23]. find [Le 58]. Finding [Ken23, Mor06, Col03]. findings [Grü78, O’G78]. Fine [Ano81c]. Finetti [Gow97]. fingerprints [Tei22]. Finite [AB23]. Finite-size [AB23]. Finland [HM24]. Finocchiaro [Sha90b]. Fire [Lea98, Cha10c]. First [Mah77, Neu71, Rat19, Haz23, Hun03, May11, Pic84, Tre07, dC02, Goo13]. fish [WH14]. Fisher [McL24, Ros17]. fisheries [Van95]. fishery [Jen07]. fishing [EHPL22]. fit [Zam03]. fitness [Cam83]. Five [Lau87, tV21]. fixing [Wal24]. Flamsteed [Hun03]. flattening [Ohn21]. Fleck [Löw08, Möß11a, Möß11b, Sci21, vdBG90]. flesh [Dem06]. flew [BS21]. Flexner [Hei24]. flood [Lea98]. floor [Fra84]. floresiensis [DD13]. flow [Bab25, BM14, Hin15, Joh21b]. flowing [Art95]. fluctuations [Val21]. fluid [Blo08]. flux [Bel90]. fluxions [Art95, Mor22]. focus [Oli92, Mis94]. fog [POW16]. Foibles [Rot22]. Folk [Kno09, Kus97]. following [Sie10]. food [HM24]. Foothold [Goo21]. Footnotes [Gat16]. for-whom [Del23]. Force [Gab71, Hat79, Sno04a, Eis21, EC25, Rui23, Smi13b]. Forces [Wat03, Ell76, Gam21b, Hei70, Pou07, Wal16]. forecasting [Eva07, Sti13]. forefather [Ose22]. foreword [Eas72]. Forging [LB94, Bow93, Jar79]. forgotten [Fay24b, Ker12, Ueb91b, Zap24, Zol93]. Form [Kra73, Neu71, And22, Gla93, GKS24, Han17, Mar24, Rup76, Wis17]. Formal [SU15, Fra94b, Fre15, Kra73, Mah77, Man12a]. Formalism [O’N86]. formalisms [Gel11a]. formalist [Bau72]. formalization [Ric03]. formation [Díe97a, McG21, Wal16]. formative [Fin02]. Formats [Vor11]. Forms [Loi16, Car25, Cra84, Fra81, Mül12]. formulas [Gro05]. formulation [Gie91a]. Forum [Ano90d, Boo23, Die24, Hus24, Ken23, MM24a, Slo24, Wes23]. Forward [Pye08, McK16]. Foscarini [Cam12]. fossil [Tur22a]. Foucault [Ano91b, Pra77, Sci21]. Foundation [TC25, Mel04, Ris90, Ris91, VK19, Hau22, Lin22a, Rot22, Sol22, Tur22b]. foundational [Bai92]. foundationalism [AP06]. foundationalist [Pas05]. Foundations [Cha20a, Cha20b, EZ13, Cos98, DN00, Dor70, Gra71, Hei74, Jan73, Ker87, KH25, Sch21b, She11, Van09, vS14, GW73, Han21, Hei75, Heu14]. Founding [Ano89g, Ano01d]. Four [Nis93, Nor20a]. Fox [Bel15, BDR13]. Fraassen [Bro11b, Kuk95, Oka00]. Fracastoro [GT05]. fractals [Mur93]. frame [Tre07, VS12]. frame-theoretic [VS12]. frames [PE24]. framework [Fei14, Ger14, Gon96, Let21, PM22, Wal22a, Zap24]. frameworks [Bal20, Bla13, Coe22b, HKMV11]. Framing [WV22]. francais [Sig90, Spa91]. France [Bor15a, Lat88, Sto25, Zin11, Sch91]. Francesco [Ede09]. Francis [May71, Cun00, Cun01b, Hor73, KL21, Lux67, Osl71]. Francisco [CSD21]. francois [Des49, Gra96, Pan06, Tre07]. Franklin [Hac91]. fraud [Rad22]. Frauhofer [Jac94]. Frederick [Liv93]. Free [Dra75, Nob23]. freedom [Bue15, Wil10]. Freind [Mar88]. French [Cra93, Le 58, BV05, Cha19a, Cro92, Das86, Des49, Dou22, Fre25, Gla86, Gla02, Hut04, Pas62, PG20, Sim17]. fresh [Cel13]. Fresnel [Saa05, Wri17]. Freud [Mic24, Yea85]. Fricker [DC09]. Friedman [Fer12, Pic22, Sch12]. Friedrich [Gam21b, Gur19, Hin15]. friends [Des49, Out96]. fringe [Gra24]. Fritz [Pes13, Sie10, Sie10]. Frontinus [Cuo00]. fruitfulness [Yap11]. full [Dra87]. Fuller [Gie91b]. Function [Bou21, Mum06, Sch06, Ano72i, BDB21, Gau81, Gla93, Han06, Lan21, Nis77, Pre06, Ver06, Rup22, Cam22]. functional [Kro06, Vae11, Woo15]. Functionalising [Lor22]. Functionalism [Pic04, Pra75, Sha94]. functions [GVD24, Hei14, Hou06, Qua75, Sch09b, VH06]. Fundamental [Bat23, Dar23, FR23, Mil23, Pal23, Ray92, Gré19, McL24, Qui17, Rad93, Kno76]. fundamentality [Isa17]. Fundamentally [Ups22]. Funding [Lin22a, Rot22, Sol22, Tur22b, Hau13, Hau22, WHMO22]. Funkenstein [PR90]. Further [Ris91, Dra76, Kon24b]. Furthering [Fri24b]. fuss [Lar03]. futility [WS21]. future [Bro73, Bud13, Jar89, Sti13]. futures [JFS15, O’N08].

G. [Pre08b]. Gadamer [Aba11]. gain [WL22]. Gaining [Goo21]. galaxies [McG21]. galaxy [McG21]. Galen [Con00, Cos98, Mow85, Ose22, Tot12]. Galilean [Cah86, Fin11, McM85, Mit72a]. Galileo [Sha90a, Sha90b, Abl15, Ano91l, Ano01b, Ari84, Ari01, AP04, Bro76, Cam15, CN83, De 00, DD69b, Dra76, Dra78, Dra86, Dra87, Dra89, Feh82, Fey74, Fin89, Fin02, Goo80, Jan15, Jar76, Kle79, LM12, LoL15, Mac73, Mar13, May11, Mer80, Mer82, Pal03, Pal05, Pru89, Rap11b, Red87, Sch15, Seg89, Wal91, Wil01, Bia94, Sha90a]. Galison [PY06]. Galton [KT21, KL21]. gambler [Wil05]. game [Cun88, Cuo06, Gra96, GY11, Weh19a, dB09, CE19, Weh19b]. gap [DK18, Joh12a, Koc06, Tur22a]. gaps [Wan02]. Garching [Wit13]. Garcia [Gra24, dC12]. gardes [Bra08]. Gardner [McA05]. Gary [Ful93]. gas [Gay77]. Gassendi [LoL15, Rog95]. Gaston [Chi08]. Gauge [Wal24, Cat22]. Gauss [Yap11]. Gavin [Ege70b]. Geminiano [Van16]. Geminus [Dra89]. Gender [Hal12, Ric21]. gene [Fal86, KD19]. Genealogy [Nik19, Mét13]. Genera [Bal20]. General [Pit22, Wal22a, Bol15, Cur92, Eve18, Ker87, Kus01, Le 58, Ryc92, Stö15, vS20, Le 58, EG80, FK90]. Generality [Pen22a, BRdS96, Mat20]. generalization [Yli19]. Generalizations [Dar96]. generalized [vdBG90]. Generalizing [Bou21, Pap24, Mee92]. generate [Rat20]. generation [Fee16, Lin22b]. Generative [Fag16, Gil83, Gro05, Ter21, Weh19a]. generativism [Lug92]. Generativist [Bai92]. genes [HM88, Sim21]. genesis [Dar21b, Ros19, Tho09, dO12]. Genetic [Kit87, Koi21, Sti13, Waj21, WL22]. Genetics [Koi21, BRdS96, Cha13, CR13, Har93, Hop94, KT21, Meu19, Rad22, Ric21, Sch74, SLP23, SLH23, Tei22, Vor14, Wat04, dCR24]. Geneva [Spa91]. Genève [Spa91, Spa91]. Genevoise [Spa91, Sig90]. genitalia [JW22]. Genomatron [Dur13]. Genome [Hee21]. genomes [Sim21]. genomic [Sti13]. genomics [Rat20]. genopolitics [Waj21]. genre [Baz88, Gro90]. gentleman [Bar22c]. genuine [Sch14b]. Geochronology [Bor25]. Geodesy [Ker12, Ohn22]. Geoffrey [Cul04, Mye89, ES06]. Geographical [dC12, MM09]. geological [Buc09]. Geology [Wu14, Bus71, FG22, Lea98, Sch69]. geometric [DBM22]. Geometrical [Gie91a, Mal90, Fer01, dJ97]. Geometry [Wri73, Cla68, Das86, Dom13, Ede09, FR24, Gra05, Gra72, Gri91, Jes99, Joh12b, Lea91, Lor20, Mar13, O’G77, Ric86, Ric03, Ros19, Slo03]. geomorphology [Ben18, Car73, Her69a, Her69b]. Georg [Hoo13]. George’ [Goo04, Con21]. Georges [Ano91b, Chi15, McC01, Mét13]. Gerd [McR72, Ano01d, Ano03f, Blo82b, Jar03b]. German [Hop94, Liv93, Car21, Gre92a, Har93, Kau08, Sie10, Ueb25, dCR24, vdB22]. Germany [Ano95c, Ahn04, Len81, LW13b, ND22, Tei22]. germs [Stu91b]. Gersonides [Gla97]. Gerstenberg [Whi74]. Geschichte [Neu71]. gestures [Alg08, Sib95]. Gesuiti [Bal92, Gia94]. get [Kle79, Kuu11]. Getting [Cun88, Swa24]. Ghiselin [Ege71]. ghost [Dou21]. Gibbs [Wei10c]. Gibson [BZ81, Ree86]. Giere [Hei75, Bro09, Sie89]. Gilbert [Geo14]. Gingras [Col09, Col08]. Giordano [Ros19]. Girolamo [Eas72]. give [CS07, Le 58]. Giving [Hoc16]. Glanvill [Bro07]. glass [Sch07a, Acu25]. Glen [Rud74]. global [Ben18, Cam18b, CL18, Cha13, Dav21b, Leh18, Lyn95, Nor22, Wal22b]. globalized [Pin14]. globe [Del22]. Glory [Sto12]. glossed [Ber93]. go [Chi16, Dar21a, WOL20a]. goal [De 24, Gri10]. goal-derived [De 24]. goals [Bau98, DCJ^{+}21, Ell13]. God [Fra93, Hat79]. Goddiksen [RGD15]. Gödel [Pas11]. gods [McN22]. goe [Sha17]. Goethe [Hen98, PB96, Rib85]. Going [PP23]. gold [Col03]. Golden [Wit24, Van04a]. Goldenberg [Chu22, Joh22, Lea22, WV22, Gol22]. Goldman [Kus01]. gone [Gol75]. Good [Iva11, Bha17, Dar21a, Fai12, Iva10, Joh22, Sar12, Goo04]. Goodman [Boy14]. Goodyear [Hin12]. Gordon [Car73]. Governance [KM22]. governing [Coe22b]. GRADE [PM22]. GRADEd [MBG21]. grain [Bou20]. grammar [Gil83]. gramophone [Ste05]. Grant [ML18]. Graph [Bou16]. Graph-based [Bou16]. graphic [Dou95]. Graphical [dC93]. graphs [Kin21]. grasp [SF25]. Grasping [LJ11, WH14]. Grassmann [Ban13]. Graves [Jan73]. gravitation [Bro16b, Che23, Con21, Hoe94, Bam96]. Gravitational [FR24, Col04, EG80, Eld25, Gin07]. gravitatione [Hen11]. Gravity [Cal21, Col04, Hen11, Koc09, Koc11b, Sch11, Duc11, Due21, MSS^{+}22, Wal22c, Col08]. Great [HLM13, CW68, Eas72, Rud74, Wra19]. greatest [Ray90]. Greece [Hei16, LS02, Cul04]. Greek [Asp13, DFT12, Fek24, Kno76, RT95]. green [Kra24]. Gregory [Liv93]. Gresham [Bie20]. Griffin [Ris91]. Griffith [HL13]. Grosseteste [KH17]. Ground [Pri19, Zan88]. groundbreaking [EB11]. grounded [GP10]. grounding [Eng18]. grounds [Wil10]. Groundwork [Nas11]. group [CM22, Fra21c, KF23, Pra75, Row11a, Sar82, Sar97a, Sar97b, SG85]. groupoid [Cat22]. groupoid-theoretical [Cat22]. Grove [Kni71, Mor91]. Growing [BK21]. growth [Fin73, Fra79a, Gla89b, GKS24, KC12, LM70, Mic71a, Sch20b]. Grünbaum [O’G78]. GRW [Mar22]. guarantee [Spe04]. guesswork [Les78]. guide [Iva10, Ste04, Top00]. guidelines [DS21, MBG21]. guiding [Kra18]. Guido [DD69b]. guise [Wei21b]. Günther [Vea72]. Guo [Cul06].

H [ML18, Mel04, Sno04a, Whi74, Mau13]. H. [Hon87a, Sch11]. Hêgemonikon [Fek12]. habit [Roc88]. habits [Pin97]. habitus [Wil05]. Hacking [Kus10, Sci21]. Haeckelii [Rup76]. Hague [McE76]. Halevi [Lan20]. Half [Mat22a, Reu22, She93, dC02]. Halley [Far07]. Hamilton [Hen84, Koe95, KVJ22, O’N86]. Hamiltonian [Pit22]. Hampshire [Bus71]. handbooks [Dun12]. handle [Han05]. handmaiden [Van95]. hands [Tur22a]. Hankel [Toa24]. Hankins [Bri73]. Hans [Fra20, MW25, Rhe25, SD25]. Hans-Jörg [MW25, Rhe25, SD25]. happen [Cam15]. happened [Dur13, Kle15, Mil19, Pen22b]. happy [Wan12]. hard [HM06, Pin97]. Hardback [Bia94, Ere91, Mis94, Rhe94a, Cul04, Cul06, Gur19, Jas04, Mel04, Mor90a, Mor90b, Pes13, Sie04, Ste04, Van18, Gui04a]. Hariot [Ale98]. Harmondsworth [Sha90a]. harmonics [Cre12, Fek12]. Harmony [Mar09]. Harriet [CM00]. Harry [Sor16]. Harvard [Bau71, Liv93, Sch91]. Harvey [Mow85]. Harwood [Hop94]. Hatfield [Ful93]. Haven [Cul04]. Health [Myk22, Pen22c, Ups22, Val22, CL22, Cre18, Fra21b, Kuk22, VABF24, Wil24b]. Healthism [Kuk22]. Heaps [Wol20b]. Heart [Mia01]. heat [Psi94, Sib95, VS12]. heavens [Has97]. Heaviside [MB22]. hedged [Kow11]. Hegel [Bei03, Suc90]. Heidegger [Cah86, Kho03]. Heideggerian [Cer19]. Heidelberg [Neu71]. Heike [GG88]. Heinrich [Nor00]. heir [tH93]. Heisenberg [Kav24]. Hele [Blo08]. Helene [Ano84g, Aba08a, Aba08b, Chi01]. helikôn [Bar09]. hell [Far07]. Helmholtz [Bok15, Cah97, Dar03, De 16, Ful90, Hat90, Hei74, Mic93, Obe15, Pat09, Tri14, Ful93]. Helmholtzian [Acu25, Dar07]. Hempel [Boy14, Hu21]. Hendry [Sce06]. Henri [Gie91a, Gow97, Stu89]. Henrich [Ham00]. Henry [Bel15, BDR13, Can71, CB92, Dor74, Tho15]. her [Saa05]. Herbert [Cab96, Smi83]. Hercules [Rem07]. Herder [HN16]. here [Kla00, Kuu11, Swa24]. hereditary [Ano91k, Sti13]. Heredity [Rhe08, LB94]. heretic [Red87, Sha90a]. heritability [Bou22a, LB22, MT22]. Heritable [LAL21]. heritage [BS13, Far97, Gra13, JW13, Jar13, LW13a, LW13b, Maa13]. Hermann [Pes13, Sie10, Dar03, Ful90, Gio16, Pes13, Sie10, Ste24a]. Hermeneutic [Jar03b, Gin88, MH84]. Hermeneutical [Aba11]. hermeneutics [Hee98]. Hero [Tyb03]. herping [Qui21]. herrings [Ste22b]. Herschel [Che03, Goo87, Nas11, Rus75, Rus78, Yeo89]. Hertz [Bar80, Eis21, Hon87a, Nor00, Pre07, Pre08a, Wil89]. Hesitancy [Chu22, Gol22, Joh22, Lea22, WV22]. Hesse [Blo75, Hal12, Hea24]. Heterodox [Brä24b]. Heuristic [Cam18a, Ans02, Cha86, Qui72]. heuristics [Bor15b, Pos71]. Hevelius [Ver07b]. Hidden [Far07, Mit22, Moy77]. hieroglyphica [Cam12]. High [Dil90, KSZ24]. high-energy [KSZ24]. Higher [Fu95]. highlanders [Bra98]. Hilbert [Mah77, Mit22, Stö15]. Hilbert-style [Mit22]. Hildesheim [Whi74]. Hintikka [Bro71, dJ97]. Hipparchikos [Sto12]. Hipparchus [Dra89]. Histology [CCB24]. Historian [Cab96, Cok15]. historians [Bou22b, Hac92, SD25]. Historical [BB23, Mun16, O’N08, Aya25, Bec71, Bor25, Car05, CL04, Cle83, Cul06, DvdB22, Dhe23, Díe97a, Díe97b, Dio18, Fra21a, Fre80a, Gam21a, Kos93, Kra81, Kus10, Lau23, Loi16, Mét13, Mit72a, Obe05, Pap24, PG20, Pit22, RDM25, Sch85, SLP23, Sti13, Swa24, Tur04, vdLBV22]. historical-critical [Pit22]. historical-philosophical [vdLBV22]. Historicism [Bei08, How90, Kuu16]. historicist [Axt93]. Historicity [ET20]. Historicizing [Gou23, Mau13]. Historico [Zam18]. Histories [O’M16, Kid17, Ruy23, OS14]. Historiographical [Cha20a, Cha20b, Aba10, AM20, Top00]. Historiography [Gol23, HK16, Buc21, Gio24b, Joh98, Kle08, Koe95, Ort08, Roc02, Rot13, Sch16, SLH23, Spe17, Tuc93, Wal75, dO12, Tos06]. historistic [Hüb80]. History [AR22, And22, Ano80c, Ano84c, Ano89a, Ano91m, Ano95c, Ano97g, Ano01a, Ano01b, Blu21a, Bol15, Buc88, Buc89, Car22, Cha20a, Cha20b, Cob11, Col16, Dre20, EG80, Gam21b, Gua13, Kal75, Kam94, LSH24, Maa13, Man80b, Mar21, May00, MW25, Nik19, Pas20, Sno04a, Stu01, Wer18, Aba08a, Aba08b, Aba11, Ait70, Alg08, Ano81c, Ano90b, Ano90d, Ano91c, Ano91d, Ano92c, Ano92d, Ano93d, Ano93e, Ano94c, Ano96a, Ano97a, Ano97b, Ano02b, Ano03a, Ano03f, Ano09f, AH15, Ara23, Bar22b, Bas03, BCC87, Ben87, Ben18, BM20, BM21, Bus71, Can98, Car03a, Car73, Car95, Cat12, Cha11, Cha44, Cha69, Chi01, Cla95, Coo20, Cor93, Cre18, Cro90, Cro91, Cul06, Den00, Edi22, Ege70b, Elw12, Fin89, Fis18, Fra93, Gam21a]. history [Gar00, Gio16, Gla89a, Gla89b, Goo04, Gow73, GT05, Hac89, Ham23, Han05, Her69a, Her69b, Hut04, Jon13, Kim95, Kin15a, Kle04, Kni71, Kra24, LL16, Llo97, Loi16, Los83, Low93, Mal01b, Mar88, May04, McN22, McO96, Mor00, Mos07, Mur93, Mye89, New10, Nun84, Pal05, Pal70, Pen15, Pra77, Rad97, Rad22, Ram08, Rei96, Ric09b, Rie14, Rit21, Rob22, RH11a, RDM25, Ryc15, Sch09a, Sch69, SW12, Sha97, Spe17, Sum13, Ter17, Tos03, Tos06, Tos07, Tri05, Tru68, Tyb22, Van18, Wal16, Wei10a, Wit13, Wra05, vdB24, SLH23, Sha90b]. history-and-philosophy-of-science [LL16]. Hobbes [Ada16, Ano91i, Ber80, Dun10, Mal01a, Sch88a]. Hobbesian [Jes99]. hobby [Muk22]. hobbyists [Muk22]. hoc [Lep75, Ray92, Sch18a, Sch24, Vot14]. hoc-ness [Sch24]. Hoedown [Mir05]. Hoefer [Ste22b]. Høffding [Fav91, Fay88]. Hölder [Mic93]. Holding [Lac15]. hole [Wei21a, Wei21b]. holes [Fur22, Per08, Tos07]. holism [Kri82, Per08, RS21a, Ver17]. Holistic [CK22]. Holography [Gre22]. Holy [Ram12a]. Homberg [Kim00]. homeostatic [Lip20]. Homing [Nee13]. Homo [DD13]. homology [Cox12, Gou23]. Honoré [Dra75]. Hooke [Gal96]. Hooker [Sie98]. Hopkins [Ano73a]. horizons [Kos88]. Horizontal [Dea12]. Horse [Bau15]. Horton [Har74]. Hosiasson [Tor24]. hours [Rat19]. Houses [Lea22]. how-possibly [VJ19]. Hoyningen [San09a]. Hoyningen-Huene [San09a]. HPS [Kuu16, Lau89]. Hubble [MM24b]. Hubris [Isa17]. Huene [San09a]. Hugo [Dar76]. Huguenot [Bai04]. Hull [Ere91, Rus89]. Human [Car25, Cre18, Fag16, Ano90d, Ano92e, BMBM20, Coh08a, DK18, Dri24, Hea24, Hol21a, Lou08, Mak08, MC25, Nob23, O’N08, Osl72, Rog72, Spe17, Stu12, Wal16, Woo71, Woo94, Yol70, Zam08]. Humanities [Bro71, Ege70b, Vea72, MSS^{+}22]. Humanity [Det22]. humans [McN22]. Humboldt [Liu22, Wis18]. Hume [Aro71, How13, How15, Pao11, Sch08a, Spa90, Wel02]. Humean [Suc90]. humility [Bre04a, Con15, Isa17, Kid11b, Kid16b]. hundred [Díe97a, Díe97b]. Husserl [Cah86, Koc11a, Mah77]. Hutchinson [Mun74]. Hutchinsonianism [Ast04]. Hutton [Ros18]. Huxleyite [Cra22]. Hyde [Jac16, Uss15]. Hydrogen [Nee13]. hydrostatic [GS02]. Hydrostatics [Cha15a, Sha74]. hyphenated [LL16]. Hypotheses [Coo23, Sch85, Ach87, Bro16a, DvdB22, Fay24a, Sch18a, Tre87, WHMO22]. hypothesi [Nye76]. hypothesis [Fre80b, Hau13, Lep75, Mor21, Par95, Sch07c, SW12, Spr10c, WHMO22, Wes72]. hypothesis-driven [WHMO22].

ICD [vdLBV22]. ICD-11 [vdLBV22]. Iconology [Mit86, Rhe94a]. idea [For89, Jac03, KC94, Kra18, Mat80, Rei97, Sha22, Whi71]. Ideal [Bro21, McH21, Bal20, Bie16, Feh82, Kit75, Sto12, Wie21]. Idealisation [Waj21]. Idealisations [Ser23]. idealism [Rab75]. Idealist [Jac20]. Idealization [Elg22, JSZ23, Ode22, Pin22, Ric22, Ter21, CK22, Jhu22, Lay78, McM85, Spa23, Wat21]. Idealizations [Rod21, Str21, Tan21]. idealized [Spa23]. Ideals [Blo07, Fed23]. Ideas [Sno04a, Boa07, Clu12b, Goo87, Hen11, Kle08, Ree86, Rus75]. identifications [Kut22]. identify [Wat23]. identifying [San22]. identities [Rie14]. identity [But93, Chi06, Cun88, vD21]. ideology [BC00, Chi15, Chu22, Gla83, May04, Mit86, Mor93, Rhe94a]. If [vdLBV22]. ignorance [Mee21]. II [CEW16, Cor11, Cro91, Díe97b, Gla89b, Gla01b, Isa20, May04, McL06b, Moy91b, Rhe92b, Row13, Ryn95b, Sar97b, Tos07, Wal22b, War93, Cha20b]. III [GT05]. illegitimate [KR22, Reu22]. illness [Tuc95]. illuminating [Cha16]. illusions [Kit93, Mal90, Vas00, Ful94]. illustrated [Lor20, Swo74]. im [Pes13, Sie10]. Image [Rhe94a, Eag07, Ger20, Mit86]. Images [Fri24a, Mal90, Mal01a, Mos07]. imaginaries [MHW22]. imaginary [Vor11]. Imagination [Bro21, McH21, Fun86, Hea24, Lor20, PR90, Wie21, Hav21]. Imitation [Gei23, Weh19a, CE19, Weh19b]. Immanent [Hen16, Car12]. immortality [Fal98]. immunology [Kam94, TC91]. iMongers [Mul14]. impact [Pan96]. impacts [Way21]. impasse [Zha22]. imperative [Ric21]. imperial [Cha13]. imperialism [BT19]. imperialist [Ale95]. impetus [Dra76]. Implications [DS19, Wat21, Wie21, Buc21, Coh08b, Fal98, Gal21, GG88, Let22, Mak03, Pot12, Qui74, Sha97, Sto10]. imply [Wil01]. impone [Bal92, Gia94]. importance [Cor01, Mer92, Ueb94]. important [Mee92]. impossibility [Ama05]. impossible [She91]. Improving [Det22, LGC21, Zam08]. Imputing [Yea85]. Imre [Fra79a]. In- [Due21]. inauthenticity [Fin11]. incentive [Hee19]. inclinations [Gow87]. including [DD69a, DD72]. inclusion [Ric21]. incoherence [Ili04]. incoherent [Ohn21]. Incommensurabilities [Dem08]. Incommensurability [Bro05, D’A00, Mal93, Bia90, Dav13, DS19, Fu95, HH90, HHOA96, Jac02, Obe05, San09b, Wan02]. Incommensurable [Rab75, Nor86]. incomplete [Fis23]. incongruous [Ver82]. Inconsistency [Mor11, Smi88, Bro90a, Dio13]. Inconsistent [Tan21, Cha10b, Díe11, Her15]. increase [Car15]. increasing [Nii14a, Row15]. Incredulity [NI03]. independence [Hud94, Wil09b]. Independent [Eld24, Gay78, Tro95]. indeterminacy [Ano91j, Cal22, Men24, Sch21b]. indeterminism [McL24]. Index [Ano71h, Ano72g, Ano73e, Ano74f, Ano75e, Ano86g, Ano98q, Ano99z, Ano00f, Ano02a, Ano05a, Ano06a, Ano80g, Ano81d, Ano82d, Ano83d, Ano84e, Ano85g, Ano87e, Ano88d, Ano89h, Ano90g, Ano91g, Ano92b, Ano95b, Ano01c, Ano04a, Gab73]. India [dC12]. Indiana [Hei75]. indicator [WL22]. indices [Ait70]. Indifference [McM93]. indifferent [Nye76]. Indigenous [Lud16, Tur22a, EHPL22]. indirect [Ell13]. indiscernibilities [Cat22]. indispensability [Bak21]. indistinguishability [CF21b, FL22]. Individual [Tra22, Can98, Kid11a, Mar04, War23]. individualism [Cur89]. Individuality [Fri14, Tra22]. individuation [Spr10b]. Induction [Ear20, Bak20, Cor21, Dav21a, LK20, McC20, Pap10, Pos71, Psi94, Red02, Rei20, ST20, She06, Stu20, Suc90, dG20]. Inductive [Fis18, Bou16, Hor73, Pas11, Szn16, tH02]. inductivism [Díe11]. inductivist [Gay78]. industry [Ara13, HL13, Pri22, Rib07, Van95]. inefficacy [Ric97]. Inertia [FR24, And94, Fre71, Gab71, Loh21]. inevitabilist [Sol08b]. Inevitability [Kid16b, Ayl19, San08, TC21]. inevitable [Kin15b, Sol08a]. infallibility [Kri80]. infer [Bog20]. Inference [CDC24, Dav21a, Lip91, Spr10c, Cha10a, Gel10, Hal05, Hen22, Hes74, Hes75, Miy23, Oka00, Pad15, Sch13c, Suá15, Wal12, Blo75, Ach92]. inferences [Gen20, PP21, Tuc20, Vez16]. inferentialism [Tan21]. infertility [Bru20]. infinite [Cor11, Nor20b, Par20b]. infinitely [Noa08]. Infinitesimal [Gio16]. Infinitesimal-Methode [Gio16]. infinity [BT09]. inflation [Ant25, Wol24]. inflection [Gal96]. influence [Abr11, Bla72, BT09, Det22, Fla83, FP99, Ful90, Kid11a, Rus75, Van86, Sno04a, Swo74]. influences [Kai92]. information [Ano05f, Ano06c, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13a, Ano13b, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17e, Ano17a, Ano17b, Ano17c, Ano17d, CT14, LL19, Lus21, PS10, Tho09, Dun22]. Information-Theoretic [Dun22]. Infra [CCB24]. Infra-molecularizing [CCB24]. Infrared [CL05b, CL05a]. infrastructure [Lyn95]. ingenious [De 18]. inherent [Dre20]. inheritance [Fab19, Sar86]. Initial [Ano01b, Ari01]. Innovation [Fis23, Lew02, Slo24, Bre04b, Gla99]. innovations [HD19]. inquiring [Jas04]. inquiry [BR19, DK22, FJ24, HK22, WC09, Wil22a]. inquisitive [Fle22]. inscrutable [Zam03]. insect [WHC23]. insects [Dhe23]. insight [Bec23, Mit72a]. Insights [MHW22, WC09, Zin11]. inspired [Rap11b]. installation [WHMO22]. instead [lC00, Rup76]. Institute [Pie97, Clu12b]. institution [BD02]. Institutional [Cur89, Hag01, Hoc87, Sum13, CL04]. institutions [PP21, Hut04]. instruction [Sto12]. instructions [Wil09b]. Instrument [Row23, TA23, Ara23, Hen23, Miy23, Pol21, Sch09a]. instrumental [Gie89, Ras96]. instrumentalism [Ara23, Dio13, Miy23, Pas11]. instrumentalist [Bau72]. Instrumentalists [Het96]. Instruments [Mer04, Ram92, Sib95, Sla02, Ben87, Bor20, DK09, Hac89, Ros19, San16b, Tau09, Tre07, Ver07a, dC93]. integrate [New10]. Integrated [AH15, BB22a, Isa14]. Integrating [Käs18, MSS^{+}22]. integration [Dre20, Lud16, Zap24]. Integrative [Jha97, Lac11, POW16]. integrity [RE19]. Intellectual [Ant25, CR13, De 96a, Jha97, Way21, AM20, Coa02, Hoc87, Kid11b, Rol16]. intelligence [Kuk01, Uss15]. intelligent [GKS24]. intelligenz [Bra95, Cah97]. intelligibility [Dom09b, Fey16a, Kub16, bAM11]. intensional [Par16]. Intensions [dL09]. intensities [Cla68, Gra72]. intensive [CL22, WL22]. intentional [VH06]. intentionality [Yea85]. Inter [Bus71, Ger14]. Inter-Disciplinary [Bus71]. inter-level [Ger14]. interaction [Mac73, Sch03a]. Interactional [CE15, PK14, Sch07a, SDC07, CERH06, CEG07, God14, Jen07, Pla20, RGD15, Rib07]. interactivist [Hoo13]. interdisciplinarity [And16, MN18]. Interdisciplinary [GYM14, Rie14, GY11, POW16]. interest [Yea82]. Interests [Sch09a, Wil22a, Yea82]. interference [Bha24]. interferon [Pie97]. Intermediate [Cha12]. International [Ano80c, Ano87d, Ano89a, Ano91m, Sno04a, War89, Zam18, Ker07]. Internationales [KMS73, Whi74]. interplay [Ara92]. interpretability [Räz24]. Interpretation [But93, Dun22, Aba08a, Aba08b, Bel90, Cah86, Gie92, How98, Jha97, Kno76, LB22, MCGP17, Pao11, Rei18, Sla22, Wan02, dJ97, vdBG90]. interpreters [Chi11]. Interpreting [Bel74, Hun21, WW04, Bak23, Rib07, Wil09b]. interrelating [Gök90]. Intersubjective [Row08b]. Intersubjectivity [Nor12]. Intertheoretic [Krü80]. intertheoretical [Wag24]. Intervening [Fri21, LC24a]. Interventionist [Fri24b, PSS18, Sta18]. interventions [LGC21, Maz22]. Interweaving [Win12]. intrinsic [Bor15b, Sol08b]. Intro [Bri73]. Introduction [Abl15, AH15, BB23, BK16, CL18, CT16, DK18, DHS14, DFT12, GN16, GYM14, Hau16, HK16, KC12, Lad19, Lau87, Man18, ME15, Nel94, PZ14, PP21, SU15, SP21, SW07, TJT16, VFS14, Ano91b, Díe97a, Díe97b, Ege70b, Gra72, Mah77, Man04, MTC17, SLH23, Sha90b, MW17, Zam18]. Introspection [Tam05]. intuition [Kra72, Kra73, Mac10, Sha03]. Intuitionism [Gil83, Pou00, BO24]. invariance [Díe97a, Pos71, Sla22, Wal24]. invariant [Gro20]. invented [Mal13]. Inventing [Sch12]. invention [Cun88, Mar15a]. inventions [Fie86]. inverse [Dor74, Fal17, Nii11]. Investigation [Bra01, PS24, Pit22]. investigations [Wan22]. invisible [Cat14, Dav20, FWW09, Goc08]. Iranian [Del22]. ironical [Ryc92]. irrational [Bia88]. irrationality [SW12]. irreconcilable [Hoc16]. irreducibility [Bue15]. irreversibility [Col90, Val21, tV21]. Isaac [Elk74, Hay04, Pal71, Ste04, BC01, Dem06, FM04, Gui04b, Sno04b, Gui04a]. ISBN [Ach92, Ano91b, Bia94, Cul04, Cul06, Ere91, Ful94, Ful93, Gie91b, Gui04a, Gur19, Heu14, Hop94, Jas04, Kam94, Liv93, ML18, Mel04, Mis94, Mul94, Pes13, Rhe94a, Rhe94b, Sch91, Sie04, Sno04a, Ste04, Van18, Zol93]. ISBN-13 [Gur19, Van18]. Islam [Nas68, Nas83, Bau71]. Island [Mos01, Blu21b]. islanders [Bra98]. isn’t [Fri03b, Räz24]. ISO [Sha94]. Isolated [Wal22a, Wal22b]. ISSN [Spa91, Zam18]. issue [Ano89b, BB23, CL18, MW17, Rit12b, Sol08b]. issues [Bro73, Col16, Cor93, MCGP17, Osb89, Rad93, Rot96]. Italia [Bal92, Gia94]. Italian [Bal92, Gat16]. Italy [Bal92, Cam12, De 96b, DD69b, Van16, Ege70a]. iteration [Ell12, Rie24]. itself [Aub08, Bor20]. Iudaicall [Clu12a]. iv [Gal71, Due22, Sim70]. Ivanova [Kid11b]. ix [Ere91, Mor90b].

J [BZ81, Bro71, Bro21, Bus71, Coo96, Cra22, Ege70b, Gal71, Hac89, Hav21, Kni71, McH21, Par95, Ree86, Rhe94a, Ros22, Rui23, Wie21]. J. [BZ81, Blo84, Coo96, Dou95, Rui23, Sut71]. James [Mis94, Sno04a, Bid11, Car89, Cro11, For75, Hud89, Oli92, Ree86, Ros18, Zap24, Mor90b]. Jane [Die24, Hus24]. Janiak [Hen14a]. Janina [Tor24]. Janus [Sak22]. Jean [Bri73, Big08, Gra96, Kin87]. Jean-Francois [Gra96]. Jeans [Gor91, Hud89]. Jed [Gui04a]. Jeffrey [Fre15]. Jesuit [Dea87, Pan96, Rap11b]. Jesuits [Bia94, Bal92, Wal91]. Jevons [Buo24, Maa99]. Jewish [Rat19]. Joelle [Rob22, Edi22]. Johann [FWW09]. Johannsen [Van86]. John [Ege70b, Jan73, Mye89, Osl72, Sch91, Swo74, Van18, AD12, Alm12, Bar22c, Blu21b, Bro15, Cam12, Che03, Clu12a, Clu12b, DB96, Feo12, Goo87, Hal05, Hol83, How04, Hun03, Joh12b, Kil99, Man12a, Mar88, Mor21, Mos01, Nas11, Par12, Pum12, Ram12a, Ram12b, Rus75, Wal06, Zwi11]. Johns [Ano73a]. Johnson [Bri73, May71]. Joint [Yos23, Ano09f]. Joints [McO01]. Jona [dOFGB22]. Jonathan [Hop94]. Jonesium [LSH24]. Jörg [MW25, Rhe25, SD25]. Josef [Nir23]. Joseph [Bas92, Bro07, Jac94]. Joule [Mor90b, Car89, For75]. jour [Buc21]. Journal [Wra19]. Journals [Ano09f]. journeys [Bud13]. Judah [Lan20]. Judea [Rat19]. Judgement [MN20]. judgements [JC20]. judging [Hoe80]. Judgment [Lev01, MJ23, Pre08b]. judgments [Kno09]. Jules [Hof13]. July [Ano04b]. junction [Vor14]. June [Ano96a, Ano22i, Ano23g, Ano24j]. just [Cus85, MM24b, Toa24]. Justice [Hag01]. Justification [Gau98, Bai90, Bai92, BP20, DC09, HH87, McK95, Sch74, Sch02a, Wat98]. justified [MJ23].

Kaibab [Mil19]. Kaila [Sla22]. Kamerlingh [GG88]. Kant [Ful93, Stu12, AA67, Buc69a, Buc69b, Car03b, Coh08b, Coh18, Coo20, Coo23, DK18, De 18, De 16, DvdB22, DiS13, Dom13, Eng18, Fal18, Fal98, Fri08, GN22, GN16, Gau16, Gon25, Gri91, Guy08, Hat90, Hei74, HN16, Hen14b, Hut21, Kle08, Kra11, Kra18, Kra72, Kra73, Lea91, LA22, Mac10, Mak03, Mak08, Mar24, Mas11, MD13, Mas13, McN15, McN22, Nas16, Ols16, Pal87a, Pip79, Pri18, Ris90, Sch13b, Sch08c, Sha98, Sha03, Sha89, Smi13b, Smi13a, Sta13, Ste03, Stu08, SM24, Wal16, Wat98, Wat03, Wat13, Win81, Zam03, Zam08, Zam16, Zha22, dJ97, vdB18, vdB24, McR72, Gal71]. Kantian [Gam21a, Bog21, Bre04a, Coh08a, Eve18, Fra20, Fri18, Ful90, Gio24b, IG98, Jar03b, Kai92, Lou08, Spa23]. Kantianism [Gio22, HHO09, Bei08].

Kantians [Pul01]. Karine [Cul06]. Karl [Gam21b, JM11]. Kate [Die24, Hus24]. Kegan [Nic72]. Keith [Slo24, Wes23, Gre24a]. Kelvin [SW89, Mor90a]. Kemeny [Wag24]. Kemp [Feh07]. Kempelen [Gei23]. Kendig [Mas24]. Kepler [KMS73, Whi74, Bai87, Bai90, Buc72, Car16, Cro91, Esc08, Fie86, FH85, Jar79, Kle83, Mal90, Mar99, Miy15, Wes72]. Kepler-Symposium [KMS73, Whi74]. Keplerian [Mal90, Mit72b]. kept [Pas05]. key [Cha12]. Keynes [Run94]. keys [CS07]. Khalidi [Pol15]. Kidd [Iva11]. Kielmeyer [Gam21b]. Kierkegaard [Kid11a]. kind [CM22, Che10, Lip20, Rat20, SLH23, Vic21]. Kinds [Buc92, McO01, Car93, Con06a, Lem21, Pol15]. kinematics [Mou81]. kinetic [Dor70]. kinetics [Ram92]. king [Wit24, Fra93]. Kinmaking [Wil22b]. Kircher [Ver07a]. Kitcher [Ful94, Jas04, McL05, Sha97, Sol95]. Klee [Goc08]. Klein [Gla93]. Kluwer [Ere91, Sie04, Sno04a, Zol93]. know [Gin07, Jas04]. Knowing [Cam15, McH21, Mor05, Haw10, Jha97, Kle04, MW17]. Knowledge [Anz19, Ara13, BS13, CP18, Del16, DS19, HL19, Hou06, Hum19, Baz88, Bir08, Blo82a, Blo07, Boa11, BK21, CT16, De 96a, De 96b, Dea03, DK09, EHPL22, Fee16, Fis85, Fri98, Gal96, Gel80, Gla89b, Gro90, Had04, Hay04, Jac94, KG13, Kim12, KC12, Kle14, Kus01, Kuu11, LM70, Lee19, Lud16, Mak03, Mar04, Meu19, Mic71a, Mir04, Mor05, Muk22, Mül12, NI03, Ram12b, RGD15, Ric03, Ric05, Rog72, Rot96, Smi84, Ste24a, Stu91a, Top04, Tur11, Ueb05b, Vet18, Wei07, Woo71, Woo94, Yea82, Zuc19, Tos06, Tos07, Fin73]. known [Cla68, Gra72]. Koch [vdBG90]. Kochan [Fee06]. Kochiras [Sch11]. Kojève [Gol75]. Konrad [Kal75]. Konstruktivismus [Pes13, Sie10]. Koyré [Ant03, Stu01]. Krafft [Whi74]. Kraft [Hei74]. Krauss [Ste22b]. Kristine [Tei22]. Kritêrion [Fek12]. Kuhn [Bir04, Kuu07, Axt93, Ben24, Bîg13, Bir02, Byc12, CL04, Dav13, Dem08, Díe07, Dom09a, Fer12, Haw96, HH90, HH06, HHL11, IG98, Kva14, Lar03, Mal93, Mas15, Möß11a, Möß11b, Obe16, Pre04, Row11a, Row13, Sci21, SS09, Sim17, Wan02, Wra07, dL09, dO12, HH95]. Kuhnian [AL16, Can98, Cor93, Pou00, Rad97, Sam17, Sch17, Wal12]. Kulturwissenschaften [Kau08]. Kunstkammer [DK09]. Kurd [Gio24b].

L [Car73, Ege70b, Gab73, Mor90b]. Labor [Kim95, Bor20]. laboratories [And13]. laboratory [Dav21b, KM22, LC21, Par95, Rhe92b, SFT23, Tri14]. lack [Cha93]. laden [Sch13a, War21]. ladenness [Bas03, Ell21, Wec86]. Lagrange [Fra90]. Lakatos [Fin73, Cor97, Fra79a, Gay76, Mot02, MH84]. Lamarckian [Par95]. land [Cuo00, Kuu09]. land-surveying [Cuo00]. landing [Pri19]. Landmark [Nas11]. landscape [Har96]. Lane [Sha90a]. Langton [Bre04a]. Language [GP10, BCC87, Coh97, Dar02, Fra21a, Ghi94, Kap16, Man80a, McO96, Mye89, Sch73, Hac12]. languages [vdB22]. Laplace [Koz15]. Laplacian [vS14]. large [Pin09]. large-scale [Pin09]. Lasinio [dOFGB22]. Lasswitz [Gio24b]. Late [Mar16, Car03a, Gio22, Hel16, Mor93, Péc11, Rup76, Sam77, Szn16, Vos07, Olm12]. late-Victorian [Mor93]. later [Ber85, Hac12, Kus16, Lau89, Mat22a, Pre98]. Latin [Bal92, Des44, Des49, FM04, GR23, Har19a, New87, New09]. Latour [Sch91, Blo99, Coh97, Kho03, Seg00]. Laudan [Los78, MH84, Sar86, Sch87, Sie90]. Launching [CD15]. Laurent [Bro73]. Lavoisier [Bas92, Lew11]. Law [Ano87d, KH20, Leh20, Sch91, BMBM20, Bur02, Car16, Dor74, Fal17, Fra05, Fre71, Gar76, Haz23, KH17, Lan20, Luc20, PMG21, Sta13, Van09, Wor05, Aga77, Bal92, Bam96]. law-first [Haz23]. Lawrence [Mul94]. Laws [Bel20, KH17, Leh06, Bai87, Bal20, Chr94, Eng18, Fek24, Gro20, Hen14b, Hut21, Jan13b, Kow11, Lan06, Le 58, McN15, Pfe12, Ray92, Rei18, Wat98, Wat12, Wri17, Zyn96]. Lay [Bal92]. Layered [Elw12]. Lead [Col03]. leadership [Sto12]. learn [FJ21, Mac10]. learned [Mee92]. learning [BG23, Har96, Rat20]. Leavitt [Ris91]. Leeds [Jon13]. Leeuwenhoek [Dav20]. Left [Ric09b, Ben24, Ric09a, Ric09b, Ueb05a, Ueb10]. legacies [AM20]. legacy [Ano86h, Aro71, Bir04, Fre15, GN22, LAL21, Isa20, McC01, Mou20, Obe16, SU15, Tur22b, Zap24]. legal [Col21, De 02, Sar89, Sha02]. legalize [Gio14]. Legem [Bal92, Gia94]. Legend [Sol95, Fra93, Ful94, Kit93]. legendary [Wan12]. legislation [Löw97]. legitimacy [Lus21]. legitimate [KR22]. legitimisation [Rem07]. legs [MT22]. Leibniz [Ant03, Ban13, Ber80, Bro16b, Cov97, Cox75, Dun10, EC25, Fre02, Fri14, Fri24a, Fut02, Kim24, Lev11, Mel91, Nob23, Qui18, Sta13, Wat12, Wil82]. Leibnizian [Win82]. Leicester [Ege70b]. Leiden [GG88]. leisure [Muk22]. Lennard [LSH24]. Lennard-Jonesium [LSH24]. Lenses [Stu01, Ruy23]. Leon [Kam94]. lequel [Le 58]. less [BK17, Bus08, Fra05, Fre19, Lau90]. lesson [Acu25, Mee92]. Lessons [Alm21, Ano81c, Ell21, JW22, Jon13, Qui21, RH11a, Sto23, Aba08a, Aba08b, D’A00, Rui23]. let [Koz15]. letter [Fin11]. letters [DD69a, DD72, GT05, HH95, HH06, May71, May11]. lettres [Pas10]. Leuschner [Eig19]. level [Bor25, Ger14, Har19b, HS23b, Row11a]. level-switching [Bor25]. Levels [Man18, Mil23]. lever [Goe72, Pal08a, Van09]. Leveraging [LF25, Elg22, Jhu22, Ode22, Pin22, Ric22]. Lewis [AP06, Fra20, Har08, Pfe12, Ric03]. Lewontin [DHL21]. Lexington [Heu14]. Leys [Nik19]. liberalism [O’N03]. Library [Ege70b, Gal71, Mun74, Rob70]. licensing [Bid14, Boe22]. Lie [Man22]. Life [Nor00, Por01, Slo19, Aub08, Bla72, Cra84, DD69a, DD72, Fra81, Gau16, Kon24a, May71, Nas16, Ols16, RH95, Rup76, Swo74, Uss15, Wel05, vdB18]. Light [Ach87, Ach90, Sha74, Bak16, Buc92, Che90, Mot02, Nee02, PY06, Rap11a, Rei18, Smi15, Wor82]. Lightman [Van18]. Lightning [Hac91]. Like [Nik19, MN18, Ose22]. Lilo [Osl71]. Lime [Mah99]. limit [SF21]. Limitations [Los77, Pin14, Wil10]. limited [Van91]. limits [Kau08, Lor20, MN18]. lines [Díe97a]. lingering [Ryc92]. Linguistic [Kap16, IG98, Ric92, Wei10b, Wit13]. linguistics [Ter21]. link [Zem06]. linkage [DHL21]. linked [Lac21]. links [Dac15]. Lipton [Ach92, FSJ08, Cha10a]. List [Ano85g, Ano88d, Ano89h, Ano91g, Ano95b, Ano87e, Ano90g, Rob70]. listening [Tri14]. listing [Ano03f]. literal [BCC87, Mye89]. literary [CL04, Kuu17, Olm12]. literature [Föl12, Ort08, Ree12]. little [Noa08]. live [Rot96]. Lives [Rad13, Bec23, Bol15, CL04, SL21]. Living [dR13, BB22a, Har96]. Lloyd [Cul04, Che90]. load [RM16]. Local [Nor22, Wal22b, Dre20, Lau23]. localism [Ayl19, Kuu11]. locality [Fag24, How85, Lau23, Mat22a, SHB94]. Locating [Dav21b, Len21, Bra15]. Locke [Ans16, BC00, Hal05, Jac03, Mat80, Whi71, Woo75, Woo71, Woo94, Yol70, Osl72, Rog72]. Lockean [Con15]. Logic [DB96, Sch74, Gla01a, Kil99, Lay78, Man92, Man80a, Nic72, Par20b, Prz69, She91, Ste05, Szn16, Ueb15, Woo11]. Logical [Bro11b, Gel80, Ueb25, Wil77, Ano72i, Hei80, Mar24, PE24, Ryc92, SU15, Tam05, Ueb05a]. logico [Vea72]. logico-philosophical [Vea72]. logicophilological [Pat69]. Logos [Sto12]. loix [Le 58]. London [Ach92, Ano73a, Bia94, Blo75, Bri73, Bus71, Cul04, Ege70b, Gab73, Gra72, Gui04a, Hei75, Hop94, Liv93, Mel04, Mis94, Mun74, Nic72, Rhe94a, Sch91, Ste04, Fre25, dC02]. long [Hay04, Len21, Sta13]. long-lost [Hay04]. Longino [Lar24, SR05]. longue [Edi22, Mel21]. look [Kle83, Kra22b, MN18, Mur89, Tro94, Tur16]. Looking [CF06, McK16, Sch07a, Acu25]. Loop [Cal21]. looseness [Zam03]. Lord [AD12, Mor90a, SW89, Sny07]. Lore [Mor01]. Lore-Abiding [Mor01]. Lorenz [Kal75]. losses [Man22]. Lost [EG78, Gre92a, Hay04, Liv93]. Lot [Del17, Fre19]. Lotteries [Nor22, Avi19]. lottery [Nor20b, Par20b]. Lotze [Mil08]. Louis [Wel05]. love [Hat10]. low [McG21, Wal22c, Van04b]. Lowys [CS78]. Lubbock [Bar22c]. Ludwig [Blo08, Fla83, Hof13]. Ludwik [Löw08, Möß11a, Möß11b, vdBG90]. Lugg [Bai92]. Lukes [Blo82c, Hes82]. lumen [RK06]. Lunar [Ale98, Ano01b, Ari84, Ari01, Wil01]. Lutheran [Kal96]. Lützen [Pre07]. lux [RK06]. Luxembourg [Osl71]. Lyell [Lau82b]. Lyme [Ell21]. Lynch [Rhe94b]. Lyon [Péc11]. Lyotard [Gra96, NI03].

M [AR22, Mor90a, Rob22]. M. [Liv09]. M.I.T [Kir73]. M51 [Kes07]. MA [Gui04a, Wit24]. Maarten [Pal09b]. MacCord [Die24, Hus24, MM24a]. Macdonald [Gab73]. Mach [Swo74, Bak23, Bla72, Fay24a, Fey84, Goe72, Hof13, Lop84, Mit93, Pal08a, Pre08a, Thé21, Hen85]. Machamer [Fey74]. Machian [Hoe94]. machine [BG23, Fri24a, Hat07, Nor20b, Ras96, Rat20, Car13]. machinery [GKS24, Moy77]. machines [Coo05, Gua01]. Maclaurin [Mor22]. Macmillan [Blo75]. macro [Dar23]. macroevolution [Bro16a]. Macromolecules [Zan88]. macroscopic [Gao21, Wol20b]. Madame [Ilt77]. Madden [Ege70b]. Madison [Ege70a, Gra72, Gro90]. Madness [AR22, Edi22, Rob22]. Magic [AG09]. magical [Ver07a]. magnete [Geo14]. magnetism [Sim70]. Magnitude [Miy17, Eva10]. magnitudes [Fay24b]. Maienschein [Die24, Hus24]. Maimon [Pri18]. Maimonides [Lan20]. mainstream [Gra24, Lar24]. maintenance [BDB21]. major [Sol08b]. make [CD15, Cul06, Don21, Jon13, Kle79]. Makes [Hav21]. Making [Cel13, Cok15, Cun01b, Maa99, PvF14, Rap11b, Ras96, Sch93, WB20, Bor15b, BM23, CR13, Cun00, Goc08, KH17, LF25, MTC17, Ste14, Top04, Tyb03, Wei07, WS21, Wil09b, dC02]. malady [GMC22]. malaria [GMC22]. Malaysian [Lee19]. Man [Maa99, Coh08b, Coo96, Dou22, Sch70, Sch88d, Ste24a, Way21, Wan12]. man-in-the-street [Ste24a]. management [HL13, JC21]. Managers [CS07]. Manchester [Mor90b, Mye89, Sum13]. Mandated [Lev01]. Mandeville [Car25]. maniere [Le 58]. manipulability [Kri18]. manipulation [LL19]. Mannheim [Blo73]. manual [Sto12]. manufacturing [Bow93]. manuscript [Buo24]. manuscripts [Gui04b]. Many [Bat23, Dar23, FR23, Mil23, Pal23, CS21, Gol97, Han21, Mor11, Sak22, Sam10, Sim17, SM24, Tri08]. Many-Body [Bat23, Dar23, FR23, Mil23, Pal23]. map [Dhe23, WHC23]. Mapping [MN18, Ale98]. maps [Ale98]. Marburg [Gio24b]. March [Ano09a, Ano11a, Ano12a, Ano14a]. Marey [Dou95, Hof13]. Margaret [Bro07, Bro11a]. margins [Hau22, Sto25]. Marian [Nic72]. Marina [Cle23, Eat23, Gol23]. Marine [Van95]. Marinus [Ose22]. Mark [Bak16, Hau22, Lin22a, Rot22, Tur22b]. markers [Fri08]. marketing [AG09]. Markets [Slo24]. Markus [San13]. Marshall [Gra72]. Martineau [CM00]. Martinus [McE76]. Marx [Yea85]. Mary [Blo75, Hal12, Hea24]. masculinity [AG09]. Mass [Bau71, Bra01, Bus71, Eas72, Elk74, Ful93, Jan73, Kir73, Liv93, Luc71, McR72, Rhe94b, Mar22, Ros17, MHW22]. Massachusetts [Sch91]. Massimi [Boo23, Ken23]. Material [Ear20, Pri22, Rei20, Bar20b, Coo96, Dav21a, JW13, LK20, LW13b, McC20, Sch70, Sch88d, ST20, Stu20, WG13, dG20]. Materialism [Whi71, Dun10, Mar16, Zam18]. materials [Kle05]. mathematica [New87, New09]. mathematicae [She87]. Mathematical [Col07a, Dun13, Gel11a, JC20, Pas11, Per91, She06, Alm12, Bak21, Bia20, BK17, Bue00, Cel08, Cor01, CP17, Cre12, Dea87, DS21, Eve18, Fie86, Fri03a, Fri24a, Gla86, Gla89b, Gla93, Gla99, Gla01a, Gla01b, Gla02, Gui04b, Hol21b, Jan00, Jes89, KC12, Lan18, Lan21, Mor22, NCW99, O’N98, Oos12, Pal09a, Pan96, Per08, Pri18, Ric86, Sha98, She93, She97, Ste04, Yap11, New87, New09]. mathematical-agricultural [Fri24a]. mathematicarum [Man92]. Mathematics [BO24, Mal91, Ada16, Ale95, Bab25, Bau72, Blo73, Bue19, Cat12, Cel13, Cor97, Cor93, Cul06, Dar03, Dom09b, Ede09, Fra90, Gla89a, Gla89b, Gla01a, Gla01b, Gro87, HS23b, Kim24, Koe95, Lan21, Mah77, Mal01a, Man92, MB22, OP18, Pal09a, Phi05, Pou00, Ris90, Rob03, RT95, Ste21, War92, War93, Win81, dJ97]. Mathematik [Neu71]. mathématique [CL05c, Cul06]. mathematischen [Neu71]. mathematization [EC25, Mal91, Pal03, Yee24]. Mathematizing [Fek12]. matrix [vL24]. matter [Cle97, DW23, Fal98, Gau16, Haz23, MSS^{+}22, Mas11, McG21, Pap77, Par20a, Sch70, Sch88d, Sie07, Vor11, Kni71]. matter-first [Haz23]. mattered [Pic22]. matters [DM13, PS10, Ryc15]. Matthew [Hav21, McH21, Wie21, Str06]. Matthews [Sie04]. mature [Díe97b]. maturity [Swi86]. Maupertuis [Wel05]. Maurice [Cra93, Sha90b]. Mauss [Blo82a]. maverick [Hee19]. Mavericks [Avi19]. Max [Bra01]. maximizing [Pin16]. Maxwell [Bok15, Cat12, Cat14, Cha73, Cha86, Che23, Col90, Dau70, Dor70, Fal17, Hei70, HG21, Mit17, Mor92, Moy77, Moy78, Nap21, Saa05, Sim70]. may [CT13, Dic08, San13, Ano72d]. Maya [Chu22, Joh22, Lea22, WV22]. Mazingira [GMC22]. Mbeki [Wei07]. McAllister [TGD05]. McGraw [DB96]. McGucken [Ano73a]. McQueen [Fag24]. me [Kuu11, SLH23, She93]. Meadows [Zap24]. Meaning [BRdS96, McO96, SBM09]. meanings [Pap77]. Measure [Lon01, Dar03, Det22]. Measurement [Bor25, Bra01, HR01, Ans16, Ban13, Bas17, Díe97a, Díe97b, Fal17, Gré19, JW22, Ker12, KH25, Luc20, MCGP17, MBC17, Mic93, MTC17, Miy17, Ohn21, Ohn22, Pap24, Péc11, RS21a, She11, Tal17, Wol20b, dCG17]. Measurements [OS14, FD21, Mit17]. Measures [Hag01, HK21, Rio15, Ste22b]. Measuring [CT14, McO01, Dav20]. Mechanical [Maa99, McL06a, Pal08b, Cha86, Cha93, Cha02, Mat19, Mit17, Pyl02, Sib95, Cox12]. Mechanicism [BR23, Eat23, Gol23]. Mechanics [DD69b, Gau16, Heu14, Rie24, TC25, vE14, Ano91i, Bab25, Bar21a, Blo08, BJ22, Coe22a, Cox12, EZ13, FJ21, Fei14, For88, HS22, Ilt77, Mar99, Men24, Mit93, Mit22, Moy78, Mul24, Pre07, Pre08a, Rei18, Sta13, Tru68, Wal22a, Wat98, Wes79, DE21, Dun22, Ait70, Ege70a, Eis21]. Mechanism [Cle23, Yli19, Ans02, Car12, Coe22a, Kon24b, Miy17, Ott08]. Mechanism-based [Yli19]. Mechanisms [Sta18, Bec16, BB22b, Cun21, Fag15, GW15, PE24, PG24, PM22, Sch13c, vE19]. Mechanistic [HK22, dR06, vEdJ16, BA10, De 11, EHPL22, Gon25, Käs18, KM14, MR19, Mat20]. Mediating [Wol20b]. Medical [Pie97, Car95, Hei24, HK21, UG20, dC12]. medically [Kon24b]. medically-unexplained [Kon24b]. medicine [Ano09f, BB23, Bir10, Cul04, Ful24, Hei24, Kle04, LS02, Mac00, Pri22, Tyb22, Wie21]. medico [De 02]. medico-legal [De 02]. Medicus [Sie10, Pes13]. medieval [Ari84, Bia94, Cla68, Cro90, Gra72, RK06, Wal91]. mediocrity [KL21]. medium [HS23b]. meet [BB23]. Meets [Kra22a, McH21, Bro11b, Cer19, JW22]. mega [Ben18]. mega-geomorphology [Ben18]. Meinong [dC10]. Melanchthonized [Kal96]. Mémoires [Spa91]. memories [Buc89]. memory [Aub08, Cha10a]. men [Cho22]. Mendel [Rad22]. Mendeleev [Bar05, NRB04, Sch08b, Sch14a]. Mendelian [Cha13, CR13, Van86]. Mendelism [Rad13, Sha21]. Mendelssohn [Fal98]. Menger [Mäk97]. Mental [Pal03, VABF24, Dra18, Gil83, Pic04]. mentalities [Bre04b, Chi01]. mentioned [Gio14]. merger [ALR21]. merit [Car14]. meso [Dar23]. mesosomes [Ras93]. messenger [Eld24]. messiness [Fed23, SFT23]. messy [JW22]. Meta [BH19, Daw22, McC21, Bar09, Daw21, Kal75, Maz22, Pap10, Psi94, ST20, Wol24]. Meta-analysis [BH19, Maz22]. Meta-empirical [Daw22, McC21, Daw21, Wol24]. meta-helikôn [Bar09]. meta-induction [Pap10, Psi94, ST20]. meta-theory [Kal75]. metadata [Ben22]. metaphilosophy [Mur93]. metaphor [HN16, Kam94, LB94, Mar97, TC91]. metaphors [CP18, Mat19, Mir89]. Metaphysical [Brä24b, Cal21, Cha20a, Cha20b, Tos07, Bab25, Cat12, Hei74, Men24, Sch21b]. Metaphysics [Buc69a, Buc69b, Ris93, Ahn04, Bro16c, Cat12, Cha19a, CL05b, CL05a, Cla79, Dom09b, DE21, Ere91, Haz23, Hul89, Ilt77, Jac23, Lev11, Mal01a, Nee96, Non94, Pri18, Stu17, Ver07a, Win82, McR72]. metatheories [Mac72]. Metchnikoff [TC91, Kam94]. meteorology [Pin09, Zwi11]. Meters [Ara13]. method [Bok15, BK17, Cat95, Cle83, Cur16, Di 11, DiS13, Dra78, Dun13, Fal17, Fal18, Gay78, Ghi69a, Ghi69b, Ghi72, Ghi84, Ghi03, GW73, Goo87, Hei75, Hor73, Lin22b, Mer80, Mor21, Ris93, Rud74, Sch07c, Sch18b, Stö15, Wan22, dC93, Ege71]. Methode [Gio16]. methodenstreit [Mäk97]. Methodological [Bro73, Buc72, DW23, Mit72b, Qui72, Tie88, Ans04, Can71, Ell12, GG88, Lac21, Leu15, Mit72a, Wra15, Zha22, Col16]. Methodologies [Fre80a, Gay76]. Methodology [Gla78, Ano81c, Ano91k, Cha73, Cra19, Dor79, Duc10, Fin81, Hol83, HG21, Lau82b, Qui72, Red80a, Sar81, Sch12, Sch85]. methods [Bor25, CM22, HC23, Joh11, KF23, Mer82, Pic03, SGO17, Ull24, Zah19]. Metre [Ker12]. metrological [Bac20]. Metzger [Aba08b, Aba08a, Ano84g, BV05, Chi01]. Meyer [Whi74]. Meyerson [BV05, Bia88]. Michael [Ege71, Ere91, Heu14, Mel04, Mun74, Rhe94b, Sie04, JM11, Jha97]. Michalos [McE76]. Michel [Ano91b]. Michela [Boo23, Ken23]. Micro [Dar23, Sch09a]. micro-history [Sch09a]. Micro-meso-macro [Dar23]. Microbes [OP18]. microbiome [FBW22, PP23]. micrometry [Dav20]. Micropolitics [Hag01]. microscope [Jac95, Jac97, Ras93]. microscopic [Wol20b]. microscopy [Jar09, Ras96]. mid [Bru20, Har19b, How04, Ric86]. mid-level [Har19b]. mid-twentieth [Bru20]. mid-Victorian [How04, Ric86]. Middle [AR22, Bat23, Dar23, Edi22, FR23, Kos93, Mil23, Pal23, PR90, Rob22, Kai19, Woo24, Fun86]. Middle-range [Kos93, Kai19]. Middlesex [Sha90a]. Mieke [Mas24]. might [And13, BR19, San22]. Migration [BT19]. migrations [Hoc87]. Military [Goo13]. Mill [Ach92, Cob11, Hol83, Los83, Men18, Pfe12]. Milton [Pic22, Sch12]. Milwaukee [Gra72]. Mind [Cha20a, Cha20b, Joh12a, Ros70, Sch70, Sch88d, Cir93, Dem06, HS23a, MM06, Mas15, Nis93, Suá12a, Kni71, Luc71]. mind-body [Dem06]. mind-dependence [Mas15]. Mind-World [Cha20a, Cha20b]. Minds [Coo05, Jas04, Tha93]. Minimal [PE24, Fan24, Sim21]. Minkowski [Sla22]. mirabilis [Pal71, Elk74]. miracle [Car91, Car93, Mee92, Men14]. Miracles [Wra18, How15, How15]. mirage [Kim95]. Mirandola [Man12a]. Mirowski [Han05, Ric05, Ueb05b, Wra05]. mirror [Sch07a, AD12]. mirrors [Kan96]. Mis [Lic21]. misconceptions [Bak23]. misconstrual [Lip20]. miseries [JFS97]. Mises [Bla24]. Misleading [Tur05a]. missed [CL04]. missing [Ekb24, MT22, Rot13]. mission [Car03a]. Mississippi [Boe22]. mistake [Ray90]. mistakes [Sch93]. Mitchell [Rhe94a]. Mixed [Bab25, Mia01, Ada16]. mixture [Nee02, WW04]. mixure [Nee08]. Mizrahi [Mar21, Miz21, Tur21]. ML [Räz24]. mobile [Fre02]. modal [Cal22, Fut02]. modality [Eng18]. Model [BCK^{+}22, Gel11b, Hin15, Isa14, Llo15, Pol21, AL11, Bol15, Cha86, DH22, Gay78, Gel11a, Gre24b, GYM14, Hac91, HS23b, HG21, JB16, JC20, JC21, Ken12, Knu11, KM14, Lip92, Mat11, O’L21, Pot12, Pri19, Sch87, Sha10, Ste14, VFHS23, Wat21, Boe22, CSD21, Sha97]. model-as-instrument [Pol21]. Model-as-replica [Pol21]. Model-based [Gel11b, Gel11a, Knu11, Pot12]. model-building [Mat11]. model-making [Ste14]. Modèle [Spa91, Sig90]. Modeling [DZ14, Yos23, BA10, Fag16, Har19b, HG21, LB22, Pac18, PvF14, PZ14]. Modelling [Knu11, KD19, Tal17, Zuc19, CSD21, Cro90, Cro91, Eve18, Gol75, HZ19, Pap20, Suá12a]. Models [BT19, Bro09, Chi11, GY11, KM14, MBC17, bAM11, Sal17, TJT16, Wei99, Bec16, BG23, Cap79, Cha10b, Con06b, De 14, Ega10, Fag16, Fan24, Fei14, Fis06, Gra11, Hin13, Jac21, KB20, Len14, Llo12, Mor11, OP18, Ord16, Pal03, Pat09, Pin09, Pol21, Por11, Spa23, VJ19, Vez16, Vor11, Win12, dOFGB22]. Modern [Abl15, Bel20, Blu21a, Cha20a, Cha20b, Gol23, Win99, Wu14, Bak16, Bie20, Boa11, Bre04b, Cam12, Car03b, Cha10c, De 96b, Fey16a, Gol75, Gue17, Hur17, Jac97, Jar79, Joh98, Joy97, Kub16, Lat93, Llo97, Mal13, Oos12, Pre98, Ram12b, Rap11a, Rog95, Rot11, Ser99, Smi15, Coh97]. modernen [Fey16b]. modernism [Bra08]. Modernity [AR22, Edi22, Rob22, Rou91]. Modes [RC17]. modest [Ega10]. modified [MSS^{+}22, Fri24b]. Modus [Ste24b]. Mogg [Tay09]. Molecular [Hei70, Abe24, Bru99b, Cam21, Fis06, dCR24]. molecularizing [CCB24]. molecules [Cat14, O’M16, Too11]. moles [Wol20b]. moment [Miy17]. momentum [And94, GG21, Joh21b]. Monadology [Smi13a]. monads [Smi13a]. Monas [Cam12]. MOND [DW23, Jac21]. MOND/dark [DW23]. monetary [Sch08a]. money [Coo96]. Monge [Gla86]. monism [Sch11]. monisms [Blo07]. Monographs [Kam94]. monomania [Sim01]. monopole [Kra81]. Monopoles [Hen83]. monopoly [Sch12]. monsters [Vot14]. Montanari [Van16]. Moon [Ale98, Ver07b]. Moore [Pre08b]. Moral [Bro21, FA13, Hav21, McH21, Buc21, Cat12, Fri08, Mor93, Phi05, Tre07, Wie21]. morals [Ste14]. Mordente [Ros19]. Moreau [Wel05]. Morgan [Phi05]. Mormann [Fri12]. morphisms [Díe97a]. morphological [FH23]. mosaic [vH91]. Moses [Lan20]. most [Pal05]. Moti [Mar21, Miz21, Tur21]. motion [Alg08, Bai87, Big08, Car15, Dom09b, Gal96, Har24, Jan15, Le 58, Lev11, Mel91, Mol82, Péc11, Rei18, Ryn95a, Ryn95b, Sch03a, Whi71]. motions [Cla68, Gra72, Cla68]. Motivational [Gio22]. motivations [Dra86]. motuum [Cla68, Gra72]. mouvement [Le 58]. move [BT19]. movement [Axt93, Gra04, Le 58]. movements [Gla97, Rol16, vE19]. Moving [Gre24b, Wil77]. Ms. [Har74]. Müller [Isa19]. Multi [JC20, Wat21, Eld24]. multi-messenger [Eld24]. Multi-model [JC20, Wat21]. multifaceted [CL04]. multimessenger [Abe22]. Multiple [Aiz18, Man18, TC21, Zin11, CF06, Ell13, Fag16, Yos23, Zah19]. multiplicity [RK06]. Munksgaard [Osl71]. Munsell [Coc14]. museum [Jon13, AD12, dS09]. Museums [WG13, Maa13, SBM09]. Musgrave [Fin73]. Music [Coh85]. must [Gio14, Spe04]. mutual [Dar23, Kri18]. my [Gin09]. Myrtle [DB96]. mysterious [Haw96]. mysticism [Mar16]. Myth [Dav03, Car12, Fin02]. mythical [CL04]. mythology [Sch12].

N [Hei75, Mor90a, Mye89, Row08a]. N-rays [Row08a]. Nagel [Sar15a]. Nambu [dOFGB22]. Napoleon [Car13]. narrate [Ros17]. narrating [Ben18]. Narrative [Bea17, Chi15, Hur17, Kin15a, Mor17, MW17, Ter17, Ram08, Rot17, Swa21, Wis17]. Narratives [Hes21, dCR24, Rou91]. Narratology [Cla95]. Narrow [Bol14]. Nash [dB09]. Nasr [Bau71]. Nathan [Cul04]. Nation [Wu14]. National [Hau22, Lin22a, Rot22, Sol22, Tur22b, Ama05, Cha13, Clu12b, Dav21b, Pie97, VK19]. Nationalism [Wu14]. Natorp [Edg08]. Natural [Bel20, BMBM20, Bou20, Buo24, Dio18, Kal96, LW94, Lip20, Liv93, Mac00, MW25, O’F10b, Pal02, PEK06, Pol15, Rus71, Wal16, Way21, Ada16, Bha15, BC01, Car03a, Car93, Cat12, Cha44, Cha69, Cle23, Con21, Coo20, Cun01a, DK18, De 96a, Dea01a, Dio13, Düp15, Ege70b, Eng18, Ful93, Gam21a, Gre92a, Gre92b, Gui04a, Hat90, Hee98, Ili04, Jac94, Joy97, KH20, Leh06, Lem21, LA22, Mas13, McC01, McL24, McN22, McO96, Mor05, Mou20, New87, NCW99, New09, O’C19, Osl92, Pum12, Sha02, Ste04, Ter17, Top04, vdB24, Fyf02, Gua13, O’F10a, O’F10b]. naturalis [New87, New09]. Naturalism [Psi97, Bir04, DS18, Men24, Sie89, Sie90, Sol95, Stu91a]. naturalist [Hoo13]. naturalistic [Ueb91a]. Naturalized [Ano91h, Fre05, Sie89, Sto09b]. naturalizing [HHL11]. naturalness [Cur92]. Nature [Bel20, Cun01b, Gre92a, HKMV11, McO01, tH00, Asp13, Bak06, Bal20, Car25, Cel08, Chr94, Col90, CE19, Cun00, De 16, Dör94, Dri24, Eng18, EZ13, Esp79, Fek24, Fin90, FP99, FWW09, Gam21b, Gau16, Gie92, Gon25, Gre22, Har08, Har96, Hut21, Jan15, KH17, KM06, Lan06, Lan20, Leh06, Leh20, Lin22b, Lou08, LW13a, LJ11, MD13, McG70, bAM11, Nap21, Nob23, Obe16, PEK06, Pal03, Sch11, Sch18b, Sch03b, Sno04a, Ste21, Vae11, Wat12, WC09, Yea05, dC02, Heu14, Liv93]. Naturelle [Spa91, Sig90]. natures [VH06, KH17]. naturphilosophie [Gow73, Sha89, Bei03, Ham00, Hen98, Slo19]. Nauenberg [Mel04]. Nautical [Dun12]. Navigating [VABF24]. navigation [Ben87, Hac89, WHC23]. nebulae [Kes07, Nas11]. nec [Ker12]. necessarily [Gin07]. Necessary [Fis85]. Necessity [Joy97, Mar21, Cat12, Dom13, Gel80, Tre87]. need [Bau98, MR19]. needlessly [Ver17]. needs [BB22a]. negative [BM14, LAL21, Len21]. neglected [SM24]. Negotiated [Hag01]. Negotiating [BM20]. neighborhood [Pes13]. Neither [Bau15]. Neo [HHO09, Bei08, Bog21, Gio24b, Lip20, Pul01]. neo-essentialist [Lip20]. neo-Kantian [Bog21, Gio24b]. Neo-Kantianism [HHO09, Bei08]. neo-Kantians [Pul01]. neoclassical [Mir89]. neoplatonism [Sam77].

Neptune [Bam96]. Nerve [Isa19]. nervous [Bra24a, Nir23]. ness [Sch24]. nest [BS21]. Netherlands [vdW04]. Network [DH22]. networks [Ara13, Eva07, KM14, Ram12b]. neuf [CL05c, Cul06]. Neumann [Mit22]. Neurath [Ben22, Cat95, PY06, Rei97, Ric09a, Ueb91a, Ueb91b, Ueb09, Ueb10, Zem06, dC18]. neurodiversity [Koi21]. neurons [Sch07a]. Neurophysiology [Nir23]. Neuroscience [Kra22a, HS23a, Hoc16, Ull24, WHMO22]. neuroscientific [CDC24, Hau21]. neutral [Hon87b, Let21, Lus21, Pic84]. neutral-weak-current [Hon87b]. never [Coh97, Kle15, Lat93]. Newman [Cha10c, Cha11]. Newton [Elk74, Gui04a, Hen14a, Pal71, Sch11, Ste04, Ach23, Ano86h, Ans04, Art95, Bai87, Bie16, Bra13, Bro16b, BC01, Dem06, Di 11, DiS13, Dom13, Duc05, Duc11, FM04, FP99, FH85, Fra05, Gro87, Gui04b, Hay04, Haz23, Hen11, Hen14a, Ili04, Jan00, Jan13a, Jan15, Jan13b, Koc09, Lay78, Man04, Mas13, McG70, Mel91, Pal87b, Par20a, Pou07, Rib85, Ryn95a, Ryn95b, Sak22, Sch11, Sch13b, Sha74, Sno04b, Spe04, Sto25, Top90, Win86, Sno04a]. Newtonian [Mel04, Cal10, Che23, Con21, DN00, Dav03, Dun13, For88, Hay04, Mas11, Smi13b]. Newtonianism [Ahn04, Bai04, Hut04, Man04, Wat13, You04, vdW04]. Nicholas [Cam22, Kra22a, Rup22, Mou20]. Nicole [Cla68, Gra72]. Niels [Hon82]. nihilism [UG20]. Niiniluoto [Row15]. Nijhoff [McE76]. nil [Hon87a]. Nineteenth [McG69, Ben74, Buc09, Bur02, Che95, Das86, Edg08, FK90, Fyf02, GW73, Gre92a, Har96, Hei75, Liv93, Lor20, Mor07, Nye76, Péc11, Pre08b, Top00, dC93, Ano73a]. Nineteenth-century [McG69, Ben74, Buc09, Bur02, Che95, FK90, Fyf02, Har96, Lor20, Nye76, Pre08b, Top00, dC93, Ano73a]. NLG180.00 [Sie04]. No [Bri73, Dar21a, Fal17, Lat90, LL19, Mje24, Mos01, Str13, Wag22, Bîg13, Cho22, Dou21, How15, How15, Men19, Wra18]. No-go [Dar21a]. no-miracles [How15, How15]. no-overlap [Bîg13]. Noble [Gay77]. noise [KL14]. nombre [Le 58]. nomenclature [Gua13]. nomic [Kui14]. nominalist [Dic06]. nomological [Lip92]. Non [Bat23, Cov97, Dar23, Fag24, FR23, Gri91, Mar22, Mil23, Pal23, RDM25, SHB94, Ahn20, Blu21b, Brä24b, Dor70, Fri14, Fut02, GID21, Gon25, Ken12, KVJ22, Reu22, Rit21, Rui23, WB20, dRM23, vL24]. non- [Brä24b, Fri14]. Non-accessible [Mar22]. non-arbitrariness [vL24]. Non-basic [Cov97]. non-causal [KVJ22]. Non-empirical [RDM25, Blu21b, Rit21, Rui23]. non-epistemic [Ahn20, GID21, Reu22]. Non-Euclidean [Gri91]. non-factual [WB20]. Non-Fundamental [Bat23, Dar23, FR23, Mil23, Pal23]. Non-locality [SHB94]. non-mechanistic [Gon25]. non-modal [Fut02]. Non-separability [Fag24]. non-separable [dRM23]. non-speculative [Dor70]. noncommutative [Kon22]. nonconservation [Fra79b]. noncontextuality [HS22]. nondiscovery [Fra79b]. nonseparability [Let22]. Nonseparable [Hea94]. nor [Bau15]. Norbert [Alg08]. Nordmann [Fri12]. normal [Ano91b, Can89, Díe07, Kri80]. Norman [Pes13]. Normative [Tur05b, Bau98, Ful93, Hat90, Sie90, Tei06]. Normativity [Dic13, Coh18, Fra06, Sch09b]. Norms [Ris98, Eig19, Fed23, PG20, PvF14, PZ14, Rot13, Sla17]. North [RH11b]. northern [Koc15a]. Norton [Bar20b, Cor21, Dav21a, Gen20]. Nortons [Ske20]. nose [Bar18, Rem07]. notations [DS21]. note [Ano70f, Ano70g, Ano72f, Ano75d, Ege70b, Har74, Mok71, Tot12, Zan82]. notebooks [Kas04]. Notes [Ano71i, Ano72h, Ano73f, Ano73g, Ano74g, Ano74h, Ano74i, Ano74j, Ano75f, Ano75g, Ano75h, Ano76b, Ano76c, Waj21, Sha90b, Sto09a]. nothing [Col03]. Notion [Mer04, Bue22, Duc05, Eis21, Kon24b, Obe05, Pre08b, Sch03b]. notorious [Ste24a]. nouvelle [Le 58]. Nova [Miy15]. Novel [Nun84, Ros18, VFS14, CD15, DM13, Goo86, Lus16, Mur89, Ree12, Sch08b, Vot14]. novelties [Rat20]. Novelty [Sch14a, Cha15a, Sch14b]. November [FSJ08, Par12]. NSF [Tur22b, Ano80f, Hau22, Rot22]. nu [Mit17]. Nuclear [Lee19, Wit13]. nucleation [Hoc87]. nucleotide [Oft22]. null [Fal17]. Nullius [Tuc12]. Number [Dar03, Kno76, Le 58, Man12a, Top90]. numbers [Díe97a, Díe97b, Kat20, O’N86]. numerical [Gra11]. Nuncio [KR08]. Nunes [Alm12]. Nursing [Fed23]. Nutter [Sch12]. NY [Ere91].

O [Bro71, Mau13]. O. [Bol14]. Oberheim [San09a]. Obesity [HM24]. object [Che03, Che10, Eag07, Sak22, Sch09a, ST20, Ste03, Thr75]. object-induction [ST20]. Objection [Del17]. objections [Bro71, DH69, dC10]. Objective [Koc15a]. Objectivity [Can98, Kid11a, Vot14, Ful94, Kit93, Leu12, Nor12, O’L20, Pin16, Rio15, Sci21, Wil22a]. Objects [AD12, Ara03, Mos07, BK21, Pal08b, SBM09, Vos07]. obscure [Mat80]. observability [Con06a]. observable [Tur05a]. observables [Kas14, Pit22]. Observation [Leh04a, MHW22, Bar21a, Gre96, Kri80, She88]. Observational [Rap80, CF21b]. observations [Ari84, Ari01, Goo86, Hud94, Sim70, Wec86, Wer80, Wil01, Ano01b]. Observatoire [Péc11]. observed [Den00, Rup76]. Occult [Par12]. Oceania [Dou22]. October [FSJ08, Ano07a, Ano22j, Ano23i, Ano24h]. odd [Sla17]. Odysseys [Bud13]. Oeconomic [Tri05]. Oersted [Sha89]. Official [Hag01]. offs [Mat11, PP23]. Oh [Kon24a]. Ohm [Hei80, Luc20]. oincidence [Ryc92]. oint [Ryc92]. OK [Mir05]. Old [Gar89, Pin97, DJ14, Mar15b, Pan99, Wei21b]. older [Set08]. Olin [Mis94]. omnipotence [Fun75]. Once [De 16]. One [Mor11, Sie00, SG85, Ant25, Col08, Cus85, Des49, Gin09, Gio14, Kra22b, Reu22, Tri08, Wag22]. one-world [Kra22b]. ongoing [Bar18]. only [Sny07]. Onnes [GG88]. ontic [Dun22, Wri15]. Ontological [Kha24, Sim14, Con06a, Fei14, HHO09, Lud16, Sch21b]. ontologically [Let21]. ontologies [Kle05, Lud16]. Ontology [Bra01, LSH24, Bor25, Bue19, Cer19, CL05b, FL22, Gli16, HM06, Mar22, Sch09b]. Openness [Dea03]. Operational [Adl22, Kra72, Let21, MB22, Moy91a, Moy91b]. Oppenheim [Wag24]. opportunism [Gay77]. opportunity [CL04]. Opposition [Sto25, Wal06]. ops [Aro18]. optica [KR08]. optical [DK09, Goo87, Hin15, Jac94, Mal90]. optics [Cro90, Cro91, Gal96, Mal01a, Sha74, Smi15, Bak16, Web22]. Optik [KR08]. optimality [ST20]. Optimisation [Mor21]. option [Fra08]. options [Ach23, Brä24b]. Opuscules [Pas10]. orbitals [Bru99b]. orbits [DJ14, Gow87, Pou07]. Order [KH20, Rhe25, Kat20, Lan20, Mou20]. ordered [Sto23]. ordering [Mor17]. Ordinary [Fra21a]. Øre [Eas72]. Oresme [Gra72, Cla68]. organic [Bro73, Fis06, Sla02]. organised [Fri03b]. organism [GN22]. organisms [AL11, BB23, BB22a, Gon25, HM88, Pra74]. Organization [BB22a, Cam22, Aub08, Gam21b, Zam03]. organized [Mar24]. orientation [GW15]. oriented [Mat11]. orienting [Woo15]. Origin [Rus71, BDB21, Gie77, Gol75]. original [Zam03]. Origines [Spa91, Sig90]. Origins [Pen22c, Alm12, Ans04, Buc69a, Buc69b, Cha10c, Fra90, Fra21c, GS02, Gur19, Har02, Kam94, McR72, Mic93, Obe05, Osb89, Pre08b, Ras96, Riv21, Sha02, TC91, Tuc20, vS14, Buc88, Moy91a]. Ørsted [LJ11]. Orta [dC12]. Orthodox [TC25]. orthodoxy [Sch21b]. Oscar [Mor93]. Oscillation [Pin09]. Other [Hau22, Kid17, Lin22a, Rot22, Sol22, CS07, Far07, Le 58, Nii11, Osb89, Tur22b]. Otto [Ben22, Ueb91b, tH93]. ought [Eli10]. our [Bec84, De 18, Har02, Lea22, RM16, Sol08a, Sti13]. Outcome [GG94, MBC17]. Outlines [Ger14, Ale98]. outlook [dL09]. Overdetermined [Lug78, Van79]. overlap [Bîg13]. Overlapping [Lud16]. Overmathematisation [dB09]. own [Eag07]. ownership [MR13]. Oxford [Ful94, Hac89, Jas04, Kam94, McR72, Mul94, Rog72]. Øystein [Eas72].

P [Kir73, Ryc92, Wig67, Wig70, Wig78]. P. [Moy91a, Moy91b]. Paduan [Dra86]. Pages [Ano22k, Ano22h, Ano22g, Ano22i, Ano22j, Ano23l, Ano23g, Ano23j, Ano23h, Ano23k, Ano23i, Ano24i, Ano24l, Ano24g, Ano24k, Ano24j, Ano24h, Ano25c, Ano25d, Swo74]. pain [Chi11, Kon24b, vdLBV22]. painting [Fin11, Har96]. paleobiogeographical [Fra81]. paleoclimate [Wat23]. paleontology [Tur05a]. Paley [Fyf02, O’F10a, O’F10b]. Palter [Elk74]. Pandemics [Tur21]. Paola [Cle23, Eat23, Gol23]. Paolo [Cam12]. Paper [Gro90, Jar17, Kle01, Feh07, Gro05]. Paperback [Ere91, Hop94, Mis94, Rhe94a, Rhe94b, Spa91, Cul04, Mel04, Sie04, Ste04, Zam18]. papers [Hes82]. Papin [Fre02]. Paradigm [lC00, FH23, Kuu17, Rad97]. Paradigms [Ced83, LL81, Möß11a, Möß11b, Rab75, Sch12, Win12]. Paradise [Ber93]. paradox [Bie20, GS02, Tei06]. paradoxall [Clu12a]. paradoxes [Boy14]. paralipomena [Kep68, KR08, Fie86]. parallactic [O’G78]. parallax [Grü78]. parallel [Gra05]. paribus [Wri17]. Paris [Bri73, Cul06, Ano84g]. Parisian [Sto25]. parity [Fra79b, Sch02a]. Parker [Cra22]. Parmenidean [MS72]. pars [KR08]. Parsons [Mac10]. Part [Ano97g, Cha20a, Cha20b, Cun01b, Gam21b, Gam21a, Iva15, Row13, Sim70, Tos07, Wal22a, Wal22b, Bru99a, Bru99b, CL05b, CL05a, CE15, CEW16, Cro90, Cun00, Díe97a, Díe97b, Gla89a, Gla89b, Gla01a, Gla01b, McL05, McL06b, Moy91a, Moy91b, O’F10a, O’F10b, Ric09a, Ric09b, Sar97a, Sar97b, Tos06, War92, War93]. Partial [Lan14, Bue97, Con06b, McL06b]. participation [MBG21, Sor16]. participatory [DCJ^{+}21, LF25]. particle [Bor15b, Che90, Che95, CS21, Fal24, Kob24, Red80b, Wal22a]. particle-mechanics [Wal22a]. particle-wave [Che95]. particulars [Den00, MS72]. parting [Nor12]. party [Ast04]. Pascal [Cha15a, Jon01]. passage [Bak16, Smi15]. Passion [Jha97]. passions [Hat07]. passive [Par20a]. past [Bud13, Car05, Cha16, CT16, Dil90, O’M16, Pye08, Smi20, Sti13, Tur04]. Pasteur [Tem86]. pasteurization [Lat88, Sch91]. pasts [JFS15]. Patchwork [Hau21]. Patent [HL13]. patenting [Bid14, Goo13]. Patents [HLM13, MR13, Ara13, Bid14]. path [BM21, Fur22, Len14, Pou07, She99]. pathogens [Tur21]. pathological [Can89, Ano91b]. paths [GKS24]. pathways [RDM25]. Patient [MBG21]. Patrizi [Ede09]. patronage [GT05, Mei21, Pum12]. Patzig [Vea72]. Paul [GT05, Nic72, Bro16c, BK16, Edg08, Goc08, Han21, HH95, HH06, Kid11a, Mir89, Mun16, Obe05, Pre98, Pre16, Sor16, vS20]. Pauling [Har08]. Paving [HS23b]. Pbk [Mul94, Fin73]. pea [Rad13]. Pedro [Alm12]. peek [Sch07a]. Peer [Atk94, Dou10, San22]. Peirce [BP20]. Pence [ML18]. Penguin [Sha90a]. Pensées [Pas62, PT03, Pas10, Jon01]. People [McO01, Mor01, Far07, HR01]. perception [BZ81, BZ86, Ful93, Gau81, Hat88, Hat90, LL81, Löw08, Obe15]. Perceptions [GMC22]. perceptual [Per91, Ree86]. Perfect [Hav21]. performances [Sam17]. performed [De 18]. period [Díe97a]. periodic [Bru07, NRB04, SW01, Sce20, Sch08b, Sch14a, Wor05]. perish [Ven22]. Perlman [Wei10c]. permanence [Toa24]. Permissible [Str21]. Perpetuum [Fre02]. Perrin [Big08]. persistence [Bou21]. persistent [Rou91]. persona [Goc08]. personalism [Dor79]. personality [Ast04]. perspectival [FBW22, Boo23, Ken23]. perspective [All96b, AL16, Aya25, BCK^{+}22, Coe22b, Cor21, Dhe23, GFE23, Hon87b, Koc24, LB22, Moy91a, Moy91b, Ric92, RDM25, Sti13, Woo15]. perspectives [Bro09, Gel11b, HBR24, Käs18, Ken23, SLP23, Fra21b, BB23]. Perspectivism [Cha10b, Boo23, Bro09, Gie06, Gie09, Gie16]. pertinent [Grü78, LM12]. pervasive [De 16]. pessimistic [Pap10, Psi94]. Peter [Cha10a, Cun01a, FSJ08, Pit22]. Petersburg [Tei06]. petit [Le 58]. Petty [McC06]. Pflüger [Sch93]. Phaenomena [Gee14]. Phaidon [Hac89]. Phaidon/Christie’s [Hac89]. phallometry [JW22]. pharmacological [Tot12]. Phase [Kob24, Lyr14, May11, Due22]. phases [Ari87, Lee19]. phenomena [And94, AB23, Cha86, KB20, Leh20, Pic84, Pin09, Por11, Rod21, Sch07b]. Phenomenological [Gio14, Por11]. Phenomenology [Rhe25, Chi08, Koc11a, Mah77]. phenomenon [Col18, Mor11, Yos23]. Phenomenotechnique [Fab19, Chi08]. phenotypic [Koi21]. phi [WC09]. Philip [Ful94, Jas04, Sol95]. Phillips [Pic22]. Philo [Cre12]. Philoponus [Osb89, WW04]. philosopher [BS21, Fra94b, Osl92, Pum12]. Philosophers [Chr94, Gar96, Lux67, Hac92, Jac94, Osl71, Pla20, Rit12a, SD25, Cle23]. philosophiæ [Des44]. Philosophical [Ano90e, Aya25, Gal71, Gue17, Hel16, Jes89, Mer04, Mur93, Zam18, Ano86h, BB23, CL04, Col02b, Dar23, Duc10, Dup94, Fri98, Gei23, Gio24a, Hau16, Hon87b, Kat92, LGC21, Mat98, PY06, Psi94, Red80b, Rus75, SLP23, Sie04, Stu12, Tyb03, Vea72, dL09, vdLBV22, Des44]. Philosophie [New87, New09, Des49]. philosophies [Fra20, Har19a]. philosophizing [McG70]. Philosophy [Aba08b, Abl15, Ano78b, Ano80f, Ano86d, Ano87d, Ano97g, Ano01a, Ano01b, BR23, Blu21a, Bor15a, Buc88, Cel13, Cha16, Cle23, Col16, Eat23, Ere91, Ful89, Gam21b, Gol23, Heu14, Kam94, Kra22a, Kuk22, Mar21, Mas13, Myk22, OHH00, Par95, Pin14, Rad97, Rie14, Rou91, Sie89, Skl92, Sto09b, Ups22, Val22, Wra05, Zol93, Ada16, Ahn04, And22, Ano90b, Ano91c, Ano91d, Ano92c, Ano92d, Ano93d, Ano93e, Ano94c, Ano96a, Ano97a, Ano97b, Ano02b, Ano03a, Ano03f, AH15, AP04, Bal92, Bau72, BCC87, BV05, Ben24, BM21, Bra15, Buc69a, Buc69b, Buc89, BC01, BRdS96, Cat12, Cha93, Cha02, Chi06, Chi08, Cir93, Cob11, Coh08a, Con21, Coo23, Cor97, Cun01a, DK18, Dea01a, Des49, DMM83, Dre20, Dun13, EZ13, Esp79]. philosophy [Fis18, Fra21a, Fra90, Fri03a, Fut02, Gam21a, Gar00, Gio24a, Gla89a, Gla89b, Gon25, Gre92b, Hal05, Hal12, Han05, HC23, Hea24, Hei14, HK16, Hen16, Hon82, Ili04, Jac97, Jar03b, Jha97, Joh11, Kal96, Kin87, Kin15a, LL16, Lea91, Leu15, LC24b, Lon05, Los77, Los83, LA22, Luc71, Mac00, Man80b, MD13, Mat19, McC01, McL06a, McR72, Mir04, Mir05, Mor22, Mor97, Mor91, Mun16, Mun74, New10, New87, NCW99, New09, O’N03, Pal09a, Pan99, Pap24, Par12, Par92, Ped20, Pre98, Pyl02, Ric03, Ric05, Ric09b, RZ21, Rog72, Rog95, Ros70, RDM25, Rus73, Rus75, Rus89, Ryc15, Sch16, Sch21a, Sha97, Sha02, Sha20, Sie04, SG13, SGO17, Sto09a, Stu89, Stu12, Suá12b, Ueb05a, Ueb05b, VK19, Win81, Woo71, Woo94, Wra07, Zam03]. philosophy [Zan88, dC18, dJ97, vS20, vdB22, Ano84c, Gui04a, Mye89, Ste04, Gie91b, Kuk22, Mul94, Myk22, Ups22, Val22]. Phlogiston [Lew11]. photo [Hin15]. photo-optical [Hin15]. photography [BDR13, Hof13, Max22, Nas11, Bel15]. photons [Bha24]. phylogenetic [ES06]. phylogenetics [Wat21]. Physical [Cha20a, Cha20b, Ohn22, Bok15, Cat12, CF21a, Col07a, Fis06, Gie91a, Gré19, Gro20, Nap21, Nee08, Noa08, Pie14, Qua75, Ros22, Ver06, Smi13a]. Physicalism [Ful18, Ben22, Nee98]. physicalist [Das86]. Physically [Bra15]. physicist [Sch88b, Sta08]. physick [Mar88]. physico [Har08]. physico-chemical [Har08]. Physics [Ada16, Ano86d, Blu21a, Cha19a, Cla79, AG13, Bab25, Blu21b, Bor15b, Bra15, BS21, Chi08, Dai19, Dar03, DE21, EB11, Fal24, Fey70, Fey16a, Fri03a, Gla20, Gow73, Hat79, Kob24, Kos99, KSZ24, Man12b, Mar13, Moy91a, Moy91b, Non94, Red80b, Rit21, Rui23, RDM25, Ryc15, Skl92, Sta08, Tul16, War92, War93, Wri73, vD21, vS20, How04, Kub16, Bat23, Dar23, FR23, Mil23, Pal23, Mul94]. Physik [Fey16b]. physiological [Gla78, Stu08]. physiology [Bra08, Isa20, Len81, Slo77, dC93, Sch93]. Physique [Sig90, Spa91, Hut04]. Pico [Man12a]. picture [Goc08, Smi13b, Smi13a]. picture-making [Goc08]. Pictures [Ste05, Alg08, Mal90]. Picturing [Ver07b]. Piecing [All96a]. Pierre [Bha15, Dio13, Dio18, Iva10, Kid11b, San16b, Stu07, Wel05]. Pierre-Sylvain [San16b]. Pieters [Lin98]. Pietro [Sha90a]. pill [Cho22]. Pisa [Seg89]. pitfalls [Llo97]. pitting [dB09]. Pittsburgh [Gur19]. più [Pal05]. Place [Sam77, Mor12, Ryn95a, Ryn95b, Tur13]. places [CF06]. plain [Cun88]. Planck [Gar76, Gor91, Kri86]. Planckian [Sch21a]. plane [Wil09b]. Planet [Bie20]. planetary [Bai87, Car15, Gal96, Gow87, Mel91]. Planets [Gow87, Kos06]. planisphere [dS09]. planning [Bak22]. plans [Clu12b]. plant [CR13, Kle05, Pri22, WL22]. plants [SOC19]. plastics [Pri22]. Plate [MS24]. Plateau [Mil19]. plates [Gab73]. platform [Shr07]. Plato [Gre96, Het96, Mou81]. platonic [Kan96, Leh20]. Platonism [Ell92]. platypus [Sla17]. play [De 18, Leh18, Reu22]. played [Gra96]. playground [Ter21]. Playing [Too11, Rei08]. plea [Cra19, Isa14]. pleading [Cur16]. please [Rot96]. plenitude [Non94]. Pliny [Gua13]. Plotinus [RK06]. Pluralism [Leu12, vS20, Cra19, DS19, Eig19, Ger14, Han21, Hel16, JC21, Kha24, Kus15, Lac11, LGC21, O’N03, PP21, San14, Stu17]. pluralist [Bar22b, Bue12, Gie92]. Pluralistic [PK14]. plurality [Ken23, Sny07]. Pluralizing [Ohn22]. plus [Ker12, Le 58]. plusieurs [Le 58]. Plutarch [Cre12, Mül12]. Pluto [Sla17]. pneumatics [Sch88a]. poetic [Tot12]. Poetics [Cla92, O’C05]. poetry [Alt12, Cox12, FM04]. Poincaré [DBM22, Gie77, Gie91a, Gow97, Grü78, Iva15, Kri86, O’G77, O’G78, Sch20a, Stu89]. Poincaréan [Mac10]. point [Cha16, Col08, Gin09, Kon24b, Lou08, Men18, Zha22]. pointer [Kas14]. points [Daw22, Dil90]. Polanyi [Jac02, JM11, Jha97]. Polar [Péc11, Ohn21]. polarization [Che03]. policy [Don21, Isa14, MR19, PG24, SL21, WHMO22]. Polish [Löw08]. polite [Tay09]. Political [Cra19, Hau22, Ueb05a, Bak22, Car03a, Cra17, D’A00, Den00, Hel16, Hol83, Lus21, McC06, O’N03, Waj21]. Politics [Ano87d, Ano91i, Bid11, BS13, Joh21a, Kid16a, McO96, Mel21, Par12, Rob22, Ste22c]. polygenic [Ser23]. polymer [Zan88]. Pope [GT05]. Popkin [Sno04a]. Popper [McE76, Bam96, Bro11b, Cat95, Díe07, Díe11, Dom09a, Gay76, JM11, Jon78, Mic71b, Row11a, Row13, Tam15, Yea85, tH93, tH02]. Popperian [Col16, Gla01a, Gla01b]. Popularizing [Kni71]. Population [BP25, Myk22, Ups22, Val22, Kuk22]. portability [Mat22a]. Portugal [Car03a]. position [Myk22]. positive [Fra90, Fri18, Sch23, VABF24]. positivism [O’N03, Pre04, Sch20a]. Positivist [Ben24, Bir04, Pao11]. possibility [Bea17, Böh80, Elw12, Fee06, Len21, Nas16, Pra75, Pri18, Swa21]. possible [Cus85, Lan14, Le 58, Par92]. possibly [Rat20, VJ19]. Post [Pre98, AL16, Ant03, Ben24, Can98, Mac13, Sam17, Vot14]. post-Copernican [Ant03]. post-Kuhnian [AL16, Can98, Sam17]. Post-modern [Pre98]. post-positivist [Ben24]. post-war [Mac13]. Posterior [Gol04]. posteriori [Gol23]. Postmodern [Lat90, Par92, Rei97]. postmodernist [NI03]. postulate [Gra05]. postwar [BK17]. potential [SFF23]. pour [Le 58]. Poverty [How90]. Povich [Lan18]. powder [Sch09a]. power [CB92, De 96b, Jac03, Mal01a, Mat80, Pol21, Ric05]. powers [Hen14b]. Pox [Lea22]. pp [Ach92, Ait70, Ano73a, Ano91b, Ano01b, Bau71, Bia94, Blo75, Bro71, Bus71, Car73, Cra93, Cul04, Cul06, Eas72, Ege70a, Ege70b, Ege71, Elk74, Ere91, Fin73, Ful94, Ful93, Gab73, Gal71, Gie91b, Gra72, Gui04a, Gur19, Hac89, Hop94, Jan73, Jas04, Kam94, Kni71, Liv93, Luc71, McE76, McR72, Mel04, Mis94, Mor90a, Mor90b, Mul94, Mun74, Mye89, Neu71, Nic72, Osl71, Osl72, PR90, Pes13, Rhe94a, Rhe94b, Sch91, Sha90a, Sha90b, Sie04, Sno04a, Spa91, Ste04, Van18, Vea72, Whi74, Zam18, Zol93]. Practical [Fis23, O’N98]. Practice [Hav21, AH15, Bas17, BC92, Cam21, Cob11, Dar02, DD13, Dre20, Duc05, Duc10, Gel11b, Gel11a, Gow73, Hau21, Hea24, IRDV24, Jes89, Lov15, Luc20, LW90, MN18, Mag96, MBC17, MBG21, Nor98, Oos12, PS24, Sce20, Sch05, Suá15, Wis13, Rhe94b]. practice-turn [Dre20]. Practices [Win99, CSD21, Mét13, Pin97, SFT23, Vet18, dCR24]. Practicing [Ras93]. Practising [Ram12a]. pragmatic [De 11, Res00, Sch87, Zha22]. pragmaticized [Bog21]. pragmatics [Ueb15]. Pragmatism [Mis94, Oli92, Bid11, Mor12]. pragmatist [Bro15, Hei24, Ruy23]. praise [Pos71]. Prandtl [Blo08]. Pre [Atr85, Dra86, Llo97, Smi13b, Tuc93, Wat03]. pre-Critical [Wat03]. pre-modern [Llo97]. pre-Newtonian [Smi13b]. pre-Paduan [Dra86]. pre-science [Tuc93]. Pre-theoretical [Atr85]. precision [Gré19, Ker12, Sib95]. precursor [Wal75]. predict [tF18, Tom22]. Predicting [BG23]. Prediction [Car14, SW01, Wor05, Wor14, Aro18, DM13, Leh04a, Len22, LSH24, Nor20a, Par14, Str21]. Predictions [VFS14, Bar05, CD15, GG94, Men18, Sch08b, Vot14]. predictive [Car14, CT13]. Predictivism [Bru07, Sch24, Bar14, Lee13]. preface [Jar03a]. prefer [Col84, FH84]. preference [Hol22]. preferences [Lis18]. prehistory [Zan88]. prejudice [Cra84, dC10]. prejudices [Pan99]. Prel [Web07]. prelude [Nee96]. presagement [Jac02]. Prescott [For75]. prescriptive [Kav24]. prescriptive-dynamical [Kav24]. presence [SBM09]. present [Bra15, Bud13, Cha16, McK16, Smi20, Tos03]. present-centred [Tos03]. presentation [Cam22]. Presenting [Sch18b]. Presentism [Ano91j, Aba08a, Aba08b, Aba11, Bar22b, Loi16]. presents [JFS15]. preservation [Jar13, LW13a]. Presocratic [Alt12]. Press [Ano73a, Bak16, Bau71, Bro71, Bus71, Cra93, Cul04, Eas72, Ege70a, Ege70b, Ege71, Elk74, Ere91, Fin73, Ful94, Ful93, Gie91b, Gra72, Gro90, Gui04a, Gur19, Hei75, Hop94, Jan73, Jas04, Kam94, Kir73, Kni71, Liv93, Luc71, ML18, McR72, Mor90a, Mor90b, Mul94, Mye89, Osl72, PR90, Rhe94a, Rhe94b, Sch91, Sha90a, Sha90b, Ste04, Swo74, Vea72]. pressure [Sha74, Wor82]. preternatural [Mac00]. previously [Del22]. previously-unknown [Del22]. Price [Cul04, Cul06, Gui04a, Jas04, ML18, Mel04, Pes13, Sie04, Sno04a, Ste04, Zam18, Gur19]. pricing [VFHS23]. Priestley [Bas92, Boa07, Rob70]. Primary [Wei07]. primer [KL02].

Primitive [Smi84, Rup76]. primitivism [Kau08]. Princeton [PR90]. Princip [Gio16]. Principal [Mee21]. principe [Le 58]. principes [Des49]. Principia [Des44, New87, New09, Ste04, NCW99, Sch13b, Ste04, Bai87, Gro87, Mel91, Pou07, Sak22]. principle [Bîg13, BB22b, Cam97, Eve18, Fay24a, Fri14, Gam21a, GS22, Gor91, Gra04, Kin21, Le 58, McM93, Pru89, Tam23, Toa24, Zam03, Zem06, Mee21, Nor85]. Principles [Bel20, Des44, DMM83, Bra13, Des49, EG78, Fan24, Kra24, New87, NCW99, New09, Ste04, Cla79]. print [Top04]. printing [Neu71]. prior [AP04, Vea72, Pat69]. priori [Dom13, Fer12, Fri02, Gol23, Hei14, Pat09, Suá12b, Ueb12, Gio22]. Priority [RS21b, HLM13, MR13]. prisoners [Het96]. private [Eva07, Mel91]. privatization [Pin14]. privilege [RS21b]. prizes [Llo97]. probabilism [Fre15]. Probabilistic [KT21, Bel90]. Probability [AP06, BM14, Gow87, Run94, Ste15, Vas18, Wal06, Wei10a, Wen23]. probably [FH88]. probe [Sch16]. Problem [Cha20a, Cha20b, Lop84, Sch11, vE14, Bou20, Bru99a, Dor79, Dou14, Eng18, Fay24b, Fin81, Fra81, Ful90, Gau76, Gou23, Gre24b, HW22, Hol21b, HM06, Joh21a, KS19, Mal01b, Mar99, Mat22a, MT22, May14, Mor97, Ohn22, Ose22, Pap24, Red02, Res00, Reu22, San11, Sar81, Sie83, Ste24a, Suc90, Van86, Wag22, Koc09]. Problem-solving [Lop84, Fin81, Sar81]. problematic [LGC21]. Problems [Hee21, Kno76, Lau82a, Mye89, Qui72, tH93, Ach90, Ano90e, Bai87, BCC87, Car22, Dic13, Gri91, Koc11b, Lau88, Lug78, Nii11, Sam10, Sie89, Van79, WW04, tV21, Cox12]. Procedures [PMG21, Wat23]. proceedings [Bus71, Sch69]. Process [Cra17, Lew02, Alg08, All96a, Bor20, Che10, Gla89a, Mil19, Tal17]. processes [Hea94, Ord16]. processing [HK22, PS10]. procrastination [Fur22]. Producing [Gal96, Vos07]. production [Muk22, Rhe92b]. Productive [SD25, Bog08]. Productivity [Cam22, MR13]. products [GY11]. professional [Sch93]. professor [Rad13, Out96]. progeneration [Wil22b]. program [Ban13, Fri18, Vae11, Van86]. programme [Alm12, Col07b, Fre79, Gel80, Gla93, Had04, Ueb91a, WHMO22, Blo07, Feh07, Fre05, Gla01a, Gla01b, Kem05, dB09]. programmes [Dor79, Gla01b, Qui72, Red80a]. progress [Aiz18, Bec84, Bir08, De 14, DHS14, Del16, Del23, Dou14, Gin88, JW21, KN89, Kui14, Lep81, Los78, Met91, Nii14a, Rap80, Ric86, Row08a, Row15, Sch88b, Sol95, Spe17, Stu16, Tam15]. prohibit [Bid14]. Project [Meu19, Dac15, Loi16, Mot02, Mou81, MHW22]. Projective [Ric86]. projects [Meu19, PY06]. promises [Hee21]. promoting [Ram12a]. promotion [VABF24]. proof [Fal98, Saa08, Ser99]. Proofs [Har12]. propagation [Phi05, Sha74]. proper [Thr75]. properties [CS21, FL22, Hon87a, Nee08, Pra74, Ryn95a, Ryn95b]. Property [Miy23, Bor15b, CR13, GG94, Lip20]. proportion [Gro87, HG05, RT95]. proportional [Ros19]. Proportionality [Oft22, Mer80, Mer82]. proposal [GL24]. propositional [Pre08b]. propositions [WH14]. prose [Alt12].

Prospective [Bro90b]. Prospects [Hus24, Kuk01, Lin22b]. protein [Rhe92a]. proteins [Nea24]. Protestantism [Gre92b]. Proto [Cun01b, Bie20, Cun00, HH06]. Proto-Ecologist [Cun01b, Cun00]. proto-scientific [Bie20]. Proto-Structure [HH06]. Protocol [Ben22, Ueb09]. prototypical [Cun21, Cam18b]. Proudfoot [Res12]. provide [VJ19]. providentialist [Den00]. provincial [Sto25]. proximate [Pao11]. Przelecki [Nic72]. Pseudo [Leh20, Cox12, Sch14b]. pseudo-Aristotelian [Cox12]. pseudo-confirmation [Sch14b]. pseudoscience [Han17, Mic24]. psy [Aro18]. psy-ops [Aro18]. Psychiatry [Ano95c, Rob22]. psychical [Noa08]. psychoanalysis [Tuc95]. Psychoanalytic [Bor92]. psychological [Fek12, Roc88]. psychologism [Edg08, Met91, Pre08b, tH93]. Psychology [Man72, Dar03, Fri08, Fri18, Guy08, Hat07, Hoc16, Kra18, Kus97, Lis18, Mak03, Pre08b, Ree86, Ric92, Rup76, Sch08c, Sha94]. psychometrics [MBC17]. psychophysical [Isa17]. Ptolemaic [CD15, Fra93]. Ptolemy [Bar09, CD15, Dra78, Fek12, Kut22, Web22]. Public [Chu22, Ger20, Gol22, Joh22, Lea22, Lin22a, Mel91, Myk22, Rot22, Sol22, Ups22, Val22, WV22, Hau22, Kuk22, LF25, MHW22, Sor16, Tur22b, ZZW20]. publication [Ano05f, Ano06c, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13a, Ano13b, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17e, Ano17a, Ano17b, Ano17c, Ano17d, Gui04b, Moy91b]. publications [Ano03f]. publicity [HLM13, Pic03, Sno04b]. Publish [Ven22]. published [Neu71]. Publishers [Ere91]. Publishing [Car73, Fyf02, Zam18, Top00]. pudding [Saa08]. pulse [Bra08]. Pure [Dou14, Uss15, Hen84, HS23b, Kra72]. purifying [BDB21]. purloined [Lew11]. purpose [Ker12, Str21]. purposefully [RM16]. pursuing [Sar12]. Pursuit [Fle22, Wil24b, Bak22, Bar22a, Fis23, HK22, Hay04, Lic21, McK95, Rui23, SD23, SW12]. pursuit-worthiness [Lic21, Rui23]. Pursuit-worthy [Wil24b]. pursuits [DK22]. Pursuitworthiness [Fur22, Sto23, HC23, Kuk01, Sha22]. pursuitworthy [San22]. Pushing [Gre24a, Kau08, Wes23, Slo24, Wes23]. put [Maa13]. Putting [Bid11, Hen22, Koc15b, Rus23, Ayl19, Pic84]. putting-up [Ayl19]. puzzle [All21]. puzzles [Zwi11]. Pyle [Cha02].

quadratic [Yap11]. quadrature [Qui18]. quaestio [Man92]. Qualitative [Cha15a, AB23, HC23, IRDV24]. qualitatum [Cla68, Gra72]. qualities [Cla68, Gra72, Per91]. quality [Leu15]. Quantification [Lev01, Por01, RC17]. Quantisation [Lin22b]. Quantitative [AB23, Joh11, SLP23]. quantities [Isa17, Par20a, She91]. quantity [Dic25]. Quantization [Car22, KC94]. Quantum [Bog21, Cal21, Cal22, Dun22, Ear20, Heu14, Rie24, Sch21b, TC25, Tas25, Wal22c, Bar18, Blu21a, BJ22, BM14, But93, Cal24, DJ14, EZ13, FJ21, For22, FL22, Gli16, HS22, Kon22, Len14, Let21, Lyr14, Men24, Mit22, Mul24, Set08, SHB94, dRM23, vS20, Mee21]. Quasi [Dic06, Gla01b, Hol21b]. quasi-empiricism [Gla01b]. quasi-erotetic [Hol21b]. Quasi-truth [Dic06]. quaternions [Koe95]. quelconque [Le 58]. queries [Ans04]. quest [GT05]. Question [Bar22a, Coh08b, Dou22, Zam08]. Questionnaire [Bra01]. Questions [Bra81]. Quetelet [CM00]. qui [Le 58]. quibus [KR08]. Quincunx [KT21]. quine [Kri82, DH69, KS19, Ver17, Bro71]. Qutb [GS22].

R [Ano91b, Ano01b, Elk74, Hin12, Mye89, Rog72, Sie04, Wit24, Des49]. R-D-et [Des49]. R. [McL24, Ros22, Sla02]. Race [Slo24, Wes23, Lin22a, LA22, Ste22c, Zam08, Mal13]. Rached [Edi22, AR22, Rob22]. racialization [Bla22]. radiation [BJ22, CL05b, CL05a, HS23c, Hud89]. Radical [Pal09a, Di 11, Fee06, Fre15]. radically [WHK14]. Radicals [Gay76]. radium [Cor21]. rainbow [Tie88]. rainfall [Leh18]. ramifications [dCG17]. Ramsey [ML18, CF06, Cru05, Pap10, Run94, Rus23]. randomizers [Nor22]. randomness [Bla24, McA03]. range [Kai19, Kos93]. Rankine [Moy77]. Rapprochement [Ano91h]. rate [Dic13, Men14, Sch17]. rates [Dun12]. Ratio [RT95, Mit17]. ratio-of-unit [Mit17]. rational [Bar18, Car95, Fum20, Krü80]. rationale [Bha15]. Rationalism [Hei16, Pre16, Sha20]. rationalist [Mou81]. rationalists [Nel94]. Rationality [Maa99, SW12, Cle83, Cur89, Díe07, Dou20, Eve18, Gie89, Hoe80, Kit87, Len22, Mow85, Nun84, PG20, Pet16, Sar82, Sar97a, Sar97b, Sie83, Sim01, SG85]. Raum [Car21, Sto10]. Rawlsian [DZ14]. ray [Sch09a, Cur92]. Raymond [Sha90a]. rays [FWW09, Hon87a, Row08a]. RCA [Ras96]. RCTs [Maz22]. Re [Axt93, Ben18, Woo15, Lau88, LL16, Mäk97, Mit17, Pal87a, Cam22]. Re-assessing [Axt93]. re-emergence [LL16]. re-evaluated [Pal87a]. re-examination [Mäk97, Mit17]. re-narrating [Ben18]. Re-orienting [Woo15]. Re-situating [Ben18]. re-visited [Lau88]. reach [Goo86, Sch05]. reaction [Mor93]. Reactionary [Del17]. reactions [Gar76]. readers [Pal87b, Rap11b, Wil09b]. Reading [Fin90, Ste22c, Bra15, Sch13b, Top00]. real [Díe97a, Fee06, Man12a, Pal08b, Rot96]. realignment [Dic25]. realisability [CF06]. Realism [Ano91k, Hei14, Heu14, Isa19, Isa20, McL06b, Pap10, Ros18, Row23, TA23, Adl22, All21, Ara23, Boo23, Bro90b, Bue19, Bus08, CD15, Cru05, Dic13, Dil90, EZ13, GL24, Gli16, Har10, Hen23, Jac95, Jar79, Kan96, Ken23, Lad98, Lan14, Lau23, Lep79, Lyo16, Man22, McL05, Mee92, Mer21, Miz12, bAM11, Pap89, Par16, Pic22, PS24, Row11b, Saa08, San08, San09b, Sie98, TC21, Tul16, Tur21, Tur04, Wra15, Elg22, Car05, Boo23, Ken23]. realist [Dun22, Mar99, Saa05, Wes72]. realistic [Bro90a, Tro94]. realists [Del19, Het96]. realities [Goc08, Noa08, Sar12]. Reality [Man18, Bog21, Fag24, Kat20, Kem07, Kim00, Kon22, Löw08, Rad93, Sib08, tH02]. Realizability [Hun18, Man18]. realization [Aiz18, Cru05, Ful18, Rad93]. realizations [Joh11]. realizers [Dra18]. really [Bro16b, Fra93, Gra11, Sam17, Wel02]. reappraisal [Dra76]. Reappraising [BK16, Pan96]. rearrangements [Fis06]. Reason [Van04a, Fri12, McN15, Mor12, Nob23, O’N98, Pal09a, Ste03, SM24, Zha22, Hac12]. Reasonable [Ach23]. Reasoning [Dar76, Bia20, Boa07, Bol14, Bou16, Bue12, Dou20, Elw12, Fis18, Gow97, KL14, Kus10, McC21, O’L21, She93, She06, Smi88, Spe04, Wan12, dJ97]. Reasons [Hei16, Fle22, Pee17]. Reassessing [Hei24, Kra24]. reassessment [Ben74, Gla93, Mot02, Sha89]. Received [Ano01a, Ano11a, Ano12a, Ano14a, Ano70b, Ano70c, Ano70d, Ano71c, Ano71d, Ano71e, Ano72d, Ano72a, Ano72b, Ano72c, Ano73b, Ano73c, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano77b, Ano78a, Ano79d, Ano79b, Ano79c, Ano80b, Ano81a, Ano82b, Ano83a, Ano84b, Ano85c, Ano85d, Ano86b, Ano86c, Ano87a, Ano87b, Ano88a, Ano89d, Ano89e, Ano90b, Ano91c, Ano91d, Ano92c, Ano92d, Ano93d, Ano93e, Ano94c, Ano96a, Ano97a, Ano97b, Ano04b, Ano07a, Ano09a, Con06b, Spr10b]. Reception [FWW09, Cam12, Fra90, Moy91b, Pan99, Whi90]. receptions [Coo05]. reciprocity [Yap11]. reckon [Wer18]. Reclaiming [Bol15]. recommendations [PM22]. reconcile [Fed23]. Reconciling [Boo23, O’M16]. reconfiguration [dC12]. reconnecting [UG20]. reconsideration [Rib85]. reconsiderations [Zam16]. reconsidered [Bro05, Fay88, Fee16, GG94, Wil89]. Reconsidering [Fri12, Gel10, Saa05]. reconstitution [Dea87]. reconstructed [Mal93]. Reconstructing [Mot02, Wan22]. Reconstruction [Wil09b, Gay76, Hei80, Krü80, Red80a, Wan02]. reconstructions [Pas20]. recording [dC93]. records [Ste05]. rectilinear [Sha74]. Red [Ste22b, EG80]. Redefining [Bra24a, Pie14]. rediscover [Koc15b]. Rediscovering [Obe16, Ueb91b, Zol93]. rediscovery [LW13b]. Redondi [Sha90a]. redshift [FR24]. reduced [Le 58, Tuc20]. Reduction [Sha18, Cat14, Mau77, Plu78, Sar15a, Wag24, FR23]. Reductionism [Pal23, Fut02, Har08]. reductionist [Waj21]. reductive [Cov97]. reductivist [Wel02]. réduits [Le 58]. redux [Sha18, Ste24b]. reef [Muk22]. refer [Man72]. Referate [KMS73, Whi74]. Reference [HHO09, Lep79, Pes10, Bou20, Koc15b, Lan14, McL05, McL06b, Sce06]. referent [Lew11]. Referential [Kle77]. referred [CS07]. Refinement [dB09]. Refinements [Moy91b]. reflected [Lau87]. reflection [Nas16, SG85]. Reflections [Jar13, LW13a, Sha03, Dar23, DW23, Kir73, Koc06, Mir05, Ris91, Wig67, Wig70, Wig78, Woo24]. reflex [Ekb24]. reflexive [Dör94]. reform [Mor91, Van04b]. reformation [De 96a]. refraction [Buc72, Car16]. Reframing [CL22]. refusal [Sch07b]. refusing [Ram92]. refutation [Fal98, Fra08, Lay78, Ueb10]. Regeneration [Die24, Hus24, Die24, Hus24, MM24a]. regions [Cat12]. Régis [Ott08, San16b]. registers [Weh19a]. regress [Col94, Col02b, Col09, Fee16, Fra94a, GG02, Jar76]. regression [KL21]. regrets [Len81]. regularities [KH20]. Regulation [Atk94, Cre18, Joh21a, KD19]. Regulative [Spa23, Gio22, Kle08, McN15]. regulatory [Coe22b, LT20, PMG21, Sch74, Sie98, SL21]. reguntur [Dun12]. Rehabilitating [McN15, Sch07b]. rehearsal [Par12]. Reichenbach [AP06, DBM22, Esp79, Fra20, Pad15, Ric05]. Reid [BZ86, Cal10, Gra05, Red02, Slo03]. reification [LB94]. Reine [Kra72]. reinterpretation [Kav24, Mac73]. Reinventing [Kra22b]. reject [Stu08, Zyn96]. rejected [Bon94, Liv09]. rejection [Col18, Slo77, Wil01]. rejoinder [Wor05]. related [MCGP17]. relation [Dem06, Fey84, Hen85, Los83, Sch16, vS20]. Relational [Heu14, Rie24, dRM23, EZ13]. relationalism [GG21]. Relations [JM11, Bro11b, Fag16, Krü80, Let22, Lin22a, Mit17, Rie14, VFHS23]. relationship [Fav91, Fay88, Sha97, Slo19, Yea82]. relationships [BS13, Dio18, Sta18]. relative [HK21, Ste22b, Ueb12]. relativism [Bla13, Can98, Cle83, Cox75, GK88, Hei16, Hel16, Mal93, Rie24, San11, San12, San14, Tos06, dRM23, Kus16]. relativisms [Gat16]. relativist [San13, Sei13a, Sei13b]. Relativity [Mar21, Miz21, Tur21, Cur92, Dar21b, Eve18, Fay24b, Gra71, Hen92, Jan73, Kat92, Ker87, Per91, Rie24, Ryc92, War89, EG80, FK90, Pit22, War92, War93]. relativized [Fer12]. relevance [Düp15, Kri18, Sam10, SFF23, vE19]. relevant [Lar24]. reliabilism [Bha17]. reliability [MM24b]. reliable [Maz22]. Relics [Bud13]. Religion [Cun01b, Dea01a, Ram08, Bla96, Bro07, Cun00, FP99, Kon24a, Mar16, Sno04a, Ste22c, Ver07a, vdW04, Cun01a]. remains [Tur22a]. Remarks [Fav91, Fin81, Bro09, Fer12]. remember [Cur18]. renaissance [Fin90, Dea03, Mac00]. Rendel [Zap24]. René [Spa91, Des49]. renewed [Zam08]. Renormalization [CM22, KF23, Fra21c]. Renouvier [Sch07c, Sch20a]. Reorienting [Hen23]. Repertoires [AL16, Sam17]. replacement [HHO09]. replica [Pol21]. replicability [MM24b]. Replication [Fee16]. replications [Ven22]. Replies [Mas24, Ris91]. Reply [AR22, Ach90, BR23, Bat23, Blo82c, Blo84, Bro21, Dea01b, Gol22, Mac84, MM24a, Miz21, Rhe25, Ric22, Row23, She22, Sol22, Val22, Bai92, Bir08, Blo82b, Bol14, Cha02, Cun01a, Eig19, Fag24, Fee06, Gen20, HHO09, Kid11b, Kuu09, Lan18, Lug92, McA05, New10, O’G78, Pyl02, Ray92, RGD15, Saa08, San13, SOC19, TGD05, Wei10c]. report [Aiz18, FB20, Hei24]. Reporting [HK21, Hur17]. reports [LA22]. represent [Gra11]. Representation [Díe97b, Hat88, Kra22a, LW90, Bol16, Dra18, Gel11b, Gel11a, Hei70, JSZ23, Knu11, Spr10b, Suá15, Tan21, Yos23, Kra22a, Rhe94b, Rup22, She22]. Representational [Pol21, Weh19a, Boe22, Mic93, Vor14]. representationalist [Ken12]. representations [Mat22b, Por11]. represented [Has97, Pal08b]. Representing [Vor11, Kes07, Knu11]. Reprint [May71, Bri73, Ege70b]. reproducibility [Wei21a]. reproduction [Bru20, Con00]. reproductive [Bru20, ND22]. Republic [Gre96]. Republican [Wu14]. reputation [Bar22c]. require [Lus21]. required [Fal17]. requirements [Stu91a]. requisites [Qui18]. research [AL16, Ant25, Bid14, Big08, Cho22, Clu12b, Coe22b, Col07b, Dav21b, Díe97a, Dor79, DCJ^{+}21, Ell12, Fey84, FBW22, Fra21b, Fre80a, Fri18, Gal21, Gla93, Gla01b, Hen85, HS23b, HK21, KM22, Koc13, KR22, Lac21, Loh21, MHW22, Meu19, ND22, PMG21, Pen22c, POW16, Qui72, Red80a, Res96, RE19, San22, Sto23, SL21, Van86, WHMO22, WS21, Wil09a, Wil24b, WL22, WHK14, Wra06, YZ19, Zah19, Pie97]. researches [Duc10, GG88, Goo87]. resistance [Bar21c]. Resisting [Psi94, Sto25]. resituation [Meu19]. resolution [Miy17, Ohn22, Row11a]. resolve [LGC21]. Resolving [Kes07]. resonance [Liu22, Sum13]. resources [Gre24b]. respective [Bal20]. respiratory [Bra24a]. Response [Ano01b, Col08, Fri12, Gin09, Nik19, Pal09b, San14, Sce05, Weh19b, Ano09f, Ari01, Buc82, Cah97, Cha11, Che90, Cur92, GL24, GP10, Goo80, Hen14a, Iva11, Jac94, Kem07, Row15, San09a, Sar86, Sce06, Sch08b, Smi84, Ste22b]. responses [Del17, Nor21a, Ueb05b]. responsibility [Fri08]. Restitution [Clu12a]. restoration [Alm21]. result [Dor74, FH88, Wil09c]. results [Kin15b, Sol08a]. retail [Dic13]. retention [O’G78]. Rethinking [Die24, Had04, Mét13, Slo19, San16b, Loi16]. retirement [Hau21]. Retrieving [Pin97]. retrograde [Car15]. retrospective [Tos03]. reunited [BO24]. reveal [Col90]. revealed [Hol22, Wil09b]. Revealing [Sol08b]. reversion [Pen22b]. Review [Ach92, Ait70, Ano73a, Ano91b, Ara23, Bak16, BR23, Bat23, Bau71, Bia94, Blo75, Bri73, Bro71, Bro21, Bus71, Cam22, Car73, Chu22, Cle23, Cra93, Cul04, Cul06, Dar23, Eas72, Eat23, Ege70a, Ege70b, Ege71, Elg22, Elk74, Ere91, Fin73, FR23, Ful94, Ful93, Gab73, Gal71, Gia94, Gie91b, Gol23, Gol22, Gra72, Gro90, Gui04a, Hac89, Hau22, Hav21, Hei75, Hen23, Heu14, Hop94, Jan73, Jas04, Jhu22, Joh22, Kam94, Kir73, Kni71, Kra22a, Kuk22, Lea22, Lin22a, Liv93, Luc71, ML18, May71, McE76, McH21, McR72, Mel04, Mil23, Mis94, Miy23, Mor90a, Mor90b, Mul94, Mun74, Mye89, Myk22, Neu71, Nic72, Ode22, Osl71, Osl72, Pal23, PR90, Pes13, Pin22, Rhe94a, Rhe94b, Ric22, Rob22, Rog72, Rot22, Row23, Rup22]. Review [Sch91, Sha90a, Sha90b, Sie04, Sno04a, Sol22, Spa91, Ste04, Swo74, Tur22b, TA23, Ups22, Val22, Van18, Vea72, Whi74, Wie21, WV22, Zam18, Zol93, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano00a, Atk94, Bar20a, Bel15, Cer19, Col08, Gin09, Gur19, Nik19, San22, Sha76, Wra19]. Reviewing [Yeo89]. Revisited [Cun01b, Ant03, Blo82a, Cha15b, CE15, CEW16, Cun00, De 16, FR24, Gre24b, Har08, Kui14, Mac10, Par14, Ueb09, Wor14]. Revisiting [PY06, Cob11]. Revitalised [Row23, TA23, Ara23, Hen23]. Revitalizing [TA23]. Reviving [McC20]. revolt [Sha20]. Revolution [Tas25, BG88, Cha12, Gre92b, Ill81, KH20, Kuu17, Mit72a, Mol82, Pan96, Pie14, Pou00, Ree86, Cha15b, Gla86, Gla02, Kle15, Kus15, Mor05, Zin11]. revolutionary [Che10, Hei80, Sie98]. Revolutions [Ill81, Kuh70, Cor93, Gon96, Kuh62, Kuh96, KH12, Sch88c, HH95, Kva14, Lar03, Rot13]. Reworking [Sib95]. Rewriting [Tas25]. Rheinberger [MW25, Rhe25, SD25]. rhetoric [Dou95]. Rhetorical [Ano91k, Cre12, O’F10a, O’F10b]. rhythms [BA10]. Riccioli [Ver07b]. Rice [Elg22, Jhu22, Ode22, Pin22]. rich [Wel05]. Richard [Eas72, Hei75, Mel04, Sno04a, Cal10, DHL21, HM88]. riches [Ter05]. Riegl [Kwa12]. Right [Wis18, Car93, Cun88, DBM22, Ueb10]. Rigor [She87, Ekb24]. Rip [Col18]. Rise [Hei16, Fra79a, Fra84, Mar09, Mat19, Nea24, Pre16]. Risk [LT20, Cur19b]. Rites [Win99]. Ritter [FWW09]. rival [Tre87, VS12]. rivalry [Con00]. River [Boe22]. RNA [Rhe92b]. road [Fra05, Jar76, Kra24]. Robert [Ano91b, BR23, Cle23, Dar23, Eat23, Ege70b, FR23, Gol23, Mil23, Mor91, Pal23, Sor16, Ans02, Cah97, Cha02, Cha16, Gol97, KH17, Sar86, Woo24]. Robino [BR23, Cle23, Eat23, Gol23]. robotics [FH23, Tam23]. robots [Bec23, Gei23]. Robustness [O’L21, Bas17, KH25, Llo15, Vez16]. rocks [Ros18]. rods [Gio14, Hac91]. Roger [Clu12a, KH17]. role [BG88, BDB21, Can98, Cha86, Chi08, Daw21, Ega10, Eve18, Gel10, GW15, Gio24a, Gun18, Hea24, HD19, HG21, How98, Jan13b, Kal75, Lak74, Lau82b, Les78, Lev11, Lis18, Llo12, Mer82, Ord16, Pac18, PM22, PvF14, Pin16, Pru89, Reu22, Rib07, RZ21, RH95, She88, Sim70, Str06, WHMO22, Wei10a, Win81]. roles [Ell13, KR22, Lov15, Sch05]. Roman [Ram12a, Cuo00, DFT12]. romance [Man80b]. romantic [Lau87, Lea98, Liu22, Car13, PB96, Zem06]. Romanticism [Ham00, O’C05]. Rome [Gia94, De 96a, De 02]. Ronald [Hei75, Ros17]. room [Pie97, Tam23]. root [RK06]. roots [Bar14, Hei24, Kai92, PM23, Why21]. Rosenblueth [Luc71]. Rosenthal [Sha90a]. rotating [CN83]. rotation [GG21, Ker87]. round [Oos12]. route [Wes79]. routes [Chi06]. Routledge [Ach92, Mis94, Nic72]. Rowan [Hen84]. Rowbottom [Ara23, Hen23, Miy23, Bir08, TA23]. Roy [Rud74]. Royal [Hol21a, Hun03, dC02]. Ructions [Mor22]. Rudolf [Psi00]. ruffians [Sny07]. rule [BMBM20, Cuo00, KVJ22, McG70]. rules [CW68, Col09, Eas72, McO96, Sla02, Spe04]. Ruse [Ere91, Mun74]. Russell [Mic93, Mil08, Pul01]. Ruth [Nik19]. Rutherford [Ach23]. ruts [Ano91h].

S [Ano91b, Haw96, Hei75, HH95, HH06, Liv09, Mor90b, Möß11a, Möß11b, Rog72, Ara13, Bar22c, Mor21, Sha17]. Sadik [Gal71]. safety [Col21, Gré19]. said [Mau90]. Saint [Goo04]. Sakurai [Rui23]. Samos [Wal75]. Samuel [Bro16b]. Samuelson [Han21, Mir89]. San [CSD21]. Sankey [HHO09, Sei13b]. satellite [Llo12]. satis [Jan13b]. save [RS21a]. Saved [Leh20, Dri24]. Saving [Feh07, Kem05, Kri18, Pal87b]. Saxony [Dra75]. say [CS07]. Scale [Tyb22, Hen21, Hol21a, PM23, Pin09]. Scales [Mil23, CL18, Dar23, Díe97a, Pac18, Riv21, Pal23]. Scaling [Wat23, AB23, KG13]. Scattering [vL24, Blu21a]. sceptic [San13, Sei13a, Sei13b]. Scepticism [Bla13, San12, Net91]. sceptics [Jar79, Slo77]. Scerri [Bar05, Sce05, Sch08b]. Schelling [Sch18b, Ste03]. Schemas [Bak20]. schematism [Lea91, Win81]. Schlick [Obe15]. Schneer [Bus71, Kni71]. Schnorr [Bla24]. Scholars [Kim12]. scholarship [DN00, Mel04]. scholastic [Ott08, San16b]. Scholium [Ryn95a, Ryn95b]. school [Gar89, Mel89, Sch88a, Gio24b]. Schopenhauer [Lau87]. Schriften [KR08]. Schrödinger [BM89, Ber85, Wes79]. Science [Aba08b, Ano78b, Ano97g, Ano97f, Ano01a, Ano01b, Bia88, Bid11, Bor15a, Car13, Cha20a, Cha20b, Chu22, Coh97, Col16, Cro92, Cul04, Cun00, Cun01b, Gam21b, Gla83, Han17, Hol21a, Jas04, Joh19, Joh21a, Joh98, Kit01, Lev01, Lin22a, Mar21, May00, Nas68, Nas83, Nik19, Pre98, Rad93, Ree12, Rot22, Row23, Sch93, Sib08, Slo19, Sol22, Spa91, Ste22c, Stu01, Suá12b, Tha93, Tos06, Tos07, TA23, VK19, Zol93, Aba11, Aiz10, AG09, All96b, And16, And22, Ano81c, Ano90b, Ano90d, Ano90f, Ano91c, Ano91d, Ano92c, Ano92d, Ano93d, Ano93e, Ano94c, Ano96a, Ano97a, Ano97b, Ano02b, Ano03a, Ano03f, Ano09f, Anz19, AH15, Ara23, Asp13, Atk94, Bak21, Bal92, Bar18, Baz88, BA10, Bel20, BCC87, BV05, Ben24, Bid13, Bid14]. science [Bla96, Boa11, BM21, Bor20, Bor25, Bow93, Bre04b, Bri73, Bro07, Bro90a, Bro16c, BH19, Bru20, Buc69a, Buc69b, Buc89, BG23, Buc09, Can98, Car05, Car21, Car95, Cha11, Cha15a, Cha16, Chi01, Chi06, Chi11, Cla95, Cob11, Coh18, Col03, CE19, CM00, Coo23, Cra17, Cra19, Cun88, Cun01a, Cur19a, Cur19b, DK18, DHS14, Dea87, Dea01a, Díe07, Don21, Dou14, Dre20, Dun13, Dur13, Ede09, Ell21, Fed23, Feh82, FB20, Fin90, Fis18, Fla83, For89, Fra20, Fra21a, Fra21b, Fre80a, Ful89, Ful94, Gam21a, Gar00, Gau98, Gie91b, GID21, Gio24a, Gio24b, Gol75, Gon96, Gra96, Gre92a, Hac91, Hal12, HC23, Han07, Har19a, Hau22, Hea24, Hee98, Hei14, Hen23, Hen84, HG05, Hoo13, Hüb80, Jac97, JFS97, Jar03b, JC20, JC21, Joh11]. science [Jon13, Joy97, KN89, Kap16, Kes07, Kin87, Kin15a, Kin15b, Kit93, Kle04, Kle16, Kni05, Koc11a, Koc15a, Koc24, Kov21, Kra11, Kra72, Kra73, Kuu11, Kwa12, LM12, Lak74, Lat90, Lau82b, LL16, LGC21, Lee13, Lee19, Les78, Leu15, LC24b, Liu22, Liv93, Llo97, LS02, Llo15, Loi16, Lon05, Los77, Los78, Los83, Low93, LF25, Lug78, LT20, Lus21, Lux67, LJ11, Maa13, MJ23, Mäk97, Mar04, Mas13, Mat98, Mau77, May04, McA96, ME15, McK95, McN22, McR72, Mir89, Mir04, Mir05, Mor00, MW17, Mos07, Mye89, Nas16, Net93, New10, NI03, Nor12, O’C05, O’C19, O’L20, O’N03, Obe16, Ort08, Pal05, Par92, Pas20, Ped20, Pet11, PS10, PS24, Pin14, Pin16, Pla20, Pol15, Pot12, Pot15, Pri18, Pyl02, Qui21, Rad97, Ram08, Rap11a, Rei96]. science [Rei97, Reu22, Ric05, Ric09b, Rie14, Rit12a, Roc02, Rog72, RH11a, Ros70, Rou91, Row11a, Row13, Sak22, San08, Sar12, Sch16, Sha97, Sie83, Sie89, Sie04, SFF23, Sig90, Sol08a, Spe17, Spr10a, Sta11, SGO17, Sto23, Sto09b, Str06, Stu89, Tam15, Top00, Tos03, Tri08, Tuc93, Tur04, Tur22b, Ueb05a, Ueb05b, Ueb15, Van95, Ver07a, WHMO22, War21, Wat23, Weh19a, WB20, Wei10a, Wel05, Wil09c, Woo71, Woo94, Wra05, Wra07, Wra18, Zam03, Zan88, Zuc19, dL09, dO12, dR98, vE19, vS20, vdB22, Psi00, Ano80f, Ano89a, Ano91m, Buc88, Cab96, Cah86, Gol23, Ara23, Tur22b, Bro21, Chu22, Elg22, Gol22, Joh22, Kra22a, Lea22, McH21, Ode22, Pin22, Ric22, Rup22, She22, WV22, Gro90, Jhu22, Luc71, Miy23, Osl71, Van18]. science [Wie21, Bau71, Bro21, Cra93, Hav21, McH21, Wie21, Sol95]. Sciences [Cra93, Hau22, Lin22a, Rap11b, Rot22, Sol22, Ano92e, CL22, Coh08a, Col07a, DS18, DK18, Düp15, Fra94b, Fri02, GN16, Gau16, HK16, JSZ23, JW13, JW21, Kid17, MN18, Mak08, Mer92, Nas16, Nii14b, Ols16, PP21, Ram12b, Ris93, Ris98, Rom16, San16a, Stu12, SM24, SD25, Tri08, Tur13, Tur22b, Van04a, YZ19, dC13, vdB18, Ano80c, Ano97g, Cro92]. Scientific [All21, Ara03, Ara23, BR19, Bir08, Con06b, CT16, Del16, Del23, Ell92, Far97, Gie89, Gie06, Gie09, Gin88, Gra13, Hen23, HH95, Kan96, Kap16, Kir73, Kuh70, Kus15, Kva14, Lar03, Loh21, LW13a, McL05, Mei21, Mel04, Nii14a, Nor00, Pad15, PMG21, PS24, Rot13, Row15, Row23, San08, San09b, Sar89, Stu17, Top00, Tos06, Tos07, TA23, Wil10, Wra06, Wra19, Ahn20, AL16, Ano86h, Ant25, Bai87, Bak22, Bar22c, Bau98, Bec84, Ben22, Ben74, Bie20, Blo75, Bol14, Bor20, Bor25, Bre04b, Bro09, Bue00, Cam21, CD15, CL04, Cat95, Cat12, Cha12, Chi15, Cob11, Coc14, Cor93, Cur89, Dar76, Dar02, DD13, De 14, Del19, DvdB22, DK22, Dor79, Dra78, Duc10, Ell12, Eve18, Föl12, Fra79a, Fri98, FA13, Ful24, Fun86, Fyf02]. scientific [Gay77, Gay78, Gel11b, Gel11a, Ger20, GW73, Gla89b, Gök90, Gow97, Gre92b, Gut73, Han05, Har10, Har19b, Har96, Har93, HK22, Hei75, Hei24, HD19, Hes74, Hes75, Hoc87, Hop94, HZ19, Hu21, Hum19, IRDV24, Jar13, Kal75, KH20, Ken23, Kid16a, Kit87, Kle77, Koc13, Koc15b, Kra18, Krü80, Kuh62, Kuh96, KH12, Lac15, Lan14, Lan21, LL16, Lep79, Lep81, Lic21, Lov15, Luc20, LW13b, Lus21, LW90, Mac75, MN20, Mar04, Mat22b, Met91, Mic71a, Mir04, Miz12, Mol82, Mor91, Mow85, bAM11, Nel94, Ner84, Par16, PR90, Pic03, Pla20, Qui72, Rap80, Red80a, Ree86, RE19, Rhe94b, Rio15, Rob70, Rol16, Rot96, Row08a, Row11b, RS21b, Rud74, Rup76, San22, Sch73, Sch05, Sch07b, Sch88c, Sch93, Sha97, SFF23, Smi88, Sol95, Sor16, SG85]. scientific [Stu16, Tab14, TC21, Tan21, Tau09, Tur21, Ueb05b, Ueb09, WG13, Wig67, Wig70, Wig78, Wil09a, Wit24, Woo15, Wri15, Yea82, Zan88, Mor05, Zin11]. scientific/intellectual [Rol16]. Scientism [Lin22a, Pee17]. Scientist [SGO17, BD02, dR22]. Scientists [BS13, Gun18, Sen01, Cla92, Col84, FH84, Goo86, Rot96, Sta11, Vas00]. scienza [Bal92, Gia94]. scissors [CF21b]. scoglio [Pal05]. scope [Fal18, Hee98, vdB18]. scores [Ser23]. Scott [Gab73]. Scottish [Can71]. Scripture [Jan15]. scrutinized [Wra15]. Scrutinizing [Kle14]. sculpture [Fin11]. Scylla [Sei13a]. sea [Fra84, FG22]. Sean [Kuk22, Myk22, Ups22]. search [Car16, Col04, Kuk01, Mar04]. seasonal [Rat19]. Second [Neu71, Bri73, Fra05, Tur16]. secondary [Per91, San16b]. Secrecy [De 96b, Dea03, Qui21]. secret [Han05, Pet11]. sections [Pou07]. security [Ama05, Lin22a]. seductive [Lip92]. seductive-nomological [Lip92]. see [BM23]. Seeing [Nor98, Löw08]. Seeking [Por11]. Seepage [O’L20]. segregation [Dar76]. Seidel [San13]. seismic [Aro18, tF18, Miy17]. selected [Bou21, GVD24]. selected-effect [Bou21]. Selection [Abe24, Bar22b, GKS24, Bou20, Bou21, BDB21, LW94, McL24, O’C19, Pra75, Ric97, Ros17, Rus71, Way21]. selections [DD69b]. selective [Oni17, Osl72, Yol70]. Self [BH19, Fri03b, dC93, Aub08, Lud16, Mak03, PB96, Rom16]. self-correct [Rom16]. Self-correction [BH19]. self-determination [Lud16]. Self-organised [Fri03b]. self-organization [Aub08]. Self-recording [dC93]. self-understanding [Mak03]. Selling [Wes23]. Selz [tH93]. semantic [Lor13, Row08a, San09b, Tho83]. semantical [AS16]. semantics [Nor98]. Semirealism [Cha98]. Semmelweis [Sch13c]. sensationalism [Bec71]. Sensations [Tri14]. Sense [Iva11, Atr85, BC00, Bha17, Blo08, BM23, Cok15, Fai12, Fri08, Gau81, Iva10, LF25, Rap11b, Red02, WB20, Whi71]. sense-idea [Whi71]. sense-making [BM23, LF25]. senses [Sib08]. sensible [Ris90, Ris91]. sensitivity [Hol21a]. sensors [Aro18, Cre18, tF18]. sensory [Isa20]. sensuous [Rob95]. Sent [Sen01]. sentences [CF06, Cru05, Pap10]. Sentiment [Jon01, Van04a]. sentinels [tF18]. separability [Fag24, How85]. separable [dRM23]. September [Ano95c, Bus71]. Serendipity [KN89]. serious [Muk22]. seriously [Col02a, Muk22, dOFGB22]. serological [vdBG90]. Service [Cam22, Myk22, Wis13]. ses [CL05c, Cul06, Des49]. Seselja [Kuu09]. set [BT09]. SETI [Kuk01]. sets [AG09, RM16, TGD05]. Setting [May00, May04]. Seventeenth [Man92, PR90, Cha15a, De 02, Dea87, Gab71, Jes89, Mal91, Sar89, Sha02, Slo77, Van16, Wil09b, Fun86]. Seventeenth-century [Man92, Cha15a, De 02, Gab71, Sar89, Sha02, Slo77, Van16, Wil09b]. seventeeth [Rem07]. several [Le 58]. Severe [WOL20a]. Sewall [Ros17]. sex [Con00]. Seymour [Meu19]. Seyyed [Bau71]. shadow [Col04, Sta13, Col08]. Shannon [Tho09]. Shao [Kat20]. shape [Pet11]. Shaped [Cha20a, Cha20b]. Shaping [Baz88, Pie97, Lin98, Gro90]. shared [Eig19, FB20]. shave [Bam96]. Shaw [Blo08]. Shea [Rup22, Cam22, Kra22a, She22]. Sheep [Ser23]. Sheridan [Sch91]. shift [EG80, Gio22, Saa05]. Shifting [Kle05, Ros17]. Ship [Bla22]. Shipwrecked [Ham00]. Shirazi [GS22]. shortcomings [Kus01]. Should [Del19, Gro20, Zyn96, Ayl19, Cor21, Jas04, Pla20]. Show [Buc09]. showcase [Maa13]. Shuchun [Cul06]. shut [Ayl19]. SI [Qui17]. Sichtbarmachung [Blo08]. sidelights [Ear20]. sidereo [KR08]. Sieroka [Pes13]. sight [Bak16, Smi15]. Sigmund [Nir23]. sign [Obe15]. sign-theory [Obe15]. signals [FL21]. significance [Gam21a, Gie77, Len07, McA96, bAM11]. Signs [Pat09, Fin90]. Sigrist [Spa91]. silence [Rot13]. Silent [Sam17]. silver [Why21, dS09]. Similarities [Bas92]. similarity [Bîg13, Gla89b]. similarity-based [Bîg13]. Simmel [Col02a, Hoo13]. Simon [Sen01, Sim01]. Simple [Räz24, Gla97]. Simples [dC12]. simplicity [Ano91k, GS22, Mar09, She88]. Simplicius [Har12]. simply [Lat90]. Simulating [Dou10, Sen01, Sim01, Wei10c]. simulation [Fal24, Lus16]. Simulations [TJT16, Bog20, GP10, KSZ24, Wis17]. sin [Gio14]. Singapore [Mel04]. single [Dor74, Oft22, Wan12]. singular [SF21]. singularities [dC02]. Sins [DK22]. Sir [Elk74, Pal71, Hay04, McC06]. Siris [Mok71]. Situated [Bar21a, McH21, Ken23]. situating [Ben18]. situation [Leu15]. Sivin [Cul04]. Six [Ueb05b]. sixteenth [DD69b, Ege70a, Rem07]. sixteenth-century [DD69b, Ege70a]. size [AB23]. skepticism [GG02, Vog10]. sketch [Ueb09]. skills [Har96]. Sklar [Mul94]. slaying [Mul14]. small [PP23]. smallest [Le 58]. Smeaton [Mor21]. Smith [Bak16, Mor90a, Blo84]. smoke [Kan96]. smoking [AD12]. Snippets [Kni05]. Social [Ano80c, Ano84a, Boh00, BG23, Cer19, Eva07, Kus02a, Leu15, Pre06, Res96, Rot22, Ano91i, Anz19, BO24, Bid14, BD02, BH19, Byc12, CM00, Cur19a, Han05, Kha24, Kus01, Lin22a, Lis18, Mir04, Mir05, Nor98, Pin97, Rad22, Rie14, Ris98, Rom16, Sch06, Tur13, Tur22b, Ueb05b, VK19, Vas00, Yee24, YZ19, dC18, Hau22, Lin22a, Rot22, Sol22, Tur22b]. Socially [Gla02, Nel94]. societal [Kus01]. Société [Spa91, Sig90]. Societies [Tha93]. Society [Ano80c, Ano84c, HLM13, dC02, BO24, Bow93, Ano87d]. sociocognitive [Gla89a]. Sociological [Gol75, BB23, Sch85, Yea82, Zha22]. sociology [Blo73, Blo82a, Blo07, Col02b, Fri98, Gar00, Gel80, Had04, KM14, Kva14, Man80b, Rie14, Smi84, Weh19a, Tos06, Tos07]. sociophilosophy [Kus97]. sociozoological [Hol21a]. Soft [Fri24b, Kra11]. software [ES06]. soil [SFF23]. Solovey [Hau22, Lin22a, Rot22, Tur22b]. solution [Hol21b, Mar99, Wag22]. solutions [PSS18]. solve [Ose22, Reu22]. solving [Fin81, Lop84, Sar81, Sie83, Van86]. Some [AM20, Buc88, Buc89, Cun88, Fer12, Gar76, Gro87, May14, Qui74, Red80b, Sim70, Spa90, Tho09, Woo24, Yea82, Bak23, Dic13, Gay76, Gri91, Hat88, MCGP17, Pin14, Pla20, Rog72, Ryc92, Sol08b, Ver82, Woo71, Woo94, Sie89]. Sommerfeld [Set08]. son [CS78]. Sons [Van18]. sont [Le 58]. Søren [Kid11a]. soul [Esc08, Fal98, Fek12, Hat07, Kra18, Web07]. sounds [Ste05]. source [KN89, Maz22, Nas16, Nee14, Wei07]. sources [Bri73, Dra86, Har02, Sim01, Sto10, Top00]. South [Coe22b, Van95]. Southern [Pin09]. Soviet [BK17]. Space [Jan00, Jan15, Ale95, Dar07, Ede09, Fay24b, For22, Gio24a, Grü78, Jac20, KC94, Mau90, Mit93, Ryn95a, Ryn95b, Sam77, Sch23, Sha03, Spa90, Win82, Car21]. space-time [KC94, Mau90]. spaces [Eva07, Szn16]. Spacetime [Kos88, Acu25, Che22, Che23, Nor21b]. spacetimes [CF21b]. spade [Tur13]. Spanish [Pet11]. spatial [Cox75, Ful93, Gao21, Hat90]. spatiotemporal [Let22, Pac18]. speaking [Joh19]. special [AL11, BB23, CL18, Cur16, Dar21b, MW17, Rit12b]. specialties [Hoc87]. specialty [Col21]. Species [McO96, Bal20, Dou22, Mag96, ND21, ND21, Rus71]. Specific [Isa19]. Specificity [vdBG90]. Spectra [Ano73a, McG69]. spectroscopes [Dör94]. spectroscopy [Ano73a, McG69]. spectrum [Koi21, Top90]. Speculation [Gow73, Swa24, Van16]. speculations [McN22]. speculative [Dor70, Mas11]. Spell [RH11b]. Spencer [Smi83]. sphere [Lus21, Tay09]. spiderweb [SFT23]. spikes [Wit24]. spin [CS21, Koc15b]. Spinoza [Bar21c, BT09].

Spinozism [Sch13b]. spirit [BJ22, LJ11, Pet15, Spe17]. spiritual [Mou20]. splendour [dR98]. Splendours [JFS97]. Splice [Rhe25, MW25, SD25]. Split [MW25, Rhe25, SD25]. spreading [Fra84]. Springer [Ait70, Neu71, Zam18]. Springer-Verlag [Ait70, Neu71]. Spuntar [Pal05]. square [Dor74, Fal17]. SSK [Wis13]. St [Cun00, Cun01b]. St. [Tei06]. stability [Ano91k, KL21, Man22]. Stadt [KMS73, Whi74]. stage [Bro09, De 18, Gal21, Gie09, Gla20]. stages [Rat19]. stake [Dhe23, Tur21]. Stance [Cha04, Mul24, Bar22a]. stand [Rot96]. Standard [Ode22, JW22, Sch20b, BCK^{+}22]. standardisation [Ker07]. Standards [LT20, Cho22, Con20]. standpoint [CK22]. standpoints [Rol16]. Stanley [Buo24]. star [Kut22]. Stark [DJ14]. starlings [LM12]. stars [Kut22, Nas11, Win86]. start [Cel13, Col18]. started [Mat22a]. Starting [Tur11]. State [Bid13, CT14, DM13, Ere91, Har19a, Jar17, Kin15b, Low93, Ama05, Blu21a]. statements [Ano72i, Ben22, Nis77, Ueb09]. States [Mee21, Pic04]. statics [De 00]. Statistical [vE14, Ari22, Det22, Hei70, Joh11, Val21]. Statistics [Hag01, BP25, Cat14, Len22]. status [Bog75, Ols16]. steel [Rib07]. Steele [Par95]. Steffens [Ham00]. Steinle [Gur19]. stellar [Eva10, Grü78]. stem [Fag16, Fis23]. stemmatic [ES06]. Stephen [Feh07]. Steps [Lat90, Dra87]. Steve [Gie91b, Rhe94b]. Steven [Blo82c, Hes82, Mac84, San14]. Stevin [Cha15a]. Sticker [Whi74]. Sticky [Slo24]. Stigler [Sch12]. Still [Cho22, Gri91, Sei13a]. Stillman [Ege70a]. stimulus [Chi16, PM23]. stimulus-error [Chi16]. stirps [Pen22a]. Stock [Win99]. Stoic [Nee02]. stone [Fra94b]. stop [Kla00]. stopped [Sta11]. stories [Bud13]. story [Buc09, Cra17, CS17, Eva10, Hee21, LB94, Pet11, Pre16]. story-telling [CS17]. Strait [Cam18b]. strange [Wor82]. Straßer [Kuu09]. stratagem [Fis18]. Strategies [RT95, Cam15, Cas24, Cov97, DFT12, Jar03b, KD19, Lac21, SFT23, Tot12]. strategy [O’F10a, O’F10b, Oni17, San13, Sei13a, Sei13b]. stratified [Elw12]. straw [Mar15b]. streamline [Hof13]. street [Ste24a]. strike [Row13]. String [vD21, Gre22, Rit21]. strings [vL24]. strong [Col94, Díe11, Fre79, Gel80, Had04, Rui23, Ver17, Blo07, Feh07, Fre05, Kem05]. strongly [Díe11]. Structural [Bue19, Gli16, Lyo16, Man22, Tul16, Adl22, Cru05, Dun22, Kro06, Lad98, LB22, MCGP17, VH06]. Structuralism [Ste21, Fre19, McK16]. structuralist [Isa20, Lor13]. Structuralists [Lad19]. Structure [Cam22, Hen21, Kat20, Kuh70, Rot13, Ano72i, Ano81c, Bla13, Blo75, Bol23, BH19, Cap79, CF06, Díe07, Ful18, Hau21, Hes74, Kuh62, Kuh96, KH12, Lem21, Lop84, Lyr14, MM09, Mer92, Mer82, Mum06, Sha17, Sol08b, Tho83, Van86, HH06, HH95, Kva14, Lar03]. Structures [DFT12, Bue97, Con06b, Gut73, JC20, Mül12]. struggle [De 16, Hoe94, LW94]. struggles [EG78]. Struggling [BM89]. STS [Koc06]. Stuart [Hol83]. Stuck [Kuu17]. Studi [Gia94, Bal92]. studied [And13]. Studies [Aba08b, Ano84a, Ano97g, Ano01b, Blu21a, Gam21b, HBR24, Zol93, All96b, Bal92, BM20, Coh97, Gam21a, Hof13, Kai19, Kin15a, Mor19, Nor20a, SGO17, Tuc12, Ueb91b, Yli19, Ano97g, Buc88, ALR21]. study [AM20, Ara92, Bas03, Bec84, Bel90, Bro11a, Chi01, Chi11, Esc08, Fal24, Fed23, Gay77, Gay78, Gei23, Gro87, Hon87a, Jen10, Kra81, LW13a, LA22, Mor92, Mor90a, Möß11a, Möß11b, Mow85, Pat69, Psi94, Ree86, Rom16, Sch24, SW89, Vea72, Yea82, YZ19, Zah19, Hee21]. Style [Hac92, Boa07, Kwa12, Mit22, Wan12, vdBG90]. Styles [Bue12, Elw12, Har93, Rit12a, Rit12b, Win12, Bol14, Koc15a, Kus10, Kwa12, Möß11a, Möß11b, Hop94]. Suárez [GL24]. subactis [Bal92, Gia94]. subject [Koc15b]. subjectivism [Isa14]. subjectivist [Sar97a, Sar97b]. Subjectivity [Koc13]. sublimated [Ano91h]. substance [Gar96, Koc11b, Sch11, Koc09]. Substances [Mau90, Wol20b]. Substantial [Hen06]. substantiality [Gio24b]. substantive [Ano91k]. substitution [Bro73]. succeeds [Ohn21]. Success [Lic21, Bar18, Bie16, Car14, CT13, Ohn21, Ros18, Tri08, Wra18]. Successful [Bol16, Luc20]. suggestion [Wil24b]. Sunspot [May11]. sunspots [Pin09]. superaddition [Con15]. Supercollider [Dom09a]. Superconducting [Dom09a]. superficial [KH25]. supermarket [Pre98]. superposition [Pru89]. Superpositions [Hof13, Gao21]. Superscientist [BD02]. Supervenience [Fut02, Zyn96]. Suppe [Lor13]. Suppes [Díe97b]. support [Kin15a, McC21]. Supporting [McK16]. suppositione [Wis84, Mer82]. Suppressing [Nor21b]. suppression [Coh97]. sure [She99]. surface [McG21]. surprising [May14]. survey [SGO17, Top00, Vet18]. surveying [Ben87, Cuo00, Hac89]. sustainability [Kra24, Sto23]. swapping [Mje24]. Sweden [SL21]. Swedish [Bru20]. switching [Bor25]. Switzerland [Zam18]. sword [Fal98]. syllogism [Pat69, Vea72]. Sylvain [San16b]. symbolic [Bec23, Sha98, Sie10]. symbolical [Mou20]. Symbolischer [Pes13, Sie10]. symbols [Fer01]. Symmetries [Wig67, Wig70, Wig78, Cat22, Cha19a, Ros22, Wal22a, Wal22b, Kir73]. Symmetry [Bor15b, Fee06, Kos99, HG05]. Symposium [Ano87d, KMS73, Whi74, Ano84g]. Syndrome [Buc21, FJ24]. synergy [SFF23]. synthesis [Raf03, Rhe92a, SF25]. synthesizing [Ueb12]. Synthetic [Sim21, KL14, dJ97]. system [Ano91l, Bra24a, Fra93, WHMO22, Wat12, Coc14]. Systematicity [Cam22, Kra18]. systematization [Pfe12, Wag24]. Systemic [Aus78]. Systems [Kra22a, Wal22a, Wal22b, Yos23, vEdJ16, Dar23].

T [Eas72, Ege71, Rhe94a, Swo74, Zam18]. T. [Cra22]. table [Bru07, NRB04, SW01, Sce20, Sch08b, Sch14a, Wil09b]. tacit [Fee16, Jha97, RGD15, Tur11]. tactics [CB92]. Tait [Moy77]. take [She93]. taken [Fra05]. takes [And22]. Taking [Col02a, Edi22, Muk22, Val21, Win99, dOFGB22]. Talbot [Bel15, BDR13]. tale [Fri24a, Ske20, Tam15]. talking [Sta11]. Taming [Stu16]. tank [Muk22]. target [Mat11]. target-oriented [Mat11]. Tartaglia [DD69b]. task [Sar97a, Sar97b]. Tauber [Kam94]. tautology [Tem86]. Taxonomic [Che95, Wan02]. Taxonomy [Wan02, Car93, Fu95]. teacher [Pan96]. Teaching [Wit13, Car21, Joh12b]. tears [Isa19, Isa20]. techne [Cox12]. Technical [Dea03, VH06, DFT12, Han06, HKMV11, Kro06, KM06, Sch06, Sto12, Vae11, Ver06, Wei07]. Technicians [Top04]. techniques [Kim12, Mat22b]. techno [Col21]. techno-legal [Col21]. Technological [Ano90f, Koc04, Lew02, Han07, McK95, Mer92, RZ21, Sch09b, Web22]. Technology [Car13, Ano09f, Bid11, Bow93, Kho03, Kle04, Koc04, Lee19, Rib07, Rob95, dC18]. technoscience [AG13, Sim21]. technosciences [MR13]. Technoscientific [Tam20, Lee19]. Tectonics [MS24]. Tekmeriodic [Har12]. telecommunications [Goo13]. teleological [Car12, McN22]. Teleology [Ano72i, GN22, Len81, Bra81, Car12, Cos98, Nis93, PE24, Why21, vdB24]. tell [Buc09, Car03b, Hee21]. telling [CS17]. Temperature [Dic25]. template [KD19]. template-based [KD19]. templates [HZ19]. temporal [Wat23]. Ten [Pee17]. tension [Dre20, HHL11, Mar97]. tensions [Smi20]. tensors [EG78]. term [Dic25, Man72]. terms [Kav24]. terrestrial [Del22]. terrible [Kon24a]. test [Coh85, EG80]. test-case [Coh85]. testability [Mar09]. Testimonial [Sch02a, DC09]. Testimony [Coa02, Haw10, Kus02b, KL02, Ser99, Sha02, Bur02, Cha10a, Gel10, Ger20, Har02, JFS15, Lip98, LA22, Tuc12, Ueb09]. Testing [Bro16a, Bur02, Gla89a, Gla89b, Hau16, Hen92, McG21, Hau13, LL16, Mor21, WHMO22, Zan88]. tests [Col21, Tro95]. text [Asp13, Eag07, HS23b, Hin12, Mit86, Pal87b, Rhe94a, Ryn95a, Wil12, Zam08]. Texts [Joh12b, Mos07, Vos07]. Textual [Kut22]. Th. [Kon24a]. Thales [She99]. Their [Mer04, Bal20, BS13, Col84, FH84, GG88, Gay76, Jac94, Mac13, Mat19, Mir04, bAM11, PS10, Pin16, Ryn95a, Ryn95b, Tei22, Ver06, Wal22a, Wal22b, Wil10]. Theism [Cal10]. Them [WV22]. themata [Het96]. themes [IG98, Pre98]. themselves [Bre04a]. theologian [Fin02]. theological [Gre92a, Liv93]. Theology [Fun86, Buo24, Cat12, Mou20, O’F10b, Pal02, PEK06, Fyf02, PR90, O’F10a, O’F10b]. theophany [Mou20]. Theorem [Hun21, Kno76, Ama05, Cam18a, Kas14, McL24, Pas11, Yap11, Bla24, How13, Lee13]. theorems [Dar21a]. theoretic [VS12, Dun22]. Theoretical [Ach86, De 24, Mac72, All96a, AS16, Atr85, Cat22, Del19, DS19, Tie88, War89, Psi00]. theorie [Böh80]. Theories [Ful93, Ros17, Rus23, Slo24, Wei99, Acu25, Adl22, Cle97, Con06b, Díe07, Gao21, GVD24, Gie92, Hat88, Hat90, Hau16, Hoc16, Kai19, Kat92, Kle77, KF23, LL16, Let21, Lor13, Mac73, Mar97, McC20, Mok71, Nor86, Prz69, Riv21, RM16, Ros19, San91, Tam15, Tei22, Wal22b, Zan82, Nic72]. theorization [Fis23]. theorizing [LA22, Roc88, Yli19]. Theory [Bas03, lC00, Due21, Ear20, Fre80b, Mag96, Mar21, Mic71a, Miz21, Nea24, Rei20, Rup76, Sch13a, Tur21, AA67, Ara92, AB23, Ari84, Bar20b, BD02, Blu21a, BJ22, Bon94, Bou22a, Bra78, Bru99a, Bru99b, Buc72, Buc92, Bue15, BT09, But93, Cal24, Cam97, Cap79, Car03b, Cas24, Cat12, CL05a, Che90, Cov97, Cur92, D’A00, Dav21a, Dav13, Daw21, DvdB22, Díe97a, Díe97b, Dor70, Dra76, DJ14, Eag04, Esc08, Eve18, Fal98, Fek12, Fey84, Fis85, Fra79a, Fum20, Gab73, Gal71, Gli16, Goe72, Gra71, Gre22, GY11, Hen85, Hen92, Hoe94, Hol22, Hüb80, Hud89, Iva10, Jac21, Jes89, Jha97, Jon78, Kal75, Kim24, Kno76, Kos93, Kra72, Kra73, Kus01, Kuu07, Lau87, Lay78, Lic21, Lin22b, Lip20, Lop84, Mal90, Mas11, MBC17, Mel91, Mer92]. theory [Mic93, Mit17, Mor92, Mow85, Moy78, Nee02, Obe15, Pat69, PM23, Pin97, Pra77, Pre08b, Psi94, Qua75, Ric05, Rit21, RH95, Ros18, Rui23, Saa05, Sar82, Sch73, Sch07b, Sch17, Sch93, ST20, Sco70, Set08, Sie07, Sim70, Sla22, Sta18, SG85, Stu16, Stu20, TC91, Tem86, Tho83, Tho09, Tuc93, Ueb09, Van86, Vas18, Vea72, War92, Wec86, Wel13, Wer80, Wes72, Whi71, Whi88, Whi90, Wil01, Wol24, Wra07, Yea82, dB09, dG20, dJ97, vD21, vL24, BCK^{+}22, Fri24b, Mah99, Rus75, Jan73, Kam94]. theory-choice [CL05a, Sch17]. theory-dependence [Wer80]. Theory-laden [Sch13a]. Theory-ladenness [Bas03, Wec86]. therapeutic [Car25, Maz22]. there [Cus85, SG13, Swa24]. thereby [Cha16]. thermodynamics [Dai19, Psi94, Val21]. thermometers [She11]. Thermoscopes [She11]. thesis [Dea01a, Grü78, Ric09a, San09b, Sar86, Ueb10, Cun01a]. thin [Fum20]. thing [Dan06, Kle14]. Things [Rhe25, dC13, Bre04a, ES12, Löw97, Mir89, Pol15]. think [Eig19, Gen20, Kle79, Pal05, Sch02b]. thinker [Ari22]. Thinking [ES12, Kra22a, Lin22a, Cha19b, Joh19, Mal13, Rit12b, Ryc92]. Third [Zha22, Fra05, McG70, Sta13]. thirteenth [Cun00, Cun01b]. Thomas [Bri73, Zol93, Ale98, Ben24, Cal10, Dac15, Dem08, Gra05, Haw96, HH95, HH06, Möß11a, Möß11b, Pao11, Red02, Sch88a, Sim17]. Thomism [Pal09a]. Thomson [Moy77, Moy78]. those [CF21b]. Thought [Gla99, Möß11a, Möß11b, Bak23, Bec23, Ben74, Buc88, Cam15, Gio14, Har93, Hop94, LC24b, McA96, Pal05, Pru89, Sch08a, Stu16, Stu20, Wan12, vdBG90]. Thought-experimentation [Gla99]. Thoughts [Lau89, Pas62, Cun01a, Dea01a]. thousands [Sch05]. threat [Ano09f]. Three [Bra13, Hen14a, Jan13a, Lee19, MT22, Sie00, Ueb15, Wil24b, Con06a, Daw22, Tam15]. threshing [Fri24a]. thrice [BD02]. Thrown [Tur22a]. tidal [Duc10]. tides [Bro76, CSD21, Mer82, Sch15]. ties [Gar89]. till [Ano14a]. Time [Ano84a, AA67, Art95, Car03b, Cov97, Dar07, For22, Fri21, Fut02, Gio24a, Hen84, KC94, Mau90, Mit93, Rat19, Ryn95a, Ryn95b, Tam20, Thé21, Tho15, Win82, Wit24, Gal71]. timelike [Mje24]. tiny [Car05, Tur04]. titles [Zam08]. tobacco [vH91]. today [Bro11a, LM12]. together [All96a, dR13]. Toine [Lin98]. tomography [Nii11]. Tonal [Isa17]. too [Saa05]. tool [BM23, Krü80]. tools [AM20, Det22, Fre80a, Goc08, Gro05, Jar17, Kle01, Sla02, Wel13]. tooth [Pie14]. topic [SM24]. Topological [Heu14, EZ13]. Toronto [WG13]. Torricelli [Pal09a]. total [Chi01]. touchstone [BG23]. tower [Goo80, Seg89]. toxicology [Bur02]. toy [AG09, Pat09]. tr [Gra72]. Tracing [Buo24, PM23, Rhe92a, SF25, Cra17, Mil19]. tracks [Axt93]. Tractatus [Cla68, Gra72]. traction [Goo21]. Trade [Mat11, PP23]. Trade-offs [Mat11, PP23]. Trading [CEG07, Jen10]. tradition [BV05, Can71, Cro90, Das86, Joh12b, SD23, Sch85]. traditional [Kim12, Stu91a, Wil01]. traditions [Aba10, Gre92a, Liv93, War93]. traduits [Des49]. tragedy [BM21]. trait [MC25]. Traité [Le 58, Le 68, Bri73]. Trans [Sch21a]. Trans-Planckian [Sch21a]. Transactionally [McH21]. Transcendental [Fri03a, Bia20, DiS13, Hon82, Sch08c]. transcontinental [Vet18]. transdisciplinary [KR22]. Transfer [HZ19, Anz19, CP18, Dai19, DS19, HL19, Hum19, Lee19, Pri19, Rhe92b, Rib07, Ull24, Zuc19]. transformation [Bai87, Bru20, Che10, Len22, Len81, Ree86, Rob95]. transformational [Gil83]. Transient [Bid13]. transition [AG13, Bel20, Lev11, Psi94]. transitions [Kob24]. Translated [Eas72, Ano91b, Des49, Sch91, Sha90a, Sha90b]. Translation [San91, Ano91j, Fis23, Gra72, Hin12, Ste04, Van04a]. transmission [Ast04]. Transparency [Qui21, Ell21]. travel [LA22]. Treatise [Le 58, CS78, Cla68, Del22, Gra72, Run94, Sim70, Sto12, Tie88, ES06]. treatments [Gio14]. tree [Why21]. trenches [Sto09b, Wag22]. trends [O’N08]. trial [Fin02, May11]. trials [Due22]. Triangulating [Roc88]. tricky [Nee13]. tried [Cha10c]. Triennial [Ano95c]. triumph [Ege71, Ghi69a, Ghi69b, Ghi72, Ghi84, Ghi03]. trivial [Cru05]. trouble [DJ14]. trouver [Le 58]. True [Ada16, Har24]. Truesdell [Ait70]. Trust [Chu22, Gol22, Joh22, Lea22, WV22, Dar21b, FA13, Har02, Joh21a, dR22]. trust-based [Joh21a]. Trusting [Fri02]. Truth [CT13, Hac12, Lep81, Sar81, Sie83, AS16, Cha10a, Dic06, Fin81, Fis85, Gat16, Gau98, Hat10, Hay04, Hüb80, Jar76, Jas04, Joh19, Kit01, Kui14, Kuu07, Lau82a, Psi97, Rad22, Sar83, Sch05, Spe04, Tos06, Tos07, Wan02]. truth-value [Wan02]. truthfully [Has97]. truthlikeness [CT13]. truths [Fun75, Sch03a]. Tswett [Liv09]. Tudor [Pum12]. tuna [Jen07]. Turbocharging [GG21]. turbulence [Gök90]. Turing [Abr11, Aiz10, CP10, GKS24]. Turing-equivalent [Aiz10]. turn [Dre20, Fra21a, IG98, JW21, Kat92, Sce20, Wit13]. turned [Mat22a]. Turner [Car05]. turning [Bir02, HG21]. turns [Kap16, Tur13]. turtle [Jen10]. Twardy [McA05]. Twelfth [Ano87d]. twelve [Kra24]. twentieth [Bru20, Chi08, Fis06, Pri22, Sla02]. twentieth-century [Chi08, Fis06, Sla02]. twin [Fra21c]. twist [Wra18]. Two [Col90, Díe97a, Fie86, HM88, Har10, HS22, HH95, Olm12, Rap11b, ST20, Bal20, De 14, Ede09, GKS24, Kuk95, O’C05, Ort08, PE24, Ske20, Vor14, VS12, Zam08]. twofold [Bak06]. Tychistic [DE21]. Tycho [Ano91l, How98, Mos01, Rem07]. Tyndall [How04]. type [SG85]. Types [Mil19, Gay76, KL21, Wit24]. typology [Ben74, HD19, Hon89].

U.S. [ND22, Vet18]. Ubaldo [DD69b]. Uebel [Zol93, Fri12]. Ugo [Gia94]. UK [Col21, MHW22, PMG21]. ultimate [Miz12, Mor97]. ultimately [Miz12]. ultra [Ker12]. Umdeutung [Kav24]. Umgebungen [Sie10, Pes13]. uncertainties [Par14, dCG17]. uncertainty [Isa14, RS21a]. Uncovering [Pin14]. Underdetermination [Miy15, Bid13, Brä24b, Dar02]. underlying [Yea82]. understand [Bec16, Che03, FD21]. Understanding [BB22a, Boh00, Cha11, Cor11, Elg22, Gij13, KG13, KB20, Mac75, McK95, RZ21, Sti13, Stu16, Tur13, Wei10c, WH14, dR13, Ano73a, Bec84, Cha12, Col07a, Del16, GP10, Hin13, Hu21, Kau08, Kon24b, Mak03, McG69, Osl72, Str13, Yol70, Lic21, Rog72, Woo71, Woo94]. Undulatory [CB92]. Unearthing [Wu14, Zem06]. unexpected [BR19]. unexplained [Kon24b]. unhappy [Tur22b]. Unification [For88, GFE23, KM14, Duc05, Gij13, Han21, MM09, Mor92]. Unified [Hus24, O’N03, Gla93, Har10, Kle16]. Uniform [Wen23]. uniformity [Cla68, Gra72]. Unifying [Bou22a, Mau77]. Union [BK17]. uniqueness [Díe97b, Tra22]. unit [Gro20, Mit17]. unit-invariant [Gro20]. unitary [Eis21]. unite [Lad19]. units [Ano91k, Ker07, Pie14]. unity [Bra13, RK06, Rei97]. universal [Coo20, ST20]. Universality [Elg22, Ode22, Pin22, Ric22, Jhu22, Kuu11]. Universals [Mäk97]. universe [Ant03, Kob24, LC24a, Osb89, Gra04]. University [Ano78b, Bak16, Bau71, Cra93, Cul04, Ege70a, Ege70b, Ege71, Ere91, Fin73, Ful94, Gra72, Gro90, Gur19, Hei75, Hop94, Jas04, Kam94, Liv93, ML18, Mor90a, Mor90b, Mul94, Mun74, Mye89, Osl72, PR90, Rhe94a, Sch91, Sha90b, Ste04, Swo74, For89, Jar13, Jon13, LW13b, Sum13, WG13]. unknown [Del22]. Unnatural [Bel20, Pap89]. unobservability [Kos88]. unpredictability [BR19, Sar15b]. Unpublished [Tal25, Ano03f, Buo24]. unrealistic [Hin13]. unreasonable [Bok15]. unrevised [Neu71]. unrigorous [MB22]. uns [Le 58]. unsuccessful [Gio16, Sch93]. unusual [DB96, WC09, dS09]. Unwarranted [Sch20b]. upon [SG85]. urban [Alm21]. urgent [Sto23]. US$13.20 [Neu71]. US$35.00 [Ste04]. US$37.99 [Zam18]. US$39.00 [Sie04]. US$65 [Gur19]. US$75.00 [Ste04]. US$98.00 [Sie04]. USA [Bak16, Max22]. Use [Sch08b, Bec71, Dan06, Gay76, Kle08, LA22, Mat19, McN15, Mor19, San13, Sch06, Sch14b, Sei13a, Sei13b, Sta18, WL22, Yea82, Zha22]. Use-novel [Sch08b]. use-novelty [Sch14b]. use-values [Mor19]. used [MJ23]. useful [Hal05]. Usefulness [Mer04]. uses [Cre12, Gro87, Mar97, Smi20, Sta08, SM24, War92]. Using [Aro18, Bec16, Wat23, ES06]. usurpation [Jac94]. utility [Rad13]. Utilizing [RZ21]. Utopian [Car13, Pas20].

V [Bro71, DH69, Nic72]. Vaccine [WV22, Chu22, Gol22, Joh22, Lea22, WV22]. vaccines [Joh22]. vacillating [Wor82]. Vacuum [Jon01, FG22]. Vailati [Pal08a]. valid [FH88, How15]. validation [JB16]. Valles [Kuk22, Myk22, Ups22, Val22]. Value [Alm21, JC21, RE19, Ans02, Bue15, DD13, Don21, Ell21, Fai12, LGC21, Lus21, Sib95, Swa24, Tor24, Wan02, War21]. value-driven [Don21]. Value-entanglement [RE19]. value-freedom [Bue15]. value-laden [War21]. value-ladenness [Ell21]. value-neutral [Lus21]. Values [Bro21, Gal21, McH21, Mil14, Nii14b, Par14, PZ14, Rol16, Ahn20, Bid13, Cho22, Con20, DCJ^{+}21, Ell13, GID21, Kha24, KR22, Lac21, ME15, Mor19, PvF14, Pin16, PP21, Reu22, Ste15, SGO17, UG20, Wie21, Wil09a, WHK14, WOL20a, Zah19]. Vandelli [Car03a]. vanishing [MS72, War22]. Variability [Gio24b]. variables [Mit22]. variations [HS23b, IG98]. Varieties [FR23, KL14, Hun18]. Variety [Abe22, Vez16]. variety-of-evidence [Vez16]. Variorum [Bia94]. vary [Col84, FH84]. vehicles [Dra18]. velocities [FD21]. Venn [Kil99, Wal06]. Venus [Ari87]. vera [Sch20b]. verba [Tuc12]. Verdict [Mos01]. verisimilitude [Nii14a, Row15]. Verlag [Ait70, Neu71, Pes13]. versa [Sie04]. versatility [Pol21]. verses [Has97]. version [Hen85]. Versions [lC00]. verssus [Cle83]. Verständlichkeit [Fey16b]. versus [Ano91k, Bai92, Cha02, Cha16, Chr94, Del16, Gré19, Hoc87, Lic21, Lyo16, WV22]. very [Hon87a, Sha22]. Vestiges [Cha44, Cha69, Ege70b]. Veterinary [Bru20]. vi [Bus71, Ege70b, Ege71]. via [PE24]. viability [Men19]. viable [SL21]. vibrates [FG22]. vice [Sie04]. Victorian [Ege70b, How04, Met91, Mor93, Ric86, Sib95]. Vienna [Ric09a, Ric09b, Ueb05a, Ueb10, Zol93, Ben24, Ueb91b]. Viète [Pan06]. View [Ode22, Bue12, Cha16, Con06b, Dun22, Edi22, For22, Fre25, HM88, Hen84, Hon89, Ken12, Kwa12, LL19, Lor13, Lou08, Men18, Nea24, Ris90, Row08a, Sar97a, Sar97b, Spr10b, Wag22, Zha22]. Viewing [Cha16]. viewpoint [Bac20]. views [Bak23, Feo12]. viii [Blo75, Bro71, Ere91, Fin73, Ful94, Mis94]. Ville [Bla24]. vindicated [Gli16]. Virtual [Tem86]. virtue [Kid11b, Llo15, Sch17, Stu07]. Virtues [Por01, Mar09, dC10]. virus [vH91]. vis [Hep10, Ilt70, Pap77]. visible [Cat14, Goc08, Slo03]. visibles [Gra05]. vision [Sar12, Thr75, Web22]. visited [Lau88]. Visual [Rem07, Bol16, Dou95, HK22, Mal90, Ver07b, WHMO22]. Visualising [BM23]. visuality [Big08]. Vital [Gam21b]. Vitalism [Ben74, Slo77]. vitality [HM24]. Vitellionem [Kep68, KR08, Fie86]. viva [Hep10, Ilt70, Pap77]. Viviani [Seg89]. Vivisection [Hol21a]. Voigt [Fri24a]. Vol [Spa91, Zol93]. vols [May71]. Volume [Ano80g, Ano81d, Ano82d, Ano83d, Ano87e, Ano90g, Ano03g, Ano71h, Ano72g, Ano73e, Ano74f, Ano75e, Ano86g, Isa17, Spa91]. voluntarism [DS19]. Vorgriechische [Neu71]. Vorlesungen [Neu71]. Vries [Dar76]. VS [Neu71, Bal20, Car05, Cus85, Dil90, Fin11, Gio14, Mal90, Nea24, Pap20, Par16, PS10, Row11a, Row13, Sor16, Tur04, Vas00].

W [Bat23, Blo84, Bro71, Moy91a, Moy91b, Osl72, Rhe94a, Rus75, Per08]. W. [DH69]. Waddington [Loi22]. Waegell [Fag24]. Wailoo [Gre24a, Wes23, Slo24]. Waiting [BM14]. waiving [MN20]. wake [She87]. Wald [Bla24]. Wales [Rei08]. Wallace [Bia94]. Wallacian [Mul24]. Wallifaction [Sch88a]. Walls [Sum13]. want [Fra93, Gin07, Han05, Jas04]. wanted [FJ21]. War [AR22, Chu22, Gol22, Joh22, Lea22, WV22, Aro18, Coo95, Dav21b, Edi22, Mac13, Rob22, Bid11, Goo13, Lin22a, Pet16]. warrant [Fri02]. wars [Ada18, Ano97f, JFS97, Ste22c]. Wartime [Mac13, Ras96]. Was [Col16, Mal13, SW12, Sha17, DBM22, Fal17, Jac16, Liv09, Nor85, Rei97, Wat04, Wis18]. wasn’t [DBM22]. Waste [Fra05]. wasted [MC25]. water [Bab25, Hin15, Nee08]. wave [Buc92, Che95, Wes79, Mil14]. wavefunction [Lor22, War22]. waves [Col04, Eld25, Gin07]. Way [Bat23, Dar23, FR23, Mil23, Pal23, Ant25, BM23, Cul04, Le 58, LS02, Sar89, She93, Tur05b, Zha22, Woo24]. Ways [Kle04, Löw08, Aiz18, Cha19b, Iva15, Nis93, Nor12, Pal87b]. weak [FD21, Hon87b, Lee13, Pic84]. weaponization [Kuk22]. Weather [Leh04b, WOL20a]. Weber [Bra01]. Weight [Mil14]. Weight-Adjusting [Mil14]. Weil [KMS73, Whi74]. Weiskopf [GP10]. well [Cur19b, Hoc16, Sto23, Yee24]. well-adapted [Cur19b]. well-being [Yee24]. well-ordered [Sto23]. Weltkenntnis [De 18]. Were [PM23, Mat22a]. West [Mal13, ND22]. western [Pre16]. Westfall [Hei75]. Westview [Gie91b]. Weyl [Pes13, Sie10, Pes13, Sie07, Sie10, Ste24a]. Wh.3469 [Sch09a]. Whatever [Dur13, Pen22b]. wheel [Kra22b]. Wheeler [Blu21b, Fur22]. Where [BB23, Rot22, Tur13, BR19, Le 58, Mat22a, Iva15]. Whewell [Ach92, And94, Bel74, Cob11, CSD21, Duc10, Fis85, Hol83, Joh11, Los83, Men18, Met91, Mor97, Rus75, San16a, Sch73, Sch88c, Sny07, Tha77]. Whewell/Mill [Cob11]. Which [CM22, Dio18, De 18, Lac21, Le 58, Vic21]. Whig [Aba08a, Aba08b]. Whipple [dS09]. Whiston [Cal10, Sno04b]. white [Sch02b]. Whitehead [Lev11, Sie00]. Whither [Lon05]. Whitman [Ste04]. Who [Aga77, CP07, Jac16, Koz15, Ach92]. whole [Fre19, LC24a]. whom [Del23]. Whorfian [IG98]. Wide [Hee21]. Wigner [Kir73, Hun21, Wig67, Wig70, Wig78]. Wilde [Mor93]. Wilderness [Alm21]. Wiley [Van18]. Wilhelm [FWW09, Kim00]. Will [Rot96, Nob23, Ric05, SLH23]. William [Ano73a, Bia94, McC06, Bel15, BDR13, Buo24, Cal10, Cro11, Fis85, Geo14, Hen84, Hol83, Jac16, Joh11, Mei21, Mor91, Mow85, O’F10a, O’F10b, Oli92, Rad13, Rei08, Rus75, Sch88c, Sno04b, Sny07, Uss15, Mis94]. Wilson [Gab73, Riv21]. Wimsatt [Nis77]. wind [Hin15]. Wisconsin [Ege70a, Gra72, Gro90]. Wise [Mor90a]. Wishful [Joh19]. Wissenschaften [Neu71]. Wissenschaftslehre [Car21]. witchcraft [Bro07]. withdrawal [VK19]. withers [Blu21a]. within [HS23b, Roc02]. without [Ano91h, AP06, CP18, Coe22a, Dac15, Eig19, Ful94, GG21, Isa19, Isa20, Kit93, KSZ24, Len81, LL19, Mal93, Mau77, ND21, Pie14, Psi97, Row15, Rup22, Str13, Ven22]. Witmer [Eas72]. Witnesses [De 02]. Wittgenstein [Ano87d, Bar80, Bar20a, Blo73, Pat09, Ste05, Wil89]. Wolf [Ser23]. Wolfe [Zam18]. Wolff [Dun13]. Wollaston [Jac16, Uss15]. women [Leu15]. won [Ach92]. Wonder [Tyb03]. Wonder-making [Tyb03]. won’t [WOL20a]. Woodward [Sla02]. Woolgar [Rhe94b]. Woolhouse [Rog72]. Wootton [Bar20a]. word [Cul04, LS02]. Words [DH69, Cun88, Bro71]. Wordsworth [Lea98]. work [Ber85, Bid11, Bla72, BK21, Cha02, Cha16, DH69, Dem08, Dio18, For75, Hel16, Kem05, Nir23, Swo74]. Working [Mor07, Mas15]. works [Ano03f, Bro71, Föl12, Smi13b]. Workshop [Ano80f, Gal96]. World [Cha20a, Cha20b, Goo13, Mel04, Win99, Bro16c, Coc14, Cus85, Kra22b, Lad19, Lan20, Mas15, MBG21, Noa08, Pal87b, Pet11, Ros17, Sch70, Sch88d, SF25, Tri08, Tro95]. worlds [Bal20, Haw96, Lan14, Sny07]. worldview [Fre21]. Worldviews [Sta08]. Worrall [Bar05, Sce05, Sch08b]. worries [Mel91]. worth [Pan96]. worthiness [Lic21, Rui23]. worthy [Bro11a, SW12, Wil24b]. would [BR19, Cam15, Mau90]. Wright [HL13, Ros17]. Writing [Jon01, Kim95, DFT12, Rhe92a, Rhe92b]. writings [Ano03f, Bid11, Coo96, Dra86, Kus16, Wat03, vS20]. Written [Dem06, Baz88, Des49, Gro90, Wil09b]. wrong [Bir02, Car91, CF06, DBM22, Yea05]. wrongdoing [RZ21]. wrote [Par12]. Würzburg [Ano95c].

X [Kam94, Ach92, Ait70, Gie91b, Gur19, Osl72, Rhe94a, Rhe94b, Sha90a, WC09, Sch09a]. x-phi [WC09]. X-ray [Sch09a]. Xenophon [Sto12]. xi [Jan73, Mul94, Zam18]. xii [Ano73a, Ege70a, Ful93, Luc71, McR72, Neu71, PR90, Sie04, Whi74, Zol93]. xiii [Gra72]. xiv [Bau71, Gab73, Kni71]. xix [Cra93, Hop94]. XIXth [Ano91m]. xvi [Sha90b, Van18]. xvii [Cul04, Sno04a, Vea72]. xviii [Kam94, Mel04]. xx [Gui04a, Swo74]. xxiv [Eas72]. xxv [Mor90a]. XXXII [Gua13].

Yale [Cul04]. Yearley [Mac84]. Years [Dra86, Díe97a, Díe97b, Dil90, Hac12, Lau89, Lau87, tH02]. Yolton [Osl72]. Yong [Kat20]. York [Ait70, Ano91b, Bri73, Bro71, Car73, Ege70b, Ere91, Ful94, Gab73, Gal71, Kni71, May71, Neu71, Vea72]. young [Ber80, MC25, Che90].

Z [Gui04a]. Zafiris [Heu14]. Zanzibar [GMC22]. Zealand [Cra22, Ste22c]. Zeeman [Ara92]. Zeno [SHB94]. Zone [Ano91b, Jen10, Pri19]. zones [CEG07]. zur [Car21, KR08]. Zurich [Pes13, Sie07].

References

Al-Azm:1967:KTT

[AA67] Sadik J. Al-Azm. Kant’s theory of time. Philosophical Library, New York, NY, USA, 1967. 84 pp. LCCN B2799.T5 A9.

Ardourel:2023:FSS

[AB23] Vincent Ardourel and Sorin Bangu. Finite-size scaling theory: Quantitative and qualitative approaches to critical phenomena. Studies in History and Philosophy of Science Part A, 100(??):99– 106, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000900.

Abadia:2008:BWHa

[Aba08a] Oscar Moro Abadía. Beyond the Whig history interpretation of history: lessons on ‘presentism’ from Hélène Metzger. Studies in History and Philosophy of Science Part A, 39(2):194– 201, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000216. See [Aba08b].

Abadia:2008:BWHb

[Aba08b] Oscar Moro Abadia. Beyond the Whig history interpretation of history: lessons on ‘presentism’ from Hélène Metzger: Studies in History and Philosophy of Science, 39(2), 194–201. Studies in History and Philosophy of Science Part A, 39(4):565, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000915. See [Aba08a].

Abadia:2010:CHT

[Aba10] Oscar Moro Abadía. Connecting historiographical traditions. Studies in History and Philosophy of Science Part A, 41(1):105– 108, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000880.

Abadia:2011:HCH

[Aba11] Oscar Moro Abadía. Hermeneutical contributions to the history of science: Gadamer on ‘presentism’. Studies in History and Philosophy of Science Part A, 42(2):372–380, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811000124X.

Abelson:2022:VEM

[Abe22] Shannon Sylvie Abelson. Variety of evidence in multimessenger astronomy. Studies in History and Philosophy of Science Part A, 94(??):133–142, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000802.

Abel:2024:SME

[Abe24] David Lynn Abel. Selection in molecular evolution. Studies in History and Philosophy of Science Part A, 107(??):54– 63, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001110.

Ablondi:2015:IGE

[Abl15] Fred Ablondi. Introduction: Galileo and early modern philosophy. Studies in History and Philosophy of Science Part A, 51 (??):69, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000217.

Abramson:2011:DIT

[Abr11] Darren Abramson. Descartes’ influence on Turing. Studies in History and Philosophy of Science Part A, 42(4):544–551, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000689.

Achinstein:1986:TD

[Ach86] Peter Achinstein. Theoretical derivations. Studies in History and Philosophy of Science Part A, 17(4):375–414, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900014.

Achinstein:1987:LH

[Ach87] Peter Achinstein. Light hypotheses. Studies in History and Philosophy of Science Part A, 18(3):293–337, September 1987. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0039368187900239. See response [Che90] and reply [Ach90].

Achinstein:1990:LPR

[Ach90] Peter Achinstein. Light problems: Reply to Chen. Studies in History and Philosophy of Science Part A, 21(4):677–684, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819090038A. See [Ach87, Che90].

Achinstein:1992:BRB

[Ach92] Peter Achinstein. Book review: Inference to the best explanation: Or, who won the Mill–Whewell debate?: Peter Lipton (London: Routledge, 1991), x + 194 pp. ISBN 0-415-05886-4 Cloth £35.00. Studies in History and Philosophy of Science Part A, 23(2):349–364, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368192900399. See [Lip91].

Achinstein:2023:DER

[Ach23] Peter Achinstein. Disregarding evidence: Reasonable options for Newton and Rutherford? Studies in History and Philosophy of Science Part A, 97(??):111–120, February 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001789.

Acuna:2025:TCL

[Acu25] Pablo Acuña. Through the convex Looking Glass: a Helmholtzian lesson for the connection between dynamics and chronogeometry in spacetime theories. Studies in History and Philosophy of Science Part A, 109(??):31–46, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001432.

Ackermann:2012:LSM

[AD12] Silke Ackermann and Louise Devoy. ‘The Lord of the smoking mirror’: Objects associated with John Dee in the British Museum. Studies in History and Philosophy of Science Part A, 43 (3):539–549, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001117.

Adams:2016:HNP

[Ada16] Marcus P. Adams. Hobbes on natural philosophy as “True Physics” and mixed mathematics. Studies in History and Philosophy of Science Part A, 56(??):43–51, April 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001600.

Adams:2018:CW

[Ada18] Fred Adams. Cognition wars. Studies in History and Philosophy of Science Part A, 68(??):20–30, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302741. See reply [SOC19].

Adlam:2022:OTS

[Adl22] Emily Adlam. Operational theories as structural realism. Studies in History and Philosophy of Science Part A, 94(??):99– 111, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000814.

Al-Gailani:2009:MSM

[AG09] Salim Al-Gailani. Magic, science and masculinity: marketing toy chemistry sets. Studies in History and Philosophy of Science Part A, 40(4):372–381, December 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000703.

Arapostathis:2013:ETP

[AG13] Stathis Arapostathis and Graeme Gooday. Electrical technoscience and physics in transition, 1880–1920. Studies in History and Philosophy of Science Part A, 44(2):202–211, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112001021.

Agassi:1971:DAA

[Aga71] Joseph Agassi. Discussion: Agassi’s alleged arbitrariness. Studies in History and Philosophy of Science Part A, 2(2):157–165, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936817190029X.

Agassi:1977:WDB

[Aga77]  Joseph Agassi. Who discovered Boyle’s Law?  Studies in

History and Philosophy of Science Part A, 8(3):189–250, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900073.

Arabatzis:2015:IIH

[AH15] Theodore Arabatzis and Don Howard. Introduction: Integrated history and philosophy of science in practice. Studies in History and Philosophy of Science Part A, 50(??):1–3, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000922.

Ahnert:2004:NEE

[Ahn04] Thomas Ahnert. Newtonianism in early Enlightenment Germany, c. 1720 to 1750: metaphysics and the critique of dogmatic philosophy. Studies in History and Philosophy of Science Part A, 35 (3):471–491, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000500.

Ahn:2020:HNE

[Ahn20] Soohyun Ahn. How non-epistemic values can be epistemically beneficial in scientific classification. Studies in History and Philosophy of Science Part A, 84(??):57–65, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812030159X.

Aiton:1970:BRB

[Ait70] E. J. Aiton. Book review: Essays in the history of mechanics: C. Truesdell, Springer-Verlag: New York, 1968. x + 384 pp., 126 figs., indices. $19.50. Studies in History and Philosophy of Science Part A, 1(3):265–273, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368170900130. See [Tru68].

Aizawa:2010:CCS

[Aiz10] Kenneth Aizawa. Computation in cognitive science: it is not all about Turing-equivalent computation. Studies in History and Philosophy of Science Part A, 41(3):227–236, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000452.

Aizawa:2018:MRM

[Aiz18] Kenneth Aizawa. Multiple realization and multiple “ways” of realization: a progress report. Studies in History and Philosophy of Science Part A, 68(??):3–9, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302716.

Ankeny:2011:WSA

[AL11] Rachel A. Ankeny and Sabina Leonelli. What’s so special about model organisms? Studies in History and Philosophy of Science Part A, 42(2):313–323, June 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001184.

Ankeny:2016:RPK

[AL16] Rachel A. Ankeny and Sabina Leonelli. Repertoires: a post-Kuhnian perspective on scientific change and collaborative research. Studies in History and Philosophy of Science Part A, 60 (??):18–28, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300449.

Alexander:1995:ISE

[Ale95] Amir Alexander. The imperialist space of Elizabethan mathematics. Studies in History and Philosophy of Science Part A, 26 (4):559–591, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000232.

Alexander:1998:LMC

[Ale98] Amir Alexander. Lunar maps and coastal outlines: Thomas Hariot’s mapping of the Moon. Studies in History and Philosophy of Science Part A, 29(3):345–368, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000119.

Algazi:2008:NEM

[Alg08] Gadi Algazi. Norbert Elias’s motion pictures: history, cinema and gestures in the process of civilization. Studies in History and Philosophy of Science Part A, 39(3):444–458, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000678.

Allchin:1996:CTC

[All96a] Douglas Allchin. Cellular and theoretical chimeras: Piecing together how cells process energy. Studies in History and Philosophy of Science Part A, 27(1):31–41, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000275.

Allwood:1996:CPS

[All96b] Carl Martin Allwood. A cognitive perspective on science studies. Studies in History and Philosophy of Science Part A, 27(4):599– 605, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368196000404.

Allzen:2021:SRE

[All21] Simon Allzén. Scientific realism and empirical confirmation: a puzzle. Studies in History and Philosophy of Science Part A, 90 (??):153–159, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001643.

Almeida:2012:ODM

[Alm12] Bruno Almeida. On the origins of Dee’s mathematical programme: the John Dee-Pedro Nunes connection. Studies in History and Philosophy of Science Part A, 43(3):460–469, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111001154.

Almassi:2021:VDU

[Alm21] Ben Almassi. Value disputes in urban ecological restoration: Lessons from the Chicago Wilderness. Studies in History and Philosophy of Science Part A, 87(??):93–100, June 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000388.

Ankeny:2021:SBC

[ALR21] Rachel A. Ankeny, James Ladyman, and Darrell Rowbottom. Studies A, B, and C merger. Studies in History and Philosophy of Science Part A, 87(??):A1, June 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000558.

Althoff:2012:PDP

[Alt12] Jochen Althoff. Presocratic discourse in poetry and prose: the case of Empedocles and Anaxagoras. Studies in History and Philosophy of Science Part A, 43(2):293–299, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111001270.

Allisson:2020:SHT

[AM20] François Allisson and Antoine Missemer. Some historiographical tools for the study of intellectual legacies. Studies in History and Philosophy of Science Part A, 84(??):132–141, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301746.

Amadae:2005:AIT

[Ama05] S. M. Amadae. Arrow’s impossibility theorem and the national security state. Studies in History and Philosophy of Science Part A, 36(4):734–743, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000749.

Anderson:1994:WFE

[And94] Ronald Anderson. The Whewell–Faraday exchange on the application of the concepts of momentum and inertia to electromagnetic phenomena. Studies in History and Philosophy of Science Part A, 25(4):577–594, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900485.

Anderson:2013:CLH

[And13] Robert G. W. Anderson. Chemistry laboratories, and how they might be studied. Studies in History and Philosophy of Science Part A, 44(4):669–675, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000459.

Andersen:2016:CIE

[And16] Hanne Andersen. Collaboration, interdisciplinarity, and the epistemology of contemporary science. Studies in History and Philosophy of Science Part A, 56(??):1–10, April 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001442.

Anderson:2022:HPS

[And22] Warwick Anderson. History and philosophy of science takes form. Studies in History and Philosophy of Science Part A, 93(??):175– 182, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000607.

Ankeny:2023:E

[Ank23] Rachel A. Ankeny. Editorial. Studies in History and Philosophy of Science Part A, 97(??):A1–A2, February 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000419.

Anonymous:1970:A

[Ano70a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 1(2):184, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900063.

Anonymous:1970:BRa

[Ano70b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 1(1):87–89, May 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900269.

Anonymous:1970:BRc

[Ano70c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 1(2):185, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900075.

Anonymous:1970:BRd

[Ano70d] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 1(3):275, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900142.

Anonymous:1970:EB

[Ano70e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 1(1):??, May 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900221.

Anonymous:1970:ENa

[Ano70f] Anonymous. Editors’ note. Studies in History and Philosophy of Science Part A, 1(1):1–2, May 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900233.

Anonymous:1970:ENb

[Ano70g] Anonymous. Editors’ note. Studies in History and Philosophy of Science Part A, 1(3):188, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900087.

Anonymous:1971:Aa

[Ano71a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 2(2):194, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900355.

Anonymous:1971:Ab

[Ano71b] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 2(3):272, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900446.

Anonymous:1971:BRa

[Ano71c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 1(4):381, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817190015X.

Anonymous:1971:BRb

[Ano71d] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 2(2):191–192, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900331.

Anonymous:1971:BRc

[Ano71e] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 2(3):295, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900483.

Anonymous:1971:C

[Ano71f] Anonymous. Corrigendum. Studies in History and Philosophy of Science Part A, 1(4):383, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900161.

Anonymous:1971:EB

[Ano71g] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 2(1):??, May 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900367.

Anonymous:1971:IV

[Ano71h] Anonymous. Index to volume 1. Studies in History and Philosophy of Science Part A, 1(4):i–vi, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900173.

Anonymous:1971:NC

[Ano71i] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 2(2):193, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900343.

Anonymous:1972:BRa

[Ano72a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 2(4):379, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900064.

Anonymous:1972:BRb

[Ano72b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 3(2):197, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817290026X.

Anonymous:1972:BRc

[Ano72c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 3(3):299, November 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817290012X.

Anonymous:1972:BRM

[Ano72d] Anonymous. Books received — May 1972. Studies in History and Philosophy of Science Part A, 3(1):101, May 1972. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368172900179.

Anonymous:1972:EB

[Ano72e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 3(1):??, May 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900131.

Anonymous:1972:EN

[Ano72f] Anonymous. Editors’ note. Studies in History and Philosophy of Science Part A, 3(3):202, November 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900088.

Anonymous:1972:IV

[Ano72g] Anonymous. Index to volume 2. Studies in History and Philosophy of Science Part A, 2(4):v–vi, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900076.

Anonymous:1972:NC

[Ano72h] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 3(2):199–200, August 1972. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368172900271.

Anonymous:1972:TLS

[Ano72i] Anonymous. Teleology and the logical structure of function statements. Studies in History and Philosophy of Science Part A, 3 (1):1–80, May 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900143.

Anonymous:1973:BRB

[Ano73a] Anonymous. Book review: Nineteenth-century spectroscopy: Development of the understanding of Spectra 1802–1897: William McGucken, Baltimore and London: The Johns Hopkins Press, 1969. xii + 233 pp. $11.00. Studies in History and Philosophy of Science Part A, 4(1):95–104, May 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900241.

See [McG69].

Anonymous:1973:BRa

[Ano73b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 3(4):397–398, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900198.

Anonymous:1973:BRc

[Ano73c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 4(1):105, May 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900253.

Anonymous:1973:EB

[Ano73d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 4(1):??, May 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900216.

Anonymous:1973:IV

[Ano73e] Anonymous. Index to volume 3. Studies in History and Philosophy of Science Part A, 3(4):i–vii, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900204.

Anonymous:1973:NCa

[Ano73f] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 4(2):208, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900058.

Anonymous:1973:NCb

[Ano73g] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 4(3):307, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900149.

Anonymous:1974:A

[Ano74a] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 4(4):396, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900132.

Anonymous:1974:BRa

[Ano74b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 4(4):393, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900119.

Anonymous:1974:BRb

[Ano74c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 5(1):94, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900223.

Anonymous:1974:BRc

[Ano74d] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 5(3):307, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900053.

Anonymous:1974:EB

[Ano74e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 5(1):??, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900156.

Anonymous:1974:IV

[Ano74f] Anonymous. Index to volume 4. Studies in History and Philosophy of Science Part A, 4(4):401–402, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900144.

Anonymous:1974:NCa

[Ano74g] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 4(4):395, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900120.

Anonymous:1974:NCb

[Ano74h] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 5(1):95, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900235.

Anonymous:1974:NCc

[Ano74i] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 5(2):203, August 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900260.

Anonymous:1974:NCd

[Ano74j] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 5(3):308, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900065.

Anonymous:1975:BRa

[Ano75a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 5(4):400, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900138.

Anonymous:1975:BRb

[Ano75b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 6(4):347–348, November 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900291.

Anonymous:1975:EB

[Ano75c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 6(1):??, April 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900011.

Anonymous:1975:EN

[Ano75d] Anonymous. Editor’s note. Studies in History and Philosophy of Science Part A, 6(1):1, April 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900023.

Anonymous:1975:IV

[Ano75e] Anonymous. Index to volume 5. Studies in History and Philosophy of Science Part A, 5(4):i–iv, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900151.

Anonymous:1975:NCa

[Ano75f] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 5(4):401, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817590014X.

Anonymous:1975:NCb

[Ano75g] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 6(1):85, April 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900059.

Anonymous:1975:NCc

[Ano75h] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 6(2):183, June 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900205.

Anonymous:1976:EB

[Ano76a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 7(1):??, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368176900017.

Anonymous:1976:NCa

[Ano76b] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 7(1):87, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368176900054.

Anonymous:1976:NCb

[Ano76c] Anonymous. Notes on contributors. Studies in History and Philosophy of Science Part A, 7(2):187, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368176900169.

Anonymous:1977:A

[Ano77a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 8(2):173, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900140.

Anonymous:1977:BR

[Ano77b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 8(1):85, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900218.

Anonymous:1977:C

[Ano77c] Anonymous. Corrigenda. Studies in History and Philosophy of Science Part A, 8(3):269, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900097.

Anonymous:1977:EB

[Ano77d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 8(1):??, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900152.

Anonymous:1978:BR

[Ano78a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 9(1):81–82, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900237.

Anonymous:1978:BUB

[Ano78b] Anonymous. Boston University: Boston Colloquium for the Philosophy of Science 1977–1978. Studies in History and Philosophy of Science Part A, 9(1):79–80, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900225.

Anonymous:1978:EB

[Ano78c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 9(1):??, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900171.

Anonymous:1978:E

[Ano78d] Anonymous. Erratum. Studies in History and Philosophy of Science Part A, 9(4):??, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900158.

Anonymous:1979:A

[Ano79a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 10(3):264, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900116.

Anonymous:1979:BRa

[Ano79b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 10(1):85–87, March 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900050.

Anonymous:1979:BRc

[Ano79c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 10(3):263–264, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900104.

Anonymous:1979:BRB

[Ano79d] Anonymous. Books received B7. Studies in History and Philosophy of Science Part A, 10(2):175, June 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900153.

Anonymous:1979:EB

[Ano79e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 10(1):??, March 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900013.

Anonymous:1980:A

[Ano80a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 11(2):174, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900242.

Anonymous:1980:BR

[Ano80b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 11(2):173–174, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900230.

Anonymous:1980:CIS

[Ano80c] Anonymous. Cheiron: The International Society for the History of Behavioral and Social Sciences. Studies in History and Philosophy of Science Part A, 11(4):350, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900138.

Anonymous:1980:E

[Ano80d] Anonymous. Editorial. Studies in History and Philosophy of Science Part A, 11(2):v, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900151.

Anonymous:1980:EB

[Ano80e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 11(1):??, March 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900011.

Anonymous:1980:NWP

[Ano80f] Anonymous. NSF Workshop on Philosophy of Science. Studies in History and Philosophy of Science Part A, 11(4):349, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368180900126.

Anonymous:1980:VCA

[Ano80g] Anonymous. Volume contents and author index to volume 11, 1980. Studies in History and Philosophy of Science Part A, 11(4): i–iv, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818090014X.

Anonymous:1981:BR

[Ano81a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 12(4):355–357, December 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368181900200.

Anonymous:1981:EB

[Ano81b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 12(1):??, March 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368181900017.

Anonymous:1981:FSH

[Ano81c] Anonymous. Fine structure history of science: Lessons for methodology. Studies in History and Philosophy of Science Part A, 12(3): 173, September 1981. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368181900078.

Anonymous:1981:VCA

[Ano81d] Anonymous. Volume 12 contents and author index. Studies in History and Philosophy of Science Part A, 12(4):i–v, December 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900212.

Anonymous:1982:A

[Ano82a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 13(2):173, June 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900243.

Anonymous:1982:BR

[Ano82b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 13(3):265, September 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900115.

Anonymous:1982:EB

[Ano82c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 13(1):??, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900012.

Anonymous:1982:VCA

[Ano82d] Anonymous. Volume 13 contents and author index. Studies in History and Philosophy of Science Part A, 13(4):i–v, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900206.

Anonymous:1983:BR

[Ano83a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 14(2):169, June 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368183900183.

Anonymous:1983:EB

[Ano83b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 14(1):??, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368183900018.

Anonymous:1983:E

[Ano83c] Anonymous. Erratum. Studies in History and Philosophy of Science Part A, 14(2):171, June 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368183900195.

Anonymous:1983:VCA

[Ano83d] Anonymous. Volume contents and author index to volume 14, 1983. Studies in History and Philosophy of Science Part A, 14(4): i–v, December 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368183900110.

Anonymous:1984:ASS

[Ano84a] Anonymous. Association for Social Studies of Time (ASSET). Studies in History and Philosophy of Science Part A, 15(3):264, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368184900256.

Anonymous:1984:BR

[Ano84b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 15(1):83–84, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900311.

Anonymous:1984:BSH

[Ano84c] Anonymous. British Society for the History of Philosophy. Studies in History and Philosophy of Science Part A, 15(3):263–264, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368184900244.

Anonymous:1984:CBD

[Ano84d] Anonymous. Commemoration of the bicentenary of the death of Dénis Diderot. Studies in History and Philosophy of Science Part A, 15(3):261–262, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900220.

Anonymous:1984:CAI

[Ano84e] Anonymous. Contents and author index. Studies in History and Philosophy of Science Part A, 15(4):iii–v, December 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368184900153.

Anonymous:1984:EB

[Ano84f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 15(1):??, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900268.

Anonymous:1984:HMS

[Ano84g] Anonymous. The Helene Metzger symposium: Paris, December 1984. Studies in History and Philosophy of Science Part A, 15(3): 263, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368184900232.

Anonymous:1985:Aa

[Ano85a] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 16(2):167, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900123.

Anonymous:1985:Ab

[Ano85b] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 16(4):399–400, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900214.

Anonymous:1985:BRa

[Ano85c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 16(2):165, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900111.

Anonymous:1985:BRb

[Ano85d] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 16(4):395–397, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900202.

Anonymous:1985:EB

[Ano85e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 16(1):??, March 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900044.

Anonymous:1985:E

[Ano85f] Anonymous. Erratum. Studies in History and Philosophy of Science Part A, 16(2):168, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900135.

Anonymous:1985:LCA

[Ano85g] Anonymous. List of contents and author index. Studies in History and Philosophy of Science Part A, 16(4):i–v, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900226.

Anonymous:1986:A

[Ano86a] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 17(2):247, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900282.

Anonymous:1986:BRa

[Ano86b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 17(1):129, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900221.

Anonymous:1986:BRb

[Ano86c] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 17(4):505, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900075.

Anonymous:1986:CPP

[Ano86d] Anonymous. Conference on physics and philosophy. Studies in History and Philosophy of Science Part A, 17(2):248, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900294.

Anonymous:1986:EB

[Ano86e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 17(1):??, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900154.

Anonymous:1986:E

[Ano86f] Anonymous. Erratum. Studies in History and Philosophy of Science Part A, 17(1):131, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900233.

Anonymous:1986:IV

[Ano86g] Anonymous. Index to volume 17, 1986. Studies in History and Philosophy of Science Part A, 17(4):i–vii, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900087.

Anonymous:1986:NPS

[Ano86h] Anonymous. Newton’s philosophical and scientific legacy. Studies in History and Philosophy of Science Part A, 17(3):373, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900142.

Anonymous:1987:BRa

[Ano87a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 18(2):253, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900203.

Anonymous:1987:BRb

[Ano87b] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 18(3):383, September 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900264.

Anonymous:1987:EB

[Ano87c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 18(1):??, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900082.

Anonymous:1987:TIW

[Ano87d] Anonymous. Twelfth International Wittgenstein Symposium: Philosophy of Law, Politics, and Society. Studies in History and Philosophy of Science Part A, 18(1):107, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900148.

Anonymous:1987:VLC

[Ano87e] Anonymous. Volume 18 list of contents and author index. Studies in History and Philosophy of Science Part A, 18(4):i–vii, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368187900070.

Anonymous:1988:BR

[Ano88a] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 19(1):139, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368188900234.

Anonymous:1988:EB

[Ano88b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 19(1):??, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368188900167.

Anonymous:1988:E

[Ano88c] Anonymous. Errata. Studies in History and Philosophy of Science Part A, 19(3):397, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368188900076.

Anonymous:1988:LCA

[Ano88d] Anonymous. List of contents and author index. Studies in History and Philosophy of Science Part A, 19(4):i–vii, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900155.

Anonymous:1989:ICH

[Ano89a] Anonymous. 18th International Congress on the History of Science. Studies in History and Philosophy of Science Part A, 20 (1):155, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818990040X.

Anonymous:1989:AI

[Ano89b] Anonymous. Anniversary issue. Studies in History and Philosophy of Science Part A, 20(1):3, March 1989. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900289.

Anonymous:1989:A

[Ano89c] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 20(2):285–286, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900083.

Anonymous:1989:BRa

[Ano89d] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 20(1):151–153, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900393.

Anonymous:1989:BRb

[Ano89e] Anonymous. Books received. Studies in History and Philosophy of Science Part A, 20(4):553–557, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900241.

Anonymous:1989:EB

[Ano89f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 20(1):??, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900265.

Anonymous:1989:FE

[Ano89g] Anonymous. Founding editors. Studies in History and Philosophy of Science Part A, 20(1):1, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900277.

Anonymous:1989:LCA

[Ano89h] Anonymous. List of contents and author index. Studies in History and Philosophy of Science Part A, 20(4):i–vii, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900253.

Anonymous:1990:A

[Ano90a] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 21(4):691, December 1990. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900416.

Anonymous:1990:BHP

[Ano90b] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 21(3):527– 530, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819090011V.

Anonymous:1990:EB

[Ano90c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 21(1):??, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819090012W.

Anonymous:1990:FHH

[Ano90d] Anonymous. Forum for history of human science. Studies in History and Philosophy of Science Part A, 21(4):691–692, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900427.

Anonymous:1990:PPE

[Ano90e] Anonymous. Philosophical problems in evolutionary biology. Studies in History and Philosophy of Science Part A, 21(1):182, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819090021Y.

Anonymous:1990:TDS

[Ano90f] Anonymous. Technological development and science in the 19th and 20th centuries. Studies in History and Philosophy of Science Part A, 21(1):181, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900209.

Anonymous:1990:VLC

[Ano90g] Anonymous. Volume 21 list of contents and author index. Studies in History and Philosophy of Science Part A, 21(4):i–vii, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900438.

Anonymous:1991:A

[Ano91a] Anonymous. Announcements. Studies in History and Philosophy of Science Part A, 22(1):197–199, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190022K.

Anonymous:1991:BRB

[Ano91b] Anonymous. Book review: The normal and the pathological: Georges Canguilhem, with an introduction by Michel Foucault, translated by Carolyn R. Fawcett in collaboration with Robert S. Cohen (New York: Zone Books, 1989), 327 pp. ISBN 0-942299-58-2 Cloth $24.50. Studies in History and Philosophy of Science Part A, 22(2):347–369, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819190031M. See [Can89].

Anonymous:1991:BHPa

[Ano91c] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 22(1):193– 196, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190021J.

Anonymous:1991:BHPb

[Ano91d] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 22(3):539– 541, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190010P.

Anonymous:1991:EB

[Ano91e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 22(1):??, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190011G.

Anonymous:1991:E

[Ano91f] Anonymous. Erratum. Studies in History and Philosophy of Science Part A, 22(1):200, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190023L.

Anonymous:1991:LCA

[Ano91g] Anonymous. List of contents and author index. Studies in History and Philosophy of Science Part A, 22(4):i–vii, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190042Q.

Anonymous:1991:NES

[Ano91h] Anonymous. Naturalized epistemology sublimated: Rapprochement without the ruts. Studies in History and Philosophy of Science Part A, 22(2):277–293, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190028Q.

Anonymous:1991:PHM

[Ano91i] Anonymous. Politics in Hobbes’ mechanics: The social as enabling. Studies in History and Philosophy of Science Part A, 22(2): 295–320, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190029R.

Anonymous:1991:PIT

[Ano91j] Anonymous. Presentism and the indeterminacy of translation. Studies in History and Philosophy of Science Part A, 22(2):321– 345, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190030V.

Anonymous:1991:RSC

[Ano91k] Anonymous. Realism and simplicity in the Castle—East debate on the stability of the hereditary units: Rhetorical devices versus substantive methodology. Studies in History and Philosophy of Science Part A, 22(2):201–221, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190024M.

Anonymous:1991:TSG

[Ano91l] Anonymous. Tycho’s system and Galileo’s dialogue. Studies in History and Philosophy of Science Part A, 22(2):259–275, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190027P.

Anonymous:1991:XIC

[Ano91m] Anonymous. XIXth International Congress of History of Science. Studies in History and Philosophy of Science Part A, 22 (2):371, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190032N.

Anonymous:1992:A

[Ano92a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 23(3):531, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290011T.

Anonymous:1992:AI

[Ano92b] Anonymous. Author index. Studies in History and Philosophy of Science Part A, 23(4):i–vii, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290022X.

Anonymous:1992:BHPa

[Ano92c] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 23(1):193– 194, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819290031Z.

Anonymous:1992:BHPb

[Ano92d] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 23(3):529– 530, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368192900104.

Anonymous:1992:CEH

[Ano92e] Anonymous. Conference on evolution and the human sciences. Studies in History and Philosophy of Science Part A, 23(4):699– 700, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819290021W.

Anonymous:1992:EB

[Ano92f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 23(1):??, March 1992. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290023Y.

Anonymous:1993:Aa

[Ano93a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 24(1):159–162, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390030N.

Anonymous:1993:Ab

[Ano93b] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 24(3):511, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390043J.

Anonymous:1993:Ac

[Ano93c] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 24(4):695, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390060W.

Anonymous:1993:BHPa

[Ano93d] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 24(1):157– 158, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390029J.

Anonymous:1993:BHPb

[Ano93e] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 24(3):509– 510, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390042I.

Anonymous:1993:EB

[Ano93f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 24(1):??, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390021B.

Anonymous:1994:Aa

[Ano94a] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 25(2):285, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900345.

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcement. Studies in History and Philosophy of Science Part A, 25(3):511, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900655.

Anonymous:1994:BHP

[Ano94c] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 25(2):281– 283, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900337.

Anonymous:1994:EB

[Ano94d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 25(1):??, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900175.

Anonymous:1995:EB

[Ano95a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 26(1):??, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195900055.

Anonymous:1995:LCA

[Ano95b] Anonymous. List of contents and author index. Studies in History and Philosophy of Science Part A, 26(4):iii–vii, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195900195.

Anonymous:1995:TCE

[Ano95c] Anonymous. The 3rd Triennial Conference of the European Association for the History of Psychiatry (EAHP): Würzburg, Germany, 11–14 September 1996. Studies in History and Philosophy of Science Part A, 26(1):173, March 1995. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195900071.

Anonymous:1996:BHP

[Ano96a] Anonymous. Books on history and philosophy of science received to June 1996. Studies in History and Philosophy of Science Part A, 27(4):627–637, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196900862.

Anonymous:1996:C

[Ano96b] Anonymous. Contents. Studies in History and Philosophy of Science Part A, 27(4):??, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368196900886.

Anonymous:1996:EB

[Ano96c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 27(1):??, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368196901107.

Anonymous:1997:BHPa

[Ano97a] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 28(2):387– 391, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197900091.

Anonymous:1997:BHPb

[Ano97b] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 28(4):??, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197900133.

Anonymous:1997:Ca

[Ano97c] Anonymous. Corrigendum. Studies in History and Philosophy of Science Part A, 28(1):??, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197882131.

Anonymous:1997:Cb

[Ano97d] Anonymous. Corrigendum. Studies in History and Philosophy of Science Part A, 28(2):??, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197827148.

Anonymous:1997:EB

[Ano97e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 28(1):??, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197900212.

Anonymous:1997:SWA

[Ano97f] Anonymous. Science wars: Apology. Studies in History and Philosophy of Science Part A, 28(4):??, December 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197872883.

Anonymous:1997:SHP

[Ano97g] Anonymous. Studies in History and Philosophy of Science Part C: Studies in History and Philosophy of Biological and Biomedical Sciences. Studies in History and Philosophy of Science Part A, 28(4):??, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368197900157.

Anonymous:1998:Aa

[Ano98a] Anonymous. Abstract. Studies in History and Philosophy of Science Part A, 29(2):299–303, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000174.

Anonymous:1998:Ab

[Ano98b] Anonymous. Abstract. Studies in History and Philosophy of Science Part A, 29(2):319–325, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000253.

Anonymous:1998:BRa

[Ano98c] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(2):305–311, June 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000277.

Anonymous:1998:BRb

[Ano98d] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(2):313–318, June 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000241.

Anonymous:1998:BRc

[Ano98e] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(3):449–457, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819800017X.

Anonymous:1998:BRd

[Ano98f] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(3):459–463, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000168.

Anonymous:1998:BRe

[Ano98g] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(3):465–479, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000065.

Anonymous:1998:BRf

[Ano98h] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(3):481–489, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000077.

Anonymous:1998:BRg

[Ano98i] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(3):491–499, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000089.

Anonymous:1998:BRh

[Ano98j] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(4):653–661, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000181.

Anonymous:1998:BRi

[Ano98k] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(4):663–672, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000193.

Anonymous:1998:BRj

[Ano98l] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(4):673–679, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819800020X.

Anonymous:1998:BRk

[Ano98m] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 29(4):681–687, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000260.

Anonymous:1998:Da

[Ano98n] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 29(4):623–637, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000302.

Anonymous:1998:Db

[Ano98o] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 29(4):639–652, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000314.

Anonymous:1998:EB

[Ano98p] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 29(1):??, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198900036.

Anonymous:1998:I

[Ano98q] Anonymous. Index. Studies in History and Philosophy of Science Part A, 29(4):??, December 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368198000478.

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(1):173–181, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000247.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(1):183–189, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000259.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(2):355–361, June 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000442.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(2):363–366, June 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000430.

Anonymous:1999:BRe

[Ano99e] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(2):367–375, June 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000454.

Anonymous:1999:BRf

[Ano99f] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(3):493–499, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000163.

Anonymous:1999:BRg

[Ano99g] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(3):501–510, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000175.

Anonymous:1999:BRh

[Ano99h] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(3):511–521, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000187.

Anonymous:1999:BRi

[Ano99i] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(3):523–530, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000466.

Anonymous:1999:BRj

[Ano99j] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(4):729–744, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000242.

Anonymous:1999:BRk

[Ano99k] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 30(4):745–749, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000254.

Anonymous:1999:Da

[Ano99l] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):113–129, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000399.

Anonymous:1999:Db

[Ano99m] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):131–136, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000405.

Anonymous:1999:Dc

[Ano99n] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):139–147, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000338.

Anonymous:1999:Dd

[Ano99o] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):149–156, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819800034X.

Anonymous:1999:De

[Ano99p] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):157–161, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000351.

Anonymous:1999:Df

[Ano99q] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):163–166, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000363.

Anonymous:1999:Dg

[Ano99r] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(1):167–171, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000375.

Anonymous:1999:Dh

[Ano99s] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(3):455–477, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000151.

Anonymous:1999:Di

[Ano99t] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):531–557, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000278.

Anonymous:1999:Dj

[Ano99u] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):651–685, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000291.

Anonymous:1999:Dk

[Ano99v] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):687–697, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000205.

Anonymous:1999:Dl

[Ano99w] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):699–720, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000217.

Anonymous:1999:Dm

[Ano99x] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):721–723, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000229.

Anonymous:1999:Dn

[Ano99y] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 30(4):725–728, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000230.

Anonymous:1999:I

[Ano99z] Anonymous. Index. Studies in History and Philosophy of Science Part A, 30(4):vii, December 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368199000369.

Anonymous:2000:BR

[Ano00a] Anonymous. Book review. Studies in History and Philosophy of Science Part A, 31(1):173–188, March 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000357.

Anonymous:2000:C

[Ano00b] Anonymous. Corrigendum. Studies in History and Philosophy of Science Part A, 31(3):??, September 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000340.

Anonymous:2000:Da

[Ano00c] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 31(1):71–86, March 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819900028X.

Anonymous:2000:Db

[Ano00d] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 31(1):87–124, March 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000382.

Anonymous:2000:Dc

[Ano00e] Anonymous. Discussion. Studies in History and Philosophy of Science Part A, 31(1):137–149, March 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000394.

Anonymous:2000:I

[Ano00f] Anonymous. Index. Studies in History and Philosophy of Science Part A, 31(4):??, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368100000364.

Anonymous:2001:BHP

[Ano01a] Anonymous. Books on history and philosophy of science received. Studies in History and Philosophy of Science Part A, 32(2):397– 399, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000024.

Anonymous:2001:IRG

[Ano01b] Anonymous. “The Initial Response to Galileo’s Lunar Observations” by R. Ariew. Studies in History and Philosophy of Science 32(3) pp. 571–581. Studies in History and Philosophy of Science Part A, 32(4):749, December 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000449. See [Ari01].

Anonymous:2001:CAI

[Ano01c] Anonymous. Contents and author index. Studies in History and Philosophy of Science Part A, 32(4):??, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000425.

Anonymous:2001:GBF

[Ano01d] Anonymous. Gerd Buchdahl (1914–2001): Founding Editor. Studies in History and Philosophy of Science Part A, 32(3):401–405, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000255.

Anonymous:2002:CAI

[Ano02a] Anonymous. 2002 contents and author index. Studies in History and Philosophy of Science Part A, 33(4):??, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000705.

Anonymous:2002:BHP

[Ano02b] Anonymous. Books in the history and philosophy of science. Studies in History and Philosophy of Science Part A, 33(4):811– 814, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000213.

Anonymous:2002:EBa

[Ano02c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 33(3):??, September 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000365.

Anonymous:2002:EBb

[Ano02d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 33(4):??, December 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810200064X.

Anonymous:2003:BHP

[Ano03a] Anonymous. Books in the history and philosophy of science. Studies in History and Philosophy of Science Part A, 34(4):829– 832, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000700.

Anonymous:2003:EBa

[Ano03b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 34(1):??, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000037.

Anonymous:2003:EBb

[Ano03c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 34(2):??, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000323.

Anonymous:2003:EBc

[Ano03d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 34(3):??, September 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000554.

Anonymous:2003:EBd

[Ano03e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 34(4):??, December 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000724.

Anonymous:2003:GBW

[Ano03f] Anonymous. Gerd Buchdahl’s writings in history and philosophy of science: a listing of publications, unpublished works, and annotated books. Studies in History and Philosophy of Science Part A, 34(1):209–227, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368102000936.

Anonymous:2003:VC

[Ano03g] Anonymous. Volume contents. Studies in History and Philosophy of Science Part A, 34(4):??, December 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000803.

Anonymous:2004:CAI

[Ano04a] Anonymous. 2004 contents and author index. Studies in History and Philosophy of Science Part A, 35(4):??, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810400113X.

Anonymous:2004:BRJ

[Ano04b] Anonymous. Books received to July 2004. Studies in History and Philosophy of Science Part A, 35(4):891–895, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000822.

Anonymous:2004:EBa

[Ano04c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 35(1):??, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000032.

Anonymous:2004:EBb

[Ano04d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 35(2):??, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000378.

Anonymous:2004:EBc

[Ano04e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 35(3):??, September 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000652.

Anonymous:2004:EBd

[Ano04f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 35(4):??, December 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104001049.

Anonymous:2004:E

[Ano04g] Anonymous. Errata. Studies in History and Philosophy of Science Part A, 35(2):413, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000251.

Anonymous:2005:CAI

[Ano05a] Anonymous. 2005 contents and author index. Studies in History and Philosophy of Science Part A, 36(4):??, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000981.

Anonymous:2005:C

[Ano05b] Anonymous. Contents. Studies in History and Philosophy of Science Part A, 36(4):721, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000932.

Anonymous:2005:EBa

[Ano05c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 36(1):??, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000038.

Anonymous:2005:EBb

[Ano05d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 36(2):??, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000269.

Anonymous:2005:EBc

[Ano05e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 36(4):??, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000890.

Anonymous:2005:EBP

[Ano05f] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 36(3):??, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000580.

Anonymous:2006:CAI

[Ano06a] Anonymous. 2006 contents and author index. Studies in History and Philosophy of Science Part A, 37(4):??, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000902.

Anonymous:2006:EB

[Ano06b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 37(3):??, September 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000550.

Anonymous:2006:EBPa

[Ano06c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 37(1):??, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000033.

Anonymous:2006:EBPb

[Ano06d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 37(2):??, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000355.

Anonymous:2006:EBPc

[Ano06e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 37(4):??, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000847.

Anonymous:2007:BRO

[Ano07a] Anonymous. Books received to October 2006. Studies in History and Philosophy of Science Part A, 38(1):283–287, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106001051.

Anonymous:2007:EBPa

[Ano07b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 38(1):??, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000039.

Anonymous:2007:EBPb

[Ano07c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 38(2):??, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000234.

Anonymous:2007:EBPc

[Ano07d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 38(3):??, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000489.

Anonymous:2007:EBPd

[Ano07e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 38(4):??, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000726.

Anonymous:2008:EBPa

[Ano08a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 39(1):??, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000046.

Anonymous:2008:EBPb

[Ano08b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 39(2):??, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000381.

Anonymous:2008:EBPc

[Ano08c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 39(3):??, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000691.

Anonymous:2008:EBPd

[Ano08d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 39(4):??, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000939.

Anonymous:2009:BRM

[Ano09a] Anonymous. Books received to March 2009. Studies in History and Philosophy of Science Part A, 40(3):332–335, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000545.

Anonymous:2009:EBPa

[Ano09b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 40(1):??, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810900003X.

Anonymous:2009:EBPb

[Ano09c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 40(2):??, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000338.

Anonymous:2009:EBPc

[Ano09d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 40(3):??, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000569.

Anonymous:2009:EBPd

[Ano09e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 40(4):??, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000764.

Anonymous:2009:JUT

[Ano09f] Anonymous. Journals under threat: a joint response from history of science, technology and medicine editors. Studies in History and Philosophy of Science Part A, 40(1):1–3, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108001064.

Anonymous:2010:EBPa

[Ano10a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 41(1):??, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000038.

Anonymous:2010:EBPb

[Ano10b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 41(2):??, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000257.

Anonymous:2010:EBPc

[Ano10c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 41(3):??, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000476.

Anonymous:2010:EBPd

[Ano10d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 41(4):??, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000683.

Anonymous:2011:BRM

[Ano11a] Anonymous. Books received to March 2011. Studies in History and Philosophy of Science Part A, 42(3):477–478, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000379.

Anonymous:2011:EBPa

[Ano11b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 42(1):??, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000033.

Anonymous:2011:EBPb

[Ano11c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 42(2):??, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000264.

Anonymous:2011:EBPc

[Ano11d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 42(3):??, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000501.

Anonymous:2011:EBPd

[Ano11e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 42(4):??, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000902.

Anonymous:2012:BRM

[Ano12a] Anonymous. Books received to March 2012. Studies in History and Philosophy of Science Part A, 43(3):430–431, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000325.

Anonymous:2012:EBPa

[Ano12b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 43(1):??, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000052.

Anonymous:2012:EBPb

[Ano12c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 43(2):??, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000246.

Anonymous:2012:EBPc

[Ano12d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 43(3):??, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811200057X.

Anonymous:2012:EBPd

[Ano12e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 43(4):??, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000921.

Anonymous:2013:EBPa

[Ano13a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 44(1):??, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000034.

Anonymous:2013:EBPb

[Ano13b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 44(2):??, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000216.

Anonymous:2014:BRT

[Ano14a] Anonymous. Books received till March 2014. Studies in History and Philosophy of Science Part A, 47(??):133–134, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000272.

Anonymous:2014:EBPa

[Ano14b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 46(??):ifc, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000387.

Anonymous:2014:EBPb

[Ano14c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 47(??):ifc, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811400065X.

Anonymous:2014:EBPc

[Ano14d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 48 (part A)(??): ifc, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001063.

Anonymous:2014:EBPd

[Ano14e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 48(??):ifc, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000983.

Anonymous:2015:EBPa

[Ano15a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 49(??):ifc, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000059.

Anonymous:2015:EBPb

[Ano15b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 50(??):ifc, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000291.

Anonymous:2015:EBPc

[Ano15c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 51(??):ifc, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000485.

Anonymous:2015:EBPd

[Ano15d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 52(??):ifc, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000783.

Anonymous:2015:EBPe

[Ano15e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 53(??):ifc, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001120.

Anonymous:2015:EBPf

[Ano15f] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 54(??):ifc, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001466.

Anonymous:2016:EBPa

[Ano16a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 55(??):ifc, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001818.

Anonymous:2016:EBPb

[Ano16b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 56(??):ifc, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811600011X.

Anonymous:2016:EBPc

[Ano16c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 57(??):ifc, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811630005X.

Anonymous:2016:EBPd

[Ano16d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 58(??):ifc, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300292.

Anonymous:2016:EBPe

[Ano16e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 59(??):ifc, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811630053X.

Anonymous:2016:EBPf

[Ano16f] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 60(??):ifc, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300917.

Anonymous:2017:EBPa

[Ano17a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Science Part A, 61(??):ifc, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300481.

Anonymous:2017:EBPb

[Ano17b] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Science Part A, 62(??):ifc, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811730095X.

Anonymous:2017:EBPc

[Ano17c] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Science Part A, 63(??):ifc, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301279.

Anonymous:2017:EBPd

[Ano17d] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Science Part A, 64(??):ifc, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117302200.

Anonymous:2017:EBP

[Ano17e] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Science Part A, 65–66(??):ifc, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302844.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial board. Studies in History and Philosophy of Science Part A, 67(??):ii, February 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300359.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial board. Studies in History and Philosophy of Science Part A, 68(??):ii, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300876.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial board. Studies in History and Philosophy of Science Part A, 69(??):ii, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301262.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 70(??):??, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302048.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 71(??):ii, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302516.

Anonymous:2018:EB

[Ano18f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 72(??):ii, December 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302991.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 73(??):ii, February 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119300445.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 74(??):ii, ???? 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119301116.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 75(??):ii, June 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368119301669.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 76(??):ii, August 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119302304.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 77(??):ii, October 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119303127.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 78(??):ii, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119303504.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 79(??):ii, February 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120300170.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 80(??):ii, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120300777.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 81(??):ii, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301230.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 82(??):ii, August 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301552.

Anonymous:2020:EBe

[Ano20e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 83(??):ii, October 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301709.

Anonymous:2020:EBf

[Ano20f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 84(??):ii, December 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301989.

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 85(??):ii, February 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812100011X.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 86(??):ii, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000261.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 87(??):ii, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000650.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 88(??):ii, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001163.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 89(??):ii, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001503.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 90(??):ii, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001977.

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 91(??):ii, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000188.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 92(??):ii, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000516.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 93(??):ii, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000760.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 94(??):ii, August 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001054.

Anonymous:2022:EBe

[Ano22e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 95(??):ii, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001431.

Anonymous:2022:EBf

[Ano22f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 96(??):ii, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001674.

Anonymous:2022:PD

[Ano22g] Anonymous. Pages 1–196 (December 2022). Studies in History and Philosophy of Science Part A, 96(??):1–196, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2022:PAb

[Ano22h] Anonymous. Pages 1–212 (August 2022). Studies in History and Philosophy of Science Part A, 94(??):1–212, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2022:PJ

[Ano22i] Anonymous. Pages 1–230 (June 2022). Studies in History and Philosophy of Science Part A, 93(??):1–230, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2022:PO

[Ano22j] Anonymous. Pages 1–236 (October 2022). Studies in History and Philosophy of Science Part A, 95(??):1–236, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2022:PAa

[Ano22k] Anonymous. Pages 1–274 (April 2022). Studies in History and Philosophy of Science Part A, 92(??):1–274, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 97(??):ii, February 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000389.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 98(??):iv, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812300064X.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 99(??):ii, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812300078X.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 100(??):ii, August 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000973.

Anonymous:2023:EBe

[Ano23e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 101(??):ii, October 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001322.

Anonymous:2023:EBf

[Ano23f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 102(??):ii, December 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001553.

Anonymous:2023:PAAc

[Ano23g] Anonymous. Pages 1–106, A1–A14 (June 2023). Studies in History and Philosophy of Science Part A, 99(??):1–106, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:PAb

[Ano23h] Anonymous. Pages 1–116 (August 2023). Studies in History and Philosophy of Science Part A, 100(??):1–116, August 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:PO

[Ano23i] Anonymous. Pages 1–70 (October 2023). Studies in History and Philosophy of Science Part A, 101(??):1–70, October 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:PAa

[Ano23j] Anonymous. Pages 1–80 (April 2023). Studies in History and Philosophy of Science Part A, 98(??):1–80, April 2023. CODEN SH-PSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:PD

[Ano23k] Anonymous. Pages 1–90 (December 2023). Studies in History and Philosophy of Science Part A, 102(??):1–90, December 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2023:PAAa

[Ano23l] Anonymous. Pages A1–A2, 1–144 (February 2023). Studies in History and Philosophy of Science Part A, 97(??):A1–, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 103(??):ii, February 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000074.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 104(??):ii, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000360.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 105(??):ii, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000955.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 106(??):ii, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812400116X.

Anonymous:2024:EBe

[Ano24e] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 107(??):ii, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001286.

Anonymous:2024:EBf

[Ano24f] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 108(??):ii, December 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001535.

Anonymous:2024:PD

[Ano24g] Anonymous. Pages 1–100 (December 2024). Studies in History and Philosophy of Science Part A, 108(??):1–100, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:PO

[Ano24h] Anonymous. Pages 1–128 (October 2024). Studies in History and Philosophy of Science Part A, 107(??):1–128, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:PAa

[Ano24i] Anonymous. Pages 1–160 (April 2024). Studies in History and Philosophy of Science Part A, 104(??):1–160, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:PJ

[Ano24j] Anonymous. Pages 1–174 (June 2024). Studies in History and Philosophy of Science Part A, 105(??):1–174, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:PF

[Ano24k] Anonymous. Pages 1–178 (February 2024). Studies in History and Philosophy of Science Part A, 103(??):1–178, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2024:PAb

[Ano24l] Anonymous. Pages 1–208 (August 2024). Studies in History and Philosophy of Science Part A, 106(??):1–208, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2025:EBa

[Ano25a] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 109(??):ii, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368125000081.

Anonymous:2025:EBb

[Ano25b] Anonymous. Editorial Board. Studies in History and Philosophy of Science Part A, 110(??):ii, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368125000391.

Anonymous:2025:PA

[Ano25c] Anonymous. Pages 1–104 (April 2025). Studies in History and Philosophy of Science Part A, 110(??):1–104, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anonymous:2025:PF

[Ano25d] Anonymous. Pages 1–146 (February 2025). Studies in History and Philosophy of Science Part A, 109(??):1–146, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

Anstey:2002:RBH

[Ans02] Peter R. Anstey. Robert Boyle and the heuristic value of mechanism. Studies in History and Philosophy of Science Part A, 33(1): 157–170, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000346.

Anstey:2004:MON

[Ans04] Peter R. Anstey. The methodological origins of Newton’s queries. Studies in History and Philosophy of Science Part A, 35(2):247– 269, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000147.

Anstey:2016:LM

[Ans16] Peter R. Anstey. Locke on measurement. Studies in History and Philosophy of Science Part A, 60(??):70–81, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300826.

Antognazza:2003:LPC

[Ant03] Maria Rosa Antognazza. Leibniz and the post-Copernican universe. Koyré revisited. Studies in History and Philosophy of Science Part A, 34(2):309–327, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000219.

Anta:2025:IIO

[Ant25] Javier Anta. Intellectual inflation: one way for scientific research to degenerate. Studies in History and Philosophy of Science Part A, 109(??):134–145, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368125000032.

Anzola:2019:KTA

[Anz19] David Anzola. Knowledge transfer in agent-based computational social science. Studies in History and Philosophy of Science Part A, 77(??):29–38, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118301195.

Atkinson:2004:GPP

[AP04] David Atkinson and Jeanne Peijnenburg. Galileo and prior philosophy. Studies in History and Philosophy of Science Part A, 35(1):115–136, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810300092X.

Atkinson:2006:PCF

[AP06] David Atkinson and Jeanne Peijnenburg. Probability without certainty: foundationalism and the Lewis–Reichenbach debate. Studies in History and Philosophy of Science Part A, 37(3):442–453, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000501.

Abi-Rached:2022:AHM

[AR22] Joelle M. Abi-Rached. ‘As.fūriyyeh: a History of Madness, Modernity, and War in the Middle East: Reply by the author Joelle M. Abi-Rached. Studies in History and Philosophy of Science Part A, 93(??):227–229, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000085. See [Edi22,

Rob22].

Arabatzis:1992:DZE

[Ara92] Theodore Arabatzis. The discovery of the Zeeman effect: a case study of the interplay between theory and experiment. Studies in History and Philosophy of Science Part A, 23(3):365–388, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819290001M.

Arabatzis:2003:BSO

[Ara03] Theodore Arabatzis. Biographies of scientific objects. Studies in History and Philosophy of Science Part A, 34(2):431– 442, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000281.

Arapostathis:2013:MPE

[Ara13] Stathis Arapostathis. Meters, patents and expertise(s): Knowledge networks in the electricity meters industry, 1880–1914. Studies in History and Philosophy of Science Part A, 44(2):234– 246, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811200101X.

Arabatzis:2023:BRI

[Ara23] Theodore Arabatzis. Book review: The instrument of science: Scientific anti-realism revitalised by Darrell Rowbottom: Cognitive instrumentalism and the history of science. Studies in History and Philosophy of Science Part A, 98(??):1–3, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000262.

Ariew:1984:GLO

[Ari84] Roger Ariew. Galileo’s lunar observations in the context of medieval lunar theory. Studies in History and Philosophy of Science Part A, 15(3):213–226, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368184900177.

Ariew:1987:PVB

[Ari87] Roger Ariew. The phases of Venus before 1610. Studies in History and Philosophy of Science Part A, 18(1):81–92, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900124.

Ariew:2001:IRG

[Ari01] Roger Ariew. The initial response to Galileo’s lunar observations. Studies in History and Philosophy of Science Part A, 32(3):571– 581, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000164. See [Ano01b].

Ariew:2022:CDS

[Ari22] André Ariew. Charles Darwin as a statistical thinker. Studies in History and Philosophy of Science Part A, 95(??):215– 223, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001194.

Aronson:1971:LHA

[Aro71] Jerrold Aronson. The legacy of Hume’s analysis of causation. Studies in History and Philosophy of Science Part A, 2(2):135– 156, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900288.

Aronova:2018:EPB

[Aro18] Elena Aronova. Earthquake prediction, biological clocks, and the cold war psy-ops: Using animals as seismic sensors in the 1970s California. Studies in History and Philosophy of Science Part A, 70 (??):38–49, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301407.

Arthur:1995:NFE

[Art95] Richard T. W. Arthur. Newton’s fluxions and equably flowing time. Studies in History and Philosophy of Science Part A, 26(2): 323–351, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819400037A.

Andreas:2016:CSA

[AS16] Holger Andreas and Georg Schiemer. A choice-semantical approach to theoretical truth. Studies in History and Philosophy of Science Part A, 58(??):1–8, August 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116000078.

Asper:2013:EBN

[Asp13] Markus Asper. Explanation between nature and text: Ancient Greek commentators on science. Studies in History and Philosophy of Science Part A, 44(1):43–50, March 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000775.

Aston:2004:PPC

[Ast04] Nigel Aston. From personality to party: the creation and transmission of Hutchinsonianism, c. 1725–1750. Studies in History and Philosophy of Science Part A, 35(3):625–644, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000585.

Atkinson:1994:RSP

[Atk94] Malcolm Atkinson. Regulation of science by ‘peer review’. Studies in History and Philosophy of Science Part A, 25(2):147– 158, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900256.

Atran:1985:PTA

[Atr85] Scott Atran. Pre-theoretical aspects of Aristotelian definition and classification of animals: The case for common sense. Studies in History and Philosophy of Science Part A, 16(2):113– 163, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818590010X.

Aubin:2008:MLI

[Aub08] David Aubin. ‘the memory of life itself’: Bénard’s cells and the cinematography of self-organization. Studies in History and Philosophy of Science Part A, 39(3):359–369, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000605.

Austin:1978:SC

[Aus78] James Austin. Systemic causation. Studies in History and Philosophy of Science Part A, 9(2):83–97, June 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900018.

Avin:2019:ML

[Avi19] Shahar Avin. Mavericks and lotteries. Studies in History and Philosophy of Science Part A, 76(??):13–23, August 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300190.

Axtell:1993:THM

[Axt93] Guy S. Axtell. In the tracks of the historicist movement: Re-assessing the Carnap–Kuhn connection. Studies in History and Philosophy of Science Part A, 24(1):119–146, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390027H.

Ayalon:2025:PCH

[Aya25] Noa Lahav Ayalon. Philosophical community from a historical perspective. Studies in History and Philosophy of Science Part A, 110(??):40–45, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368125000123.

Aylward:2019:ADT

[Ayl19] Alex Aylward. Against defaultism and towards localism in the contingency/inevitability conversation: Or, why we should shut up about putting-up. Studies in History and Philosophy of Science Part A, 74(??):30–41, April 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368117303291.

Bechtel:2010:DME

[BA10] William Bechtel and Adele Abrahamsen. Dynamic mechanistic explanation: computational modeling of circadian rhythms as an exemplar for cognitive science. Studies in History and Philosophy of Science Part A, 41(3):321–333, September 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811000035X.

Babes:2025:MMM

[Bab25] Ovidiu Babes. Mixed mathematics and metaphysical physics: Descartes and the mechanics of the flow of water. Studies in History and Philosophy of Science Part A, 109(??):58–71, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001584.

Baccelli:2020:BMV

[Bac20] Jean Baccelli. Beyond the metrological viewpoint. Studies in History and Philosophy of Science Part A, 80(??):56–61, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301985.

Baetu:2024:EAC

[Bae24] Tudor M. Baetu. Extrapolating animal consciousness. Studies in History and Philosophy of Science Part A, 104(??):150– 159, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000256.

Baigrie:1987:KLP

[Bai87] Brian S. Baigrie. Kepler’s laws of planetary motion, before and after Newton’s Principia: An essay on the transformation of scientific problems. Studies in History and Philosophy of Science Part A, 18 (2):177–208, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368187900173. See [New87, NCW99,

New09].

Baigrie:1990:JKE

[Bai90] Brian S. Baigrie. The justification of Kepler’s ellipse. Studies in History and Philosophy of Science Part A, 21(4):633–664, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900368. See reply [Lug92] and

response [Bai92].

Baigrie:1992:GVF

[Bai92] Brian S. Baigrie. Generativist versus foundational justification: a reply to Andrew Lugg. Studies in History and Philosophy of Science Part A, 23(3):503–508, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290007S.

See [Bai90, Lug92].

Baillon:2004:EEC

[Bai04] Jean-François Baillon. Early eighteenth-century Newtonianism: the Huguenot contribution. Studies in History and Philosophy of Science Part A, 35(3):533–548, September 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000548.

Baker:2006:TNA

[Bak06] Lynne Rudder Baker. On the twofold nature of artefacts. Studies in History and Philosophy of Science Part A, 37(1):132– 136, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105001172.

Baker:2016:BRS

[Bak16] Tawrin Baker. Book review: From sight to light: the passage from ancient to modern optics, A. Mark Smith. University of Chicago Press, USA (2015). Studies in History and Philosophy of Science Part A, 59(??):115–120, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116000212.

Baker:2020:SI

[Bak20] Alan Baker. Schemas for induction. Studies in History and Philosophy of Science Part A, 82(??):114–119, August 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119301232.

Baker:2021:CIM

[Bak21] Alan Baker. Circularity, indispensability, and mathematical explanation in science. Studies in History and Philosophy of Science Part A, 88(??):156–163, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000820.

Baker:2022:PEP

[Bak22] Erik Baker. From planning to entrepreneurship: On the political economy of scientific pursuit. Studies in History and Philosophy of Science Part A, 92(??):27–35, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000139.

Bakshi:2023:CSM

[Bak23] Kabir S. Bakshi. Clarifying some misconceptions in interpreting Ernst Mach’s views on thought experiments. Studies in History and Philosophy of Science Part A, 97(??):58–67, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001807.

Baldini:1992:LIS

[Bal92] Ugo Baldini. Legem impone subactis: studi su filosofia e scienza dei Gesuiti in Italia, 1540–1632. (Latin and Italian) [Lay after the conquest of the Law: studies of philosophy and science of the Jesuits in Italy], volume 3 of Collana dell’Istituto di filosofia (Università G. D’Annunzio di Chieti). Nuova serie. Bulzoni, Roma, Italy, 1992. ISBN 88-7119-503-5. 601 pp. LCCN BX3737 .B35 1992 (LC).

Balaban:2020:GSV

[Bal20] Oded Balaban. Genera and species vs. laws of nature two epistemic frameworks and their respective ideal worlds. Studies in History and Philosophy of Science Part A, 81(??):6–15, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300748.

Bamford:1996:PHC

[Bam96] Greg Bamford. Popper and his commentators on the discovery of Neptune: a close shave for the Law of Gravitation? Studies in History and Philosophy of Science Part A, 27(2):207– 232, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000453.

Murad:2011:MSR

[bAM11] Mohd Hazim Shah bin Abdul Murad. Models, scientific realism, the intelligibility of nature, and their cultural significance. Studies in History and Philosophy of Science Part A, 42(2):253– 261, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110001123.

Banks:2013:EMC

[Ban13] Erik C. Banks. Extension and measurement: a constructivist program from Leibniz to Grassmann. Studies in History and Philosophy of Science Part A, 44(1):20–31, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000696.

Barker:1980:HW

[Bar80] Peter Barker. Hertz and Wittgenstein. Studies in History and Philosophy of Science Part A, 11(3):243–256, September 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818090028X.

Barnes:2005:MPC

[Bar05] Eric Barnes. On Mendeleev’s predictions: comment on Scerri and Worrall. Studies in History and Philosophy of Science Part A, 36 (4):801–812, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000695. See [SW01,

Sce05, Wor05, Sch08b].

Barker:2009:PMH

[Bar09] Andrew Barker. Ptolemy and the meta-helikôn. Studies in History and Philosophy of Science Part A, 40(4):344–351, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000673.

Barnes:2014:RP

[Bar14] Eric Christian Barnes. The roots of predictivism. Studies in History and Philosophy of Science Part A, 45(??):46–53, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000988.

Barwich:2018:HRA

[Bar18] Ann-Sophie Barwich. How to be rational about empirical success in ongoing science: The case of the quantum nose and its critics. Studies in History and Philosophy of Science Part A, 69(??): 40–51, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301632.

Barker:2020:ERW

[Bar20a] Peter Barker. Essay review, Wootton and Wittgenstein. Studies in History and Philosophy of Science Part A, 84(??):95–98, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301643.

Bartha:2020:NMT

[Bar20b] Paul Bartha. Norton’s material theory of analogy. Studies in History and Philosophy of Science Part A, 82(??):104–113, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119301347.

Barrett:2021:SOB

[Bar21a] Jeffrey A. Barrett. Situated observation in Bohmian mechanics. Studies in History and Philosophy of Science Part A, 88(??):345– 357, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000935.

Barrett:2021:CA

[Bar21b] Thomas William Barrett. The curvature argument. Studies in History and Philosophy of Science Part A, 88(??):30–40, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000522.

Barry:2021:SRB

[Bar21c] Galen Barry. Spinoza on the resistance of bodies. Studies in History and Philosophy of Science Part A, 86(??):56–67, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000145.

Barseghyan:2022:QPE

[Bar22a] Hakob Barseghyan. Question pursuit as an epistemic stance. Studies in History and Philosophy of Science Part A, 94(??):112– 120, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000826.

Barseghyan:2022:SPP

[Bar22b] Hakob Barseghyan. Selection, presentism, and pluralist history. Studies in History and Philosophy of Science Part A, 92(??): 60–70, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000036.

Barton:2022:SRJ

[Bar22c] Ruth Barton. The scientific reputation(s) of John Lubbock, Darwinian gentleman. Studies in History and Philosophy of Science Part A, 95(??):185–203, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000942.

Basu:1992:SDB

[Bas92] Prajit K. Basu. Similarities and dissimilarities between Joseph Priestley’s and Antoine Lavoisier’s chemical beliefs. Studies in History and Philosophy of Science Part A, 23(3):445–469, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290004P.

Basu:2003:TLE

[Bas03] Prajit K. Basu. Theory-ladenness of evidence: a case study from history of chemistry. Studies in History and Philosophy of Science Part A, 34(2):351–368, June 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000220.

Basso:2017:ARM

[Bas17] Alessandra Basso. The appeal to robustness in measurement practice. Studies in History and Philosophy of Science Part A, 65–66(??):57–66, October/December 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117300596.

Batterman:2023:BRM

[Bat23] Robert W. Batterman. Book review: A Middle Way: a Non-Fundamental Approach to Many-Body Physics by Robert W. Batterman: Reply by the Author. Studies in History and Philosophy of Science Part A, 97(??):130–132, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001315. See [Dar23, FR23, Mil23,

Pal23, Woo24].

Baumer:1971:BRB

[Bau71] William H. Baumer. Book review: Science and civilization in Islam: Seyyed Nasr, Harvard University Press: Cambridge, Mass., 1968. xiv + 384 pp. $8.95. Studies in History and Philosophy of Science Part A, 2(2):183–190, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817190032X.

See [Nas68, Nas83].

Baum:1972:IFE

[Bau72] Robert J. Baum. The instrumentalist and formalist elements of Berkeley’s philosophy of mathematics. Studies in History and Philosophy of Science Part A, 3(2):119–134, August 1972. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368172900192.

Baumslag:1998:CSG

[Bau98] David Baumslag. Choosing scientific goals: The need for a normative approach. Studies in History and Philosophy of Science Part A, 29(1):81–96, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368197000113.

Baum:2015:NDN

[Bau15] Emily Baum. Neither donkey nor horse. Studies in History and Philosophy of Science Part A, 49(??):108–111, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114001137.

Bazerman:1988:SWK

[Baz88] Charles Bazerman. Shaping written knowledge: the genre and activity of the experimental article in science. Rhetoric of the human sciences. University of Wisconsin Press, Madison, WI, USA, 1988. ISBN 0-299-11690-5, 0-299-11694-8 (paperback). xi + 356 pp. LCCN T11 .B375 1988. US$40.00; US$17.50.

Bich:2022:ONO

[BB22a] Leonardo Bich and William Bechtel. Organization needs organization: Understanding integrated control in living organisms. Studies in History and Philosophy of Science Part A, 93(??):96– 106, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000565.

Bollhagen:2022:DAD

[BB22b] Andrew Bollhagen and William Bechtel. Discovering autoinhibition as a design principle for the control of biological mechanisms. Studies in History and Philosophy of Science Part A, 95(??):145– 157, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001224.

Baedke:2023:WOM

[BB23] Jan Baedke and Tatjana Buklijas. Where organisms meet the environment: Introduction to the special issue ‘What counts as environment in biology and medicine: Historical, philosophical and sociological perspectives’. Studies in History and Philosophy of Science Part A, 99(??):A4–A9, June 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001467.

Ben-Chaim:1992:EEP

[BC92] Michael Ben-Chaim. The empiric experience and the practice of autonomy. Studies in History and Philosophy of Science Part A, 23(4):533–555, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819290012U.

Ben-Chaim:2000:LIC

[BC00] Michael Ben-Chaim. Locke’s ideology of ‘common sense’. Studies in History and Philosophy of Science Part A, 31(3):473–501, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000157.

Buchwald:2001:INN

[BC01] Jed Z. Buchwald and I. Bernard Cohen. Isaac Newton’s natural philosophy. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02477-2 (hardcover). xx + 354 pp. LCCN Q143.N495 I8 2001.

Benjamin:1987:FLP

[BCC87] Andrew E. Benjamin, G. N. Cantor, and J. R. R. (John R. R.) Christie, editors. The Figural and the literal: problems of language in the history of science and philosophy, 1630–1800. Manchester University Press, Manchester, UK, 1987. ISBN 0-7190-1486-7. LCCN P301 .F54 1987.

Bechtle:2022:BSM

[BCK^{+}22] Philip Bechtle, Cristin Chall, Martin King, Michael Krämer, Peter Mättig, and Michael Stöltzner. Bottoms up: the Standard Model Effective Field Theory from a model perspective. Studies in History and Philosophy of Science Part A, 92(??):129– 143, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000140.

Bishop:2002:TTT

[BD02] Michael A. Bishop and Stephen M. Downes. The theory theory thrice over: the child as scientist, superscientist or social institution? Studies in History and Philosophy of Science Part A, 33(1):117–132, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000292.

Brunet:2021:RPS

[BDB21] Tyler D. P. Brunet, W. Ford Doolittle, and Joseph P. Bielawski. The role of purifying selection in the origin and maintenance of complex function. Studies in History and Philosophy of Science Part A, 87(??):125–135, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000364.

Brusius:2013:WHF

[BDR13] Mirjam Brusius, Katrina Dean, and Chitra Ramalingam, editors. William Henry Fox Talbot: beyond photography, volume 23 of Studies in British art. The Yale Center for British Art, The Paul Mellon Centre for Studies in British Art, New Haven, CT, USA, 2013. ISBN 0-300-17934-0 (hardcover). viii + 308 pp. LCCN TR140.T3 W55 2013.

Beatty:2017:NPN

[Bea17] John Beatty. Narrative possibility and narrative explanation. Studies in History and Philosophy of Science Part A, 62(??): 31–41, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300687.

Beckman:1971:UHE

[Bec71] T. A. Beckman. On the use of historical examples in Agassi’s ‘sensationalism’. Studies in History and Philosophy of Science Part A, 1(4):293–309, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368171900100.

Bechtel:1984:EOU

[Bec84] William Bechtel. The evolution of our understanding of the cell: a study in the dynamics of scientific progress. Studies in History and Philosophy of Science Part A, 15(4):309–356, December 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368184900141.

Bechtel:2016:UCM

[Bec16] William Bechtel. Using computational models to discover and understand mechanisms. Studies in History and Philosophy of Science Part A, 56(??):113–121, April 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001429.

Becker:2023:ALA

[Bec23] Joffrey Becker. Artificial lives, analogies and symbolic thought: an anthropological insight on robots and AI. Studies in History and Philosophy of Science Part A, 99(??):89–96, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000729.

Beiser:2003:HN

[Bei03] Frederick Beiser. Hegel and Naturphilosophie. Studies in History and Philosophy of Science Part A, 34(1):135–147, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000833.

Beiser:2008:HNK

[Bei08] Fred Beiser. Historicism and neo-Kantianism. Studies in History and Philosophy of Science Part A, 39(4):554–564, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000903.

Belsey:1974:IW

[Bel74] Andrew Belsey. Interpreting Whewell. Studies in History and Philosophy of Science Part A, 5(1):49–58, May 1974. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368174900181.

Beller:1990:BPI

[Bel90] Mara Beller. Born’s probabilistic interpretation: a case study of ‘concepts in flux’. Studies in History and Philosophy of Science Part A, 21(4):563–588, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900335.

Beller:1992:BBC

[Bel92] Mara Beller. The birth of Bohr’s complementarity: The context and the dialogues. Studies in History and Philosophy of Science Part A, 23(1):147–180, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368192900296.

Belknap:2015:WHF

[Bel15] Geoffrey Belknap. William Henry Fox Talbot. Beyond Photography: a review. Studies in History and Philosophy of Science Part A, 51(??):105–107, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000035.

Belkind:2020:UAT

[Bel20] Ori Belkind. Unnatural acts: the transition from natural principles to laws of nature in early modern science. Studies in History and Philosophy of Science Part A, 81(??):62–73, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301110.

Benton:1974:VNC

[Ben74] E. Benton. Vitalism in nineteenth-century scientific thought: a typology and reassessment. Studies in History and Philosophy of Science Part A, 5(1):17–48, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817490017X.

Bennett:1987:DCH

[Ben87] J. A. (James A.) Bennett. The divided circle: a history of instruments for astronomy, navigation, and surveying. Phaidon, Christie’s, Oxford, UK, 1987. ISBN 0-7148-8038-8. 224 pp. LCCN QB85.8 .B46 1987.

Benson:2018:RSF

[Ben18] Etienne S. Benson. Re-situating fieldwork and re-narrating disciplinary history in global mega-geomorphology. Studies in History and Philosophy of Science Part A, 70(??):18–27, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301389.

Bentley:2022:PSP

[Ben22] Joseph Bentley. Protocol statements, physicalism, and metadata: Otto Neurath on scientific evidence. Studies in History and Philosophy of Science Part A, 96(??):125–134, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812200139X.

Bentley:2024:PPP

[Ben24] Joseph Bentley. Positivist or post-positivist philosophy of science? The left Vienna Circle and Thomas Kuhn. Studies in History and Philosophy of Science Part A, 107(??):107–117, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001237.

Bernstein:1980:CHY

[Ber80] Howard R. Bernstein. Conatus, Hobbes, and the young Leibniz. Studies in History and Philosophy of Science Part A, 11(1):

25–37, March 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900035.

Bertotti:1985:LWS

[Ber85] Bruno Bertotti. The later work of E. Schrödinger. Studies in History and Philosophy of Science Part A, 16(2):83–100, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900081.

Bernal:1993:PG

[Ber93] Martin Bernal. Paradise glossed. Studies in History and Philosophy of Science Part A, 24(4):669–675, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390057Q.

Barker:1988:RCC

[BG88] Peter Barker and Bernard R. Goldstein. The role of comets in the Copernican revolution. Studies in History and Philosophy of Science Part A, 19(3):299–319, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368188900027.

Buchholz:2023:PEM

[BG23] Oliver Buchholz and Thomas Grote. Predicting and explaining with machine learning models: Social science as a touchstone. Studies in History and Philosophy of Science Part A, 102(??):60– 69, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001425.

Bruner:2019:SCS

[BH19] Justin P. Bruner and Bennett Holman. Self-correction in science: Meta-analysis, bias and social structure. Studies in History and Philosophy of Science Part A, 78(??):93–97, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811830164X.

Bhakthavatsalam:2015:RBP

[Bha15] Sindhuja Bhakthavatsalam. The rationale behind Pierre Duhem’s natural classification. Studies in History and Philosophy of Science Part A, 51(??):11–21, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368115000254.

Bhakthavatsalam:2017:DGS

[Bha17] Sindhuja Bhakthavatsalam. Duhemian good sense and agent reliabilism. Studies in History and Philosophy of Science Part A, 64(??):22–29, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300450.

Bhatta:2024:CAI

[Bha24] Varun S. Bhatta. The controversy about interference of photons. Studies in History and Philosophy of Science Part A, 106(??):146– 154, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812400102X.

Biagioli:1988:MSI

[Bia88] Mario Biagioli. Meyerson: Science and the “irrational”. Studies in History and Philosophy of Science Part A, 19(1):5–42, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900180.

Biagioli:1990:AI

[Bia90] Mario Biagioli. The anthropology of incommensurability. Studies in History and Philosophy of Science Part A, 21(2):183– 209, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819090022Z.

Biagoli:1994:BRB

[Bia94] Mario Biagoli. Book review: Galileo, the Jesuits, and the medieval Aristotle: William A. Wallace, (London: Variorum, 1991), 350 pp. ISBN 0-86078-297-2 Hardback £45.00. Studies in History and Philosophy of Science Part A, 25(4):637–646, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900531. See [Wal91].

Biagioli:2020:ECT

[Bia20] Francesca Biagioli. Ernst Cassirer’s transcendental account of mathematical reasoning. Studies in History and Philosophy of Science Part A, 79(??):30–40, February 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300797.

Biddle:2011:PPW

[Bid11] Justin Biddle. Putting pragmatism to work in the Cold War: Science, technology, and politics in the writings of James B. Conant. Studies in History and Philosophy of Science Part A, 42(4):552– 561, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000628.

Biddle:2013:SFT

[Bid13] Justin Biddle. State of the field: Transient underdetermination and values in science. Studies in History and Philosophy of Science Part A, 44(1):124–133, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811200074X.

Biddle:2014:CPP

[Bid14] Justin B. Biddle. Can patents prohibit research? On the social epistemology of patenting and licensing in science. Studies in History and Philosophy of Science Part A, 45(??):14– 23, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113001301.

Biener:2016:NIE

[Bie16] Zvi Biener. Newton and the ideal of exegetical success. Studies in History and Philosophy of Science Part A, 60(??):82–87, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300371.

Bienias:2020:EGA

[Bie20] Barbara Bienias. Edward Gresham’s Astrostereon, or A Discourse of the Falling of the Planet (1603), the Copernican paradox, and the construction of early modern proto-scientific discourse.

Studies in History and Philosophy of Science Part A, 82(??):44– 56, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119302006.

Bigg:2008:EAV

[Big08] Charlotte Bigg. Evident atoms: visuality in Jean Perrin’s Brownian motion research. Studies in History and Philosophy of Science Part A, 39(3):312–322, September 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000563.

Bigu:2013:SBA

[Bîg13] Dragos Bîgu. A similarity-based approach of Kuhn’s nooverlap principle and anomalies. Studies in History and Philosophy of Science Part A, 44(3):330–338, September 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000320.

Bird:2002:KWT

[Bir02] Alexander Bird. Kuhn’s wrong turning. Studies in History and Philosophy of Science Part A, 33(3):443–463, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000286.

Bird:2004:KNP

[Bir04] Alexander Bird. Kuhn, naturalism, and the positivist legacy. Studies in History and Philosophy of Science Part A, 35(2):337– 356, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000184.

Bird:2008:SPA

[Bir08] Alexander Bird. Scientific progress as accumulation of knowledge: a reply to Rowbottom. Studies in History and Philosophy of Science Part A, 39(2):279–281, June 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000332. See [Row08a].

Bird:2010:EAE

[Bir10] Alexander Bird. Eliminative abduction: examples from medicine. Studies in History and Philosophy of Science Part A, 41(4):345– 352, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000646.

Blum:2022:BQM

[BJ22] Alexander S. Blum and Martin Jähnert. The birth of quantum mechanics from the spirit of radiation theory. Studies in History and Philosophy of Science Part A, 91(??):125–147, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001813.

Brown:2016:IRP

[BK16] Matthew J. Brown and Ian James Kidd. Introduction: Reappraising Paul Feyerabend. Studies in History and Philosophy of Science Part A, 57(??):1–8, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001636.

Boldyrev:2017:CME

[BK17] Ivan Boldyrev and Olessia Kirtchik. The cultures of mathematical economics in the postwar Soviet Union: More than a method, less than a discipline. Studies in History and Philosophy of Science Part A, 63(??):1–10, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117300857.

Bursten:2021:GKE

[BK21] Julia R. S. Bursten and Catherine Kendig. Growing knowledge: Epistemic objects in agricultural extension work. Studies in History and Philosophy of Science Part A, 88(??):85– 91, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000339.

Blackmore:1972:EMH

[Bla72] John Thomas Blackmore. Ernst Mach: his work, life, and influence. University of California Press, Berkeley, CA, USA, 1972. ISBN 0-520-01849-4. xx + 414 pp. LCCN B3303.

Blackwell:1996:ASR

[Bla96] Richard J. Blackwell. Authority in science and in religion. Studies in History and Philosophy of Science Part A, 27(1):143– 148, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000186.

Bland:2013:SRS

[Bla13] Steven Bland. Scepticism, relativism, and the structure of epistemic frameworks. Studies in History and Philosophy of Science Part A, 44(4):539–544, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000599. See response

[San14].

Blakley:2022:SFC

[Bla22] Christopher M. Blakley. Ship fever, confinement, and the racialization of disease. Studies in History and Philosophy of Science Part A, 95(??):96–103, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001091.

Blando:2024:WSM

[Bla24] Francesca Zaffora Blando. From Wald to Schnorr: von Mises’ definition of randomness in the aftermath of Ville’s Theorem. Studies in History and Philosophy of Science Part A, 106(??):196– 207, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001018.

Bloor:1973:WMS

[Blo73] David Bloor. Wittgenstein and Mannheim on the sociology of mathematics. Studies in History and Philosophy of Science Part A, 4(2):173–191, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368173900034.

Bloor:1975:BRB

[Blo75] David Bloor. Book review: The structure of scientific inference: Mary Hesse, London: Macmillan, 1974. viii + 309 pp. £5.95. Studies in History and Philosophy of Science Part A, 5(4):382– 395, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900114. See [Hes74].

Bloor:1982:DMR

[Blo82a] David Bloor. Durkheim and Mauss revisited: Classification and the sociology of knowledge. Studies in History and Philosophy of Science Part A, 13(4):267–297, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900127.

See response [Buc82], comments [Luk82, Blo82c, Hes82], and reply [Blo82b].

Bloor:1982:RGB

[Blo82b] David Bloor. A reply to Gerd Buchdahl. Studies in History and Philosophy of Science Part A, 13(4):305–311, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900140. See [Blo82a, Buc82, Blo82c,

Hes82].

Bloor:1982:RSL

[Blo82c] David Bloor. Reply to steven lukes. Studies in History and Philosophy of Science Part A, 13(4):319–323, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900164. See [Blo82a, Buc82, Blo82b,

Luk82, Hes82].

Bloor:1984:RJW

[Blo84] David Bloor. Reply to J. W. Smith. Studies in History and Philosophy of Science Part A, 15(3):245–249, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368184900207. See [Smi84].

Bloor:1999:AL

[Blo99] David Bloor. Anti-Latour. Studies in History and Philosophy of Science Part A, 30(1):81–112, March 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368198000387. See [Kem05, Feh07, Kem07, Blo07].

Bloor:2007:IMR

[Blo07] David Bloor. Ideals and monisms: recent criticisms of the Strong Programme in the sociology of knowledge. Studies in History and Philosophy of Science Part A, 38(1):210–234, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000987.

Bloor:2008:SCS

[Blo08] David Bloor. Sichtbarmachung, common sense and construction in fluid mechanics: the cases of Hele–Shaw and Ludwig Prandtl. Studies in History and Philosophy of Science Part A, 39(3):349– 358, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000599.

Blum:2021:ESA

[Blu21a] Alexander S. Blum. Erratum to “The state is not abolished, it withers away: How quantum field theory became a theory of scattering” [Studies in History and Philosophy of Modern Physics 60 (2017) 46–80]. Studies in History and Philosophy of Science Part A, 85(??):220, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000029.

Blum:2021:JWD

[Blu21b] Alexander S. Blum. John Wheeler’s Desert Island : the conservatism of non-empirical physics. Studies in History and Philosophy of Science Part A, 90(??):219–225, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001667.

Ben-Menahem:1989:SCS

[BM89] Yemina Ben-Menahem. Struggling with causality: Schrödinger’s case. Studies in History and Philosophy of Science Part A, 20 (3):307–334, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900101.

Bracken:2014:WQB

[BM14] A. J. Bracken and G. F. Melloy. Waiting for the quantum bus: the flow of negative probability. Studies in History and Philosophy of Science Part A, 48 (part A)(??):13–19, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000999.

Bolinska:2020:NHC

[BM20] Agnes Bolinska and Joseph D. Martin. Negotiating history: Contingency, canonicity, and case studies. Studies in History and Philosophy of Science Part A, 80(??):37–46, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119300019.

Bolinska:2021:TCP

[BM21] Agnes Bolinska and Joseph D. Martin. The tragedy of the canon; or, path dependence in the history and philosophy of science. Studies in History and Philosophy of Science Part A, 89(??):63– 73, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001096.

Boumans:2023:DYS

[BM23] Marcel Boumans and Mary S. Morgan. Do you see it this way? Visualising as a tool of sense-making. Studies in History and Philosophy of Science Part A, 101(??):30–39, October 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001085.

Ben-Menahem:2020:RLN

[BMBM20] Hanina Ben-Menahem and Yemima Ben-Menahem. The rule of law: Natural, human, and divine. Studies in History and Philosophy of Science Part A, 81(??):46–54, June 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300803.

Bar-On:2024:MSR

[BO24] Kati Kish Bar-On. Mathematics and society reunited: the social aspects of Brouwer’s intuitionism. Studies in History and Philosophy of Science Part A, 108(??):28–37, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001407.

Boantza:2007:CAI

[Boa07] Victor D. Boantza. Collecting airs and ideas: Priestley’s style of experimental reasoning. Studies in History and Philosophy of Science Part A, 38(3):506–522, September 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000349.

Boantza:2011:ECK

[Boa11] Victor D. Boantza. From experimental to corporate knowledge in early modern science. Studies in History and Philosophy of Science Part A, 42(4):613–617, December 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000665.

Boesch:2022:CER

[Boe22] Brandon Boesch. A concrete example of representational licensing: the Mississippi River Basin Model. Studies in History and Philosophy of Science Part A, 92(??):36–44, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000024.

Bogaard:1975:SCB

[Bog75] Paul A. Bogaard. The status of complex bodies in Epicurean atomism. Studies in History and Philosophy of Science Part A, 6(4):315–329, November 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368175900266.

Bogen:2008:CPA

[Bog08] Jim Bogen. Causally productive activities. Studies in History and Philosophy of Science Part A, 39(1):112–123, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810700091X.

Boge:2020:HIE

[Bog20] Florian J. Boge. How to infer explanations from computer simulations. Studies in History and Philosophy of Science Part A, 82(??):25–33, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301888.

Boge:2021:QRP

[Bog21] Florian J. Boge. Quantum reality: a pragmaticized neo-Kantian approach. Studies in History and Philosophy of Science Part A, 87(??):101–113, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000431.

Bohme:1980:PCT

[Böh80] Gernot Böhme. On the possibility of ‘closed theories’. Studies in History and Philosophy of Science Part A, 11(2):163– 172, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900229.

Bohlin:2000:SUD

[Boh00] Ingemar Bohlin. A social understanding of delegation. Studies in History and Philosophy of Science Part A, 31(4):731–750, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000273.

Bokulich:2015:MHU

[Bok15] Alisa Bokulich. Maxwell, Helmholtz, and the unreasonable effectiveness of the method of physical analogy. Studies in History and Philosophy of Science Part A, 50(??):28–37, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000909.

Bolduc:2014:NBS

[Bol14] Jean-Sébastien Bolduc. Narrow and broad styles of scientific reasoning: a reply to O. Bueno. Studies in History and Philosophy of Science Part A, 47(??):104–110, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000260. See [Bue12].

Boldyrev:2015:HGE

[Bol15] Ivan Boldyrev. History and (general) equilibrium: Reclaiming lives behind a model. Studies in History and Philosophy of Science Part A, 54(??):127–131, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001077.

Bolinska:2016:SVE

[Bol16] Agnes Bolinska. Successful visual epistemic representation. Studies in History and Philosophy of Science Part A, 56(??):153– 160, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811500134X.

Bolinska:2023:EED

[Bol23] Agnes Bolinska. Epistemic expression in the determination of biomolecular structure. Studies in History and Philosophy of Science Part A, 100(??):107–115, August 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000894.

Bonk:1994:WBT

[Bon94] Thomas Bonk. Why has de Broglie’s theory been rejected? Studies in History and Philosophy of Science Part A, 25(3):375– 396, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900582.

Boon:2023:BFP

[Boo23] Mieke Boon. Book forum: Perspectival Realism by Michela Massimi: Reconciling perspectivism and realism. Studies in History and Philosophy of Science Part A, 102(??):45–47, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001371.

Borossa:1992:PB

[Bor92] Julia Borossa. Psychoanalytic battles. Studies in History and Philosophy of Science Part A, 23(4):681–689, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368192900193.

Bordoni:2015:BBS

[Bor15a] Stefano Bordoni. On the borderline between science and philosophy: a debate on determinism in France around 1880. Studies in History and Philosophy of Science Part A, 49(??):27–35, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114001125.

Borrelli:2015:MIP

[Bor15b] Arianna Borrelli. The making of an intrinsic property: “symmetry heuristics” in early particle physics. Studies in History and Philosophy of Science Part A, 50(??):59–70, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000879.

Borg:2020:ASS

[Bor20] George Borg. On “the application of science to science itself:” chemistry, instruments, and the scientific labor process. Studies in History and Philosophy of Science Part A, 79(??):41–56, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118300529.

Borg:2025:MDL

[Bor25] George Borg. Measurement, decomposition and level-switching in historical science: Geochronology and the ontology of scientific methods. Studies in History and Philosophy of Science Part A, 109 (??):123–131, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124001687.

Boumans:2016:GBI

[Bou16] Marcel Boumans. Graph-based inductive reasoning. Studies in History and Philosophy of Science Part A, 59(??):1–10, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300139.

Bourrat:2020:NSR

[Bou20] Pierrick Bourrat. Natural selection and the reference grain problem. Studies in History and Philosophy of Science Part A, 80 (??):1–8, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300079.

Bourrat:2021:FPS

[Bou21] Pierrick Bourrat. Function, persistence, and selection: Generalizing the selected-effect account of function adequately. Studies in History and Philosophy of Science Part A, 90(??):61–67, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001412.

Bourrat:2022:UHE

[Bou22a] Pierrick Bourrat. Unifying heritability in evolutionary theory. Studies in History and Philosophy of Science Part A, 91(??):201– 210, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001758.

Bouterse:2022:CH

[Bou22b] Jeroen Bouterse. Contingentism for historians. Studies in History and Philosophy of Science Part A, 96(??):27–34, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001157.

Bowker:1993:CSF

[Bow93] Geoffrey C. Bowker. Constructing science, forging technology and manufacturing society. Studies in History and Philosophy of Science Part A, 24(1):147–155, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390028I.

Boyce:2014:EGH

[Boy14] Kenneth Boyce. On the equivalence of Goodman’s and Hempel’s paradoxes. Studies in History and Philosophy of Science Part A, 45(??):32–42, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113001313.

Bellucci:2020:PJA

[BP20] Francesco Bellucci and Ahti-Veikko Pietarinen. Peirce on the justification of abduction. Studies in History and Philosophy of Science Part A, 84(??):12–19, December 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119300950.

Bertoldi:2025:PBS

[BP25] Nicola Bertoldi and Charles H. Pence. “Population” in biology and statistics. Studies in History and Philosophy of Science Part A, 109(??):1–11, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124001699.

Bedessem:2019:SAU

[BR19] Baptiste Bedessem and Stéphanie Ruphy. Scientific autonomy and the unpredictability of scientific inquiry: the unexpected might not be where you would expect. Studies in History and Philosophy of Science Part A, 73(??):1–7, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117302406.

Banchetti-Robino:2023:BRC

[BR23] Marina Paola Banchetti-Robino. Book review: The Chemical Philosophy of Robert Boyle: Mechanicism, Chymical Atoms, and Emergence by Marina Paola Banchetti-Robino: Reply by the Author. Studies in History and Philosophy of Science Part A, 97 (??):141–144, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001637.

Brandon:1978:AET

[Bra78] Robert N. Brandon. Adaptation and evolutionary theory. Studies in History and Philosophy of Science Part A, 9(3):181–206, September 1978. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368178900055.

Brandon:1981:BTQ

[Bra81] Robert N. Brandon. Biological teleology: Questions and explanations. Studies in History and Philosophy of Science Part A, 12(2): 91–105, June 1981. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368181900157.

Brain:1995:BI

[Bra95] Robert M. Brain. Bürgerliche intelligenz. Studies in History and Philosophy of Science Part A, 26(4):617–635, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000240.

Bravo:1998:AAH

[Bra98] Michael T. Bravo. The anti-anthropology of highlanders and islanders. Studies in History and Philosophy of Science Part A, 29 (3):369–389, September 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368198000211.

Brain:2001:OQM

[Bra01] Robert Michael Brain. The ontology of the questionnaire: Max Weber on measurement and mass investigation. Studies in History and Philosophy of Science Part A, 32(4):647–684, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000267.

Brain:2008:PME

[Bra08] Robert Michael Brain. The pulse of modernism: experimental physiology and aesthetic avant-gardes circa 1900. Studies in History and Philosophy of Science Part A, 39(3):393–417, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000642.

Brading:2013:TPU

[Bra13] Katherine Brading. Three principles of unity in Newton. Studies in History and Philosophy of Science Part A, 44(3):408–415, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000805.

Brading:2015:PLP

[Bra15] Katherine Brading. Physically locating the present: a case of reading physics as a contribution to philosophy. Studies in History and Philosophy of Science Part A, 50(??):13–19, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000855.

Bradley:2024:RDC

[Bra24a] James Bradley. Redefining a discovery: Charles Bell, the respiratory nervous system and the birth of the emotions. Studies in History and Philosophy of Science Part A, 106(??):12– 20, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000566.

Brautigam:2024:HUM

[Brä24b] Maren Bräutigam. Heterodox underdetermination: Metaphysical options for discernibility and (non-)entanglement. Studies in History and Philosophy of Science Part A, 103(??):77– 84, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001632.

Burian:1996:AGM

[BRdS96] Richard M. Burian, Robert C. Richardson, and Wim J. Van der Steen. Against generality: Meaning in genetics and philosophy. Studies in History and Philosophy of Science Part A, 27(1):1– 29, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000348.

Breitenbach:2004:LTT

[Bre04a] Angela Breitenbach. Langton on things in themselves: a critique of Kantian humility. Studies in History and Philosophy of Science Part A, 35(1):137–148, March 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000918.

Brennan:2004:BMS

[Bre04b] Andrew Brennan. The birth of modern science: culture, mentalities and scientific innovation. Studies in History and Philosophy of Science Part A, 35(2):199–225, June 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000135.

Briggs:1973:BRB

[Bri73] J. Morton Briggs, Jr. Book review: Traité de dynamique: Jean d’Alembert. A reprint of the second edition, Paris, 1758. Intro. by Thomas K. Hankins; The sources of science, No. 72, Johnson Reprint Corporation, New York and London, 1968. Studies in History and Philosophy of Science Part A, 3(4):386–396, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900186. See [Le 58, Le 68].

Brody:1971:BRB

[Bro71] B. A. Brody. Book review: Words and objections: Essays on the works of W. V. O. Quine: D. Davidson and J. Hintikka, eds., Humanities Press: New York, 1969. viii + 366 pp. $14.50. Studies in History and Philosophy of Science Part A, 2(2):167– 175, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900306. See [DH69].

Brooke:1973:CSF

[Bro73] John Hedley Brooke. Chlorine substitution and the future of organic chemistry: Methodological issues in the Laurent–Berzelius correspondence (1843–1844). Studies in History and Philosophy of Science Part A, 4(1):47–94, May 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817390023X.

Brown:1976:GET

[Bro76] Harold I. Brown. Galileo, the elements, and the tides. Studies in History and Philosophy of Science Part A, 7(4):337–351, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936817690011X.

Brown:1990:HRA

[Bro90a] Bryson Brown. How to be realistic about inconsistency in science. Studies in History and Philosophy of Science Part A, 21(2): 281–294, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900276.

Brown:1990:PR

[Bro90b] Harold I. Brown. Prospective realism. Studies in History and Philosophy of Science Part A, 21(2):211–242, June 1990. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368190900232.

Brown:2005:IR

[Bro05] Harold I. Brown. Incommensurability reconsidered. Studies in History and Philosophy of Science Part A, 36(1):149–169, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104001244.

Broad:2007:MCJ

[Bro07] Jacqueline Broad. Margaret Cavendish and Joseph Glanvill: science, religion, and witchcraft. Studies in History and Philosophy of Science Part A, 38(3):493–505, September 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000350.

Brown:2009:MPS

[Bro09] Matthew J. Brown. Models and perspectives on stage: remarks on Giere’s scientific perspectivism. Studies in History and Philosophy of Science Part A, 40(2):213–220, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000156. See [Gie06, Gie09].

Broad:2011:MCW

[Bro11a] Jacqueline Broad. Is Margaret Cavendish worthy of study today? Studies in History and Philosophy of Science Part A, 42 (3):457–461, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111000197.

Brown:2011:VFM

[Bro11b] Harold I. Brown. Van Fraassen meets Popper: Logical relations and cognitive abilities. Studies in History and Philosophy of Science Part A, 42(2):381–385, June 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000148.

Brown:2015:JDP

[Bro15] Matthew J. Brown. John Dewey’s pragmatist alternative to the belief–acceptance dichotomy. Studies in History and Philosophy of Science Part A, 53(??):62–70, October 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000722.

Bromham:2016:THM

[Bro16a] Lindell Bromham. Testing hypotheses in macroevolution. Studies in History and Philosophy of Science Part A, 55(??):47–59, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001065.

Brown:2016:DSC

[Bro16b] Gregory Brown. Did Samuel Clarke really disavow action at a distance in his correspondence with Leibniz?: Newton, Clarke, and Bentley on gravitation and action at a distance. Studies in History and Philosophy of Science Part A, 60(??):38–47, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300474.

Brown:2016:AWP

[Bro16c] Matthew J. Brown. The abundant world: Paul Feyerabend’s metaphysics of science. Studies in History and Philosophy of Science Part A, 57(??):142–154, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001752.

Brown:2021:SMI

[Bro21] Matthew J. Brown. Book review: Science and Moral Imagination: a New Ideal for Values in Science by Matthew J. Brown: Reply by the author. Studies in History and Philosophy of Science Part A, 89(??):301–303, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812100042X.

Brush:1999:DTCa

[Bru99a] Stephen G. Brush. Dynamics of theory change in chemistry: Part 1. The benzene problem 1865–1945. Studies in History and Philosophy of Science Part A, 30(1):21–79, March 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198000272.

Brush:1999:DTCb

[Bru99b] Stephen G. Brush. Dynamics of theory change in chemistry: Part 2. Benzene and molecular orbitals, 1945–1980. Studies in History and Philosophy of Science Part A, 30(2):263–302, June 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198000284.

Brush:2007:PPT

[Bru07] Stephen G. Brush. Predictivism and the periodic table. Studies in History and Philosophy of Science Part A, 38(1):256– 259, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106001026.

Bruno:2020:DCR

[Bru20] Karl Bruno. Disciplining cattle reproduction: Veterinary reproductive science, bull infertility, and the mid-twentieth century transformation of Swedish dairy cattle breeding. Studies in History and Philosophy of Science Part A, 84(??):106–118, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301667.

Boudia:2013:STC

[BS13] Soraya Boudia and Sébastien Soubiran. Scientists and their cultural heritage: Knowledge, politics and ambivalent relationships. Studies in History and Philosophy of Science Part A, 44(4):643– 651, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000447.

Brading:2021:HPF

[BS21] Katherine Brading and Marius Stan. How physics flew the philosophers’ nest. Studies in History and Philosophy of Science Part A, 88(??):312–320, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000947.

Bussotti:2009:ISC

[BT09] Paolo Bussotti and Christian Tapp. The influence of Spinoza’s concept of infinity on Cantor’s set theory. Studies in History and Philosophy of Science Part A, 40(1):25–35, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810800109X.

Bradley:2019:MMM

[BT19] Seamus Bradley and Karim P. Y. Thébault. Models on the move: Migration and imperialism. Studies in History and Philosophy of Science Part A, 77(??):81–92, October 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117303047.

Buchdahl:1969:MPSa

[Buc69a] Gerd Buchdahl. Metaphysics and the philosophy of science; the classical origins, Descartes to Kant. Basil Blackwell, Oxford, UK, 1969. ISBN 0-631-11720-2. xii + 714 pp. LCCN BD313 .B77.

Buchdahl:1969:MPSb

[Buc69b] Gerd Buchdahl. Metaphysics and the philosophy of science; the classical origins: Descartes to Kant. MIT Press, Cambridge, MA, USA, 1969. ISBN 0-262-02057-2. xii + 714 pp. LCCN BD111 .B85.

Buchdahl:1972:MAK

[Buc72] Gerd Buchdahl. Methodological aspects of Kepler’s theory of refraction. Studies in History and Philosophy of Science Part A, 3 (3):265–298, November 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368172900118.

Buchdahl:1982:ERD

[Buc82] Gerd Buchdahl. Editorial response to David Bloor. Studies in History and Philosophy of Science Part A, 13(4):299–304, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368182900139. See [Blo82a, Blo82b,

Blo82c, Hes82].

Buchdahl:1988:SHP

[Buc88] Gerd Buchdahl. Studies in History and Philosophy of Science. Origins and aims: Some ‘birthday thoughts’. Studies in History and Philosophy of Science Part A, 19(1):1–3, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900179.

Buchdahl:1989:HPS

[Buc89] Gerd Buchdahl. History and philosophy of science: Some anecdotal memories. Studies in History and Philosophy of Science Part A, 20(1):5–8, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900290.

Buchwald:1992:KWT

[Buc92] Jed Z. Buchwald. Kinds and the wave theory of light. Studies in History and Philosophy of Science Part A, 23(1):39–74, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368192900263.

Buckland:2009:STD

[Buc09] Adelene Buckland. Show and tell: the dramatic story of nineteenth-century geological science. Studies in History and Philosophy of Science Part A, 40(1):114–117, March 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108001192.

Buchanan:2021:SDJ

[Buc21] Roderick D. Buchanan. Syndrome du jour: the historiography and moral implications of Diagnosing Darwin. Studies in History and Philosophy of Science Part A, 90(??):86–101, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001400.

Bud:2013:EOR

[Bud13] Robert Bud. Embodied odysseys: Relics of stories about journeys through past, present, and future. Studies in History and Philosophy of Science Part A, 44(4):639–642, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000460.

Bueno:1997:EAP

[Bue97] Otávio Bueno. Empirical adequacy: a partial structures approach. Studies in History and Philosophy of Science Part A, 28(4):585– 610, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197000125.

Bueno:2000:ESC

[Bue00] Otávio Bueno. Empiricism, scientific change and mathematical change. Studies in History and Philosophy of Science Part A, 31 (2):269–296, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368199000370.

Bueno:2012:SRP

[Bue12] Otávio Bueno. Styles of reasoning: a pluralist view. Studies in History and Philosophy of Science Part A, 43(4):657–665, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000507. See reply [Bol14].

Bueter:2015:IVF

[Bue15] Anke Bueter. The irreducibility of value-freedom to theory assessment. Studies in History and Philosophy of Science Part A, 49(??):18–26, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811400096X.

Bueno:2019:SRM

[Bue19] Otávio Bueno. Structural realism, mathematics, and ontology. Studies in History and Philosophy of Science Part A, 74(??): 4–9, ???? 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118303327.

Bueter:2022:BEN

[Bue22] Anke Bueter. Bias as an epistemic notion. Studies in History and Philosophy of Science Part A, 91(??):307–315, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121002028.

Buono:2024:TED

[Buo24] Eleonora Buono. Tracing the evidence of design: Natural theology through an unpublished manuscript by William Stanley Jevons. Studies in History and Philosophy of Science Part A, 105(??): 74–84, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812400044X.

Burney:2002:TTT

[Bur02] Ian A. Burney. Testing testimony: toxicology and the law of evidence in early nineteenth-century England. Studies in History and Philosophy of Science Part A, 33(2):289–314, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810200002X.

Bush:1971:BRB

[Bus71] Rachel Bush. Book review: Toward a history of geology: proceedings of the New Hampshire Inter-Disciplinary Conference on the History of Geology, September 7–12, 1967: Cecil J. Schneer (ed.), MIT Press: Cambridge, Mass., and London, 1969. vi + 469 pp. $22.50. Studies in History and Philosophy of Science Part A, 2 (2):176–182, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368171900318. See [Sch69].

Busch:2008:ERC

[Bus08] Jacob Busch. Eclectic realism — a cake less filling. Studies in History and Philosophy of Science Part A, 39(2):270– 272, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000307. See reply [Saa08].

Butterfield:1993:IIQ

[But93] Jeremy Butterfield. Interpretation and identity in quantum theory. Studies in History and Philosophy of Science Part A, 24(3): 443–476, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390037K.

Bensaude-Vincent:2005:CFT

[BV05] Bernadette Bensaude-Vincent. Chemistry in the French tradition of philosophy of science: Duhem, Meyerson, Metzger and Bachelard. Studies in History and Philosophy of Science Part A, 36(4):627– 649, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000798.

Bycroft:2012:KES

[Byc12] Michael Bycroft. Kuhn’s evolutionary social epistemology. Studies in History and Philosophy of Science Part A, 43(3):425–429, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000362.

Ben-Zeev:1981:JJG

[BZ81] Aaron Ben-Zeev. J. J. Gibson and the ecological approach to perception. Studies in History and Philosophy of Science Part A, 12 (2):107–139, June 1981. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368181900169.

Ben-Zeev:1986:RDA

[BZ86] Aaron Ben-Zeev. Reid’s direct approach to perception. Studies in History and Philosophy of Science Part A, 17(1):99–114, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900208.

Cabral:1996:HBC

[Cab96] Regis Cabral. Herbert Butterfield (1900–1979) as a Christian Historian of Science. Studies in History and Philosophy of Science Part A, 27(4):547–564, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368196000076.

Cabrera:2020:EEC

[Cab20] Frank Cabrera. Evidence and explanation in Cicero’s On Divination. Studies in History and Philosophy of Science Part A, 82(??): 34–43, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301451.

Cahoone:1986:IGS

[Cah86] Lawrence E. Cahoone. The interpretation of Galilean Science: Cassirer contrasted with Husserl and Heidegger. Studies in History and Philosophy of Science Part A, 17(1):1–21, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900166.

Cahan:1997:HBI

[Cah97] David Cahan. On Helmholtz and ‘Bürgerliche intelligenz’: a response to Robert Brain. Studies in History and Philosophy of Science Part A, 28(3):521–532, September 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197863364.

Callergaard:2010:TRN

[Cal10] Robert Callergård. Thomas Reid’s Newtonian theism: his differences with the classical arguments of Richard Bentley and William Whiston. Studies in History and Philosophy of Science Part A, 41(2):109–119, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811000021X.

Calamari:2021:MCL

[Cal21] Martin Calamari. The metaphysical challenge of loop quantum gravity. Studies in History and Philosophy of Science Part A, 86 (??):68–83, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000169.

Calosi:2022:QMI

[Cal22] Claudio Calosi. Quantum modal indeterminacy. Studies in History and Philosophy of Science Part A, 95(??):177–184, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001261.

Calderon:2024:CAA

[Cal24] Francisco Calderón. The causal axioms of algebraic quantum field theory: a diagnostic. Studies in History and Philosophy of Science Part A, 104(??):98–108, April 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000268.

Campbell:1983:AF

[Cam83] Margaret Campbell. Adaptation and fitness. Studies in History and Philosophy of Science Part A, 14(1):59–65, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368183900055.

Campbell:1997:EEP

[Cam97] Sue Campbell. Emotion as an explanatory principle in early evolutionary theory. Studies in History and Philosophy of Science Part A, 28(3):453–473, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368196000180.

Campbell:2012:RJD

[Cam12] Andrew Campbell. The reception of John Dee’s Monas hieroglyphica in early modern Italy: the case of Paolo Antonio Foscarini (c 1562–1616). Studies in History and Philosophy of Science Part A, 43(3):519–529, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811100121X.

Camilleri:2015:KWW

[Cam15] Kristian Camilleri. Knowing what would happen: the epistemic strategies in Galileo’s thought experiments. Studies in History and Philosophy of Science Part A, 54(??):102–112, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001028.

Campos:2018:HAA

[Cam18a] Daniel G. Campos. Heuristic analogy in Ars Conjectandi: From Archimedes’ De Circuli Dimensione to Bernoulli’s theorem. Studies in History and Philosophy of Science Part A, 67(??):44–53, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301620.

Camprubi:2018:EDG

[Cam18b] Lino Camprubí. Experiencing deep and global currents at a ‘Prototypical Strait’, 1870s and 1980s. Studies in History and Philosophy of Science Part A, 70(??):1–5, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301365.

Camacho:2021:BDA

[Cam21] M. Polo Camacho. Beyond descriptive accuracy: the central dogma of molecular biology in scientific practice. Studies in History and Philosophy of Science Part A, 86(??):20–26, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000017.

Camp:2022:BRR

[Cam22] Elisabeth Camp. Book review: Re presentation in Cognitive Function by Nicholas Shea: Organization and structure in the service of systematicity and productivity. Studies in History and Philosophy of Science Part A, 92(??):264–266, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000790.

Cantor:1971:HBS

[Can71] G. N. Cantor. Henry Brougham and the Scottish methodological tradition. Studies in History and Philosophy of Science Part A, 2(1):69–89, May 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368171900380.

Canguilhem:1989:NP

[Can89] Georges Canguilhem. The normal and the pathological. Zone Books, New York, NY, USA, 1989. ISBN 0-942299-58-2 (hardcover), 0-942299-59-0 (paperback). 327 pp. LCCN RB112 .C313 1989. Translated by Carolyn R. Fawcett in collaboration with Robert S. Cohen, with an introduction by Michel Foucault.

Caneva:1998:ORI

[Can98] Kenneth L. Caneva. Objectivity, relativism, and the individual: a role for a post-Kuhnian history of science. Studies in History and Philosophy of Science Part A, 29(3):327–344, September 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198000120.

Caplan:1979:DDM

[Cap79] Arthur L. Caplan. Darwinism and deductivist models of theory structure. Studies in History and Philosophy of Science Part A, 10 (4):341–353, December 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368179900190.

Carnap:1921:RBZ

[Car21] Rudolf Carnap. Der Raum: Ein Beitrag zur Wissenschaftslehre. (German) [Space: a contribution to science teaching]. PhD thesis, Universität Göttingen, Göttingen, Germany, 1921. 87 pp.

Carozzi:1973:BRB

[Car73] Albert V. Carozzi. Book review: The Earth in decay: A history of British geomorphology 1578–1878: Gordon L. Davies, New York: American Elsevier Publishing Company Inc., 1969. 390 pp. $16. Studies in History and Philosophy of Science Part A, 4(3):289– 299, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900125. See [Her69a, Her69b].

Cardwell:1989:JJB

[Car89] D. S. L. (Donald Stephen Lowell) Cardwell. James Joule: a biography. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-3025-0. ix + 333 pp. LCCN QC16.J68 C368 1989.

Carrier:1991:WWM

[Car91] Martin Carrier. What is wrong with the miracle argument? Studies in History and Philosophy of Science Part A, 22(1):23– 36, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190013I.

Carrier:1993:WRM

[Car93] Martin Carrier. What is right with the miracle argument: Establishing a taxonomy of natural kinds. Studies in History and Philosophy of Science Part A, 24(3):391–409, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390035I.

Carter:1995:TRH

[Car95] K. Codell Carter. Toward a rational history of medical science. Studies in History and Philosophy of Science Part A, 26(3):493– 502, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819500007A.

Cardoso:2003:NHP

[Car03a] José Luís Cardoso. From natural history to political economy: the enlightened mission of Domenico Vandelli in late eighteenth-century Portugal. Studies in History and Philosophy of Science Part A, 34(4):781–803, December 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000645.

Carrier:2003:HTC

[Car03b] Martin Carrier. How to tell causes from effects: Kant’s causal theory of time and modern approaches. Studies in History and Philosophy of Science Part A, 34(1):59–71, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000845.

Carman:2005:EDC

[Car05] Christián C. Carman. The electrons of the dinosaurs and the center of the Earth: comments on D. D. Turner’s ‘The past vs. the tiny: historical science and the abductive arguments for realism’. Studies in History and Philosophy of Science Part A, 36(1):171– 173, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104001256.

Carlin:2012:BTM

[Car12]  Laurence Carlin.  Boyle’s teleological mechanism and the

myth of immanent teleology. Studies in History and Philosophy of Science Part A, 43(1):54–63, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000653.

Carnino:2013:RMU

[Car13] Guillaume Carnino. The Romantic Machine: Utopian science and technology after Napoleon. Studies in History and Philosophy of Science Part A, 44(4):608–612, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000174.

Carrier:2014:PCC

[Car14] Martin Carrier. Prediction in context: On the comparative epistemic merit of predictive success. Studies in History and Philosophy of Science Part A, 45(??):97–102, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811300099X.

Carman:2015:PIB

[Car15] Christián Carlos Carman. The planetary increase of brightness during retrograde motion: an explanandum constructed ad explanantem. Studies in History and Philosophy of Science Part A, 54(??):90–101, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001363.

Cardona:2016:KAS

[Car16] Carlos Alberto Cardona. Kepler: Analogies in the search for the law of refraction. Studies in History and Philosophy of Science Part A, 59(??):22–35, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300164.

Carosso:2022:QHP

[Car22] Andrea Carosso. Quantization: History and problems. Studies in History and Philosophy of Science Part A, 96(??):35–50, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001339.

Carvalho:2025:HNT

[Car25] Cláudio Alexandre S. Carvalho. Human nature and therapeutic forms in B. Mandeville. Studies in History and Philosophy of Science Part A, 110(??):76–87, April 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368125000160.

Castellani:2024:CST

[Cas24] Elena Castellani. Convergence strategies for theory assessment. Studies in History and Philosophy of Science Part A, 104(??): 78–87, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000190.

Cat:1995:PND

[Cat95] Jordi Cat. The Popper–Neurath debate and Neurath’s attack on scientific method. Studies in History and Philosophy of Science Part A, 26(2):219–250, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819400038B.

Cat:2012:RPM

[Cat12] Jordi Cat. Into the ‘regions of physical and metaphysical chaos’: Maxwell’s scientific metaphysics and natural philosophy of action (agency, determinacy and necessity from theology, moral philosophy and history to mathematics, theory and experiment). Studies in History and Philosophy of Science Part A, 43(1):91– 104, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001002.

Cat:2014:MCS

[Cat14] Jordi Cat. Maxwell’s color statistics: From reduction of visible errors to reduction to invisible molecules. Studies in History and Philosophy of Science Part A, 48(??):60–75, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000582.

Catren:2022:GSI

[Cat22] Gabriel Catren. On gauge symmetries, indiscernibilities, and groupoid-theoretical equalities. Studies in History and Philosophy of Science Part A, 91(??):244–261, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001795.

Chen:1992:CAP

[CB92] Xiang Chen and Peter Barker. Cognitive appraisal and power: David Brewster, Henry Brougham, and the tactics of the emission — undulatory controversy during the early 1850s. Studies in History and Philosophy of Science Part A, 23(1):75–101, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368192900274.

Campbell:2024:HAI

[CCB24] Jonah Campbell, Alberto Cambrosio, and Mark Basik. Histology agnosticism: Infra-molecularizing disease? Studies in History and Philosophy of Science Part A, 104(??):14–22, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000116.

Carman:2015:DPM

[CD15] Christián Carman and José Díez. Did Ptolemy make novel predictions? Launching Ptolemaic astronomy into the scientific realism debate. Studies in History and Philosophy of Science Part A, 52(??):20–34, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000436.

Coraci:2024:IBN

[CDC24] Davide Coraci, Igor Douven, and Gustavo Cevolani. Inference to the best neuroscientific explanation. Studies in History and Philosophy of Science Part A, 107(??):33–42, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001079.

Collins:2015:ERP

[CE15] Harry Collins and Robert Evans. Expertise revisited, Part I — interactional expertise. Studies in History and Philosophy of Science Part A, 54(??):113–123, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811500093X.

Collins:2019:IGN

[CE19] Harry Collins and Robert Evans. The Imitation Game and the nature of science. Studies in History and Philosophy of Science Part A, 76(??):86–90, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118303169.

Cedarbaum:1983:P

[Ced83] Daniel Goldman Cedarbaum. Paradigms. Studies in History and Philosophy of Science Part A, 14(3):173–213, September 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368183900122.

Collins:2007:TZI

[CEG07] Harry Collins, Robert Evans, and Mike Gorman. Trading zones and interactional expertise. Studies in History and Philosophy of Science Part A, 38(4):657–666, December 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810700060X.

Cellucci:2008:NME

[Cel08] Carlo Cellucci. The nature of mathematical explanation. Studies in History and Philosophy of Science Part A, 39(2):202– 210, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000228.

Cellucci:2013:PMM

[Cel13] Carlo Cellucci. Philosophy of mathematics: Making a fresh start. Studies in History and Philosophy of Science Part A, 44(1):32– 42, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000738.

Cerbone:2019:ERS

[Cer19] David R. Cerbone. Essay review: Social epistemology meets Heideggerian ontology. Studies in History and Philosophy of Science Part A, 76(??):94–97, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302954.

Collins:2006:EIE

[CERH06] Harry Collins, Rob Evans, Rodrigo Ribeiro, and Martin Hall. Experiments with interactional expertise. Studies in History and Philosophy of Science Part A, 37(4):656–674, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810600080X.

Collins:2016:ERP

[CEW16] Harry Collins, Robert Evans, and Martin Weinel. Expertise revisited, Part II: Contributory expertise. Studies in History and Philosophy of Science Part A, 56(??):103–110, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000928.

Cei:2006:LSA

[CF06] Angelo Cei and Steven French. Looking for structure in all the wrong places: Ramsey sentences, multiple realisability, and structure. Studies in History and Philosophy of Science Part A, 37 (4):633–655, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000793.

Chen:2021:AAP

[CF21a] Lu Chen and Tobias Fritz. An algebraic approach to physical fields. Studies in History and Philosophy of Science Part A, 89(??):188– 201, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001278.

Cinti:2021:CSE

[CF21b] Enrico Cinti and Vincenzo Fano. Careful with those scissors, Eugene! Against the observational indistinguishability of spacetimes. Studies in History and Philosophy of Science Part A, 89(??):103– 113, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001084.

Chambers:1844:VNH

[Cha44] Robert Chambers. Vestiges of the natural history of creation. J. Churchill, London, UK, 1844. vi + 390 pp. LCCN QH363 .C4 1844.

Chambers:1969:VNH

[Cha69] Robert Chambers. Vestiges of the natural history of creation. The Victorian library. Humanities Press, New York, NY, USA, 1969. ISBN 0-7185-5001-3. 390 pp. LCCN QH363 .C4 1970. With an introduction by Gavin de Beer.

Chalmers:1973:MMH

[Cha73] A. F. Chalmers. Maxwell’s methodology and his application of it to electromagnetism. Studies in History and Philosophy of Science Part A, 4(2):107–164, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900010.

Chalmers:1986:HRM

[Cha86] Alan F. Chalmers. The heuristic role of Maxwell’s mechanical model of electromagnetic phenomena. Studies in History and Philosophy of Science Part A, 17(4):415–427, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900026.

Chalmers:1993:LEB

[Cha93] Alan Chalmers. The lack of excellency of Boyle’s mechanical philosophy. Studies in History and Philosophy of Science Part A, 24 (4):541–564, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390052L.

Chakravartty:1998:S

[Cha98] Anjan Chakravartty. Semirealism. Studies in History and Philosophy of Science Part A, 29(3):391–408, September 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000132.

Chalmers:2002:EVM

[Cha02] Alan Chalmers. Experiment versus mechanical philosophy in the work of Robert Boyle: a reply to Anstey and Pyle. Studies in History and Philosophy of Science Part A, 33(1):187– 193, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810100036X.

Chakravartty:2004:ES

[Cha04] Anjan Chakravartty. The empirical stance. Studies in History and Philosophy of Science Part A, 35(1):173–184, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000888.

Chakravartty:2010:EIT

[Cha10a] Anjan Chakravartty. Explanation, inference, testimony, and truth: essays dedicated to the memory of Peter Lipton. Studies in History and Philosophy of Science Part A, 41(4):335–336, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000592.

Chakravartty:2010:PIM

[Cha10b] Anjan Chakravartty. Perspectivism, inconsistent models, and contrastive explanation. Studies in History and Philosophy of Science Part A, 41(4):405–412, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000622.

Chalmers:2010:BOM

[Cha10c] Alan F. Chalmers. Boyle and the origins of modern chemistry: Newman tried in the fire. Studies in History and Philosophy of Science Part A, 41(1):1–10, March 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000879. See reply [New10].

Chalmers:2011:UST

[Cha11] Alan Chalmers. Understanding science through its history: a response to Newman. Studies in History and Philosophy of Science Part A, 42(1):150–153, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000853.

Chalmers:2012:ICE

[Cha12] Alan Chalmers. Intermediate causes and explanations: the key to understanding the scientific revolution. Studies in History and Philosophy of Science Part A, 43(4):551–562, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000416.

Charnley:2013:EEB

[Cha13] Berris Charnley. Experiments in empire-building: Mendelian genetics as a national, imperial, and global agricultural enterprise. Studies in History and Philosophy of Science Part A, 44(2):292– 300, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001033.

Chalmers:2015:QNS

[Cha15a] Alan F. Chalmers. Qualitative novelty in seventeenth-century science: Hydrostatics from Stevin to Pascal. Studies in History and Philosophy of Science Part A, 51(??):1–10, June 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000023.

Chang:2015:CRR

[Cha15b] Hasok Chang. The Chemical Revolution revisited. Studies in History and Philosophy of Science Part A, 49(??):91–98, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114001101.

Chalmers:2016:VPS

[Cha16] Alan Chalmers. Viewing past science from the point of view of present science, thereby illuminating both: Philosophy versus experiment in the work of Robert Boyle. Studies in History and Philosophy of Science Part A, 55(??):27–35, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000989.

Chakravartty:2019:PMD

[Cha19a] Anjan Chakravartty. Physics, metaphysics, dispositions, and symmetries — à la French. Studies in History and Philosophy of Science Part A, 74(??):10–15, ???? 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118303339.

Champagne:2019:DAW

[Cha19b] Marc Champagne. Diagrams and alien ways of thinking. Studies in History and Philosophy of Science Part A, 75(??):12–22, June 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0039368118300281.

Chatzigeorgiou:2020:HMWa

[Cha20a] Konstantinos Chatzigeorgiou. How the mind-world problem shaped the history of science: a historiographical analysis of Edwin Arthur Burtt’s The Metaphysical Foundations of Modern Physical Science Part I. Studies in History and Philosophy of Science Part A, 83 (??):121–132, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119303383.

Chatzigeorgiou:2020:HMWb

[Cha20b] Konstantinos Chatzigeorgiou. How the mind-world problem shaped the history of science: a historiographical analysis of Edwin Arthur Burtt’s The Metaphysical Foundations of Modern Physical Science Part II. Studies in History and Philosophy of Science Part A, 83 (??):133–143, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120300078.

Chen:1990:YLP

[Che90] Xiang Chen. Young and Lloyd on the particle theory of light: a response to Achinstein. Studies in History and Philosophy of Science Part A, 21(4):665–676, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900379.

See [Ach87, Ach90].

Chen:1995:TCP

[Che95] Xiang Chen. Taxonomic changes and the particle-wave debate in early nineteenth-century Britain. Studies in History and Philosophy of Science Part A, 26(2):251–271, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000014.

Chen:2003:WDJ

[Che03] Xiang Chen. Why did John Herschel fail to understand polarization? The differences between object and event concepts. Studies in History and Philosophy of Science Part A, 34(3):491–513, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810300044X.

Chen:2010:DKR

[Che10] Xiang Chen. A different kind of revolutionary change: transformation from object to process concepts. Studies in History and Philosophy of Science Part A, 41(2):182–191, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000208.

Chen:2022:CWE

[Che22] Lu Chen. Can we “effectivize” spacetime? Studies in History and Philosophy of Science Part A, 95(??):75–83, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001017.

Chen:2023:NGM

[Che23] Elliott D. Chen. Newtonian gravitation in Maxwell spacetime. Studies in History and Philosophy of Science Part A, 102(??):22– 30, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001048.

Chimisso:2001:HMH

[Chi01] Cristina Chimisso. Hélène Metzger: the history of science between the study of mentalities and total history. Studies in History and Philosophy of Science Part A, 32(2):203–241, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000061.

Chimisso:2006:IRP

[Chi06] Cristina Chimisso. The identity and routes of philosophy of science. Studies in History and Philosophy of Science Part A, 37(2): 353–360, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000318.

Chimisso:2008:PPR

[Chi08] Cristina Chimisso. From phenomenology to phenomenotechnique: the role of early twentieth-century physics in Gaston Bachelard’s philosophy. Studies in History and Philosophy of Science Part A, 39 (3):384–392, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000630.

Chin:2011:MIC

[Chi11] Chuanfei Chin. Models as interpreters (with a case study from pain science). Studies in History and Philosophy of Science Part A, 42(2):303–312, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110001172.

Chimisso:2015:NEG

[Chi15] Cristina Chimisso. Narrative and epistemology: Georges Canguilhem’s concept of scientific ideology. Studies in History and Philosophy of Science Part A, 54(??):64–73, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001090.

Chirimuuta:2016:WSE

[Chi16] M. Chirimuuta. Why the “stimulus-error” did not go away. Studies in History and Philosophy of Science Part A, 56(??): 33–42, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001557.

ChoGlueck:2022:SNP

[Cho22] Christopher ChoGlueck. Still no pill for men? Double standards and demarcating values in biomedical research. Studies in History and Philosophy of Science Part A, 91(??):66–76, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001874.

Christie:1994:PVC

[Chr94] Maureen Christie. Philosophers versus chemists concerning ‘laws of nature’. Studies in History and Philosophy of Science Part A, 25(4):613–629, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900507.

Chung:2022:BRV

[Chu22] Ryoa Chung. Book review: Vaccine Hesitancy: Public Trust, Expertise, and the War on Science by Maya Goldenberg: Science, ideology, and the democratic ethos. Studies in History and Philosophy of Science Part A, 96(??):188–190, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001509. See [Joh22, WV22, Lea22] and

reply [Gol22].

Cirera:1993:CPM

[Cir93] Ramon Cirera. Carnap’s philosophy of mind. Studies in History and Philosophy of Science Part A, 24(3):351–358, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390033G.

Carrillo:2022:HIA

[CK22] Natalia Carrillo and Tarja Knuuttila. Holistic idealization: an artifactual standpoint. Studies in History and Philosophy of Science Part A, 91(??):49–59, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001655.

Castelao-Lawless:2004:KMO

[CL04] Teresa Castelão-Lawless. Kuhn’s missed opportunity and the multifaceted lives of Bachelard: mythical, institutional, historical, philosophical, literary, scientific. Studies in History and Philosophy of Science Part A, 35(4):873–881, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000895.

Chang:2005:IMR

[CL05a] Hasok Chang and Sabina Leonelli. Infrared metaphysics: radiation and theory-choice. Part 2. Studies in History and Philosophy of Science Part A, 36(4):687–706, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000816.

Chang:2005:IME

[CL05b] Hasok Chang and Sabina Leonelli. Infrared metaphysics: the elusive ontology of radiation. Part 1. Studies in History and Philosophy of Science Part A, 36(3):477–508, September 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000397.

Chemla:2005:NCC

[CL05c] Karine Chemla and Geoffrey Lloyd. Les neuf chapitres: le classique mathématique de la Chine ancienne et ses commentaires. Dunod, Paris, France, 2005. ISBN 2-10-049589-5. 1216 pp. LCCN ????

Camprubi:2018:SEI

[CL18] Lino Camprubí and Philipp Lehmann. The scales of experience: Introduction to the special issue experiencing the global environment. Studies in History and Philosophy of Science Part A, 70 (??):ii, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301353.

Canali:2022:RED

[CL22] Stefano Canali and Sabina Leonelli. Reframing the environment in data-intensive health sciences. Studies in History and Philosophy of Science Part A, 93(??):203–214, June 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000656.

Clagett:1968:NOM

[Cla68] Marshall Clagett, editor. Nicole Oresme and the medieval geometry of qualities and motions; a treatise on the uniformity and difformity of intensities known as Tractatus de configurationibus qualitatum et motuum: Medieval geometry of qualities and motions. The University of Wisconsin publications in medieval science. University of Wisconsin Press, Madison, WI, USA, 1968. xiii + 713 pp. LCCN QA32 .O6813. Edited with an introduction, English translation, and commentary by Marshall Clagett.

Clarke:1979:PMD

[Cla79] Desmond Clarke. Physics and metaphysics in Descartes’ Principles. Studies in History and Philosophy of Science Part A, 10(2): 89–112, June 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900128. See [Des44, Des49,

DMM83].

Clark:1992:PS

[Cla92] William Clark. Poetics for scientists. Studies in History and Philosophy of Science Part A, 23(1):181–192, March 1992. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936819290030A.

Clark:1995:NHS

[Cla95] William Clark. Narratology and the history of science. Studies in History and Philosophy of Science Part A, 26(1):1–71, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194000299.

Clendinnen:1983:RMV

[Cle83] F. John Clendinnen. The rationality of method verssus historical relativism. Studies in History and Philosophy of Science Part A, 14(1):23–38, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368183900031.

Clericuzio:1997:ATM

[Cle97] Antonio Clericuzio. Alchemical theories of matter. Studies in History and Philosophy of Science Part A, 28(2):369–375, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368196000209.

Clericuzio:2023:BRC

[Cle23] Antonio Clericuzio. Book review: The Chemical Philosophy of Robert Boyle: Mechanism, Chymical Atoms, and Emergence by Marina Paola Banchetti-Robino: the agreement and the disagreement of chymists with natural philosophers. Studies in History and Philosophy of Science Part A, 97(??):133–134, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001601.

Clucas:2012:PRI

[Clu12a] Stephen Clucas. ‘this paradoxall Restitution Iudaicall’: the apocalyptic correspondence of John Dee and Roger Edwardes. Studies in History and Philosophy of Science Part A, 43(3):509–518, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001191.

Clulee:2012:JDI

[Clu12b] Nicholas H. Clulee. John Dee’s ideas and plans for a national research institute. Studies in History and Philosophy of Science Part A, 43(3):437–448, September 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001130.

Cooper:2000:DAB

[CM00] Brian P. Cooper and Margueritte S. Murphy. The death of the author at the birth of social science: The cases of Harriet Martineau and Adolphe Quetelet. Studies in History and Philosophy of Science Part A, 31(1):1–36, March 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000345.

Castellani:2022:RGM

[CM22] Elena Castellani and Emilia Margoni. Renormalization group methods: Which kind of explanation? Studies in History and Philosophy of Science Part A, 95(??):158–166, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001145.

Chalmers:1983:GDE

[CN83] Alan Chalmers and Richard Nicholas. Galileo on the dissipative effect of a rotating earth. Studies in History and Philosophy of Science Part A, 14(4):315–340, December 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368183900109.

Coady:2002:TIA

[Coa02] C. A. J. Coady. Testimony and intellectual autonomy. Studies in History and Philosophy of Science Part A, 33(2):355– 372, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000043.

Cobb:2011:HSP

[Cob11] Aaron D. Cobb. History and scientific practice in the construction of an adequate philosophy of science: revisiting a Whewell/ Mill debate. Studies in History and Philosophy of Science Part A, 42(1):85–93, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110001019.

Cochrane:2014:MCS

[Coc14] Sally Cochrane. The Munsell Color System: a scientific compromise from the world of art. Studies in History and Philosophy of Science Part A, 47(??):26–41, September 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000235.

Coelho:2022:CEM

[Coe22a] Ricardo Lopes Coelho. Comment on Eisenthal’s ‘mechanics without mechanisms’. Studies in History and Philosophy of Science Part A, 95(??):104–107, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001170.

Coetser:2022:AEP

[Coe22b] Yolandi M. Coetser. An African ethical perspective on South Africa’s regulatory frameworks governing animals in research. Studies in History and Philosophy of Science Part A, 92(??):119– 128, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000310.

Cohen:1985:MTC

[Coh85] H. F. Cohen. Music as a test-case. Studies in History and Philosophy of Science Part A, 16(4):351–378, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900172.

Cohen:1997:SSL

[Coh97] Sande Cohen. Science studies and language suppression — a critique of Bruno Latour’s We have never been modern. Studies in History and Philosophy of Science Part A, 28(2):339– 361, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000179. See [Lat93].

Cohen:2008:KPH

[Coh08a] Alix A. Cohen. Kantian philosophy and the human sciences. Studies in History and Philosophy of Science Part A, 39(4):459– 461, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000794.

Cohen:2008:KAQ

[Coh08b] Alix A. Cohen. Kant’s answer to the question ‘what is man?’ and its implications for anthropology. Studies in History and Philosophy of Science Part A, 39(4):506–514, December 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000848.

Cohen:2018:KSN

[Coh18] Alix Cohen. Kant on science and normativity. Studies in History and Philosophy of Science Part A, 71(??):6–12, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811730016X.

Coko:2015:EBH

[Cok15] Klodian Coko. Epistemology of a believing historian: Making sense of Duhem’s anti-atomism. Studies in History and Philosophy of Science Part A, 50(??):71–82, April 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000818.

Collins:1984:WDS

[Col84] H. M. Collins. When do scientists prefer to vary their experiments? Studies in History and Philosophy of Science Part A, 15(2):169–174, June 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368184900050.

Collier:1990:TFM

[Col90] John D. Collier. Two faces of Maxwell’s demon reveal the nature of irreversibility. Studies in History and Philosophy of Science Part A, 21(2):257–268, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900254.

Collins:1994:SCE

[Col94] H. M. Collins. A strong confirmation of the experimenters’ regress. Studies in History and Philosophy of Science Part A, 25(3):493– 503, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900639.

Coleman:2002:TSS

[Col02a] Martin Coleman. Taking Simmel seriously in evolutionary epistemology. Studies in History and Philosophy of Science Part A, 33 (1):55–74, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000280.

Collins:2002:ERP

[Col02b] H. M. Collins. The experimenter’s regress as philosophical sociology. Studies in History and Philosophy of Science Part A, 33(1): 149–156, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000334.

Collins:2003:LGS

[Col03] H. M. Collins. Lead into gold: the science of finding nothing. Studies in History and Philosophy of Science Part A, 34(4):661– 691, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000657.

Collins:2004:GSS

[Col04] H. M. (Harry M.) Collins. Gravity’s shadow: the search for gravitational waves. University of Chicago Press, Chicago, IL, USA and London, UK, 2004. ISBN 0-226-11377-9 (hardcover), 0-226-11378-7 (paperback). xxiii + 870 pp. LCCN QC179

.C65 2004. URL http://www.loc.gov/catdir/description/
uchi052/2003023823.html;   http://www.loc.gov/catdir/enhancements/
fy0614/2003023823-b.html;   http://www.loc.gov/catdir/
toc/ecip0410/2003023823.html.

Collins:2007:MUP

[Col07a] Harry Collins. Mathematical understanding and the physical sciences. Studies in History and Philosophy of Science Part A, 38 (4):667–685, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000581.

Collins:2007:NPR

[Col07b] Harry Collins. A new programme of research? Studies in History and Philosophy of Science Part A, 38(4):615–620, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000623.

Collins:2008:ROP

[Col08] Harry Collins. Response to one point in Gingras’s review of Gravity’s shadow. Studies in History and Philosophy of Science Part A, 39(1):151–153, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000957. See [Gin07].

Collins:2009:GRR

[Col09] Harry Collins. Gingras and the rules regress. Studies in History and Philosophy of Science Part A, 40(1):113, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108001180.

Collodel:2016:WFP

[Col16] Matteo Collodel. Was Feyerabend a Popperian? Methodological issues in the history of the philosophy of science. Studies in History and Philosophy of Science Part A, 57(??):27– 56, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000977.

Colaco:2018:RIS

[Col18] David Colaço. Rip it up and start again: the rejection of a characterization of a phenomenon. Studies in History and Philosophy of Science Part A, 72(??):32–40, December 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811730211X.

Coles:2021:ETL

[Col21] Anne-Marie Coles. Emergence of a techno-legal specialty: Animal tests to assess chemical safety in the UK, 1945–1960. Studies in History and Philosophy of Science Part A, 90(??):131–139, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001370.

Connell:2000:AGS

[Con00] Sophia M. Connell. Aristotle and Galen on sex difference and reproduction: a new approach to an ancient rivalry. Studies in History and Philosophy of Science Part A, 31(3):405–427, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000078.

Contessa:2006:CEO

[Con06a] Gabriele Contessa. Constructive empiricism, observability and three kinds of ontological commitment. Studies in History and Philosophy of Science Part A, 37(3):454–468, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000513.

Contessa:2006:SMP

[Con06b] Gabriele Contessa. Scientific models, partial structures and the new received view of theories. Studies in History and Philosophy of Science Part A, 37(2):370–377, June 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000331.

Connolly:2015:LSL

[Con15] Patrick J. Connolly. Lockean superaddition and Lockean humility. Studies in History and Philosophy of Science Part A, 51(??):

53–61, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000278.

Conix:2020:ECE

[Con20] Stijn Conix. Enzyme classification and the entanglement of values and epistemic standards. Studies in History and Philosophy of Science Part A, 84(??):37–45, December 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119302043.

Connolly:2021:CGG

[Con21] Patrick J. Connolly. Causation and gravitation in George Cheyne’s Newtonian natural philosophy. Studies in History and Philosophy of Science Part A, 85(??):145–154, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301874.

Cooter:1995:DW

[Coo95] Roger Cooter. Discourses on war. Studies in History and Philosophy of Science Part A, 26(4):637–647, December 1995. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936819500016X.

Cook:1996:MMA

[Coo96] Harold J. Cook. A material man: The alchemy of money in J. J. Becher’s writings. Studies in History and Philosophy of Science Part A, 27(3):387–396, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000496.

Cook:2005:MME

[Coo05] Simon Cook. Minds, machines and economic agents: Cambridge receptions of Boole and Babbage. Studies in History and Philosophy of Science Part A, 36(2):331–350, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810500018X.

Cooper:2020:KUC

[Coo20] Andrew Cooper. Kant’s universal conception of natural history. Studies in History and Philosophy of Science Part A, 79(??):77– 86, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301584.

Cooper:2023:HKP

[Coo23] Andrew Cooper. Hypotheses in Kant’s philosophy of science. Studies in History and Philosophy of Science Part A, 99(??):97– 105, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000668.

Corry:1993:KIS

[Cor93] Leo Corry. Kuhnian issues, scientific revolutions and the history of mathematics. Studies in History and Philosophy of Science Part A, 24(1):95–117, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390026G.

Corfield:1997:ALP

[Cor97] David Corfield. Assaying Lakatos’s philosophy of mathematics. Studies in History and Philosophy of Science Part A, 28(1):99– 121, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000027.

Corfield:2001:IMC

[Cor01] David Corfield. The importance of mathematical conceptualisation. Studies in History and Philosophy of Science Part A, 32 (3):507–533, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000073.

Corfield:2011:UII

[Cor11] David Corfield. Understanding the infinite II: Coalgebra. Studies in History and Philosophy of Science Part A, 42(4):571–579, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811100077X.

Corvini:2021:WIW

[Cor21] Pat Corvini. What induction is (and what it should not be): a concepts-centric perspective on Norton’s radium chloride example. Studies in History and Philosophy of Science Part A, 86(??): 27–34, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120302053.

Cosans:1998:EFG

[Cos98] Christopher E. Cosans. The experimental foundations of Galen’s teleology. Studies in History and Philosophy of Science Part A, 29 (1):63–80, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000052.

Cover:1997:NBT

[Cov97] J. A. Cover. Non-basic time and reductive strategies: Leibniz’s theory of time. Studies in History and Philosophy of Science Part A, 28(2):289–318, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000167.

Cox:1975:DLS

[Cox75] Chana B. Cox. A defence of Leibniz’s spatial relativism. Studies in History and Philosophy of Science Part A, 6(2):87–111, June 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900163.

Coxhead:2012:CEP

[Cox12] Michael A. Coxhead. A close examination of the pseudo-Aristotelian Mechanical Problems: the homology between mechanics and poetry as techne. Studies in History and Philosophy of Science Part A, 43(2):300–306, June 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001269.

Collins:2007:WBC

[CP07] Harry Collins and Trevor Pinch. Who is to blame for the Challenger explosion? Studies in History and Philosophy of Science Part A, 38(1):254–255, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106001014.

Copeland:2010:DET

[CP10] B. Jack Copeland and Diane Proudfoot. Deviant encodings and Turing’s analysis of computability. Studies in History and Philosophy of Science Part A, 41(3):247–252, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000427.

Craver:2017:DDM

[CP17] Carl F. Craver and Mark Povich. The directionality of distinctively mathematical explanations. Studies in History and Philosophy of Science Part A, 63(??):31–38, June 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116301108. See reply [Lan18].

Castellane:2018:KTK

[CP18] Ariane Castellane and Cédric Paternotte. Knowledge transfer without knowledge? The case of agentive metaphors in biology. Studies in History and Philosophy of Science Part A, 72(??):49– 58, December 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301213.

Charnley:2013:IPP

[CR13] Berris Charnley and Gregory Radick. Intellectual property, plant breeding and the making of Mendelian genetics. Studies in History and Philosophy of Science Part A, 44(2):222–233, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112001045.

Craw:1984:CPD

[Cra84] Robin C. Craw. ‘conservative prejudice’ in the debate over disjunctively distributed life forms. Studies in History and Philosophy of Science Part A, 15(2):131–140, June 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900037.

Crawford:1993:BRB

[Cra93] Elisabeth Crawford. Book review: Science under control: The French Academy Sciences, 1795–1914: Maurice Crosland, (Cambridge: Cambridge University Press, 1992), xix + 454 pp. Cloth £60.00. Studies in History and Philosophy of Science Part A, 24(2):305–312, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390050T. See [Cro92].

Crasnow:2017:PTP

[Cra17] Sharon Crasnow. Process tracing in political science: What’s the story? Studies in History and Philosophy of Science Part A, 62 (??):6–13, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300699.

Crasnow:2019:PSM

[Cra19] Sharon Crasnow. Political science methodology: a plea for pluralism. Studies in History and Philosophy of Science Part A, 78 (??):40–47, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117303035.

Crane:2022:BDD

[Cra22] Rosi Crane. ‘A better day dawned for biology’: T. J. Parker, New Zealand Huxleyite. Studies in History and Philosophy of Science Part A, 91(??):262–269, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121002119.

Creese:2012:RUM

[Cre12] David Creese. Rhetorical uses of mathematical harmonics in Philo and Plutarch. Studies in History and Philosophy of Science Part A, 43(2):258–269, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001312.

Creager:2018:HBC

[Cre18] Angela N. H. Creager. Human bodies as chemical sensors: a history of biomonitoring for environmental health and regulation. Studies in History and Philosophy of Science Part A, 70(??):58– 69, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301432.

Crombie:1990:EMA

[Cro90] A. C. Crombie. Expectation, modelling and assent in the history of optics: Part I. Alhazen and the medieval tradition. Studies in History and Philosophy of Science Part A, 21(4):605–632, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900357.

Crombie:1991:EMA

[Cro91] A. C. Crombie. Expectation, modelling and assent in the history of optics — II. Kepler and Descartes. Studies in History and Philosophy of Science Part A, 22(1):89–115, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190016L.

Crosland:1992:SUC

[Cro92] Maurice P. Crosland. Science under control: the French Academy of Sciences, 1795–1914. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41373-7. xvii + 454 pp.

LCCN Q46 .C77 1992. URL http://www.loc.gov/catdir/
description/cam024/91028748.html;   http://www.loc.gov/
catdir/samples/cam031/91028748.html;   http://www.loc.
gov/catdir/toc/cam028/91028748.html.

Croce:2011:WJA

[Cro11] Paul J. Croce. William James: in the academy but not of it. Studies in History and Philosophy of Science Part A, 42(1):206– 209, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000877.

Cruse:2005:RSS

[Cru05] Pierre Cruse. Ramsey sentences, structural realism and trivial realization. Studies in History and Philosophy of Science Part A, 36 (3):557–576, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000518.

Chaucer:1978:TAA

[CS78] Geoffrey Chaucer and Walter William Skeat. A treatise on the astrolabe, addressed to his son Lowys, A.D. 1391. Meridian, Amsterdam, The Netherlands, 1978. lxix + vii + 119 pp. LCCN ????

Collins:2007:TGY

[CS07] Harry Collins and Gary Sanders. They give you the keys and say ‘drive it!’ Managers, referred expertise, and other expertises. Studies in History and Philosophy of Science Part A, 38(4):621– 641, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000593.

Currie:2017:DST

[CS17] Adrian Currie and Kim Sterelny. In defence of story-telling. Studies in History and Philosophy of Science Part A, 62(??): 14–21, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300705.

Corti:2021:HMP

[CS21] Alberto Corti and Marco Sanchioni. How many properties of spin does a particle have? Studies in History and Philosophy of Science Part A, 90(??):111–121, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001382.

Cristalli:2021:CMP

[CSD21] Claudia Cristalli and Julia Sánchez-Dorado. Colligation in modelling practices: From Whewell’s tides to the San Francisco Bay Model. Studies in History and Philosophy of Science Part A, 85 (??):1–15, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120302016.

Cevolani:2013:TME

[CT13] Gustavo Cevolani and Luca Tambolo. Truth may not explain predictive success, but truthlikeness does. Studies in History and Philosophy of Science Part A, 44(4):590–593, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000393.

Crupi:2014:SFM

[CT14] Vincenzo Crupi and Katya Tentori. State of the field: Measuring information and confirmation. Studies in History and Philosophy of Science Part A, 47(??):81–90, September 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811400048X.

Currie:2016:ISK

[CT16] Adrian Currie and Derek Turner. Introduction: Scientific knowledge of the deep past. Studies in History and Philosophy of Science Part A, 55(??):43–46, February 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001326.

Cullen:2004:BRB

[Cul04] Christopher Cullen. Book review: The way and the word. Science and medicine in early China and Greece: Geoffrey Lloyd & Nathan Sivin; Yale University Press, New Haven & London, 2002, pp. xvii + 348, Price £25.00 hardback, ISBN 0-300-09297-0, Price £14.50 paperback, ISBN 0-300-10160-0. Studies in History and Philosophy of Science Part A, 35(2):357– 362, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000196. See [LS02].

Cullen:2006:ERC

[Cul06] Christopher Cullen. Essay review: Can we make the history of mathematics historical? The case of ancient China. Les neuf chapitres: Le classique mathématique de la Chine ancienne et ses commentaires, Karine Chemla & Guo Shuchun; Dunod, Paris, 2004, pp. 1140, Price £80 hardback, ISBN 2-10-0495895. Studies in History and Philosophy of Science Part A, 37(3):515–525, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000525. See [CL05c].

Cunningham:1988:GGR

[Cun88] Andrew Cunningham. Getting the game right: Some plain words on the identity and invention of science. Studies in History and Philosophy of Science Part A, 19(3):365–389, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900052.

Cunningham:2000:SRT

[Cun00] Andrew Cunningham. Science and religion in the thirteenth century revisited: the making of St Francis the proto-ecologist: Part 1: creature not nature. Studies in History and Philosophy of Science Part A, 31(4):613–643, December 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000224.

Cunningham:2001:RPD

[Cun01a] Andrew Cunningham. A reply to Peter Dear’s ‘Religion, science and natural philosophy: thoughts on Cunningham’s thesis’. Studies in History and Philosophy of Science Part A, 32(2):387– 391, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000048.

Cunningham:2001:SRT

[Cun01b] Andrew Cunningham. Science and religion in the Thirteenth Century revisited: the making of St Francis the proto-ecologist: Part 2: Nature not creature. Studies in History and Philosophy of Science Part A, 32(1):69–98, March 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000236.

Cunningham:2021:PCM

[Cun21] Bryon Cunningham. A prototypical conceptualization of mechanisms. Studies in History and Philosophy of Science Part A, 85 (??):79–91, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301801.

Cuomo:1995:FC

[Cuo95] Serafina Cuomo. A favourable conjuncture. Studies in History and Philosophy of Science Part A, 26(4):667–672, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000194.

Cuomo:2000:DRF

[Cuo00] Serafina Cuomo. Divide and rule: Frontinus and Roman landsurveying. Studies in History and Philosophy of Science Part A, 31(2):189–202, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810000011X.

Cuomo:2006:BG

[Cuo06] Serafina Cuomo. A beautiful game. Studies in History and Philosophy of Science Part A, 37(2):337–343, June 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810600029X.

Curtis:1989:IIE

[Cur89] Ronald Curtis. Institutional individualism and the emergence of scientific rationality. Studies in History and Philosophy of Science Part A, 20(1):77–113, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900356.

Curry:1992:NCC

[Cur92] Charles Curry. The naturalness of the cosmological constant in the general theory of relativity: a response to Ray. Studies in History and Philosophy of Science Part A, 23(4):657–660, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819290016Y. See [Ray90, Ray92].

Currie:2016:EAC

[Cur16] Adrian Currie. Ethnographic analogy, the comparative method, and archaeological special pleading. Studies in History and Philosophy of Science Part A, 55(??):84–94, February 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811500103X.

Curry:2018:CCC

[Cur18] Devin Sanchez Curry. Cartesian critters can’t remember. Studies in History and Philosophy of Science Part A, 69(??):72– 85, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302376.

Currie:2019:CCS

[Cur19a] Adrian Currie. Creativity, conservativeness and the social epistemology of science. Studies in History and Philosophy of Science Part A, 76(??):1–4, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301808.

Currie:2019:ERC

[Cur19b] Adrian Currie. Existential risk, creativity and well-adapted science. Studies in History and Philosophy of Science Part A, 76(??): 39–48, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117303278.

Cushing:1985:TJO

[Cus85] James T. Cushing. Is there just one possible world? Contingency vs the bootstrap. Studies in History and Philosophy of Science Part A, 16(1):31–48, March 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368185900068.

Cardano:1968:GAR

[CW68] Girolamo Cardano and T. Richard Witmer. The great art; or, The rules of algebra. MIT Press, Cambridge, MA, USA, 1968. xxiv + 267 pp. LCCN QA35 .C2661.

DAgostino:2000:ICL

[D’A00] Fred D’Agostino. Incommensurability and commensuration: lessons from (and to) ethico-political theory. Studies in History and Philosophy of Science Part A, 31(3):429–447, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000133.

Dacey:2015:AAL

[Dac15] Mike Dacey. Associationism without associative links: Thomas Brown and the associationist project. Studies in History and Philosophy of Science Part A, 54(??):31–40, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000941.

Dais:2019:DTT

[Dai19] Photis Dais. The double transfer of thermodynamics: From physics to chemistry and from Europe to America. Studies in History and Philosophy of Science Part A, 77(??):54–63, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300189.

Dancy:2006:TU

[Dan06] Jonathan Dancy. The thing to use. Studies in History and Philosophy of Science Part A, 37(1):58–61, March 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105001093.

Darden:1976:RSC

[Dar76] Lindley Darden. Reasoning in scientific change: Charles Darwin, Hugo de Vries, and the discovery of segregation. Studies in History and Philosophy of Science Part A, 7(2):127–169, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368176900145.

Darden:1996:GB

[Dar96] Lindley Darden. Generalizations in biology. Studies in History and Philosophy of Science Part A, 27(3):409–419, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819500050X.

Darling:2002:CDU

[Dar02] Karen Merikangas Darling. The complete Duhemian underdetermination argument: scientific language and practice. Studies in History and Philosophy of Science Part A, 33(3):511–533, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000274.

Darrigol:2003:NMH

[Dar03] Olivier Darrigol. Number and measure: Hermann von Helmholtz at the crossroads of mathematics, physics, and psychology. Studies in History and Philosophy of Science Part A, 34(3):515–573, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000438.

Darrigol:2007:HAS

[Dar07] Olivier Darrigol. A Helmholtzian approach to space and time. Studies in History and Philosophy of Science Part A, 38(3):528– 542, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000374.

Dardashti:2021:NGT

[Dar21a] Radin Dardashti. No-go theorems: What are they good for? Studies in History and Philosophy of Science Part A, 86(??): 47–55, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000054.

Darrigol:2021:CWT

[Dar21b] Olivier Darrigol. Can we trust Einstein’s accounts of the genesis of special relativity? Studies in History and Philosophy of Science Part A, 89(??):138–154, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000492.

Darrigol:2023:BRM

[Dar23] Olivier Darrigol. Book review: A Middle Way: a Non-Fundamental Approach to Many-Body Physics by Robert Batterman: Micromeso-macro: Batterman’s philosophical reflections on the mutual (in)dependence of scales in many-body systems. Studies in History and Philosophy of Science Part A, 97(??):121–122, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001273. See [Bat23].

Daston:1986:PTE

[Das86] Lorraine J. Daston. The physicalist tradition in early nineteenth century French geometry. Studies in History and Philosophy of Science Part A, 17(3):269–295, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900105.

Daub:1970:MD

[Dau70] Edward E. Daub. Maxwell’s demon. Studies in History and Philosophy of Science Part A, 1(3):213–227, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368170900105.

Davies:2003:NM

[Dav03] E. B. Davies. The Newtonian myth. Studies in History and Philosophy of Science Part A, 34(4):763–780, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000451.

Davies:2013:KIT

[Dav13] Alex Davies. Kuhn on incommensurability and theory choice. Studies in History and Philosophy of Science Part A, 44(4):571– 579, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113001064.

Davis:2020:AVL

[Dav20] Ian M. Davis. Antoni van Leeuwenhoek and measuring the invisible: the context of 16th and 17th century micrometry. Studies in History and Philosophy of Science Part A, 83(??):75– 85, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119303577.

Davey:2021:IBE

[Dav21a] Kevin Davey. Inference to the best explanation and Norton’s material theory of induction. Studies in History and Philosophy of Science Part A, 85(??):137–144, February 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301862.

Davies:2021:LCW

[Dav21b] Gail Davies. Locating the ‘culture wars’ in laboratory animal research: national constitutions and global competition. Studies in History and Philosophy of Science Part A, 89(??):177– 187, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001266.

Dawid:2021:RME

[Daw21] Richard Dawid. The role of meta-empirical theory assessment in the acceptance of atomism. Studies in History and Philosophy of Science Part A, 90(??):50–60, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001333.

Dawid:2022:MEC

[Daw22] Richard Dawid. Meta-empirical confirmation: Addressing three points of criticism. Studies in History and Philosophy of Science Part A, 93(??):66–71, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000371.

Dalton:1996:JDM

[DB96] Thomas C. Dalton and Victor W. Bergenn. John Dewey, Myrtle McGraw and logic: An unusual collaboration in the 1930s. Studies in History and Philosophy of Science Part A, 27(1):69– 107, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000259.

deBruin:2009:OGT

[dB09] Boudewijn de Bruin. Overmathematisation in game theory: pitting the Nash Equilibrium Refinement Programme against the Epistemic Programme. Studies in History and Philosophy of Science Part A, 40(3):290–300, September 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810900048X.

Duerr:2022:WRW

[DBM22] Patrick M. Duerr and Yemima Ben-Menahem. Why Reichenbach wasn’t entirely wrong, and Poincaré was almost right, about geometric conventionalism. Studies in History and Philosophy of Science Part A, 96(??):154–173, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812200156X.

deChadarevian:1993:GMD

[dC93] Soraya de Chadarevian. Graphical method and discipline: Selfrecording instruments in nineteenth-century physiology. Studies in History and Philosophy of Science Part A, 24(2):267– 291, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390048O.

daCosta:2002:MEF

[dC02] Palmira Fontes da Costa. The making of extraordinary facts: authentication of singularities of nature at the Royal Society of London in the first half of the eighteenth century. Studies in History and Philosophy of Science Part A, 33(2):265–288, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000122.

Douven:2009:FTJ

[DC09] Igor Douven and Stefaan E. Cuypers. Fricker on testimonial justification. Studies in History and Philosophy of Science Part A, 40 (1):36–44, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108001106.

deCourtenay:2010:EVA

[dC10] Nadine de Courtenay. The epistemological virtues of assumptions: towards a coming of age of Boltzmann and Meinong’s objections to ‘the prejudice in favour of the actual’? Studies in History and Philosophy of Science Part A, 41(1):41–57, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810900096X.

daCosta:2012:GER

[dC12] Palmira Fontes da Costa. Geographical expansion and the reconfiguration of medical authority: Garcia de Orta’s Colloquies on the Simples and Drugs of India (1563). Studies in History and Philosophy of Science Part A, 43(1):74–81, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000793.

deChadarevian:2013:TAR

[dC13] Soraya de Chadarevian. Things and the archives of recent sciences. Studies in History and Philosophy of Science Part A, 44 (4):634–638, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000472.

daCunha:2018:CDE

[dC18] Ivan Ferreira da Cunha. Constructing dystopian experience: a Neurath–Cartwrightian approach to the philosophy of social technology. Studies in History and Philosophy of Science Part A, 72 (??):41–48, December 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117303096.

deCourtenay:2017:EMU

[dCG17] Nadine de Courtenay and Fabien Grégis. The evaluation of measurement uncertainties and its epistemological ramifications. Studies in History and Philosophy of Science Part A, 65–66(??):21– 32, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301383.

Dunlap:2021:DVG

[DCJ^{+}21] Lucas Dunlap, Amanda Corris, Melissa Jacquart, Zvi Biener, and Angela Potochnik. Divergence of values and goals in participatory research. Studies in History and Philosophy of Science Part A, 88(??):284–291, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000996.

deCock:2024:NCP

[dCR24] Wessel de Cock and Carsten Reinhardt. Narratives of contingency and practices of comparing in the emergence of German molecular genetics (1958–1968). Studies in History and Philosophy of Science Part A, 107(??):118–127, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001080.

Darwin:1969:LLC

[DD69a] Charles Darwin and Sir Francis Darwin. The life and letters of Charles Darwin, including an autobiographical chapter, volume 102 of The Sources of science. Johnson Reprint Corp., New York, NY, USA, 1969. ???? pp. LCCN QH31 .D2A2 1969.

Drake:1969:MSC

[DD69b] Stillman Drake and I. E. (Israel Edward) Drabkin. Mechanics in sixteenth-century Italy; selections from Tartaglia, Benedetti, Guido Ubaldo, and Galileo, volume 13 of University of Wisconsin publications in medieval science. University of Wisconsin Press, Madison, WI, USA, 1969. ISBN 0-299-05100-5. xii + 428 pp. LCCN QA802 .D713.

Darwin:1972:LLC

[DD72] Charles Darwin and Sir Francis Darwin. The life and letters of Charles Darwin; including an autobiographical chapter, volume 17– 18 of The Works of Charles Darwin. AMS Press, New York, NY, USA, 1972. ISBN 0-404-08417-6 (vol. 1). ???? pp. LCCN QH31.D2 A2 1972.

DeCruz:2013:VED

[DD13] Helen De Cruz and Johan De Smedt. The value of epistemic disagreement in scientific practice. the case of Homo floresiensis. Studies in History and Philosophy of Science Part A, 44(2): 169–177, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000162.

DeRenzi:1996:CCI

[De 96a] Silvia De Renzi. Courts and conversions: Intellectual battles and natural knowledge in counter-reformation Rome. Studies in History and Philosophy of Science Part A, 27(4):429–449, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000585.

DeRenzi:1996:SPK

[De 96b] Silvia De Renzi. Secrecy, power and knowledge in early modern Italy. Studies in History and Philosophy of Science Part A, 27 (3):397–407, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000569.

DeGroot:2000:AAS

[De 00] J. De Groot. Aspects of Aristotelian statics in Galileo’s dynamics. Studies in History and Philosophy of Science Part A, 31(4):645– 664, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000194.

DeRenzi:2002:WBM

[De 02] Silvia De Renzi. Witnesses of the body: medico-legal cases in seventeenth-century Rome. Studies in History and Philosophy of Science Part A, 33(2):219–242, June 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000055.

DeWinter:2011:PAM

[De 11] Jan De Winter. A pragmatic account of mechanistic artifact explanation. Studies in History and Philosophy of Science Part A, 42 (4):602–609, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111000732.

DeLanghe:2014:CTM

[De 14] Rogier De Langhe. A comparison of two models of scientific progress. Studies in History and Philosophy of Science Part A, 46 (??):94–99, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000211.

DeKock:2016:HKR

[De 16] Liesbet De Kock. Helmholtz’s Kant revisited (once more). The all-pervasive nature of Helmholtz’s struggle with Kant’s Anschauung. Studies in History and Philosophy of Science Part A, 56(??): 20–32, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001570.

DeBianchi:2018:SWO

[De 18] Silvia De Bianchi. The stage on which our ingenious play is performed: Kant’s epistemology of weltkenntnis. Studies in History and Philosophy of Science Part A, 71(??):58–66, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300158.

DeBenedetto:2024:TCG

[De 24] Matteo De Benedetto. Theoretical concepts as goal-derived concepts. Studies in History and Philosophy of Science Part A, 107 (??):82–91, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124001249.

Duerr:2021:PMT

[DE21] Patrick M. Duerr and Alexander Ehmann. The physics and metaphysics of Tychistic Bohmian Mechanics. Studies in History and Philosophy of Science Part A, 90(??):168–183, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001540.

Dear:1987:JMS

[Dea87] Peter Dear. Jesuit mathematical science and the reconstitution of experience in the early seventeenth century. Studies in History and Philosophy of Science Part A, 18(2):133–175, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900161.

Dear:2001:RSN

[Dea01a] Peter Dear. Religion, science and natural philosophy: thoughts on Cunningham’s thesis. Studies in History and Philosophy of Science Part A, 32(2):377–386, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000036.

Dear:2001:RAC

[Dea01b] Peter Dear. Reply to Andrew Cunningham. Studies in History and Philosophy of Science Part A, 32(2):393–395, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810100005X.

Dear:2003:OSA

[Dea03] Peter Dear. Openness, secrecy, authorship: Technical arts and the culture of knowledge from antiquity to the Renaissance. Studies in History and Philosophy of Science Part A, 34(4):821–828, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000682.

Dear:2012:HEE

[Dea12] Peter Dear. Horizontal explanation in the enlightenment. Studies in History and Philosophy of Science Part A, 43(1):221– 223, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811100104X.

Dellsen:2016:SPK

[Del16] Finnur Dellsén. Scientific progress: Knowledge versus understanding. Studies in History and Philosophy of Science Part A, 56(??): 72–83, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116000042.

Dellsen:2017:RRB

[Del17] Finnur Dellsén. Reactionary responses to the Bad Lot Objection. Studies in History and Philosophy of Science Part A, 61(??):32– 40, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300486.

Dellsen:2019:SSR

[Del19] Finnur Dellsén. Should scientific realists embrace theoretical conservatism? Studies in History and Philosophy of Science Part A, 76(??):30–38, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117303187.

Delgarm:2022:PUI

[Del22] Majid Heydari Delgarm. A previously-unknown Iranian treatise on a terrestrial globe. Studies in History and Philosophy of Science Part A, 95(??):204–214, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812200108X.

Dellsen:2023:SPW

[Del23] Finnur Dellsén. Scientific progress: By-whom or for-whom? Studies in History and Philosophy of Science Part A, 97(??):20–28, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812200173X.

Dempsey:2006:WFI

[Dem06] Liam Dempsey. Written in the flesh: Isaac Newton on the mindbody relation. Studies in History and Philosophy of Science Part A, 37(3):420–441, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000495.

Demir:2008:IWT

[Dem08] Ipek Demir. Incommensurabilities in the work of Thomas Kuhn. Studies in History and Philosophy of Science Part A, 39(1):133– 142, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000933.

Denis:2000:EOP

[Den00] Andy Denis. Epistemology, observed particulars and providentialist assumptions: the fact in the history of political economy. Studies in History and Philosophy of Science Part A, 31(2):353– 361, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000054.

Descartes:1644:PPL

[Des44] René Descartes. Principia philosophiæ. (Latin) [Philosophical Principles]. Ludovicum Elzevirium, Amstelodami, The Netherlands, 1644. xxii + 310 pp.

Descartes:1649:PPE

[Des49] René Descartes. Les principes de la philosophie, escrits en Latin par R-D-et traduits en françois par un de ses amis. (French) [The principles of philosophy, written in Latin by René Descartes and translated to French by one of his friends]. ????, Paris, France, 1649. Translation by l’abbé Picot from Latin to French.

Dethier:2022:CST

[Det22] Corey Dethier. Calibrating statistical tools: Improving the measure of Humanity’s influence on the climate. Studies in History and Philosophy of Science Part A, 94(??):158–166, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000917.

Doody:2012:SSA

[DFT12] Aude Doody, Sabine Föllinger, and Liba Taub. Structures and strategies in ancient Greek and Roman technical writing: an introduction. Studies in History and Philosophy of Science Part A, 43(2):233–236, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001324.

deGrefte:2020:EBM

[dG20] Job de Grefte. Epistemic benefits of the material theory of induction. Studies in History and Philosophy of Science Part A, 84 (??):99–105, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301655.

Davidson:1969:WOE

[DH69] Donald Davidson and Jaakko Hintikka. Words and objections. Essays on the work of W. V. Quine. Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1969. vii + 366 pp. LCCN B945.Q54 D38 1969.

Douven:2022:NEB

[DH22] Igor Douven and Rainer Hegselmann. Network effects in a bounded confidence model. Studies in History and Philosophy of Science Part A, 94(??):56–71, August 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812200070X.

Dhein:2023:CMD

[Dhe23] Kelle Dhein. The cognitive map debate in insects: a historical perspective on what is at stake. Studies in History and Philosophy of Science Part A, 98(??):62–79, April 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001856.

Dietrich:2021:RLC

[DHL21] Michael R. Dietrich, Oren Harman, and Ehud Lamm. Richard Lewontin and the “complications of linkage”. Studies in History and Philosophy of Science Part A, 88(??):237–244, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000856.

DeLanghe:2014:IPS

[DHS14] Rogier De Langhe, Stephan Hartmann, and Jan Sprenger. Introduction: the progress of science. Studies in History and Philosophy of Science Part A, 46(??):54, June 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811400017X.

DiFate:2011:NRE

[Di 11] Victor Joseph Di Fate. Is Newton a ‘radical empiricist’ about method? Studies in History and Philosophy of Science Part A, 42(1):28–36, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110001032.

Dicken:2006:CCE

[Dic06] Paul Dicken. Can the constructive empiricist be a nominalist? quasi-truth, commitment and consistency. Studies in History and Philosophy of Science Part A, 37(2):191–209, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000240.

Dicken:2008:CMA

[Dic08] Paul Dicken. Conditions may apply. Studies in History and Philosophy of Science Part A, 39(2):290–293, June 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000368.

Dicken:2013:NBR

[Dic13] Paul Dicken. Normativity, the base-rate fallacy, and some problems for retail realism. Studies in History and Philosophy of Science Part A, 44(4):563–570, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000964.

Dickinson:2025:TCC

[Dic25] Jon Dickinson. Temperature changes: the conceptual realignment of a quantity term. Studies in History and Philosophy of Science Part A, 109(??):47–57, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001675.

Diez:1997:HYNa

[Díe97a] JoséA. Díez. A hundred years of numbers. an historical introduction to measurement theory 1887–1990: Part I: The formation period. two lines of research: Axiomatics and real morphisms, scales and invariance. Studies in History and Philosophy of Science Part A, 28(1):167–185, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000143.

Diez:1997:HYNb

[Díe97b] JoséA. Díez. A hundred years of numbers. an historical introduction to measurement theory 1887–1990: Part II: Suppes and the mature theory. representation and uniqueness. Studies in History and Philosophy of Science Part A, 28(2):237– 265, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000155.

Diez:2007:FST

[Díe07] Jose Díez. Falsificationism and the structure of theories: the Popper–Kuhn controversy about the rationality of normal science. Studies in History and Philosophy of Science Part A, 38(3):543– 554, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000386.

Diez:2011:PSI

[Díe11] José A. Díez. On Popper’s strong inductivism (or strongly inconsistent anti-inductivism). Studies in History and Philosophy of Science Part A, 42(1):105–116, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001020.

Dietrich:2024:BFW

[Die24] Michael R. Dietrich. Book forum: What is Regeneration? by Jane Maienschein and Kate MacCord: Rethinking regeneration. Studies in History and Philosophy of Science Part A, 103(??):176– 177, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000037.

Dilworth:1990:EVR

[Dil90] Craig Dilworth. Empiricism vs. realism: High points in the debate during the past 150 years. Studies in History and Philosophy of Science Part A, 21(3):431–462, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819090004R.

Dion:2013:PDI

[Dio13] Sonia Maria Dion. Pierre Duhem and the inconsistency between instrumentalism and natural classification. Studies in History and Philosophy of Science Part A, 44(1):12–19, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000726.

Dion:2018:NCP

[Dio18] Sonia Maria Dion. Natural classification and Pierre Duhem’s historical work: Which relationships? Studies in History and Philosophy of Science Part A, 69(??):34–39, June 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301747.

Dipert:2006:CED

[Dip06] Randall R. Dipert. Coherence and engineering design. Studies in History and Philosophy of Science Part A, 37(1):152– 158, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105001196.

DiSalle:2013:TMN

[DiS13] Robert DiSalle. The transcendental method from Newton to Kant. Studies in History and Philosophy of Science Part A, 44(3):448– 456, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000817.

deJong:1997:KTG

[dJ97] Willem R. de Jong. Kant’s theory of geometrical reasoning and the analytic-synthetic distinction. On Hintikka’s interpretation of Kant’s philosophy of mathematics. Studies in History and Philosophy of Science Part A, 28(1):141–166, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368196000064.

Duncan:2014:TOS

[DJ14] Anthony Duncan and Michel Janssen. The trouble with orbits: the Stark effect in the old and the new quantum theory. Studies in History and Philosophy of Science Part A, 48 (part A)(??):68– 83, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000847.

Dupre:2009:IKC

[DK09] Sven Dupré and Michael Korey. Inside the Kunstkammer: the circulation of optical knowledge and instruments at the Dresden Court. Studies in History and Philosophy of Science Part A, 40 (4):405–420, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810900065X.

DeBianchi:2018:IKP

[DK18] Silvia De Bianchi and Katharina Kraus. Introduction to Kant’s philosophy of science: Bridging the gap between the natural and the human sciences. Studies in History and Philosophy of Science Part A, 71(??):1–5, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118302097.

DiMarco:2022:SIH

[DK22] Marina DiMarco and Kareem Khalifa. Sins of inquiry: How to criticize scientific pursuits. Studies in History and Philosophy of Science Part A, 92(??):86–96, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121002089.

deLucas:2009:IBS

[dL09] Juan V. Mayoral de Lucas. Intensions, belief and science: Kuhn’s early philosophical outlook (1940–1945). Studies in History and Philosophy of Science Part A, 40(2):175–184, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000223.

Douglas:2013:SFW

[DM13] Heather Douglas and P. D. Magnus. State of the field: Why novel prediction matters. Studies in History and Philosophy of Science Part A, 44(4):580–589, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000198.

Descartes:1983:PP

[DMM83] René Descartes, Valentine Rodger Miller, and Reese P. Miller. Principles of philosophy, volume 24 of Synthese historical library: texts and studies in the history of logic and philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1451-7. xxviii + 325 pp. LCCN B1863.E53 M54 1983.

Dalitz:2000:FNS

[DN00] R. H. (Richard Henry) Dalitz and Michael Nauenberg, editors. The foundations of Newtonian scholarship. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-3920-3 (hardcover), 981-02-4044-9. xviii + 242 pp. LCCN QC16.N7 F68 2000. URL http://www.loc.gov/

catdir/toc/fy042/99088303.html.

deOliveira:2012:KGN

[dO12] J. C. Pinto de Oliveira. Kuhn and the genesis of the “new historiography of science”. Studies in History and Philosophy of Science Part A, 43(1):115–121, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001026.

deOlano:2022:TAS

[dOFGB22] Pablo Ruiz de Olano, James D. Fraser, Rocco Gaudenzi, and Alexander S. Blum. Taking approximations seriously: the cases of the Chew and Nambu–Jona–Lasinio models. Studies in History and Philosophy of Science Part A, 93(??):82–95, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000449.

Domondon:2009:KPS

[Dom09a] Andrew T. Domondon. Kuhn, Popper, and the Superconducting Supercollider. Studies in History and Philosophy of Science Part A, 40(3):301–314, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000478.

Domski:2009:IMC

[Dom09b] Mary Domski. The intelligibility of motion and construction: Descartes’ early mathematics and metaphysics, 1619–1637. Studies in History and Philosophy of Science Part A, 40(2):119– 130, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000193.

Domski:2013:KNP

[Dom13] Mary Domski. Kant and Newton on the a priori necessity of geometry. Studies in History and Philosophy of Science Part A, 44 (3):438–447, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000829.

Donhauser:2021:HMV

[Don21] Justin Donhauser. How to make value-driven climate science for policy more ethical. Studies in History and Philosophy of Science Part A, 89(??):31–40, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000984.

Dorling:1970:MAA

[Dor70] Jon Dorling. Maxwell’s attempts to arrive at non-speculative foundations for the kinetic theory. Studies in History and Philosophy of Science Part A, 1(3):229–248, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900117.

Dorling:1974:HCD

[Dor74] Jon Dorling. Henry Cavendish’s deduction of the electrostatic inverse square law from the result of a single experiment. Studies in History and Philosophy of Science Part A, 4(4):327–348, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368174900089.

Dorling:1979:BPM

[Dor79] Jon Dorling. Bayesian personalism, the methodology of scientific research programmes, and Duhem’s problem. Studies in History and Philosophy of Science Part A, 10(3):177–187, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368179900062.

Dorries:1994:BSR

[Dör94] Matthias Dörries. Balances, spectroscopes, and the reflexive nature of experiment. Studies in History and Philosophy of Science Part A, 25(1):1–36, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900183.

Douard:1995:JMV

[Dou95] John W. Douard. E.-J. Marey’s visual rhetoric and the graphic decomposition of the body. Studies in History and Philosophy of Science Part A, 26(2):175–204, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819400039C.

Douven:2010:SPD

[Dou10] Igor Douven. Simulating peer disagreements. Studies in History and Philosophy of Science Part A, 41(2):148–157, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000221.

Douglas:2014:PSP

[Dou14] Heather Douglas. Pure science and the problem of progress. Studies in History and Philosophy of Science Part A, 46(??): 55–63, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000132.

Douven:2020:ERE

[Dou20] Igor Douven. The ecological rationality of explanatory reasoning. Studies in History and Philosophy of Science Part A, 79(??):1– 14, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118303285.

Dougherty:2021:AAN

[Dou21] John Dougherty. I ain’t afraid of no ghost. Studies in History and Philosophy of Science Part A, 88(??):70–84, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000698.

Douglas:2022:DFS

[Dou22] Bronwen Douglas. Darwin and the French: the species question and ‘man’ in Oceania. Studies in History and Philosophy of Science Part A, 91(??):168–180, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001734.

deRegt:1998:ESS

[dR98] Henk W. de Regt. Explaining the splendour of science. Studies in History and Philosophy of Science Part A, 29(1):155– 165, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197000265.

deRidder:2006:MAE

[dR06] Jeroen de Ridder. Mechanistic artefact explanation. Studies in History and Philosophy of Science Part A, 37(1):81–96, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105001123.

deRegt:2013:UEL

[dR13] Henk W. de Regt. Understanding and explanation: Living apart together? Studies in History and Philosophy of Science Part A, 44 (3):505–509, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112001124.

deRidder:2022:HTS

[dR22] Jeroen de Ridder. How to trust a scientist. Studies in History and Philosophy of Science Part A, 93(??):11–20, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000346.

Drake:1975:FFA

[Dra75] Stillman Drake. Free fall from Albert of Saxony to Honoré Fabri. Studies in History and Philosophy of Science Part A, 5(4):347– 366, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900072.

Drake:1976:FRI

[Dra76] Stillman Drake. A further reappraisal of impetus theory: Buridan, Benedetti, and Galileo. Studies in History and Philosophy of Science Part A, 7(4):319–336, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368176900108.

Drake:1978:PGS

[Dra78] Stillman Drake. Ptolemy, Galileo, and scientific method. Studies in History and Philosophy of Science Part A, 9(2):99–115, June 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936817890002X.

Drake:1986:GPP

[Dra86] Stillman Drake. Galileo’s pre-Paduan writings: Years, sources, motivations. Studies in History and Philosophy of Science Part A, 17 (4):429–448, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368186900038.

Drake:1987:GSF

[Dra87] Stillman Drake. Galileo’s steps to full Copernicanism, and back. Studies in History and Philosophy of Science Part A, 18(1):93– 105, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368187900136.

Drake:1989:HGG

[Dra89] Stillman Drake. Hipparchus–Geminus–Galileo. Studies in History and Philosophy of Science Part A, 20(1):47–56, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900332.

Drayson:2018:RVM

[Dra18] Zoe Drayson. The realizers and vehicles of mental representation. Studies in History and Philosophy of Science Part A, 68(??): 80–87, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118300098.

Dresow:2020:HPS

[Dre20] Max Dresow. History and philosophy of science after the practiceturn: From inherent tension to local integration. Studies in History and Philosophy of Science Part A, 82(??):57–65, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119301554.

Driscoll:2024:CHN

[Dri24] Catherine Driscoll. Can human nature be saved? Studies in History and Philosophy of Science Part A, 103(??):39–45, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001620.

deRonde:2023:RQE

[dRM23] Christian de Ronde and César Massri. Relational quantum entanglement beyond non-separable and contextual relativism. Studies in History and Philosophy of Science Part A, 97(??):68–78, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001728.

deSoysa:2009:USC

[dS09] Kemal de Soysa. An unusual silver celestial planisphere in the Whipple Museum. Studies in History and Philosophy of Science Part A, 40(4):421–430, December 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000739.

Dawes:2018:NS

[DS18] Gregory W. Dawes and Tiddy Smith. The naturalism of the sciences. Studies in History and Philosophy of Science Part A, 67 (??):22–31, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117300389.

Donhauser:2019:KTT

[DS19] Justin Donhauser and Jamie Shaw. Knowledge transfer in theoretical ecology: Implications for incommensurability, voluntarism, and pluralism. Studies in History and Philosophy of Science Part A, 77(??):11–20, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118301730.

Dutz:2021:BGD

[DS21] Jonah Dutz and Dirk Schlimm. Babbage’s guidelines for the design of mathematical notations. Studies in History and Philosophy of Science Part A, 88(??):92–101, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000315.

Ducheyne:2005:NNP

[Duc05] Steffen Ducheyne. Newton’s notion and practice of unification. Studies in History and Philosophy of Science Part A, 36(1):61– 78, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104001207.

Ducheyne:2010:WTR

[Duc10] Steffen Ducheyne. Whewell’s tidal researches: scientific practice and philosophical methodology. Studies in History and Philosophy of Science Part A, 41(1):26–40, March 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000971.

Ducheyne:2011:NAD

[Duc11] Steffen Ducheyne. Newton on action at a distance and the cause of gravity. Studies in History and Philosophy of Science Part A, 42(1):154–159, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000828.

Duerr:2021:TEC

[Due21] Patrick M. Duerr. Theory (in-)equivalence and conventionalism in f(R) gravity. Studies in History and Philosophy of Science Part A, 88(??):10–29, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000510.

Due:2022:PIT

[Due22] Austin Due. Are ‘phase IV’ trials exploratory or confirmatory experiments? Studies in History and Philosophy of Science Part A, 95(??):126–133, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812200111X.

Duncan:2010:LHM

[Dun10] Stewart Duncan. Leibniz on Hobbes’s materialism. Studies in History and Philosophy of Science Part A, 41(1):11–18, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000922.

Dunsch:2012:ARR

[Dun12] Boris Dunsch. Arte rates reguntur: Nautical handbooks in antiquity? Studies in History and Philosophy of Science Part A, 43(2):270–283, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811100135X.

Dunlop:2013:MMN

[Dun13] Katherine Dunlop. Mathematical method and Newtonian science in the philosophy of Christian Wolff. Studies in History and Philosophy of Science Part A, 44(3):457–469, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000830.

Dunlap:2022:ITI

[Dun22] Lucas Dunlap. Is the Information-Theoretic Interpretation of Quantum Mechanics an ontic structural realist view? Studies in History and Philosophy of Science Part A, 91(??):41–48, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001837.

Dupre:1994:PBB

[Dup94] John Dupré. The philosophical basis of biological classification. Studies in History and Philosophy of Science Part A, 25(2):271– 279, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900329.

Duppe:2015:BCB

[Düp15] Till Düppe. Border cases between autonomy and relevance: Economic sciences in Berlin — a natural experiment. Studies in History and Philosophy of Science Part A, 51(??):22–32, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000205.

Durant:2013:SHG

[Dur13] John Durant. “Whatever happened to the Genomatron?” documenting a 21st century science. Studies in History and Philosophy of Science Part A, 44(4):676–682, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000484.

Demarest:2022:KTS

[DvdB22] Boris Demarest and Hein van den Berg. Kant’s theory of scientific hypotheses in its historical context. Studies in History and Philosophy of Science Part A, 92(??):12–19, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000115.

Duerr:2023:MRM

[DW23] Patrick M. Duerr and William J. Wolf. Methodological reflections on the MOND/dark matter debate. Studies in History and Philosophy of Science Part A, 101(??):1–23, October 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000936.

Diekmann:2014:MFR

[DZ14] Sven Diekmann and Sjoerd D. Zwart. Modeling for fairness: a Rawlsian approach. Studies in History and Philosophy of Science Part A, 46(??):46–53, June 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113001234.

Eagle:2004:CTC

[Eag04] Antony Eagle. A causal theory of chance? Studies in History and Philosophy of Science Part A, 35(4):883–890, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000901.

Eagleton:2007:COA

[Eag07] Catherine Eagleton. ‘Chaucer’s own astrolabe’: text, image and object. Studies in History and Philosophy of Science Part A, 38(2): 303–326, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000155. See [CS78].

Earman:1990:BB

[Ear90] John Earman. Bayes’ Bayesianism. Studies in History and Philosophy of Science Part A, 21(3):351–370, September 1990. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936819090001O.

Earman:2020:QSM

[Ear20] John Earman. Quantum sidelights on the material theory of induction. Studies in History and Philosophy of Science Part A, 82(??):9–16, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301268.

Easton:1972:BRB

[Eas72] Joy B. Easton. Book review: The great art or the rules of algebra: Girolamo Cardano, Translated and edited by T. Richard Witmer, with a foreword by Øystein Øre. MIT Press: Cambridge, Mass., 1968. xxiv + 267 pp. $10.00. Studies in History and Philosophy of Science Part A, 2(4):363–368, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368172900040. See [CW68].

Eaton:2023:BRC

[Eat23] William Eaton. Book review: The Chemical Philosophy of Robert Boyle: Mechanicism, Chymical Atoms, and Emergence by Marina Paola Banchetti-Robino-Robino: Chymical Emergence in the Philosophy of Robert Boyle. Studies in History and Philosophy of Science Part A, 97(??):135–136, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001613.

El-Bizri:2011:GPA

[EB11] Nader El-Bizri. The groundbreaking physics of Averroës. Studies in History and Philosophy of Science Part A, 42(1):210– 214, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000889.

Elawani:2025:AEM

[EC25] Jeffrey Elawani and Filippo Costantini. The art of estimation and the mathematization of force in Leibniz. Studies in History and Philosophy of Science Part A, 110(??):65–75, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812400164X.

Edelheit:2009:FPT

[Ede09] Amos Edelheit. Francesco Patrizi’s two books on space: geometry, mathematics, and dialectic beyond Aristotelian science. Studies in History and Philosophy of Science Part A, 40(3):243–257, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000508.

Edgar:2008:PNE

[Edg08] Scott Edgar. Paul Natorp and the emergence of anti-psychologism in the nineteenth century. Studies in History and Philosophy of Science Part A, 39(1):54–65, March 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000866.

Edington:2022:JAR

[Edi22] Claire Edington. Joelle Abi-Rached. ’Asfūriyyeh: a history of madness, modernity and war in the Middle East: Taking the longue durée view. Studies in History and Philosophy of Science Part A, 93(??):225–226, June 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000097. See reply [AR22].

Earman:1978:LTE

[EG78] John Earman and Clark Glymour. Lost in the tensors: Einstein’s struggles with covariance principles 1912–1916. Studies in History and Philosophy of Science Part A, 9(4):251–278, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368178900080.

Earman:1980:GRS

[EG80] John Earman and Clark Glymour. The gravitational red shift as a test of General Relativity: History and analysis. Studies in History and Philosophy of Science Part A, 11(3):175–214, September 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900254.

Egan:2010:CMM

[Ega10] Frances Egan. Computational models: a modest role for content. Studies in History and Philosophy of Science Part A, 41(3):253– 259, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000415.

Egerton:1970:BRBa

[Ege70a] Frank N. Egerton. Book review: Mechanics in sixteenth-century Italy: Stillman Drake and I. E. Drabkin. University of Wisconsin Press: Madison, 1969. xii + 428 pp. $12.50. Studies in History and Philosophy of Science Part A, 1(2):161–175, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936817090004X. See [DD69b].

Egerton:1970:BRBb

[Ege70b] Frank N. Egerton. Book review: Vestiges of the natural history of creation: Robert Chambers, (London: John Churchill, 1844). vi + 390 pp. Fascimile reprint with introduction by Gavin de Beer, pp. 8–36, and bibliographical note by J. L. Madden, pp. 37–38. Leicester and New York: The Victorian Library of Leicester University Press and Humanities Press, 1969. 50s. Studies in History and Philosophy of Science Part A, 1(2):176– 183, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368170900051. See [Cha44, Cha69].

Egerton:1971:BRB

[Ege71] Frank N. Egerton. Book review: The triumph of the Darwinian method: Michael T. Ghiselin, Berkeley and Los Angeles: University of California Press, 1969. vi + 287 pp. $7.50. Studies in History and Philosophy of Science Part A, 2(3):281–286, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936817190046X. See [Ghi69a, Ghi69b,

Ghi72, Ghi84, Ghi03].

El-Hani:2022:BDB

[EHPL22] Charbel N. El-Hani, Luana Poliseli, and David Ludwig. Beyond the divide between indigenous and academic knowledge: Causal and mechanistic explanations in a Brazilian fishing community. Studies in History and Philosophy of Science Part A, 91(??):296– 306, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001783.

Eigi:2019:HTA

[Eig19] Jaana Eigi. How to think about shared norms and pluralism without circularity: a reply to Anna Leuschner. Studies in History and Philosophy of Science Part A, 75(??):51–56, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368118302413. See [Leu12].

Eisenthal:2021:HEU

[Eis21] Joshua Eisenthal. Hertz’s Mechanics and a unitary notion of force. Studies in History and Philosophy of Science Part A, 90(??):226– 234, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001606.

Ekbote:2024:ERC

[Ekb24] Anand Ekbote. Euclidean rigor and the curious case of the (missing) reflex angle. Studies in History and Philosophy of Science Part A, 108(??):10–18, December 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001225.

Elder:2024:IEM

[Eld24] Jamee Elder. Independent evidence in multi-messenger astrophysics. Studies in History and Philosophy of Science Part A, 104(??):119–129, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000165.

Elder:2025:DDG

[Eld25] Jamee Elder. On the “direct detection” of gravitational waves. Studies in History and Philosophy of Science Part A, 110(??): 1–12, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368125000020.

Elgin:2022:BRL

[Elg22] Catherine Z. Elgin. Book review: Leveraging Distortions: Explanation, Idealization, and Universality in Science, by Collin Rice: Universality, Understanding, and Realism. Studies in History and Philosophy of Science Part A, 95(??):228–229, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000875. See [Ode22, Jhu22, Pin22] and

reply [Ric22].

Eliasmith:2010:HWO

[Eli10] Chris Eliasmith. How we ought to describe computation in the brain. Studies in History and Philosophy of Science Part A, 41 (3):313–320, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000336.

Elkana:1974:BRB

[Elk74] Yehuda Elkana. Book review: The Annus mirabilis of Sir Isaac Newton: 1666–1966: R. Palter, ed., Cambridge, Mass.: MIT Press, 1971. 351 pp. $15.00. Studies in History and Philosophy of Science Part A, 5(1):87–93, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368174900211. See [Pal71].

Ellis:1976:EF

[Ell76] Brian Ellis. The existence of forces. Studies in History and Philosophy of Science Part A, 7(2):171–185, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368176900157.

Ellis:1992:SP

[Ell92] Brian Ellis. Scientific Platonism. Studies in History and Philosophy of Science Part A, 23(4):665–679, December 1992. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368192900182.

Elliott:2012:EMI

[Ell12] Kevin C. Elliott. Epistemic and methodological iteration in scientific research. Studies in History and Philosophy of Science Part A, 43(2):376–382, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000039.

Elliott:2013:DVI

[Ell13] Kevin C. Elliott. Douglas on values: From indirect roles to multiple goals. Studies in History and Philosophy of Science Part A, 44 (3):375–383, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000423.

Elliott:2021:VLT

[Ell21] Kevin C. Elliott. The value-ladenness of transparency in science: Lessons from Lyme disease. Studies in History and Philosophy of Science Part A, 88(??):1–9, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000327.

Elwick:2012:LHS

[Elw12] James Elwick. Layered history: Styles of reasoning as stratified conditions of possibility. Studies in History and Philosophy of Science Part A, 43(4):619–627, December 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000465.

Engelhard:2018:PGN

[Eng18] Kristina Engelhard. The problem of grounding natural modality in Kant’s account of empirical laws of nature. Studies in History and Philosophy of Science Part A, 71(??):24–34, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300237.

Ereshefsky:1991:BRB

[Ere91] Marc Ereshefsky. Book review: The metaphysics of evolution: David Hull, (Albany, NY: State University of New York Press, 1989), viii + 331 pp., ISBN 0-7914-0211-8 Hardback $73.50, Paperback $24.95. Michael Ruse (ed.), What the Philosophy of Biology Is: Essays Dedicated to David Hull (Dordrecht: Kluwer Academic Publishers, 1989), ix + 337 pp., ISBN 90-247-3778-8 Hardback Dfl 180.00/$99.00/£59.00. Studies in History and Philosophy of Science Part A, 22(3):525–532, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190008G. See [Hul89, Rus89].

Eagleton:2006:CCG

[ES06] Catherine Eagleton and Matthew Spencer. Copying and conflation in Geoffrey Chaucer’s Treatise on the astrolabe: a stemmatic analysis using phylogenetic software. Studies in History and Philosophy of Science Part A, 37(2):237–268, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000161. See [CS78].

Easterby-Smith:2012:TTT

[ES12] Sarah Easterby-Smith. Thinking through things. Studies in History and Philosophy of Science Part A, 43(1):208–212, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000781.

Escobar:2008:KTS

[Esc08] Jorge M. Escobar. Kepler’s theory of the soul: a study on epistemology. Studies in History and Philosophy of Science Part A, 39(1): 15–41, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000842.

Esposito:1979:RPN

[Esp79] Joseph L. Esposito. Reichenbach’s philosophy of nature. Studies in History and Philosophy of Science Part A, 10(3):189–200, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900074.

Ereshefsky:2020:HE

[ET20] Marc Ereshefsky and Derek Turner. Historicity and explanation. Studies in History and Philosophy of Science Part A, 80(??): 47–55, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118302693.

Evans:2007:SNP

[Eva07] Robert Evans. Social networks and private spaces in economic forecasting. Studies in History and Philosophy of Science Part A, 38(4):686–697, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000611.

Evans:2010:ACS

[Eva10] Michael S. Evans. Achieving continuity: a story of stellar magnitude. Studies in History and Philosophy of Science Part A, 41(1): 86–94, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000934.

Everett:2018:KAM

[Eve18] Jonathan Everett. A Kantian account of mathematical modelling and the rationality of scientific theory change: the role of the equivalence principle in the development of general relativity. Studies in History and Philosophy of Science Part A, 71(??):

45–57, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300195.

Epperson:2013:FRR

[EZ13] Michael Epperson and Elias Zafiris. Foundations of relational realism: a topological approach to quantum mechanics and the philosophy of nature. Lexington Books, Lanham, MD, USA, 2013. ISBN 0-7391-8032-0 (hardcover), 0-7391-8033-9 (electronic). xx + 419 pp. LCCN QC174.17.M35 E67 2013.

Frost-Arnold:2013:MTS

[FA13] Karen Frost-Arnold. Moral trust & scientific collaboration. Studies in History and Philosophy of Science Part A, 44(3):301–310, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000319.

Fabry:2019:PBC

[Fab19] Lucie Fabry. Phenomenotechnique: Bachelard’s critical inheritance of conventionalism. Studies in History and Philosophy of Science Part A, 75(??):34–42, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368118301523.

Fagan:2015:CEB

[Fag15] Melinda Bonnie Fagan. Collaborative explanation and biological mechanisms. Studies in History and Philosophy of Science Part A, 52(??):67–78, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000394.

Fagan:2016:GMH

[Fag16] Melinda Bonnie Fagan. Generative models: Human embryonic stem cells and multiple modeling relations. Studies in History and Philosophy of Science Part A, 56(??):122–134, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001417.

Faglia:2024:NSL

[Fag24] Paolo Faglia. Non-separability, locality and criteria of reality: a reply to Waegell and McQueen. Studies in History and Philosophy of Science Part A, 106(??):43–53, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000529.

Fairweather:2012:EVG

[Fai12] Abrol Fairweather. The epistemic value of good sense. Studies in History and Philosophy of Science Part A, 43(1):139– 146, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001075.

Falk:1986:WG

[Fal86] Raphael Falk. What is a gene? Studies in History and Philosophy of Science Part A, 17(2):133–173, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900245.

Falkenstein:1998:DES

[Fal98] Lorne Falkenstein. A double edged sword? Kant’s refutation of Mendelssohn’s proof of the immortality of the soul and its implications for his theory of matter. Studies in History and Philosophy of Science Part A, 29(4):561–588, December 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198000223.

Falconer:2017:NAM

[Fal17] Isobel Falconer. No actual measurement … was required: Maxwell and Cavendish’s null method for the inverse square law of electrostatics. Studies in History and Philosophy of Science Part A, 65–66(??):74–86, October/December 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301188.

Falkenburg:2018:KSA

[Fal18] Brigitte Falkenburg. Kant and the scope of the analytic method. Studies in History and Philosophy of Science Part A, 71(??):13– 23, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300122.

Falkenburg:2024:CSD

[Fal24] Brigitte Falkenburg. Computer simulation in data analysis: a case study from particle physics. Studies in History and Philosophy of Science Part A, 105(??):99–108, June 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000530.

Fang:2024:DPM

[Fan24] Wei Fang. Design principles as minimal models. Studies in History and Philosophy of Science Part A, 105(??):50–58, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000517.

Fara:1997:SH

[Far97] Patricia Fara. Scientific heritage. Studies in History and Philosophy of Science Part A, 28(1):187–195, March 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368196000015.

Fara:2007:HDH

[Far07] Patricia Fara. Hidden depths: Halley, hell and other people. Studies in History and Philosophy of Science Part A, 38(3):570–583, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000398.

Faulkner:2008:CWA

[Fau08] Paul Faulkner. Can we agree to disagree? Studies in History and Philosophy of Science Part A, 39(2):282–285, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000344.

Favrholdt:1991:RBH

[Fav91] David Favrholdt. Remarks on the Bohr–Høffding relationship. Studies in History and Philosophy of Science Part A, 22(3):399– 414, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190002A.

Faye:1988:BHR

[Fay88] Jan Faye. The Bohr–Høffding relationship reconsidered. Studies in History and Philosophy of Science Part A, 19(3):321–346, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900039.

Fay:2024:MPM

[Fay24a] Jonathan Fay. Mach’s principle and Mach’s hypotheses. Studies in History and Philosophy of Science Part A, 103(??):58– 68, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001644.

Fay:2024:RMD

[Fay24b] Jonathan Fay. On the relativity of magnitudes: Delboeuf’s forgotten contribution to the 19th century problem of space. Studies in History and Philosophy of Science Part A, 106(??):165– 176, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001031.

Fernandez-Beanato:2020:CDS

[FB20] Damian Fernandez-Beanato. Cicero’s demarcation of science: a report of shared criteria. Studies in History and Philosophy of Science Part A, 83(??):97–102, October 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119301815.

Formosinho:2022:EBM

[FBW22] Joana Formosinho, Adam Bencard, and Louise Whiteley. Environmentality in biomedicine: microbiome research and the perspectival body. Studies in History and Philosophy of Science Part A, 91 (??):148–158, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001825.

Fankhauser:2021:HUW

[FD21] Johannes Fankhauser and Patrick M. Dürr. How (not) to understand weak measurements of velocities. Studies in History and Philosophy of Science Part A, 85(??):16–29, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812030203X.

Fedyk:2023:NSS

[Fed23] Mark Fedyk. Nursing science as the study of how to reconcile behavioral messiness with clinical norms and ideals. Studies in History and Philosophy of Science Part A, 99(??):37– 45, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001832.

Feenberg:2006:SAR

[Fee06] Andrew Feenberg. Symmetry, asymmetry, and the real possibility of radical change: reply to Kochan. Studies in History and Philosophy of Science Part A, 37(4):721–727, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000823. See [Koc06].

Feest:2016:ERR

[Fee16] Uljana Feest. The experimenters’ regress reconsidered: Replication, tacit knowledge, and the dynamics of knowledge generation. Studies in History and Philosophy of Science Part A, 58(??):34– 45, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300036.

Feher:1982:GDI

[Feh82] Martha Fehér. Galileo and the demonstrative ideal of science. Studies in History and Philosophy of Science Part A, 13(2):87– 110, June 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368182900218.

Feher:2007:SSP

[Feh07] Márta Fehér. Saving the Strong Programme: a critique of Stephen Kemp’s recent paper. Studies in History and Philosophy of Science Part A, 38(1):235–240, March 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368106000999. See [Blo99, Kem05, Kem07].

Feintzeig:2014:COM

[Fei14] Benjamin Feintzeig. Can the ontological models framework accommodate Bohmian mechanics? Studies in History and Philosophy of Science Part A, 48 (part A)(??):59–67, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000823.

Feke:2012:MSD

[Fek12] Jacqueline Feke. Mathematizing the soul: the development of Ptolemy’s psychological theory from On the Kritêrion and Hêgemonikon to the Harmonics. Studies in History and Philosophy of Science Part A, 43(4):585–594, December 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000428.

Feke:2024:AGL

[Fek24] Jacqueline Feke. Ancient Greek laws of nature. Studies in History and Philosophy of Science Part A, 107(??):92–106, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812400058X.

Feola:2012:EAC

[Feo12] Vittoria Feola. Elias Ashmole’s collections and views about John Dee. Studies in History and Philosophy of Science Part A, 43 (3):530–538, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001221.

Ferraro:2001:ASG

[Fer01] Giovanni Ferraro. Analytical symbols and geometrical figures in eighteenth-century calculus. Studies in History and Philosophy of Science Part A, 32(3):535–555, September 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000097.

Ferrari:2012:BCK

[Fer12] Massimo Ferrari. Between Cassirer and Kuhn. some remarks on Friedman’s relativized a priori. Studies in History and Philosophy of Science Part A, 43(1):18–26, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000847.

Feyerabend:1970:DCP

[Fey70] Paul Feyerabend. In defence of classical physics. Studies in History and Philosophy of Science Part A, 1(1):59–85, May 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368170900257.

Feyerabend:1974:MG

[Fey74] Paul Feyerabend. Machamer on Galileo. Studies in History and Philosophy of Science Part A, 5(3):297–304, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936817490003X.

Feyerabend:1984:MTR

[Fey84] Paul K. Feyerabend. Mach’s theory of research and its relation to Einstein. Studies in History and Philosophy of Science Part A, 15(1):1–22, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936818490027X.

Feyerabend:2016:CIM

[Fey16a] Paul K. Feyerabend. The concept of intelligibility in modern physics (1948). Studies in History and Philosophy of Science Part A, 57(??):64–66, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001764.

Feyerabend:2016:BVM

[Fey16b] Paul K. Feyerabend. Der Begriff der Verständlichkeit in der modernen Physik (1948). Studies in History and Philosophy of Science Part A, 57(??):67–69, June 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001727.

Furlan:2022:EVA

[FG22] Stefano Furlan and Rocco Gaudenzi. The earth vibrates with analogies: the Dirac sea and the geology of the vacuum. Studies in History and Philosophy of Science Part A, 93(??):163– 174, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000590.

Franklin:1984:WDS

[FH84] Allan Franklin and Colin Howson. Why do scientists prefer to vary their experiments? Studies in History and Philosophy of Science Part A, 15(1):51–62, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900293.

Franklin:1985:NKB

[FH85] Allan Franklin and Colin Howson. Newton and Kepler, a Bayesian approach. Studies in History and Philosophy of Science Part A, 16 (4):379–385, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368185900184.

Franklin:1988:IPV

[FH88] Allan Franklin and Colin Howson. It probably is a valid experimental result: a Bayesian approach to the epistemology of experiment. Studies in History and Philosophy of Science Part A, 19 (4):419–427, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936818890009X.

Freyberg:2023:MPR

[FH23] Sascha Freyberg and Helmut Hauser. The morphological paradigm in robotics. Studies in History and Philosophy of Science Part A, 100(??):1–11, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123000742.

Field:1986:TMI

[Fie86] J. V. Field. Two mathematical inventions in Kepler’s ‘Ad vitellionem paralipomena’. Studies in History and Philosophy of Science Part A, 17(4):449–468, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818690004X.

See [Kep68, KR08].

Finocchiaro:1973:BRB

[Fin73] Maurice A. Finocchiaro. Book review: Criticism and the growth of knowledge: I. Lakatos and A. Musgrave. Cambridge University Press: Cambridge, 1970. viii + 282 pp. £1 (pbk.). Studies in History and Philosophy of Science Part A, 3(4):357–372, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900162. See [LM70].

Finocchiaro:1981:RTP

[Fin81] Maurice A. Finocchiaro. Remarks on truth, problem-solving, and methodology. Studies in History and Philosophy of Science Part A, 12(3):261–268, September 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368181900108.

Finocchiaro:1989:GAD

[Fin89] Maurice A. Finocchiaro, editor. The Galileo affair: a documentary history. California studies in the history of science. University of California Press, Berkeley, CA, USA, 1989. ISBN 0-520-06360-0, 0-520-06662-6 (paperback). xvi + 382 pp. LCCN QB36.G2 G319 1989.

Findlen:1990:ESR

[Fin90] Paula Findlen. Empty signs? reading the book of nature in renaissance science. Studies in History and Philosophy of Science Part A, 21(3):511–518, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090008V.

Finocchiaro:2002:GBT

[Fin02] Maurice A. Finocchiaro. Galileo as a ‘bad theologian’: a formative myth about Galileo’s trial. Studies in History and Philosophy of Science Part A, 33(4):753–791, December 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000304.

Finocchiaro:2011:GAI

[Fin11] Maurice A. Finocchiaro. Galilean argumentation and the inauthenticity of the Cigoli letter on painting vs. sculpture. Studies in History and Philosophy of Science Part A, 42(4):492–508, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811100063X.

Fisch:1985:NCT

[Fis85] Menachem Fisch. Necessary and contingent truth in William Whewell’s antithetical theory of knowledge. Studies in History and Philosophy of Science Part A, 16(4):275–314, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900147.

Fisher:2006:AME

[Fis06] Grant Fisher. The autonomy of models and explanation: anomalous molecular rearrangements in early twentieth-century physical organic chemistry. Studies in History and Philosophy of Science Part A, 37(4):562–584, December 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000768.

Fisher:2018:IRC

[Fis18] Amy A. Fisher. Inductive reasoning in the context of discovery: Analogy as an experimental stratagem in the history and philosophy of science. Studies in History and Philosophy of Science Part A, 69(??):23–33, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301711.

Fisher:2023:PPS

[Fis23] Grant Fisher. Practical pursuit in stem cell biology: Innovation, translation, and incomplete theorization. Studies in History and Philosophy of Science Part A, 97(??):1–12, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001595.

Faye:2021:WBW

[FJ21] Jan Faye and Rasmus Jaksland. What Bohr wanted Carnap to learn from quantum mechanics. Studies in History and Philosophy of Science Part A, 88(??):110–119, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000741.

Fischer:2024:BEE

[FJ24] Enno Fischer and Saana Jukola. Bodies of evidence: the ‘Excited Delirium Syndrome’ and the epistemology of cause-of-death inquiry. Studies in History and Philosophy of Science Part A, 104 (??):38–47, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000189.

Farwell:1990:EAN

[FK90] Ruth Farwell and Christopher Knee. The end of the absolute: a nineteenth-century contribution to General Relativity. Studies in History and Philosophy of Science Part A, 21(1):91– 121, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900162.

Fallis:2021:ADC

[FL21] Don Fallis and Peter J. Lewis. Animal deception and the content of signals. Studies in History and Philosophy of Science Part A, 87(??):114–124, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000352.

Fortin:2022:EIQ

[FL22] Sebastian Fortin and Olimpia Lombardi. Entanglement and indistinguishability in a quantum ontology of properties. Studies in History and Philosophy of Science Part A, 91(??):234– 243, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121002004.

Flamm:1983:LBH

[Fla83] D. Flamm. Ludwig Boltzmann and his influence on science. Studies in History and Philosophy of Science Part A, 14(4):255–278, December 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368183900080.

Fleisher:2022:PIR

[Fle22] Will Fleisher. Pursuit and inquisitive reasons. Studies in History and Philosophy of Science Part A, 94(??):17–30, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000681.

Fara:2004:INA

[FM04] Patricia Fara and David Money. Isaac Newton and Augustan Anglo–Latin poetry. Studies in History and Philosophy of Science Part A, 35(3):549–571, September 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810400055X.

Follinger:2012:ABW

[Föl12] Sabine Föllinger. Aristotle’s biological works as scientific literature. Studies in History and Philosophy of Science Part A, 43(2): 237–244, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001361.

Forrester:1975:CCE

[For75] John Forrester. Chemistry and the conservation of energy: The work of James Prescott Joule. Studies in History and Philosophy of Science Part A, 6(4):273–313, November 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900254.

Forster:1988:UEC

[For88] Malcolm R. Forster. Unification, explanation, and the composition of causes in Newtonian mechanics. Studies in History and Philosophy of Science Part A, 19(1):55–101, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900209.

Forgan:1989:ASI

[For89] Sophie Forgan. The architecture of science and the idea of a university. Studies in History and Philosophy of Science Part A, 20 (4):405–434, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900174.

Forgione:2022:FST

[For22] Marco Forgione. Feynman’s space–time view in quantum electrodynamics. Studies in History and Philosophy of Science Part A, 93(??):136–148, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000577.

Force:1999:NRC

[FP99] James E. Force and Richard H. (Richard Henry) Popkin, editors. Newton and religion: context, nature, and influence, volume 161 of International archives of the history of ideas = Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5744-2. xvii + 325 pp. LCCN

BX4827.N48 N48 1999. URL http://www.loc.gov/catdir/
enhancements/fy0821/99015083-d.html;   http://www.loc.
gov/catdir/enhancements/fy0821/99015083-t.html.

Franklin:2023:BRM

[FR23] Alexander Franklin and Katie Robertson. Book review: A Middle Way: a Non-Fundamental Approach to Many-Body Physics by Robert Batterman: Autonomy and Varieties of Reduction. Studies in History and Philosophy of Science Part A, 97(??):123– 125, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001285. See [Bat23].

Fankhauser:2024:GRR

[FR24] Johannes Fankhauser and James Read. Gravitational redshift revisited: Inertia, geometry, and charge. Studies in History and Philosophy of Science Part A, 108(??):19–27, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812400133X.

Frankel:1979:CCD

[Fra79a] Henry Frankel. The career of continental drift theory: An application of Imre Lakatos’ analysis of scientific growth to the rise of drift theory. Studies in History and Philosophy of Science Part A, 10(1):21–66, March 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368179900037.

Franklin:1979:DNP

[Fra79b] Allan Franklin. The discovery and nondiscovery of parity nonconservation. Studies in History and Philosophy of Science Part A, 10 (3):201–257, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368179900086.

Frankel:1981:PDP

[Fra81] Henry Frankel. The paleobiogeographical debate over the problem of disjunctively distributed life forms. Studies in History and Philosophy of Science Part A, 12(3):211–259, September 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900091.

Frankel:1984:BBA

[Fra84] Henry Frankel. Biogeography, before and after the rise of sea floor spreading. Studies in History and Philosophy of Science Part A, 15(2):141–168, June 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368184900049.

Fraser:1990:LAM

[Fra90] Craig G. Fraser. Lagrange’s analytical mathematics, its Cartesian origins and reception in Comte’s positive philosophy. Studies in History and Philosophy of Science Part A, 21(2):243– 256, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900243.

Franssen:1993:DKA

[Fra93] Maarten Franssen. Did King Alfonso of Castile really want to advise God against the Ptolemaic system? The legend in history. Studies in History and Philosophy of Science Part A, 24(3):313– 325, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390031E.

Franklin:1994:HAE

[Fra94a] Allan Franklin. How to avoid the experimenters’ regress. Studies in History and Philosophy of Science Part A, 25(3):463– 491, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900620.

Franklin:1994:FSD

[Fra94b] J. Franklin. The formal sciences discover the philosophers’ stone. Studies in History and Philosophy of Science Part A, 25(4):513– 533, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900450.

Fraser:2005:TLN

[Fra05] Doreen L. Fraser. The third law in Newton’s waste book (or, the road less taken to the second law). Studies in History and Philosophy of Science Part A, 36(1):43–60, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104001190.

Franssen:2006:NA

[Fra06] Maarten Franssen. The normativity of artefacts. Studies in History and Philosophy of Science Part A, 37(1):42–57, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105001081.

Franklin:2008:FOC

[Fra08] Allan Franklin. Is failure an option? contingency and refutation. Studies in History and Philosophy of Science Part A, 39(2): 242–252, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000265.

Franco:2020:HRC

[Fra20] Paul L. Franco. Hans Reichenbach’s and C. I. Lewis’s Kantian philosophies of science. Studies in History and Philosophy of Science Part A, 80(??):62–71, April 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811830267X.

Franco:2021:OLP

[Fra21a] Paul L. Franco. Ordinary language philosophy, explanation, and the historical turn in philosophy of science. Studies in History and Philosophy of Science Part A, 90(??):77–85, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001436.

Frank:2021:WEE

[Fra21b] David M. Frank. What is the environment in environmental health research? Perspectives from the ethics of science. Studies in History and Philosophy of Science Part A, 88(??):172– 180, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000923.

Fraser:2021:TOR

[Fra21c] James D. Fraser. The twin origins of renormalization group concepts. Studies in History and Philosophy of Science Part A, 89 (??):114–128, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001126.

Frege:1971:LI

[Fre71] G. Frege. On the law of inertia. Studies in History and Philosophy of Science Part A, 2(3):195–212, November 1971. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368171900410.

Freundlich:1978:DC

[Fre78] Yehudah Freundlich. In defence of Copenhagenism. Studies in History and Philosophy of Science Part A, 9(3):151–179, September 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900043.

Freudenthal:1979:HSD

[Fre79] Gad Freudenthal. How strong is Dr. Bloor’s ‘strong programme’? Studies in History and Philosophy of Science Part A, 10(1):67– 83, March 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900049.

Freundlich:1980:MST

[Fre80a] Yehudah Freundlich. Methodologies of science as tools for historical research. Studies in History and Philosophy of Science Part A, 11 (4):257–266, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368180900060.

Freundlich:1980:TEB

[Fre80b] Yehudah Freundlich. Theory evaluation and the bootstrap hypothesis. Studies in History and Philosophy of Science Part A, 11 (4):267–277, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368180900072.

Freudenthal:2002:PML

[Fre02] Gideon Freudenthal. Perpetuum mobile: the Leibniz–Papin controversy. Studies in History and Philosophy of Science Part A, 33 (3):573–637, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000401.

Freedman:2005:NEW

[Fre05] Karyn L. Freedman. Naturalized epistemology, or what the Strong Programme can’t explain. Studies in History and Philosophy of Science Part A, 36(1):135–148, March 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104001232.

French:2015:EFE

[Fre15] Christopher F. French. Explicating formal epistemology: Carnap’s legacy as Jeffrey’s radical probabilism. Studies in History and Philosophy of Science Part A, 53(??):33–42, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000618.

French:2019:DES

[Fre19] Steven French. Defending eliminative structuralism and a whole lot more (or less). Studies in History and Philosophy of Science Part A, 74(??):22–29, ???? 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118303340.

Freestone:2021:CDW

[Fre21] Jamie Milton Freestone. Contemporary Darwinism as a worldview. Studies in History and Philosophy of Science Part A, 90 (??):68–76, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001242.

French:2025:FVL

[Fre25] Steven French. A French view of London. Studies in History and Philosophy of Science Part A, 110(??):30–39, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368125000111.

Friedman:1998:SSK

[Fri98] Michael Friedman. On the sociology of scientific knowledge and its philosophical agenda. Studies in History and Philosophy of Science Part A, 29(2):239–271, June 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000216.

Fricker:2002:TSP

[Fri02] Elizabeth Fricker. Trusting others in the sciences: a priori or empirical warrant? Studies in History and Philosophy of Science Part A, 33(2):373–383, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368102000067.

Friedman:2003:TPM

[Fri03a] Michael Friedman. Transcendental philosophy and mathematical physics. Studies in History and Philosophy of Science Part A, 34 (1):29–43, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000857.

Frigg:2003:SOC

[Fri03b] Roman Frigg. Self-organised criticality-what it is and what it isn’t. Studies in History and Philosophy of Science Part A, 34(3):613– 632, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000463.

Frierson:2008:EPC

[Fri08] Patrick Frierson. Empirical psychology, common sense, and Kant’s empirical markers for moral responsibility. Studies in History and Philosophy of Science Part A, 39(4):473–482, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000812.

Friedman:2012:RDR

[Fri12] Michael Friedman. Reconsidering the dynamics of reason: Response to Ferrari, Mormann, Nordmann, and Uebel. Studies in History and Philosophy of Science Part A, 43(1):47–53, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000872.

Friebe:2014:IDN

[Fri14] Cord Friebe. Individuality, distinguishability, and (non-)entanglement: a defense of Leibniz’s principle. Studies in History and Philosophy of Science Part A, 48 (part A)(??):89–98, November 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000860.

Frierson:2018:TRP

[Fri18] Patrick R. Frierson. Towards a research program in Kantian positive psychology. Studies in History and Philosophy of Science Part A, 71(??):89–98, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116301315.

Friend:2021:ITD

[Fri21] Toby Friend. Intervening on time derivatives. Studies in History and Philosophy of Science Part A, 89(??):74–83, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001060.

Friedman:2024:TTM

[Fri24a] Michael Friedman. A tale of a threshing machine: Images of the Voigt–Leibniz mathematical-agricultural machine at the beginning of the 18th century. Studies in History and Philosophy of Science Part A, 105(??):17–31, June 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000220.

Friend:2024:SCF

[Fri24b] Toby Friend. Soft control: Furthering the case for Modified Interventionist Theory. Studies in History and Philosophy of Science Part A, 108(??):93–100, December 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001365.

F-S:2008:PLO

[FSJ08] M. F.-S. and N. J. Peter Lipton (9^{th} October 1954–25^{th} November 2007). Studies in History and Philosophy of Science Part A, 39 (1):1, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000022.

Fu:1995:HTH

[Fu95] Daiwie Fu. Higher taxonomy and higher incommensurability. Studies in History and Philosophy of Science Part A, 26(2):273– 294, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000025.

Fuller:1989:PSD

[Ful89] Steve Fuller. Philosophy of science and its discontents. Westview Press, Boulder, CO, USA, 1989. ISBN 0-8133-0611-6. x + 188 pp. LCCN Q175 .F925 1989.

Fullinwider:1990:HHP

[Ful90] S. P. Fullinwider. Hermann von Helmholtz: The problem of Kantian influence. Studies in History and Philosophy of Science Part A, 21(1):41–55, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090014Y.

Fullinwider:1993:BRB

[Ful93] S. P. Fullinwider. Book review: The natural and the normative: Theories of spatial perception from Kant to Helmholtz: Gary Hatfield, (Cambridge, Mass.: MIT Press, 1990), xii + 366 pp. ISBN 0-262-08086-9 Cloth $35.00. Studies in History and Philosophy of Science Part A, 24(3):485–491, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390039M. See [Hat90].

Fuller:1994:BRB

[Ful94] Steve Fuller. Book review: The advancement of science: Science without legend, objectivity without illusions: Philip Kitcher (Oxford and New York: Oxford University Press, 1993), viii + 421 pp. ISBN 0-19-504628-5. Studies in History and Philosophy of Science Part A, 25(2):251–261, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900302. See [Kit93].

Fuller:2018:PRS

[Ful18] Gary Fuller. Physicalism, realization, and structure. Studies in History and Philosophy of Science Part A, 68(??):31– 36, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302728.

Fuller:2024:DSM

[Ful24] Jonathan Fuller. Demarcating scientific medicine. Studies in History and Philosophy of Science Part A, 106(??):177–185, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000992.

Fumagalli:2020:HTR

[Fum20] Roberto Fumagalli. How thin rational choice theory explains choices. Studies in History and Philosophy of Science Part A, 83(??):63–74, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301943.

Funkenstein:1975:DET

[Fun75] Amos Funkenstein. Descartes, eternal truths, and the divine omnipotence. Studies in History and Philosophy of Science Part A, 6(3):185–199, October 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368175900229.

Funkenstein:1986:TSI

[Fun86] Amos Funkenstein. Theology and the scientific imagination from the Middle Ages to the Seventeenth Century. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08408-4. xii + 421 pp. LCCN BT130 .F86 1986. US$39.50.

Furlan:2022:PBD

[Fur22] Stefano Furlan. Pursuitworthiness between daring conservatism and procrastination: Wheeler and the path towards black holes. Studies in History and Philosophy of Science Part A, 96(??):174– 185, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001571.

Futch:2002:SNM

[Fut02] Michael J. Futch. Supervenience and (non-modal) reductionism in Leibniz’s philosophy of time. Studies in History and Philosophy of Science Part A, 33(4):793–810, December 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000316.

Frercks:2009:RDN

[FWW09] Jan Frercks, Heiko Weber, and Gerhard Wiesenfeldt. Reception and discovery: the nature of Johann Wilhelm Ritter’s invisible rays. Studies in History and Philosophy of Science Part A, 40(2): 143–156, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810900020X.

Fyfe:2002:PCP

[Fyf02] Aileen Fyfe. Publishing and the classics: Paley’s Natural Theology and the nineteenth-century scientific canon. Studies in History and Philosophy of Science Part A, 33(4):729–751, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000328. See [Pal02, PEK06].

Gabbey:1971:FIS

[Gab71] Alan Gabbey. Force and inertia in seventeenth-century dynamics. Studies in History and Philosophy of Science Part A, 2(1): 1–67, May 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900379.

Gabbey:1973:BRB

[Gab73] Alan Gabbey. Book review: The conflict between atomism and conservation theory 1644–1860: Wilson L. Scott. Macdonald: London and Elsevier: New York, 1970. xiv + 312 pp., 3 plates, 6 figs., index. £5.00 ($16.00). Studies in History and Philosophy of Science Part A, 3(4):373–385, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368173900174. See [Sco70].

Gale:1971:BRB

[Gal71] Richard M. Gale. Book review: Kant’s theory of time: Sadik J. Al-Azm, Philosophical Library, Inc.: New York, 1967. iv + 84 pp. $3.95. Studies in History and Philosophy of Science Part A, 2 (1):95–96, May 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900409. See [AA67].

Gal:1996:PKW

[Gal96] Ofer Gal. Producing knowledge in the workshop: Hooke’s ‘inflection’ from optics to planetary motion. Studies in History and Philosophy of Science Part A, 27(2):181–205, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000356.

Galbraith:2021:VES

[Gal21] Jude Galbraith. Values in early-stage climate engineering: the ethical implications of “doing the research”. Studies in History and Philosophy of Science Part A, 86(??):103–113, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000182.

Gambarotto:2021:CKP

[Gam21a] Andrea Gambarotto. Corrigendum to “The ‘Kantian principle’ for natural history and its historical significance studies in history and philosophy of science part C: Studies in history and philosophy of biological and biomedical science” [64 (2017) 22– 27]. Studies in History and Philosophy of Science Part A, 85(??): 219, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120302107.

Gambarotto:2021:CVF

[Gam21b] Andrea Gambarotto. Corrigendum to “Vital forces and organization: Philosophy of nature and biology in Karl Friedrich Kielmeyer”

[Studies in History and Philosophy of Science Part C: Studies in History and Philosophy of Biological and Biomedical Science 48 (2014) 12–20]. Studies in History and Philosophy of Science Part A, 85(??):218, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120302028.

Gao:2021:EMS

[Gao21] Shan Gao. Existence of macroscopic spatial superpositions in collapse theories. Studies in History and Philosophy of Science Part A, 86(??):1–5, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120302089.

Garber:1976:SRP

[Gar76] Elizabeth Garber. Some reactions to Planck’s law, 1900–1914. Studies in History and Philosophy of Science Part A, 7(2):89– 126, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900133.

Garber:1989:OST

[Gar89] Daniel Garber. Old school ties. Studies in History and Philosophy of Science Part A, 20(4):531–539, December 1989. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368189900228.

Garber:1996:PS

[Gar96] Daniel Garber. Philosophers of substance. Studies in History and Philosophy of Science Part A, 27(3):421–427, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000399.

Gare:2000:ABH

[Gar00] Arran Gare. Aleksandr Bogdanov’s history, sociology and philosophy of science. Studies in History and Philosophy of Science Part A, 31(2):231–248, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000029.

Gattei:2016:FTR

[Gat16] Stefano Gattei. Feyerabend, truth, and relativisms: Footnotes to the Italian debate. Studies in History and Philosophy of Science Part A, 57(??):87–95, June 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001740.

Gaukroger:1976:BPE

[Gau76] Stephen W. Gaukroger. Bachelard and the problem of epistemological analysis. Studies in History and Philosophy of Science Part A, 7(3):189–244, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368176900066.

Gaukroger:1981:AFS

[Gau81] Stephen Gaukroger. Aristotle on the function of sense perception. Studies in History and Philosophy of Science Part A, 12(1): 75–89, March 1981. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368181900066.

Gaukroger:1998:JTD

[Gau98] Stephen Gaukroger. Justification, truth, and the development of science. Studies in History and Philosophy of Science Part A, 29(1): 97–112, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197000204.

Gaukroger:2016:KNM

[Gau16] Stephen Gaukroger. Kant and the nature of matter: Mechanics, chemistry, and the life sciences. Studies in History and Philosophy of Science Part A, 58(??):108–114, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116000261.

Gay:1976:RTC

[Gay76] Hannah Gay. Radicals and types: a critical comparison of the methodologies of Popper and Lakatos and their use in the reconstruction of some 19th century chemistry. Studies in History and Philosophy of Science Part A, 7(1):1–51, ????

1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368176900029.

Gay:1977:NGC

[Gay77] Hannah Gay. Noble gas compounds: a case study in scientific conservatism and opportunism. Studies in History and Philosophy of Science Part A, 8(1):61–70, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817790019X.

Gay:1978:ACA

[Gay78] Hannah Gay. The asymmetric carbon atom: (a) a case study of independent discovery; (b) an inductivist model for scientific method. Studies in History and Philosophy of Science Part A, 9 (3):207–238, September 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368178900067.

Gee:2014:CAA

[Gee14] Teri Gee. Cultural alterations of Aratus’s Phaenomena. Studies in History and Philosophy of Science Part A, 48(??):42–45, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811400079X.

Geiszler:2023:IAR

[Gei23] Lukas Geiszler. Imitation in automata and robots: a philosophical case study on Kempelen. Studies in History and Philosophy of Science Part A, 100(??):22–31, August 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000845.

Gellatly:1980:LNS

[Gel80] Angus Gellatly. Logical necessity and the strong programme for the sociology of knowledge. Studies in History and Philosophy of Science Part A, 11(4):325–339, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900102.

Gelfert:2010:RRI

[Gel10] Axel Gelfert. Reconsidering the role of inference to the best explanation in the epistemology of testimony. Studies in History and Philosophy of Science Part A, 41(4):386–396, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000567.

Gelfert:2011:MFS

[Gel11a] Axel Gelfert. Mathematical formalisms in scientific practice: From denotation to model-based representation. Studies in History and Philosophy of Science Part A, 42(2):272–286, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001147.

Gelfert:2011:MBR

[Gel11b] Axel Gelfert. Model-based representation in scientific practice: New perspectives. Studies in History and Philosophy of Science Part A, 42(2):251–252, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001111.

Genta:2020:HTA

[Gen20] Benjamin S. Genta. How to think about analogical inferences: a reply to Norton. Studies in History and Philosophy of Science Part A, 82(??):17–24, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301359.

Georgescu:2014:DDW

[Geo14] Laura Georgescu. The diagrammatic dimension of William Gilbert’s De magnete. Studies in History and Philosophy of Science Part A, 47(??):18–25, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000478.

Gervais:2014:FIL

[Ger14] Raoul Gervais. A framework for inter-level explanations: Outlines for a new explanatory pluralism. Studies in History and Philosophy of Science Part A, 48(??):1–9, December 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000636.

Gerken:2020:PST

[Ger20] Mikkel Gerken. Public scientific testimony in the scientific image. Studies in History and Philosophy of Science Part A, 80(??): 90–101, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119300640.

Gebharter:2023:UEC

[GFE23] Alexander Gebharter and Christian J. Feldbacher-Escamilla. Unification and explanation from a causal perspective. Studies in History and Philosophy of Science Part A, 99(??):28–36, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001820.

Gavroglu:1988:HKO

[GG88] Kostas Gavroğlu and Yorgos Goudaroulis. Heike Kamerlingh Onnes’ researches at Leiden and their methodological implications. Studies in History and Philosophy of Science Part A, 19(2): 243–274, June 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900283.

Ghirardi:1994:OPP

[GG94] GianCarlo Ghirardi and Renata Grassi. Outcome predictions and property attribution: the EPR argument reconsidered. Studies in History and Philosophy of Science Part A, 25(3):397– 423, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900590.

Godin:2002:ERS

[GG02] Benoît Godin and Yves Gingras. The experimenters’ regress: from skepticism to argumentation. Studies in History and Philosophy of Science Part A, 33(1):133–148, March 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000322.

Gomes:2021:AMR

[GG21] Henrique Gomes and Sean Gryb. Angular momentum without rotation: Turbocharging relationalism. Studies in History and Philosophy of Science Part A, 88(??):138–155, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000704.

Ghiselin:1969:TDMa

[Ghi69a] Michael T. Ghiselin. The triumph of the Darwinian method. University of Chicago Press, Chicago, IL, USA and London, UK, 1969. ISBN 0-226-29024-7. xxi + 287 pp. LCCN ????

Ghiselin:1969:TDMb

[Ghi69b] Michael T. Ghiselin. The triumph of the Darwinian method. University of California Press, Berkeley, CA, USA, 1969. vi + 287 pp. LCCN ????

Ghiselin:1972:TDM

[Ghi72] Michael T. Ghiselin. The triumph of the Darwinian method. University of California Press, Berkeley, CA, USA, 1972. ISBN 0-520-02247-5. LCCN ????

Ghiselin:1984:TDM

[Ghi84] Michael T. Ghiselin. The triumph of the Darwinian method. University of Chicago Press, Chicago, IL, USA and London, UK, 1984. ISBN 0-226-29024-7. xxi + 287 pp. LCCN QH31.

Ghiselin:1994:ELE

[Ghi94] Michael T. Ghiselin. Evolving the language of evolution. Studies in History and Philosophy of Science Part A, 25(2):263– 269, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900310.

Ghiselin:2003:TDM

[Ghi03] Michael T. Ghiselin. The triumph of the Darwinian method. Dover Publications, Inc., New York, NY, USA, 2003. ISBN 0-486-43274-2 (paperback). xii + 287 pp. LCCN QH31.D2

G5 2003. URL http://catdir.loc.gov/catdir/enhancements/
fy0614/2003048487-d.html.

Giagioli:1994:BRB

[Gia94] Mario Giagioli. Book review: ‘Legem impone subactis’: Studi su filosofia e scienza dei gesuiti in Italia, 1540–1632: Ugo Baldini, (Rome: Bulzoni, 1992). Studies in History and Philosophy of Science Part A, 25(4):637–646, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900523.

See [Bal92].

Gillette:2021:DES

[GID21] Kinley Gillette, S. Andrew Inkpen, and C. Tyler DesRoches. Does environmental science crowd out non-epistemic values? Studies in History and Philosophy of Science Part A, 87(??):81– 92, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000170.

Giedymin:1977:OSP

[Gie77] Jerzy Giedymin. On the origin and significance of Poincaré’s conventionalism. Studies in History and Philosophy of Science Part A, 8(4):271–301, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936817790022X.

Giere:1989:SRI

[Gie89] Ronald N. Giere. Scientific rationality as instrumental rationality. Studies in History and Philosophy of Science Part A, 20(3): 377–384, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900137.

Giedymin:1991:GPC

[Gie91a] Jerzy Giedymin. Geometrical and physical conventionalism of Henri Poincaré in epistemological formulation. Studies in History and Philosophy of Science Part A, 22(1):1–22, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190012H.

Giere:1991:BRB

[Gie91b] Ronald N. Giere. Book review: Philosophy of science and its discontents: Steve Fuller, (Boulder: Westview Press, 1989), x + 188 pp., ISBN 0-8133-0611-6 Cloth. Studies in History and Philosophy of Science Part A, 22(3):515–523, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190007F. See [Ful89].

Giedymin:1992:CPC

[Gie92] Jerzy Giedymin. Conventionalism, the pluralist conception of theories and the nature of interpretation. Studies in History and Philosophy of Science Part A, 23(3):423–443, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290003O.

Giere:2006:SP

[Gie06] Ronald N. Giere. Scientific perspectivism. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-29212-6 (hardcover). 151 pp. LCCN Q175 .G48899 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0666/2006002489-b.
html;   http://www.loc.gov/catdir/enhancements/fy0666/
2006002489-d.html;   http://www.loc.gov/catdir/toc/ecip067/
2006002489.html.

Giere:2009:SPB

[Gie09] Ronald N. Giere. Scientific perspectivism: behind the stage door. Studies in History and Philosophy of Science Part A, 40(2):221– 223, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000168. See [Bro09].

Giere:2016:FP

[Gie16] Ronald N. Giere. Feyerabend’s perspectivism. Studies in History and Philosophy of Science Part A, 57(??):137–141, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001685.

Gijsbers:2013:UEU

[Gij13] Victor Gijsbers. Understanding, explanation, and unification. Studies in History and Philosophy of Science Part A, 44(3):516–522, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001136.

Gil:1983:ITG

[Gil83] David Gil. Intuitionism, transformational generative grammar and mental acts. Studies in History and Philosophy of Science Part A, 14(3):231–254, September 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368183900146.

Ginev:1988:SPH

[Gin88] Dimiter Ginev. Scientific progress and the hermeneutic circle. Studies in History and Philosophy of Science Part A, 19(3):391– 395, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900064.

Gingras:2007:EYD

[Gin07] Yves Gingras. Everything you did not necessarily want to know about gravitational waves. and why. Studies in History and Philosophy of Science Part A, 38(1):268–282, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810600104X. See [Col04] and comment

[Col08].

Gingras:2009:RCA

[Gin09] Yves Gingras. Response to Collins about ‘one point’ that is absent from my review of his book. Studies in History and Philosophy of Science Part A, 40(1):112, March 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108001179.

Giovanelli:2014:OML

[Gio14] Marco Giovanelli. ‘But one must not legalize the mentioned sin’: Phenomenological vs. dynamical treatments of rods and clocks in Einstein’s thought. Studies in History and Philosophy of Science Part A, 48 (part A)(??):20–44, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000963.

Giovanelli:2016:HCP

[Gio16] Marco Giovanelli. Hermann Cohen’s Das Princip der Infinitesimal-Methode: the history of an unsuccessful book. Studies in History and Philosophy of Science Part A, 58(??):9–23, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116000091.

Giovanelli:2022:MKC

[Gio22] Marco Giovanelli. Motivational Kantianism: Cassirer’s late shift towards a regulative conception of the a priori. Studies in History and Philosophy of Science Part A, 95(??):118–125, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001169.

Giovanelli:2024:PCA

[Gio24a] Marco Giovanelli. The philosophical coming of age of science. Euler’s role in Cassirer’s early philosophy of space and time. Studies in History and Philosophy of Science Part A, 108(??):55–63, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001390.

Giovanelli:2024:VSK

[Gio24b] Marco Giovanelli. Variability and substantiality. Kurd Lasswitz, the Marburg School and the neo-Kantian historiography of science. Studies in History and Philosophy of Science Part A, 106 (??):155–164, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124001043.

Gregersen:1988:AER

[GK88] Frans Gregersen and Simo Køppe. Against epistemological relativism. Studies in History and Philosophy of Science Part A, 19 (4):447–487, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368188900118.

Greif:2024:SGF

[GKS24] Hajo Greif, Adam P. Kubiak, and Paweł Stacewicz. Selection, growth and form. Turing’s two biological paths towards intelligent machinery. Studies in History and Philosophy of Science Part A, 106(??):126–135, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000657.

Gentile:2024:CBR

[GL24] Nélida Gentile and Susana Lucero. On compatibility between realism and fictionalism: a response to Suárez’ proposal. Studies in History and Philosophy of Science Part A, 103(??):168– 175, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001711.

Glas:1978:MEP

[Gla78] E. Glas. Methodology and the emergence of physiological chemistry. Studies in History and Philosophy of Science Part A, 9 (4):291–312, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368178900109.

Glas:1983:BSB

[Gla83] Eduard Glas. Bio-science between experiment and ideology, 1835– 1850. Studies in History and Philosophy of Science Part A, 14(1): 39–57, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368183900043.

Glas:1986:DMC

[Gla86] Eduard Glas. On the dynamics of mathematical change in the case of Monge and the French Revolution. Studies in History and Philosophy of Science Part A, 17(3):249–268, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900099.

Glas:1989:TPMa

[Gla89a] Eduard Glas. Testing the philosophy of mathematics in the history of mathematics: Part I: The sociocognitive process of conceptual change. Studies in History and Philosophy of Science Part A, 20(1):115–131, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900368.

Glas:1989:TPMb

[Gla89b] Eduard Glas. Testing the philosophy of mathematics in the history of mathematics: Part II: The similarity between mathematical and scientific growth of knowledge. Studies in History and Philosophy of Science Part A, 20(2):157–174, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900010.

Glas:1993:FFR

[Gla93] Eduard Glas. From form to function: a reassessment of Felix Klein’s unified programme of mathematical research, education and development. Studies in History and Philosophy of Science Part A, 24(4):611–631, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390055O.

Glasner:1997:GSC

[Gla97] Ruth Glasner. Gersonides on simple and composite movements. Studies in History and Philosophy of Science Part A, 28(4):545– 584, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000295.

Glas:1999:TEM

[Gla99] Eduard Glas. Thought-experimentation and mathematical innovation. Studies in History and Philosophy of Science Part A, 30(1): 1–19, March 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000417.

Glas:2001:PPMa

[Gla01a] Eduard Glas. The ‘Popperian Programme’ and mathematics: Part I: the fallibilist logic of mathematical discovery. Studies in History and Philosophy of Science Part A, 32(1):119– 137, March 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000376.

Glas:2001:PPMb

[Gla01b] Eduard Glas. The ‘Popperian Programme’ and mathematics: Part II: From quasi-empiricism to mathematical research programmes. Studies in History and Philosophy of Science Part A, 32(2):355– 376, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000388.

Glas:2002:SCM

[Gla02] Eduard Glas. Socially conditioned mathematical change: the case of the French Revolution. Studies in History and Philosophy of Science Part A, 33(4):709–728, December 2002. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000298.

Glasner:2020:ESE

[Gla20] Ruth Glasner. An early stage in the evolution of Aristotle’s physics. Studies in History and Philosophy of Science Part A, 81 (??):24–31, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301006.

Glick:2016:OQF

[Gli16] David Glick. The ontology of quantum field theory: Structural realism vindicated? Studies in History and Philosophy of Science Part A, 59(??):78–86, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300243.

Graboyes:2022:MMM

[GMC22] Melissa Graboyes, Judith Meta, and Rhaine Clarke. Mazingira and the malady of malaria: Perceptions of malaria as an environmental disease in contemporary Zanzibar. Studies in History and Philosophy of Science Part A, 95(??):134–144, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001108.

Gaukroger:2016:IKE

[GN16] Stephen Gaukroger and Dalia Nassar. Introduction: Kant and the empirical sciences. Studies in History and Philosophy of Science Part A, 58(??):55–56, August 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116000285.

Gambarotto:2022:TOK

[GN22] Andrea Gambarotto and Auguste Nahas. Teleology and the organism: Kant’s controversial legacy for contemporary biology. Studies in History and Philosophy of Science Part A, 93(??):

47–56, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812200036X.

Gockel:2008:PKP

[Goc08] Bettina Gockel. Paul Klee’s picture-making and persona: tools for making invisible realities visible. Studies in History and Philosophy of Science Part A, 39(3):418–433, September 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000654.

Goddiksen:2014:CIC

[God14] Mads Goddiksen. Clarifying interactional and contributory expertise. Studies in History and Philosophy of Science Part A, 47 (??):111–117, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000557. See reply [RGD15].

Goe:1972:ATL

[Goe72] George Goe. Archimedes’ theory of the lever and Mach’s critique. Studies in History and Philosophy of Science Part A, 2(4):329– 345, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900027.

Gokalp:1990:ISF

[Gök90] Iskender Gökalp. The interrelating of scientific fields: The case of turbulence and combustion. Studies in History and Philosophy of Science Part A, 21(3):413–429, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819090003Q.

Goldman:1975:AKO

[Gol75] Steven Louis Goldman. Alexander Kojève on the origin of modern science: Sociological modelling gone awry. Studies in History and Philosophy of Science Part A, 6(2):113–124, June 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900175.

Golinski:1997:RBC

[Gol97] Jan Golinski. Robert Boyle’s coat of many colours. Studies in History and Philosophy of Science Part A, 28(1):209– 217, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936819788212X.

Goldin:2004:ACD

[Gol04] Owen Goldin. Atoms, complexes, and demonstration: Posterior analytics 96b15-25. Studies in History and Philosophy of Science Part A, 35(4):707–727, December 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000834.

Goldenberg:2022:BRV

[Gol22] Maya J. Goldenberg. Book review: Vaccine Hesitancy: Public Trust, Expertise, and the War on Science by Maya Goldenberg: Reply by the author. Studies in History and Philosophy of Science Part A, 96(??):121–124, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368122001546. See [Joh22, Chu22, WV22, Lea22].

Goldberg:2023:BRC

[Gol23] Benjamin Goldberg. Book review: The Chemical Philosophy of Robert Boyle: Mechanicism, Chymical Atoms, and Emergence by Marina Paola Banchetti-Robino: a priori, a posteriori, and the Historiography of Early Modern Science. Studies in History and Philosophy of Science Part A, 97(??):137–140, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001625.

Gonzalez:1996:TNF

[Gon96] Wenceslao J. González. Towards a new framework for revolutions in science. Studies in History and Philosophy of Science Part A, 27 (4):607–625, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000372.

Gonzalez:2025:BOK

[Gon25] Juan Carlos González. Believing in organisms: Kant’s nonmechanistic philosophy of nature. Studies in History and Philosophy of Science Part A, 109(??):109–119, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001651.

Gooding:1978:CEB

[Goo78] David Gooding. Conceptual and experimental bases of Faraday’s denial of electrostatic action at a distance. Studies in History and Philosophy of Science Part A, 9(2):117–149, June 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900031.

Goosens:1980:GRT

[Goo80] William K. Goosens. Galileo’s response to the tower argument. Studies in History and Philosophy of Science Part A, 11(3):215– 227, September 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900266.

Gooding:1986:HDS

[Goo86] David Gooding. How do scientists reach agreement about novel observations? Studies in History and Philosophy of Science Part A, 17(2):205–230, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368186900269.

Good:1987:JHO

[Goo87] Gregory Good. John Herschel’s optical researches and the development of his ideas on method and causality. Studies in History and Philosophy of Science Part A, 18(1):1–41, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900094.

Gooday:2004:CGH

[Goo04] Graeme Gooday. Cry ‘Good for history, Cambridge and Saint George’? Studies in History and Philosophy of Science Part A, 35 (4):861–872, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000937.

Gooday:2013:CPM

[Goo13] Graeme Gooday. Combative patenting: Military entrepreneurship in First World War telecommunications. Studies in History and Philosophy of Science Part A, 44(2):247–258, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112001057.

Goodwin:2021:GTF

[Goo21] William Goodwin. Gaining traction: Foothold concepts and exemplars in conceptual change. Studies in History and Philosophy of Science Part A, 90(??):145–152, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001448.

Gorham:1991:PPJ

[Gor91] Geoffrey Gorham. Planck’s principle and Jeans’s conversion. Studies in History and Philosophy of Science Part A, 22(3):471–497, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190005D.

Gouvea:2023:HHP

[Gou23] Devin Y. Gouvêa. Historicizing the homology problem. Studies in History and Philosophy of Science Part A, 99(??):56– 66, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000055.

Gower:1973:SPH

[Gow73] Barry Gower. Speculation in physics: The history and practice of naturphilosophie. Studies in History and Philosophy of Science Part A, 3(4):301–356, February 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900150.

Gower:1987:PPD

[Gow87] Barry Gower. Planets and probability: Daniel Bernoulli on the inclinations of the planetary orbits. Studies in History and Philosophy of Science Part A, 18(4):441–454, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900021.

Gower:1997:HPB

[Gow97] B. S. Gower. Henri Poincaré and Bruno de Finetti: Conventions and scientific reasoning. Studies in History and Philosophy of Science Part A, 28(4):657–679, December 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000137.

Gibbs:2010:LUG

[GP10] Raymond W. Gibbs, Jr. and Marcus Perlman. Language understanding is grounded in experiential simulations: a response to Weiskopf. Studies in History and Philosophy of Science Part A, 41 (3):305–308, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000361.

Gonzalez-Rivadeneira:2023:BAL

[GR23] Tania I. González-Rivadeneira. The ‘biocultural approach’ in Latin American ethnobiology. Studies in History and Philosophy of Science Part A, 101(??):24–29, October 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001024.

Graves:1971:CFC

[Gra71] John Cowperthwaite Graves. The conceptual foundations of contemporary relativity theory. MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-07040-5. ix + 361 pp. LCCN QC6 .G687. Foreword by John Archibald Wheeler.

Grant:1972:BRB

[Gra72] Edward Grant. Book review: Nicole Oresme and the medieval geometry of qualities and motions. A treatise on the uniformity and difformity of intensities known as ‘tractatus de configurationibus qualitatum et motuum’: Marshall Clagett (ed. and tr.), edited with an introduction, English translation and commentary by Marshall Clagett. University of Wisconsin Press: Madison, Milwaukee, 1968; and London, 1969. xiii + 713 pp. £7.75.

Studies in History and Philosophy of Science Part A, 3(2):167– 182, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900222. See [Cla68].

Gray:1996:GSP

[Gra96] Chris Hables Gray. The game of science: As played by Jean-François Lyotard. Studies in History and Philosophy of Science Part A, 27(3):367–380, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000577.

Granada:2004:ACB

[Gra04] Miguel A. Granada. Aristotle, Copernicus, Bruno: centrality, the principle of movement and the extension of the Universe. Studies in History and Philosophy of Science Part A, 35(1):91– 114, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000931.

Grandi:2005:TRG

[Gra05] Giovanni B. Grandi. Thomas Reid’s geometry of visibles and the parallel postulate. Studies in History and Philosophy of Science Part A, 36(1):79–103, March 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104001219.

Gramelsberger:2011:WDN

[Gra11] Gabriele Gramelsberger. What do numerical (climate) models really represent? Studies in History and Philosophy of Science Part A, 42(2):296–302, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110001160.

Granato:2013:SHB

[Gra13] Marcus Granato. Scientific heritage in Brazil. Studies in History and Philosophy of Science Part A, 44(4):690–699, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000502.

Gradowski:2024:FMG

[Gra24] Laura Gradowski. From fringe to mainstream: the Garcia effect. Studies in History and Philosophy of Science Part A, 103(??):114– 122, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812300170X.

Gregory:1992:NLN

[Gre92a] Frederick Gregory. Nature lost?: natural science and the German theological traditions of the nineteenth century. Harvard University Press, Cambridge, MA, USA, 1992. ISBN 0-674-60483-0. viii + 341 pp. LCCN BL245 .G74 1992.

Grell:1992:PNP

[Gre92b] Ole Peter Grell. Protestantism, natural philosophy, and the scientific revolution. Studies in History and Philosophy of Science Part A, 23(3):519–527, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290009U.

Gregory:1996:AOP

[Gre96] Andrew Gregory. Astronomy and observation in Plato’s Republic. Studies in History and Philosophy of Science Part A, 27(4):451– 471, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368196000623.

Gregis:2019:AAM

[Gré19] Fabien Grégis. Assessing accuracy in measurement: the dilemma of safety versus precision in the adjustment of the fundamental physical constants. Studies in History and Philosophy of Science Part A, 74(??):42–55, April 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368117302972.

Greif:2021:AAB

[Gre21] Hajo Greif. Adaptation and its analogues: Biological categories for biosemantics. Studies in History and Philosophy of Science Part A, 90(??):298–307, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001722.

Greenspan:2022:HAS

[Gre22] Lauren Greenspan. Holography, application, and string theory’s changing nature. Studies in History and Philosophy of Science Part A, 94(??):72–86, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000723.

Greene:2024:PCK

[Gre24a] Jeremy Greene. Pushing cool by Keith wailoo: Big data and bigger disparities. Studies in History and Philosophy of Science Part A, 103(??):3–4, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123001395.

Greenwood:2024:PCR

[Gre24b] Samara Greenwood. The problem of context revisited: Moving beyond the resources model. Studies in History and Philosophy of Science Part A, 105(??):126–137, June 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000608.

Griffin:1991:NEG

[Gri91] Nicholas Griffin. Non-Euclidean geometry: Still some problems for Kant. Studies in History and Philosophy of Science Part A, 22(4):661–663, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190038T.

See [Lea91, Gri91].

Grimm:2010:GE

[Gri10] Stephen R. Grimm. The goal of explanation. Studies in History and Philosophy of Science Part A, 41(4):337–344, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000610.

Grosholz:1987:SUP

[Gro87] Emily Grosholz. Some uses of proportion in Newton’s Principia, book I: a case study in applied mathematics. Studies in History and Philosophy of Science Part A, 18(2):209–220, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900185. See [New87, NCW99, New09].

Gross:1990:BRB

[Gro90] Alan G. Gross. Book review: Shaping written knowledge: The genre and activity of the experimental article in science: Charles Bazerman, (Madison: University of Wisconsin Press, 1988). Paper $17.50. Studies in History and Philosophy of Science Part A, 21(2):341–349, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900313. See [Baz88].

Grosholz:2005:BFG

[Gro05] Emily R. Grosholz. Berzelian formulas as generative paper tools. Studies in History and Philosophy of Science Part A, 36(2):411– 417, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000233.

Grozier:2020:SPL

[Gro20] Jim Grozier. Should physical laws be unit-invariant? Studies in History and Philosophy of Science Part A, 80(??):9–18, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300420.

Grunbaum:1978:PTA

[Grü78] Adolf Grünbaum. Poincaré’s thesis that any and all stellar parallax findings are compatible with the Euclideanism of the pertinent astronomical 3-space. Studies in History and Philosophy of Science Part A, 9(4):313–318, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900110.

See reply [O’G78].

Gaukroger:2002:HPO

[GS02] Stephen Gaukroger and John Schuster. The hydrostatic paradox and the origins of Cartesian dynamics. Studies in History and Philosophy of Science Part A, 33(3):535–572, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000262.

Gamini:2022:PSQ

[GS22] Amir-Mohammad Gamini and Mohammad-Mahdi Sadrforati. The principle of simplicity for Qut.b al-Din Shirazi. Studies in History and Philosophy of Science Part A, 91(??):60–65, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001916.

Granada:2005:CFD

[GT05] Miguel A. Granada and Dario Tessicini. Copernicus and Fracastoro: the dedicatory letters to Pope Paul III, the history of astronomy, and the quest for patronage. Studies in History and Philosophy of Science Part A, 36(3):431–476, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810500049X.

Guala:2001:BEM

[Gua01] Francesco Guala. Building economic machines: The FCC auctions. Studies in History and Philosophy of Science Part A, 32 (3):453–477, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000085.

Guasparri:2013:ENC

[Gua13] Andrea Guasparri. Explicit nomenclature and classification in Pliny’s Natural History XXXII. Studies in History and Philosophy of Science Part A, 44(3):347–353, September 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000368.

Guerrini:2017:PBE

[Gue17] Anita Guerrini. Philosophical bodies in early modern Europe. Studies in History and Philosophy of Science Part A, 61(??):61– 65, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116301017.

Guicciardini:2004:BRB

[Gui04a] Niccolò Guicciardini. Book review: Isaac Newton’s natural philosophy: Jed. Z. Buchwald, & I. Bernard Cohen (Eds.); MIT Press, Cambridge, MA & London, 2001, pp. xx + 354, Price £32.95 US $50.00, ISBN 0-262-02477-2 hardback. Studies in History and Philosophy of Science Part A, 35(3):670–674, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000627. See [BC01].

Guicciardini:2004:INP

[Gui04b] Niccolò Guicciardini. Isaac Newton and the publication of his mathematical manuscripts. Studies in History and Philosophy of Science Part A, 35(3):455–470, September 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000494.

Gundersen:2018:SED

[Gun18] Torbjørn Gundersen. Scientists as experts: A distinct role? Studies in History and Philosophy of Science Part A, 69(??): 52–59, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811730208X.

Guralp:2019:EEE

[Gur19] Genco Guralp. Exploratory experimentation: Essay review of Exploratory experiments: Ampére, Faraday, and the origins of electrodynamics, by Friedrich Steinle, Friedrich Steinle. University of Pittsburgh Press, Pittsburgh (2016), pp. x+494, price US$65 hardback, ISBN-13: 978-0-8229-4450-8. Studies in History and Philosophy of Science Part A, 73(??):72–76, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118302231.

Gutting:1973:CSS

[Gut73] Gary Gutting. Conceptual structures and scientific change. Studies in History and Philosophy of Science Part A, 4(3):209–230, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900083.

Guyer:2008:PKA

[Guy08] Paul Guyer. The psychology of Kant’s aesthetics. Studies in History and Philosophy of Science Part A, 39(4):483–494, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000824.

Garcia-Valdecasas:2024:BFC

[GVD24] Miguel García-Valdecasas and Terrence W. Deacon. Biological functions are causes, not effects: a critique of selected effects theories. Studies in History and Philosophy of Science Part A, 103 (??):20–28, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123001607.

Giere:1973:FSM

[GW73] Ronald N. Giere and Richard S. Westfall, editors. Foundations of scientific method: the nineteenth century. Indiana University Press, Bloomington, IN, USA, 1973. ISBN 0-253-32400-9. LCCN Q175 .F72 1973.

Gervais:2015:ROE

[GW15] Raoul Gervais and Erik Weber. The role of orientation experiments in discovering mechanisms. Studies in History and Philosophy of Science Part A, 54(??):46–55, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001089.

Grune-Yanoff:2011:MPI

[GY11] Till Grüne-Yanoff. Models as products of interdisciplinary exchange: Evidence from evolutionary game theory. Studies in History and Philosophy of Science Part A, 42(2):386–397, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001251.

Grune-Yanoff:2014:IIM

[GYM14] Till Grüne-Yanoff and Uskali Mäki. Introduction: Interdisciplinary model exchanges. Studies in History and Philosophy of Science Part A, 48(??):52–59, December 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000752.

Hacking:1989:BRB

[Hac89] Ian Hacking. Book review: The divided circle: A history of instruments for astronomy, navigation and surveying: J. A. Bennett (Phaidon/Christie’s: Oxford, 1987), $224 pp. Cloth £45.00. Studies in History and Philosophy of Science Part A, 20(2):265– 270, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818990006X. See [Ben87].

Hackmann:1991:LRM

[Hac91] Willem Hackmann. Lightning rods and model experiments: Franklin’s science comes of age. Studies in History and Philosophy of Science Part A, 22(4):679–684, December 1991. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936819190041P.

Hacking:1992:SHP

[Hac92] Ian Hacking. ‘style’ for historians and philosophers. Studies in History and Philosophy of Science Part A, 23(1):1–20, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290024Z.

Hacking:2012:LTR

[Hac12] Ian Hacking. ‘Language, Truth and Reason’ 30 years later. Studies in History and Philosophy of Science Part A, 43(4):599–609, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000441.

Haddock:2004:RSP

[Had04] Adrian Haddock. Rethinking the “strong programme” in the sociology of knowledge. Studies in History and Philosophy of Science Part A, 35(1):19–40, March 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000979.

Haggerty:2001:NMI

[Hag01] Kevin D. Haggerty. Negotiated measures: The institutional micropolitics of official criminal justice statistics. Studies in History and Philosophy of Science Part A, 32(4):705–722, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000188.

Halabi:2005:UAJ

[Hal05] Selman Halabi. A useful anachronism: John Locke, the corpuscular philosophy, and inference to the best explanation. Studies in History and Philosophy of Science Part A, 36(2):241– 259, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000166.

Hallberg:2012:GPS

[Hal12] Margareta Hallberg. Gender and philosophy of science: the case of Mary Hesse. Studies in History and Philosophy of Science Part A, 43(2):333–340, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001397.

Hamm:2000:SRH

[Ham00] E. P. Hamm. Shipwrecked romanticism? Henrich Steffens and the career of Naturphilosophie. Studies in History and Philosophy of Science Part A, 31(3):509–536, September 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000212.

Hamid:2023:AHE

[Ham23] Nabeel Hamid. Anthropology and history in the early Dilthey. Studies in History and Philosophy of Science Part A, 100(??):90– 98, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000924.

Hands:2005:YWS

[Han05] D. Wade Hands. You want the social? you can’t handle the social! Mirowski on the secret history of scientific philosophy. Studies in History and Philosophy of Science Part A, 36(4):726– 733, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000737.

Hansson:2006:DTF

[Han06] Sven Ove Hansson. Defining technical function. Studies in History and Philosophy of Science Part A, 37(1):19–22, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105001056.

Hansson:2007:WTS

[Han07]  Sven Ove Hansson. What is technological science?  Stud-

ies in History and Philosophy of Science Part A, 38(3):523–527, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000362.

Hansson:2017:SDF

[Han17] Sven Ove Hansson. Science denial as a form of pseudoscience. Studies in History and Philosophy of Science Part A, 63(??): 39–47, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300681.

Hands:2021:MFU

[Han21] D. Wade Hands. The many faces of unification and pluralism in economics: the case of Paul Samuelson’s Foundations of Economic Analysis. Studies in History and Philosophy of Science Part A, 88(??):209–219, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000911.

Harre:1974:NMH

[Har74] Rom Harré. A note on Ms. Horton’s defence of Bacon. Studies in History and Philosophy of Science Part A, 5(3):305–306, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368174900041.

Harwood:1993:SST

[Har93] Jonathan Harwood. Styles of scientific thought: the German genetics community, 1900–1933. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1993. ISBN 0-226-31881-8 (hardcover), 0-226-31882-6 (paper). xix + 423 pp. LCCN QH428.2.G3 H37 1993. URL http://www.loc.

gov/catdir/description/uchi051/92015321.html;   http://
www.loc.gov/catdir/enhancements/fy0609/92015321-b.html;
 http://www.loc.gov/catdir/toc/uchi051/92015321.html.

Hartley:1996:LAN

[Har96] Beryl Hartley. The living academies of nature: scientific experiment in learning and communicating the new skills of early nineteenthcentury landscape painting. Studies in History and Philosophy of Science Part A, 27(2):149–180, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000224.

Harris:2002:COS

[Har02] Paul L. Harris. Checking our sources: the origins of trust in testimony. Studies in History and Philosophy of Science Part A, 33(2): 315–333, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000079.

Harris:2008:CRR

[Har08] Martha L. Harris. Chemical reductionism revisited: Lewis, Pauling and the physico-chemical nature of the chemical bond. Studies in History and Philosophy of Science Part A, 39(1):78– 90, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810700088X.

Harker:2010:TAS

[Har10] David Harker. Two arguments for scientific realism unified. Studies in History and Philosophy of Science Part A, 41(2):192– 202, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811000018X.

Harari:2012:STP

[Har12] Orna Harari. Simplicius on Tekmeriodic proofs. Studies in History and Philosophy of Science Part A, 43(2):366–375, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111001439.

Harding:2019:SFL

[Har19a] Sandra Harding. State of the field: Latin American decolonial philosophies of science. Studies in History and Philosophy of Science Part A, 78(??):48–63, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302541.

Harnagel:2019:MLA

[Har19b] Audrey Harnagel. A mid-level approach to modeling scientific communities. Studies in History and Philosophy of Science Part A, 76(??):49–59, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300049.

Harkema:2024:BTM

[Har24] Scott Harkema. Berkeley on true motion. Studies in History and Philosophy of Science Part A, 105(??):165–174, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000633.

Haskell:1997:AHT

[Has97] Yasmin Haskell. All the heavens, truthfully represented, it can enclose with its verses. Studies in History and Philosophy of Science Part A, 28(4):681–697, December 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000186.

Hatfield:1979:FGD

[Hat79] Gary C. Hatfield. Force (God) in Descartes’ physics. Studies in History and Philosophy of Science Part A, 10(2):113– 140, June 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936817990013X.

Hatfield:1988:RCS

[Hat88] Gary Hatfield. Representation and content in some (actual) theories of perception. Studies in History and Philosophy of Science Part A, 19(2):175–214, June 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818890026X.

Hatfield:1990:NNT

[Hat90] Gary C. (Gary Carl) Hatfield. The natural and the normative: theories of spatial perception from Kant to Helmholtz. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-08086-9. xii + 366 pp. LCCN BF469 .H34 1990.

Hatfield:2007:PSD

[Hat07] Gary Hatfield. The passions of the soul and Descartes’s machine psychology. Studies in History and Philosophy of Science Part A, 38(1):1–35, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106001129.

Hattiangadi:2010:LT

[Hat10] Anandi Hattiangadi. The love of truth. Studies in History and Philosophy of Science Part A, 41(4):422–432, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000580.

Haufe:2013:WDF

[Hau13] Chris Haufe. Why do funding agencies favor hypothesis testing? Studies in History and Philosophy of Science Part A, 44(3):363– 374, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811300037X.

Haufe:2016:ITP

[Hau16] Chris Haufe. Introduction: Testing philosophical theories. Studies in History and Philosophy of Science Part A, 59(??):68– 73, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811630022X.

Haueis:2021:DCC

[Hau21] Philipp Haueis. The death of the cortical column? Patchwork structure and conceptual retirement in neuroscientific practice. Studies in History and Philosophy of Science Part A, 85(??):101– 113, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301825.

Hauptmann:2022:BRS

[Hau22] Emily Hauptmann. Book review: Social science for what? Battles over public funding for the “Other Sciences” at the National Science Foundation by Mark Solovey: On the margins of the margins: Political science at the NSF. Studies in History and Philosophy of Science Part A, 94(??):206–207, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000401. See [Lin22a, Rot22, Tur22b]

and reply [Sol22].

Havstad:2021:BRS

[Hav21] Joyce C. Havstad. Book review: Science and Moral Imagination by Matthew J. Brown: Practice makes perfect. Studies in History and Philosophy of Science Part A, 89(??):297– 298, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000406.

Hawley:1996:TKM

[Haw96] Katherine Hawley. Thomas S. Kuhn’s mysterious worlds. Studies in History and Philosophy of Science Part A, 27(2):291– 300, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000380.

Hawley:2010:TKH

[Haw10] Katherine Hawley. Testimony and knowing how. Studies in History and Philosophy of Science Part A, 41(4):397–404, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000609.

Haycock:2004:LLT

[Hay04] David Boyd Haycock. ‘the long-lost truth’: Sir Isaac Newton and the Newtonian pursuit of ancient knowledge. Studies in History and Philosophy of Science Part A, 35(3):605–623, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000573.

Hazelwood:2023:NLF

[Haz23]  Caleb Hazelwood.  Newton’s “law-first” epistemology and

“matter-first” metaphysics. Studies in History and Philosophy of Science Part A, 101(??):40–47, October 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001061.

Hesketh:2024:UDA

[HBR24] Ian Hesketh, Ruth Barton, and Evelleen Richards. Down under Darwin: Australasian perspectives on Darwin studies. Studies in History and Philosophy of Science Part A, 103(??):69– 76, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001619.

Hangel:2023:PQM

[HC23] Nora Hangel and Christopher ChoGlueck. On the pursuitworthiness of qualitative methods in empirical philosophy of science. Studies in History and Philosophy of Science Part A, 98(??): 29–39, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001868.

Herfeld:2019:DSI

[HD19] Catherine Herfeld and Malte Doehne. The diffusion of scientific innovations: a role typology. Studies in History and Philosophy of Science Part A, 77(??):64–80, October 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117303321.

Healey:1994:NPC

[Hea94] Richard Healey. Nonseparable processes and causal explanation. Studies in History and Philosophy of Science Part A, 25(3):337– 374, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900574.

Heap:2024:MHR

[Hea24] Jodie Lee Heap. Mary Hesse on the role of the human imagination in the philosophy and practice of science. Studies in History and Philosophy of Science Part A, 107(??):73–81, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001067.

Heelan:1998:SHN

[Hee98] Patrick A. Heelan. The scope of hermeneutics in natural science. Studies in History and Philosophy of Science Part A, 29(2):

273–298, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000028.

Heesen:2019:CIM

[Hee19] Remco Heesen. The credit incentive to be a maverick. Studies in History and Philosophy of Science Part A, 76(??):5–12, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811730314X.

Heeney:2021:PPH

[Hee21] Catherine Heeney. Problems and promises: How to tell the story of a Genome Wide Association Study? Studies in History and Philosophy of Science Part A, 89(??):1–10, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000868.

Heimann:1970:MFS

[Hei70] P. M. Heimann. Molecular forces, statistical representation and Maxwell’s demon. Studies in History and Philosophy of Science Part A, 1(3):189–211, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900099.

Heimann:1974:HKM

[Hei74] P. M. Heimann. Helmholtz and Kant: The metaphysical foundations of Über die Erhaltung der Kraft. Studies in History and Philosophy of Science Part A, 5(3):205–238, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368174900016.

Heimann:1975:BRB

[Hei75] P. M. Heimann. Book review: Foundations of scientific method: the nineteenth century: Ronald N. Giere and Richard S. Westfall, eds., Bloomington and London: Indiana University Press, 1973. $10.00. Studies in History and Philosophy of Science Part A, 5 (4):397–399, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368175900126. See [GW73].

Heidelberger:1980:TLR

[Hei80] Michael Heidelberger. Towards a logical reconstruction of revolutionary change: The case of Ohm as an example. Studies in History and Philosophy of Science Part A, 11(2):103– 121, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900175.

Heis:2014:RFP

[Hei14] Jeremy Heis. Realism, functions, and the a priori: Ernst Cassirer’s philosophy of science. Studies in History and Philosophy of Science Part A, 48(??):10–19, December 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000776.

Heit:2016:RRF

[Hei16] Helmut Heit. Reasons for relativism: Feyerabend on the ‘rise of rationalism’ in ancient Greece. Studies in History and Philosophy of Science Part A, 57(??):70–78, June 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001673.

Heinbokel:2024:PRS

[Hei24] Timm Heinbokel. The pragmatist roots of scientific medicine: Reassessing Abraham Flexner’s report on medical education. Studies in History and Philosophy of Science Part A, 106(??):186– 195, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000670.

Heller:2016:BRP

[Hel16] Lisa Heller. Between relativism and pluralism: Philosophical and political relativism in Feyerabend’s late work. Studies in History and Philosophy of Science Part A, 57(??):96–105, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001715.

Hendry:1983:MBD

[Hen83] John Hendry. Monopoles before Dirac. Studies in History and Philosophy of Science Part A, 14(1):81–87, March 1983. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368183900079.

Hendry:1984:EWR

[Hen84] John Hendry. The evolution of William Rowan Hamilton’s view of algebra as the science of pure time. Studies in History and Philosophy of Science Part A, 15(1):63–81, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818490030X.

Hentschel:1985:FVM

[Hen85] Klaus Hentschel. On Feyerabend’s version of ‘Mach’s theory of research and its relation to Einstein’. Studies in History and Philosophy of Science Part A, 16(4):387–394, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900196.

Hentschel:1992:EAT

[Hen92] Klaus Hentschel. Einstein’s attitude towards experiments: Testing relativity theory 1907–1927. Studies in History and Philosophy of Science Part A, 23(4):593–624, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290014W.

Henderson:1998:GN

[Hen98] Fergus Henderson. Goethe’s ‘Naturphilosophie’. Studies in History and Philosophy of Science Part A, 29(1):143–153, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368197000198.

Hendry:2006:SC

[Hen06] Robin Findlay Hendry. Substantial confusion. Studies in History and Philosophy of Science Part A, 37(2):322–336, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000288.

Henry:2011:GGD

[Hen11] John Henry. Gravity and De gravitatione: the development of Newton’s ideas on action at a distance. Studies in History and Philosophy of Science Part A, 42(1):11–27, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001044.

Henry:2014:NAD

[Hen14a] John Henry. Newton and action at a distance between bodies — a response to Andrew Janiak’s “Three concepts of causation in Newton”. Studies in History and Philosophy of Science Part A, 47 (??):91–97, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811400020X. See [Jan13a].

Henschen:2014:KCL

[Hen14b] Tobias Henschen. Kant on causal laws and powers. Studies in History and Philosophy of Science Part A, 48(??):20–29, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000788.

Hendry:2016:IP

[Hen16] Robin Findlay Hendry. Immanent philosophy of X. Studies in History and Philosophy of Science Part A, 55(??):36–42, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001016.

Hendry:2021:SSE

[Hen21] Robin Findlay Hendry. Structure, scale and emergence. Studies in History and Philosophy of Science Part A, 85(??):44–53, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301631.

Henderson:2022:PIB

[Hen22] Leah Henderson. Putting inference to the best explanation into context. Studies in History and Philosophy of Science Part A, 94(??):167–176, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000905.

Henderson:2023:BRI

[Hen23] Leah Henderson. Book review: The instrument of science: Scientific anti-realism revitalised by Darrell Rowbottom: Reorienting the scientific realism debate. Studies in History and Philosophy of Science Part A, 98(??):4–6, April 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000249.

Hepburn:2010:EVV

[Hep10] Brian Hepburn. Euler, vis viva, and equilibrium. Studies in History and Philosophy of Science Part A, 41(2):120–127, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000233.

HerriesDavies:1969:EDHa

[Her69a] G. L. Herries Davies. The Earth in decay: a history of British geomorphology, 1578–1878. An Oldbourne book; [History of science library]. Macdonald and Co., London, 1969. ISBN 0-356-02676-0. xvi + 390 pp. LCCN GB401 .H38 1969b.

HerriesDavies:1969:EDHb

[Her69b] Gordon L. Herries Davies. The Earth in decay; a history of British geomorphology, 1578–1878. American Elsevier Pub. Co., New York, NY, USA, 1969. ISBN 0-444-19701-X. xvi + 390 pp. LCCN GB401 .H38 1969.

Hersch:2015:EEI

[Her15] Gil Hersch. Experimental economics’ inconsistent ban on deception. Studies in History and Philosophy of Science Part A, 52(??): 13–19, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000461.

Hesse:1974:SSI

[Hes74] Mary B. Hesse. The structure of scientific inference. University of California Press, Berkeley, CA, USA, 1974. ISBN 0-520-02582-2. vii + 309 pp. LCCN Q175 .H432.

Hesse:1975:BSI

[Hes75] Mary Hesse. Bayesianism and scientific inference. Studies in History and Philosophy of Science Part A, 5(4):367–370, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900084.

Hesse:1982:CPD

[Hes82] Mary Hesse. Comments on the papers of David Bloor and Steven Lukes. Studies in History and Philosophy of Science Part A, 13(4):325–331, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900176.

See [Blo82a, Buc82, Blo82b, Luk82, Blo82c].

Hesketh:2021:NCD

[Hes21] Ian Hesketh. Narratives of Charles Darwin Down Under. Studies in History and Philosophy of Science Part A, 88(??):303– 311, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812100090X.

Hetherington:1996:PEI

[Het96] Norriss S. Hetherington. Plato and Eudoxus: Instrumentalists, realists, or prisoners of themata? Studies in History and Philosophy of Science Part A, 27(2):271–289, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000461.

Heunen:2014:FRR

[Heu14] Chris Heunen. Book review: Foundations of Relational Realism: a Topological Approach to Quantum Mechanics and the Philosophy of Nature, Michael Epperson, Elias Zafiris. Lexington Books (2013), 419pp., ISBN: 978-0-7391-8032-7. Studies in History and Philosophy of Science Part A, 48 (part A)(??):99–100, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000653.

Hon:2005:PBB

[HG05] Giora Hon and Bernard R. Goldstein. From proportion to balance: the background to symmetry in science. Studies in History and Philosophy of Science Part A, 36(1):1–21, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104001177.

Hon:2021:MRT

[HG21] Giora Hon and Bernard R. Goldstein. Maxwell’s role in turning the concept of model into the methodology of modeling. Studies in History and Philosophy of Science Part A, 88(??):321– 333, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000443.

Hoyningen-Huene:1987:CDC

[HH87] Paul Hoyningen-Huene. Context of discovery and context of justification. Studies in History and Philosophy of Science Part A, 18 (4):501–515, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368187900057.

Hoyningen-Huene:1990:KCI

[HH90] Paul Hoyningen-Huene. Kuhn’s conception of incommensurability. Studies in History and Philosophy of Science Part A, 21(3): 481–492, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090006T.

Hoyningen-Huene:1995:TLP

[HH95] Paul Hoyningen-Huene. Two letters of Paul Feyerabend to Thomas S. Kühn on a draft of the Structure of Scientific Revolutions. Studies in History and Philosophy of Science Part A, 26(3):353–387, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000058. See [Kuh62, Kuh70,

Kuh96, KH12].

Halonen:2000:AE

[HH00] Ilpo Halonen and Jaakko Hintikka. Aristotelian explanations. Studies in History and Philosophy of Science Part A, 31(1):125– 136, March 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368199000321.

Hoyningen-Huene:2006:MLP

[HH06] Paul Hoyningen-Huene. More letters by Paul Feyerabend to Thomas S. Kuhn on Proto-Structure. Studies in History and Philosophy of Science Part A, 37(4):610–632, December 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000781.

Hoyningen-Huene:2011:NKE

[HHL11] Paul Hoyningen-Huene and Simon Lohse. On naturalizing Kuhn’s essential tension. Studies in History and Philosophy of Science Part A, 42(1):215–218, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000890.

Hoyningen-Huene:2009:ROR

[HHO09] Paul Hoyningen-Huene and Eric Oberheim. Reference, ontological replacement and neo-Kantianism: a reply to Sankey. Studies in History and Philosophy of Science Part A, 40(2):203– 209, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810900017X. See [San09b].

Hoyningen-Huene:1996:I

[HHOA96] Paul Hoyningen-Huene, Eric Oberheim, and Hanne Andersen. On incommensurability. Studies in History and Philosophy of Science Part A, 27(1):131–141, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000208.

Hine:2012:ANT

[Hin12] Harry Hine. Aetna: a new translation based on the text of F. R. D. Goodyear. Studies in History and Philosophy of Science Part A, 43(2):316–325, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001282.

Hindriks:2013:EUU

[Hin13] Frank Hindriks. Explanation, understanding, and unrealistic models. Studies in History and Philosophy of Science Part A, 44(3): 523–531, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112001148.

Hinterwaldner:2015:MBW

[Hin15] Inge Hinterwaldner. Model building with wind and water: Friedrich Ahlborn’s photo-optical flow analysis. Studies in History and Philosophy of Science Part A, 49(??):1–17, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000934.

Hendry:2016:IHP

[HK16] Robin Findlay Hendry and Ian James Kidd. Introduction: Historiography and the philosophy of the sciences. Studies in History and Philosophy of Science Part A, 55(??):1–2, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000965.

Hoefer:2021:MEM

[HK21] Carl Hoefer and Alexander Krauss. Measures of effectiveness in medical research: Reporting both absolute and relative measures. Studies in History and Philosophy of Science Part A, 88 (??):280–283, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000960. See response

[Ste22b].

Haueis:2022:MIS

[HK22] Philipp Haueis and Lena Kästner. Mechanistic inquiry and scientific pursuit: the case of visual processing. Studies in History and Philosophy of Science Part A, 93(??):123–135, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000589.

Houkes:2011:DNC

[HKMV11] Wybo Houkes, Peter Kroes, Anthonie Meijers, and Pieter E. Vermaas. Dual-nature and collectivist frameworks for technical artefacts: a constructive comparison. Studies in History and Philosophy of Science Part A, 42(1):198–205, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000816.

Hopwood-Lewis:2013:GBS

[HL13] Jonathan Hopwood-Lewis. Griffith Brewer, “The Wright brothers’ Boswell”: Patent management and the British aviation industry, 1903–1914. Studies in History and Philosophy of Science Part A, 44(2):259–268, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112001069.

Herfeld:2019:KTC

[HL19] Catherine Herfeld and Chiara Lisciandra. Knowledge transfer and its contexts. Studies in History and Philosophy of Science Part A, 77(??):1–10, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118303273.

Hopwood-Lewis:2013:PPP

[HLM13] Jonathan Hopwood-Lewis and Christine MacLeod. Patents, publicity and priority: the Aeronautical Society of Great Britain, 1897– 1919. Studies in History and Philosophy of Science Part A, 44(2): 212–221, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001070.

Hampe:1988:TCR

[HM88] M. Hampe and S. R. Morgan. Two consequences of Richard Dawkins’ view of genes and organisms. Studies in History and Philosophy of Science Part A, 19(1):119–138, March 1988. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368188900222.

Houkes:2006:OAH

[HM06] Wybo Houkes and Anthonie Meijers. The ontology of artefacts: the hard problem. Studies in History and Philosophy of Science Part A, 37(1):118–131, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105001160.

Hyrkas:2024:OVF

[HM24] Eve-Riina Hyrkäs and Mikko Myllykangas. Obesity and the vitality of food in Finland, ca. 1950–1970. Studies in History and Philosophy of Science Part A, 106(??):99–108, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000980.

Helbig:2016:MEK

[HN16] Daniela Helbig and Dalia Nassar. The metaphor of epigenesis: Kant, Blumenbach and Herder. Studies in History and Philosophy of Science Part A, 58(??):98–107, August 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300152.

Hoch:1987:IVI

[Hoc87] Paul K. Hoch. Institutional versus intellectual migrations in the nucleation of new scientific specialties. Studies in History and Philosophy of Science Part A, 18(4):481–500, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900045.

Hochstein:2016:GCA

[Hoc16] Eric Hochstein. Giving up on convergence and autonomy: Why the theories of psychology and neuroscience are codependent as well as irreconcilable. Studies in History and Philosophy of Science Part A, 56(??):135–144, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001399.

Hoering:1980:JR

[Hoe80] Walter Hoering. On judging rationality. Studies in History and Philosophy of Science Part A, 11(2):123–136, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368180900187.

Hoefer:1994:ESM

[Hoe94] Carl Hoefer. Einstein’s struggle for a Machian gravitation theory. Studies in History and Philosophy of Science Part A, 25(3): 287–335, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900566.

Hoffmann:2013:SLM

[Hof13] Christoph Hoffmann. Superpositions: Ludwig Mach and Étienne-Jules Marey’s studies in streamline photography. Studies in History and Philosophy of Science Part A, 44(1):1–11, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000714.

Hollander:1983:WWJ

[Hol83] Samuel Hollander. William Whewell and John Stuart Mill on the methodology of political economy. Studies in History and Philosophy of Science Part A, 14(2):127–168, June 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368183900171.

Holmes:2021:SSS

[Hol21a] Tarquin Holmes. Science, sensitivity and the sociozoological scale: Constituting and complicating the human–animal boundary at the 1875 Royal Commission on Vivisection and beyond. Studies in History and Philosophy of Science Part A, 90(??):194–207, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812100159X.

Holmes:2021:DME

[Hol21b] Travis L. Holmes. Distinctively mathematical explanation and the problem of directionality: a quasi-erotetic solution. Studies in History and Philosophy of Science Part A, 87(??):13– 21, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000200.

Holmes:2022:HRP

[Hol22] Travis Holmes. How revealed preference theory can be explanatory. Studies in History and Philosophy of Science Part A, 91 (??):20–27, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001710.

Homberg:1702:EC

[Hom02] Wilhelm Homberg. Essais de chimie. Histoire de l’Académie royale des sciences avec les Mémoires de Mathématique & de Physique, Mémoires, ??(??):33–52, 1702.

Honner:1982:TPN

[Hon82] John Honner. The transcendental philosophy of Niels Bohr. Studies in History and Philosophy of Science Part A, 13(1):1– 29, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368182900024.

Hon:1987:HHE

[Hon87a] Giora Hon. H. Hertz: ‘The electrostatic and electromagnetic properties of the cathode rays are either nil or very feeble.’ (1883) a case-study of an experimental error. Studies in History and Philosophy of Science Part A, 18(3):367–382, September 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900252.

Hones:1987:NWC

[Hon87b] Michael J. Hones. The neutral-weak-current experiments: a philosophical perspective. Studies in History and Philosophy of Science Part A, 18(2):221–251, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900197.

Hon:1989:TTE

[Hon89] Giora Hon. Towards a typology of experimental errors: An epistemological view. Studies in History and Philosophy of Science Part A, 20(4):469–504, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900204.

Hooker:2013:GSN

[Hoo13] Cliff Hooker. Georg Simmel and naturalist interactivist epistemology of science. Studies in History and Philosophy of Science Part A, 44(3):311–317, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000344.

Hopwood:1994:BRB

[Hop94] Nick Hopwood. Book review: Styles of scientific thought: The German genetics community 1900–1933: Jonathan Harwood (Chicago and London: University of Chicago Press, 1993), xix + 423 pp.

ISBN 0-226-31881-8 Cloth $74.75/£51.95, ISBN 0-226-31882-6 Paperback $27.50/£17.95. Studies in History and Philosophy of Science Part A, 25(2):237–250, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900299.

See [Har93].

Horton:1973:DFB

[Hor73] Mary Horton. In defence of Francis Bacon: a criticism of the critics of the inductive method. Studies in History and Philosophy of Science Part A, 4(3):241–278, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368173900101.

Houkes:2006:KAF

[Hou06] Wybo Houkes. Knowledge of artefact functions. Studies in History and Philosophy of Science Part A, 37(1):102–113, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105001147.

Howard:1985:ELS

[How85] Don Howard. Einstein on locality and separability. Studies in History and Philosophy of Science Part A, 16(3):171–201, September 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368185900019.

Howson:1990:PH

[How90] Colin Howson. The poverty of historicism. Studies in History and Philosophy of Science Part A, 21(1):173–179, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900195.

Howell:1998:RBI

[How98] Kenneth J. Howell. The role of biblical interpretation in the cosmology of Tycho Brahe. Studies in History and Philosophy of Science Part A, 29(4):515–537, December 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000326.

Howard:2004:PFJ

[How04] Jill Howard. ‘Physics and fashion’: John Tyndall and his audiences in mid-Victorian Britain. Studies in History and Philosophy of Science Part A, 35(4):729–758, December 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000846.

Howson:2013:HT

[How13] Colin Howson. Hume’s Theorem. Studies in History and Philosophy of Science Part A, 44(3):339–346, September 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000186.

Howson:2015:DHN

[How15] Colin Howson. David Hume’s no-miracles argument begets a valid No-Miracles Argument. Studies in History and Philosophy of Science Part A, 54(??):41–45, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001041.

Hamm:2001:MPP

[HR01] E. P. Hamm and Alan W. Richardson. Measurement of the people, by the people, and for the people. Studies in History and Philosophy of Science Part A, 32(4):607–612, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000176.

Hofer-Szabo:2022:TCN

[HS22] Gábor Hofer-Szabó. Two concepts of noncontextuality in quantum mechanics. Studies in History and Philosophy of Science Part A, 93(??):21–29, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000437.

Hemmo:2023:MBC

[HS23a] Meir Hemmo and Orly Shenker. Is the mind in the brain in contemporary computational neuroscience? Studies in History and Philosophy of Science Part A, 100(??):64–80, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000870.

Heuer:2023:PCR

[HS23b] Karl Heuer and Deniz Sarikaya. Paving the cowpath in research within pure mathematics: a medium level model based on text driven variations. Studies in History and Philosophy of Science Part A, 100(??):39–46, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123000869.

Hubert:2023:AAE

[HS23c] Mario Hubert and Charles T. Sebens. Absorbing the arrow of electromagnetic radiation. Studies in History and Philosophy of Science Part A, 99(??):10–27, June 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812300002X.

Hu:2021:HSU

[Hu21] Xingming Hu. Hempel on scientific understanding. Studies in History and Philosophy of Science Part A, 88(??):164–171, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812100073X.

Hubner:1980:CTH

[Hüb80] Kurt Hübner. The concept of truth in a historistic theory of science. Studies in History and Philosophy of Science Part A, 11(2): 145–151, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900205.

Hudson:1989:JJR

[Hud89] Rob Hudson. James Jeans and radiation theory. Studies in History and Philosophy of Science Part A, 20(1):57–76, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900344.

Hudson:1994:BIC

[Hud94] Robert G. Hudson. Background independence and the causation of observations. Studies in History and Philosophy of Science Part A, 25(4):595–612, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900493.

Hull:1989:ME

[Hul89] David L. Hull. The metaphysics of evolution. SUNY series in philosophy and biology. State University of New York Press, Albany, 1989. ISBN 0-7914-0211-8, 0-7914-0212-6 (paperback). viii + 331 pp. LCCN QH371 .H85 1989.

Humphreys:2019:KTA

[Hum19] Paul Humphreys. Knowledge transfer across scientific disciplines. Studies in History and Philosophy of Science Part A, 77(??):112– 119, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302674.

Hunter:2003:CJF

[Hun03] Michael Hunter. The correspondence of John Flamsteed, first Astronomer Royal. Studies in History and Philosophy of Science Part A, 34(4):805–820, December 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000694.

Huneman:2018:RVE

[Hun18] Philippe Huneman. Realizability and the varieties of explanation. Studies in History and Philosophy of Science Part A, 68(??): 37–50, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118300086.

Hunt:2021:IWE

[Hun21] Josh Hunt. Interpreting the Wigner–Eckart theorem. Studies in History and Philosophy of Science Part A, 87(??):28– 43, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000078.

Hurwitz:2017:NCM

[Hur17] Brian Hurwitz. Narrative constructs in modern clinical case reporting. Studies in History and Philosophy of Science Part A, 62(??): 65–73, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300717.

Huss:2024:BFW

[Hus24] John E. Huss. Book forum: What is Regeneration? by Jane Maienschein and Kate MacCord: (prospects for unified regeneration). Studies in History and Philosophy of Science Part A, 104 (??):1–2, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000013.

Hutton:2004:EDC

[Hut04] Sarah Hutton. Émilie du Châtelet’s Institutions de physique as a document in the history of French Newtonianism. Studies in History and Philosophy of Science Part A, 35(3):515–531, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000536.

Hutton:2021:KCL

[Hut21] James Hutton. Kant, causation and laws of nature. Studies in History and Philosophy of Science Part A, 86(??):93– 102, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000030.

Holman:2022:NDP

[HW22] Bennett Holman and Torsten Wilholt. The new demarcation problem. Studies in History and Philosophy of Science Part A, 91 (??):211–220, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001886.

Houkes:2019:TTS

[HZ19] Wybo Houkes and Sjoerd D. Zwart. Transfer and templates in scientific modelling. Studies in History and Philosophy of Science Part A, 77(??):93–100, October 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302698.

Irzik:1998:WVK

[IG98] Gürol Irzik and Teo Grünberg. Whorfian variations on Kantian themes: Kuhn’s linguistic turn. Studies in History and Philosophy of Science Part A, 29(2):207–221, June 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819800003X.

Iliffe:2004:ACD

[Ili04] Rob Iliffe. Abstract considerations: disciplines and the incoherence of Newton’s natural philosophy. Studies in History and Philosophy of Science Part A, 35(3):427–454, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000524.

Illy:1981:RR

[Ill81] József Illy. Revolutions in a revolution. Studies in History and Philosophy of Science Part A, 12(3):175–210, September 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818190008X.

Iltis:1970:DVV

[Ilt70] Carolyn Iltis. D’Alembert and the vis viva controversy. Studies in History and Philosophy of Science Part A, 1(2):135–144, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368170900026.

Iltis:1977:MDC

[Ilt77] Carolyn Iltis. Madame du Châtelet’s metaphysics and mechanics. Studies in History and Philosophy of Science Part A, 8(1): 29–48, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900176.

Ivanova:2024:BEQ

[IRDV24] Milena Ivanova, Bridget Ritz, Marcela Duque, and Brandon Vaidyanathan. Beauty in experiment: a qualitative analysis of aesthetic experiences in scientific practice. Studies in History and Philosophy of Science Part A, 104(??):3–11, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000128.

Isaac:2014:MUP

[Isa14] Alistair M. C. Isaac. Model uncertainty and policy choice: a plea for integrated subjectivism. Studies in History and Philosophy of Science Part A, 47(??):42–50, September 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000545.

Isaac:2017:HHT

[Isa17] Alistair M. C. Isaac. Hubris to humility: Tonal volume and the fundamentality of psychophysical quantities. Studies in History and Philosophy of Science Part A, 65–66(??):99–111, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301589.

Isaac:2019:RTM

[Isa19] Alistair M. C. Isaac. Realism without tears I: Müller’s doctrine of specific nerve energies. Studies in History and Philosophy of Science Part A, 78(??):83–92, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811830133X.

Isaac:2020:RTI

[Isa20] Alistair M. C. Isaac. Realism without tears II: the structuralist legacy of sensory physiology. Studies in History and Philosophy of Science Part A, 79(??):15–29, February 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301201.

Ivanova:2010:PDG

[Iva10] Milena Ivanova. Pierre Duhem’s good sense as a guide to theory choice. Studies in History and Philosophy of Science Part A, 41 (1):58–64, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000958. See reply [Kid11b]

and response [Iva11].

Ivanova:2011:GSC

[Iva11] Milena Ivanova. ‘good sense’ in context: a response to Kidd. Studies in History and Philosophy of Science Part A, 42(4):610– 612, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000707. See [Iva10, Kid11b].

Ivanova:2015:CAW

[Iva15] Milena Ivanova. Conventionalism about what? Where Duhem and Poincaré part ways. Studies in History and Philosophy of Science Part A, 54(??):80–89, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001338.

Jackson:1994:AKE

[Jac94] Myles W. Jackson. Artisanal knowledge and experimental natural philosophers: The British response to Joseph Frauhofer and the Bavarian usurpation of their optical empire. Studies in History and Philosophy of Science Part A, 25(4):549–575, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900477.

Jacquette:1995:CAC

[Jac95] Dale Jacquette. Color and Armstrong’s color realism under the microscope. Studies in History and Philosophy of Science Part A, 26(3):389–406, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000069.

Jacquette:1997:MEM

[Jac97] Dale Jacquette. The microscope in early modern science and philosophy. Studies in History and Philosophy of Science Part A, 28 (2):377–386, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000210.

Jacobs:2002:PPI

[Jac02] Struan Jacobs. Polanyi’s presagement of the incommensurability concept. Studies in History and Philosophy of Science Part A, 33(1):105–116, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000310.

Jacovides:2003:LCI

[Jac03] Michael Jacovides. Locke’s construction of the idea of power. Studies in History and Philosophy of Science Part A, 34(2):329– 350, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000341.

Jackson:2016:WWW

[Jac16] Catherine M. Jackson. Who was William Hyde Wollaston? Studies in History and Philosophy of Science Part A, 58(??):50– 54, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811600008X.

Jacobs:2020:DCI

[Jac20] Caspar Jacobs. Du Châtelet: Idealist about extension, bodies and space. Studies in History and Philosophy of Science Part A, 82(??): 66–74, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301827.

Jacquart:2021:CMD

[Jac21] Melissa Jacquart. ΛCDM and MOND: a debate about models or theory? Studies in History and Philosophy of Science Part A, 89(??):226–234, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812100100X.

Jacobs:2023:MFB

[Jac23] Caspar Jacobs. The metaphysics of fibre bundles. Studies in History and Philosophy of Science Part A, 97(??):34–43, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001777.

Janis:1973:BRB

[Jan73] Allen I. Janis. Book review: The conceptual foundations of contemporary relativity theory: John Cowperthwaite Graves, Cambridge, Mass.: The MIT Press, 1971. xi + 361 pp. $15.00. Studies in History and Philosophy of Science Part A, 4(3):300–306, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900137. See [Gra71].

Janiak:2000:SAM

[Jan00] Andrew Janiak. Space, atoms and mathematical divisibility in Newton. Studies in History and Philosophy of Science Part A, 31(2):203–230, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000030.

Janiak:2013:TCC

[Jan13a] Andrew Janiak. Three concepts of causation in Newton. Studies in History and Philosophy of Science Part A, 44(3):396–407, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000842. See response [Hen14a].

Jansson:2013:NSE

[Jan13b] Lina Jansson. Newton’s “satis est”: a new explanatory role for laws. Studies in History and Philosophy of Science Part A, 44 (4):553–562, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000617.

Janiak:2015:SMN

[Jan15] Andrew Janiak. Space and motion in nature and scripture: Galileo, Descartes, Newton. Studies in History and Philosophy of Science Part A, 51(??):89–99, June 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000242.

Jardine:1976:GRT

[Jar76] N. Jardine. Galileo’s road to truth and the demonstrative regress. Studies in History and Philosophy of Science Part A, 7(4):277– 318, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900091.

Jardine:1979:FMR

[Jar79] Nicholas Jardine. The forging of modern realism: Clavius and Kepler against the sceptics. Studies in History and Philosophy of Science Part A, 10(2):141–173, June 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368179900141.

Jardine:1989:DF

[Jar89] N. Jardine. A dip into the future. Studies in History and Philosophy of Science Part A, 20(1):15–18, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900319.

Jardine:2003:EP

[Jar03a] Nick Jardine. Editorial preface. Studies in History and Philosophy of Science Part A, 34(1):1–4, March 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000821.

Jardine:2003:HSG

[Jar03b] Nick Jardine. Hermeneutic strategies in Gerd Buchdahl’s Kantian philosophy of science. Studies in History and Philosophy of Science Part A, 34(1):183–208, March 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000869.

Jardine:2009:BBB

[Jar09] Boris Jardine. Between the Beagle and the barnacle: Darwin’s microscopy, 1837–1854. Studies in History and Philosophy of Science Part A, 40(4):382–395, December 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000715.

Jardine:2013:RPR

[Jar13] Nicholas Jardine. Reflections on the preservation of recent scientific heritage in dispersed university collections. Studies in History and Philosophy of Science Part A, 44(4):735–743, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000514.

Jardine:2017:SFP

[Jar17] Boris Jardine. State of the field: Paper tools. Studies in History and Philosophy of Science Part A, 64(??):53–63, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301875.

Jasanoff:2004:BRB

[Jas04] Sheila Jasanoff. Book review: What inquiring minds should want to know: Science, truth and democracy, Philip Kitcher; Oxford University Press, Oxford, 2001, pp. 256, Price £22.50 hardback, ISBN 0-19-514583-6. Studies in History and Philosophy of Science Part A, 35(1):149–157, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000906. See [Kit01].

Jebeile:2016:EAM

[JB16] Julie Jebeile and Anouk Barberousse. Empirical agreement in model validation. Studies in History and Philosophy of Science Part A, 56(??):168–174, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001351.

Jebeile:2020:MME

[JC20] Julie Jebeile and Michel Crucifix. Multi-model ensembles in climate science: Mathematical structures and expert judgements. Studies in History and Philosophy of Science Part A, 83(??):44– 52, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301864.

Jebeile:2021:VMM

[JC21] Julie Jebeile and Michel Crucifix. Value management and model pluralism in climate science. Studies in History and Philosophy of Science Part A, 88(??):120–127, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812100087X.

Jenkins:2007:BIE

[Jen07] Lekelia D. Jenkins. Bycatch: interactional expertise, dolphins and the US tuna fishery. Studies in History and Philosophy of Science Part A, 38(4):698–712, December 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000635.

Jenkins:2010:ETZ

[Jen10] Lekelia D. Jenkins. The evolution of a trading zone: a case study of the turtle excluder device. Studies in History and Philosophy of Science Part A, 41(1):75–85, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000946.

Jesseph:1989:PTM

[Jes89] Douglas M. Jesseph. Philosophical theory and mathematical practice in the seventeenth century. Studies in History and Philosophy of Science Part A, 20(2):215–244, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900046.

Jesseph:1999:DFH

[Jes99] Douglas M. Jesseph. The decline and fall of Hobbesian geometry. Studies in History and Philosophy of Science Part A, 30(3): 425–453, September 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368199000114.

Jardine:1997:SMS

[JFS97] Nick Jardine and Marina Frasca-Spada. Splendours and miseries of the science wars. Studies in History and Philosophy of Science Part A, 28(2):219–235, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368197000022.

Jardine:2015:PPF

[JFS15] Nicholas Jardine and Marina Frasca-Spada. The pasts, presents, and futures of testimony. Studies in History and Philosophy of Science Part A, 52(??):95–100, August 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000448.

Jha:1997:NIM

[Jha97] S. R. Jha. A new interpretation of Michael Polanyi’s theory of tacit knowing: Integrative philosophy with ‘intellectual passions’. Studies in History and Philosophy of Science Part A, 28(4):611– 631, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000131.

Jhun:2022:BRL

[Jhu22] Jennifer S. Jhun. Book review: Leveraging distortions: explanation, idealization, and universality in science by Collin Rice: applications in economics. Studies in History and Philosophy of Science Part A, 95(??):226–227, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000863. See [Ode22,

Elg22, Pin22] and reply [Ric22].

Jacobs:2011:RBK

[JM11] Struan Jacobs and Phil Mullins. Relations between Karl Popper and Michael Polanyi. Studies in History and Philosophy of Science Part A, 42(3):426–435, September 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000185.

Johns:1998:SBM

[Joh98] Adrian Johns. Science and the book in modern cultural historiography. Studies in History and Philosophy of Science Part A, 29(2):167–194, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368198000041.

Johnson:2011:QRP

[Joh11] Kent Johnson. Quantitative realizations of philosophy of science: William Whewell and statistical methods. Studies in History and Philosophy of Science Part A, 42(3):399–409, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000203.

John:2012:MG

[Joh12a] Stephen John. Mind the gap. Studies in History and Philosophy of Science Part A, 43(1):218–220, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000884.

Johnston:2012:JDG

[Joh12b] Stephen Johnston. John Dee on geometry: Texts, teaching and the Euclidean tradition. Studies in History and Philosophy of Science Part A, 43(3):470–479, September 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001166.

John:2019:STD

[Joh19] Stephen John. Science, truth and dictatorship: Wishful thinking or wishful speaking? Studies in History and Philosophy of Science Part A, 78(??):64–72, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811830181X.

John:2021:SPR

[Joh21a] Stephen John. Science, politics and regulation: the trustbased approach to the demarcation problem. Studies in History and Philosophy of Science Part A, 90(??):1–9, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001229.

Johns:2021:EFM

[Joh21b] Oliver Davis Johns. Is electromagnetic field momentum due to the flow of field energy? Studies in History and Philosophy of Science Part A, 88(??):358–366, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000819.

John:2022:BRV

[Joh22] Stephen John. Book review: Vaccine Hesitancy: Public Trust, Expertise, and the War on Science by Maya Goldenberg: So, are the vaccines any good or not? Studies in History and Philosophy of Science Part A, 96(??):186–187, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001522. See [Chu22, WV22, Lea22] and

reply [Gol22].

Jones:1978:PTA

[Jon78] Gary E. Jones. Popper and theory appraisal. Studies in History and Philosophy of Science Part A, 9(3):239–249, September 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900079.

Jones:2001:WSB

[Jon01] Matthew L. Jones. Writing and sentiment: Blaise Pascal, the vacuum, and the Pensées. Studies in History and Philosophy of Science Part A, 32(1):139–181, March 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368100000248. See [Pas62, PT03, Pas10].

Jones:2013:HMU

[Jon13] Claire L. Jones. How to make a university history of science museum: Lessons from Leeds. Studies in History and Philosophy of Science Part A, 44(4):716–724, December 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113000526.

Joy:1997:NCN

[Joy97] Lynn S. Joy. Necessity, contingency, and the natural in modern science. Studies in History and Philosophy of Science Part A, 28(1):197–202, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000402.

Jacquart:2023:IRE

[JSZ23] Melissa Jacquart, Elay Shech, and Martin Zach. Idealization, representation, and explanation in the sciences. Studies in History and Philosophy of Science Part A, 99(??):A10–A14, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001352.

Jardine:2013:RMH

[JW13] Nicholas Jardine and Lydia Wilson. Recent material heritage of the sciences. Studies in History and Philosophy of Science Part A, 44(4):632–633, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000496.

Justus:2021:ATC

[JW21] James Justus and Samantha Wakil. The algorithmic turn in conservation biology: Characterizing progress in ethically-driven sciences. Studies in History and Philosophy of Science Part A, 88 (??):181–192, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000777.

Jackson:2022:WSM

[JW22] Rebecca L. Jackson and Merlin Wassermann. When standard measurement meets messy genitalia: Lessons from 20th century phallometry and cervimetry. Studies in History and Philosophy of Science Part A, 95(??):37–49, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000954.

Kaiser:1992:MRC

[Kai92] David Kaiser. More roots of complementarity: Kantian aspects and influences. Studies in History and Philosophy of Science Part A, 23(2):213–239, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368192900333.

Kaidesoja:2019:BMR

[Kai19] Tuukka Kaidesoja. Building middle-range theories from case studies. Studies in History and Philosophy of Science Part A, 78(??): 23–31, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117303023.

Kalikow:1975:HKL

[Kal75] Theo J. Kalikow. History of Konrad Lorenz’s ethological theory, 1927–1939: The role of meta-theory, theory, anomaly and new discoveries in a scientific ‘evolution’. Studies in History and Philosophy of Science Part A, 6(4):331–341, November 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900278.

Kallinen:1996:NPM

[Kal96] Maija Kallinen. Natural philosophy “Melanchthonized”, or how to create a Lutheran discipline? Studies in History and Philosophy of Science Part A, 27(3):381–386, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000526.

Kamminga:1994:BRB

[Kam94] Harmke Kamminga. Book review: Metchnikoff and the origins of immunology: From metaphor to theory: Alfred I. Tauber and Leon Chernyak Monographs on the History and Philosophy of Biology (Oxford: Oxford University Press, 1991), xviii + 247 pp. ISBN 0-19-506447-X Cloth £35.00. Studies in History and Philosophy of Science Part A, 25(1):131–145, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900248. See [TC91].

Kantorovich:1996:SRD

[Kan96] Aharon Kantorovich. Scientific realism: Darwinian smoke and platonic mirrors. Studies in History and Philosophy of Science Part A, 27(2):301–309, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000410.

Kaplan:2016:LTS

[Kap16] Judith Kaplan. Linguistic turns: Scientific Babel, the language of science, and the science of language. Studies in History and Philosophy of Science Part A, 60(??):88–91, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300395.

Kassell:2004:AHN

[Kas04] Lauren Kassell. An alchemist and his notebooks. Studies in History and Philosophy of Science Part A, 35(4):845–849, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000913.

Kastner:2014:EPO

[Kas14] Ruth E. Kastner. ‘Einselection’ of pointer observables: the new H-theorem? Studies in History and Philosophy of Science Part A, 48 (part A)(??):56–58, November 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981400063X.

Kastner:2018:IME

[Käs18] Lena Kästner. Integrating mechanistic explanations through epistemic perspectives. Studies in History and Philosophy of Science Part A, 68(??):68–79, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300256.

Katsumori:1992:TRE

[Kat92] Makoto Katsumori. The theories of relativity and Einstein’s philosophical turn. Studies in History and Philosophy of Science Part A, 23(4):557–592, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819290013V.

Katz:2020:SNS

[Kat20] Sophia Katz. Structure and numbers: Shao Yong on the order of reality. Studies in History and Philosophy of Science Part A, 81 (??):16–23, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301037.

Kaufmann:2008:PLU

[Kau08] Doris Kaufmann. ‘pushing the limits of understanding’: the discourse on primitivism in German Kulturwissenschaften, 1880–1930. Studies in History and Philosophy of Science Part A, 39(3):434– 443, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000666.

Kaveh:2024:RHU

[Kav24] Shahin Kaveh. A reinterpretation of Heisenberg’s Umdeutung in prescriptive-dynamical terms. Studies in History and Philosophy of Science Part A, 107(??):43–53, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001092.

Knusel:2020:UCP

[KB20] Benedikt Knüsel and Christoph Baumberger. Understanding climate phenomena with data-driven models. Studies in History and Philosophy of Science Part A, 84(??):46–56, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301606.

Kragh:1994:TAS

[KC94] Helge Kragh and Bruno Carazza. From time atoms to spacetime quantization: the idea of discrete time, ca. 1925–1936. Studies in History and Philosophy of Science Part A, 25(3):437– 462, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900612.

Kjeldsen:2012:GMK

[KC12] Tinne Hoff Kjeldsen and Jessica Carter. The growth of mathematical knowledge — introduction of convex bodies. Studies in History and Philosophy of Science Part A, 43(2):359– 365, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001427.

Knuuttila:2019:MGR

[KD19] Tarja Knuuttila and Vivette García Deister. Modelling gene regulation: (de)compositional and template-based strategies. Studies in History and Philosophy of Science Part A, 77(??):101– 111, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302686.

Kemp:2005:SSP

[Kem05] Stephen Kemp. Saving the Strong Programme? A critique of David Bloor’s recent work. Studies in History and Philosophy of Science Part A, 36(4):707–720, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368105000804. See [Blo99], critique [Feh07], and response [Kem07].

Kemp:2007:CAR

[Kem07] Stephen Kemp. Concepts, anomalies and reality: a response to Bloor and Fehér. Studies in History and Philosophy of Science Part A, 38(1):241–253, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106001002.

Kennedy:2012:NRV

[Ken12] Ashley Graham Kennedy. A non representationalist view of model explanation. Studies in History and Philosophy of Science Part A, 43(2):326–332, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001403.

Kendig:2023:BFP

[Ken23] Catherine Kendig. Book forum: Perspectival Realism by Michela Massimi: Finding realism in a plurality of situated scientific perspectives. Studies in History and Philosophy of Science Part A, 102(??):84–86, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812300136X.

Kepler:1968:AVP

[Kep68] Johannes Kepler. Ad Vitellionem paralipomena. Culture et Civilisation, Bruxelles, Belgium, 1968. ???? pp. LCCN QC381 .K39 1968.

Kerszberg:1987:RRE

[Ker87] Pierre Kerszberg. The relativity of rotation in the early foundations of general relativity. Studies in History and Philosophy of Science Part A, 18(1):53–79, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900112.

Kershaw:2007:IEU

[Ker07] Michael Kershaw. The international electrical units: a failure in standardisation? Studies in History and Philosophy of Science Part A, 38(1):108–131, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106001099.

Kershaw:2012:NPU

[Ker12] Michael Kershaw. The ‘nec plus ultra’ of precision measurement: Geodesy and the forgotten purpose of the Metre Convention. Studies in History and Philosophy of Science Part A, 43 (4):563–576, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000398.

Kessler:2007:RNS

[Kes07] Elizabeth A. Kessler. Resolving the nebulae: the science and art of representing M51. Studies in History and Philosophy of Science Part A, 38(2):477–491, June 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000210.

Koberinski:2023:RGM

[KF23] Adam Koberinski and Doreen Fraser. Renormalization group methods and the epistemology of effective field theories. Studies in History and Philosophy of Science Part A, 98(??):14– 28, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000031.

Khalifa:2013:UEK

[KG13] Kareem Khalifa and Michael Gadomski. Understanding as explanatory knowledge: the case of Bjorken scaling. Studies in History and Philosophy of Science Part A, 44(3):384–392, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000435.

Kuhn:2012:SSR

[KH12] Thomas S. Kuhn and Ian Hacking. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, fourth edition, 2012. ISBN 0-226-45811-3 (hardcover), 0-226-45812-1 (paperback), 0-226-45814-8 (e-book). xlvi + 217 pp. LCCN Q175 .K95 2012. With an introductory essay by Ian Hacking.

Kedar:2017:NLM

[KH17] Yael Kedar and Giora Hon. ‘Natures’ and ‘laws’: the making of the concept of law of nature — Robert Grosseteste (c.

1168–1253) and Roger Bacon (1214/1220–1292). Studies in History and Philosophy of Science Part A, 61(??):21–31, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116301145.

Kedar:2020:LON

[KH20] Yael Kedar and Giora Hon. Law and order natural regularities before the scientific revolution. Studies in History and Philosophy of Science Part A, 81(??):1–5, June 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120300571.

Keyser:2025:BSM

[KH25] Vadim Keyser and Hannah Howland. Bolstering superficial measurement robustness with community-based data foundations. Studies in History and Philosophy of Science Part A, 110(??): 19–29, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368125000019.

Khalidi:2024:OPS

[Kha24] Muhammad Ali Khalidi. Ontological pluralism and social values. Studies in History and Philosophy of Science Part A, 104(??): 61–67, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000219.

Khong:2003:AEH

[Kho03] Lynnette Khong. Actants and enframing: Heidegger and Latour on technology. Studies in History and Philosophy of Science Part A, 34(4):693–704, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000669.

Kidd:2011:OAI

[Kid11a] Ian James Kidd. Objectivity, abstraction, and the individual: The influence of Søren Kierkegaard on Paul Feyerabend. Studies in History and Philosophy of Science Part A, 42(1):125– 134, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110001056.

Kidd:2011:PDE

[Kid11b] Ian James Kidd. Pierre Duhem’s epistemic aims and the intellectual virtue of humility: a reply to Ivanova. Studies in History and Philosophy of Science Part A, 42(1):185–189, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000865. See [Iva10, Iva11].

Kidd:2016:FPE

[Kid16a] Ian James Kidd. Feyerabend on politics, education, and scientific culture. Studies in History and Philosophy of Science Part A, 57(??):121–128, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001697.

Kidd:2016:ICE

[Kid16b] Ian James Kidd. Inevitability, contingency, and epistemic humility. Studies in History and Philosophy of Science Part A, 55(??):12– 19, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000990.

Kidd:2017:OHO

[Kid17] Ian James Kidd. Other histories, other sciences. Studies in History and Philosophy of Science Part A, 61(??):57–60, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300031.

Kilic:1999:JVE

[Kil99] Berna Eden Kiliç. John Venn’s evolutionary logic of chance. Studies in History and Philosophy of Science Part A, 30(4):559– 585, December 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368199000266.

Kim:1995:LMW

[Kim95] Mi Gyung Kim. Labor and mirage: Writing the history of chemistry. Studies in History and Philosophy of Science Part A, 26(1): 155–165, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819400030D.

Kim:2000:CAD

[Kim00] Mi Gyung Kim. Chemical analysis and the domains of reality: Wilhelm Homberg’s Essais de chimie, 1702–1709. Studies in History and Philosophy of Science Part A, 31(1):37–69, March 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368199000333. See [Hom02].

Kim:2012:SKT

[Kim12] Yung Sik Kim. Scholars, knowledge, and techniques in traditional China. Studies in History and Philosophy of Science Part A, 43(1): 228–231, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001233.

Kim:2024:DML

[Kim24] Jun Young Kim. Divine mathematics: Leibniz’s combinatorial theory of compossibility. Studies in History and Philosophy of Science Part A, 106(??):60–69, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000438.

King:1987:JBP

[Kin87] Peter King. Jean Buridan’s philosophy of science. Studies in History and Philosophy of Science Part A, 18(2):109–132, June 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818790015X.

Kinzel:2015:NEH

[Kin15a] Katherina Kinzel. Narrative and evidence. how can case studies from the history of science support claims in the philosophy of science? Studies in History and Philosophy of Science Part A, 49(??):48–57, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114001149.

Kinzel:2015:SFR

[Kin15b] Katherina Kinzel. State of the field: Are the results of science contingent or inevitable? Studies in History and Philosophy of Science Part A, 52(??):55–66, August 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000734.

Kinney:2021:CPC

[Kin21] David Kinney. Curie’s principle and causal graphs. Studies in History and Philosophy of Science Part A, 87(??):22–27, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000224.

Kirschenmann:1973:BRB

[Kir73] Peter Kirschenmann. Book review: Symmetries and reflections: Scientific essays: E. P. Wigner Cambridge, Mass.: M.I.T. Press, 1970. Studies in History and Philosophy of Science Part A, 4(2): 193–207, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900046. See [Wig67, Wig70,

Wig78].

Kirkham:2009:BNA

[Kir09] Georgiana Kirkham. Is biotechnology the new alchemy? Studies in History and Philosophy of Science Part A, 40(1):70–80, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108001131.

Kitcher:1975:BIA

[Kit75] Philip Kitcher. Bolzano’s ideal of algebraic analysis. Studies in History and Philosophy of Science Part A, 6(3):229–269, October 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900242.

Kitchener:1987:GEE

[Kit87] Richard F. Kitchener. Genetic epistemology, equilibration and the rationality of scientific change. Studies in History and Philosophy of Science Part A, 18(3):339–366, September 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900240.

Kitcher:1993:ASS

[Kit93] Philip Kitcher. The advancement of science: science without legend, objectivity without illusions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-504628-5. viii + 421

pp. LCCN Q175 .K533 1993. URL http://www.loc.gov/catdir/
enhancements/fy0637/92019532-d.html.

Kitcher:2001:STD

[Kit01] Philip Kitcher. Science, truth, and democracy. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514583-6. xiii + 219 pp.

LCCN Q175.5 .K525 2002. URL http://www.loc.gov/catdir/
enhancements/fy0639/2001036144-d.html.

Kusch:2002:TP

[KL02] Martin Kusch and Peter Lipton. Testimony: a primer. Studies in History and Philosophy of Science Part A, 33(2):209– 217, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000031.

Knuuttila:2014:VNA

[KL14] Tarja Knuuttila and Andrea Loettgers. Varieties of noise: Analogical reasoning in synthetic biology. Studies in History and Philosophy of Science Part A, 48(??):76–88, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000612.

Krashniak:2021:FGR

[KL21] Adam Krashniak and Ehud Lamm. Francis Galton’s regression towards mediocrity and the stability of types. Studies in History and Philosophy of Science Part A, 86(??):6–19, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120302090.

Klagge:2000:DHS

[Kla00] James C. Klagge. The difficulty here is: to stop. Studies in History and Philosophy of Science Part A, 31(3):551–557, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000182.

Kleiner:1977:RDS

[Kle77] Scott A. Kleiner. Referential divergence in scientific theories. Studies in History and Philosophy of Science Part A, 8(2):87– 109, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900103.

Kleiner:1979:FGD

[Kle79] Scott A. Kleiner. Feyerabend, Galileo and Darwin: How to make the best out of what you have — or think you can get. Studies in History and Philosophy of Science Part A, 10(4):285–309, December 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900177.

Kleiner:1983:NLK

[Kle83] Scott A. Kleiner. A new look at Kepler and abductive argument. Studies in History and Philosophy of Science Part A, 14(4):279– 313, December 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368183900092.

Klein:2001:PTE

[Kle01] Ursula Klein. Paper tools in experimental cultures. Studies in History and Philosophy of Science Part A, 32(2):265–302, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000103.

Klein:2004:WKN

[Kle04] Ursula Klein. Ways of knowing. A new history of science, technology and medicine. Studies in History and Philosophy of Science Part A, 35(1):159–172, March 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810300089X.

Klein:2005:SOC

[Kle05] Ursula Klein. Shifting ontologies, changing classifications: plant materials from 1700 to 1830. Studies in History and Philosophy of Science Part A, 36(2):261–329, June 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000178.

Kleingeld:2008:KHU

[Kle08] Pauline Kleingeld. Kant on historiography and the use of regulative ideas. Studies in History and Philosophy of Science Part A, 39 (4):523–528, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000861.

Kletzl:2014:STK

[Kle14] Sebastian Kletzl. Scrutinizing thing knowledge. Studies in History and Philosophy of Science Part A, 47(??):118–123, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000569.

Klein:2015:RNH

[Kle15] Ursula Klein. A Revolution that never happened. Studies in History and Philosophy of Science Part A, 49(??):80–90, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114001113.

Klev:2016:CUS

[Kle16] Ansten Klev. Carnap on unified science. Studies in History and Philosophy of Science Part A, 59(??):53–67, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300206.

Kroes:2006:DNT

[KM06] Peter Kroes and Anthonie Meijers. The dual nature of technical artefacts. Studies in History and Philosophy of Science Part A, 37(1):1–4, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105001032.

Kuorikoski:2014:UMD

[KM14] Jaakko Kuorikoski and Caterina Marchionni. Unification and mechanistic detail as drivers of model construction: Models of networks in economics and sociology. Studies in History and Philosophy of Science Part A, 48(??):97–104, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000600.

Kirk:2022:GEC

[KM22] Robert G. W. Kirk and Dmitriy Myelnikov. Governance, expertise, and the ‘culture of care’: the changing constitutions of laboratory animal research in Britain, 1876–2000. Studies in History and Philosophy of Science Part A, 93(??):107–122, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000553.

Krafft:1973:IKS

[KMS73] F. Krafft, K. Meyer, and B. Sticker, editors. Internationales Kepler-Symposium Weil der Stadt 1971. Referate und diskussionen, volume 1 of Arbor scientiarum Reihe A. H. A. Gerstenberg, Hildesheim, West Germany, 1973. ISBN 3-8067-0155-5. LCCN QB36.K4.

Kantorovich:1989:SSE

[KN89] Aharon Kantorovich and Yuval Ne’eman. Serendipity as a source of evolutionary progress in science. Studies in History and Philosophy of Science Part A, 20(4):505–529, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900216.

Knight:1971:BRP

[Kni71] David M. Knight. Book review: Popularizing the history of chemistry: C. J. Schneer, Mind and matter. Grove Press: New York, 1970. xiv + 305 pp. $8.50. Studies in History and Philosophy of Science Part A, 1(4):363–368, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900136.

See [Sch70, Sch88d].

Knight:2005:SS

[Kni05] David Knight. Snippets of science. Studies in History and Philosophy of Science Part A, 36(3):618–625, September 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000555.

Knorr:1976:PIG

[Kno76] W. Knorr. Problems in the interpretation of Greek number theory: Euclid and the ‘Fundamental Theorem of Arithmetic’. Studies in History and Philosophy of Science Part A, 7(4):353– 368, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900121.

Knobe:2009:FJC

[Kno09] Joshua Knobe. Folk judgments of causation. Studies in History and Philosophy of Science Part A, 40(2):238–242, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000302.

Knuuttila:2011:MRA

[Knu11] Tarja Knuuttila. Modelling and representing: An artefactual approach to model-based representation. Studies in History and Philosophy of Science Part A, 42(2):262–271, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001135.

Koberinski:2024:PTB

[Kob24] Adam Koberinski. Phase transitions and the birth of early universe particle physics. Studies in History and Philosophy of Science Part A, 105(??):59–73, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000499.

Kochan:2004:TDD

[Koc04] Jeff Kochan. Technological democracy or democratic technology? Studies in History and Philosophy of Science Part A, 35(2):401– 412, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810400024X.

Kochan:2006:FSC

[Koc06] Jeff Kochan. Feenberg and STS: counter-reflections on bridging the gap. Studies in History and Philosophy of Science Part A, 37 (4):702–720, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000537. See reply [Fee06].

Kochiras:2009:GNS

[Koc09] Hylarie Kochiras. Gravity and Newton’s Substance Counting Problem. Studies in History and Philosophy of Science Part A, 40(3): 267–280, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810900051X. See discussion

[Sch11].

Kochan:2011:HPS

[Koc11a] Jeff Kochan. Husserl and the phenomenology of science. Studies in History and Philosophy of Science Part A, 42(3):467–471, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000173.

Kochiras:2011:GCS

[Koc11b] Hylarie Kochiras. Gravity’s cause and substance counting: contextualizing the problems. Studies in History and Philosophy of Science Part A, 42(1):167–184, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000841.

Kochan:2013:SES

[Koc13] Jeff Kochan. Subjectivity and emotion in scientific research. Studies in History and Philosophy of Science Part A, 44(3):354–362, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000381.

Kochan:2015:OSN

[Koc15a] Jeff Kochan. Objective styles in northern field science. Studies in History and Philosophy of Science Part A, 52(??):1–12, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000424.

Kochan:2015:PSC

[Koc15b] Jeff Kochan. Putting a spin on circulating reference, or how to rediscover the scientific subject. Studies in History and Philosophy of Science Part A, 49(??):103–107, February 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000946.

Kochan:2024:ASE

[Koc24] Jeff Kochan. Animism and science in European perspective. Studies in History and Philosophy of Science Part A, 103(??):46– 57, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001589.

Koetsier:1995:EHM

[Koe95] Teun Koetsier. Explanation in the historiography of mathematics: The case of Hamilton’s quaternions. Studies in History and Philosophy of Science Part A, 26(4):593–616, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000267.

Koi:2021:GNS

[Koi21] Polaris Koi. Genetics on the neurodiversity spectrum: Genetic, phenotypic and endophenotypic continua in autism and ADHD. Studies in History and Philosophy of Science Part A, 89(??):52– 62, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001072.

Kong:2022:TNQ

[Kon22] Otto C. W. Kong. Towards noncommutative quantum reality. Studies in History and Philosophy of Science Part A, 92(??):186– 195, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000334.

Konashev:2024:OHB

[Kon24a] Mikhail B. Konashev. “Oh, how beautiful life is and how terrible death is!” (Th. Dobzhansky and religion). Studies in History and Philosophy of Science Part A, 107(??):25–32, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001109.

Konsman:2024:ENM

[Kon24b] Jan Pieter Konsman. Expanding the notion of mechanism to further understanding of biopsychosocial disorders? Depression and medically-unexplained pain as cases in point. Studies in History and Philosophy of Science Part A, 103(??):123–136, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001693.

Kosso:1988:SHU

[Kos88] Peter Kosso. Spacetime horizons and unobservability. Studies in History and Philosophy of Science Part A, 19(2):161– 173, June 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900258.

Kosso:1993:MRT

[Kos93] Peter Kosso. Middle-range theory in historical archaeology. Studies in History and Philosophy of Science Part A, 24(2):163– 184, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390044K.

Kosso:1999:SAP

[Kos99] Peter Kosso. Symmetry arguments in physics. Studies in History and Philosophy of Science Part A, 30(3):479–492, September 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368199000126.

Kosso:2006:DEP

[Kos06] Peter Kosso. Detecting extrasolar planets. Studies in History and Philosophy of Science Part A, 37(2):224–236, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000227.

Kovaka:2021:ECS

[Kov21] Karen Kovaka. Evaluating community science. Studies in History and Philosophy of Science Part A, 88(??):102–109, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000686.

Kowalenko:2011:EHL

[Kow11] Robert Kowalenko. The epistemology of hedged laws. Studies in History and Philosophy of Science Part A, 42(3):445–452, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000240.

Koznjak:2015:WLD

[Koz15] Boris Koznjak. Who let the demon out? Laplace and Boscovich on determinism. Studies in History and Philosophy of Science Part A, 51(??):42–52, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000370.

Kepler:2008:SOA

[KR08] Johannes Kepler and Rolf Riekher. Schriften zur Optik 1604– 1611: Ad Vitellionem paralipomena quibus astronomiae pars optica; Dissertatio cum Nuncio sidereo; Dioptrice, volume 198 of Ostwalds Klassiker der exakten Wissenschaften. Verlag Harri Deutsch, Frankfurt, Germany, 2008. ISBN 3-8171-3198-4 (paperback). ISSN 0232-3419. 588 pp. LCCN QC353 .K385 2008.

Koskinen:2022:DBL

[KR22] Inkeri Koskinen and Kristina Rolin. Distinguishing between legitimate and illegitimate roles for values in transdisciplinary research. Studies in History and Philosophy of Science Part A, 91 (??):191–198, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121002016.

Krausser:1972:OCR

[Kra72] Peter Krausser. The operational conception of ‘Reine anschauung’ (pure intuition) in Kant’s theory of experience and science. Studies in History and Philosophy of Science Part A, 3(1): 81–87, May 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900155.

Krausser:1973:FIF

[Kra73] Peter Krausser. ‘form of intuition’ and ‘formal intuition’ in Kant’s theory of experience and science. Studies in History and Philosophy of Science Part A, 4(3):279–287, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368173900113.

Kragh:1981:CMH

[Kra81] Helge Kragh. The concept of the monopole. A historical and analytical case-study. Studies in History and Philosophy of Science Part A, 12(2):141–172, June 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368181900170.

Kraus:2011:KSS

[Kra11] Katharina T. Kraus. Kant and the ‘soft sciences’. Studies in History and Philosophy of Science Part A, 42(4):618–624, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000756.

Kraus:2018:SGI

[Kra18] Katharina T. Kraus. The soul as the ‘guiding idea’ of psychology: Kant on scientific psychology, systematicity, and the idea of the soul. Studies in History and Philosophy of Science Part A, 71(??):77–88, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117300225.

Krakauer:2022:BRR

[Kra22a] John W. Krakauer. Book review: Representation in Cognitive Science by Nicholas Shea: But is it thinking? The philosophy of representation meets systems neuroscience. Studies in History and Philosophy of Science Part A, 92(??):267– 269, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000789.

Krasnodebski:2022:RWC

[Kra22b] Marcin Krasnodebski. Reinventing the wheel: a critical look at one-world and circular chemistries. Studies in History and Philosophy of Science Part A, 96(??):112–120, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001364.

Krasnodebski:2024:BRS

[Kra24] Marcin Krasnodebski. The bumpy road to sustainability: Reassessing the history of the twelve principles of green chemistry. Studies in History and Philosophy of Science Part A, 103(??):85– 94, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812300167X.

Krips:1980:AIN

[Kri80] H. Krips. Aristotle on the infallibility of normal observation. Studies in History and Philosophy of Science Part A, 11(1):79– 86, March 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900059.

Krips:1982:EHD

[Kri82] H. Krips. Epistemological holism: Duhem or quine? Studies in History and Philosophy of Science Part A, 13(3):251–264, September 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368182900103.

Krips:1986:APP

[Kri86] H. Krips. Atomism, Poincaré and Planck. Studies in History and Philosophy of Science Part A, 17(1):43–63, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818690018X.

Krickel:2018:SMM

[Kri18] Beate Krickel. Saving the mutual manipulability account of constitutive relevance. Studies in History and Philosophy of Science Part A, 68(??):58–67, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300074.

Kroes:2006:CSF

[Kro06] Peter Kroes. Coherence of structural and functional descriptions of technical artefacts. Studies in History and Philosophy of Science Part A, 37(1):137–151, March 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105001184.

Kruger:1980:IRT

[Krü80] Lorenz Krüger. Intertheoretic relations as a tool for the rational reconstruction of scientific development. Studies in History and Philosophy of Science Part A, 11(2):89–101, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368180900163.

Kashyap:2019:DQP

[KS19] Abhishek Kashyap and Vikram S. Sirola. The Duhem–Quine problem for equiprobable conjuncts. Studies in History and Philosophy of Science Part A, 75(??):43–50, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368118300657.

Kramer:2024:EHE

[KSZ24] Michael Krämer, Gregor Schiemann, and Christian Zeitnitz. Experimental high-energy physics without computer simulations. Studies in History and Philosophy of Science Part A, 106(??):37– 42, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000426.

Kaplan:2021:GQP

[KT21] Jonathan Michael Kaplan and Eric Turkheimer. Galton’s quincunx: Probabilistic causation in developmental behavior genetics. Studies in History and Philosophy of Science Part A, 88(??):60– 69, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000455.

Kuby:2016:FCI

[Kub16] Daniel Kuby. Feyerabend’s ‘The concept of intelligibility in modern physics’ (1948). Studies in History and Philosophy of Science Part A, 57(??):57–63, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001648.

Kuhn:1962:SSR

[Kuh62] Thomas S. Kuhn. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, 1962. xv + 172 pp. LCCN Q175 .K95.

Kuhn:1970:SSR

[Kuh70] Thomas S. Kuhn. The Structure of Scientific Revolutions. International encyclopedia of unified science. Foundations of the unity of science, v. 2, no. 2. University of Chicago Press, Chicago, IL, USA and London, UK, second enlarged edition, 1970. ISBN 0-226-45803-2 (hardcover), 0-226-45804-0 (paperback), 0-226-45807-5. xii + 210 pp. LCCN Q175 .K95 1970.

Kuhn:1996:SSR

[Kuh96] Thomas S. Kuhn. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA and London, UK, third edition, 1996. ISBN 0-226-45807-5 (hardcover), 0-226-45808-3 (paperback). xiv + 212 pp. LCCN Q175 .K95 1996. URL http://

www.loc.gov/catdir/description/uchi051/96013195.html;
http://www.loc.gov/catdir/toc/uchi051/96013195.html.

Kuipers:2014:EPN

[Kui14] Theo A. F. Kuipers. Empirical progress and nomic truth approximation revisited. Studies in History and Philosophy of Science Part A, 46(??):64–72, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000156.

Kukla:1995:TAB

[Kuk95] Andre Kukla. The two antirealisms of Bas van Fraassen. Studies in History and Philosophy of Science Part A, 26(3):431–454, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819500010K.

Kukla:2001:SPP

[Kuk01] André Kukla. SETI: On the prospects and pursuitworthiness of the search for extraterrestrial intelligence. Studies in History and Philosophy of Science Part A, 32(1):31–67, March 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810000025X.

Kukla:2022:PPH

[Kuk22] Quill R Kukla. Book review: Philosophy of population health: Philosophy for a new public health era by Sean Valles: Healthism and the weaponization of “health”. Studies in History and Philosophy of Science Part A, 91(??):316–319, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121002041.

Kusch:1997:SFP

[Kus97] Martin Kusch. The sociophilosophy of folk psychology. Studies in History and Philosophy of Science Part A, 28(1):1–25, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936819600012X.

Kusch:2001:GTS

[Kus01] Martin Kusch. ‘A general theory of societal knowledge’?: Aspirations and shortcomings of Alvin Goldman’s social epistemology. Studies in History and Philosophy of Science Part A, 32(1):183– 192, March 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000315.

Kusch:2002:SCW

[Kus02a] Martin Kusch. The social construction of what? Studies in History and Philosophy of Science Part A, 33(3):639–647, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000371.

Kusch:2002:TCE

[Kus02b] Martin Kusch. Testimony in communitarian epistemology. Studies in History and Philosophy of Science Part A, 33(2):335– 354, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000080.

Kusch:2010:HHE

[Kus10] Martin Kusch. Hacking’s historical epistemology: a critique of styles of reasoning. Studies in History and Philosophy of Science Part A, 41(2):158–173, June 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000191.

Kusch:2015:SPC

[Kus15] Martin Kusch. Scientific pluralism and the Chemical Revolution. Studies in History and Philosophy of Science Part A, 49(??):69– 79, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811400082X.

Kusch:2016:RFL

[Kus16] Martin Kusch. Relativism in Feyerabend’s later writings. Studies in History and Philosophy of Science Part A, 57(??):106– 113, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001703.

Kutrovatz:2022:AIS

[Kut22] Gábor Kutrovátz. Anatomical identifications of stars: Textual descriptions in Ptolemy’s star catalogue. Studies in History and Philosophy of Science Part A, 91(??):94–102, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001898.

Kuukkanen:2007:KCT

[Kuu07] Jouni-Matti Kuukkanen. Kuhn, the correspondence theory of truth and coherentist epistemology. Studies in History and Philosophy of Science Part A, 38(3):555–566, September 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810700043X.

Kuukkanen:2009:CDC

[Kuu09] Jouni-Matti Kuukkanen. Closing the door to cloud-cuckoo land: a reply to Seselja and Straßer. Studies in History and Philosophy of Science Part A, 40(3):328–331, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000454. See [SS09].

Kuukkanen:2011:KGM

[Kuu11] Jouni-Matti Kuukkanen. I am knowledge. Get me out of here! On localism and the universality of science. Studies in History and Philosophy of Science Part A, 42(4):590–601, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000690.

Kuukkanen:2016:HFH

[Kuu16] Jouni-Matti Kuukkanen. Historicism and the failure of HPS. Studies in History and Philosophy of Science Part A, 55(??):3– 11, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000953.

Kuukkanen:2017:ALR

[Kuu17] Jouni-Matti Kuukkanen. An aging literary revolution: Stuck with the paradigm. Studies in History and Philosophy of Science Part A, 64(??):67–70, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301954.

Kvasz:2014:KSS

[Kva14] Ladislav Kvasz. Kuhn’s Structure of Scientific Revolutions between sociology and epistemology. Studies in History and Philosophy of Science Part A, 46(??):78–84, June 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000181.

Koliofotis:2022:HRN

[KVJ22] Vaios Koliofotis and Philippe Verreault-Julien. Hamilton’s rule: a non-causal explanation? Studies in History and Philosophy of Science Part A, 92(??):109–118, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001850.

Kwa:2012:SVS

[Kwa12] Chunglin Kwa. An ‘ecological’ view of styles of science and of art: Alois Riegl’s explorations of the style concept. Studies in History and Philosophy of Science Part A, 43(4):610–618, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000453.

Lu-Adler:2022:KUT

[LA22] Huaping Lu-Adler. Kant’s use of travel reports in theorizing about race — a case study of how testimony features in natural philosophy. Studies in History and Philosophy of Science Part A, 91 (??):10–19, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812100176X.

Lacey:2011:IP

[Lac11] Hugh Lacey. Integrative pluralism. Studies in History and Philosophy of Science Part A, 42(1):219–222, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000907.

Lacey:2015:HEC

[Lac15] Hugh Lacey. ‘Holding’ and ‘endorsing’ claims in the course of scientific activities. Studies in History and Philosophy of Science Part A, 53(??):89–95, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000692.

Lacey:2021:MSA

[Lac21] Hugh Lacey. The methodological strategies of agroecological research and the values with which they are linked. Studies in History and Philosophy of Science Part A, 88(??):292–302, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000972.

Ladyman:1998:WSR

[Lad98] James Ladyman. What is structural realism? Studies in History and Philosophy of Science Part A, 29(3):409–424, September 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198801295.

Ladyman:2019:ISW

[Lad19] James Ladyman. Introduction: Structuralists of the world unite. Studies in History and Philosophy of Science Part A, 74(??): 1–3, ???? 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118303315.

Lakatos:1974:RCE

[Lak74] Imre Lakatos. The role of crucial experiments in science. Studies in History and Philosophy of Science Part A, 4(4):309–325, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368174900077.

Goff:2021:HCE

[LAL21] Anne Le Goff, Patrick Allard, and Hannah Landecker. Heritable changeability: Epimutation and the legacy of negative definition in epigenetic concepts. Studies in History and Philosophy of Science Part A, 86(??):35–46, April 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120302077.

Lange:2006:FLN

[Lan06] Marc Lange. Farewell to laws of nature? Studies in History and Philosophy of Science Part A, 37(2):361–369, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810600032X.

Landig:2014:PRS

[Lan14] Anders Landig. Partial reference, scientific realism and possible worlds. Studies in History and Philosophy of Science Part A, 47 (??):1–9, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000247.

Lange:2018:RCP

[Lan18] Marc Lange. A reply to Craver and Povich on the directionality of distinctively mathematical explanations. Studies in History and Philosophy of Science Part A, 67(??):85–88, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117302418. See [CP17].

Langermann:2020:MMJ

[Lan20] Y. Tzvi Langermann. Moses Maimonides and Judah Halevi on order and law in the world of nature, and beyond. Studies in History and Philosophy of Science Part A, 81(??):39–45, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300712.

Lange:2021:WCM

[Lan21] Marc Lange. What could mathematics be for it to function in distinctively mathematical scientific explanations? Studies in History and Philosophy of Science Part A, 87(??):44–53, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000157.

Larvor:2003:WDK

[Lar03] Brendan Larvor. Why did Kuhn’s Structure of Scientific Revolutions cause a fuss? Studies in History and Philosophy of Science Part A, 34(2):369–390, June 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368103000232. See [Kuh62, Kuh70, Kuh96, KH12].

Lari:2024:WCR

[Lar24] Teemu Lari. What counts as relevant criticism? Longino’s critical contextual empiricism and the feminist criticism of mainstream economics. Studies in History and Philosophy of Science Part A, 104(??):88–97, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000141.

Latour:1988:PF

[Lat88] Bruno Latour. The pasteurization of France. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-65760-8 (hardcover), 0-674-65761-6 (paperback). 273 pp. LCCN QR22.F8 L3813 1988. Translated by Alan Sheridan and John Law.

Latour:1990:PNS

[Lat90] Bruno Latour. Postmodern? no, simply amodern! steps towards an anthropology of science. Studies in History and Philosophy of Science Part A, 21(1):145–171, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900184.

Latour:1993:WNB

[Lat93] Bruno Latour. We have never been modern. Harvard University Press, Cambridge, MA, USA, 1993. ISBN 0-674-94838-6, 0-674-94839-4 (paperback). ix + 157 pp. LCCN Q175.5 .L3513 1993.

Laudan:1982:PTC

[Lau82a] Larry Laudan. Problems, truth, and consistency. Studies in History and Philosophy of Science Part A, 13(1):73–80, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818290005X.

Laudan:1982:RML

[Lau82b] Rachel Laudan. The role of methodology in Lyell’s science. Studies in History and Philosophy of Science Part A, 13(3):215–249, September 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368182900097.

Lauxtermann:1987:FDY

[Lau87] P. F. H. Lauxtermann. Five decisive years: Schopenhauer’s epistemology as reflected in his theory of colour: Introduction: Schopenhauer as an enlightened romantic. Studies in History and Philosophy of Science Part A, 18(3):271–291, September 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368187900227.

Laudan:1988:CPR

[Lau88] Larry Laudan. Conceptual problems re-visited. Studies in History and Philosophy of Science Part A, 19(4):531–534, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900143.

Laudan:1989:THY

[Lau89] Larry Laudan. Thoughts on HPS: 20 years later. Studies in History and Philosophy of Science Part A, 20(1):9–13, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900307.

Laudan:1990:ALE

[Lau90]  Larry Laudan. Aim-less epistemology?  Studies in His-

tory and Philosophy of Science Part A, 21(2):315–322, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900298.

Laudisa:2023:HWD

[Lau23] Federico Laudisa. How and when did locality become ‘local realism’? A historical and critical analysis (1963–1978). Studies in History and Philosophy of Science Part A, 97(??):44–57, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001753.

Laymon:1978:NEC

[Lay78] Ronald Laymon. Newton’s experimentum crucis and the logic of idealization and theory refutation. Studies in History and Philosophy of Science Part A, 9(1):51–77, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900213.

Lopez-Beltran:1994:FHM

[LB94] Carlos López-Beltrán. Forging heredity: From metaphor to cause, a reification story. Studies in History and Philosophy of Science Part A, 25(2):211–235, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900280.

Lu:2022:CIH

[LB22] Qiaoying Lu and Pierrick Bourrat. On the causal interpretation of heritability from a structural causal modeling perspective. Studies in History and Philosophy of Science Part A, 94(??):87– 98, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000796.

Chen:2000:TVI

[lC00] Ruey lin Chen. Theory versions instead of articulations of a paradigm. Studies in History and Philosophy of Science Part A, 31 (3):449–471, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000121.

Leconte-Chevillard:2021:ECL

[LC21] Gauvain Leconte-Chevillard. Experimentation in the cosmic laboratory. Studies in History and Philosophy of Science Part A, 90 (??):265–274, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001618.

Leconte-Chevillard:2024:ECI

[LC24a] Gauvain Leconte-Chevillard. Experimentation in cosmology: Intervening on the whole universe. Studies in History and Philosophy of Science Part A, 106(??):136–145, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001006.

Levy:2024:BTE

[LC24b] Arnon Levy and Adrian Currie. Bringing thought experiments back into the philosophy of science. Studies in History and Philosophy of Science Part A, 105(??):149–157, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000554.

LeRonddAlembert:1758:TDL

[Le 58] Jean Le Rond d’Alembert. Traité de dynamique, dans lequel les loix de l’équilibre and du mouvement des corps sont réduits au plus petit nombre possible, and démontrées d’une maniere nouvelle, and où l’on donne un principe général pour trouver le mouvement de plusieurs corps qui agissent les uns sur les autres d’une maniere quelconque. (French) [Treatise on dynamics, in which the laws of balance and the movement of the body are reduced to the smallest number possible, and demonstrated in a new way, and where we give a general principle to find the motion of several bodies which act on each other in any way]. David, Paris, 1758. xl + 272 + xiii pp. LCCN QA804 .A35.

LeRonddAlembert:1968:TD

[Le 68] Jean Le Rond d’Alembert. Traité de dynamique, volume 72 of The sources of science. Johnson Reprint Corp., New York, NY, USA, 1968. xxxvi + 272 + xiii pp. LCCN QA804 .A35 1968. With a new introductory and bibliography by Thomas L. Hankins.

Leavitt:1991:KSH

[Lea91] Frank J. Leavitt. Kant’s schematism and his philosophy of geometry. Studies in History and Philosophy of Science Part A, 22 (4):647–659, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819190037S. See comments

[Gri91] and reply [Ris91].

Leask:1998:FFW

[Lea98] Nigel Leask. Fire or flood? Wordsworth and romantic geology. Studies in History and Philosophy of Science Part A, 29(1):117– 127, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000325.

Leach:2022:BRV

[Lea22] Joan Leach. Book review: Vaccine Hesitancy: Public Trust, Expertise, and the War on Science by Maya Goldenberg: a Pox on all our Houses. Studies in History and Philosophy of Science Part A, 96 (??):193–195, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001510. See [Joh22,

Chu22, WV22] and reply [Gol22].

Lee:2013:ATW

[Lee13] Wang-Yen Lee. Akaike’s Theorem and weak predictivism in science. Studies in History and Philosophy of Science Part A, 44 (4):594–599, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811300040X.

Lee:2019:NST

[Lee19] Clarissa Ai Ling Lee. Nuclear science and technology in the Malaysian context: Three phases of technoscientific knowledge transfer (ETTLG). Studies in History and Philosophy of Science Part A, 77(??):130–140, October 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302431.

Lehoux:2004:OPA

[Leh04a] Daryn Lehoux. Observation and prediction in ancient astrology. Studies in History and Philosophy of Science Part A, 35(2):227– 246, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000955.

Lehoux:2004:WWW

[Leh04b] Daryn Lehoux. Weather, when and why? Studies in History and Philosophy of Science Part A, 35(4):835–843, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000202.

Lehoux:2006:LNN

[Leh06] Daryn Lehoux. Laws of nature and natural laws. Studies in History and Philosophy of Science Part A, 37(4):527–549, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106000744.

Lehmann:2018:ARP

[Leh18] Philipp Lehmann. Average rainfall and the play of colors: Colonial experience and global climate data. Studies in History and Philosophy of Science Part A, 70(??):28–37, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301390.

Lehoux:2020:SPL

[Leh20] Daryn Lehoux. Saved by the phenomena: Law and nature in Cicero and the (pseudo?) Platonic Epinomis. Studies in History and Philosophy of Science Part A, 81(??):55–61, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301031.

Lemeire:2021:CSN

[Lem21] Olivier Lemeire. The causal structure of natural kinds. Studies in History and Philosophy of Science Part A, 85(??):200– 207, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301928.

Lenoir:1981:TRT

[Len81] Timothy Lenoir. Teleology without regrets. the transformation of physiology in Germany: 1790–1847. Studies in History and Philosophy of Science Part A, 12(4):293–354, December 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900194.

Lennon:2007:SBC

[Len07] Thomas M. Lennon. The significance of the Barrovian Case. Studies in History and Philosophy of Science Part A, 38(1):36– 55, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106001117.

Lenhard:2014:DMC

[Len14] Johannes Lenhard. Disciplines, models, and computers: the path to computational quantum chemistry. Studies in History and Philosophy of Science Part A, 48(??):89–96, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000533.

Lennox:2021:ANL

[Len21] James G. Lennox. Accentuate the negative: Locating possibility in Darwin’s ‘long argument’. Studies in History and Philosophy of Science Part A, 87(??):147–157, June 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000066.

Lenhard:2022:TBS

[Len22] Johannes Lenhard. A transformation of Bayesian statistics: Computation, prediction, and rationality. Studies in History and Philosophy of Science Part A, 92(??):144–151, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000279.

Leplin:1975:CAH

[Lep75] Jarrett Leplin. The concept of an ad hoc hypothesis. Studies in History and Philosophy of Science Part A, 5(4):309–345, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900060.

Leplin:1979:RSR

[Lep79] Jarrett Leplin. Reference and scientific realism. Studies in History and Philosophy of Science Part A, 10(4):265–284, December 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368179900165.

Leplin:1981:TSP

[Lep81] Jarrett Leplin. Truth and scientific progress. Studies in History and Philosophy of Science Part A, 12(4):269–291, December 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900182.

Lesher:1978:RGS

[Les78] J. H. Lesher. On the role of guesswork in science. Studies in History and Philosophy of Science Part A, 9(1):19–33, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900195.

Letertre:2021:OFQ

[Let21] Laurie Letertre. The operational framework for quantum theories is both epistemologically and ontologically neutral. Studies in History and Philosophy of Science Part A, 89(??):129– 137, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001205.

Letertre:2022:CNI

[Let22] Laurie Letertre. Causal nonseparability and its implications for spatiotemporal relations. Studies in History and Philosophy of Science Part A, 95(??):64–74, October 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001005.

Leuschner:2012:POE

[Leu12] Anna Leuschner. Pluralism and objectivity: Exposing and breaking a circle. Studies in History and Philosophy of Science Part A, 43(1):

191–198, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001415. See reply [Eig19].

Leuschner:2015:SEA

[Leu15] Anna Leuschner. Social exclusion in academia through biases in methodological quality evaluation: On the situation of women in science and philosophy. Studies in History and Philosophy of Science Part A, 54(??):56–63, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001107.

Levy:2001:QMS

[Lev01] Ed Levy. Quantification, mandated science and judgment. Studies in History and Philosophy of Science Part A, 32(4):723–737, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810100019X.

Levanon:2011:CTR

[Lev11] Tamar Levanon. The concept of transition and its role in Leibniz’s and Whitehead’s metaphysics of motion. Studies in History and Philosophy of Science Part A, 42(2):352–361, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001226.

Lewens:2002:TIE

[Lew02] Tim Lewens. Technological innovation as an evolutionary process Darwinnovation! Studies in History and Philosophy of Science Part A, 33(1):195–203, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000383.

Lewowicz:2011:PLP

[Lew11] Lucía Lewowicz. Phlogiston, Lavoisier and the purloined referent. Studies in History and Philosophy of Science Part A, 42(3):436– 444, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000161.

Lucken:2025:LPS

[LF25] Collin Lucken and Tim Elmo Feiten. Leveraging participatory sense-making and public engagement with science for AI democratization. Studies in History and Philosophy of Science Part A, 110(??):55–64, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368125000135.

Laursen:2021:IPD

[LGC21] Bethany K. Laursen, Chad Gonnerman, and Stephen J. Crowley. Improving philosophical dialogue interventions to better resolve problematic value pluralism in collaborative environmental science. Studies in History and Philosophy of Science Part A, 87 (??):54–71, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000194.

Lichtenstein:2021:MUS

[Lic21] Eli I. Lichtenstein. (Mis)Understanding scientific disagreement: Success versus pursuit-worthiness in theory choice. Studies in History and Philosophy of Science Part A, 85(??):166–175, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301898.

Lindenmann:1998:TPS

[Lin98] Jean Lindenmann. On Toine Pieters’ ‘shaping a new biological factor’. Studies in History and Philosophy of Science Part A, 29(1): 113–116, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936819700023X.

Lindee:2022:BRS

[Lin22a] Susan Lindee. Book review: Social Science for What? Battles over Public Funding for the “Other Sciences” at the National Science Foundation by Mark Solovey: Scientism, race relations and national security: Thinking about the social sciences in the Cold War. Studies in History and Philosophy of Science Part A, 94(??): 201–203, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000395. See [Rot22, Hau22,

Tur22b] and reply [Sol22].

Linnemann:2022:QMG

[Lin22b] Niels Linnemann. Quantisation as a method of generation: the nature and prospects of theory changes through quantisation. Studies in History and Philosophy of Science Part A, 92(??):209– 223, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001746.

Lipton:1991:IBE

[Lip91] Peter Lipton. Inference to the best explanation. Philosophical issues in science series. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1991. ISBN 0-415-05886-4 (hardcover). x + 194 pp. LCCN Q175 .L556 1991.

Lipton:1992:SNM

[Lip92] Peter Lipton. The seductive-nomological model. Studies in History and Philosophy of Science Part A, 23(4):691–698, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368192900207.

Lipton:1998:ET

[Lip98] Peter Lipton. The epistemology of testimony. Studies in History and Philosophy of Science Part A, 29(1):1–31, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368197000228.

Lipski:2020:NDN

[Lip20] Joachim Lipski. Natural diversity: a neo-essentialist misconstrual of homeostatic property cluster theory in natural kind debates. Studies in History and Philosophy of Science Part A, 82(??):94– 103, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119302183.

Lisciandra:2018:RPB

[Lis18] Chiara Lisciandra. The role of psychology in behavioral economics: the case of social preferences. Studies in History and Philosophy of Science Part A, 72(??):11–21, December 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118300244.

Liu:2022:HDR

[Liu22] Xuansong Liu. Humboldt, Darwin, and romantic resonance in science. Studies in History and Philosophy of Science Part A, 92(??): 196–208, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000309.

Livingston:1993:BRB

[Liv93] James C. Livingston. Book review: Nature lost? Natural science and the German theological traditions of the nineteenth century: Frederick Gregory, (Cambridge, Mass. and London: Harvard University Press, 1992), 341 pp. ISBN 0-674-60483-0 cloth £31.95. Studies in History and Philosophy of Science Part A, 24(2):293–303, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390049P. See [Gre92a].

Livengood:2009:WWM

[Liv09] Jonathan Livengood. Why was M. S. Tswett’s chromatographic adsorption analysis rejected? Studies in History and Philosophy of Science Part A, 40(1):57–69, March 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810800112X.

Lynning:2011:GSN

[LJ11] Kristine Hays Lynning and Anja Skaar Jacobsen. Grasping the spirit in nature: Anschauung in Ørsted’s epistemology of science and beauty. Studies in History and Philosophy of Science Part A, 42(1):45–57, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811000097X.

Livengood:2020:DMI

[LK20] Jonathan Livengood and Daniel Z. Korman. Debunking material induction. Studies in History and Philosophy of Science Part A, 84(??):20–27, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301372.

Lane:1981:PP

[LL81] N. R. Lane and S. A. Lane. Paradigms and perception. Studies in History and Philosophy of Science Part A, 12(1):47–60, March 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900042.

Laudan:2016:REH

[LL16] Larry Laudan and Rachel Laudan. The re-emergence of hyphenated history-and-philosophy-of-science and the testing of theories of scientific change. Studies in History and Philosophy of Science Part A, 59(??):74–77, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300267.

Lopez:2019:NCM

[LL19] Cristian Ariel López and Olimpia Iris Lombardi. No communication without manipulation: a causal-deflationary view of information. Studies in History and Philosophy of Science Part A, 73 (??):34–43, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117302765.

Lloyd:1997:CHP

[Llo97] Geoffrey Lloyd. The comparative history of pre-modern science: The pitfalls and the prizes. Studies in History and Philosophy of Science Part A, 28(2):363–368, June 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368196000192.

Lloyd:2012:RCE

[Llo12] Elisabeth A. Lloyd. The role of ‘complex’ empiricism in the debates about satellite data and climate models. Studies in History and Philosophy of Science Part A, 43(2):390–401, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000192.

Lloyd:2015:MRC

[Llo15] Elisabeth A. Lloyd. Model robustness as a confirmatory virtue: the case of climate science. Studies in History and Philosophy of Science Part A, 49(??):58–68, February 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114001150.

Lakatos:1970:CGK

[LM70] Imre Lakatos and Alan Musgrave, editors. Criticism and the growth of knowledge. Cambridge University Press, Cambridge, UK, 1970. ISBN 0-521-07826-1. LCCN Q175 .I514 1965. URL http://

www.loc.gov/catdir/enhancements/fy0642/78105496-d.html;
 http://www.loc.gov/catdir/enhancements/fy0642/78105496-
t.html.

Lacey:2012:ESG

[LM12] Hugh Lacey and Pablo R. Mariconda. The eagle and the starlings: Galileo’s argument for the autonomy of science — how pertinent is it today? Studies in History and Philosophy of Science Part A, 43(1):122–131, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001038.

Lohse:2021:SIA

[Loh21] Simon Lohse. Scientific inertia in animal-based research in biomedicine. Studies in History and Philosophy of Science Part A, 89(??):41–51, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001035.

Loison:2016:FPH

[Loi16] Laurent Loison. Forms of presentism in the history of science. Rethinking the project of historical epistemology. Studies in History and Philosophy of Science Part A, 60(??):29–37, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300516.

Loison:2022:EAC

[Loi22] Laurent Loison. The environment: an ambiguous concept in Waddington’s biology. Studies in History and Philosophy of Science Part A, 91(??):181–190, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812100203X.

LoLordo:2015:CEG

[LoL15] Antonia LoLordo. Copernicus, Epicurus, Galileo, and Gassendi. Studies in History and Philosophy of Science Part A, 51(??): 82–88, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000230.

Longino:2001:WDW

[Lon01] Helen E. Longino. What do we measure when we measure aggression? Studies in History and Philosophy of Science Part A, 32 (4):685–704, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000206.

Longino:2005:WPS

[Lon05]  Helen E. Longino. Whither philosophy of science?  Stud-

ies in History and Philosophy of Science Part A, 36(4):774–778, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000774.

Loparic:1984:PST

[Lop84] Zeljko Loparić. Problem-solving and theory structure in Mach. Studies in History and Philosophy of Science Part A, 15(1):23– 49, March 1984. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368184900281.

Lorenzano:2013:SCS

[Lor13] Pablo Lorenzano. The semantic conception and the structuralist view of theories: a critique of Suppe’s criticisms. Studies in History and Philosophy of Science Part A, 44(4):600–607, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000927.

Lorenat:2020:DIL

[Lor20] Jemma Lorenat. Drawing on the imagination: the limits of illustrated figures in nineteenth-century geometry. Studies in History and Philosophy of Science Part A, 82(??):75–87, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119302171.

Lorenzetti:2022:FW

[Lor22] Lorenzo Lorenzetti. Functionalising the wavefunction. Studies in History and Philosophy of Science Part A, 96(??):141–153, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001583.

Losee:1977:LEP

[Los77] John Losee. Limitations of an evolutionist philosophy of science. Studies in History and Philosophy of Science Part A, 8(4):349– 352, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900255.

Losee:1978:LPS

[Los78] John Losee. Laudan on progress in science. Studies in History and Philosophy of Science Part A, 9(4):333–340, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900146.

Losee:1983:WMR

[Los83] John Losee. Whewell and Mill on the relation between philosophy of science and history of science. Studies in History and Philosophy of Science Part A, 14(2):113–126, June 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818390016X.

Louden:2008:AKP

[Lou08] Robert B. Louden. Anthropology from a Kantian point of view: toward a cosmopolitan conception of human nature. Studies in History and Philosophy of Science Part A, 39(4):515–522, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810800085X.

Love:2015:CEE

[Lov15] Alan C. Love. Collaborative explanation, explanatory roles, and scientific explaining in practice. Studies in History and Philosophy of Science Part A, 52(??):88–94, August 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000382.

Low:1993:HEA

[Low93] Morris F. Low. The history of East Asian science: State of the art. Studies in History and Philosophy of Science Part A, 24(4):677– 686, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390058R.

Lowy:1997:LT

[Löw97] Ilana Löwy. The legislation of things. Studies in History and Philosophy of Science Part A, 28(3):533–543, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368197863376.

Lowy:2008:WSL

[Löw08] Ilana Löwy. Ways of seeing: Ludwik Fleck and Polish debates on the perception of reality, 1890–1947. Studies in History and Philosophy of Science Part A, 39(3):375–383, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000629.

Lloyd:2002:WWS

[LS02] G. E. R. (Geoffrey Ernest Richard) Lloyd and Nathan Sivin. The way and the word: science and medicine in early China and Greece. Yale University Press, New Haven, CT, USA, 2002. ISBN 0-300-09297-0, 0-300-10160-0 (paperback). xvii + 348 pp. LCCN Q127.C5 L6 2002.

Lenhard:2024:CPH

[LSH24] Johannes Lenhard, Simon Stephan, and Hans Hasse. A child of prediction. on the history, ontology, and computation of the Lennard-Jonesium. Studies in History and Philosophy of Science Part A, 103 (??):105–113, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123001668.

Lujan:2020:SEC

[LT20] José Luis Luján and Oliver Todt. Standards of evidence and causality in regulatory science: Risk and benefit assessment. Studies in History and Philosophy of Science Part A, 80(??): 82–89, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118302462.

Luchins:1971:BRB

[Luc71] Abraham S. Luchins. Book review: Mind and brain: a philosophy of science: Arturo Rosenblueth Cambridge, Mass.: MIT Press, 1970. xii + 128 pp. $5.95. Studies in History and Philosophy of Science Part A, 2(3):287–294, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368171900471. See [Ros70].

Luchetti:2020:SMB

[Luc20] Michele Luchetti. From successful measurement to the birth of a law: Disentangling coordination in Ohm’s scientific practice. Studies in History and Philosophy of Science Part A, 84(??):119– 131, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301758.

Ludwig:2016:OOI

[Lud16] David Ludwig. Overlapping ontologies and indigenous knowledge. from integration to ontological self-determination. Studies in History and Philosophy of Science Part A, 59(??):36– 45, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300188.

Lugg:1978:OPS

[Lug78] Andrew Lugg. Overdetermined problems in science. Studies in History and Philosophy of Science Part A, 9(1):1–18, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900183.

Lugg:1992:WGR

[Lug92] Andrew Lugg. What generativism is not: a reply to Brian Baigrie. Studies in History and Philosophy of Science Part A, 23(3):499– 501, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819290006R. See [Bai90, Bai92].

Lukes:1982:CDB

[Luk82] Steven Lukes. Comments on David Bloor. Studies in History and Philosophy of Science Part A, 13(4):313–318, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900152. See [Blo82a, Buc82, Blo82b,

Blo82c, Blo82c, Hes82].

Lusk:2016:CSF

[Lus16] Greg Lusk. Computer simulation and the features of novel empirical data. Studies in History and Philosophy of Science Part A, 56(??):145–152, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001430.

Lusk:2021:DDR

[Lus21] Greg Lusk. Does democracy require value-neutral science? analyzing the legitimacy of scientific information in the political sphere. Studies in History and Philosophy of Science Part A, 90(??):102– 110, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001254.

Luxembourg:1967:FBD

[Lux67] Lilo K. Luxembourg. Francis Bacon and Dénis Diderot: Philosophers of science. Munksgaard, København, Danmark, 1967. 127 pp. LCCN B1198 .L8.

Lynch:1990:RSP

[LW90] Michael Lynch and Steve Woolgar, editors. Representation in scientific practice. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-62076-6. x + 365 pp. LCCN Q222 .R46 1990.

Lennox:1994:NSS

[LW94] James G. Lennox and Bradley E. Wilson. Natural selection and the struggle for existence. Studies in History and Philosophy of Science Part A, 25(1):65–80, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900205.

Lourenco:2013:SHR

[LW13a] Marta C. Lourenço and Lydia Wilson. Scientific heritage: Reflections on its nature and new approaches to preservation, study and access. Studies in History and Philosophy of Science Part A, 44 (4):744–753, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000538.

Ludwig:2013:RSC

[LW13b] David Ludwig and Cornelia Weber. A rediscovery of scientific collections as material heritage? The case of university collections in Germany. Studies in History and Philosophy of Science Part A, 44 (4):652–659, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811300054X.

Lynch:1995:BGI

[Lyn95] Michael Lynch. Building a global infrastructure. Studies in History and Philosophy of Science Part A, 26(1):167–172, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194000325.

Lyons:2016:SRV

[Lyo16] Timothy D. Lyons. Structural realism versus deployment realism: a comparative evaluation. Studies in History and Philosophy of Science Part A, 59(??):95–105, October 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300255.

Lyre:2014:BPQ

[Lyr14] Holger Lyre. Berry phase and quantum structure. Studies in History and Philosophy of Science Part A, 48 (part A)(??):45– 51, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000975.

Maas:1999:MRJ

[Maa99] Harro Maas. Mechanical rationality: Jevons and the making of economic man. Studies in History and Philosophy of Science Part A, 30(4):587–619, December 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368199000308.

Maas:2013:HPB

[Maa13] Ad Maas. How to put a black box in a showcase: History of science museums and recent heritage. Studies in History and Philosophy of Science Part A, 44(4):660–668, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000551.

Mackinnon:1972:TEM

[Mac72] Edward Mackinnon. Theoretical entities and metatheories. Studies in History and Philosophy of Science Part A, 3(2):105–117, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368172900180.

Machamer:1973:FGI

[Mac73] Peter K. Machamer. Feyerabend and Galileo: The interaction of theories, and the reinterpretation of experience. Studies in History and Philosophy of Science Part A, 4(1):1–46, May 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368173900228.

Machamer:1975:USC

[Mac75] Peter K. Machamer. Understanding scientific change. Studies in History and Philosophy of Science Part A, 5(4):373–381, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900102.

MacKenzie:1984:RSY

[Mac84] Donald MacKenzie. Reply to Steven Yearley. Studies in History and Philosophy of Science Part A, 15(3):251–259, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368184900219. See [Yea82].

Maclean:2000:NPR

[Mac00] Ian Maclean. Natural and preternatural in Renaissance philosophy and medicine. Studies in History and Philosophy of Science Part A, 31(2):331–342, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000066.

MacDougall:2010:PIR

[Mac10] Margaret MacDougall. Poincaréan intuition revisited: what can we learn from Kant and Parsons? Studies in History and Philosophy of Science Part A, 41(2):138–147, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000178.

MacLeod:2013:SDB

[Mac13] Christine MacLeod. “A delicate business”: Wartime airplane designs and their post-war evaluation, 1919–1924. Studies in History and Philosophy of Science Part A, 44(2):269–279, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112001082.

Magnus:1996:TPE

[Mag96] David Magnus. Theory, practice, and epistemology in the development of species concepts. Studies in History and Philosophy of Science Part A, 27(4):521–545, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819600009X.

Mahnke:1977:HHF

[Mah77] Dietrich Mahnke. From Hilbert to Husserl: First introduction to phenomenology, especially that of formal mathematics. Studies in History and Philosophy of Science Part A, 8(1):71– 84, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900206.

Maher:1999:CBT

[Mah99] Patrick Maher. The confirmation of Black’s Theory of Lime. Studies in History and Philosophy of Science Part A, 30(2):335– 353, June 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936819800048X.

Maki:1997:UMR

[Mäk97] Uskali Mäki. Universals and the methodenstreit: a re-examination of Carl Menger’s conception of economics as an exact science. Studies in History and Philosophy of Science Part A, 28(3):475– 495, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000118.

Makkreel:2003:CKD

[Mak03] Rudolf A. Makkreel. The cognition-knowledge distinction in Kant and Dilthey and the implications for psychology and selfunderstanding. Studies in History and Philosophy of Science Part A, 34(1):149–164, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368102000870.

Makkreel:2008:KDH

[Mak08] Rudolf A. Makkreel. Kant and the development of the human and cultural sciences. Studies in History and Philosophy of Science Part A, 39(4):546–553, December 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000897.

Malet:1990:KIG

[Mal90] Antoni Malet. Keplerian illusions: Geometrical pictures vs optical images in Kepler’s visual theory. Studies in History and Philosophy of Science Part A, 21(1):1–40, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819090013X.

Malet:1991:MMS

[Mal91] Antoni Malet. Mathematics and mathematization in the seventeenth century. Studies in History and Philosophy of Science Part A, 22(4):673–678, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190040Y.

Malone:1993:KRI

[Mal93] Michael E. Malone. Kuhn reconstructed: Incommensurability without relativism. Studies in History and Philosophy of Science Part A, 24(1):69–93, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390025F.

Malet:2001:PIM

[Mal01a] Antoni Malet. The power of images: mathematics and metaphysics in Hobbes’s optics. Studies in History and Philosophy of Science Part A, 32(2):303–333, June 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810000039X.

Malzkorn:2001:DDC

[Mal01b] Wolfgang Malzkorn. Defining disposition concepts: a brief history of the problem. Studies in History and Philosophy of Science Part A, 32(2):335–353, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810000042X.

Mallon:2013:WRT

[Mal13] Ron Mallon. Was Race thinking invented in the modern West? Studies in History and Philosophy of Science Part A, 44(1):77– 88, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000763.

Mandelbaum:1972:WDT

[Man72] Maurice Mandelbaum. To what does the term ‘psychology’ refer? Studies in History and Philosophy of Science Part A, 2(4):347– 361, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900039.

Manier:1980:DLL

[Man80a] Edward Manier. Darwin’s language and logic. Studies in History and Philosophy of Science Part A, 11(4):305–323, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368180900096.

Manier:1980:HPS

[Man80b] Edward Manier. History, philosophy and sociology of biology: a family romance. Studies in History and Philosophy of Science Part A, 11(1):1–24, March 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368180900023.

Mancosu:1992:ALE

[Man92] Paolo Mancosu. Aristotelian logic and Euclidean mathematics: Seventeenth-century developments of the quaestio de certitudine mathematicarum. Studies in History and Philosophy of Science Part A, 23(2):241–265, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368192900344.

Manson:2002:EC

[Man02] Neil Campbell Manson. Epistemic consciousness. Studies in History and Philosophy of Science Part A, 33(3):425–441, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000250.

Mandelbrote:2004:NNI

[Man04] Scott Mandelbrote. Newton and Newtonianism: an introduction. Studies in History and Philosophy of Science Part A, 35(3):415– 425, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000482.

Mandosio:2012:BPD

[Man12a] Jean-Marc Mandosio. Beyond Pico della Mirandola: John Dee’s ‘formal numbers’ and ‘real cabala’. Studies in History and Philosophy of Science Part A, 43(3):489–497, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811100118X.

Manning:2012:AFD

[Man12b] Gideon Manning. Analogy and falsification in Descartes’ physics. Studies in History and Philosophy of Science Part A, 43(2):402– 411, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000210.

Manafu:2018:IMR

[Man18] Alex Manafu. Introduction: Multiple realizability and levels of reality. Studies in History and Philosophy of Science Part A, 68 (??):1–2, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811830030X.

Manero:2022:SLS

[Man22] Jorge Manero. Structural losses, structural realism and the stability of Lie algebras. Studies in History and Philosophy of Science Part A, 91(??):28–40, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001801.

Martin:1988:EJF

[Mar88] R. J. J. Martin. Explaining John Freind’s history of physick. Studies in History and Philosophy of Science Part A, 19(4):399– 418, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900088.

Marcos:1997:TBA

[Mar97] Alfredo Marcos. The tension between Aristotle’s theories and uses of metaphor. Studies in History and Philosophy of Science Part A, 28(1):123–139, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000106.

Martens:1999:KSP

[Mar99] Rhonda Martens. Kepler’s solution to the problem of a realist celestial mechanics. Studies in History and Philosophy of Science Part A, 30(3):377–394, September 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000138.

Marcos:2004:TSI

[Mar04] Alfredo Marcos. Towards a science of the individual: the Aristotelian search for scientific knowledge of individual entities. Studies in History and Philosophy of Science Part A, 35(1):73–89, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000943.

Martens:2009:HSA

[Mar09] Rhonda Martens. Harmony and simplicity: aesthetic virtues and the rise of testability. Studies in History and Philosophy of Science Part A, 40(3):258–266, September 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000533.

Marshall:2013:GDA

[Mar13] Douglas Bertrand Marshall. Galileo’s defense of the application of geometry to physics in the Dialogue. Studies in History and Philosophy of Science Part A, 44(2):178–187, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000150.

Martin:2015:IA

[Mar15a] Craig Martin. The invention of atmosphere. Studies in History and Philosophy of Science Part A, 52(??):44–54, August 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000679.

Martin:2015:NSO

[Mar15b] Joseph D. Martin. New straw for the old broom. Studies in History and Philosophy of Science Part A, 54(??):138–143, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001387.

Martin:2016:LFM

[Mar16] Eric C. Martin. Late Feyerabend on materialism, mysticism, and religion. Studies in History and Philosophy of Science Part A, 57(??):129–136, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001776.

Martin:2021:RTM

[Mar21] Joseph D. Martin. The relativity of theory by Moti Mizrahi: On the necessity of history in philosophy of science. Studies in History and Philosophy of Science Part A, 87(??):170– 172, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000480.

Mariani:2022:NAM

[Mar22] Cristian Mariani. Non-accessible mass and the ontology of GRW. Studies in History and Philosophy of Science Part A, 91(??):270– 279, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001928.

Marre:2024:KLF

[Mar24] Thomas Marré. Kant on the logical form of organized being. Studies in History and Philosophy of Science Part A, 108(??):46– 54, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001389.

Massimi:2011:KDT

[Mas11] Michela Massimi. Kant’s dynamical theory of matter in 1755, and its debt to speculative Newtonian experimentalism. Studies in History and Philosophy of Science Part A, 42(4):525–543, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000720.

Massimi:2013:PNS

[Mas13] Michela Massimi. Philosophy of natural science from Newton to Kant. Studies in History and Philosophy of Science Part A, 44 (3):393–395, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000854.

Massimi:2015:WNW

[Mas15] Michela Massimi. ‘Working in a new world’: Kuhn, constructivism, and mind-dependence. Studies in History and Philosophy of Science Part A, 50(??):83–89, April 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000892.

Massimi:2024:RMB

[Mas24] Michela Massimi. Replies to Mieke Boon and Catherine Kendig. Studies in History and Philosophy of Science Part A, 103(??):16– 19, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001358.

Mattern:1980:LAP

[Mat80] R. M. Mattern. Locke on active power and the obscure idea of active power from bodies. Studies in History and Philosophy of Science Part A, 11(1):39–77, March 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368180900047.

Matthews:1998:CSE

[Mat98] Michael R. Matthews, editor. Constructivism in science education: a philosophical examination. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-5033-2. xii + 234 pp. LCCN Q181 .C612 1998.

Matthewson:2011:TOM

[Mat11] John Matthewson. Trade-offs in model-building: a more target-oriented approach. Studies in History and Philosophy of Science Part A, 42(2):324–333, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001196.

Matthiessen:2019:RCM

[Mat19] Dana Matthiessen. The rise of cryptographic metaphors in Boyle and their use for the mechanical philosophy. Studies in History and Philosophy of Science Part A, 73(??):8–21, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117302522.

Matthewson:2020:DGM

[Mat20] John Matthewson. Detail and generality in mechanistic explanation. Studies in History and Philosophy of Science Part A, 80(??): 28–36, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300304.

Matthews:2022:HCL

[Mat22a] Lucas J. Matthews. Half a century later and we’re back where we started: How the problem of locality turned in to the problem of portability. Studies in History and Philosophy of Science Part A, 91(??):1–9, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001771.

Matthiessen:2022:ETA

[Mat22b] Dana Matthiessen. Empirical techniques and the accuracy of scientific representations. Studies in History and Philosophy of Science Part A, 94(??):143–157, August 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000838.

Maull:1977:USR

[Mau77] Nancy L. Maull. Unifying science without reduction. Studies in History and Philosophy of Science Part A, 8(2):143–162, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900127.

Maudlin:1990:SST

[Mau90] Tim Maudlin. Substances and space-time: What Aristotle would have said to Einstein. Studies in History and Philosophy of Science Part A, 21(4):531–561, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900324.

Mauskopf:2013:H

[Mau13] Seymour H. Mauskopf. Historicizing H_{2}O. Studies in History and Philosophy of Science Part A, 44(4):623–630, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000940.

Maxwell:2022:EPB

[Max22] Anne Maxwell. Eugenics and photography in Britain, the USA and Australia 1870–1940. Studies in History and Philosophy of Science Part A, 92(??):71–85, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812200005X.

Mayr:1971:BRB

[May71] Ernst Mayr. Book review: The life and letters of Charles Darwin: Charles Darwin ed., Francis Darwin. Johnson Reprint Corporation: New York, 1969. 3 vols. 50. Studies in History and Philosophy of Science Part A, 2(3):273–280, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368171900458. See [DD69a, DD72].

Mayer:2000:SDC

[May00] Anna-K Mayer. Setting up a discipline: Conflicting agendas of the Cambridge History of Science Committee, 1936–1950. Studies in History and Philosophy of Science Part A, 31(4):665–689, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000261.

Mayer:2004:SDI

[May04] Anna-K. Mayer. Setting up a discipline, II: British history of science and “the end of ideology”, 1931–1948. Studies in History and Philosophy of Science Part A, 35(1):41–72, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000967.

Mayer:2011:CGS

[May11] Thomas F. Mayer. The censoring of Galileo’s Sunspot Letters and the first phase of his trial. Studies in History and Philosophy of Science Part A, 42(1):1–10, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001068.

Mayo:2014:SSF

[May14] D. Mayo. Some surprising facts about (the problem of) surprising facts: (from the Dusseldorf Conference, February 2011). Studies in History and Philosophy of Science Part A, 45(??):79– 86, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113001015.

Maziarz:2022:MAR

[Maz22] Mariusz Maziarz. Is meta-analysis of RCTs assessing the efficacy of interventions a reliable source of evidence for therapeutic decisions? Studies in History and Philosophy of Science Part A, 91 (??):159–167, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001849.

McCullough-Benner:2022:AUM

[MB22] Colin McCullough-Benner. Applying unrigorous mathematics: Heaviside’s operational calculus. Studies in History and Philosophy of Science Part A, 91(??):113–124, February 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001904.

McClimans:2017:COM

[MBC17] Leah McClimans, John Browne, and Stefan Cano. Clinical outcome measurement: Models, theory, psychometrics and practice. Studies in History and Philosophy of Science Part A, 65–66(??):67– 73, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301619.

Mercuri:2021:PPC

[MBG21] Mathew Mercuri, Brian S. Baigrie, and Amiram Gafni. Patient participation in the clinical encounter and clinical practice guidelines: the case of patients’ participation in a GRADEd world. Studies in History and Philosophy of Science Part A, 85(??):192– 199, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301916.

Meneganzin:2025:WYC

[MC25] Andra Meneganzin and Adrian Currie. Not wasted on the young: Childhood, trait complexes and human behavioral ecology. Studies in History and Philosophy of Science Part A, 109(??):12– 20, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001663.

McAllister:1996:EST

[McA96] James W. McAllister. The evidential significance of thought experiment in science. Studies in History and Philosophy of Science Part A, 27(2):233–250, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000445.

McAllister:2003:ARE

[McA03] James W. McAllister. Algorithmic randomness in empirical data. Studies in History and Philosophy of Science Part A, 34(3):633– 646, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000475. See replies [TGD05,

McA05].

McAllister:2005:ACE

[McA05] James W. McAllister. Algorithmic compression of empirical data: reply to Twardy, Gardner, and Dowe. Studies in History and Philosophy of Science Part A, 36(2):403–410, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000221. See [McA03, TGD05].

McClellan:2001:LGC

[McC01] Chris McClellan. The legacy of Georges Cuvier in Auguste Comte’s natural philosophy. Studies in History and Philosophy of Science Part A, 32(1):1–29, March 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000418.

McCormick:2006:APA

[McC06] Ted McCormick. Alchemy in the political arithmetic of Sir William Petty (1623–1687). Studies in History and Philosophy of Science Part A, 37(2):290–307, June 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000185.

McCaskey:2020:RMT

[McC20] John P. McCaskey. Reviving material theories of induction. Studies in History and Philosophy of Science Part A, 83(??):1– 7, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301062.

McCoy:2021:MES

[McC21] C. D. McCoy. Meta-empirical support for eliminative reasoning. Studies in History and Philosophy of Science Part A, 90(??):15– 29, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001369.

McEvoy:1976:BRB

[McE76] J. G. McEvoy. Book review: The Popper–Carnap controversy: Alex C. Michalos, The Hague: Martinus Nijhoff, 1971. 124 pp $25.50. Studies in History and Philosophy of Science Part A, 7 (1):63–85, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900042. See [Mic71b].

McGucken:1969:NCS

[McG69] William McGucken. Nineteenth-century spectroscopy; development of the understanding of spectra, 1802–1897. The Johns Hopkins University Press, Baltimore, MD, USA, 1969. ISBN 0-8018-1059-0. xii + 233 pp. LCCN QC451 .M24 1969.

McGuire:1970:AAN

[McG70] J. E. McGuire. Atoms and the ‘analogy of nature’: Newton’s third rule of philosophizing. Studies in History and Philosophy of Science Part A, 1(1):3–58, May 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368170900245.

McGaugh:2021:TGF

[McG21] Stacy S. McGaugh. Testing galaxy formation and dark matter with low surface brightness galaxies. Studies in History and Philosophy of Science Part A, 88(??):220–236, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000728.

Mari:2017:SIM

[MCGP17] Luca Mari, Paolo Carbone, Alessandro Giordani, and Dario Petri. A structural interpretation of measurement and some related epistemological issues. Studies in History and Philosophy of Science Part A, 65–66(??):46–56, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301942.

McHugh:2021:BRS

[McH21] Nancy Arden McHugh. Book review: Science and Moral Imagination: a New Ideal for Values in Science by Matthew J. Brown: Moral Imagination and Transactionally Situated Knowing: Author Meets Critics. Studies in History and Philosophy of Science Part A, 89(??):295–296, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812100039X.

McKinney:1995:BJP

[McK95] William J. McKinney. Between justification and pursuit: Understanding the technological essence of science. Studies in History and Philosophy of Science Part A, 26(3):455–468, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819500011B.

McKenzie:2016:LFB

[McK16] Kerry McKenzie. Looking forward, not back: Supporting structuralism in the present. Studies in History and Philosophy of Science Part A, 59(??):87–94, October 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300231.

McLeish:2005:SRB

[McL05] Christina McLeish. Scientific realism bit by bit: Part I. Kitcher on reference. Studies in History and Philosophy of Science Part A, 36 (4):668–686, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000828.

McLaughlin:2006:MPA

[McL06a] P. McLaughlin. Mechanical philosophy and artefact explanation. Studies in History and Philosophy of Science Part A, 37(1):97– 101, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105001135.

McLeish:2006:RBB

[McL06b] Christina McLeish. Realism bit by bit: Part II. disjunctive partial reference. Studies in History and Philosophy of Science Part A, 37(2):171–190, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000239.

McLoone:2024:RFI

[McL24] Brian McLoone. R. A. Fisher, indeterminism, and the fundamental theorem of natural selection. Studies in History and Philosophy of Science Part A, 105(??):120–125, June 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000487.

McMullin:1985:GI

[McM85] Ernan McMullin. Galilean idealization. Studies in History and Philosophy of Science Part A, 16(3):247–273, September 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900032.

McMullin:1993:IPA

[McM93] Ernan McMullin. Indifference principle and anthropic principle in cosmology. Studies in History and Philosophy of Science Part A, 24(3):359–389, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390034H.

McNulty:2015:RRU

[McN15] Michael Bennett McNulty. Rehabilitating the regulative use of reason: Kant on empirical and chemical laws. Studies in History and Philosophy of Science Part A, 54(??):1–10, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000758.

McNulty:2022:SGS

[McN22] Michael Bennett McNulty. A science for gods, a science for humans: Kant on teleological speculations in natural history. Studies in History and Philosophy of Science Part A, 94(??):47– 55, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812200067X.

McOuat:1996:SRM

[McO96] Gordon R. McOuat. Species, rules and meaning: The politics of language and the ends of definitions in 19th century natural history. Studies in History and Philosophy of Science Part A, 27 (4):473–519, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000607.

McOuat:2001:CNJ

[McO01] Gordon McOuat. From cutting nature at its joints to measuring it: New kinds and new kinds of people in biology. Studies in History and Philosophy of Science Part A, 32(4):613–645, December 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000279.

McRae:1972:BRB

[McR72] Robert McRae. Book review: Metaphysics and the philosophy of science. The classical origins: Descartes to Kant: Gerd Buchdahl Blackwell: Oxford and MIT Press: Cambridge, Mass., 1969. xii + 714 pp. £6.25; $15.00. Studies in History and Philosophy of Science Part A, 3(1):89–99, May 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368172900167. See [Buc69a,

Buc69b].

Massimi:2013:CEK

[MD13] Michela Massimi and Silvia De Bianchi. Cartesian echoes in Kant’s philosophy of nature. Studies in History and Philosophy of Science Part A, 44(3):481–492, September 2013. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000866.

McKaughan:2015:ICA

[ME15] Daniel J. McKaughan and Kevin C. Elliott. Introduction: Cognitive attitudes and values in science. Studies in History and Philosophy of Science Part A, 53(??):57–61, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000710.

Meehl:1992:MAR

[Mee92] Paul E. Meehl. The miracle argument for realism: An important lesson to be learned by generalizing from Carrier’s counterexamples. Studies in History and Philosophy of Science Part A, 23(2):267–282, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368192900355.

Meehan:2021:SII

[Mee21] Alexander Meehan. States of ignorance and ignorance of states: Examining the Quantum Principal Principle. Studies in History and Philosophy of Science Part A, 90(??):160–167, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001539.

Meiring:2021:SPA

[Mei21] H. Meiring. Scientific patronage in the age of Darwin: the curious case of William Boyd Dawkins. Studies in History and Philosophy of Science Part A, 89(??):267–282, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001102.

Meli:1989:FCH

[Mel89] Domenico Bertoloni Meli. Federico Commandino and his school. Studies in History and Philosophy of Science Part A, 20(3):397– 403, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368189900162.

Meli:1991:PCP

[Mel91] D. Bertoloni Meli. Public claims, private worries: Newton’s Principia and Leibniz’s theory of planetary motion. Studies in History and Philosophy of Science Part A, 22(3):415–449, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190003B. See [New87, NCW99, New09].

Meli:2004:BRB

[Mel04] Domenico Bertoloni Meli. Book review: The foundation of Newtonian scholarship: Richard H. Dalitz, & Michael Nauenberg (Eds.); World Scientific, Singapore & London, 2000, pp. xviii + 242, Price £44.00 hardback, ISBN 981-02-3920-3, Price £29.00 paperback, ISBN 981-02-4044-9. Studies in History and Philosophy of Science Part A, 35(3):667–669, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000639. See [DN00].

Meloni:2021:PEB

[Mel21] Maurizio Meloni. The politics of environments before the environment: Biopolitics in the longue durée. Studies in History and Philosophy of Science Part A, 88(??):334–344, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000959.

Menke:2014:DMA

[Men14] Cornelis Menke. Does the miracle argument embody a base rate fallacy? Studies in History and Philosophy of Science Part A, 45(??):103–108, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113001027.

Menke:2018:WMD

[Men18] Cornelis Menke. The Whewell–Mill debate on predictions, from Mill’s point of view. Studies in History and Philosophy of Science Part A, 69(??):60–71, June 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302030.

Menon:2019:VNA

[Men19] Tushar Menon. On the viability of the No Alternatives Argument. Studies in History and Philosophy of Science Part A, 76(??):69– 75, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301109.

Menon:2024:ANM

[Men24] Tushar Menon. On algebraic naturalism and metaphysical indeterminacy in quantum mechanics. Studies in History and Philosophy of Science Part A, 105(??):1–16, June 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000244.

Mertz:1980:GMC

[Mer80] Donald W. Mertz. On Galileo’s method of causal proportionality. Studies in History and Philosophy of Science Part A, 11(3): 229–242, September 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368180900278.

Mertz:1982:CSG

[Mer82] Donald W. Mertz. The concept of structure in Galileo: Its role in the methods of proportionality and ex suppositione as applied to the tides. Studies in History and Philosophy of Science Part A, 13(2):111–131, June 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936818290022X.

Mertens:1992:CST

[Mer92] Joost Mertens. The conceptual structure of the technological sciences and the importance of action theory. Studies in History and Philosophy of Science Part A, 23(2):333–348, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368192900388.

Mertens:2004:PIN

[Mer04] Joost Mertens. Philosophical instruments: Notion displayers, black boxes, and their usefulness. Studies in History and Philosophy of Science Part A, 35(4):851–859, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000925.

Merritt:2021:CR

[Mer21] David Merritt. Cosmological realism. Studies in History and Philosophy of Science Part A, 88(??):193–208, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000753.

Metcalfe:1991:WDP

[Met91] John F. Metcalfe. Whewell’s developmental psychologism: a Victorian account of scientific progress. Studies in History and Philosophy of Science Part A, 22(1):117–139, March 1991. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936819190017M.

Methot:2013:GCE

[Mét13] Pierre-Olivier Méthot. On the genealogy of concepts and experimental practices: Rethinking Georges Canguilhem’s historical epistemology. Studies in History and Philosophy of Science Part A, 44(1):112–123, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000386.

Meunier:2019:PKR

[Meu19] Robert Meunier. Project knowledge and its resituation in the design of research projects: Seymour Benzer’s behavioral genetics, 1965–1974. Studies in History and Philosophy of Science Part A, 77(??):39–53, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300980.

Murphy:1984:LLH

[MH84] Anthony C. Murphy and R. E. Hendrick. Lakatos, Laudan and the hermeneutic circle. Studies in History and Philosophy of Science Part A, 15(2):119–130, June 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900025.

McGlacken:2022:CIP

[MHW22] Renelle McGlacken and Pru Hobson-West. Critiquing imaginaries of ‘the public’ in UK dialogue around animal research: Insights from the Mass Observation Project. Studies in History and Philosophy of Science Part A, 91(??):280–287, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121002090.

Mialet:2001:WAB

[Mia01] Hélène Mialet. We have always been mixed up: Aristotle at the heart of the ‘composite age’. Studies in History and Philosophy of Science Part A, 32(1):193–202, March 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000285.

Michalos:1971:DTA

[Mic71a] Alex C. Michalos. Discussion: Theory appraisal and the growth of scientific knowledge. Studies in History and Philosophy of Science Part A, 1(4):353–361, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900124.

Michalos:1971:PCC

[Mic71b] Alex C. Michalos. The Popper–Carnap controversy. Martinus Nijhoff, The Hague, The Netherlands, 1971. ISBN 90-247-5127-6. 134 pp. LCCN Q175 .M59.

Michell:1993:ORT

[Mic93] Joel Michell. The origins of the representational theory of measurement: Helmholtz, Hölder, and Russell. Studies in History and Philosophy of Science Part A, 24(2):185–206, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390045L.

Michael:2024:FBP

[Mic24] Michael T. Michael. Freud, bullshit, and pseudoscience. Studies in History and Philosophy of Science Part A, 108(??):64–72, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001353.

Milkov:2008:RDL

[Mil08] Nikolay Milkov. Russell’s debt to Lotze. Studies in History and Philosophy of Science Part A, 39(2):186–193, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000204.

Miller:2014:CWW

[Mil14] Boaz Miller. Catching the WAVE: the Weight-Adjusting Account of Values and Evidence. Studies in History and Philosophy of Science Part A, 47(??):69–80, September 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000193.

Millstein:2019:TEC

[Mil19] Roberta L. Millstein. Types of experiments and causal process tracing: What happened on the Kaibab Plateau in the 1920s. Studies in History and Philosophy of Science Part A, 78(??):98– 104, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301900.

Miller:2023:BRM

[Mil23] Michael E. Miller. Book review: A Middle Way: a Non-Fundamental Approach to Many-Body Physics by Robert Batterman: From Scales to Levels. Studies in History and Philosophy of Science Part A, 97(??):126–127, February 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001297. See [Bat23].

Mirowski:1989:HDT

[Mir89] Philip Mirowski. How not to do things with metaphors: Paul Samuelson and the science of neoclassical economics. Studies in History and Philosophy of Science Part A, 20(2):175–191, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900022.

Mirowski:2004:SDS

[Mir04] Philip Mirowski. The scientific dimensions of social knowledge and their distant echoes in 20th-century American philosophy of science. Studies in History and Philosophy of Science Part A, 35(2): 283–326, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000160.

Mirowski:2005:HOC

[Mir05] Philip Mirowski. Hoedown at the OK Corral: more reflections on the ‘social’ in current philosophy of science. Studies in History and Philosophy of Science Part A, 36(4):790–800, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000853.

Misak:1994:BRB

[Mis94] Cheryl Misak. Book review: William James: Pragmatism in focus: Doris Olin (ed.) (London: Routledge, 1992), viii + 251 pp. ISBN 0-415-04057-4 Paperback £12.99, ISBN 0-415-04056-6 Hardback £40.00. Studies in History and Philosophy of Science Part A, 25 (1):123–129, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819490023X. See [Oli92].

Mittelstrass:1972:GRH

[Mit72a] Jürgen Mittelstrass. The Galilean revolution: The historical fate of a methodological insight. Studies in History and Philosophy of Science Part A, 2(4):297–328, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900015.

Mittelstrass:1972:MEK

[Mit72b] Jürgen Mittelstrass. Methodological elements of Keplerian astronomy. Studies in History and Philosophy of Science Part A, 3(3): 203–232, November 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936817290009X.

Mitchell:1986:IIT

[Mit86] W. J. T. (William John Thomas) Mitchell. Iconology: image, text, ideology. University of Chicago Press, Chicago, IL, USA and London, UK, 1986. ISBN 0-226-53229-1 (paperback), 0-226-53228-3 (hardcover). x + 226 pp. LCCN N7565 .M5 1986. US$22.00.

Mitchell:1993:MMA

[Mit93] Sam Mitchell. Mach’s mechanics and absolute space and time. Studies in History and Philosophy of Science Part A, 24(4):565– 583, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390053M.

Mitchell:2017:WNR

[Mit17] Daniel Jon Mitchell. What’s nu? A re-examination of Maxwell’s ‘ratio-of-units’ argument, from the mechanical theory of the electromagnetic field to ‘On the elementary relations between electrical measurements’. Studies in History and Philosophy of Science Part A, 65–66(??):87–98, October/December 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301607.

Mitsch:2022:HSA

[Mit22] Chris Mitsch. Hilbert-style axiomatic completion: On von Neumann and hidden variables in quantum mechanics. Studies in History and Philosophy of Science Part A, 95(??):84–95, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000978.

Miyake:2015:UDK

[Miy15] Teru Miyake. Underdetermination and decomposition in Kepler’s Astronomia Nova. Studies in History and Philosophy of Science Part A, 50(??):20–27, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000867.

Miyake:2017:MMM

[Miy17] Teru Miyake. Magnitude, moment, and measurement: the seismic mechanism controversy and its resolution. Studies in History and Philosophy of Science Part A, 65–66(??):112–120, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301590.

Miyake:2023:BRI

[Miy23] Teru Miyake. Book review: The instrument of science by Darrell Rowbottom: Property instrumentalism and inference chains. Studies in History and Philosophy of Science Part A, 98(??): 12–13, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000250.

Mizrahi:2012:WUA

[Miz12] Moti Mizrahi. Why the ultimate argument for scientific realism ultimately fails. Studies in History and Philosophy of Science Part A, 43(1):132–138, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001051.

Mizrahi:2021:RTM

[Miz21] Moti Mizrahi. The relativity of theory by Moti Mizrahi: Reply by the author. Studies in History and Philosophy of Science Part A, 87(??):173–174, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000467.

Majszak:2023:EJC

[MJ23] Mason Majszak and Julie Jebeile. Expert judgment in climate science: How it is used and how it can be justified. Studies in History and Philosophy of Science Part A, 100(??):32– 38, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000857.

Mjelva:2024:DCE

[Mje24] Jørn Kløvfjell Mjelva. Delayed-choice entanglement swapping experiments: No evidence for timelike entanglement. Studies in History and Philosophy of Science Part A, 105(??):138–148, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000542.

Mariscal:2018:BRC

[ML18] Carlos Mariscal and Alexander Lerner. Book review: Chance in Evolution, Grant Ramsey, Charles H. Pence (Eds.). University of Chicago Press, Chicago (2016), 359, Price $45.00 cloth, ISBN: 978-0-226-40188-1. Studies in History and Philosophy of Science Part A, 69(??):97–100, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300487.

Machamer:2006:DCM

[MM06] Peter Machamer and J. E. McGuire. Descartes’s changing mind. Studies in History and Philosophy of Science Part A, 37(3):398– 419, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000483.

Maki:2009:SEU

[MM09] Uskali Mäki and Caterina Marchionni. On the structure of explanatory unification: the case of geographical economics. Studies in History and Philosophy of Science Part A, 40(2):185– 195, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000211.

Maienschein:2024:BFW

[MM24a] Jane Maienschein and Kate MacCord. Book forum: What is Regeneration? By Jane Maienschein and Kate MacCord: Reply by the authors. Studies in History and Philosophy of Science Part A, 104(??):12–13, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000025.

Matarese:2024:WRM

[MM24b] Vera Matarese and C. D. McCoy. When “replicability” is more than just “reliability”: the Hubble constant controversy. Studies in History and Philosophy of Science Part A, 107(??):1– 10, October 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001122.

MacLeod:2018:WDI

[MN18] Miles MacLeod and Michiru Nagatsu. What does interdisciplinarity look like in practice: Mapping interdisciplinarity and its limits in the environmental sciences. Studies in History and Philosophy of Science Part A, 67(??):74–84, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301863.

Marcoci:2020:JAS

[MN20] Alexandru Marcoci and James Nguyen. Judgement aggregation in scientific collaborations: the case for waiving expertise. Studies in History and Philosophy of Science Part A, 84(??):66–74, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301588.

Moked:1971:NBC

[Mok71] Gabriel Moked. A note on Berkeley’s corpuscularian theories in Siris. Studies in History and Philosophy of Science Part A, 2 (3):257–271, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368171900434.

Molland:1982:AMF

[Mol82] A. G. Molland. The atomisation of motion: a facet of the scientific revolution. Studies in History and Philosophy of Science Part A, 13(1):31–54, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368182900036.

Morus:1990:BRBa

[Mor90a] Iwan Rhys Morus. Book review: Energy & Empire: A biographical study of Lord Kelvin: Smith, C. and Wise, M. N., (Cambridge: Cambridge University Press, 1989), xxv + 866 pp., hardback £60.00. Studies in History and Philosophy of Science Part A, 21(3):519–525, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900106. See [SW89].

Morus:1990:BRBb

[Mor90b] Iwan Rhys Morus. Book review: James Joule: A biography: Cardwell, D. S. L., (Manchester: Manchester University Press, 1989), ix + 333 pp., hardback £35.00. Studies in History and Philosophy of Science Part A, 21(3):519–525, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819090009W. See [Car89].

Morus:1991:CCW

[Mor91] Iwan Rhys Morus. Correlation and control: William Robert Grove and the construction of a new philosophy of scientific reform. Studies in History and Philosophy of Science Part A, 22(4):589– 621, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190035Q.

Morrison:1992:STU

[Mor92] Margaret Morrison. A study in theory unification: The case of Maxwell’s electromagnetic theory. Studies in History and Philosophy of Science Part A, 23(1):103–145, March 1992. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368192900285.

Morris:1993:OWE

[Mor93] Andrew R. Morris. Oscar Wilde and the eclipse of Darwinism aestheticism, degeneration, and moral reaction in late-Victorian ideology. Studies in History and Philosophy of Science Part A, 24 (4):513–540, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390051K.

Morrison:1997:WUP

[Mor97] Margaret Morrison. Whewell on the ultimate problem of philosophy. Studies in History and Philosophy of Science Part A, 28 (3):417–437, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000283.

Moran:2000:ACH

[Mor00] Bruce T. Moran. Alchemy, chemistry and the history of science. Studies in History and Philosophy of Science Part A, 31(4):711– 720, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000327.

Morton:2001:LAP

[Mor01] Adam Morton. Lore-abiding people. Studies in History and Philosophy of Science Part A, 32(3):601–606, September 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000140.

Moran:2005:KHK

[Mor05] Bruce T. Moran. Knowing how and knowing that: artisans, bodies, and natural knowledge in the Scientific Revolution. Studies in History and Philosophy of Science Part A, 36(3):577–585, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810500052X.

Morton:2006:FC

[Mor06] Adam Morton. Finding the corkscrew. Studies in History and Philosophy of Science Part A, 37(1):114–117, March 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105001159.

Morus:2007:WNC

[Mor07] Iwan Rhys Morus. Working out in the nineteenth century. Studies in History and Philosophy of Science Part A, 38(3):605–609, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000441.

Morrison:2011:OPM

[Mor11] Margaret Morrison. One phenomenon, many models: Inconsistency and complementarity. Studies in History and Philosophy of Science Part A, 42(2):342–351, June 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001214.

Mormann:2012:PPD

[Mor12] Thomas Mormann. A place for pragmatism in the dynamics of reason? Studies in History and Philosophy of Science Part A, 43(1): 27–37, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000859.

Morgan:2017:NOE

[Mor17] Mary S. Morgan. Narrative ordering and explanation. Studies in History and Philosophy of Science Part A, 62(??):86– 97, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300730.

Morgan:2019:EUV

[Mor19] Mary S. Morgan. Exemplification and the use-values of cases and case studies. Studies in History and Philosophy of Science Part A, 78(??):5–13, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300153.

Morris:2021:EEJ

[Mor21] Andrew M. A. Morris. English engineer John Smeaton’s experimental method(s): Optimisation, hypothesis testing and exploratory experimentation. Studies in History and Philosophy of Science Part A, 89(??):283–294, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001059.

Moriarty:2022:RFM

[Mor22] Clare Marie Moriarty. Ructions over fluxions: Maclaurin’s draft, The Analyst controversy and Berkeley’s anti-mathematical philosophy. Studies in History and Philosophy of Science Part A, 96 (??):77–86, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001236.

Mosley:2001:JDV

[Mos01] Adam Mosley. John Donne’s verdict on Tycho Brahe: No astronomer is an island? Studies in History and Philosophy of Science Part A, 32(3):583–600, September 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000152.

Mosley:2007:OTI

[Mos07] Adam Mosley. Objects, texts and images in the history of science. Studies in History and Philosophy of Science Part A, 38(2): 289–302, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000143.

Mossner:2011:TSPa

[Möß11a] Nicola Mößner. Thought styles and paradigms — a comparative study of Ludwik Fleck and Thomas S. Kuhn. Studies in History and Philosophy of Science Part A, 42(2):362–371, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001238.

Mossner:2011:TSPb

[Möß11b] Nicola Mößner. Thought styles and paradigms: a comparative study of Ludwik Fleck and Thomas S. Kuhn. Studies in History and Philosophy of Science Part A, 42(3):416–425, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811100015X.

Motterlini:2002:RLR

[Mot02] Matteo Motterlini. Reconstructing Lakatos: a reassessment of Lakatos’ epistemological project in the light of the Lakatos archive. Studies in History and Philosophy of Science Part A, 33(3):487– 509, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000249.

Mourelatos:1981:AKP

[Mou81] Alexander P. D. Mourelatos. Astronomy and kinematics in Plato’s project of rationalist explanation. Studies in History and Philosophy of Science Part A, 12(1):1–32, March 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368181900029.

Moulin:2020:BNO

[Mou20] Isabelle Moulin. Beauty as natural order. the legacy of antiquity to Bonaventure’s symbolical theology and Nicholas of Cusa’s spiritual theophany. Studies in History and Philosophy of Science Part A, 81(??):32–38, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118300682.

Mowry:1985:GTW

[Mow85] Bryan Mowry. From Galen’s theory to William Harvey’s theory: a case study in the rationality of scientific theory change. Studies in History and Philosophy of Science Part A, 16(1):49– 82, March 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818590007X.

Moyer:1977:EDH

[Moy77] Donald Franklin Moyer. Energy, dynamics, hidden machinery: Rankine, Thomson and Tait, Maxwell. Studies in History and Philosophy of Science Part A, 8(3):251–268, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900085.

Moyer:1978:CMF

[Moy78] Donald Franklin Moyer. Continuum mechanics and field theory: Thomson and Maxwell. Studies in History and Philosophy of Science Part A, 9(1):35–50, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368178900201.

Moyer:1991:PWBa

[Moy91a] Albert E. Moyer. P. W. Bridgman’s operational perspective on physics. Part I: Origins and development. Studies in History and Philosophy of Science Part A, 22(2):237–258, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190026O.

Moyer:1991:PWBb

[Moy91b] Albert E. Moyer. P. W. Bridgman’s operational perspective on physics. Part II: Refinements, publication, and reception. Studies in History and Philosophy of Science Part A, 22(3):373–397, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368191900019.

MacLeod:2013:COT

[MR13] Christine MacLeod and Gregory Radick. Claiming ownership in the technosciences: Patents, priority and productivity. Studies in History and Philosophy of Science Part A, 44(2):188– 201, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001100.

Marchionni:2019:WME

[MR19] Caterina Marchionni and Samuli Reijula. What is mechanistic evidence, and why do we need it for evidence-based policy? Studies in History and Philosophy of Science Part A, 73(??):54– 63, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118300311.

Madden:1972:PPV

[MS72] Edward H. Madden and Mendel Sachs. Parmenidean particulars and vanishing elements. Studies in History and Philosophy of Science Part A, 3(2):151–166, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900210.

Miyata-Sturm:2024:ACD

[MS24] Mariona E. Miyata-Sturm. Aesthetic considerations in the development of plate tectonics. Studies in History and Philosophy of Science Part A, 108(??):1–9, December 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000463.

Martens:2022:IDM

[MSS^{+}22] Niels C. M. Martens, Miguel Ángel Carretero Sahuquillo, Erhard Scholz, Dennis Lehmkuhl, and Michael Krämer. Integrating dark matter, modified gravity, and the humanities. Studies in History and Philosophy of Science Part A, 91(??):A1– A5, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812100131X.

Matthews:2022:TLM

[MT22] Lucas J. Matthews and Eric Turkheimer. Three legs of the missing heritability problem. Studies in History and Philosophy of Science Part A, 93(??):183–191, June 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000632.

Mitchell:2017:MME

[MTC17] Daniel Jon Mitchell, Eran Tal, and Hasok Chang. The making of measurement: Editors’ introduction. Studies in History and Philosophy of Science Part A, 65–66(??):1–7, October/ December 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117302339.

Muka:2022:THS

[Muk22] Samantha Muka. Taking hobbyists seriously: the reef tank hobby and knowledge production in serious leisure. Studies in History and Philosophy of Science Part A, 93(??):192–202, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000619.

Muller:1994:BRB

[Mul94] F. A. Muller. Book review: Philosophy of physics: Lawrence Sklar, (Oxford: Oxford University Press, 1992), xi + 246 pp. ISBN 0-19-875138-9. Pbk. £11.95. Studies in History and Philosophy of Science Part A, 25(3):505–509, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900647.

See [Skl92].

Muller:2012:DSF

[Mül12] Alexander Müller. Dialogic structures and forms of knowledge in Plutarch’s ‘The E at Delphi’. Studies in History and Philosophy of Science Part A, 43(2):245–249, June 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001336.

Muller:2014:SI

[Mul14] F. A. Muller. The slaying of the iMongers. Studies in History and Philosophy of Science Part A, 48 (part A)(??):52–55, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001002.

Mulder:2024:CSD

[Mul24] Ruward Mulder. The classical stance: Dennett’s criterion in Wallacian quantum mechanics. Studies in History and Philosophy of Science Part A, 107(??):11–24, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001055.

Mumford:2006:FSC

[Mum06] Stephen Mumford. Function, structure, capacity. Studies in History and Philosophy of Science Part A, 37(1):76–80, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105001111.

Munson:1974:BRB

[Mun74] Ronald Munson. Book review: The philosophy of biology: Michael Ruse, London: Hutchinson University Library, 1973. 231 pp. £3.00. Studies in History and Philosophy of Science Part A, 5 (1):73–85, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936817490020X. See [Rus73].

Munevar:2016:HAP

[Mun16] Gonzalo Munévar. Historical antecedents to the philosophy of Paul Feyerabend. Studies in History and Philosophy of Science Part A, 57(??):9–16, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001624.

Murphy:1989:ALN

[Mur89] Nancey Murphy. Another look at novel facts. Studies in History and Philosophy of Science Part A, 20(3):385–388, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900149.

Murphy:1993:PFH

[Mur93] Nancey Murphy. Philosophical fractals: Or, history as metaphilosophy. Studies in History and Philosophy of Science Part A, 24(3):501–508, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390041H.

Morgan:2017:NSN

[MW17] Mary S. Morgan and M. Norton Wise. Narrative science and narrative knowing. Introduction to special issue on narrative science. Studies in History and Philosophy of Science Part A, 62 (??):1–5, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300729.

Muller-Wille:2025:SSH

[MW25] Staffan Müller-Wille. Split and Splice by Hans-Jörg Rheinberger: a natural history of experimentation. Studies in History and Philosophy of Science Part A, 109(??):120–122, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001626. See reply [Rhe25].

Myers:1989:BRB

[Mye89] Greg Myers. Book review: The figural and the literal: Problems of language in the history of science and philosophy: Andrew E. Benjamin, Geoffrey N. Cantor and John R. R. Christie, editors, (Manchester University Press: Manchester, 1987), 229 pp., Cloth £27.50. Studies in History and Philosophy of Science Part A, 20 (2):271–284, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368189900071. See [BCC87].

Mykhalovskiy:2022:PPH

[Myk22] Eric Mykhalovskiy. Book review: Philosophy of Population Health: Philosophy for a New Public Health Era by Sean Valles: Critique and philosophy of population health from the position of service. Studies in History and Philosophy of Science Part A, 91(??):199– 200, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121002077.

Nappo:2021:DNM

[Nap21] Francesco Nappo. The double nature of Maxwell’s physical analogies. Studies in History and Philosophy of Science Part A, 89 (??):212–225, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001199.

Nasr:1968:SCI

[Nas68] Seyyed Hossein Nasr. Science and civilization in Islam. Harvard University Press, Cambridge, MA, USA, 1968. xiv + 384 pp. LCCN Q125 .N17. With a preface by Giorgio de Santillana.

Nasr:1983:SCI

[Nas83] Seyyed Hossein Nasr. Science and civilization in Islam. Suhail Academy, Lahore, Pakistan, 1983. ISBN ???? xiv + 384 pp. LCCN Q124.97 .N37 1983. Rs130.00. With a preface by Giorgio de Santillana.

Nasim:2011:LGS

[Nas11] Omar W. Nasim. The ‘landmark’ and ‘groundwork’ of stars: John Herschel, photography and the drawing of nebulae. Studies in History and Philosophy of Science Part A, 42(1):67–84, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000981.

Nassar:2016:ARS

[Nas16] Dalia Nassar. Analogical reflection as a source for the science of life: Kant and the possibility of the biological sciences. Studies in History and Philosophy of Science Part A, 58(??):57– 66, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116000273.

Newton:1999:PMP

[NCW99] Sir Isaac Newton, I. Bernard Cohen, and Anne Miller Whitman. The Principia: mathematical principles of natural philosophy. University of California Press, Berkeley, CA, USA, 1999. ISBN 0-520-08816-6, 0-520-08817-4. xvii + 966 pp. LCCN QA803

.N413 1999. URL http://www.loc.gov/catdir/bios/ucal052/
99010278.html;   http://www.loc.gov/catdir/description/
ucal041/99010278.html.

Novick:2021:SS

[ND21] Aaron Novick and W. Ford Doolittle. ‘Species’ without species. Studies in History and Philosophy of Science Part A, 87(??): 72–80, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000376.

Nemec:2022:ERB

[ND22] Birgit Nemec and Heather Dron. The environments of reproductive and birth defects research in the U.S. and West Germany (c. 1955–1975). Studies in History and Philosophy of Science Part A, 95(??):50–63, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812200098X.

Neal:2024:TVE

[Nea24] Jacob P. Neal. Theory vs. experiment: the rise of the dynamic view of proteins. Studies in History and Philosophy of Science Part A, 106(??):86–98, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000578.

Needham:1996:ACP

[Nee96] Paul Needham. Aristotelian chemistry: a prelude to Duhemian metaphysics. Studies in History and Philosophy of Science Part A, 27(2):251–269, June 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819500047X.

Needham:1998:DP

[Nee98] Paul Needham. Duhem’s physicalism. Studies in History and Philosophy of Science Part A, 29(1):33–62, March 1998. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368198000144.

Needham:2002:DTM

[Nee02] Paul Needham. Duhem’s theory of mixture in the light of the Stoic challenge to the Aristotelian conception. Studies in History and Philosophy of Science Part A, 33(4):685–708, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000237.

Needham:2008:WMB

[Nee08] Paul Needham. Is water a mixure? bridging the distinction between physical and chemical properties. Studies in History and Philosophy of Science Part A, 39(1):66–77, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000878.

Needham:2013:HBH

[Nee13] Paul Needham. Hydrogen bonding: Homing in on a tricky chemical concept. Studies in History and Philosophy of Science Part A, 44(1):51–65, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000349.

Needham:2014:SCB

[Nee14] Paul Needham. The source of chemical bonding. Studies in History and Philosophy of Science Part A, 45(??):1–13, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113001209.

Nelson:1994:HCS

[Nel94] Alan Nelson. How could scientific facts be socially constructed?: Introduction: The dispute between constructivists and rationalists. Studies in History and Philosophy of Science Part A, 25(4): 535–547, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900469.

Nersessian:1984:ACS

[Ner84] Nancy J. Nersessian. Aether/or: The creation of scientific concepts. Studies in History and Philosophy of Science Part A, 15 (3):175–212, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368184900165.

Neto:1991:FS

[Net91] José R. Maia Neto. Feyerabend’s scepticism. Studies in History and Philosophy of Science Part A, 22(4):543–555, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190033O.

Neto:1993:FAS

[Net93] JoséR. Maia Neto. Feyerabend on the authority of science. Studies in History and Philosophy of Science Part A, 24(4):687– 694, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390059S.

Neugebauer:1971:BRB

[Neu71] O. Neugebauer. Book review: Babylonian algebra: Form VS. content: Vorlesungen über Geschichte der antiken mathematischen Wissenschaften, Band I: Vorgriechische Mathematik. Second, unrevised printing, Springer-Verlag Berlin–Heidelberg–New York, 1969. (First published, 1934.) xii + 212 pp. US$13.20. Studies in History and Philosophy of Science Part A, 1(4):369– 380, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900148.

Newton:1687:PNP

[New87] Isaac Newton. Philosophie naturalis principia mathematica. (Latin) [Mathematical principles of natural philosophy]. Societatis regiae, London, UK, 1687. 510 pp. LCCN ????

Newton:2009:PNP

[New09] Isaac Newton. Philosophie naturalis principia mathematica. (Latin) [Mathematical principles of natural philosophy]. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 2009. 316 pp.

LCCN ???? URL http://www.gutenberg.org/etext/28233.

Newman:2010:HIH

[New10] William R. Newman. How not to integrate the history and philosophy of science: a reply to Chalmers. Studies in History and Philosophy of Science Part A, 41(2):203–213, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000142. See [Cha10c].

Nola:2003:ITL

[NI03] Robert Nola and Gürol Irzik. Incredulity towards Lyotard: a critique of a postmodernist account of science and knowledge. Studies in History and Philosophy of Science Part A, 34(2):391– 421, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000244.

Nicholas:1972:BRB

[Nic72] John M. Nicholas. Book review: The logic of empirical theories: Marian Przelecki, London: Routledge & Kegan Paul, 1969. v + 108 pp. £1.87. Studies in History and Philosophy of Science Part A, 3(2):194–195, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368172900258. See [Prz69].

Niiniluoto:2011:ATO

[Nii11] Ilkka Niiniluoto. Abduction, tomography, and other inverse problems. Studies in History and Philosophy of Science Part A, 42(1): 135–139, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811000107X.

Niiniluoto:2014:SPI

[Nii14a] Ilkka Niiniluoto. Scientific progress as increasing verisimilitude. Studies in History and Philosophy of Science Part A, 46(??):73–77, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000144. See response [Row15].

Niiniluoto:2014:VDS

[Nii14b] Ilkka Niiniluoto. Values in design sciences. Studies in History and Philosophy of Science Part A, 46(??):11–15, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113001246.

Nikopoulos:2019:ERW

[Nik19] Dr James Nikopoulos. Essay review: Why can’t science be more like history: a response to Ruth Leys’ The Ascent of Affect. Genealogy and Critique. Studies in History and Philosophy of Science Part A, 75(??):57–61, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368118300827.

Niro:2023:CNE

[Nir23] Leonardo Niro. The conservation of nervous energy: Neurophysiology and energy conservation in the work of Sigmund Exner and Josef Breuer. Studies in History and Philosophy of Science Part A, 102(??):1–11, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123001036.

Nissen:1977:WFS

[Nis77] Lowell Nissen. Wimsatt on function statements. Studies in History and Philosophy of Science Part A, 8(4):341–347, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900243.

Nissen:1993:FWE

[Nis93] Lowell Nissen. Four ways of eliminating mind from teleology. Studies in History and Philosophy of Science Part A, 24(1):27– 48, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390023D.

Noakes:2008:WIL

[Noa08] Richard Noakes. The ‘world of the infinitely little’: connecting physical and psychical realities circa 1900. Studies in History and Philosophy of Science Part A, 39(3):323–334, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000575.

Noble:2023:ARF

[Nob23] Christopher P. Noble. Automata, reason, and free will: Leibniz’s critique of Descartes on animal and human nature. Studies in History and Philosophy of Science Part A, 100(??):56– 63, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000912.

Nonnoi:1994:AED

[Non94] Giancarlo Nonnoi. Against emptiness: Descartes’s physics and metaphysics of plenitude. Studies in History and Philosophy of Science Part A, 25(1):81–96, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900213.

Norton:1985:WWE

[Nor85] John Norton. What was Einstein’s Principle of Equivalence? Studies in History and Philosophy of Science Part A, 16(3):203–246, September 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368185900020.

Nordmann:1986:CIT

[Nor86] Alfred Nordmann. Comparing incommensurable theories. Studies in History and Philosophy of Science Part A, 17(2):231– 246, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900270.

Norman:1998:SSS

[Nor98] Andrew Norman. Seeing, semantics and social epistemic practice. Studies in History and Philosophy of Science Part A, 29(4):501– 513, December 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000296.

Nordmann:2000:HHS

[Nor00] Alfred Nordmann. Heinrich Hertz: Scientific biography and experimental life. Studies in History and Philosophy of Science Part A, 31 (3):537–549, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000145.

Nordmann:2012:APW

[Nor12] Alfred Nordmann. Another parting of the ways: Intersubjectivity and the objectivity of science. Studies in History and Philosophy of Science Part A, 43(1):38–46, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000860.

Northcott:2020:BDP

[Nor20a] Robert Northcott. Big data and prediction: Four case studies. Studies in History and Philosophy of Science Part A, 81(??):96– 104, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119300652.

Norton:2020:HBI

[Nor20b] John D. Norton. How NOT to build an infinite lottery machine. Studies in History and Philosophy of Science Part A, 82(??):1– 8, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811830325X.

Norton:2021:AR

[Nor21a] John D. Norton. Author’s responses. Studies in History and Philosophy of Science Part A, 85(??):114–126, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301837.

Norton:2021:SSE

[Nor21b] Joshua Norton. Suppressing spacetime emergence. Studies in History and Philosophy of Science Part A, 88(??):50–59, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000534.

Norton:2022:LBA

[Nor22] John D. Norton. Lotteries, bookmaking and ancient randomizers: Local and global analyses of chance. Studies in History and Philosophy of Science Part A, 95(??):108–117, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001182.

Niaz:2004:AMC

[NRB04] Mansoor Niaz, María A. Rodríguez, and Angmary Brito. An appraisal of Mendeleev’s contribution to the development of the periodic table. Studies in History and Philosophy of Science Part A, 35(2):271–282, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000159.

Nunan:1984:NFB

[Nun84] Richard Nunan. Novel facts, Bayesian rationality, and the history of continental drift. Studies in History and Philosophy of Science Part A, 15(4):267–307, December 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818490013X.

Nye:1976:NCA

[Nye76] Mary Jo Nye. The nineteenth-century atomic debates and the dilemma of an ‘indifferent hypothesis’. Studies in History and Philosophy of Science Part A, 7(3):245–268, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368176900078.

Oberheim:2005:HOC

[Obe05] Eric Oberheim. On the historical origins of the contemporary notion of incommensurability: Paul Feyerabend’s assault on conceptual conservativism. Studies in History and Philosophy of Science Part A, 36(2):363–390, June 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000208.

Oberdan:2015:HSE

[Obe15] Thomas Oberdan. From Helmholtz to Schlick: the evolution of the sign-theory of perception. Studies in History and Philosophy of Science Part A, 52(??):35–43, August 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811500045X.

Oberheim:2016:REL

[Obe16] Eric Oberheim. Rediscovering Einstein’s legacy: How Einstein anticipates Kuhn and Feyerabend on the nature of science. Studies in History and Philosophy of Science Part A, 57(??):17– 26, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811500165X.

OConnor:2005:PES

[O’C05] Ralph O’Connor. The poetics of earth science: ‘romanticism’ and the two cultures. Studies in History and Philosophy of Science Part A, 36(3):607–617, September 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000543.

OConnor:2019:NSC

[O’C19] Cailin O’Connor. The natural selection of conservative science. Studies in History and Philosophy of Science Part A, 76(??):24– 29, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117303205.

Odenbaugh:2022:BRL

[Ode22] Jay Odenbaugh. Book review: Leveraging Distortions: Explanation, Idealization, and Universality in Science by Collin Rice: a Defense of the “Standard View”. Studies in History and Philosophy of Science Part A, 95(??):224–225, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000851. See [Jhu22, Elg22, Pin22] and

reply [Ric22].

OFlaherty:2010:RSWa

[O’F10a] Niall O’Flaherty. The rhetorical strategy of William Paley’s Natural theology (1802): Part 1, William Paley’s Natural Theology in context. Studies in History and Philosophy of Science Part A, 41(1): 19–25, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000909. See [Pal02, PEK06].

OFlaherty:2010:RSWb

[O’F10b] Niall O’Flaherty. The rhetorical strategy of William Paley’s Natural theology (1802): Part 2, William Paley’s natural theology and the challenge of atheism. Studies in History and Philosophy of Science Part A, 41(2):128–137, June 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368110000166. See [Pal02, PEK06].

Oftedal:2022:PSN

[Oft22] Gry Oftedal. Proportionality of single nucleotide causation. Studies in History and Philosophy of Science Part A, 93(??):215– 222, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000644.

OGorman:1977:PCA

[O’G77] F. P. O’Gorman. Poincaré’s conventionalism of applied geometry. Studies in History and Philosophy of Science Part A, 8(4): 303–340, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900231.

OGorman:1978:PRE

[O’G78] F. P. O’Gorman. Poincaré’s retention of Euclid on apparently adverse parallactic findings: a reply to A. Grünbaum. Studies in History and Philosophy of Science Part A, 9(4):319–321, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368178900122. See [Grü78].

Oberheim:2000:FEP

[OHH00] Eric Oberheim and Paul Hoyningen-Huene. Feyerabend’s early philosophy. Studies in History and Philosophy of Science Part A, 31 (2):363–375, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000078.

Ohnesorge:2021:HIM

[Ohn21] Miguel Ohnesorge. How incoherent measurement succeeds: Coordination and success in the measurement of the Earth’s polar flattening. Studies in History and Philosophy of Science Part A, 88(??):245–262, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000893.

Ohnesorge:2022:PMP

[Ohn22] Miguel Ohnesorge. Pluralizing measurement: Physical geodesy’s measurement problem and its resolution. Studies in History and Philosophy of Science Part A, 96(??):51–67, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812200125X.

Okasha:2000:VFC

[Oka00] Samir Okasha. Van Fraassen’s critique of inference to the best explanation. Studies in History and Philosophy of Science Part A, 31(4):691–710, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000169.

OLoughlin:2020:SOC

[O’L20] Ryan O’Loughlin. Seepage, objectivity, and climate science. Studies in History and Philosophy of Science Part A, 81(??): 74–81, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119300871.

OLoughlin:2021:RRC

[O’L21] Ryan O’Loughlin. Robustness reasoning in climate model comparisons. Studies in History and Philosophy of Science Part A, 85(??):34–43, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120302065.

Olin:1992:WJP

[Oli92] Doris Olin, editor. William James: Pragmatism, in focus. Routledge philosophers in focus series. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1992. ISBN 0-415-04056-6, 0-415-04057-4 (paperback). viii + 251 pp. LCCN B945.J23

P7339 1992. URL http://www.loc.gov/catdir/enhancements/
fy0909/91035744-d.html.

Olmos:2012:TLE

[Olm12] Paula Olmos. Two literary encyclopaedias from Late Antiquity. Studies in History and Philosophy of Science Part A, 43(2):284– 292, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001294.

Olson:2016:KAS

[Ols16] Michael J. Olson. Kant on anatomy and the status of the life sciences. Studies in History and Philosophy of Science Part A, 58(??):77–84, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811600025X.

OMalley:2016:HMR

[O’M16] Maureen A. O’Malley. Histories of molecules: Reconciling the past. Studies in History and Philosophy of Science Part A, 55(??):69– 83, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001314.

ONeill:1986:FHC

[O’N86] John O’Neill. Formalism, Hamilton and complex numbers. Studies in History and Philosophy of Science Part A, 17(3):351–372, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900130.

ONeill:1998:PRM

[O’N98] John O’Neill. Practical reason and mathematical argument. Studies in History and Philosophy of Science Part A, 29(2):195– 205, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000090.

ONeill:2003:USP

[O’N03] John O’Neill. Unified science as political philosophy: positivism, pluralism and liberalism. Studies in History and Philosophy of Science Part A, 34(3):575–596, September 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000487.

ONeill:2008:HTH

[O’N08] Onora O’Neill. Historical trends and human futures. Studies in History and Philosophy of Science Part A, 39(4):529–534, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000873.

Onishi:2017:DSC

[Oni17] Yukinori Onishi. Defending the selective confirmation strategy. Studies in History and Philosophy of Science Part A, 64(??):1– 10, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301796.

Oosterhoff:2012:EMM

[Oos12] Richard J. Oosterhoff. Early modern mathematical practice in the round. Studies in History and Philosophy of Science Part A, 43(1): 224–227, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001087.

OMalley:2018:MMM

[OP18] Maureen A. O’Malley and Emily C. Parke. Microbes, mathematics, and models. Studies in History and Philosophy of Science Part A, 72(??):1–10, December 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301528.

Ordorica:2016:ERA

[Ord16] Sergio Armando Gallegos Ordorica. The explanatory role of abstraction processes in models: the case of aggregations. Studies in History and Philosophy of Science Part A, 56(??):161– 167, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001405.

Ortolano:2008:LST

[Ort08] Guy Ortolano. The literature and the science of ‘two cultures’ historiography. Studies in History and Philosophy of Science Part A, 39(1):143–150, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000945.

Okon:2014:MAC

[OS14] Elias Okon and Daniel Sudarsky. Measurements according to Consistent Histories. Studies in History and Philosophy of Science Part A, 48 (part A)(??):7–12, November 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000951.

Osborne:1989:POU

[Osb89] Catherine Osborne. Philoponus on the origins of the universe and other issues. Studies in History and Philosophy of Science Part A, 20(3):389–395, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900150.

Osen:2022:MAG

[Ose22] Elana Osen. Marinus of Alexandria: Galen’s anatomical forefather, or: How do you solve a problem like Marinus? Studies in History and Philosophy of Science Part A, 92(??):224– 238, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301850.

Osler:1971:BRB

[Osl71] Margaret J. Osler. Book review: Francis Bacon and Dénis Diderot: philosophers of science: Lilo K. Luxembourg, Munksgaard: Copenhagen, 1967. 127 pp. $6.00. Studies in History and Philosophy of Science Part A, 2(1):91–95, May 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900392.

See [Lux67].

Osler:1972:BRB

[Osl72] Margaret J. Osler. Book review: Locke and the compass of human understanding. A selective commentary on the ‘essay’: John W. Yolton, Cambridge University Press, 1970. x + 234 pp. $10.00. Studies in History and Philosophy of Science Part A, 3(2):189– 194, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900246. See [Yol70].

Osler:1992:DNP

[Osl92] Margaret J. Osler. Descartes, natural philosopher. Studies in History and Philosophy of Science Part A, 23(3):509–518, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290008T.

Ott:2008:RSM

[Ott08] Walter Ott. Régis’s scholastic mechanism. Studies in History and Philosophy of Science Part A, 39(1):2–14, March 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000830.

Outram:1996:PBH

[Out96] Dorinda Outram. Professor Branestawm and his friends. Studies in History and Philosophy of Science Part A, 27(1):109– 114, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000291.

Pacchetti:2018:RSS

[Pac18] Marina Baldissera Pacchetti. A role for spatiotemporal scales in modeling. Studies in History and Philosophy of Science Part A, 67(??):14–21, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116301327.

Padovani:2015:RCS

[Pad15] Flavia Padovani. Reichenbach on causality in 1923: Scientific inference, coordination, and confirmation. Studies in History and Philosophy of Science Part A, 53(??):3–11, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000643.

Paley:1802:NTE

[Pal02] William Paley. Natural theology, or, Evidences of the existence and attributes of the Deity. R. Faulder, London, UK, 1802. xii + 586 pp. LCCN ????

Palter:1970:AHE

[Pal70] Robert Palter. An approach to the history of early astronomy. Studies in History and Philosophy of Science Part A, 1(2):93– 133, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368170900014.

Palter:1971:AMS

[Pal71] Robert Palter, editor. The Annus mirabilis of Sir Isaac Newton, 1666–1966. MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-16035-8. LCCN QC16.N7 A57.

Palmquist:1987:KCR

[Pal87a] Stephen Palmquist. Kant’s cosmogony re-evaluated. Studies in History and Philosophy of Science Part A, 18(3):255–269, September 1987. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368187900215.

Palter:1987:SNT

[Pal87b] Robert Palter. Saving Newton’s text: Documents, readers, and the ways of the world. Studies in History and Philosophy of Science Part A, 18(4):385–439, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936818790001X.

Palmieri:2003:MMG

[Pal03] Paolo Palmieri. Mental models in Galileo’s early mathematization of nature. Studies in History and Philosophy of Science Part A, 34(2):229–264, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000256.

Palmieri:2005:SSP

[Pal05] Paolo Palmieri. ‘spuntar lo scoglio più duro’: did Galileo ever think the most beautiful thought experiment in the history of science? Studies in History and Philosophy of Science Part A, 36(2):

223–240, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000154.

Palmieri:2008:EBE

[Pal08a] Paolo Palmieri. The empirical basis of equilibrium: Mach, Vailati, and the lever. Studies in History and Philosophy of Science Part A, 39(1):42–53, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000854.

Palmieri:2008:MOR

[Pal08b] Paolo Palmieri. Mechanical objects, represented and real. Studies in History and Philosophy of Science Part A, 39(1):154– 159, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000969.

Palmieri:2009:RMT

[Pal09a] Paolo Palmieri. Radical mathematical Thomism: beings of reason and divine decrees in Torricelli’s philosophy of mathematics. Studies in History and Philosophy of Science Part A, 40(2): 131–142, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000235.

Palmieri:2009:RMV

[Pal09b] Paolo Palmieri. Response to Maarten Van Dyck’s commentary. Studies in History and Philosophy of Science Part A, 40(3):319– 321, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000491. See [Van09].

Palacios:2023:BRM

[Pal23] Patricia Palacios. Book review: A Middle Way: a Non-Fundamental Approach to Many-Body Physics by Robert Batterman: Reductionism and the Autonomy of Scales. Studies in History and Philosophy of Science Part A, 97(??):128–129, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001303. See [Bat23].

Pantin:1996:CWR

[Pan96] Isabelle Pantin. Is Clavius worth reappraising? The impact of a Jesuit mathematical teacher on the eve of the astronomical revolution. Studies in History and Philosophy of Science Part A, 27 (4):593–598, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000542.

Pantin:1999:NPO

[Pan99] Isabelle Pantin. New philosophy and old prejudices: Aspects of the reception of Copernicanism in a divided Europe. Studies in History and Philosophy of Science Part A, 30(2):237– 262, June 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000491.

Panza:2006:FVB

[Pan06] Marco Panza. François Viète: between analysis and cryptanalysis. Studies in History and Philosophy of Science Part A, 37(2): 269–289, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000173.

Paoletti:2011:CPE

[Pao11] Cristina Paoletti. Causes as proximate events: Thomas Brown and the Positivist interpretation of Hume on causality. Studies in History and Philosophy of Science Part A, 42(1):37–44, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110001081.

Papineau:1977:VVC

[Pap77] David Papineau. The vis viva controversy: Do meanings matter? Studies in History and Philosophy of Science Part A, 8(2): 111–142, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900115.

Papineau:1989:UAR

[Pap89] David Papineau. An unnatural anti-realism. Studies in History and Philosophy of Science Part A, 20(1):133–138, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818990037X.

Papineau:2010:RRS

[Pap10] David Papineau. Realism, Ramsey sentences and the pessimistic meta-induction. Studies in History and Philosophy of Science Part A, 41(4):375–385, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000579.

Papayannopoulos:2020:CMA

[Pap20] Philippos Papayannopoulos. Computing and modelling: Analog vs. analogue. Studies in History and Philosophy of Science Part A, 83(??):103–120, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301177.

Papale:2024:PMP

[Pap24] François Papale. From the philosophy of measurement to the philosophy of classification: Generalizing the problem of coordination and historical coherentism. Studies in History and Philosophy of Science Part A, 106(??):1–11, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000682.

Parusnikova:1992:PPS

[Par92] Zuzana Parusnikova. Is a postmodern philosophy of science possible? Studies in History and Philosophy of Science Part A, 23(1): 21–37, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368192900252.

Parascandola:1995:PLD

[Par95] Mark Parascandola. Philosophy in the laboratory: The debate over evidence for E. J. Steele’s Lamarckian hypothesis. Studies in History and Philosophy of Science Part A, 26(3):469–492, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819500012C.

Parry:2012:OPP

[Par12] Glyn Parry. Occult philosophy and politics: Why John Dee wrote his Compendious rehearsal in November 1592. Studies in History and Philosophy of Science Part A, 43(3):480–488, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001178.

Parker:2014:VUC

[Par14] Wendy Parker. Values and uncertainties in climate prediction, revisited. Studies in History and Philosophy of Science Part A, 46 (??):24–30, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113001258.

Park:2016:ESR

[Par16] Seungbae Park. Extensional scientific realism vs. intensional scientific realism. Studies in History and Philosophy of Science Part A, 59(??):46–52, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300176.

Parker:2020:NAP

[Par20a] Adwait A. Parker. Newton on active and passive quantities of matter. Studies in History and Philosophy of Science Part A, 84 (??):1–11, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120300674.

Parker:2020:CIL

[Par20b] Matthew W. Parker. Comparative infinite lottery logic. Studies in History and Philosophy of Science Part A, 84(??):28–36, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119301426.

Pascal:1662:PFT

[Pas62] Blaise Pascal. Pensées. (French) [Thoughts]. ????, ????, 1662. ???? pp.

Paseau:2005:WFF

[Pas05] Alexander Paseau. What the foundationalist filter kept out. Studies in History and Philosophy of Science Part A, 36(1):191– 201, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104001281.

Pascal:2010:POL

[Pas10] Blaise Pascal. Pensées: Opuscules et lettres. Éd. Classiques Garnier, Paris, France, 2010. 807 pp. LCCN ????

Paseau:2011:MIG

[Pas11] Alexander Paseau. Mathematical instrumentalism, Gödel’s theorem, and inductive evidence. Studies in History and Philosophy of Science Part A, 42(1):140–149, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001093.

Paskins:2020:HSU

[Pas20] Matthew Paskins. History of science and its utopian reconstructions. Studies in History and Philosophy of Science Part A, 81(??): 82–95, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118302905.

Patzig:1969:ATS

[Pat69] Günther Patzig. Aristotle’s theory of the syllogism. A logicophilological study of Book A of the Prior Analytics. Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1969. xvii + 219 pp. LCCN B440 .P313.

Patton:2009:STM

[Pat09] Lydia Patton. Signs, toy models, and the a priori: from Helmholtz to Wittgenstein. Studies in History and Philosophy of Science Part A, 40(3):281–289, September 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000521.

Pratt:1996:GAR

[PB96] Vernon Pratt and Isis Brook. Goethe’s archetype and the Romantic concept of the self. Studies in History and Philosophy of Science Part A, 27(3):351–365, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000488.

Perez-Escobar:2024:MLT

[PE24] José Antonio Pérez-Escobar. Minimal logical teleology in artifacts and biology connects the two domains and frames mechanisms via epistemic circularity. Studies in History and Philosophy of Science Part A, 104(??):23–37, April 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000104.

Pecontal:2011:PMM

[Péc11] Emmanuel Pécontal. Polar motion measurement at the Observatoire de Lyon in the late nineteenth century. Studies in History and Philosophy of Science Part A, 42(1):94–104, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000993.

Peden:2020:BEP

[Ped20] William Peden. The Bayesian Era in the philosophy of science. Studies in History and Philosophy of Science Part A, 80(??):123– 127, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368119303668.

Peels:2017:TRE

[Pee17] Rik Peels. Ten reasons to embrace scientism. Studies in History and Philosophy of Science Part A, 63(??):11–21, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300875.

Paley:2006:NTE

[PEK06] William Paley, Matthew Eddy, and David M. Knight. Natural theology: or, evidence of the existence and attributes of the deity, collected from the appearances of nature. Oxford world’s classics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-280584-3 (paperback). xxxvii + 342 pp. LCCN BL183 .P35 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0640/2005026316-d.
html;   http://www.loc.gov/catdir/enhancements/fy0726/
2005026316-b.html;   http://www.loc.gov/catdir/toc/ecip0519/

2005026316.html. Edited with an introduction and notes by Matthew D. Eddy and David Knight.

Pence:2015:EHC

[Pen15] Charles H. Pence. The early history of chance in evolution. Studies in History and Philosophy of Science Part A, 50(??): 48–58, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000843.

Pence:2022:SCG

[Pen22a] Charles H. Pence. Of stirps and chromosomes: Generality through detail. Studies in History and Philosophy of Science Part A, 94 (??):177–190, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000966.

Pence:2022:WHR

[Pen22b] Charles H. Pence. Whatever happened to reversion? Studies in History and Philosophy of Science Part A, 92(??):97– 108, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000292.

Penkler:2022:CBC

[Pen22c] Michael Penkler. Caring for biosocial complexity. articulations of the environment in research on the developmental origins of health and disease. Studies in History and Philosophy of Science Part A, 93(??):1–10, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000358.

Pereboom:1991:MEP

[Per91] Derk Pereboom. Mathematical expressibility, perceptual relativity, and secondary qualities. Studies in History and Philosophy of Science Part A, 22(1):63–88, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190015K.

Peressini:2008:CHM

[Per08] Anthony Peressini. Confirmational holism and its mathematical (w)holes. Studies in History and Philosophy of Science Part A, 39(1):102–111, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000908.

Pessian:2010:RBE

[Pes10] Arash Pessian. Reference to the best explanation. Studies in History and Philosophy of Science Part A, 41(4):363–374, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811000066X.

Pesic:2013:ERH

[Pes13] Peter Pesic. Essay review: Hermann Weyl’s neighborhood: Umgebungen: Symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus, by Norman Sieroka; Chronos Verlag, Zurich, 2010, pp. 411, Price EUR 43,00, hardback, ISBN 978-3-0340-1006-1. Studies in History and Philosophy of Science Part A, 44(1):150–153, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000751.

Peterson:2011:SWS

[Pet11] Heather R. Peterson. The shape of the world: the story of Spanish expansion and the secret science of cosmography. Studies in History and Philosophy of Science Part A, 42(1):223– 226, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000919.

Peterson:2015:BBT

[Pet15] Erik L. Peterson. The baubles of biotech, or, that’s the spirit. Studies in History and Philosophy of Science Part A, 54(??):124– 126, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000904.

Pettit:2016:DCW

[Pet16] Michael Pettit. Deflating Cold War rationality. Studies in History and Philosophy of Science Part A, 58(??):46–49, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116000029.

Pfeifer:2012:MLL

[Pfe12] Jessica Pfeifer. Mill and Lewis on laws, experimentation, and systematization. Studies in History and Philosophy of Science Part A, 43(1):172–181, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001257.

Pena-Guzman:2020:FHE

[PG20] David M. Peña-Guzmán. French historical epistemology: Discourse, concepts, and the norms of rationality. Studies in History and Philosophy of Science Part A, 79(??):68–76, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301626.

Perez-Gonzalez:2024:EME

[PG24] Saúl Pérez-González. Evidence of mechanisms in evidence-based policy. Studies in History and Philosophy of Science Part A, 103 (??):95–104, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123001656.

Phillips:2005:AMP

[Phi05] Christopher Phillips. Augustus De Morgan and the propagation of moral mathematics. Studies in History and Philosophy of Science Part A, 36(1):105–133, March 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104001220.

Pickering:1984:APP

[Pic84] Andy Pickering. Against putting the phenomena first: The discovery of the weak neutral current. Studies in History and Philosophy of Science Part A, 15(2):85–117, June 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368184900013.

Piccinini:2003:EDP

[Pic03] Gualtiero Piccinini. Epistemic divergence and the publicity of scientific methods. Studies in History and Philosophy of Science Part A, 34(3):597–612, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000499.

Piccinini:2004:FCM

[Pic04] Gualtiero Piccinini. Functionalism, computationalism, and mental states. Studies in History and Philosophy of Science Part A, 35 (4):811–833, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000883.

Picchio:2022:WRA

[Pic22] Marcos Picchio. When the “realism of assumptions” mattered: Milton Friedman’s critique of the Phillips curve. Studies in History and Philosophy of Science Part A, 94(??):8–16, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000693.

Pieters:1997:SNB

[Pie97] Toine Pieters. Shaping a new biological factor, ‘the interferon’, in room 215 of the National Institute for Medical Research, 1956/ 57. Studies in History and Philosophy of Science Part A, 28(1): 27–73, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000040.

Pietsch:2014:RTC

[Pie14] Wolfgang Pietsch. A revolution without tooth and claw — redefining the physical base units. Studies in History and Philosophy of Science Part A, 46(??):85–93, June 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000168.

Pinch:1997:OHH

[Pin97] Trevor Pinch. Old habits die hard: Retrieving practices from social theory. Studies in History and Philosophy of Science Part A, 28(1):203–208, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000534.

Pincock:2009:SSO

[Pin09] Christopher Pincock. From sunspots to the Southern Oscillation: confirming models of large-scale phenomena in meteorology. Studies in History and Philosophy of Science Part A, 40(1):

45–56, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108001118.

Pinto:2014:PSG

[Pin14] Manuela Fernández Pinto. Philosophy of science for globalized privatization: Uncovering some limitations of critical contextual empiricism. Studies in History and Philosophy of Science Part A, 47 (??):10–17, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000259.

Pinto:2016:DVT

[Pin16] Manuela Fernández Pinto. Democratic values and their role in maximizing the objectivity of science. Studies in History and Philosophy of Science Part A, 59(??):121–124, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300012.

Pincock:2022:BRL

[Pin22] Christopher Pincock. Book review: Leveraging Distortions: Explanation, Idealization, and Universality in Science by Collin Rice: the Counterfactual Account of Explanation. Studies in History and Philosophy of Science Part A, 95(??):230–232, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000887.

Pippin:1979:KEC

[Pip79] Robert B. Pippin. Kant on empirical concepts. Studies in History and Philosophy of Science Part A, 10(1):1–19, March 1979. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368179900025.

Pitts:2022:PBO

[Pit22] J. Brian Pitts. Peter Bergmann on observables in Hamiltonian General Relativity: a historical-critical investigation. Studies in History and Philosophy of Science Part A, 95(??):1–27, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000930.

Plaisance:2014:PAI

[PK14] Kathryn S. Plaisance and Eric B. Kennedy. A pluralistic approach to interactional expertise. Studies in History and Philosophy of Science Part A, 47(??):60–68, September 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000594.

Plaisance:2020:BAI

[Pla20] Kathryn S. Plaisance. The benefits of acquiring interactional expertise: Why (some) philosophers of science should engage scientific communities. Studies in History and Philosophy of Science Part A, 83(??):53–62, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119302055.

Pluhar:1978:ER

[Plu78] Evelyn B. Pluhar. Emergence and reduction. Studies in History and Philosophy of Science Part A, 9(4):279–289, December 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368178900092.

Perillat:2022:CRR

[PM22] Lucie Perillat and Mathew Mercuri. Clinical recommendations: the role of mechanisms in the GRADE framework. Studies in History and Philosophy of Science Part A, 96(??):1–9, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001327.

Perkins-McVey:2023:WSE

[PM23] Matthew Perkins-McVey. Were the scale of excitability a circle: Tracing the roots of the disease theory of alcoholism through Brunonian stimulus dependence. Studies in History and Philosophy of Science Part A, 99(??):46–55, June 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000699.

Palmer:2021:ECA

[PMG21] Alexandra Palmer, Reuben Message, and Beth Greenhough. Edge cases in animal research law: Constituting the regulatory borderlands of the UK’s Animals (Scientific Procedures) Act. Studies in History and Philosophy of Science Part A, 90(??):122–130, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001461.

Politi:2015:NKC

[Pol15] Vincenzo Politi. Natural kinds, causes and domains: Khalidi on how science classifies things. Studies in History and Philosophy of Science Part A, 54(??):132–137, December 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115001375.

Polat:2021:MRM

[Pol21] Bican Polat. Model-as-replica, model-as-instrument: Representational power and contextual versatility in animal models. Studies in History and Philosophy of Science Part A, 89(??):19– 30, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001047.

Porter:2001:VDQ

[Por01] Theodore M. Porter. On the virtues and disadvantage of quantification for democratic life. Studies in History and Philosophy of Science Part A, 32(4):739–747, December 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000218.

Portides:2011:SRP

[Por11] Demetris Portides. Seeking representations of phenomena: Phenomenological models. Studies in History and Philosophy of Science Part A, 42(2):334–341, June 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110001202.

Post:1971:CIH

[Pos71] H. R. Post. Correspondence, invariance and heuristics: In praise of conservative induction. Studies in History and Philosophy of Science Part A, 2(3):213–255, November 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900422.

Potochnik:2012:FIM

[Pot12] Angela Potochnik. Feminist implications of model-based science. Studies in History and Philosophy of Science Part A, 43(2):383– 389, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000027.

Potochnik:2015:DAS

[Pot15] Angela Potochnik. The diverse aims of science. Studies in History and Philosophy of Science Part A, 53(??):71–80, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000680.

Pourciau:2000:IFK

[Pou00] Bruce Pourciau. Intuitionism as a (failed) Kuhnian revolution in mathematics. Studies in History and Philosophy of Science Part A, 31(2):297–329, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368100000108.

Pourciau:2007:CFC

[Pou07] Bruce Pourciau. From centripetal forces to conic orbits: a path through the early sections of Newton’s Principia. Studies in History and Philosophy of Science Part A, 38(1):56–83, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106001130. See [New87, NCW99, New09].

Piso:2016:FCI

[POW16] Zachary Piso, Michael O’Rourke, and Kathleen C. Weathers. Out of the fog: Catalyzing integrative capacity in interdisciplinary research. Studies in History and Philosophy of Science Part A, 56 (??):84–94, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116000030.

Piso:2021:IVP

[PP21] Zachary Piso and Viorel Pâslaru. Introduction to values and pluralism in the environmental sciences: From inferences to institutions. Studies in History and Philosophy of Science Part A, 90 (??):140–144, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812100145X.

Parke:2023:GBG

[PP23] Emily C. Parke and Anya Plutynski. Going big by going small: Trade-offs in microbiome explanations of cancer. Studies in History and Philosophy of Science Part A, 97(??):101–110, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001844.

Perez-Ramos:1990:BRB

[PR90] A. Pérez-Ramos. Book review: Theology and the scientific imagination from the Middle Ages to the Seventeenth Century: Amos Funkenstein, (Princeton University Press: Princeton, 1986), xii + 421 pp., Cloth $49.50. Studies in History and Philosophy of Science Part A, 21(2):323–339, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819090030C. See [Fun86].

Pratt:1974:EPO

[Pra74] Vernon Pratt. Explaining the properties of organisms. Studies in History and Philosophy of Science Part A, 5(1):1–15, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368174900168.

Pratt:1975:FPG

[Pra75] Vernon Pratt. Functionalism and the possibility of group selection. Studies in History and Philosophy of Science Part A, 5(4):371– 372, February 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900096.

Pratt:1977:FHC

[Pra77] Vernon Pratt. Foucault & the history of classification theory. Studies in History and Philosophy of Science Part A, 8(2):163– 171, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368177900139.

Preston:1998:SSP

[Pre98] John Preston. Science as supermarket: ‘post-modern’ themes in Paul Feyerabend’s later philosophy of science. Studies in History and Philosophy of Science Part A, 29(3):425–447, September 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368198000156.

Preston:2004:BKP

[Pre04] John Preston. Bird, Kuhn, and positivism. Studies in History and Philosophy of Science Part A, 35(2):327–335, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000172.

Preston:2006:SCA

[Pre06] Beth Preston. Social context and artefact function. Studies in History and Philosophy of Science Part A, 37(1):37–41, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810500107X.

Preston:2007:LHM

[Pre07] John Preston. Lützen on Hertz’s mechanics. Studies in History and Philosophy of Science Part A, 38(1):260–267, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368106001038.

Preston:2008:MHM

[Pre08a] John Preston. Mach and Hertz’s mechanics. Studies in History and Philosophy of Science Part A, 39(1):91–101, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000891.

Preti:2008:OCN

[Pre08b] Consuelo Preti. On the origins of the contemporary notion of propositional content: anti-psychologism in nineteenth-century psychology and G. E. Moore’s early theory of judgment. Studies in History and Philosophy of Science Part A, 39(2):176– 185, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000198.

Preston:2016:RWR

[Pre16] John Preston. The rise of western rationalism: Paul Feyerabend’s story. Studies in History and Philosophy of Science Part A, 57(??): 79–86, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001739.

Pringe:2018:MCK

[Pri18] Hernán Pringe. Maimon’s criticism of Kant’s doctrine of mathematical cognition and the possibility of metaphysics as a science. Studies in History and Philosophy of Science Part A, 71(??): 35–44, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300092.

Price:2019:LZG

[Pri19] Justin Price. The landing zone — ground for model transfer in chemistry. Studies in History and Philosophy of Science Part A, 77(??):21–28, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118301729.

Prinz:2022:HBM

[Pri22] Benjamin Prinz. How blood met plastics, plant and animal extracts: Material encounters between medicine and industry in the twentieth century. Studies in History and Philosophy of Science Part A, 92(??):45–55, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000073.

Prudovsky:1989:CSP

[Pru89] Gad Prudovsky. The confirmation of the superposition principle: On the role of a constructive thought experiment in Galileo’s discorsi. Studies in History and Philosophy of Science Part A, 20 (4):453–468, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368189900198.

Przelecki:1969:LET

[Prz69] Marian Przelecki. The logic of empirical theories. Monographs in modern logic series. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1969. ISBN 0-7100-6230-3. v + 108 pp. LCCN BC91 .P7.

Piccinini:2010:CVI

[PS10] Gualtiero Piccinini and Andrea Scarantino. Computation vs. information processing: why their difference matters to cognitive science. Studies in History and Philosophy of Science Part A, 41 (3):237–246, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000440.

Pils:2024:SRS

[PS24] Raimund Pils and Philipp Schoenegger. Scientific realism, scientific practice, and science communication: an empirical investigation of academics and science communicators. Studies in History and Philosophy of Science Part A, 105(??):85–98, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000475.

Psillos:1994:PST

[Psi94] Stathis Psillos. A philosophical study of the transition from the caloric theory of heat to thermodynamics: Resisting the pessimistic meta-induction. Studies in History and Philosophy of Science Part A, 25(2):159–190, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900264.

Psillos:1997:NT

[Psi97] Stathis Psillos. Naturalism without truth? Studies in History and Philosophy of Science Part A, 28(4):699–713, December 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368196000313.

Psillos:2000:RCT

[Psi00] Stathis Psillos. Rudolf Carnap’s ‘Theoretical Concepts in Science’. Studies in History and Philosophy of Science Part A, 31(1):151– 172, March 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936819900031X.

Polger:2018:DIS

[PSS18] Thomas W. Polger, Lawrence A. Shapiro, and Reuben Stern. In defense of interventionist solutions to exclusion. Studies in History and Philosophy of Science Part A, 68(??):51–57, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300268.

Pascal:2003:P

[PT03] Blaise Pascal and W. F. (William Finlayson) Trotter. Pensées. Dover philosophical classics. Dover Publications, Inc., New York, NY, USA, 2003. ISBN 0-486-43255-6 (paperback). xix + 297 pp.

LCCN B1901.P42 E5 2003. URL http://www.loc.gov/catdir/
enhancements/fy0614/2003055277-d.html.

Pulkkinen:2001:RNK

[Pul01] Jarmo Pulkkinen. Russell and the neo-Kantians. Studies in History and Philosophy of Science Part A, 32(1):99–117, March 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000406.

Pumfrey:2012:JDP

[Pum12] Stephen Pumfrey. John Dee: the patronage of a natural philosopher in Tudor England. Studies in History and Philosophy of Science Part A, 43(3):449–459, September 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001142.

Peschard:2014:MAC

[PvF14] Isabelle F. Peschard and Bas C. van Fraassen. Making the abstract concrete: the role of norms and values in experimental modeling. Studies in History and Philosophy of Science Part A, 46 (??):3–10, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811300126X.

Potochnik:2006:RGA

[PY06] Angela Potochnik and Audrey Yap. Revisiting Galison’s ‘Aufbau/Bauhaus’ in light of Neurath’s philosophical projects. Studies in History and Philosophy of Science Part A, 37(3):469–488, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000197.

Pyenson:2008:FP

[Pye08] Lewis Pyenson. Forward into the past. Studies in History and Philosophy of Science Part A, 39(2):211–219, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810800023X.

Pyle:2002:BSM

[Pyl02] Andrew Pyle. Boyle on science and the mechanical philosophy: a reply to Chalmers. Studies in History and Philosophy of Science Part A, 33(1):171–186, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000358.

Peterson:2014:IVN

[PZ14] Martin Peterson and Sjoerd D. Zwart. Introduction: Values and norms in modeling. Studies in History and Philosophy of Science Part A, 46(??):1–2, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113001271.

Quay:1975:EFP

[Qua75] Paul M. Quay. The estimative functions of physical theory. Studies in History and Philosophy of Science Part A, 6(2):125– 157, June 1975. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368175900187.

Quinn:1972:MAH

[Qui72] Philip Quinn. Methodological appraisal and heuristic advice: Problems in the methodology of scientific research programmes. Studies in History and Philosophy of Science Part A, 3(2):135– 149, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900209.

Quinn:1974:SEI

[Qui74] Philip L. Quinn. Some epistemic implications of ‘crucial experiments’. Studies in History and Philosophy of Science Part A, 5 (1):59–72, May 1974. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368174900193.

Quinn:2017:AAN

[Qui17] Terry Quinn. From artefacts to atoms — a new SI for 2018 to be based on fundamental constants. Studies in History and Philosophy of Science Part A, 65–66(??):8–20, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301851.

Quintana:2018:LRE

[Qui18] Federico Raffo Quintana. Leibniz on the requisites of an exact arithmetical quadrature. Studies in History and Philosophy of Science Part A, 67(??):65–73, February 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301553.

Quinn:2021:TSC

[Qui21] Aleta Quinn. Transparency and secrecy in citizen science: Lessons from herping. Studies in History and Philosophy of Science Part A, 85(??):208–217, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301941.

Rabb:1975:IPC

[Rab75] J. Douglas Rabb. Incommensurable paradigms and critical idealism. Studies in History and Philosophy of Science Part A, 6 (4):343–346, November 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936817590028X.

Radder:1993:SRR

[Rad93] Hans Radder. Science, realization and reality: The fundamental issues. Studies in History and Philosophy of Science Part A, 24(3):327–349, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390032F.

Radder:1997:PHS

[Rad97] Hans Radder. Philosophy and history of science: Beyond the Kuhnian paradigm. Studies in History and Philosophy of Science Part A, 28(4):633–655, December 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368197000150.

Radick:2013:PPL

[Rad13] Gregory Radick. The professor and the pea: Lives and afterlives of William Bateson’s campaign for the utility of Mendelism. Studies in History and Philosophy of Science Part A, 44(2):280– 291, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001094.

Radick:2022:MFS

[Rad22] Gregory Radick. Mendel the fraud? A social history of truth in genetics. Studies in History and Philosophy of Science Part A, 93 (??):39–46, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121002120.

Raftopoulos:2003:CAS

[Raf03] Athanassios Raftopoulos. Cartesian analysis and synthesis. Studies in History and Philosophy of Science Part A, 34(2):265– 308, June 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000268.

Ralley:2006:AAA

[Ral06] Robert Ralley. Alchemical artisans, artisanal alchemy. Studies in History and Philosophy of Science Part A, 37(2):344– 352, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000306.

Ramsey:1992:REB

[Ram92] Jeffry L. Ramsey. On refusing to be an epistemologically black box: Instruments in chemical kinetics during the 1920s and ’30s. Studies in History and Philosophy of Science Part A, 23(2):283– 304, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368192900366.

Rampelt:2008:RNB

[Ram08] Jason M. Rampelt. Religion and narrative building in the history of science. Studies in History and Philosophy of Science Part A, 39(2):286–289, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000356.

Rampling:2012:JDA

[Ram12a] Jennifer M. Rampling. John Dee and the alchemists: Practising and promoting English alchemy in the Holy Roman Empire. Studies in History and Philosophy of Science Part A, 43(3):498– 508, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001208.

Rampling:2012:JDS

[Ram12b] Jennifer M. Rampling. John Dee and the sciences: early modern networks of knowledge. Studies in History and Philosophy of Science Part A, 43(3):432–436, September 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001129.

Rapp:1980:ODS

[Rap80] Friedrich Rapp. Observational data and scientific progress. Studies in History and Philosophy of Science Part A, 11(2):153– 162, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900217.

Raphael:2011:CNL

[Rap11a] Renée J. Raphael. Casting new light on Catholic censorship and early modern science. Studies in History and Philosophy of Science Part A, 42(3):453–456, September 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000227.

Raphael:2011:MSD

[Rap11b] Renée Jennifer Raphael. Making sense of Day 1 of the two new sciences: Galileo’s Aristotelian-inspired agenda and his Jesuit readers.

Studies in History and Philosophy of Science Part A, 42(4):479– 491, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000744.

Rasmussen:1993:FAM

[Ras93] Nicolas Rasmussen. Facts, artifacts, and mesosomes: Practicing epistemology with the electron microscope. Studies in History and Philosophy of Science Part A, 24(2):227–265, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390047N.

Rasmussen:1996:MMI

[Ras96] Nicolas Rasmussen. Making a machine instrumental: RCA and the wartime origins of biological electron microscopy in America, 1940–1945. Studies in History and Philosophy of Science Part A, 27 (3):311–349, September 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368195000429.

Ratzon:2019:JTF

[Rat19] Eshbal Ratzon. Jewish time: First stages of seasonal hours in Judea. Studies in History and Philosophy of Science Part A, 75 (??):23–33, June 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S003936811830044X.

Ratti:2020:WKN

[Rat20] Emanuele Ratti. What kind of novelties can machine learning possibly generate? The case of genomics. Studies in History and Philosophy of Science Part A, 83(??):86–96, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119302924.

Ray:1990:CCE

[Ray90] Christopher Ray. The cosmological constant: Einstein’s greatest mistake? Studies in History and Philosophy of Science Part A, 21 (4):589–604, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900346. See response

[Cur92] and reply [Ray92].

Ray:1992:FLA

[Ray92] Christopher Ray. Fundamental laws and ad hoc decisions: a reply to Curry. Studies in History and Philosophy of Science Part A, 23(4):661–664, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290017Z.

See [Ray90, Cur92].

Raz:2024:MIS

[Räz24] Tim Räz. ML interpretability: Simple isn’t easy. Studies in History and Philosophy of Science Part A, 103(??):159–167, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001723.

Ruthenberg:2017:AMC

[RC17] Klaus Ruthenberg and Hasok Chang. Acidity: Modes of characterization and quantification. Studies in History and Philosophy of Science Part A, 65–66(??):121–131, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301073.

RuizdeOlano:2025:NEP

[RDM25] Pablo Ruiz de Olano, Richard Dawid, and C. D. McCoy. Nonempirical physics from a historical perspective: New pathways in history and philosophy of physics. Studies in History and Philosophy of Science Part A, 110(??):13–16, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368125000044.

Resnik:2019:VEI

[RE19] David B. Resnik and Kevin C. Elliott. Value-entanglement and the integrity of scientific research. Studies in History and Philosophy of Science Part A, 75(??):1–11, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S003936811830092X.

Redhead:1980:BRM

[Red80a] M. L. G. Redhead. A Bayesian reconstruction of the methodology of scientific research programmes. Studies in History and Philosophy of Science Part A, 11(4):341–347, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368180900114.

Redhead:1980:SPA

[Red80b] M. L. G. Redhead. Some philosophical aspects of particle physics. Studies in History and Philosophy of Science Part A, 11(4):279– 304, December 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900084.

Redondi:1987:GHG

[Red87] Pietro Redondi. Galileo heretic = Galileo eretico. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08451-3. x + 356 + 16 pp. LCCN Q125.2 .R4313 1987. US$29.95.

Redekop:2002:TRP

[Red02] Benjamin W. Redekop. Thomas Reid and the problem of induction: from common experience to common sense. Studies in History and Philosophy of Science Part A, 33(1):35–57, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810100022X.

Reed:1986:JJG

[Ree86] Edward S. Reed. James J. Gibson’s revolution in perceptual psychology: a case study of the transformation of scientific ideas. Studies in History and Philosophy of Science Part A, 17(1):65– 98, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900191.

Reeves:2012:SLN

[Ree12] Eileen Reeves. Science and literature: a novel approach. Studies in History and Philosophy of Science Part A, 43(3):421–424, September 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000374.

Reingold:1996:BAH

[Rei96] Nathan Reingold. Between American history and history of science. Studies in History and Philosophy of Science Part A, 27(1):

115–129, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000305.

Reisch:1997:HPW

[Rei97] George A. Reisch, Jr. How postmodern was Neurath’s idea of unity of science? Studies in History and Philosophy of Science Part A, 28 (3):439–451, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000271.

Reid:2008:WWC

[Rei08] Francis Lucian Reid. William Wales (ca. 1734–1798): playing the astronomer. Studies in History and Philosophy of Science Part A, 39(2):170–175, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000186.

Reichenberger:2018:EDC

[Rei18] Andrea Reichenberger. Émilie Du Châtelet’s interpretation of the laws of motion in the light of 18th century mechanics. Studies in History and Philosophy of Science Part A, 69(??):1–11, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300177.

Reiss:2020:WDI

[Rei20] Julian Reiss. What are the drivers of induction? Towards a material theory. Studies in History and Philosophy of Science Part A, 83(??):8–16, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119301414.

Remmert:2007:VLA

[Rem07] Volker R. Remmert. Visual legitimisation of astronomy in the sixteenth and seventeeth centuries: Atlas, Hercules and Tycho’s nose. Studies in History and Philosophy of Science Part A, 38(2): 327–362, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000167.

Resnik:1989:AE

[Res89] David B. Resnik. Adaptationist explanations. Studies in History and Philosophy of Science Part A, 20(2):193–213, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900034.

Resnik:1996:SEE

[Res96] David Resnik. Social epistemology and the ethics of research. Studies in History and Philosophy of Science Part A, 27(4):565– 586, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819600043X.

Resnik:2000:PAD

[Res00] David B. Resnik. A pragmatic approach to the demarcation problem. Studies in History and Philosophy of Science Part A, 31(2): 249–267, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368100000042.

Rescorla:2012:CPC

[Res12] Michael Rescorla. Copeland and Proudfoot on computability. Studies in History and Philosophy of Science Part A, 43(1):199– 202, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001063.

Reutlinger:2022:WDN

[Reu22] Alexander Reutlinger. When do non-epistemic values play an epistemically illegitimate role in science? How to solve one half of the new demarcation problem. Studies in History and Philosophy of Science Part A, 92(??):152–161, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000280.

Reyes-Galindo:2015:BTK

[RGD15] Luis I. Reyes-Galindo and Tiago Ribeiro Duarte. Bringing tacit knowledge back to contributory and interactional expertise: a reply to Goddiksen. Studies in History and Philosophy of Science Part A, 49(??):99–102, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000958. See [God14].

Roll-Hansen:1995:RTE

[RH95] Nils Roll-Hansen. The role of theory in experimental life. Studies in History and Philosophy of Science Part A, 26(4):673–679, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000178.

Roll-Hansen:2011:LHS

[RH11a] Nils Roll-Hansen. Lessons from the history of science. Studies in History and Philosophy of Science Part A, 42(3):462–466, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000380.

Roll-Hansen:2011:SN

[RH11b] Nils Roll-Hansen. The Spell of the North. Studies in History and Philosophy of Science Part A, 42(1):227–230, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000920.

Rheinberger:1992:EDWa

[Rhe92a] Hans-Jörg Rheinberger. Experiment, difference, and writing: I. tracing protein synthesis. Studies in History and Philosophy of Science Part A, 23(2):305–331, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368192900377.

Rheinberger:1992:EDWb

[Rhe92b] Hans-Jörg Rheinberger. Experiment, difference, and writing: II. The laboratory production of transfer RNA. Studies in History and Philosophy of Science Part A, 23(3):389–422, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819290002N.

Rheinberger:1994:BRBa

[Rhe94a] Hans-Jörg Rheinberger. Book review: Iconology: Image, text, ideology: W. J. T. Mitchell, (Chicago and London: University of Chicago Press, 1987), x + 226 pp. ISBN 0-226-53228-3 Hardback, ISBN 0-226-53229-1 Paperback £8.95. Studies in History and Philosophy of Science Part A, 25(4):647–654, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900558. See [Mit86].

Rheinberger:1994:BRBb

[Rhe94b] Hans-Jörg Rheinberger. Book review: Representation in scientific practice: Michael Lynch and Steve Woolgar (eds), (Cambridge, Mass.: MIT Press, 1990), x + 365 pp. ISBN 0-262-62076-6 Paperback $16.95 /£14.95. Studies in History and Philosophy of Science Part A, 25(4):647–654, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819490054X.

See [LW90].

Rheinberger:2008:HEA

[Rhe08] Hans-Jörg Rheinberger. Heredity and its entities around 1900. Studies in History and Philosophy of Science Part A, 39(3):370– 374, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000617.

Rheinberger:2025:ROT

[Rhe25] Hans-Jörg Rheinberger. Reply: an order of things? Split and Splice. A Phenomenology of Experimentation by Hans-Jörg Rheinberger: Reply by the author. Studies in History and Philosophy of Science Part A, 110(??):17–18, April 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368124001560. See [SF25, SD25, MW25].

Ribe:1985:GCN

[Rib85] Neil M. Ribe. Goethe’s critique of Newton: a reconsideration. Studies in History and Philosophy of Science Part A, 16(4):315– 335, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368185900159.

Ribeiro:2007:RIE

[Rib07] Rodrigo Ribeiro. The role of interactional expertise in interpreting: the case of technology transfer in the steel industry. Studies in History and Philosophy of Science Part A, 38(4):713–721, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000647.

Richards:1986:PGM

[Ric86] Joan L. Richards. Projective geometry and mathematical progress in mid-Victorian Britain. Studies in History and Philosophy of Science Part A, 17(3):297–325, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900117.

Richards:1992:APE

[Ric92] Graham Richards. The absence of psychology in the eighteenth century: a linguistic perspective. Studies in History and Philosophy of Science Part A, 23(2):195–211, June 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368192900322.

Richards:1997:DIA

[Ric97] Richard A. Richards. Darwin and the inefficacy of artificial selection. Studies in History and Philosophy of Science Part A, 28(1): 75–97, March 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000088.

Richardson:2003:GKL

[Ric03] Alan Richardson. The geometry of knowledge: Lewis, Becker, Carnap and the formalization of philosophy in the 1920s. Studies in History and Philosophy of Science Part A, 34(1):165– 182, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000882.

Richardson:2005:RDM

[Ric05] Alan Richardson. Reichenbach’s disease and Mirowski’s theory of knowledge? Or, will to power as philosophy of science. Studies in History and Philosophy of Science Part A, 36(4):744–753, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000750.

Richardson:2009:LVCa

[Ric09a] Sarah S. Richardson. The Left Vienna Circle, Part 1. Carnap, Neurath, and the Left Vienna Circle thesis. Studies in History and Philosophy of Science Part A, 40(1):14–24, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108001088.

Richardson:2009:LVCb

[Ric09b] Sarah S. Richardson. The Left Vienna Circle, Part 2. The Left Vienna Circle, disciplinary history, and feminist philosophy of science. Studies in History and Philosophy of Science Part A, 40(2): 167–174, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000259.

Richmond:2021:IIG

[Ric21] Marsha L. Richmond. The imperative for inclusion: a gender analysis of genetics. Studies in History and Philosophy of Science Part A, 90(??):247–264, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001631.

Rice:2022:BRL

[Ric22] Collin Rice. Book review: Leveraging Distortions: Explanation, Idealization, and Universality in Science by Collin Rice: Reply by the Author. Studies in History and Philosophy of Science Part A, 95(??):233–235, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000899.

Riesch:2014:PHS

[Rie14] Hauke Riesch. Philosophy, history and sociology of science: Interdisciplinary relations and complex social identities. Studies in History and Philosophy of Science Part A, 48(??):30–37, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000910.

Riedel:2024:RQM

[Rie24] Timotheus Riedel. Relational quantum mechanics, quantum relativism, and the iteration of relativity. Studies in History and Philosophy of Science Part A, 104(??):109–118, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000232.

Riordan:2015:OSM

[Rio15] Sally Riordan. The objectivity of scientific measures. Studies in History and Philosophy of Science Part A, 50(??):38– 47, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000831.

Risjord:1990:SFM

[Ris90] Mark Risjord. The sensible foundation for mathematics: a defense of Kant’s view. Studies in History and Philosophy of Science Part A, 21(1):123–143, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900173.

Risjord:1991:FRS

[Ris91] Mark Risjord. Further reflections on the sensible foundation: Replies to Leavitt and Griffin. Studies in History and Philosophy of Science Part A, 22(4):665–672, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190039U.

See [Lea91, Gri91].

Risjord:1993:MME

[Ris93] Mark Risjord. Metaphysics, method, and the exact sciences. Studies in History and Philosophy of Science Part A, 24(3):493– 499, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390040Q.

Risjord:1998:NES

[Ris98] Mark Risjord. Norms and explanation in the social sciences. Studies in History and Philosophy of Science Part A, 29(2):223– 237, June 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000053.

Ritchie:2012:SPS

[Rit12a] Jack Ritchie. Styles for philosophers of science. Studies in History and Philosophy of Science Part A, 43(4):649–656, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000490.

Ritchie:2012:STS

[Rit12b] Jack Ritchie. Styles of thinking: the special issue. Studies in History and Philosophy of Science Part A, 43(4):595–598, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811200043X.

Ritson:2021:CDA

[Rit21] Sophie Ritson. Constraints and divergent assessments of fertility in non-empirical physics in the history of the string theory controversy. Studies in History and Philosophy of Science Part A, 90(??):39–49, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001321.

Rivat:2021:DSA

[Riv21] Sébastien Rivat. Drawing scales apart: the origins of Wilson’s conception of effective field theories. Studies in History and Philosophy of Science Part A, 90(??):321–338, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001692.

Raizman-Kedar:2006:PCU

[RK06] Yael Raizman-Kedar. Plotinus’s conception of unity and multiplicity as the root to the medieval distinction between lux and lumen. Studies in History and Philosophy of Science Part A, 37 (3):379–397, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000471.

Rochefort-Maranda:2016:HWL

[RM16] Guillaume Rochefort-Maranda. How we load our data sets with theories and why we do so purposefully. Studies in History and Philosophy of Science Part A, 60(??):1–6, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368116300437.

Robinson:1970:PLS

[Rob70] Eric Robinson. Priestley’s library of scientific books: a new list. Studies in History and Philosophy of Science Part A, 1(2):145– 160, August 1970. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368170900038.

Roberts:1995:DSC

[Rob95] Lissa Roberts. The death of the sensuous chemist: The ‘new’ chemistry and the transformation of sensuous technology. Studies in History and Philosophy of Science Part A, 26(4):503–529, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000135.

Robson:2003:AM

[Rob03] Eleanor Robson. Ancient mathematics. Studies in History and Philosophy of Science Part A, 34(2):423–429, June 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810300027X.

Robcis:2022:BRA

[Rob22] Camille Robcis. Book review: ‘As.fūriyyeh: a history of madness, modernity, and war in the Middle East by Joelle M. Abi-Rached: Psychiatry as politics. Studies in History and Philosophy of Science Part A, 93(??):223–224, June 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000103. See reply [AR22].

Rochowiak:1988:DPT

[Roc88] Daniel Rochowiak. Darwin’s psychological theorizing: Triangulating on habit. Studies in History and Philosophy of Science Part A, 19(2):215–241, June 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368188900271.

Rochberg:2002:CBA

[Roc02] Francesca Rochberg. A consideration of Babylonian astronomy within the historiography of science. Studies in History and Philosophy of Science Part A, 33(4):661–684, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000225.

Rodriguez:2021:IAE

[Rod21] Quentin Rodriguez. Idealizations and analogies: Explaining critical phenomena. Studies in History and Philosophy of Science Part A, 89(??):235–247, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001230.

Rogers:1972:BRB

[Rog72] G. A. J. Rogers. Book review: Locke’s philosophy of science and knowledge. A consideration of some aspects of ‘an essay concerning human understanding’: R. S. Woolhouse, Oxford: Basil Blackwell, 1971. £2.75. Studies in History and Philosophy of Science Part A, 3(2):183–189, August 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368172900234.

See [Woo71, Woo94].

Rogers:1995:GBM

[Rog95] G. A. J. Rogers. Gassendi and the birth of modern philosophy. Studies in History and Philosophy of Science Part A, 26(4):681– 687, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000216.

Rolin:2016:VSS

[Rol16] Kristina Rolin. Values, standpoints, and scientific/intellectual movements. Studies in History and Philosophy of Science Part A, 56(??):11–19, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001569.

Romero:2016:CBS

[Rom16] Felipe Romero. Can the behavioral sciences self-correct? A social epistemic study. Studies in History and Philosophy of Science Part A, 60(??):55–69, December 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300802.

Rosenblueth:1970:MBP

[Ros70] Arturo Rosenblueth. Mind and brain; a philosophy of science. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-18041-3. xii + 128 pp. LCCN Q175 .R55.

Rosales:2017:TNW

[Ros17] Alirio Rosales. Theories that narrate the world: Ronald A. Fisher’s mass selection and Sewall Wright’s shifting balance. Studies in History and Philosophy of Science Part A, 62(??):22– 30, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300742.

Rossetter:2018:RRN

[Ros18] Thomas Rossetter. Realism on the rocks: Novel success and James Hutton’s theory of the earth. Studies in History and Philosophy of Science Part A, 67(??):1–13, February 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117300390.

Rossini:2019:NTN

[Ros19] Paolo Rossini. New theories for new instruments: Fabrizio Mordente’s proportional compass and the genesis of Giordano Bruno’s atomist geometry. Studies in History and Philosophy of Science Part A, 76(??):60–68, August 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301912.

Rosochotsky:2022:RJB

[Ros22] Aviram Rosochotsky. R. J. Boscovich on physical symmetries. Studies in History and Philosophy of Science Part A, 93(??):149– 162, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000462.

Roth:1996:WRS

[Rot96] Paul A. Roth. Will the real scientists please stand up? Dead ends and live issues in the explanation of scientific knowledge. Studies in History and Philosophy of Science Part A, 27(1):43– 68, March 1996. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000364.

Rothman:2011:DAC

[Rot11] Aviva Rothman. Defining astronomical community in early modern Europe. Studies in History and Philosophy of Science Part A, 42(1):231–234, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000932.

Roth:2013:SNM

[Rot13] Paul A. Roth. The silence of the norms: the missing historiography of The Structure of Scientific Revolutions. Studies in History and Philosophy of Science Part A, 44(4):545–552, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000605.

Roth:2017:ENE

[Rot17] Paul A. Roth. Essentially narrative explanations. Studies in History and Philosophy of Science Part A, 62(??):42–50, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300754.

Roth:2022:BRS

[Rot22] Paul A. Roth. Book review: Social Science for What? Battles over Public Funding for the “Other Sciences” at the National Science Foundation by Mark Solovey: Where’s the Beef? Foibles of Social Science Funding at NSF. Studies in History and Philosophy of Science Part A, 94(??):204–205, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000383. See [Lin22a, Hau22, Tur22b]

and reply [Sol22].

Rouse:1991:PSP

[Rou91] Joseph Rouse. Philosophy of science and the persistent narratives of modernity. Studies in History and Philosophy of Science Part A, 22(1):141–162, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819190018N.

Rouse:2005:ED

[Rou05] Joseph Rouse. Epistemological derangement. Studies in History and Philosophy of Science Part A, 36(4):835–847, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000683.

Rowbottom:2008:RSV

[Row08a] Darrell P. Rowbottom. N-rays and the semantic view of scientific progress. Studies in History and Philosophy of Science Part A, 39 (2):277–278, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000320. See reply [Bir08].

Rowbottom:2008:IC

[Row08b] Darrell Patrick Rowbottom. Intersubjective corroboration. Studies in History and Philosophy of Science Part A, 39(1):124– 132, March 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000921.

Rowbottom:2011:KVP

[Row11a] Darrell P. Rowbottom. Kuhn vs. Popper on criticism and dogmatism in science: a resolution at the group level. Studies in History and Philosophy of Science Part A, 42(1):117–124, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811000110X.

Rowbottom:2011:WBS

[Row11b] Darrell P. Rowbottom. What’s at the bottom of scientific realism? Studies in History and Philosophy of Science Part A, 42 (4):625–628, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811100080X.

Rowbottom:2013:KVP

[Row13] Darrell P. Rowbottom. Kuhn vs. Popper on criticism and dogmatism in science, part II: How to strike the balance. Studies in History and Philosophy of Science Part A, 44(2):161– 168, June 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112001161.

Rowbottom:2015:SPI

[Row15] Darrell P. Rowbottom. Scientific progress without increasing verisimilitude: In response to Niiniluoto. Studies in History and Philosophy of Science Part A, 51(??):100–104, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000199. See [Nii14a].

Rowbottom:2023:BRI

[Row23] Darrell P. Rowbottom. Book review: The Instrument of Science: Scientific Anti-Realism Revitalised by Darrell Rowbottom: Reply by the Author. Studies in History and Philosophy of Science Part A, 98(??):9–11, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123000286.

Ritson:2021:HUC

[RS21a] Sophie Ritson and Kent Staley. How uncertainty can save measurement from circularity and holism. Studies in History and Philosophy of Science Part A, 85(??):155–165, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301886.

Rubin:2021:PPS

[RS21b] Hannah Rubin and Mike D. Schneider. Priority and privilege in scientific discovery. Studies in History and Philosophy of Science Part A, 89(??):202–211, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001217.

Rusnock:1995:SCC

[RT95] Paul Rusnock and Paul Thagard. Strategies for conceptual change: Ratio and proportion in classical Greek mathematics.

Studies in History and Philosophy of Science Part A, 26(1):107– 131, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194000347.

Rudwick:1974:DGR

[Rud74] Martin Rudwick. Darwin and Glen Roy: a “great failure” in scientific method? Studies in History and Philosophy of Science Part A, 5(2):97–185, August 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368174900247.

RuizdeOlano:2023:CPW

[Rui23] Pablo Ruiz de Olano. Confirmation, or pursuit-worthiness? Lessons from J. J. Sakurai’s 1960 theory of the strong force for the debate on non-empirical physics. Studies in History and Philosophy of Science Part A, 99(??):77–88, June 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123000717.

Runde:1994:KAR

[Run94] Jochen Runde. Keynes after Ramsey: In defence of a treatise on probability. Studies in History and Philosophy of Science Part A, 25(1):97–121, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194900221.

Rupke:1976:BHP

[Rup76] Nicolaas A. Rupke. Bathybius Haeckelii and the psychology of scientific discovery: Theory instead of observed data controlled the late 19th century ‘discovery’ of a primitive form of life. Studies in History and Philosophy of Science Part A, 7(1):53– 62, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900030.

Rupert:2022:BRR

[Rup22] Robert D. Rupert. Book review: Representation in Cognitive Science by Nicholas Shea: Content without Function. Studies in History and Philosophy of Science Part A, 92(??):260– 263, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000807.

Ruse:1971:NSO

[Rus71] Michael Ruse. Natural selection in The Origin of Species. Studies in History and Philosophy of Science Part A, 1(4):311–351, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900112.

Ruse:1973:PB

[Rus73] Michael Ruse. The philosophy of biology. Philosophy. Hutchinson University Library, London, UK, 1973. ISBN 0-09-115220-8, 0-09-115221-6 (paperback). 231 pp. LCCN QH331 .R879.

Ruse:1975:DDP

[Rus75] Michael Ruse. Darwin’s debt to philosophy: An examination of the influence of the philosophical ideas of John F. W. Herschel and William Whewell on the development of Charles Darwin’s Theory of Evolution. Studies in History and Philosophy of Science Part A, 6(2):159–181, June 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368175900199.

Ruse:1978:DH

[Rus78] Michael Ruse. Darwin and Herschel. Studies in History and Philosophy of Science Part A, 9(4):323–331, December 1978. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368178900134.

Ruse:1989:WPB

[Rus89] Michael Ruse, editor. What the philosophy of biology is: essays dedicated to David Hull, volume 32 of Nijhoff international philosophy series. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 90-247-3778-8. viii + 337 pp. LCCN QH331 .W466 1989.

Rushing:2023:PDR

[Rus23] Bruce Rushing. Putting the “Decision” in Ramsey’s “Theories”. Studies in History and Philosophy of Science Part A, 102(??):48– 59, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001450.

Ruyant:2023:CHT

[Ruy23] Dr Quentin Ruyant. Consistent histories through pragmatist lenses. Studies in History and Philosophy of Science Part A, 98(??): 40–48, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000018.

Ryckman:1992:POC

[Ryc92] T. A. Ryckman. “P(oint)-c(oincidence) thinking”: The ironical attachment of logical empiricism to general relativity (and some lingering consequences). Studies in History and Philosophy of Science Part A, 23(3):471–497, September 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290005Q.

Ryckman:2015:WHM

[Ryc15] Thomas Ryckman. Why history matters to philosophy of physics. Studies in History and Philosophy of Science Part A, 50(??):4– 12, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000880.

Rynasiewicz:1995:TPCa

[Ryn95a] Robert Rynasiewicz. By their properties, causes and effects: Newton’s Scholium on time, space, place and motion — I. The text. Studies in History and Philosophy of Science Part A, 26(1):133– 153, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194000358.

Rynasiewicz:1995:TPCb

[Ryn95b] Robert Rynasiewicz. By their properties, causes and effects: Newton’s Scholium on time, space, place and motion — II. The context. Studies in History and Philosophy of Science Part A, 26(2): 295–321, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819400049F.

Roe:2021:URW

[RZ21] Sarah M. Roe and Elyse Zavar. Understanding the role of wrongdoing in technological disasters: Utilizing ecofeminist philosophy to examine commemoration. Studies in History and Philosophy of Science Part A, 87(??):158–167, June 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000340.

Saatsi:2005:RFM

[Saa05] Juha Saatsi. Reconsidering the Fresnel–Maxwell theory shift: how the realist can have her cake and EAT it too. Studies in History and Philosophy of Science Part A, 36(3):509–538, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000506.

Saatsi:2008:ERP

[Saa08] Juha Saatsi. Eclectic realism — the proof of the pudding: a reply to Busch. Studies in History and Philosophy of Science Part A, 39(2):273–276, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000319. See [Bus08].

Sakellariou:2022:COS

[Sak22] Vassilis Sakellariou. Constituting the ‘object’ of science in Newton’s Principia: the many faces of Janus. Studies in History and Philosophy of Science Part A, 95(??):28–36, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000991.

Salis:2017:MEM

[Sal17] Fiora Salis. Models and exploratory models. Studies in History and Philosophy of Science Part A, 63(??):58–61, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301085.

Sambursky:1977:PSL

[Sam77] Shmuel Sambursky. Place and space in late neoplatonism. Studies in History and Philosophy of Science Part A, 8(3):173–187, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900061.

Samuels:2010:CCM

[Sam10] Richard Samuels. Classical computationalism and the many problems of cognitive relevance. Studies in History and Philosophy of Science Part A, 41(3):280–293, September 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000385.

Sample:2017:SPR

[Sam17] Matthew Sample. Silent performances: Are “repertoires” really post-Kuhnian? Studies in History and Philosophy of Science Part A, 61(??):51–56, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300814.

Sankey:1991:TFB

[San91] Howard Sankey. Translation failure between theories. Studies in History and Philosophy of Science Part A, 22(2):223– 236, June 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190025N.

Sankey:2008:SRI

[San08] Howard Sankey. Scientific realism and the inevitability of science. Studies in History and Philosophy of Science Part A, 39(2): 259–264, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000289.

Sankey:2009:CDR

[San09a] Howard Sankey. A curious disagreement: response to Hoyningen-Huene and Oberheim. Studies in History and Philosophy of Science Part A, 40(2):210–212, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368109000181.

Sankey:2009:SRS

[San09b] Howard Sankey. Scientific realism and the semantic incommensurability thesis. Studies in History and Philosophy of Science Part A, 40(2):196–202, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000260. See reply [HHO09].

Sankey:2011:ERP

[San11] Howard Sankey. Epistemic relativism and the problem of the criterion. Studies in History and Philosophy of Science Part A, 42 (4):562–570, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111000768.

Sankey:2012:SRA

[San12] Howard Sankey. Scepticism, relativism and the argument from the criterion. Studies in History and Philosophy of Science Part A, 43(1):182–190, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111001373.

Sankey:2013:HER

[San13] Howard Sankey. How the epistemic relativist may use the sceptic’s strategy: a reply to Markus Seidel. Studies in History and Philosophy of Science Part A, 44(1):140–144, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000787.

Sankey:2014:RPR

[San14] Howard Sankey. On relativism and pluralism: Response to Steven Bland. Studies in History and Philosophy of Science Part A, 47 (??):98–103, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114000223. See [Bla13].

Sandoz:2016:WCS

[San16a] Raphaël Sandoz. Whewell on the classification of the sciences. Studies in History and Philosophy of Science Part A, 60(??):48– 54, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300784.

Sangiacomo:2016:SCA

[San16b] Andrea Sangiacomo. From secondary causes to artificial instruments: Pierre-Sylvain Régis’s rethinking of scholastic accounts of causation. Studies in History and Philosophy of Science Part A, 60(??):7–17, December 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116300462.

Santana:2022:WCR

[San22] Carlos Santana. Why citizen review might beat peer review at identifying pursuitworthy scientific research. Studies in History and Philosophy of Science Part A, 92(??):20–26, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000127.

Sarkar:1981:TPS

[Sar81] Husain Sarkar. Truth, problem-solving and methodology. Studies in History and Philosophy of Science Part A, 12(1):61–73, March 1981. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368181900054.

Sarkar:1982:TGR

[Sar82] Husain Sarkar. A theory of group rationality. Studies in History and Philosophy of Science Part A, 13(1):55–72, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900048.

Sarkar:1983:DT

[Sar83] Husain Sarkar. In defence of truth. Studies in History and Philosophy of Science Part A, 14(1):67–79, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368183900067.

Sargent:1986:RBB

[Sar86] Rose-Mary Sargent. Robert Boyle’s Baconian inheritance: a response to Laudan’s Cartesian thesis. Studies in History and Philosophy of Science Part A, 17(4):469–486, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900051.

Sargent:1989:SEL

[Sar89] Rose-Mary Sargent. Scientific experiment and legal expertise: The way of experience in seventeenth-century England. Studies in History and Philosophy of Science Part A, 20(1):19–45, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900320.

Sarkar:1997:TGRa

[Sar97a] Husain Sarkar. The task of group rationality: The subjectivist’s view — Part I. Studies in History and Philosophy of Science Part A, 28(2):267–288, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368196000258.

Sarkar:1997:TGRb

[Sar97b] Husain Sarkar. The task of group rationality: The subjectivist’s view — Part II. Studies in History and Philosophy of Science Part A, 28(3):497–520, September 1997. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936819600026X.

Sargent:2012:BBV

[Sar12] Rose-Mary Sargent. From Bacon to Banks: The vision and the realities of pursuing science for the common good. Studies in History and Philosophy of Science Part A, 43(1):82–90, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000999.

Sarkar:2015:NR

[Sar15a] Sahotra Sarkar. Nagel on reduction. Studies in History and Philosophy of Science Part A, 53(??):43–56, October 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000667.

Sartenaer:2015:EED

[Sar15b] Olivier Sartenaer. Emergent evolutionism, determinism and unpredictability. Studies in History and Philosophy of Science Part A, 51(??):62–68, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000412.

Soderqvist:2009:BMC

[SBM09] Thomas Söderqvist, Adam Bencard, and Camilla Mordhorst. Between meaning culture and presence effects: contemporary biomedical objects as a challenge to museums. Studies in History and Philosophy of Science Part A, 40(4):431–438, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000740.

Scerri:2005:RBC

[Sce05] Eric R. Scerri. Response to Barnes’s critique of Scerri and Worrall. Studies in History and Philosophy of Science Part A, 36(4):813– 816, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000701. See [SW01, Bar05,

Wor05, Sch08b].

Scerri:2006:CRE

[Sce06] Eric R. Scerri. On the continuity of reference of the elements: a response to Hendry. Studies in History and Philosophy of Science Part A, 37(2):308–321, June 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106000276.

Scerri:2020:PTT

[Sce20] Eric R. Scerri. The periodic table and the turn to practice. Studies in History and Philosophy of Science Part A, 79(??):87– 93, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118302280.

Schneer:1969:THG

[Sch69] Cecil J. Schneer, editor. Toward a history of geology; proceedings. MIT Press, Cambridge, MA, USA, 1969. ISBN 0-262-19058-3. LCCN QE11 .N532 1967.

Schneer:1970:MMM

[Sch70] Cecil J. Schneer. Mind and matter: man’s changing concepts of the material world. Grove Press, New York, NY, USA, 1970. ISBN ???? xiv + 305 pp. LCCN ????

Schagrin:1973:WTS

[Sch73] Morton L. Schagrin. Whewell’s theory of scientific language. Studies in History and Philosophy of Science Part A, 4(3):231–240, November 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900095.

Schaffner:1974:LDJ

[Sch74] Kenneth Schaffner. Logic of discovery and justification in regulatory genetics. Studies in History and Philosophy of Science Part A, 4(4):349–385, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368174900090.

Schmaus:1985:HHA

[Sch85] Warreb Schmaus. Hypotheses and historical analysis in Durkheim’s sociological methodology: a Comtean tradition. Studies in History and Philosophy of Science Part A, 16(1):1–30, March 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368185900056.

Schabas:1987:ALP

[Sch87] Margaret Schabas. An anomaly for Laudan’s pragmatic model. Studies in History and Philosophy of Science Part A, 18(1):43– 52, March 1987. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368187900100.

Schaffer:1988:WTH

[Sch88a] Simon Schaffer. Wallifaction: Thomas Hobbes on school divinity and experimental pneumatics. Studies in History and Philosophy of Science Part A, 19(3):275–298, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368188900015.

Scheibe:1988:PCP

[Sch88b] Erhard Scheibe. The physicists’ conception of progress. Studies in History and Philosophy of Science Part A, 19(2):141– 159, June 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900246.

Schipper:1988:WWC

[Sch88c] Frits Schipper. William Whewell’s conception of scientific revolutions. Studies in History and Philosophy of Science Part A, 19(1): 43–53, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900192.

Schneer:1988:MMM

[Sch88d] Cecil J. Schneer. Mind and matter: man’s changing concepts of the material world. History of science and technology reprint series. Iowa State University Press, Ames, IA, USA, 1988. ISBN 0-8138-0013-7. xiii + 305 pp. LCCN QD11 .S29 1988.

Schaffer:1991:BRB

[Sch91] Simon Schaffer. Book review: The pasteurization of France: Bruno Latour, translated by Alan Sheridan and John Law (Cambridge, Massachusetts and London: Harvard University Press, 1988), 273 pp. ISBN 0-674-65760-8 Cloth £23.95. Studies in History and Philosophy of Science Part A, 22(1):174–192, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190020S. See [Lat88].

Schlich:1993:MMS

[Sch93] Thomas Schlich. Making mistakes in science: Eduard Pflüger, his scientific and professional concept of Physiology, and his unsuccessful theory of diabetes (1903–1910). Studies in History and Philosophy of Science Part A, 24(3):411–441, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390036J.

Schmitt:2002:TJP

[Sch02a] Frederick F. Schmitt. Testimonial justification: the parity argument. Studies in History and Philosophy of Science Part A, 33(2): 385–406, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000110.

Schwitzgebel:2002:WDW

[Sch02b] Eric Schwitzgebel. Why did we think we dreamed in black and white? Studies in History and Philosophy of Science Part A, 33 (4):649–660, December 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810200033X.

Schmaltz:2003:CCB

[Sch03a] Tad M. Schmaltz. Cartesian causation: body-body interaction, motion, and eternal truths. Studies in History and Philosophy of Science Part A, 34(4):737–762, December 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368103000670.

Schummer:2003:NNC

[Sch03b] Joachim Schummer. The notion of nature in chemistry. Studies in History and Philosophy of Science Part A, 34(4):705–736, December 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000505.

Schickore:2005:TTE

[Sch05] Jutta Schickore. ‘through thousands of errors we reach the truth’ — but how? On the epistemic roles of error in scientific practice. Studies in History and Philosophy of Science Part A, 36 (3):539–556, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000385.

Scheele:2006:FUT

[Sch06] Marcel Scheele. Function and use of technical artefacts: social conditions of function ascription. Studies in History and Philosophy of Science Part A, 37(1):23–36, March 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105001068.

Schilhab:2007:IET

[Sch07a] Theresa Schilhab. Interactional expertise through the looking glass: a peek at mirror neurons. Studies in History and Philosophy of Science Part A, 38(4):741–747, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000659.

Schindler:2007:RTR

[Sch07b] Samuel Schindler. Rehabilitating theory: refusal of the ‘bottomup’ construction of scientific phenomena. Studies in History and Philosophy of Science Part A, 38(1):160–184, March 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368106001063.

Schmaus:2007:RMH

[Sch07c] Warren Schmaus. Renouvier and the method of hypothesis. Studies in History and Philosophy of Science Part A, 38(1):132– 148, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106001087.

Schabas:2008:HMT

[Sch08a] Margaret Schabas. Hume’s monetary thought experiments. Studies in History and Philosophy of Science Part A, 39(2):161– 169, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000174.

Schindler:2008:UNP

[Sch08b] Samuel Schindler. Use-novel predictions and Mendeleev’s periodic table: response to Scerri and Worrall (2001). Studies in History and Philosophy of Science Part A, 39(2):265–269, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000290. See [SW01, Bar05, Sce05,

Bar05, Wor05].

Schmidt:2008:KTE

[Sch08c] Claudia M. Schmidt. Kant’s transcendental and empirical psychology of cognition. Studies in History and Philosophy of Science Part A, 39(4):462–472, December 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000800.

Scheffler:2009:IIM

[Sch09a] Robin Wolfe Scheffler. Interests and instrument: a micro-history of object Wh.3469 (X-ray powder diffraction camera, ca. 1940). Studies in History and Philosophy of Science Part A, 40(4):396– 404, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000727.

Schyfter:2009:BAC

[Sch09b] Pablo Schyfter. The bootstrapped artefact: a collectivist account of technological ontology, functions, and normativity. Studies in History and Philosophy of Science Part A, 40(1):102– 111, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108001167.

Schliesser:2011:NSM

[Sch11] Eric Schliesser. Newton’s substance monism, distant action, and the nature of Newton’s empiricism: discussion of H. Kochiras, “Gravity and Newton’s Substance Counting Problem”. Studies in History and Philosophy of Science Part A, 42(1):160– 166, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811000083X. See [Koc09].

Schliesser:2012:IPM

[Sch12] Eric Schliesser. Inventing paradigms, monopoly, methodology, and mythology at ‘Chicago’: Nutter, Stigler, and Milton Friedman. Studies in History and Philosophy of Science Part A, 43(1):160– 171, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001105.

Schindler:2013:TLE

[Sch13a] Samuel Schindler. Theory-laden experimentation. Studies in History and Philosophy of Science Part A, 44(1):89–101, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000684.

Schliesser:2013:RNE

[Sch13b] Eric Schliesser. On reading Newton as an Epicurean: Kant, Spinozism and the changes to the Principia. Studies in History and Philosophy of Science Part A, 44(3):416–428, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000878.

Scholl:2013:CIM

[Sch13c] Raphael Scholl. Causal inference, mechanisms, and the Semmelweis case. Studies in History and Philosophy of Science Part A, 44(1):

66–76, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000350.

Schindler:2014:NCM

[Sch14a] Samuel Schindler. Novelty, coherence, and Mendeleev’s periodic table. Studies in History and Philosophy of Science Part A, 45(??): 62–69, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113001039.

Schurz:2014:BPC

[Sch14b] Gerhard Schurz. Bayesian pseudo-confirmation, use-novelty, and genuine confirmation. Studies in History and Philosophy of Science Part A, 45(??):87–96, March 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113001040.

Schmaltz:2015:GDC

[Sch15] Tad M. Schmaltz. Galileo and Descartes on Copernicanism and the cause of the tides. Studies in History and Philosophy of Science Part A, 51(??):70–81, June 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000229.

Schickore:2016:EEP

[Sch16] Jutta Schickore. “Exploratory experimentation” as a probe into the relation between historiography and philosophy of science. Studies in History and Philosophy of Science Part A, 55(??):20– 26, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001004.

Schindler:2017:KTC

[Sch17] Samuel Schindler. Kuhnian theory-choice and virtue convergence: Facing the base rate fallacy. Studies in History and Philosophy of Science Part A, 64(??):30–37, August 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300723.

Schindler:2018:CCA

[Sch18a] Samuel Schindler. A coherentist conception of ad hoc hypotheses. Studies in History and Philosophy of Science Part A, 67(??):54– 64, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301036.

Schmid:2018:SMD

[Sch18b] Jelscha Schmid. Schelling’s method of Darstellung: Presenting nature through experiment. Studies in History and Philosophy of Science Part A, 69(??):12–22, June 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301425.

Schmaus:2020:PCC

[Sch20a] Warren Schmaus. From positivism to conventionalism: Comte, Renouvier, and Poincaré. Studies in History and Philosophy of Science Part A, 80(??):102–109, April 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119300263.

Scholl:2020:UAC

[Sch20b] Raphael Scholl. Unwarranted assumptions: Claude Bernard and the growth of the vera causa standard. Studies in History and Philosophy of Science Part A, 82(??):120–130, August 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118303261.

Schneider:2021:TPP

[Sch21a] Mike D. Schneider. Trans-Planckian philosophy of cosmology. Studies in History and Philosophy of Science Part A, 90(??):184– 193, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001576.

Schroeren:2021:QMI

[Sch21b] David Schroeren. Quantum metaphysical indeterminacy and the ontological foundations of orthodoxy. Studies in History and Philosophy of Science Part A, 90(??):235–246, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001424.

Schneider:2023:ESP

[Sch23] Mike D. Schneider. Empty space and the (positive) cosmological constant. Studies in History and Philosophy of Science Part A, 100(??):12–21, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123000882.

Schindler:2024:PAA

[Sch24] Samuel Schindler. Predictivism and avoidance of ad hoc-ness: an empirical study. Studies in History and Philosophy of Science Part A, 104(??):68–77, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000177.

Sciortino:2021:EOF

[Sci21] Luca Sciortino. The emergence of objectivity: Fleck, Foucault, Kuhn and Hacking. Studies in History and Philosophy of Science Part A, 88(??):128–137, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000881.

Scott:1970:CBA

[Sco70] Wilson L. (Wilson Ludlow) Scott. The conflict between atomism and conservation theory, 1644–1860. History of science library. Macdonald and Co., London, UK, 1970. ISBN 0-444-19699-4. xiv + 312 + 5 pp. LCCN QC173 .S292.

Sanchez-Dorado:2023:CPE

[SD23] Julia Sánchez-Dorado. Creativity, pursuit and epistemic tradition. Studies in History and Philosophy of Science Part A, 100(??):81– 89, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000833.

Suarez-Diaz:2025:SSH

[SD25] Edna Suárez-Díaz. Split and splice by Hans-Jörg Rheinberger: Productive differentiation for historians and philosophers of the sciences. Studies in History and Philosophy of Science Part A, 109(??):132–133, February 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001638. See reply [Rhe25].

Selinger:2007:IEE

[SDC07] Evan Selinger, Hubert Dreyfus, and Harry Collins. Interactional expertise and embodiment. Studies in History and Philosophy of Science Part A, 38(4):722–740, December 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000660.

Segre:1989:GVT

[Seg89] Michael Segre. Galileo, Viviani and the Tower of Pisa. Studies in History and Philosophy of Science Part A, 20(4):435–451, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368189900186.

Seguin:2000:BLF

[Seg00] Eve Seguin. Bloor, Latour, and the field. Studies in History and Philosophy of Science Part A, 31(3):503–508, September 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000200.

Seidel:2013:SCE

[Sei13a] Markus Seidel. Scylla and Charybdis of the epistemic relativist: Why the epistemic relativist still cannot use the sceptic’s strategy. Studies in History and Philosophy of Science Part A, 44(1): 145–149, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000799.

Seidel:2013:WER

[Sei13b] Markus Seidel. Why the epistemic relativist cannot use the sceptic’s strategy. A comment on Sankey. Studies in History and Philosophy of Science Part A, 44(1):134–139, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000404.

Sent:2001:SSS

[Sen01] Esther-Mirjam Sent. Sent simulating Simon simulating scientists. Studies in History and Philosophy of Science Part A, 32 (3):479–500, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000127. See comments

[Sim01].

Serjeantson:1999:TPE

[Ser99] R. W. Serjeantson. Testimony and proof in early-modern England. Studies in History and Philosophy of Science Part A, 30(2): 195–236, June 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000508.

Serpico:2023:WSC

[Ser23] Davide Serpico. A wolf in sheep’s clothing: Idealisations and the aims of polygenic scores. Studies in History and Philosophy of Science Part A, 102(??):72–83, December 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001449.

Seth:2008:CQA

[Set08] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the older quantum theory. Studies in History and Philosophy of Science Part A, 39(3):335–348, September 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000587.

Steeger:2021:CLS

[SF21] Jeremy Steeger and Benjamin H. Feintzeig. Is the classical limit “singular”? Studies in History and Philosophy of Science Part A, 88(??):263–279, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000716.

Scott-Fordsmand:2025:TWT

[SF25] Helene Scott-Fordsmand. Tracing the world through grasp and synthesis. Studies in History and Philosophy of Science Part A, 109 (??):106–108, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124001572. See reply [Rhe25].

Sigl:2023:CAR

[SFF23] Lisa Sigl, Ruth Falkenberg, and Maximilian Fochler. Changing articulations of relevance in soil science: Diversity and (potential) synergy of epistemic commitments in a scientific discipline. Studies in History and Philosophy of Science Part A, 97(??):79– 90, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001819.

Scott-Fordsmand:2023:ADM

[SFT23] Helene Scott-Fordsmand and Karin Tybjerg. Approaching diagnostic messiness through spiderweb strategies: Connecting epistemic practices in the clinic and the laboratory. Studies in History and Philosophy of Science Part A, 102(??):12–21, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001073.

Stefanov:1985:ODS

[SG85] Anguel Stefanov and Dimiter Ginev. One dimension of the scientific type of rationality (a reflection upon the theory of group rationality). Studies in History and Philosophy of Science Part A, 16(2):101–111, June 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368185900093.

Slaveva-Griffin:2013:TPA

[SG13] Svetla Slaveva-Griffin. Is there philosophy after Aristotle? Studies in History and Philosophy of Science Part A, 44(1):154– 159, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000702.

Steel:2017:SAS

[SGO17] Daniel Steel, Chad Gonnerman, and Michael O’Rourke. Scientists’ attitudes on science and values: Case studies and survey methods in philosophy of science. Studies in History and Philosophy of Science Part A, 63(??):22–30, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301048.

Shapiro:1974:LPR

[Sha74] Alan E. Shapiro. Light, pressure, and rectilinear propagation: Descartes’ celestial optics and Newton’s hydrostatics. Studies in History and Philosophy of Science Part A, 5(3):239–296, November 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368174900028.

Shaffer:1976:ER

[Sha76] E. S. Shaffer. Essay review. Studies in History and Philosophy of Science Part A, 7(3):269–275, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936817690008X.

Shanahan:1989:KNO

[Sha89] Timothy Shanahan. Kant, naturphilosophie, and Oersted’s discovery of electromagnetism: a reassessment. Studies in History and Philosophy of Science Part A, 20(3):287–305, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368189900095.

Sharratt:1990:BRBa

[Sha90a] Michael Sharratt. Book review: Galileo heretic (Galileo eretico): by Pietro Redondi, translated by Raymond Rosenthal (Harmondsworth, Middlesex: Allen Lane, The Penguin Press, 1988), pp. x + 356, £17.95. Studies in History and Philosophy of Science Part A, 21(4):685, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090039B. See [Red87].

Sharratt:1990:BRBb

[Sha90b] Michael Sharratt. Book review: The Galileo affair: A documentary history: edited and translated with an introduction and notes by Maurice A. Finocchiaro (Berkeley: University of California Press, 1989), pp. xvi + 382, £8.95. Studies in History and Philosophy of Science Part A, 21(4):685–690, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819090040F. See [Fin89].

Shapiro:1994:BIF

[Sha94] Lawrence A. Shapiro. Behavior, ISO functionalism, and psychology. Studies in History and Philosophy of Science Part A, 25(2): 191–209, April 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900272.

Shanahan:1997:KCC

[Sha97] Timothy Shanahan. Kitcher’s Compromise: a critical examination of the Compromise Model of scientific closure, and its implications for the relationship between history and philosophy of science. Studies in History and Philosophy of Science Part A, 28(2): 319–338, June 1997. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368196000234.

Shabel:1998:KSC

[Sha98] Lisa Shabel. Kant on the ‘symbolic construction’ of mathematical concepts. Studies in History and Philosophy of Science Part A, 29 (4):589–621, December 1998. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368198000235.

Shapiro:2002:TSC

[Sha02] Barbara J. Shapiro. Testimony in seventeenth-century English natural philosophy: legal origins and early development. Studies in History and Philosophy of Science Part A, 33(2):243– 263, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000109.

Shabel:2003:RKC

[Sha03] Lisa Shabel. Reflections on Kant’s concept (and intuition) of space. Studies in History and Philosophy of Science Part A, 34(1): 45–57, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000894.

Shagrir:2010:BAM

[Sha10] Oron Shagrir. Brains as analog-model computers. Studies in History and Philosophy of Science Part A, 41(3):271–279, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000397.

Shaw:2017:WFA

[Sha17] Jamie Shaw. Was Feyerabend an anarchist? The structure(s) of ‘anything goes’. Studies in History and Philosophy of Science Part A, 64(??):11–21, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117300572.

Shapiro:2018:RR

[Sha18] Lawrence Shapiro. Reduction redux. Studies in History and Philosophy of Science Part A, 68(??):10–19, ???? 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302704.

Shaw:2020:RAR

[Sha20] Jamie Shaw. The revolt against rationalism: Feyerabend’s critical philosophy. Studies in History and Philosophy of Science Part A, 80(??):110–122, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118302863.

Shan:2021:BMB

[Sha21] Yafeng Shan. Beyond Mendelism and biometry. Studies in History and Philosophy of Science Part A, 89(??):155–163, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001308.

Shaw:2022:VIP

[Sha22] Jamie Shaw. On the very idea of pursuitworthiness. Studies in History and Philosophy of Science Part A, 91(??):103– 112, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812100193X.

Squires:1994:NLA

[SHB94] E. J. Squires, L. Hardy, and H. R. Brown. Non-locality from an analogue of the quantum Zeno effect. Studies in History and Philosophy of Science Part A, 25(3):425–435, June 1994. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368194900604.

Sherry:1987:WBA

[She87] David Sherry. The wake of Berkeley’s analyst: Rigor mathematicae? Studies in History and Philosophy of Science Part A, 18 (4):455–480, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368187900033.

Shelton:1988:ROS

[She88] Jim Shelton. The role of observation and simplicity in Einstein’s epistemology. Studies in History and Philosophy of Science Part A, 19(1):103–118, March 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368188900210.

Sherry:1991:LIQ

[She91] David Sherry. The logic of impossible quantities. Studies in History and Philosophy of Science Part A, 22(1):37–62, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190014J.

Sherry:1993:DTM

[She93] David Sherry. Don’t take me half the way: On Berkeley on mathematical reasoning. Studies in History and Philosophy of Science Part A, 24(2):207–225, June 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390046M.

Sherry:1997:ME

[She97] David Sherry. On mathematical error. Studies in History and Philosophy of Science Part A, 28(3):393–416, September 1997. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368196000246.

Sherry:1999:TSP

[She99] David Sherry. Thales’s sure path. Studies in History and Philosophy of Science Part A, 30(4):621–650, December 1999. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368199000199.

Sherry:2006:MRI

[She06] David Sherry. Mathematical reasoning: induction, deduction and beyond. Studies in History and Philosophy of Science Part A, 37 (3):489–504, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000203.

Sherry:2011:TTF

[She11] David Sherry. Thermoscopes, thermometers, and the foundations of measurement. Studies in History and Philosophy of Science Part A, 42(4):509–524, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111000616.

Shea:2022:RCS

[She22] Nicholas Shea. Representation in Cognitive Science by Nicholas Shea: Reply by the author. Studies in History and Philosophy of Science Part A, 92(??):270–273, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000765.

Shrager:2007:EBC

[Shr07] Jeff Shrager. The evolution of BioBike: Community adaptation of a biocomputing platform. Studies in History and Philosophy of Science Part A, 38(4):642–656, December 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000672.

Sibum:1995:RMV

[Sib95] Heinz Otto Sibum. Reworking the mechanical value of heat: Instruments of precision and gestures of accuracy in early Victorian England. Studies in History and Philosophy of Science Part A, 26(1):73–106, March 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368194000369.

Sibum:2008:SCS

[Sib08] H. Otto Sibum. Science and the changing senses of reality circa 1900. Studies in History and Philosophy of Science Part A, 39 (3):295–297, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810800054X.

Siegel:1983:TPS

[Sie83] Harvey Siegel. Truth, problem solving and the rationality of science. Studies in History and Philosophy of Science Part A, 14(2): 89–112, June 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368183900158.

Siegel:1989:PSN

[Sie89] Harvey Siegel. Philosophy of science naturalized? Some problems with Giere’s naturalism. Studies in History and Philosophy of Science Part A, 20(3):365–375, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900125.

Siegel:1990:LNN

[Sie90] Harvey Siegel. Laudan’s normative naturalism. Studies in History and Philosophy of Science Part A, 21(2):295–313, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368190900287.

Siegel:1998:HRR

[Sie98] Harvey Siegel. Hooker’s revolutionary regulatory realism. Studies in History and Philosophy of Science Part A, 29(1):129– 141, March 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368197000162.

Sieroka:2000:OWT

[Sie00] Norman Sieroka. One Whitehead, not three. Studies in History and Philosophy of Science Part A, 31(4):721–730, December 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000297.

Siegel:2004:BRB

[Sie04] Harvey Siegel. Book review: The bearing of philosophy of science on science education, and vice versa: the case of constructivism: Constructivism in science education: a philosophical examination, Michael R. Matthews (Ed.); Dordrecht: Kluwer, 1998, pp. xii + 234, Price US$98.00 £59.00 NLG180.00 hardback, ISBN 0-7923-5033-2; US$39.00 paperback, ISBN 0-7923-4924-5. Studies in History and Philosophy of Science Part A, 35(1):185– 198, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000876. See [Mat98].

Sieroka:2007:WAT

[Sie07] Norman Sieroka. Weyl’s ‘agens theory’ of matter and the Zürich Fichte. Studies in History and Philosophy of Science Part A, 38(1): 84–107, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106001105.

Sieroka:2010:USK

[Sie10] Norman Sieroka. Umgebungen: symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus. (German) [Environments: symbolic constructivism following Hermann Weyl and Fritz Medicus], volume 8 of Legierungen. Chronos Verlag, Zürich, Switzerland, 2010. ISBN 3-0340-1006-0. 411 pp. LCCN ???? EUR

43.00; SFR 58.00. URL http://www.chronos-verlag.ch/php/
book_latest-new.php?book=978-3-0340-1006-1.

Sigrist:1990:OSP

[Sig90] René Sigrist. Les origines de la Société de Physique et d’Histoire Naturelle (1790–1822): la science genevoise face au modèle français. Société de physique et d’histoire naturelle, Geneva, Switzerland, 1990. ISSN 0252-7960. viii + 236 pp. LCCN ????

Simpson:1970:SOM

[Sim70] Thomas K. Simpson. Some observations on Maxwell’s treatise on electricity and magnetism: On the role of the ‘dynamical theory of the electromagnetic field’ in part IV of the treatise. Studies in History and Philosophy of Science Part A, 1(3):249–263, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368170900129.

Simon:2001:SSH

[Sim01] Herbert Simon. On simulating Simon: His monomania, and its sources in bounded rationality. Studies in History and Philosophy of Science Part A, 32(3):501–505, September 2001. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368101000139. See [Sen01].

Simpson:2014:OAC

[Sim14] William M. R. Simpson. Ontological aspects of the Casimir effect. Studies in History and Philosophy of Science Part A, 48 (part A) (??):84–88, November 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1355219814000859.

Simons:2017:MET

[Sim17] Massimiliano Simons. The many encounters of Thomas Kuhn and French epistemology. Studies in History and Philosophy of Science Part A, 61(??):41–50, February 2017. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300693.

Simons:2021:SBT

[Sim21] Massimiliano Simons. Synthetic biology as a technoscience: the case of minimal genomes and essential genes. Studies in History and Philosophy of Science Part A, 85(??):127–136, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301849.

Skeels:2020:TTN

[Ske20] Patrick Skeels. A tale of two Nortons. Studies in History and Philosophy of Science Part A, 83(??):28–35, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119300780.

Sklar:1992:PP

[Skl92] Lawrence Sklar. Philosophy of physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-875138-9 (paperback). xi + 246 pp. LCCN QC6 .S579 1992.

Svard:2021:FAR

[SL21] Per-Anders Svärd and Helena Tinnerholm Ljungberg. Fetal and animal research in Sweden: the construction of viable lives in regulatory policy debates, 1970–1980. Studies in History and Philosophy of Science Part A, 89(??):248–256, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812100128X.

Slater:2002:IRR

[Sla02] Leo B. Slater. Instruments and rules: R. B. Woodward and the tools of twentieth-century organic chemistry. Studies in History and Philosophy of Science Part A, 33(1):1–33, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368101000243.

Slater:2017:PPO

[Sla17] Matthew H. Slater. Pluto and the platypus: an odd ball and an odd duck — on classificatory norms. Studies in History and Philosophy of Science Part A, 61(??):1–10, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300018.

Slavov:2022:KIE

[Sla22] Matias Slavov. Kaila’s interpretation of Einstein–Minkowski invariance theory. Studies in History and Philosophy of Science Part A, 93(??):57–65, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000474.

Shan:2023:HWK

[SLH23] Yafeng Shan, Ehud Lamm, and Oren Harman. ‘History will be kind to me’: an introduction to new directions in the historiography of genetics. Studies in History and Philosophy of Science Part A, 99(??):A1–A3, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001479.

Sloan:1977:DSR

[Slo77] Phillip R. Sloan. Descartes, the sceptics, and the rejection of vitalism in seventeenth-century physiology. Studies in History and Philosophy of Science Part A, 8(1):1–28, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900164.

Slowik:2003:CRG

[Slo03] Edward Slowik. Conventionalism in Reid’s ‘geometry of visibles’. Studies in History and Philosophy of Science Part A, 34(3):467– 489, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368103000517.

Sloan:2019:LSN

[Slo19] Phillip R. Sloan. Life science and Naturphilosophie: Rethinking the relationship. Studies in History and Philosophy of Science Part A, 76(??):98–100, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119300196.

Sloane:2024:BFP

[Slo24] Mona Sloane. Book forum: Pushing Cool by Keith Wailoo: Sticky theories of race, markets, and innovation. Studies in History and Philosophy of Science Part A, 103(??):1–2, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001413.

Serpico:2023:NHP

[SLP23] Davide Serpico, Kate E. Lynch, and Theodore M. Porter. New historical and philosophical perspectives on quantitative genetics. Studies in History and Philosophy of Science Part A, 97(??):29– 33, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001765.

Sturm:2024:KMU

[SM24] Thomas Sturm and Rudolf Meer. Kant on the many uses of reason in the sciences: a neglected topic. Studies in History and Philosophy of Science Part A, 106(??):54–59, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000645.

Smith:1983:HSE

[Smi83] C. U. M. Smith. Herbert Spencer’s epigenetic epistemology. Studies in History and Philosophy of Science Part A, 14(1):1– 22, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818390002X.

Smith:1984:PCS

[Smi84] Joseph Wayne Smith. Primitive classification and the sociology of knowledge: a response to Bloor. Studies in History and Philosophy of Science Part A, 15(3):237–243, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368184900190. See reply [Blo84].

Smith:1988:ISR

[Smi88] Joel M. Smith. Inconsistency and scientific reasoning. Studies in History and Philosophy of Science Part A, 19(4):429–445, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900106.

Smith:2013:KPM

[Smi13a] Sheldon Smith. Kant’s picture of monads in the Physical Monadology. Studies in History and Philosophy of Science Part A, 44(1):102–111, March 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112001173.

Smith:2013:DKP

[Smi13b] Sheldon R. Smith. Does Kant have a pre-Newtonian picture of force in the balance argument? An account of how the balance argument works. Studies in History and Philosophy of Science Part A, 44 (3):470–480, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936811200088X.

Smith:2015:SLP

[Smi15] A. Mark Smith. From sight to light: the passage from ancient to modern optics. University of Chicago Press, Chicago, IL, USA and London, UK, 2015. ISBN 0-226-17476-X (hardcover), 0-226-17493-X (e-book). xi + 457 pp. LCCN QC352 .S57 2014.

Smith:2020:ETP

[Smi20] Eden T. Smith. Examining tensions in the past and present uses of concepts. Studies in History and Philosophy of Science Part A, 84(??):84–94, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301618.

Snobelen:2004:BRB

[Sno04a] Stephen David Snobelen. Book review: Newton and religion: context, nature and influence: James E. Force, & Richard H. Popkin (Eds.); International Archives of the History of Ideas; Kluwer Academic, Dordrecht, 1999, pp. xvii + 325, Price $169.00, ISBN 0-7923-5744-2. Studies in History and Philosophy of Science Part A, 35(3):674–680, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000615. See [FP99].

Snobelen:2004:WWI

[Sno04b] Stephen David Snobelen. William Whiston, Isaac Newton and the crisis of publicity. Studies in History and Philosophy of Science Part A, 35(3):573–603, September 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000561.

Snyder:2007:LOR

[Sny07] Laura J. Snyder. ‘Lord only of the ruffians and fiends’? William Whewell and the plurality of worlds debate. Studies in History and Philosophy of Science Part A, 38(3):584–592, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000404.

Segundo-Ortin:2019:PCR

[SOC19] Miguel Segundo-Ortin and Paco Calvo. Are plants cognitive? A reply to Adams. Studies in History and Philosophy of Science Part A, 73(??):64–71, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368118302383. See [Ada18].

Solomon:1995:LNS

[Sol95] Miriam Solomon. Legend naturalism and scientific progress: An essay on Philip Kitcher’s The advancement of science. Studies in History and Philosophy of Science Part A, 26(2):205– 218, June 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000036. See [Kit93].

Soler:2008:ROS

[Sol08a] Léna Soler. Are the results of our science contingent or inevitable? Studies in History and Philosophy of Science Part A, 39(2):221– 229, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000241.

Soler:2008:RAS

[Sol08b] Léna Soler. Revealing the analytical structure and some intrinsic major difficulties of the contingentist/inevitabilist issue. Studies in History and Philosophy of Science Part A, 39(2):230– 241, June 2008. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368108000253.

Solovey:2022:BRS

[Sol22] Mark Solovey. Book review: Social Science for What? Battles over Public Funding for the “Other Sciences” at the National Science Foundation by Mark Solovey: Reply by the Author. Studies in History and Philosophy of Science Part A, 94(??):210– 211, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000425. See [Lin22a, Rot22,

Hau22, Tur22b].

Sorgner:2016:CEP

[Sor16] Helene Sorgner. Challenging expertise: Paul Feyerabend vs. Harry Collins & Robert Evans on democracy, public participation and scientific authority: Paul Feyerabend vs. Harry Collins & Robert Evans on scientific authority and public participation. Studies in History and Philosophy of Science Part A, 57(??):114– 120, June 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001661.

Shech:2021:I

[SP21] Elay Shech and Wendy S. Parker. Introduction. Studies in History and Philosophy of Science Part A, 85(??):30–33, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120302041.

Spada:1990:SFH

[Spa90] Marina Frasca Spada. Some features of Hume’s conception of space. Studies in History and Philosophy of Science Part A, 21 (3):371–411, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090002P.

Spary:1991:BRB

[Spa91] Emma Spary. Book review: Les Origines de la Société de Physique et d’Histoire Naturelle (1790–1822): La Science Genevoise Face au Modèle Français: René Sigrist: Mémoires de la Société de Physique et d’Histoire Naturelle de Genève, Vol. 45, Bicentenary volume (Geneva: Société de Physique et d’Histoire Naturelle de Gèneve, 1990), 236 pp. Paperback ISSN 0252-7960. Studies in History and Philosophy of Science Part A, 22(3):533–538, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190009H. See [Sig90].

Spary:1995:CC

[Spa95] Emma Spary. Colonising cultures. Studies in History and Philosophy of Science Part A, 26(4):649–656, December 1995. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368195000283.

Spagnesi:2023:RIK

[Spa23] Lorenzo Spagnesi. Regulative idealization: a Kantian approach to idealized models. Studies in History and Philosophy of Science Part A, 99(??):1–9, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368123000420.

Spencer:2004:DNR

[Spe04] Quayshawn Spencer. Do Newton’s rules of reasoning guarantee truth … must they? Studies in History and Philosophy of Science Part A, 35(4):759–782, December 2004. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368104000858.

Spelda:2017:HSP

[Spe17] Daniel Spelda. The history of science as the progress of the human spirit: the historiography of astronomy in the eighteenth century. Studies in History and Philosophy of Science Part A, 63(??): 48–57, June 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811630111X.

Sprevak:2005:CC

[Spr05] Mark D. Sprevak. The Chinese carnival. Studies in History and Philosophy of Science Part A, 36(1):203–209, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104001293.

Sprevak:2010:CCS

[Spr10a] Mark Sprevak. Computation and cognitive science. Studies in History and Philosophy of Science Part A, 41(3):223–226, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000439.

Sprevak:2010:CIR

[Spr10b] Mark Sprevak. Computation, individuation, and the received view on representation. Studies in History and Philosophy of Science Part A, 41(3):260–270, September 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000403.

Sprevak:2010:IHE

[Spr10c] Mark Sprevak. Inference to the hypothesis of extended cognition. Studies in History and Philosophy of Science Part A, 41(4):353– 362, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000658.

Solomon:2005:CCL

[SR05] Miriam Solomon and Alan Richardson. A critical context for Longino’s critical contextual empiricism. Studies in History and Philosophy of Science Part A, 36(1):211–222, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810400130X.

Seselja:2009:KCE

[SS09] Dunja Seselja and Christian Straßer. Kuhn and coherentist epistemology. Studies in History and Philosophy of Science Part A, 40 (3):322–327, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000466. See reply [Kuu09].

Schurz:2020:MTO

[ST20] Gerhard Schurz and Paul Thorn. The material theory of object-induction and the universal optimality of meta-induction: Two complementary accounts. Studies in History and Philosophy of Science Part A, 82(??):88–93, August 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119301281.

Staley:2008:WPE

[Sta08] Richard Staley. Worldviews and physicists’ experience of disciplinary change: on the uses of ‘classical’ physics. Studies in History and Philosophy of Science Part A, 39(3):298–311, September 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108000551.

Stanley:2011:HSS

[Sta11] Matthew Stanley. How scientists stopped talking about science. Studies in History and Philosophy of Science Part A, 42(1):235– 239, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000944.

Stan:2013:KTL

[Sta13] Marius Stan. Kant’s third law of mechanics: the long shadow of Leibniz. Studies in History and Philosophy of Science Part A, 44 (3):493–504, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000891.

Statham:2018:MIT

[Sta18] Georgie Statham. Mechanisms, the interventionist theory, and the ability to use causal relationships. Studies in History and Philosophy of Science Part A, 72(??):22–31, December 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936811730242X.

Steigerwald:2003:DRE

[Ste03] Joan Steigerwald. The dynamics of reason and its elusive object in Kant, Fichte and Schelling. Studies in History and Philosophy of Science Part A, 34(1):111–134, March 2003. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368102000900.

Stewart:2004:BRB

[Ste04] Ian G. Stewart. Book review: The principia: mathematical principles of natural philosophy: Isaac Newton; a new translation by I. Bernard Cohen and Anne Whitman; with a guide to Newton’s Principia by I. Bernard Cohen; University of California Press, Berkeley, Los Angeles, & London, 1999, pp. 1025, Price £60.00 US$75.00 hardback, ISBN 0-520-08816-6, Price £24.95 US$35.00 paperback, ISBN 0-520-08817-4. Studies in History and Philosophy of Science Part A, 35(3):665–667, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104000603. See [New87, NCW99, New09].

Sterrett:2005:PSW

[Ste05] Susan G. Sterrett. Pictures of sounds: Wittgenstein on gramophone records and the logic of depiction. Studies in History and Philosophy of Science Part A, 36(2):351–362, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000191.

Sterrett:2014:MMM

[Ste14] S. G. Sterrett. The morals of model-making. Studies in History and Philosophy of Science Part A, 46(??):31–45, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113001283.

Steel:2015:AVP

[Ste15] Daniel Steel. Acceptance, values, and probability. Studies in History and Philosophy of Science Part A, 53(??):81–88, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000709.

Stemeroff:2021:SCM

[Ste21] Noah Stemeroff. Structuralism and the conformity of mathematics and nature. Studies in History and Philosophy of Science Part A, 86(??):84–92, April 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000042.

Stegenga:2022:EE

[Ste22a] Jacob Stegenga. Evidence of effectiveness. Studies in History and Philosophy of Science Part A, 91(??):288–295, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000012.

Stegenga:2022:RHA

[Ste22b] Jacob Stegenga. Red herrings about relative measures: a response to Hoefer and Krauss. Studies in History and Philosophy of Science Part A, 92(??):56–59, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000048. See [HK21].

Stenhouse:2022:RDD

[Ste22c] John Stenhouse. Reading Darwin during the new Zealand wars: Science, religion, politics and race, 1835-1900. Studies in History and Philosophy of Science Part A, 96(??):87–99, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001340.

Stemeroff:2024:NMS

[Ste24a] Noah Stemeroff. The notorious man-in-the-street: Hermann Weyl and the problem of knowledge. Studies in History and Philosophy of Science Part A, 104(??):48–60, April 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000153.

Stephens:2024:MDR

[Ste24b] Christopher Stephens. Modus Darwin redux. Studies in History and Philosophy of Science Part A, 108(??):73–83, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124001419.

Stillwell:2013:GCH

[Sti13] Devon Stillwell. Genetic counseling in historical perspective: Understanding our hereditary past and forecasting our genomic future. Studies in History and Philosophy of Science Part A, 44 (4):618–622, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000939.

Stotz:2009:EPB

[Sto09a] Karola Stotz. Experimental philosophy of biology: notes from the field. Studies in History and Philosophy of Science Part A, 40(2): 233–237, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368109000272.

Stotz:2009:PTN

[Sto09b] Karola Stotz. Philosophy in the trenches: from naturalized to experimental philosophy (of science). Studies in History and Philosophy of Science Part A, 40(2):225–226, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000284.

Stone:2010:SIC

[Sto10] Abraham D. Stone. On the sources and implications of Carnap’s Der Raum. Studies in History and Philosophy of Science Part A, 41(1):65–74, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000910. See [Car21].

Stoll:2012:GAX

[Sto12] Oliver Stoll. For the glory of Athens: Xenophon’s Hipparchikos <Logos>, a technical treatise and instruction manual on ideal leadership. Studies in History and Philosophy of Science Part A, 43(2): 250–257, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111001348.

Stoltzner:2015:HAM

[Stö15] Michael Stöltzner. Hilbert’s axiomatic method and Carnap’s general axiomatics. Studies in History and Philosophy of Science Part A, 53(??):12–22, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115000655.

Stojanovic:2023:PUR

[Sto23] Milutin Stojanovic. Pursuitworthiness in urgent research: Lessons on well-ordered science from sustainability science. Studies in History and Philosophy of Science Part A, 98(??):49–61, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000043.

Storni:2025:RNP

[Sto25] Marco Storni. Resisting Newton in provincial France, 1750s–1770s: Opposition from the margins to the Parisian academic community. Studies in History and Philosophy of Science Part A, 109(??):21– 30, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001596.

Strevens:2006:RME

[Str06] Michael Strevens. The role of the Matthew effect in science. Studies in History and Philosophy of Science Part A, 37(2):159– 170, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000252.

Strevens:2013:NUE

[Str13] Michael Strevens. No understanding without explanation. Studies in History and Philosophy of Science Part A, 44(3):510–515, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811200115X.

Strevens:2021:PIP

[Str21] Michael Strevens. Permissible idealizations for the purpose of prediction. Studies in History and Philosophy of Science Part A, 85 (??):92–100, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301813.

Stump:1989:HPP

[Stu89] David Stump. Henri Poincaré’s philosophy of science. Studies in History and Philosophy of Science Part A, 20(3):335–363, September 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368189900113.

Stump:1991:FNT

[Stu91a] David Stump. Fallibilism, naturalism and the traditional requirements for knowledge. Studies in History and Philosophy of Science Part A, 22(3):451–469, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819190004C.

Sturdy:1991:GNE

[Stu91b] Steve Sturdy. The germs of a new enlightenment. Studies in History and Philosophy of Science Part A, 22(1):163–173, March 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190019O.

Stump:2001:HST

[Stu01] James B. Stump. History of science through Koyré’s lenses. Studies in History and Philosophy of Science Part A, 32(2):243– 263, June 2001. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368101000115.

Stump:2007:PDV

[Stu07] David J. Stump. Pierre Duhem’s virtue epistemology. Studies in History and Philosophy of Science Part A, 38(1):149– 159, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106001075.

Sturm:2008:WDK

[Stu08] Thomas Sturm. Why did Kant reject physiological explanations in his anthropology? Studies in History and Philosophy of Science Part A, 39(4):495–505, December 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000836.

Sturm:2012:WPA

[Stu12] Thomas Sturm. What’s philosophical about kant’s philosophy of the human sciences? Studies in History and Philosophy of Science Part A, 43(1):203–207, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001245.

Stuart:2016:TTT

[Stu16] Michael T. Stuart. Taming theory with thought experiments: Understanding and scientific progress. Studies in History and Philosophy of Science Part A, 58(??):24–33, August 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116300024.

Stump:2017:SPM

[Stu17] David J. Stump. Scientific pluralism and metaphysics. Studies in History and Philosophy of Science Part A, 64(??):64– 66, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117301838.

Stuart:2020:MTI

[Stu20] Michael T. Stuart. The material theory of induction and the epistemology of thought experiments. Studies in History and Philosophy of Science Part A, 83(??):17–27, October 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368119301141.

Sarkar:2015:IFE

[SU15] Sahotra Sarkar and Thomas Uebel. Introduction: Formal epistemology and the legacy of logical empiricism. Studies in History and Philosophy of Science Part A, 53(??):1–2, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000631.

Suarez:2012:AMM

[Suá12a] Mauricio Suárez. The ample modelling mind. Studies in History and Philosophy of Science Part A, 43(1):213–217, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000811.

Suarez:2012:SPP

[Suá12b] Mauricio Suárez. Science, philosophy and the a priori. Studies in History and Philosophy of Science Part A, 43(1):1–6, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000823.

Suarez:2015:DRI

[Suá15] Mauricio Suárez. Deflationary representation, inference, and practice. Studies in History and Philosophy of Science Part A, 49 (??):36–47, February 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368114001095.

Suchting:1990:HHP

[Suc90] W. A. Suchting. Hegel and the Humean problem of induction. Studies in History and Philosophy of Science Part A, 21(3):493– 510, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819090007U.

Sumner:2013:WRI

[Sum13] James Sumner. Walls of resonance: Institutional history and the buildings of the University of Manchester. Studies in History and Philosophy of Science Part A, 44(4):700–715, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000563.

Sutton:1971:JFD

[Sut71] M. A. Sutton. J. F. Daniell and the Boscovichean atom. Studies in History and Philosophy of Science Part A, 1(4):277–292, February 1971. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368171900094.

Smith:1989:EEB

[SW89] Crosbie Smith and M. Norton Wise. Energy and empire: a biographical study of Lord Kelvin. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-26173-2. xxvi + 866 pp.

LCCN QC16.K3 S65 1989. URL http://www.loc.gov/catdir/
description/cam041/88025685.html;   http://www.loc.gov/
catdir/toc/cam041/88025685.html.

Scerri:2001:PPT

[SW01] Eric R. Scerri and John Worrall. Prediction and the periodic table. Studies in History and Philosophy of Science Part A, 32(3): 407–452, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000231. See discussion

[Bar05, Sce05, Bar05, Wor05, Sch08b].

Snyder:2007:I

[SW07] Laura J. Snyder and Thomas P. Weber. Introduction. Studies in History and Philosophy of Science Part A, 38(3):567–569, September 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000416.

Seselja:2012:RIH

[SW12] Dunja Seselja and Erik Weber. Rationality and irrationality in the history of continental drift: Was the hypothesis of continental drift worthy of pursuit? Studies in History and Philosophy of Science Part A, 43(1):147–159, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111001099.

Swaim:2021:WNP

[Swa21] Daniel G. Swaim. What is narrative possibility? Studies in History and Philosophy of Science Part A, 89(??):257–266, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001291.

Swaim:2024:GHT

[Swa24] Daniel G. Swaim. Getting from here to there: the contingency of historical evidence and the value of speculation. Studies in History and Philosophy of Science Part A, 106(??):118– 125, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000591.

Swijtink:1986:DMC

[Swi86] Zeno G. Swijtink. D’Alembert and the maturity of chances. Studies in History and Philosophy of Science Part A, 17(3):327–349, September 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900129.

Swoboda:1974:BRB

[Swo74] Wolfram Swoboda. Book review: Ernst Mach: His life, work, and influence: John T. Blackmore, Berkeley: University of California Press, 1972. xx + 414 pages, illustrated. Studies in History and Philosophy of Science Part A, 5(2):187–201, August 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368174900259. See [Bla72].

Sznajder:2016:WCS

[Szn16] Marta Sznajder. What conceptual spaces can do for Carnap’s late inductive logic. Studies in History and Philosophy of Science Part A, 56(??):62–71, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368115001788.

Turner:2023:BRI

[TA23] Derek Turner and Ahmed AboHamad. Book review: The Instrument of Science: Scientific Anti-Realism Revitalised by Darrell Rowbottom: Revitalizing antirealism even more. Studies in History and Philosophy of Science Part A, 98(??):7–8, April 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000274.

Taborsky:2014:CSC

[Tab14]  Paul Taborsky. Is complexity a scientific concept?  Stud-

ies in History and Philosophy of Science Part A, 47(??):51–59, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000570.

Tal:2017:CMM

[Tal17] Eran Tal. Calibration: Modelling the measurement process. Studies in History and Philosophy of Science Part A, 65–66(??):33– 45, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117302108.

Talbot:2025:UAD

[Tal25] Chris Talbot. An unpublished article by David Bohm. Studies in History and Philosophy of Science Part A, 110(??):88– 103, April 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001420.

Tamminga:2005:ICC

[Tam05] Allard Tamminga. Introspection and change in Carnap’s logical behaviourism. Studies in History and Philosophy of Science Part A, 36(4):650–667, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810500083X.

Tambolo:2015:TTT

[Tam15] Luca Tambolo. A tale of three theories: Feyerabend and Popper on progress and the aim of science. Studies in History and Philosophy of Science Part A, 51(??):33–41, June 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115000266.

Tamborini:2020:TAD

[Tam20] Marco Tamborini. Technoscientific approaches to deep time. Studies in History and Philosophy of Science Part A, 79(??):57–67, February 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301705.

Tamborini:2023:ERB

[Tam23] Marco Tamborini. The elephant in the room: the biomimetic principle in bio-robotics and embodied AI. Studies in History and Philosophy of Science Part A, 97(??):13–19, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001741.

Tan:2021:III

[Tan21] Peter Tan. Inconsistent idealizations and inferentialism about scientific representation. Studies in History and Philosophy of Science Part A, 89(??):11–18, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001011.

Taschetto:2025:RQR

[Tas25] Diana Taschetto. Rewriting the quantum “Revolution”. Studies in History and Philosophy of Science Part A, 109(??):72– 88, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001614.

Taub:2009:SI

[Tau09] Liba Taub. On scientific instruments. Studies in History and Philosophy of Science Part A, 40(4):337–343, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000661.

Taylor:2009:MCS

[Tay09] Katie Taylor. Mogg’s celestial sphere (1813): the construction of polite astronomy. Studies in History and Philosophy of Science Part A, 40(4):360–371, December 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368109000697.

Tauber:1991:MOI

[TC91] Alfred I. Tauber and Leon Chernyak. Metchnikoff and the origins of immunology: from metaphor to theory. Monographs on the history and philosophy of biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506447-X. xviii + 247 pp. LCCN QR180.72.M48 T38 1991. URL http://www.loc.

gov/catdir/enhancements/fy0603/90026245-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0603/90026245-t.html.

Tambolo:2021:MDI

[TC21] Luca Tambolo and Gustavo Cevolani. Multiple discoveries, inevitability, and scientific realism. Studies in History and Philosophy of Science Part A, 90(??):30–38, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121001357.

Taschetto:2025:DDF

[TC25] Diana Taschetto and Ricardo Correa da Silva. The dual dynamical foundation of orthodox quantum mechanics. Studies in History and Philosophy of Science Part A, 109(??):89–105, February 2025. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001602.

Teira:2006:NDS

[Tei06] David Teira. On the normative dimension of the St. Petersburg paradox. Studies in History and Philosophy of Science Part A, 37(2):210–223, June 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000215.

Teicher:2022:KBT

[Tei22] Amir Teicher. Kristine Bonnevie’s theories on the genetics of fingerprints, and their application in Germany. Studies in History and Philosophy of Science Part A, 92(??):162–176, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000061.

Temple:1986:PTF

[Tem86] Dennis Temple. Pasteur’s theory of fermentation: a “virtual tautology”? Studies in History and Philosophy of Science Part A, 17 (4):487–503, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368186900063.

Terrall:2005:ER

[Ter05] Mary Terrall. An embarrassment of riches. Studies in History and Philosophy of Science Part A, 36(1):185–190, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810400127X.

Terrall:2017:NNH

[Ter17] Mary Terrall. Narrative and natural history in the eighteenth century. Studies in History and Philosophy of Science Part A, 62(??): 51–64, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300766.

Terzian:2021:CPI

[Ter21] Giulia Terzian. Chomsky in the playground: Idealization in generative linguistics. Studies in History and Philosophy of Science Part A, 87(??):1–12, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000212.

Fan:2018:CAP

[tF18] Fa ti Fan. Can animals predict earthquakes?: Bio-sentinels as seismic sensors in communist China and beyond. Studies in History and Philosophy of Science Part A, 70(??):50–57, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301419.

Twardy:2005:EDS

[TGD05] Charles Twardy, Steve Gardner, and David L. Dowe. Empirical data sets are algorithmically compressible: reply to McAllister? Studies in History and Philosophy of Science Part A, 36(2):391–402, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810500021X. See [McA03,

McA05].

terHark:1993:PPP

[tH93] Michel ter Hark. Problems and psychologism: Popper as the heir to Otto Selz. Studies in History and Philosophy of Science Part A, 24(4):585–609, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819390054N.

teHeesen:2000:BN

[tH00] Anke te Heesen. Boxes in nature. Studies in History and Philosophy of Science Part A, 31(3):381–403, September 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000170.

terHark:2002:BAR

[tH02] Michel ter Hark. Between autobiography and reality: Popper’s inductive years. Studies in History and Philosophy of Science Part A, 33(1):79–103, March 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000309.

Thagard:1977:DW

[Tha77] Paul R. Thagard. Darwin and Whewell. Studies in History and Philosophy of Science Part A, 8(4):353–356, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368177900267.

Thagard:1993:SMS

[Tha93] Paul Thagard. Societies of minds: Science as distributed computing. Studies in History and Philosophy of Science Part A, 24(1): 49–67, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390024E.

Thebault:2021:MT

[Thé21] Karim P. Y. Thébault. On Mach on time. Studies in History and Philosophy of Science Part A, 89(??):84–102, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001114.

Thompson:1983:SET

[Tho83] Paul Thompson. The structure of evolutionary theory: a semantic approach. Studies in History and Philosophy of Science Part A, 14 (3):215–229, September 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368183900134.

Thomsen:2009:SEC

[Tho09] Samuel W. Thomsen. Some evidence concerning the genesis of Shannon’s information theory. Studies in History and Philosophy of Science Part A, 40(1):81–91, March 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108001143.

Thomas:2015:HMD

[Tho15] Emily Thomas. Henry More and the development of absolute time. Studies in History and Philosophy of Science Part A, 54(??):11– 19, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811500076X.

Thrane:1975:POV

[Thr75] Gary Thrane. The proper object of vision. Studies in History and Philosophy of Science Part A, 6(1):3–41, April 1975. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0039368175900035.

Tiemersma:1988:MTA

[Tie88] Douwe Tiemersma. Methodological and theoretical aspects of Descartes’ treatise on the rainbow. Studies in History and Philosophy of Science Part A, 19(3):347–364, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368188900040.

Thomson-Jones:2016:IMS

[TJT16] Martin Thomson-Jones and Adam Toon. Introduction: Models and simulations 6. Studies in History and Philosophy of Science Part A, 56(??):111–112, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368116000066.

Toader:2024:BCP

[Toa24] Iulian D. Toader. Is Bohr’s correspondence principle just Hankel’s principle of permanence? Studies in History and Philosophy of Science Part A, 103(??):137–145, February 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001681.

Tomczyk:2022:DEP

[Tom22] Hannah Tomczyk. Did Einstein predict Bose–Einstein condensation? Studies in History and Philosophy of Science Part A, 93(??): 30–38, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000450.

Toon:2011:PM

[Too11] Adam Toon. Playing with molecules. Studies in History and Philosophy of Science Part A, 42(4):580–589, December 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000641.

Topper:1990:NNC

[Top90] David Topper. Newton on the number of colours in the spectrum. Studies in History and Philosophy of Science Part A, 21(2): 269–279, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368190900265.

Topham:2000:SPR

[Top00] Jonathan R. Topham. Scientific publishing and the reading of science in nineteenth-century Britain: a historiographical survey and guide to sources. Studies in History and Philosophy of Science Part A, 31(4):559–612, December 2000. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368100000303.

Topham:2004:TPM

[Top04] Jonathan R. Topham. Technicians of print and the making of natural knowledge. Studies in History and Philosophy of Science Part A, 35(2):391–400, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000238.

Torsell:2024:JHV

[Tor24] Christian Torsell. Janina Hosiasson and the value of evidence. Studies in History and Philosophy of Science Part A, 106(??):31– 36, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936812400061X.

Tosh:2003:ARE

[Tos03] Nick Tosh. Anachronism and retrospective explanation: in defence of a present-centred history of science. Studies in History and Philosophy of Science Part A, 34(3):647–659, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000529.

Tosh:2006:STH

[Tos06] Nick Tosh. Science, truth and history, Part I. Historiography, relativism and the Sociology of Scientific Knowledge. Studies in History and Philosophy of Science Part A, 37(4):675–701, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000811.

Tosh:2007:STH

[Tos07] Nick Tosh. Science, truth and history, part II. metaphysical bolt-holes for the Sociology of Scientific Knowledge? Studies in History and Philosophy of Science Part A, 38(1):185– 209, March 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368106000975.

Totelin:2012:EPN

[Tot12] Laurence M. V. Totelin. And to end on a poetic note: Galen’s authorial strategies in the pharmacological books. Studies in History and Philosophy of Science Part A, 43(2):307–315, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111001300.

Trappes:2022:IDU

[Tra22] Rose Trappes. Individual differences, uniqueness, and individuality in behavioural ecology. Studies in History and Philosophy of Science Part A, 96(??):18–26, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122001212.

Treitel:1987:CCN

[Tre87] Jonathan Treitel. Confirmation as competition: The necessity for dummy rival hypotheses. Studies in History and Philosophy of Science Part A, 18(4):517–525, December 1987. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368187900069.

Tresch:2007:DFF

[Tre07] John Tresch. The daguerreotype’s first frame: François Arago’s moral economy of instruments. Studies in History and Philosophy of Science Part A, 38(2):445–476, June 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000209.

Tribe:2005:OH

[Tri05] Keith Tribe. Oeconomic history. Studies in History and Philosophy of Science Part A, 36(3):586–597, September 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000531.

Trizio:2008:HMS

[Tri08] Emiliano Trizio. How many sciences for one world? contingency and the success of science. Studies in History and Philosophy of Science Part A, 39(2):253–258, June 2008. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108000277.

Trippett:2014:SLH

[Tri14] David Trippett. Sensations of listening in Helmholtz’s laboratory. Studies in History and Philosophy of Science Part A, 47(??):124– 132, September 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000028.

Trout:1994:RLB

[Tro94] J. D. Trout. A realistic look backward. Studies in History and Philosophy of Science Part A, 25(1):37–64, February 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900191.

Trout:1995:DTI

[Tro95] J. D. Trout. Diverse tests on an independent world. Studies in History and Philosophy of Science Part A, 26(3):407–429, September 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000047.

Truesdell:1968:EHM

[Tru68] C. (Clifford) Truesdell. Essays in the history of mechanics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. 383 pp. LCCN QC122 .T7.

Tucker:1993:THP

[Tuc93] Aviezer Tucker. A theory of historiography as a pre-science. Studies in History and Philosophy of Science Part A, 24(4):633– 667, October 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390056P.

Tucker:1995:IP

[Tuc95] Aviezer Tucker. The illness of psychoanalysis. Studies in History and Philosophy of Science Part A, 26(4):657–665, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368195000151.

Tucker:2012:NVR

[Tuc12] Aviezer Tucker. Nullius in verba: Recent studies in the epistemology of testimony. Studies in History and Philosophy of Science Part A, 43(2):412–419, June 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368112000209.

Tucker:2020:ICC

[Tuc20] Aviezer Tucker. The inferences of common causes reduced to common origins. Studies in History and Philosophy of Science Part A, 81(??):105–115, June 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119300706.

Tulodziecki:2016:SRB

[Tul16] Dana Tulodziecki. Structural realism beyond physics. Studies in History and Philosophy of Science Part A, 59(??):106– 114, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300279.

Turner:2004:PVT

[Tur04] Derek D. Turner. The past vs. the tiny: historical science and the abductive arguments for realism. Studies in History and Philosophy of Science Part A, 35(1):1–17, March 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368103000980.

Turner:2005:MOA

[Tur05a] Derek D. Turner. Misleading observable analogues in paleontology. Studies in History and Philosophy of Science Part A, 36(1): 175–183, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104001268.

Turner:2005:NAW

[Tur05b] Stephen P. Turner. Normative all the way down. Studies in History and Philosophy of Science Part A, 36(2):419–429, June 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000245.

Turner:2011:STK

[Tur11] Stephen P. Turner. Starting with tacit knowledge, ending with Durkheim? Studies in History and Philosophy of Science Part A, 42(3):472–476, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368111000215.

Turner:2013:WEE

[Tur13] Stephen Turner. Where explanation ends: Understanding as the place the spade turns in the social sciences. Studies in History and Philosophy of Science Part A, 44(3):532–538, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112001112.

Turner:2016:SLC

[Tur16] Derek D. Turner. A second look at the colors of the dinosaurs. Studies in History and Philosophy of Science Part A, 55(??):60– 68, February 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001053.

Turnbull:2021:RTM

[Tur21] Margaret Greta Turnbull. The relativity of theory by Moti Mizrahi: Pandemics and pathogens: What’s at stake in the debate over scientific realism?. Studies in History and Philosophy of Science Part A, 87(??):168–169, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000479.

Turnbull:2022:TFG

[Tur22a] Paul Turnbull. ‘Thrown into the fossil gap’: Indigenous Australian ancestral bodily remains in the hands of early Darwinian anatomists, c. 1860–1916. Studies in History and Philosophy of Science Part A, 92(??):1–11, April 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121002107.

Turner:2022:BRS

[Tur22b] Stephen Turner. Book review: Social science for what? Battles over public Funding for the “other sciences ” at the National Science Foundation by Mark Solovey: NSF’s unhappy legacy in American social science. Studies in History and Philosophy of Science Part A, 94(??):208–209, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122000413. See [Lin22a,

Rot22, Hau22] and reply [Sol22].

teVrugt:2021:FPI

[tV21] Michael te Vrugt. The five problems of irreversibility. Studies in History and Philosophy of Science Part A, 87(??):136– 146, June 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121000509.

Tybjerg:2003:WMP

[Tyb03] Karin Tybjerg. Wonder-making and philosophical wonder in Hero of Alexandria. Studies in History and Philosophy of Science Part A, 34(3):443–466, September 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368103000530.

Tybjerg:2022:SHM

[Tyb22] Karin Tybjerg. Scale in the history of medicine. Studies in History and Philosophy of Science Part A, 91(??):221–233, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001394.

Uebel:1991:NPN

[Ueb91a] Thomas E. Uebel. Neurath’s programme for naturalistic epistemology. Studies in History and Philosophy of Science Part A, 22 (4):623–646, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819190036R.

Uebel:1991:RFV

[Ueb91b] Thomas E. (Thomas Ernst) Uebel, editor. Rediscovering the forgotten Vienna Circle: Austrian studies on Otto Neurath and the Vienna Circle, volume 133 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346. xii + 326 pp. LCCN Q174 .B67

vol. 133. URL https://link.springer.com/book/10.1007/978-
94-011-3182-7.

Uebel:1994:IBA

[Ueb94] Thomas E. Uebel. The importance of being Austrian. Studies in History and Philosophy of Science Part A, 25(4):631– 636, August 1994. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368194900515.

Uebel:2005:PPS

[Ueb05a] Thomas Uebel. Political philosophy of science in logical empiricism: the Left Vienna Circle. Studies in History and Philosophy of Science Part A, 36(4):754–773, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000762.

Uebel:2005:SDS

[Ueb05b] Thomas Uebel. The social dimension of scientific knowledge and its distinct echo in philosophy of science: Six responses to Mirowski. Studies in History and Philosophy of Science Part A, 36(4):723– 725, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368105000841.

Uebel:2009:NPS

[Ueb09] Thomas Uebel. Neurath’s protocol statements revisited: sketch of a theory of scientific testimony. Studies in History and Philosophy of Science Part A, 40(1):4–13, March 2009. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368108001076.

Uebel:2010:WRA

[Ueb10] Thomas Uebel. What’s right about Carnap, Neurath and the Left Vienna Circle thesis: a refutation. Studies in History and Philosophy of Science Part A, 41(2):214–221, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000130.

Uebel:2012:SRP

[Ueb12] Thomas Uebel. De-synthesizing the relative a priori. Studies in History and Philosophy of Science Part A, 43(1):7–17, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368111000835.

Uebel:2015:TCC

[Ueb15] Thomas Uebel. Three challenges to the complementarity of the logic and the pragmatics of science. Studies in History and Philosophy of Science Part A, 53(??):23–32, October 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936811500062X.

Uebel:2025:LEA

[Ueb25] Thomas Uebel. Logical empiricist anti-exceptionalism in its Austro–German context. Studies in History and Philosophy of Science Part A, 110(??):46–54, April 2025. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368125000172.

Upshur:2020:CMN

[UG20] Ross Upshur and Maya J. Goldenberg. Countering medical nihilism by reconnecting facts and values. Studies in History and Philosophy of Science Part A, 84(??):75–83, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812030162X.

Ullis:2024:DTA

[Ull24] Dzintra Ullis. Development and transfer of automated methods in neuroscience: the DADTA. Studies in History and Philosophy of Science Part A, 106(??):109–117, August 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124000669.

Upshur:2022:PPH

[Ups22] Ross Upshur. Book review: Philosophy of Population Health: Philosophy for a New Public Health Era by Sean Valles: Fundamentally correct. Studies in History and Philosophy of Science Part A, 91 (??):320–321, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121002053.

Usselman:2015:PIL

[Uss15] Melvyn C. Usselman. Pure intelligence: the life of William Hyde Wollaston. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2015. ISBN 0-226-24573-X (hardcover), 0-226-24587-X (e-book). xv + 413 pp. LCCN Q143.W795 U87 2015.

Varga:2024:MHP

[VABF24] Somogy Varga, Martin Marchmann Andersen, Anke Bueter, and Anna Paldam Folker. Mental health promotion and the positive concept of health: Navigating dilemmas. Studies in History and Philosophy of Science Part A, 105(??):32–40, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368124000414.

Vaesen:2011:FBD

[Vae11] Krist Vaesen. The functional bias of the dual nature of technical artefacts program. Studies in History and Philosophy of Science Part A, 42(1):190–197, March 2011. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000804.

Valente:2021:TST

[Val21] Giovanni Valente. Taking up statistical thermodynamics: Equilibrium fluctuations and irreversibility. Studies in History and Philosophy of Science Part A, 85(??):176–184, February 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812030193X.

Valles:2022:PPH

[Val22] Sean A. Valles. Book review: Philosophy of Population Health: Philosophy for a New Public Health Era by Sean Valles: Reply by the author. Studies in History and Philosophy of Science Part A, 91(??):322–323, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121002065.

VanDerVet:1979:OPA

[Van79] Paul Van Der Vet. Overdetermined problems and anomalies. Studies in History and Philosophy of Science Part A, 10(3):259–261, September 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368179900098.

VanBalen:1986:IJD

[Van86] Gerrit A. M. Van Balen. The influence of Johannsen’s discoveries on the constraint-structure of the Mendelian research program. an example of conceptual problem solving in evolutionary theory. Studies in History and Philosophy of Science Part A, 17 (2):175–204, June 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368186900257.

VanBrakel:1991:LBC

[Van91] J. Van Brakel. The limited belief in chance. Studies in History and Philosophy of Science Part A, 22(3):499–513, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819190006E.

VanSittert:1995:HIM

[Van95] Lance Van Sittert. ‘the handmaiden of industry’: Marine science and fisheries development in South Africa 1895–1939. Studies in History and Philosophy of Science Part A, 26(4):531–558, December 1995. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368195000143.

VanDamme:2004:RSE

[Van04a] Stephane Van Damme. Reason and sentiment: the enlightenment, golden age of the translation of the sciences? Studies in History and Philosophy of Science Part A, 35(2):383– 389, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000226.

VandenBroecke:2004:ARC

[Van04b] Steven Vanden Broecke. Astrological reform, Calvinism, and Cartesianism: Copernican astronomy in the Low Countries, 1550–1650. Studies in History and Philosophy of Science Part A, 35(2):363– 381, June 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368104000214.

VanDyck:2009:EFL

[Van09] Maarten Van Dyck. On the epistemological foundations of the law of the lever. Studies in History and Philosophy of Science Part A, 40(3):315–318, September 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000442. See response

[Pal09b].

Vanzo:2016:ESS

[Van16] Alberto Vanzo. Experiment and speculation in seventeenthcentury Italy: the case of Geminiano Montanari. Studies in History and Philosophy of Science Part A, 56(??):52–61, April 2016. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368115001612.

VanDamme:2018:BRC

[Van18] Stéphane Van Damme. Book review: A Companion to the history of science, Bernard Lightman (Ed.). John Wiley and Sons and Blackwell, Chichester (2016), xvi + 601 pp. ISBN-13: 978-1-118-62077-9. £120.00 (hardback). Studies in History and Philosophy of Science Part A, 69(??):90–96, June 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118300475.

Vasco:2000:ISV

[Vas00] Carlos E. Vasco. The illusions of scientists vs. the illusions of social epistemologists. Studies in History and Philosophy of Science Part A, 31(2):343–351, June 2000. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936810000008X.

Vasudevan:2018:CDC

[Vas18] Anubav Vasudevan. Chance, determinism and the classical theory of probability. Studies in History and Philosophy of Science Part A, 67(??):32–43, February 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301693.

vanDongen:2021:STE

[vD21] Jeroen van Dongen. String theory, Einstein, and the identity of physics: Theory assessment in absence of the empirical. Studies in History and Philosophy of Science Part A, 89(??):164– 176, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001023.

vandenBerg:2018:KSA

[vdB18] Hein van den Berg. Kant and the scope of analogy in the life sciences. Studies in History and Philosophy of Science Part A, 71 (??):67–76, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300146.

vandenBerg:2022:ALE

[vdB22] Hein van den Berg. Animal languages in eighteenth-century German philosophy and science. Studies in History and Philosophy of Science Part A, 93(??):72–81, June 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000541.

vandenBerg:2024:ETA

[vdB24] Hein van den Berg. Explanation, teleology, and analogy in natural history and comparative anatomy around 1800: Kant and Cuvier. Studies in History and Philosophy of Science Part A, 105(??): 109–119, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000451.

vandenBelt:1990:SEK

[vdBG90] Henk van den Belt and Bart Gremmen. Specificity in the era of Koch and Ehrlich: a generalized interpretation of Ludwik Fleck’s ‘serological’ thought style. Studies in History and Philosophy of Science Part A, 21(3):463–479, September 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819090005S.

vanderLinden:2022:IIC

[vdLBV22] Rik van der Linden, Timo Bolt, and Mario Veen. ‘If it can’t be coded, it doesn’t exist’. A historical-philosophical analysis of the new ICD-11 classification of chronic pain. Studies in History and Philosophy of Science Part A, 94(??):121–132, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812200084X.

vanderWall:2004:NRN

[vdW04] Ernestine G. E. van der Wall. Newtonianism and religion in the Netherlands. Studies in History and Philosophy of Science Part A, 35(3):493–514, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000512.

vanEnk:2014:BDP

[vE14] Steven J. van Enk. The Brandeis dice problem and statistical mechanics. Studies in History and Philosophy of Science Part A, 48 (part A)(??):1–6, November 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000914.

vanEck:2019:CRC

[vE19] Dingmar van Eck. Constitutive relevance in cognitive science: the case of eye movements and cognitive mechanisms. Studies in History and Philosophy of Science Part A, 73(??):44–53, February 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811730331X.

Veatch:1972:BRB

[Vea72] Henry Veatch. Book review: Aristotle’s theory of the syllogism: a logico-philosophical study of Book A of the prior analytics: Günther Patzig, New York: Humanities Press, 1969. xvii + 215 pp. $14.25. Studies in History and Philosophy of Science Part A, 2(4):369– 378, February 1972. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368172900052. See [Pat69].

vanEck:2016:MEC

[vEdJ16] Dingmar van Eck and Huib Looren de Jong. Mechanistic explanation, cognitive systems demarcation, and extended cognition. Studies in History and Philosophy of Science Part A, 59(??):11– 21, October 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116300140.

Ventura:2022:PBP

[Ven22] Rafael Ventura. Publish without bias or perish without replications. Studies in History and Philosophy of Science Part A, 96 (??):10–17, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368122001248.

Verdon:1982:DAS

[Ver82] Michel Verdon. Durkheim and Aristotle: Of some incongruous congruences. Studies in History and Philosophy of Science Part A, 13(4):333–352, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368182900188.

Vermaas:2006:PCE

[Ver06] Pieter E. Vermaas. The physical connection: engineering function ascriptions to technical artefacts and their components. Studies in History and Philosophy of Science Part A, 37(1):62–75, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810500110X.

Vermeir:2007:AKM

[Ver07a] Koen Vermeir. Athanasius Kircher’s magical instruments: an essay on ‘science’, ‘religion’ and applied metaphysics. Studies in History and Philosophy of Science Part A, 38(2):363–400, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000179.

Vertesi:2007:PMH

[Ver07b] Janet Vertesi. Picturing the Moon: Hevelius’s and Riccioli’s visual debate. Studies in History and Philosophy of Science Part A, 38(2):401–421, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368107000180.

Verhaegh:2017:QNS

[Ver17] Sander Verhaegh. Quine’s ‘needlessly strong’ holism. Studies in History and Philosophy of Science Part A, 61(??):11–20, February 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116301248.

Vetter:2018:ECK

[Vet18] Jeremy Vetter. Experiential and cosmopolitan knowledge: the transcontinental field practices of the U.S. Bureau of Biological Survey. Studies in History and Philosophy of Science Part A, 70 (??):6–17, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118301377.

Vezer:2016:CME

[Vez16] Martin A. Vezér. Computer models and the evidence of anthropogenic climate change: an epistemology of variety-of-evidence inferences and robustness analysis. Studies in History and Philosophy of Science Part A, 56(??):95–102, April 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116000054.

Vergara-Fernandez:2023:DMR

[VFHS23] Melissa Vergara-Fernández, Conrad Heilmann, and Marta Szymanowska. Describing model relations: the case of the capital asset pricing model (CAPM) family in financial economics. Studies in History and Philosophy of Science Part A, 97(??):91– 100, February 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001790.

Votsis:2014:INP

[VFS14] Ioannis Votsis, Ludwig Fahrbach, and Gerhard Schurz. Introduction: Novel predictions. Studies in History and Philosophy of Science Part A, 45(??):43–45, March 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113001003.

vanHelvoort:1991:WVC

[vH91] Ton van Helvoort. What is a virus? The case of tobacco mosaic disease. Studies in History and Philosophy of Science Part A, 22 (4):557–588, December 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819190034P.

Vermaas:2006:TFD

[VH06] Pieter E. Vermaas and Wybo Houkes. Technical functions: a drawbridge between the intentional and structural natures of technical artefacts. Studies in History and Philosophy of Science Part A, 37(1):5–18, March 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105001044.

Viciana:2021:ACW

[Vic21] Hugo Viciana. Animal culture: But of which kind? Studies in History and Philosophy of Science Part A, 90(??):208–218, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001680.

Verreault-Julien:2019:HCM

[VJ19] Philippe Verreault-Julien. How could models possibly provide how-possibly explanations? Studies in History and Philosophy of Science Part A, 73(??):22–33, February 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117301231.

Vaesen:2019:NSF

[VK19] Krist Vaesen and Joel Katzav. The National Science Foundation and philosophy of science’s withdrawal from social concerns. Studies in History and Philosophy of Science Part A, 78(??):73– 82, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811830116X.

vanLeeuwen:2024:MTS

[vL24] Robert van Leeuwen. From S-matrix theory to strings: Scattering data and the commitment to non-arbitrariness. Studies in History and Philosophy of Science Part A, 104(??):130– 149, April 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000207.

Vogel:2010:BES

[Vog10] Jonathan Vogel. BonJour on explanation and skepticism. Studies in History and Philosophy of Science Part A, 41(4):413–421, December 2010. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000634.

Vorms:2011:RIM

[Vor11] Marion Vorms. Representing with imaginary models: Formats matter. Studies in History and Philosophy of Science Part A, 42(2): 287–295, June 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110001159.

Vorms:2014:BCG

[Vor14] Marion Vorms. The birth of classical genetics as the junction of two disciplines: Conceptual change as representational change. Studies in History and Philosophy of Science Part A, 48(??):105– 116, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000624.

Voskuhl:2007:POP

[Vos07] Adelheid Voskuhl. Producing objects, producing texts: accounts of android automata in late eighteenth-century Europe. Studies in History and Philosophy of Science Part A, 38(2):422– 444, June 2007. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368107000192.

Votsis:2014:OCP

[Vot14] Ioannis Votsis. Objectivity in confirmation: Post hoc monsters and novel predictions. Studies in History and Philosophy of Science Part A, 45(??):70–78, March 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368113001052.

Votsis:2012:FTA

[VS12] Ioannis Votsis and Gerhard Schurz. A frame-theoretic analysis of two rival conceptions of heat. Studies in History and Philosophy of Science Part A, 43(1):105–114, March 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001014.

vanStrien:2014:OFL

[vS14] Marij van Strien. On the origins and foundations of Laplacian determinism. Studies in History and Philosophy of Science Part A, 45(??):24–31, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113001325.

vanStrien:2020:PAQ

[vS20] Marij van Strien. Pluralism and anarchism in quantum physics: Paul Feyerabend’s writings on quantum physics in relation to his general philosophy of science. Studies in History and Philosophy of Science Part A, 80(??):72–81, April 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118303364.

Wagner:2022:NOS

[Wag22] Wendy E. Wagner. No one solution to the “new demarcation problem”?: a view from the trenches. Studies in History and Philosophy of Science Part A, 92(??):177–185, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000322.

Wagner:2024:CSK

[Wag24] Gerhard Wagner. On the concept of systematization in the Kemeny–Oppenheim approach to intertheoretical reduction. Studies in History and Philosophy of Science Part A, 103(??):29– 38, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001437.

Wajzer:2021:IGE

[Waj21] Mateusz Wajzer. Idealisation, genetic explanations and political behaviours: Notes on the anti-reductionist critique of genopolitics. Studies in History and Philosophy of Science Part A, 90 (??):275–284, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001709.

Wall:1975:APH

[Wal75] Byron Emerson Wall. Anatomy of a precursor: the historiography of Aristarchos of Samos. Studies in History and Philosophy of Science Part A, 6(3):201–228, October 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368175900230.

Wallace:1991:GJM

[Wal91] William A. Wallace. Galileo, the Jesuits, and the medieval Aristotle, volume CS346 of Collected studies series. Ashgate, Aldershot, Hampshire, England, 1991. ISBN 0-86078-297-2. ???? pp. LCCN QB36.G2 W36 1991. US$87.95.

Wall:2006:JVO

[Wal06] Byron E. Wall. John Venn’s opposition to probability as degree of belief. Studies in History and Philosophy of Science Part A, 37 (4):550–561, December 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000756.

Walker:2012:KDI

[Wal12] David Walker. A Kuhnian defence of inference to the best explanation. Studies in History and Philosophy of Science Part A, 43(1): 64–73, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368111000719.

Waldow:2016:NHF

[Wal16] Anik Waldow. Natural history and the formation of the human being: Kant on active forces. Studies in History and Philosophy of Science Part A, 58(??):67–76, August 2016. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368116000248.

Wallace:2022:ISTa

[Wal22a] David Wallace. Isolated systems and their symmetries, part I: General framework and particle-mechanics examples. Studies in History and Philosophy of Science Part A, 92(??):239– 248, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000255.

Wallace:2022:ISTb

[Wal22b] David Wallace. Isolated systems and their symmetries, part II: Local and global symmetries of field theories. Studies in History and Philosophy of Science Part A, 92(??):249–259, April 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000267.

Wallace:2022:QGL

[Wal22c] David Wallace. Quantum gravity at low energies. Studies in History and Philosophy of Science Part A, 94(??):31–46, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000620.

Wallace:2024:GIT

[Wal24] David Wallace. Gauge invariance through gauge fixing. Studies in History and Philosophy of Science Part A, 108(??):38–45, December 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124001341.

Wang:2002:TTV

[Wan02] Xinli Wang. Taxonomy, truth-value gaps and incommensurability: a reconstruction of Kuhn’s taxonomic interpretation of incommensurability. Studies in History and Philosophy of Science Part A, 33 (3):465–485, September 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368101000395.

Wanderer:2012:HTS

[Wan12] Jeremy Wanderer. ‘The happy thought of a single man’: On the legendary beginnings of a style of reasoning. Studies in History and Philosophy of Science Part A, 43(4):640–648, December 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368112000489.

Wang:2022:FDM

[Wan22] Yafeng Wang. Feature dependence: a method for reconstructing actual causes in engineering failure investigations. Studies in History and Philosophy of Science Part A, 96(??):100–111, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001376.

Warwick:1989:IRE

[War89] Andrew Warwick. International relativity: The establishment of a theoretical discipline. Studies in History and Philosophy of Science Part A, 20(1):139–149, March 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368189900381.

Warwick:1992:CMC

[War92] Andrew Warwick. Cambridge mathematics and Cavendish physics: Cunningham, Campbell and Einstein’s Relativity 1905–1911: Part I: The uses of theory. Studies in History and Philosophy of Science Part A, 23(4):625–656, December 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819290015X.

Warwick:1993:CMC

[War93] Andrew Warwick. Cambridge mathematics and Cavendish physics: Cunningham, Campbell and Einstein’s Relativity 1905–1911: Part II: Comparing traditions in Cambridge physics. Studies in History and Philosophy of Science Part A, 24(1):1–25, March 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936819390022C.

Ward:2021:VLS

[War21] Zina B. Ward. On value-laden science. Studies in History and Philosophy of Science Part A, 85(??):54–62, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368120301783.

Warrier:2022:CVW

[War22] Niranjana Warrier. The case of the vanishing wavefunction. Studies in History and Philosophy of Science Part A, 96(??):135–140, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001558.

Ward:2023:EID

[War23] Zina B. Ward. Explaining individual differences. Studies in History and Philosophy of Science Part A, 101(??):61–70, October 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001097.

Watkins:1998:KJL

[Wat98] Eric Watkins. Kant’s justification of the laws of mechanics. Studies in History and Philosophy of Science Part A, 29(4):539–560, December 1998. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368198000429.

Watkins:2003:FCK

[Wat03] Eric Watkins. Forces and causes in Kant’s early pre-critical writings. Studies in History and Philosophy of Science Part A, 34(1): 5–27, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000912.

Waters:2004:WWC

[Wat04]  C. Kenneth Waters. What was classical genetics?  Stud-

ies in History and Philosophy of Science Part A, 35(4):783–809, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936810400086X.

Watson:2012:LLN

[Wat12] Joshua L. Watson. Leibniz on the laws of nature and the best deductive system. Studies in History and Philosophy of Science Part A, 43(4):577–584, December 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000519.

Watkins:2013:EKA

[Wat13] Eric Watkins. The early Kant’s (anti-)Newtonianism. Studies in History and Philosophy of Science Part A, 44(3):429–437, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368112000908.

Watkins:2021:MMA

[Wat21] Aja Watkins. Multi-model approaches to phylogenetics: Implications for idealization. Studies in History and Philosophy of Science Part A, 90(??):285–297, December 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936812100162X.

Watkins:2023:SPC

[Wat23] Aja Watkins. Scaling procedures in climate science: Using temporal scaling to identify a paleoclimate analogue. Studies in History and Philosophy of Science Part A, 102(??):31–44, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123001383.

Way:2021:NSA

[Way21] Amy Way. Natural selection and the ‘antiquity of man’: Intellectual impacts in the Australian colonies. Studies in History and Philosophy of Science Part A, 90(??):308–320, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001679.

Weinberger:2020:MSN

[WB20] Naftali Weinberger and Seamus Bradley. Making sense of nonfactual disagreement in science. Studies in History and Philosophy of Science Part A, 83(??):36–43, October 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368119300937.

Weinberg:2009:XPU

[WC09] Jonathan M. Weinberg and Stephen Crowley. The x-phi(les): unusual insights into the nature of inquiry. Studies in History and Philosophy of Science Part A, 40(2):227–232, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000296.

Weber:2007:CDP

[Web07] Thomas P. Weber. Carl du Prel (1839–1899): explorer of dreams, the soul, and the cosmos. Studies in History and Philosophy of Science Part A, 38(3):593–604, September 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000428.

Webster:2022:POD

[Web22] Colin Webster. Ptolemy’s Optics, double-vision, and the technological afterimage. Studies in History and Philosophy of Science Part A, 94(??):191–200, August 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368122000929.

Weckert:1986:TLO

[Wec86] John Weckert. The theory-ladenness of observations. Studies in History and Philosophy of Science Part A, 17(1):115– 127, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818690021X.

Wehrens:2019:ESS

[Weh19a] Rik Wehrens. Experimentation in the sociology of science: Representational and generative registers in the imitation game. Studies in History and Philosophy of Science Part A, 76(??):76– 85, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368118300608.

Wehrens:2019:IGR

[Weh19b] Rik Wehrens. The Imitation Game: Response to Collins and Evans. Studies in History and Philosophy of Science Part A, 76(??):91– 93, August 2019. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936811930072X.

Weinert:1999:TMC

[Wei99] Friedel Weinert. Theories, models and constraints. Studies in History and Philosophy of Science Part A, 30(2):303–333, June 1999. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936819800051X.

Weinel:2007:PSK

[Wei07] Martin Weinel. Primary source knowledge and technical decisionmaking: Mbeki and the AZT debate. Studies in History and Philosophy of Science Part A, 38(4):748–760, December 2007. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368107000684.

Weinert:2010:RPA

[Wei10a] Friedel Weinert. The role of probability arguments in the history of science. Studies in History and Philosophy of Science Part A, 41(1):95–104, March 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368109000892.

Weiskopf:2010:ECL

[Wei10b] Daniel A. Weiskopf. Embodied cognition and linguistic comprehension. Studies in History and Philosophy of Science Part A, 41 (3):294–304, September 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000373.

Weiskopf:2010:USR

[Wei10c] Daniel A. Weiskopf. Understanding is not simulating: a reply to Gibbs and Perlman. Studies in History and Philosophy of Science Part A, 41(3):309–312, September 2010. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368110000348.

Weinstein:2021:CRE

[Wei21a] Galina Weinstein. Coincidence and reproducibility in the EHT black hole experiment. Studies in History and Philosophy of Science Part A, 85(??):63–78, February 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368120301795.

Weinstein:2021:EBH

[Wei21b] Galina Weinstein. Is the EHT black hole experiment a new experiment in the guise of an old experiment? Studies in History and Philosophy of Science Part A, 88(??):41–49, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121000546.

Welbourne:2002:HRR

[Wel02] Michael Welbourne. Is Hume really a reductivist? Studies in History and Philosophy of Science Part A, 33(2):407–423, June 2002. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000092.

Wellman:2005:RLS

[Wel05] Kathleen Wellman. A rich life in science: the case of Pierre Louis Moreau de Maupertuis. Studies in History and Philosophy of Science Part A, 36(3):598–606, September 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000567.

Welch:2013:NTT

[Wel13] John R. Welch. New tools for theory choice and theory diagnosis. Studies in History and Philosophy of Science Part A, 44(3): 318–329, September 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000332.

Wenmackers:2023:UPC

[Wen23] Sylvia Wenmackers. Uniform probability in cosmology. Studies in History and Philosophy of Science Part A, 101(??):48– 60, October 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001103.

Werth:1980:TDO

[Wer80] R. Werth. On the theory-dependence of observations. Studies in History and Philosophy of Science Part A, 11(2):137– 143, June 1980. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368180900199.

Werrett:2018:HR

[Wer18] Simon Werrett. History to reckon with. Studies in History and Philosophy of Science Part A, 72(??):59–62, December 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301572.

Westman:1972:KTH

[Wes72] Robert S. Westman. Kepler’s theory of hypothesis and the ‘realist dilemma’. Studies in History and Philosophy of Science Part A, 3 (3):233–264, November 1972. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368172900106.

Wessels:1979:SRW

[Wes79] Linda Wessels. Schrödinger’s route to wave mechanics. Studies in History and Philosophy of Science Part A, 10(4):311–340, December 1979. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL https://www.math.utah.edu/pub/
bibnet/authors/s/schroedinger-erwin.bib;   http://www.
sciencedirect.com/science/article/pii/0039368179900189.

West:2023:BFP

[Wes23] E. James West. Book forum: Pushing Cool by Keith Wailoo: (Pushing Cool, Selling Race). Studies in History and Philosophy of Science Part A, 102(??):70–71, December 2023. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001401.

Weidenhammer:2013:MSM

[WG13] Erich Weidenhammer and Ari Gross. Museums and scientific material culture at the University of Toronto. Studies in History and Philosophy of Science Part A, 44(4):725–734, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000575.

Wilkenfeld:2014:UBG

[WH14] Daniel A. Wilkenfeld and Jennifer K. Hellmann. Understanding beyond grasping propositions: a discussion of chess and fish. Studies in History and Philosophy of Science Part A, 48(??):46– 51, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368114000806.

Wehner:2023:CMD

[WHC23] Rüdiger Wehner, Thierry Hoinville, and Holk Cruse. On the ‘cognitive map debate’ in insect navigation. Studies in History and Philosophy of Science Part A, 102(??):87–89, December 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936812300105X.

White:1971:MCM

[Whi71] P. J. White. Materialism and the concept of motion in Locke’s theory of sense-idea causation. Studies in History and Philosophy of Science Part A, 2(2):97–134, August 1971. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368171900276.

Whiteside:1974:BRB

[Whi74] D. T. Whiteside. Book review: Internationales Kepler-Symposium Weil der Stadt 1971. Referate und diskussionen: F. Krafft, K. Meyer and B. Sticker, ed. H. A. Gerstenberg: Hildesheim, 1973. xii + 490 pp. DM 160. Studies in History and Philosophy of Science Part A, 4(4):387–392, February 1974. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368174900107.

See [KMS73].

Whitt:1988:CDT

[Whi88] L. A. Whitt. Conceptual dimensions of theory appraisal. Studies in History and Philosophy of Science Part A, 19(4):517–529, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368188900131.

Whitt:1990:AAA

[Whi90] L. A. Whitt. Atoms or affinities? The ambivalent reception of Daltonian theory. Studies in History and Philosophy of Science Part A, 21(1):57–89, March 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936819090015Z.

Winsberg:2014:AVR

[WHK14] Eric Winsberg, Bryce Huebner, and Rebecca Kukla. Accountability and values in radically collaborative research. Studies in History and Philosophy of Science Part A, 46(??):16–23, June 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113001295.

Wainer:2022:PTH

[WHMO22] Juan Manuel Garrido Wainer, Natalia Hirmas-Montecinos, and Nicolás Trujillo Osorio. The policy of testing hypotheses in Chilean science. The role of a hypothesis-driven research funding programme in the installation of a hypothesis-driven experimental system in visual neuroscience. Studies in History and Philosophy of Science Part A, 96(??):68–76, December 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122001388.

Whyte:2021:RST

[Why21] Jennifer Whyte. The roots of the silver tree: Boyle, alchemy, and teleology. Studies in History and Philosophy of Science Part A, 85 (??):185–191, February 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368120301904.

Wieten:2021:BRS

[Wie21] Sarah Wieten. Book review: Science and moral imagination: a new ideal for values in science by Matthew J. Brown: Implications for values in medicine. Studies in History and Philosophy of Science Part A, 89(??):299–300, October 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368121000418.

Wigner:1967:SRS

[Wig67] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Wigner:1970:SRS

[Wig70] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-73021-9. viii + 280 pp. LCCN Q171 .W65 1970.

Wigner:1978:SRS

[Wig78] Eugene Paul Wigner. Symmetries and reflections: scientific essays of Eugene P. Wigner. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1978. ISBN 0-313-20107-2. viii + 280 pp. LCCN Q171 .W65 1978.

Williamson:1977:LEE

[Wil77] Robert B. Williamson. Logical economy in Einstein’s “On the electrodynamics of moving bodies”. Studies in History and Philosophy of Science Part A, 8(1):49–60, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900188.

Wilson:1982:LA

[Wil82] Catherine Wilson. Leibniz and atomism. Studies in History and Philosophy of Science Part A, 13(3):175–199, September 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900073.

Wilson:1989:HBW

[Wil89] Andrew D. Wilson. Hertz, Boltzmann and Wittgenstein reconsidered. Studies in History and Philosophy of Science Part A, 20(2): 245–263, June 1989. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368189900058.

Wilson:2001:GLO

[Wil01] Fred Wilson. Galileo’s lunar observations: do they imply the rejection of traditional lunar theory? Studies in History and Philosophy of Science Part A, 32(3):557–570, September 2001. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368100000352.

Williams:2005:CGH

[Wil05] Lambert Williams. Cardano and the gambler’s habitus. Studies in History and Philosophy of Science Part A, 36(1):23–41, March 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368104001189.

Wilholt:2009:BVS

[Wil09a] Torsten Wilholt. Bias and values in scientific research. Studies in History and Philosophy of Science Part A, 40(1):92–101, March 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368108001155.

Willmoth:2009:RIW

[Wil09b] Frances Willmoth. ‘reconstruction’ and interpreting written instructions: what making a seventeenth-century plane table revealed about the independence of readers. Studies in History and Philosophy of Science Part A, 40(4):352–359, December 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000685.

Wills:2009:ESC

[Wil09c] Ian Wills. Edison and science: a curious result. Studies in History and Philosophy of Science Part A, 40(2):157–166, June 2009. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368109000247.

Wilholt:2010:SFG

[Wil10] Torsten Wilholt. Scientific freedom: its grounds and their limitations. Studies in History and Philosophy of Science Part A, 41(2):174–181, June 2010. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368110000154.

Wilson:2012:WT

[Wil12]  Adrian Wilson. What is a text?  Studies in History and

Philosophy of Science Part A, 43(2):341–358, June 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111001385.

Wilholt:2022:EIO

[Wil22a] Torsten Wilholt. Epistemic interests and the objectivity of inquiry. Studies in History and Philosophy of Science Part A, 91(??):86– 93, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001862.

Wilson:2022:KPE

[Wil22b] Robert A. Wilson. Kinmaking, progeneration, and ethnography. Studies in History and Philosophy of Science Part A, 91(??):77– 85, February 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368121001588.

Wilhelm:2024:EC

[Wil24a] Isaac Wilhelm. Explanatory circles. Studies in History and Philosophy of Science Part A, 108(??):84–92, December 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001377.

Wilkenfeld:2024:PWR

[Wil24b] Daniel A. Wilkenfeld. Pursuit-worthy research in health: Three examples and a suggestion. Studies in History and Philosophy of Science Part A, 107(??):64–72, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001213.

Winterbourne:1981:CRS

[Win81] A. T. Winterbourne. Construction and the role of schematism in Kant’s philosophy of mathematics. Studies in History and Philosophy of Science Part A, 12(1):33–46, March 1981. CODEN SHPSB5.

ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368181900030.

Winterbourne:1982:MLS

[Win82] A. T. Winterbourne. On the metaphysics of Leibnizian space and time. Studies in History and Philosophy of Science Part A, 13 (3):201–214, September 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368182900085.

Winkler:1986:BNS

[Win86] Kenneth P. Winkler. Berkeley, Newton and the stars. Studies in History and Philosophy of Science Part A, 17(1):23–42, March 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368186900178.

Wintroub:1999:TSE

[Win99] Michael Wintroub. Taking stock at the end of the world: Rites of distinction and practices of collecting in Early Modern Europe. Studies in History and Philosophy of Science Part A, 30(3):395– 424, September 1999. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S003936819900014X.

Winther:2012:ICS

[Win12] Rasmus Grønfeldt Winther. Interweaving categories: Styles, paradigms, and models. Studies in History and Philosophy of Science Part A, 43(4):628–639, December 2012. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368112000477.

Wisan:1984:AES

[Wis84] Winifred Lovell Wisan. On argument ex suppositione falsa. Studies in History and Philosophy of Science Part A, 15(3):227–236, September 1984. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368184900189.

Wisnioski:2013:DES

[Wis13] Matthew Wisnioski. Design enigmas: SSK in the service of practice. Studies in History and Philosophy of Science Part A, 44 (4):613–617, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368113000411.

Wise:2017:NFS

[Wis17] M. Norton Wise. On the narrative form of simulations. Studies in History and Philosophy of Science Part A, 62(??):74– 85, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368117300778.

Wise:2018:AHW

[Wis18] M. Norton Wise. Afterward: Humboldt was right. Studies in History and Philosophy of Science Part A, 70(??):70–81, ???? 2018. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368118301444.

Wittje:2013:GNE

[Wit13] Roland Wittje. The Garching nuclear egg: Teaching contemporary history beyond the linguistic turn. Studies in History and Philosophy of Science Part A, 44(4):683–689, December 2013. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368113000587.

Witteveen:2024:GSS

[Wit24] Joeri Witteveen. Golden spikes, scientific types, and the ma(r)king of deep time. Studies in History and Philosophy of Science Part A, 106(??):70–85, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S003936812400013X.

Williamson:2022:AAD

[WL22] Hugh F. Williamson and Sabina Leonelli. Accelerating agriculture: Data-intensive plant breeding and the use of genetic gain as an indicator for agricultural research and development. Studies in History and Philosophy of Science Part A, 95(??):167– 176, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122001200.

Wolff:2019:WE

[Wol19]  J. E. Wolff. Why eliminativism?  Studies in History and

Philosophy of Science Part A, 74(??):16–21, ???? 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302279.

Winsberg:2020:SWE

[WOL20a] Eric Winsberg, Naomi Oreskes, and Elisabeth Lloyd. Severe weather event attribution: Why values won’t go away. Studies in History and Philosophy of Science Part A, 84(??):142–149, December 2020. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368120301734.

Wolff:2020:HMM

[Wol20b] J. E. Wolff. Heaps of moles? — Mediating macroscopic and microscopic measurement of chemical substances. Studies in History and Philosophy of Science Part A, 80(??):19–27, April 2020. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300043.

Wolf:2024:CIM

[Wol24] William J. Wolf. Cosmological inflation and meta-empirical theory assessment. Studies in History and Philosophy of Science Part A, 103(??):146–158, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368123001735.

Woolhouse:1971:LPS

[Woo71] R. S. Woolhouse. Locke’s philosophy of science and knowledge: a consideration of some aspects of An essay concerning human understanding. Blackwell, Oxford, UK, 1971. ISBN 0-631-12960-X. xi + 204 pp. LCCN ????

Wood:1975:BCL

[Woo75] Neal Wood. The Baconian character of Locke’s ‘essay’. Studies in History and Philosophy of Science Part A, 6(1):43–84, April 1975. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368175900047.

Woolhouse:1994:LPS

[Woo94] R. S. Woolhouse. Locke’s philosophy of science and knowledge: a consideration of some aspects of An essay concerning human understanding. Modern revivals in philosophy. Gregg Revivals, Aldershot, Hampshire, England, 1994. ISBN 0-7512-0282-7. xi + 204 pp. LCCN ????

Woods:2011:RDA

[Woo11] John Woods. Recent developments in abductive logic. Studies in History and Philosophy of Science Part A, 42(1):240– 244, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110000956.

Woody:2015:ROD

[Woo15] Andrea I. Woody. Re-orienting discussions of scientific explanation: a functional perspective. Studies in History and Philosophy of Science Part A, 52(??):79–87, August 2015. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368115000400.

Woodward:2024:SRR

[Woo24] James Woodward. Some reflections on Robert Batterman’s A middle way. Studies in History and Philosophy of Science Part A, 106(??):21–30, August 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000694. See [Bat23].

Worrall:1982:PLS

[Wor82] John Worrall. The pressure of light: The strange case of the vacillating ‘crucial experiment’. Studies in History and Philosophy of Science Part A, 13(2):133–171, June 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368182900231.

Worrall:2005:PPL

[Wor05] John Worrall. Prediction and the ‘periodic law’: a rejoinder to Barnes. Studies in History and Philosophy of Science Part A, 36 (4):817–826, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368105000713. See [SW01,

Bar05, Sce05, Sch08b].

Worrall:2014:PAR

[Wor14] John Worrall. Prediction and accommodation revisited. Studies in History and Philosophy of Science Part A, 45(??):54– 61, March 2014. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368113000976.

Wray:2005:PSA

[Wra05] K. Brad Wray. Philosophy of science after Mirowski’s history of the philosophy of science. Studies in History and Philosophy of Science Part A, 36(4):779–789, December 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368105000786.

Wray:2006:SAA

[Wra06] K. Brad Wray. Scientific authorship in the age of collaborative research. Studies in History and Philosophy of Science Part A, 37 (3):505–514, September 2006. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368106000264.

Wray:2007:CDT

[Wra07] K. Brad Wray. The cognition dimension of theory change in Kuhn’s philosophy of science. Studies in History and Philosophy of Science Part A, 38(3):610–613, September 2007. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368107000453.

Wray:2015:MDR

[Wra15] K. Brad Wray. The methodological defense of realism scrutinized. Studies in History and Philosophy of Science Part A, 54(??):74– 79, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115001302.

Wray:2018:NTN

[Wra18] K. Brad Wray. A new twist to the No Miracles Argument for the success of science. Studies in History and Philosophy of Science Part A, 69(??):86–89, June 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302601.

Wray:2019:ERA

[Wra19] K. Brad Wray. Essay review: Another great 19th century creation: the scientific journal. Studies in History and Philosophy of Science Part A, 75(??):62–64, June 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0039368118302255.

Wright:1973:AEG

[Wri73] Larry Wright. The astronomy of Eudoxus: Geometry or physics? Studies in History and Philosophy of Science Part A, 4(2):165– 172, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368173900022.

Wright:2015:OCS

[Wri15] Cory Wright. The ontic conception of scientific explanation. Studies in History and Philosophy of Science Part A, 54(??):20–30, December 2015. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368115000746.

Wright:2017:FLC

[Wri17] Aaron Sidney Wright. Fresnel’s laws, ceteris paribus. Studies in History and Philosophy of Science Part A, 64(??):38– 52, August 2017. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116301182.

Weiss:2021:FDM

[WS21] David J. Weiss and James Shanteau. The futility of decision making research. Studies in History and Philosophy of Science Part A, 90 (??):10–14, December 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121001345.

Wu:2014:UNM

[Wu14] Shellen X. Wu. Unearthing the nation: Modern geology and nationalism in Republican China. Studies in History and Philosophy of Science Part A, 48(??):38–41, December 2014. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368114000764.

Wilson:2022:BRV

[WV22] Yolonda Wilson and Lou Vinarcsik. Book review: Vaccine Hesitancy: Public Trust, Expertise, and the War on Science by Maya Goldenberg: Vaccine Hesitancy and the Failure of “Us” versus “Them” Framing. Studies in History and Philosophy of Science Part A, 96(??):191–192, December 2022. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0039368122001534. See [Joh22, Chu22, Lea22] and reply [Gol22].

Wood:2004:IAM

[WW04] Rega Wood and Michael Weisberg. Interpreting Aristotle on mixture: problems about elemental composition from Philoponus to Cooper. Studies in History and Philosophy of Science Part A, 35 (4):681–706, December 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000871.

Wysocki:2023:DBD

[Wys23] Tomasz Wysocki. The delusive benefit of the doubt. Studies in History and Philosophy of Science Part A, 100(??):47– 55, August 2023. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123000730.

Yap:2011:GQR

[Yap11] Audrey Yap. Gauss’ quadratic reciprocity theorem and mathematical fruitfulness. Studies in History and Philosophy of Science Part A, 42(3):410–415, September 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368111000239.

Yearley:1982:RBE

[Yea82] Steven Yearley. The relationship between epistemological and sociological cognitive interests: Some ambiguities underlying the use of interest theory in the study of scientific knowledge. Studies in History and Philosophy of Science Part A, 13(4):353–388, December 1982. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936818290019X. See reply [Mac84].

Yearley:1985:IIP

[Yea85] Steven Yearley. Imputing intentionality: Popper, demarcation and Darwin, Freud and Marx. Studies in History and Philosophy of Science Part A, 16(4):337–350, December 1985. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368185900160.

Yearley:2005:WEN

[Yea05] Steven Yearley. The wrong end of nature. Studies in History and Philosophy of Science Part A, 36(4):827–834, December 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368105000671.

Yee:2024:EMS

[Yee24] Adrian K. Yee. Edgeworth’s mathematization of social well-being. Studies in History and Philosophy of Science Part A, 103(??):5– 15, February 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368123001590.

Yeo:1989:RHD

[Yeo89] Richard Yeo. Reviewing Herschel’s discourse. Studies in History and Philosophy of Science Part A, 20(4):541–552, December 1989. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/003936818990023X.

Ylikoski:2019:MBT

[Yli19] Petri Ylikoski. Mechanism-based theorizing and generalization from case studies. Studies in History and Philosophy of Science Part A, 78(??):14–22, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117303175.

Yolton:1970:LCH

[Yol70] John W. Yolton. Locke and the compass of human understanding: a selective commentary on the Essay. Cambridge University Press, Cambridge, UK, 1970. ISBN 0-521-07838-5. x + 234

pp. LCCN B1294 .Y64 1970. URL http://www.loc.gov/catdir/
enhancements/fy1001/76112477-d.html;   http://www.loc.
gov/catdir/enhancements/fy1001/76112477-t.html.

Yoshida:2023:JRM

[Yos23] Yoshinari Yoshida. Joint representation: Modeling a phenomenon with multiple biological systems. Studies in History and Philosophy of Science Part A, 99(??):67–76, June 2023. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368123000705.

Young:2004:NEE

[You04] Brian Young. Newtonianism and the enthusiasm of Enlightenment. Studies in History and Philosophy of Science Part A, 35 (3):645–663, September 2004. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368104000597.

Ylikoski:2019:CSR

[YZ19] Petri Ylikoski and Julie Zahle. Case study research in the social sciences. Studies in History and Philosophy of Science Part A, 78(??):1–4, December 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368119303334.

Zahle:2019:DEV

[Zah19] Julie Zahle. Data, epistemic values, and multiple methods in case study research. Studies in History and Philosophy of Science Part A, 78(??):32–39, December 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118301742.

Zammito:2003:IPO

[Zam03] John H. Zammito. ‘this inscrutable principle of an original organization’: epigenesis and ‘looseness of fit’ in Kant’s philosophy of science. Studies in History and Philosophy of Science Part A, 34(1):73–109, March 2003. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368102000924.

Zammito:2008:TTT

[Zam08] John H. Zammito. A text of two titles: Kant’s ‘A renewed attempt to answer the question: “is the human race continually improving?”’. Studies in History and Philosophy of Science Part A, 39 (4):535–545, December 2008. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368108000885.

Zammito:2016:EKR

[Zam16] John H. Zammito. Epigenesis in Kant: Recent reconsiderations. Studies in History and Philosophy of Science Part A, 58(??):85– 97, August 2016. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368116000224.

Zammito:2018:MHP

[Zam18] John H. Zammito. Book review: Materialism: A Historico– Philosophical Introduction, Charles T. Wolfe. Springer International Publishing, Switzerland (2016), pp. xi + 134. Price US$37.99 paperback, ISSN 2211-4548. Studies in History and Philosophy of Science Part A, 67(??):89–96, February 2018. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368117302595.

Zandvoort:1982:NCT

[Zan82] Henk Zandvoort. A note on closed theories. Studies in History and Philosophy of Science Part A, 13(1):81–86, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368182900061.

Zandvoort:1988:MDS

[Zan88] H. Zandvoort. Macromolecules, dogmatism, and scientific change: The prehistory of polymer chemistry as testing ground for philosophy of science. Studies in History and Philosophy of Science Part A, 19(4):489–515, December 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/003936818890012X.

Zappala:2024:FID

[Zap24] María Alejandra Petino Zappala. A framework for the integration of development and evolution: the forgotten legacy of James Meadows Rendel. Studies in History and Philosophy of Science Part A, 105(??):41–49, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368124000505.

Zellmer:2024:DCD

[Zel24] Jacob Zellmer. Descartes on certainty in deduction. Studies in History and Philosophy of Science Part A, 105(??):158– 164, June 2024. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368124000621.

Zemplen:2006:DNP

[Zem06] Gábor Á. Zemplén. The development of the Neurath principle: unearthing the Romantic link. Studies in History and Philosophy of Science Part A, 37(4):585–609, December 2006. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S003936810600077X.

Zhavoronkov:2022:KPU

[Zha22] Alexey Zhavoronkov. Kant’s pragmatic use of reason from a sociological point of view: Third way or methodological impasse? Studies in History and Philosophy of Science Part A, 94(??):1– 7, August 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000711.

Zinsser:2011:MBN

[Zin11] Judith P. Zinsser. Multiple beginnings: new insights on the Scientific Revolution and Enlightenment in France. Studies in History and Philosophy of Science Part A, 42(1):245–249, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368110000968.

Zolo:1993:BRB

[Zol93] Danilo Zolo. Book review: Rediscovering the forgotten Vienna Circle: Thomas E. Uebel (ed.), Boston Studies in the Philosophy of Science, Vol. 133 (Dordrecht: Kluwer, 1991), xii + 326 pp. ISBN 0-7923-1276-7 Cloth Dfl. 175.00/$99.00/£59.00. Studies in History and Philosophy of Science Part A, 24(3):477– 484, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819390038L. See [Ueb91b].

Zuchowski:2019:MKT

[Zuc19] Lena Zuchowski. Modelling and knowledge transfer in complexity science. Studies in History and Philosophy of Science Part A, 77 (??):120–129, October 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117302443.

Zwier:2011:JDP

[Zwi11] Karen R. Zwier. John Dalton’s puzzles: from meteorology to chemistry. Studies in History and Philosophy of Science Part A, 42(1): 58–66, March 2011. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368110001007.

Zynda:1996:SWR

[Zyn96] Lyle Zynda. Should we reject supervenience analyses of laws, chance, and causation? Studies in History and Philosophy of Science Part A, 27(4):587–592, December 1996. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368195000593.

Zhu:2020:WDC

[ZZW20] Jing Zhu, Mingjun Zhang, and Michael Weisberg. Why does the Chinese public accept evolution? Studies in History and Philosophy of Science Part A, 81(??):116–124, June 2020. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368118302437.